A Bibliography on Pattern Matching, Regular Expressions, and String Matching

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

and

David Salomon Professor Emeritus Department of Computer Science California State University Northridge Northridge, CA, USA

05 November 2025 Version 1.228

Title word cross-reference

(l, d) [Tan14]. 1 [Mun07]. 2 [ASG99, BSM^{+}07, BZ98, CR95b, KPR97, KD15, LT90b, SHCY93, Via02, Via04]. $29.95 [Ano97a]. 3 [BSM^{+}07, CJ93, LT90b, TCCK90]. $65.00 [Ano97b]. ^{(1)} [Sid99, Sid02]. ^{2} [Ram94]. ^{33} [BGFK15a, BGFK15b]. ^{4} [ZLN11]. m [Jan23]. U [Jan23]. c [WD99]. c^{n} [Rob79]. D [Sid99, CGK08, CDEK95, GKP19, dRL95]. d^{(1)} [Sid95, Sid00]. δ [CIL^{+}03]. ϵ [Gef03, HM98, HSW01, Lif03]. GREP^{(1)} [Sid00]. K [COZ09, LVN87, ATX21, ALP04, CLZ^{+}15, CKP^{+}21, CW18a, CW18b, CN21, FWW13a, FWW13b, FGKU15, GG86, GU16, GN19, GGF13, Gra15, GL89, HT17, JL93, JL96, JN24, KRS19a, KRS19b, KRS23, KVX12, LV88, NYuR15, NR17, QQC^{+}13a, QQC^{+}13b, WD99]. L_{1} [LP08, LP11]. L_{2} [LP11]. L_{∞} [LP11]. L_{p} [WC14]. μ [DJ96]. N [ML96a, ML96b, KST94, KWL07, KWLL08, Wag74]. O(modT mod ^{3}) [KKM^{+}85]. O(log log n) [BG90]. O(n^{2} log m) [CNS18]. O(n⌈m/w⌉) [GF08]. O(n log^{2}(n)) [HM98]. O(n log^{3} m) [CH97b]. O(n log n) [Gef03]. O(s^{2}) [CZ01]. ω [BdFED^{+}20, EJ23]. ω^{n} [Cho78]. ωT [BdFED^{+}20]. q [HKN14, KB22, KPA10, STK06, Sal12, ST95, ST96b, ST96c, ST04, Ukk92]. r [Pol13]. ρ [CFK07]. t [Liu14]. x^{m}y^{n} = z^{p} [NC92].

-automaton [COZ09]. -bit [KVX12]. -calculus [CFK07]. -Cube [ML96a, ML96b]. -D [SHCY93, BZ98]. -dependent [Jan23]. -Dimensional [CDEK95, CR95b, KPR97, CGK08]. -dimensions [dRL95]. -expression [COZ09]. -formulae [WD99]. -frame [KB22]. -Free [HM98, HSW01, Gef03, Lif03]. -Gram [ST95, HKN14, KST94, KWL07, KWLL08, KPA10, Sal12, KWL07]. -Gram/2L-approximation [KWL07]. -Grams [ST04, STK06, Ukk92]. -Interval [Via02, Via04]. -Like [HK11]. -M [Ram94]. -Matching [CIL^{+}03]. -means [WD99]. -Mismatch [ATX21, JN24]. -mismatches [GGF13, Gra15]. -motif [Tan14]. -norm [WC14]. -Regular [EJ23, BdFED^{+}20, BdFED^{+}20]. -Samples [ST96b, ST96c]. -Separated [Pol13]. -Statistics [Liu14, Jan23]. -tapes [Cho78]. -time [CNS18]. -TM [BGFK15a, BGFK15b]. -transformation [Sid95, Sid99, Sid00].

.NET [AS04, SM04, Stu07].

’08 [ACM08].

1 [KJS17]. 1003.1-2001 [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d]. 10th [PC99]. 11th [GS00, Kap92]. 12th [AL01, Bun94]. 13th [AT02]. 14th [AAC^{+}01, BYCC03]. 15th [SMD04]. 16th [ACP05]. 17th [LV06]. 18th [MZ07]. 1969 [ACM69]. 1971 [FGR72]. 1974 [ACM74]. 1976 [ACM76]. 1978 [Win78]. 1979 [Ng79]. 1981 [ACM81]. 1984 [ACM84]. 1986 [ACM86]. 1987 [DT87]. 1990 [ACM90b, Cro92a, NEH90]. 1991 [ACM91]. 1992 [ACM92a, ACM92c, ACM92d, Sto92]. 1993 [ACM93a, ACM93b]. 1994 [ACM94a, ACM94d, SW94]. 1995 [ACM95a, ACM95b, ACM95c]. 1997 [ACM97b, ACM97c]. 1999 [ACM99a, ACM99b, AOV^{+}99]. 19th [ABB93, FL08].

2001 [AAC^{+}01]. 2002 [B^{+}02]. 2004 [SC04]. 2005 [SC05]. 2006 [ACM06]. 2007 [ACM07]. 2008 [ACM08, LL08, Neu10]. 2011 [NH11, SM09, SM11]. 20th [KU09]. 21st [ACM94a, AP10]. 22nd [ACM95a, GM11]. 23rd [KS12a]. 27th [KP15]. 2L-approximation [KWL07]. 2nd [OND98]. 2nd-order [OND98].

3 [Ruc15]. 3.0 [BWN08]. 30th [IEE89]. 31st [IEE90, KLB12]. 32 [Gro91a]. 33rd [IEE92]. 34th [IEE93]. 36th [IEE95a]. 38th [IEE97]. 39th [IEE98].

4 [Ano12, Bro77]. 40th [ACM08]. 4th [Apo93a].

5 [B^{+}05]. 50th [IEE09]. 512 [CGT25]. 5PM [BEM^{+}12, BEM^{+}13]. 5th [CG94b, NH11].

6 [IEE01a, IEE01g, IEE01e, IEE01b, IEE01h, IEE01f]. 68 [HK77]. 6th [BGNP94, GU95].

7 [Rob99a, Rob99b]. ’79 [Ng79]. 7th [DMVT13, HM96, Hwa85, Win78].

’86 [Cha86, CVP86]. ’87 [ACM87, Ano87]. ’88 [IEE88]. 8th [AH97].

’90 [ACM90a, WN90]. 90k [Gro91a]. ’92 [ACM92b]. ’93 [ABB93, Bao93, SC93]. ’94 [ACM94a, ACM94c, BGG^{+}94, SW94]. ’95 [ACM95a, ACM95b, IEE95b, Lev95, SC95]. ’96 [SC96]. ’97 [ACM97a]. 978 [Ano12]. 978-1-4493-1943-4 [Ano12]. ’98 [ACM98, SC98]. ’99 [SC99]. 9th [FC98, FJ92].

A-DFA [BC13b]. A. [AFI98, Pet95]. Aarhus [AH97]. absence [AGH^{+}17]. absolute [ZZJC20]. Abstract [CDL95, Gon02, HOS85b, JO97, LM02, Pre99, AG06, BC93, Chl08, CM95, GPN96, GV00, HOS85a, Pie08, Zei08]. Abstracting [JSH09]. abstraction [Wad87]. Abstractions [Jok90, MNS07]. Accelerated [ZCH23, BYB^{+}24, MLC08, SR16]. Accelerating [BBK12, GÁSÁ^{+}13, KBN09, LLCC13, LLL^{+}24, TT82, VSP08]. Acceleration [GWX^{+}23, SALP20]. Accelerator [JLK^{+}20, KYG19, TT22, TLC15, WGL^{+}21, ZL18, TLLL09]. Accelerators [HKL^{+}14, LHMH91, MGW14]. acceptors [TIT83]. Access [Fal85, JDXD13, MR11, JSH09, KT14, KCK93, KFG15]. accesses [DSv94]. accessible [SBR^{+}07]. According [PV91]. account [Fil21, KSVJ15]. Accurate [BAM^{+}24, ZS17, STKD20]. achieve [AK08]. Acid [CCL87, LVN87]. ACM [ACM69, ACM74, ACM83, ACM84, ACM86, ACM87, ACM89, ACM90a, ACM90b, ACM91, ACM92a, ACM92c, ACM92b, ACM92d, ACM93a, ACM93b, ACM94a, ACM94c, ACM94d, ACM95a, ACM95b, ACM97b, ACM97a, ACM97c, ACM98, ACM99a, ACM99b, ACM00, ACM06, ACM08, DGBH93, FMA02, HF13, KLB12, LL08, Len11, SW94, Sto92]. ACM-SIAM [ACM97b]. ACM/SIGAPP [DGBH93]. Acquisition [BZ98]. across [DCM15, SHS14]. action [Han92]. actions [CK08]. ActionScript [BWN08]. active [BDMT16]. activity [BWG12]. ActorSpace [AC93]. Ada [Wes97, TBD22, Wes97, WT88, WT89]. Ada/Tcl [Wes97]. Adapted [RJK79]. Adapting [CFG12, DS04, LRV13]. Adaptive [CW84, HMW^{+}23, JP73, LDI98, NdMM02b, SRR92, SRR95, SW09, HRN^{+}15]. Adding [Sha88b]. Addition [PP06]. address [AAK^{+}09, AAB^{+}09, AEK^{+}11, ZZJC20]. addressable [LMT16]. Addressing [RJK79, Lau01]. Adjacencies [LJZZ13]. Adjeroh [Neu10]. Administration [Sar02]. Advanced [B^{+}07]. Adversarial [TD18]. Adversaries [HL10, HT14]. Advisor [Mu 95, MuT95, Mun95]. Affine [DN77, VS01]. affordable [NE93]. afind [GN01]. Against [Bun95, HL10, LA12, AV23, BSTU08, MPW21]. Aggregations [CvW18]. aggressive [Dai09]. Agrep [MW92b, GN01, WM92b]. Ahead [Yan95, ZBST14]. Aho [CW13, NK07, PLL10, RKM21, TM05b, TZH^{+}13, TVCM12]. Aid [AC75]. Aided [KP15]. Alberto [Ano97b]. Albuquerque [ACM92a]. Algebra [KN12, LRSV18, SS93a, BFS00, Coo86, Fat15, KMMPN85]. Algebraic [ACM94b, Bro93, Cha86, Hea71, Lev95, McI85b, Ng79, WN90, BD98, Fat15, McI85a, OR11]. Algebras [CM95]. Algol [Bro77, HK77]. Algol-based [Bro77]. Algorithm [AR00, ATX21, HR02, AJS92, ACD01, BST^{+}03, BYKZ^{+}92, BYN96, Bar81, BC13b, Ben94, Bir10, BM77, Bra94, Bre93, BL16, CF06, CLP98, CF85, Col94a, CH02, CGH^{+}98, DNM00, FL12a, Gal79, GP90, Gal95, GC01, IMP01, IST05, IS86, KRS97, KST94, Kül10, KV15, KZ02, LY17, LLCC13, LJH^{+}17, LLC17, LDI98, LJZZ13, LCL06, Man75, MW94, MR11, MUHT96, McC76, MOSZ18, MNS84, ML96a, ML96b, Mye92, Mye98, NW13, NBY01, Nel96, Ode14, OR12, PS10, Pou93, PK85, Rai92, RKM21, RPE81, Sad96, Ski98, Smi91b, Smi94, SW09, Sun90a, Sun90b, Tak86a, Tak86b, VB12, VB98, WPKL13, Wat96, WT89, WMM95, Yam01, Alb89, AR13, AGW13, BYR92, BPMA02, BGJ89, BBM21, BG90, Bre96, BC95, Cha93a, CLS95, CW13, CDC96, CNPS15, CNS18, Col90, CR91]. algorithm [Dai09, DR06, DS04, Der95, Dow91, Gal92, GBY90a, GBY90b, GO80, GLS92, Han93, HFS05, HR03, IKX15, IP96, II86, ISHY88, IA80, KKM^{+}85, KR89, KST92, Kim99, KKR^{+}13, KIH15, LV86b, LVN87, Lec92, Lee82, Liu81, LHCK04, Maa06, MBY91, Men89, Mis03, MKSiA98, MS95, Mor90, Mye99, NRO12, Neb06, PLL10, PS90, Per94, Ryt80, SW90, Sch88a, SS94, SGYM00, Sto02, Tak96b, Tak93, TA90, TZYH14, TJD^{+}17, TM05b, TU88, TTO^{+}22, Tho68, TJMC20, WW03, Wat03, WT88, Yam19, YT03, YHV^{+}15, ZC99, dB93]. algorithme [Alb89]. Algorithmic [ABBH^{+}16]. Algorithms [ACM97b, AHU74, ALR08, iA94, ADLM96, BY96a, BLP94, Bak96, BS97, BH02, BJM79, BCFL12, CFG12, CFP19, CL92, CL97, CHL14, Chu95, CHZ06, CLR90, CR92, CCG^{+}94, DB86, FL12b, FMMS20, FRU^{+}20, FBY92, Gal76b, GG97, GS85, GIG77, GK86, Gus97, HUN^{+}19, Hig86, HSTS01, ISNH94, JTU96, KKP16a, KKP16b, KTU87, KR81a, KR81b, KR87, KP99c, Kha16, KMT^{+}01, KTY^{+}18, Lab12a, Lab12b, Lec95, LLLC17, LT16, LS94, Lut02, MR11, MP05, Med23, Mut97, Ott94, Par96, Pol13, PRT^{+}16, RS98, SV94, SN92, Sed83, Sed90, Sed92, Sim94, Ste94, Tar81a, VG01, YD95, ZZ12, de 82, ALP04, Apo93b, AG97, ADLM01, ARS16, BYF96, Bak93, Bar84a, BYB^{+}24, CMO^{+}08, CDDM05, CX20, CHL^{+}25, CLT07, CWZ10, CCG^{+}93, Col91, CT96, CR94, CL96, DC94, DV21, ECSS88]. algorithms [Gal75, Gal76a, Gal84, GG13, HIEH22, HR01, HOK18a, HOK18b, HTX17, Ind98, JHU^{+}24, JU91, KB22, KN00, KM13, KM84, Lec07, MAC14, MW92b, Mha05, MM07, MR13, NR02, Par98, PDC94, PS88, QLY07, Sal12, Sch91a, Sch91b, SZ05, SN24, Tan14, THG17, Val09, VHL^{+}12, WZ96, A^{+}08, Len93, Ano97b]. Aligned [LSTW^{+}17, SN94]. Alignment [BLP94, Ben94, BAM^{+}24, BDFW94, HPM94, JWZ94, KK08, LPT12, LPR^{+}08, Pol13, RND97, RFD23, CLT07, FSL^{+}15, NT20]. Alignments [Cha94]. All-Against-All [LA12, BSTU08]. Alley [DNM00, Nel96]. Allocation [VSM87, YD95]. Allowing [FNU02, CCF13, WM92a]. Allows [Man94, Man97]. Almost [CGPS13b, GR99, FS13, LMM17]. Almost-linear [CGPS13b]. Almost-optimal [GR99]. Alphabet [AFM94, ABF94a, CR95b, KR94, KRR17, KMRY20, SJNS19, TP97, AFI98, AGM05, GP92, KMR21]. Alphabet-Friendly [SJNS19]. Alphabet-Independent [CR95b, KR94, GP92]. Alphabet-Invariant [KMRY20, KMR21]. Alphabets [Bre94a, CLP98, Fre02, KT06, KST12, NR21, Ris16, STK10, Cro92b, Fre03, YHV^{+}15]. Alternating [BL16]. Alternative [Bar81, JWZ94, AP90, Fat15]. Alto [IEE93, IEE98]. always [LMM17, LLS^{+}20]. Amar [Neu10]. AMASS [KS99]. Ambiguity [MGH93, SL17]. ambiguous [BBM21, NdMM02a]. American [NEH90]. amino [LVN87]. Analyses [KPP19, WHZ^{+}17]. Analysing [HH93a, HH93b, HH93c]. Analysis [AHU74, AR13, AJS92, BBH^{+}87, FO76, KR92, Les95, Liu88, LS94, Med23, Nip98, Par96, Par98, PRT^{+}16, SJ13, SCFC94, SP16, Sca11, SWZ01, WCW82, Yan95, ZWH^{+}21, DSv94, GLS07, GBY90a, GBY90b, HV93, KT14, MAC14, MP09, MLM^{+}08, NA90, ORPF13, PIR17, PM23, SW93, Sid95, THQ19, TPT13]. analyst [ZV97]. analytic [SHvR^{+}16]. analytical [Bar84a]. analyze [CFM00]. Analyzer [LS79, ZGE85]. Analyzing [HSTS01, MR82, MNNS12]. Ancestry [FK16]. Ancient [SP16]. ands [Edw07]. animate [BYF96]. Annotated [Sak21, GGN06, RH81]. annotating [AMRV16]. annotation [YCJK08]. Annual [ACM81, ACM87, ACM92a, ACM93a, ACM97b, ACM00, ACM08, AP10, AH97, AT02, FC98, FL08, FJ92, GM11, HM96, IEE90, IEE92, IEE93, IEE95a, IEE97, IEE98, IEE09, KS12a, KU09, Len93, LV06, MZ07, NEH90, ACM74, ACM76, ACM84, ACM86, ACM89, ACM90b, ACM91, ACM92d, ACM93b, ACM94d, ACM95c, ACM97c, ACM99b, AL01, Apo92, Apo93a, ACP05, BYCC03, CG94b, DT87, GU95, GS00, IEE89, IEE95b, PC99, SMD04]. Anomaly [GKW^{+}10]. Answer [KKSL01, ADT15]. Answering [CJR^{+}21, KKSL01, ZCÖZ12, AL08, CDL08, CKC07]. Answers [GN19, Ano92a]. Antidictionaries [STSA99]. Antisymmetric [Gil70]. Antonio [IEE94b]. Ants [FR17, Joh01]. Any [Bur11, PW93]. Apostolico [Ano97b]. Application [AV23, GPP04, GT90, Hud89, IK83, MKF91, MGW14, NA90, WKA94, Akl78, CFK07, Fat15, GÁSÁ^{+}13, Lau00, Man76, MW94, MM03, PA10, SHS14, Sid95, Sid99, TIAY90, WKR09, dLFM07]. application-database [SHS14]. application-specific [WKR09]. Applications [Bak96, BM00, Brz62, CL94, DMVT13, FK16, Gia93, GV05, HSTS01, Hui92, HN05, IEE94b, IEE95b, IIK08, KRRW24, MHKR12, NR03, Pol13, Sch95, AS04, Bak93, B^{+}07, Che96, CX20, FMdB99, FG99, FLSS93a, FLSS93b, GV00, Ind97, Jan23, KKP92, RTO15, Sid00, SVM14, SR16, SWZ01, SALP20, VHL^{+}12, WYA^{+}07]. Applied [DGBH93, PDL98, DGBH93]. Applying [AK08, SdM01]. Approach [ABF94a, CFM17, CCH09, Cox09, DC94, FKRV15, FL12b, GN19, IMR08, KTSA99, KS99, LP13, LN19, Lut02, NR98, NR99b, RM88, Reg92, Sha93, Tar81b, AP13, B^{+}05, BYG92, BSTU08, BG91a, BCD14, BYK22, FS22, FMdB99, GPR95b, GW92, Goo05, HLN09, KD15, LBK08, LMS21, MF96, Mus03, dSOMY15, PP85, PSK17, SVS97, SD91a, SD91b, SSLL21, Sri93]. Approaches [BM08, vNG01, FBMA05, MR13]. Approximate [Aku94, Aku95, AAK^{+}09, AEK^{+}11, ACD01, BYP92, BYN96, BYN97, BYN98, BYN99, BCP02, BH02, BPPR20, BM00, BK93g, Bun95, CJM12, CLS^{+}10, CL90, CL92, CM94, CL94, CCH09, CN02, CH02, CIM^{+}02, EMC96, FPT22, FNU02, Fre06, Fu96, GP90, GIMV03, GGF13, HD80, HLS07, HT17, HM00, HHLS06, HN02, HN05, IMP01, JTU96, KYG19, KH25, KM92, KM95a, KST16, LSW08, LH03, LP11, LLW^{+}15, MW92a, MW92b, Mel95, MM02, MIH17, MM96, MM89, Mye98, MOG98, NW13, Nav98, NBY99a, NBY99b, NBY01, Nav04a, NRS18, OM88, PMD01, Par96, PW95, Phi94, PP09, RDRF24, RNOM09, Sad96, STK10, ST95, ST96b, ST96c, ST04, TT20, Tak94, TU93, Ukk92, Ukk93, UW93, VRD01, Wri94, WM92b, WMM95, ZMAB03, van14a, van14b, AGW13, BYP96, BLPL92]. approximate [BFG09, CRV06, DLF^{+}15, DC94, FCLST07, FS22, FN04, HOK18a, HOK18b, HLS^{+}11, HTX17, HFN05, Hyy08, JU91, KST92, KWL07, KNT11, LV86b, LV89, LG16, LT97, LLL13, MW94, MM03, MM07, MBH20, Mus05, Mye99, Nak14, NBY99c, Nav00, NKT^{+}01, Nav01a, NF04, NC06, Par98, Sad93, SY23, SW90, TLS16, WC14, ZA17, ZD95]. Approximated [PW93]. Approximately [Cob94, KRS19a, KRS19b, KRS23, Mye95]. Approximating [TY97]. Approximation [ADLM96, ADLM01, BLP94, CM08, Huc21, LJZZ13, KR89, KWL07, TU88]. April [ACM74, ACM84, ACM90a, AGS93e, Apo92, DMVT13, SC93, SC96, SC98, SC02]. Arabic [JZAA19, Kul11, Mus03, Mus05, ZA87]. arbiters [SMT^{+}86]. Arbitrary [Nav04a, WMGS19, YH92]. arc [GGN06]. arc-annotated [GGN06]. Architectural [CL09, GSL17, IS90]. Architecture [AWD^{+}18, BYHT18, BTC06, CG87, CF85, CDC^{+}23, HKL^{+}14, KL02, LHCH93, Lee09, PLL08, SRV^{+}19, TS05, YP12, ZCH23, dLBHC22, FKSB06, KRL87, MM07, NNSP22, TYNM86, ZV97]. Architectures [OWP16, TVCM12, San09]. Arden [LHCH93]. area [SV87]. Ariel [Han92]. Arithmetic [Hwa85, KP96a, MHKR12, CS22, KP96b, MP88]. Arizona [Apo92, ACM97a]. arm [NHN^{+}20]. Armenian [Gue87]. Array [CPW88, GHK^{+}91, LK90, WBA83, DK13, LK88, ME97, MM07]. Arrays [AOK02, ABM08, GV05, MM93, Neu10, PP06, Shi96, Bak78, CR91, DSv94, GV00, HHLS06]. Arrhythmia [ARM^{+}19]. arrivals [SWZ01]. Art [DGBH93, Knu05, Ruc15, WDG^{+}14]. Artificial [IEE94b, ZGY^{+}16]. ASCII [Pol01]. Asia [IEE94a]. Asia-Pacific [IEE94a]. Asilomar [CG94b]. Aspects [FJ92]. Assembly [KS99, MW92b, Sno01, FL71]. assertional [PS90]. Assertions [MC24, Jay92]. assessment [HBRV10]. assignments [LSO17, Mor02]. assist [KKM^{+}06]. Assisting [XPZ^{+}23]. Associated [CLST^{+}13]. Association [DT87]. Associative [Eke95, YIAS89]. Associative-Commutative [Eke95]. Assuring [YDW18]. Astrid [STKD20]. Astronautics [WKA94]. Asymmetric [QWX^{+}13, CGPS13a]. Asymptotic [Jan23, KC11]. Asynchronous [KFG15, LPJ23]. ATARI [TSM88]. Athens [Len11]. Atlanta [ACM83, ACM99b, IEE09]. Atlas [TMV^{+}01]. Atoms [Les79]. Attacks [ABBH^{+}16, AV23, PM23, VS11]. attention [TJGY22]. attribute [Har02]. Attributed [CTF^{+}98, Gro92]. attribution [Far92a, Far92b]. Audio [DCM15]. Audit [KNMH00]. Augmented [OS11, SA96]. August [ABB93, BGNP94, B^{+}02, FGR72, IEE95b, WN90]. authenticated [PPTT15]. Authentication [RMK^{+}14]. Authorizing [WYA^{+}07]. autocompletion [XQW^{+}13]. Automata [AWD^{+}18, Ant95, Cha02a, CLOZ04, Coo72, CDPP23, DKA^{+}15, DMVT13, DM11, EZYA23, FL12a, Ghi62, Gol93, GWX^{+}23, GH13, GH15, HIRS17, HU79, HU92, HMU01, HMU07, HSW97, HSW01, KPR97, KPR00, KV15, LKM23, LT16, LPJ23, Loh10, MS98, MHKR12, MHT09, MY60, Mel95, MSZ17, Moh97, MR09b, NR98, NWE97, Ned98, OS11, Pet92, RS59, RSG^{+}19, SRV^{+}19, SM56, SM74, She59, SSSS10, Sim94, UW93, dLBHC22, BYCG94, BYBDS09, BdFED^{+}20, BDFR08, BS86, BT21, BH96, BK93f, Cho78, CR87, FS22, GOMSJVGP08, Gef03, HW07, HM00, HR00, Kle56, Lau00, LSO17, Lei81, NR00, NWE99, NdMM02a, NK07, Pet94, Ryt89, SY72, SBR^{+}07, SH85, THL^{+}20, VW11, YH91, YB13, ZZH16]. automata-driven [NWE99]. automata-like [YH91]. Automata-Processing [AWD^{+}18]. Automated [BY92, BY96b, Bun94, CS18, Kap92, PPPdG20, RDRF24, Zha96, Rém17, BY91, Bun94]. Automaten [SM74]. Automatentheorie [HU92]. Automatic [BA06, BDD^{+}14, DKA^{+}15, DMWW77, KP93, OES25, PHXD19, RTO15, SWY75, SMS15, WKR09, Ear74, HA90, MR09a, PS93a]. Automatically [CGP^{+}08, Kuk92, Mor02, SdM01]. Automating [Kah06, VR18]. Automaton [BP11, CZOdlH17, CZ01, GJ16, LY86, MOSZ18, Prů17, Ant96, BYG96, COZ09, Hur84, Lei80, TLLL07, TLLL09, ZC99]. automaton-matching [TLLL09]. Autonomously [KT23]. Auxiliary [CZW15]. availability [LLL12]. Average [BYR92, BLP18, BMMR19, FN04, HKN14, Mon17, NF04, Sal12, SCFC94, Bar22, CGR99, GFG11, Quo92]. Average-Case [BLP18, HKN14, SCFC94, Bar22, Quo92]. Average-optimal [FN04]. Avoiding [AGM19, Fos89, Thi93]. AVX [CGT25]. AVX-512 [CGT25]. AVX2 [CGT25]. aware [MBP22, WOQ^{+}07]. Awk [AKW85, Dou91, Mis89a, Mis89b, Bar90]. AWK-like [Mis89a, Mis89b]. Axiom [Ano68, JT94]. axiomatization [HN11]. axioms [CD18].

B [Pet95, FG99, FV16]. B-Tree [FV16, FG99]. B.E. [Sca11]. Back [GH15, ESL89, FS19, Lar98]. back-references [FS19]. background [RH81]. Backing [BAP06]. backreferences [BvdM23]. Backtracking [FKP77, MNR^{+}23]. backward [Sal12]. Bad [Len93, PRT^{+}16, MLM^{+}08]. Baeza [Hyy08]. Baeza-Yates [Hyy08]. Balanced [EJ23]. Balanced-by-Construction [EJ23]. Balancing [GL19, MM02, MM03]. Baltimore [ACM90b]. Banff [A^{+}08]. Barcelona [LV06]. Bark’ [Gue87]. Base [IEE01a, IEE01b, ZZJC20]. Based [AOK02, AV23, BYHT18, BL16, CFP19, CDM11, CZCD09, D’A98, DS19, EGP14, EZYA23, FYJ^{+}17, FL12a, FMP20, GMAS22, GR96, GKW^{+}10, HH83, JZAA19, Kid09, KKSL01, KL02, KVX12, KNMH00, LLLC17, Liu14, LT09, LS94, Lut02, MU02, MGF97, MNR^{+}23, Mye98, ND02, OWP16, PS10, PP06, Sad96, SF01, SvS14, SL17, SXZ^{+}25, TMV^{+}01, TK07, VB23, WPKL13, WMGS19, WD99, WZU14, YC22, Yun12, AFI98, Agu23, ASM17, Ano96a, ARS16, BKLE18, BDMT16, BC06, Bro77, BFS00, CW13, CFCK22, CLP95, CK08, DLF^{+}15, Far92a, Far92b, Gan89a, GN19, Gre88, Ier09, IIK08, II08, JHU^{+}24, KB22, KS07, KN00, LLL12, LHCK04, Lus94, MLC08, Mye99, NE93, NRO12, Ode14, PD12, PJMR14, PRT^{+}16, PKK18, Rus92, SBB19, SNM^{+}13, SZ05, SSYW19, SMW^{+}23, THQ19, TM04, TP07a, TP07b]. based [TPT13, TN13, TN15, VRC24, WL15b, WSW16, WHZ^{+}17, YT03, ZV97, ZZ12, ZZJC20, ZYX^{+}12, dH05, Ano96b, NTS93]. Bases [AAC^{+}01, B^{+}02, Gon83, ABB93]. Batched [Man86]. Bayesian [SD91a, SD91b]. BC [LL08]. Be [Cox07, LY86, EMT23, PW93, Sch91a, Sch91b, AK09b]. Beach [HM96, CVP86, IEE97]. Beam [DMWW77]. beams [NA90]. Beating [ZGY^{+}16]. beats [THG17]. Beautiful [OW07]. Beauty [FvGGM90]. BEG [ESL89]. Beginning [JNS08]. Behavior [CDM11, Wei83, LYWL08]. Behaviour [BMMR19]. behind [CA20]. Beijing [ACM07, Bao93]. Bell [Neu10]. Benchmark [Ano10, KL02]. Berkeley [ACM86]. Best [Ben94, BK93g, LG16]. best-effort [LG16]. Best-First [BK93g]. Better [Gon02, TWZ23, Roe25]. better-directed [Roe25]. Between [Nag21, Pre99, BGJ01, LSV08, Rot91, SN94]. Beyond [BdFED^{+}20]. bible [Blu08, BWN08]. Bibliographic [AC75, CPW88]. Bibliography [Bee13, RH81]. bidirectional [KH06, KST16]. big [Sch81]. BIIIG [PJMR14]. Bilbao [DMVT13]. billion [SWW^{+}12]. Binary [BFKL13, GHLW15, Ayc15, BPMA02, Bra90, Gef03, GG13, GS06, KBN09, LMT16]. Binding [LSTW^{+}17, Wes97]. binoid [HWJ03, HSJ04]. binoids [HSJ04]. Bioinformatics [Gon02, KPP19, SJNS19, SNB^{+}19]. Biological [CL94, LJH^{+}17, Lut02, MSS^{+}19, YDDB15, ECSS88, HIEH22, IHEH23, NR02]. Biology [Gus97, Lab12a, Lab12b, Val09]. Biometrics [Rei03]. Bipartite [Eke95, FS13]. birthday [FvGGM90]. bisimulation [BCG07]. Bit [CF06, CHL14, GF08, HT17, HN02, HN05, KKK11, Kül10, LKM23, Mye98, NW13, NR98, NH11, RD17, TBS06, TLS16, WSVS22, ASM17, AAK^{+}09, CL09, GS93b, GS93c, HFN05, Hyy08, KVX12, KIH15, Mye99, NR00, SBB19]. Bit-Coded [RD17, NH11]. Bit-Parallel [CF06, CHL14, HT17, HN02, HN05, Kül10, NR98, GF08, TLS16, Hyy08, KIH15]. bit-parallelism [ASM17, HFN05, NR00]. Bit-Serial [WSVS22, GS93b, GS93c]. Bit-Split [KKK11, TBS06]. Bit-Vector [Mye98, NW13, Mye99]. Bitcoin [VD17]. bits [AEK^{+}11]. Bitstream [Bur11]. Bitwise [KF91]. Bivariate [Liu14]. biXid [KH06]. BIXSAN [VS11]. Blackboard [DJ96]. BLIM [Kül10]. Blindfolded [EF13]. Block [LT97, Chr96, HFFA09]. block-interchanges [Chr96]. blockchain [SN24]. Blocked [FTJ95]. Blocking [Bon07]. blocks [MBH20]. Bloom [MA12, ZS17]. Bluetooth [LTL04]. BNF [Man06]. board [AAB^{+}86]. Body [JM85, BWG12]. body-worn [BWG12]. Bolt [WZL^{+}23]. bonds [PS89]. BonXai [MNNS12]. Book [Ano97a, Ano97b, Ano12, Gue87, Hig95, Hum97a, Hum97b, Neu10, Pux97, Sal01, Tal81, Uma97]. Bookshelf [BF97, Lut02]. Boolean [SJ13]. Boomerang [BFP^{+}08]. Boost [BBvdM21]. Boosting [CDL^{+}15, LLW^{+}15]. both [JDXD13]. bottom [BDB90, CPT92, Cha87]. bottom-up [BDB90, CPT92, Cha87]. Bound [BG92, GK86, Les94, BCKM15, BG91b, GZ10b, Ind98, Lif03, LP08]. boundaries [SHS14]. Bounded [Dur94, LLC17, NR03, SD95, BFK^{+}03, CR95a, FK96, KNT11]. Bounds [CM08, Col94a, CHPZ95, CH97a, GG91, GG92, GHW05, KS94, BCT93, CJPS13, Col90, CH92, SASU13, Shi97, SV87]. boxed [CNS18]. boxed-mesh [CNS18]. Boyer [Tal81, AR13, AG86, BYR92, BYCG94, BYBDS09, BPMA02, Ber00, BKM95, CFG12, Col90, Col94a, DR06, EMC96, Gal79, Gol90, GO80, HA90, Kau92, KP96a, KP96b, KBN09, Kun95, Lec92, Men89, NT05, Pie95, Rai92, Rus85, Rus92, Ryt80, STK10, Sch88a, Smi94, Sto02, Tak96b, TU93, TJMC20, WW03, WW93, Woo86]. Brain [TMV^{+}01]. brains [Joh01]. Branching [Dur94, BK86]. Bridging [BGJ01]. British [ACM92d, ACM08, MG94]. Browser [NTS93, VS11]. Bruijn [Sri88]. Brute [GHK14]. Brute-force [GHK14]. BSR [Sto96]. BST [SCFC94]. BUFFERS [Duf82]. bugs [WBS22]. Building [BC06, GHK^{+}91, SK17, DF00]. Burrows [Neu10, ABM08, DGG^{+}19, ZMAB03]. buses [KRL87]. business [BJK^{+}12, FvGGM90, PJMR14]. buts [Edw07]. BWT [AMB^{+}02, BPMA02, FBMA05]. BWT-transformed [AMB^{+}02]. Bypassing [PLT14]. byte [TMK^{+}02].

C [Gue87, Sal01, Bar84b, Car77, Eck89, Fre78, GS93b, GS93c, Hol84, KOI94, Mad01, Mis89a, Mis89b, OK94, Sab76, Sed90, Sed92, SDS14, Str13, SM04, Stu07, Wea94]. CA [Ano97a, CG94b, KP15, ACM92b]. Cache [BTTF02, FV16, HLS^{+}11, TLC15, FGG^{+}08]. Cache-Oblivious [FV16, HLS^{+}11]. cache-obliviously [FGG^{+}08]. caches [HFFA09]. Caching [PS10, RT17]. CADE [Bun94, Kap92]. CADE-11 [Kap92]. CADE-12 [Bun94]. Calculation [SD95]. calculi [Dow91, Dow93]. Calculus [For02, CFK07, Tej20]. California [ACM69, ACM86, ACM92c, ACM93b, ACM95a, ACM95b, HM96, IEE93, Sto92, IEE98, USE92]. Call [Jon07, MCP17, Roe25]. Call-pattern [Jon07]. calls [FF08]. CAM [WSVS22]. Camera [LT90b]. Can [Cal00, Cox07, DKA^{+}15, JGMP22, Sch91a, Sch91b]. Canada [ACM92d, ACM94d, ACM08, A^{+}08, GS00, LL08, MZ07, Lev95, MG94]. Candidates [MUHT96]. Cannot [JL96, LY86, PW93, EMT23, GGL94, JL93]. can’t [LLS^{+}20]. Capabilities [Cal00, Fri97b]. Captions [GR96]. Capture [Sch22]. Capturing [MCF^{+}11, MCF^{+}14]. Car [KK02]. Care [Aku94, Aku95]. Cares [BL94, CEPR10, KR97, MBY91, NR17]. Carolina [ACM93a, IEE89]. Carte [TL12a, TL12b]. Cartesian [F^{+}23, KH25]. Cascaded [GC01]. CASCON [BGG^{+}94]. Case [BLP18, CCL87, FTJ95, Gal79, JM85, PV91, SP16, Shi92, Bar22, BG91a, BGWXP22, Duf82, Fen01b, GF08, HKN14, KT90, Pep91, Quo92, SCFC94, SKS96]. Cases [ALLL98a, ALLL98b, BAC06, PPPdG20]. Categorical [LT90a, TG96]. Categories [Kul11]. Caucasus [Sal01]. CAV [KP15]. Cell [AK09b, Sca11]. Cell/B.E. [Sca11]. Center [ACM89, Bao93, Hoa77]. Central [Liu14]. Centric [TLC15, BYK22]. Certified [RD17]. CGI [Han01]. CGOOD [TG96]. Chain [MNS10, LBK08, SMT^{+}86]. chaining [KFG15]. Chalmers [AJ89]. change [Joh94a]. changes [UIM22]. channels [AGH^{+}17]. Chaos [ZGY^{+}16]. Character [Car77, CLP95, Dav73, HZ13, HH93a, HK77, MBP22, MR11, NR03, PRT^{+}16, TMK^{+}02, TJD^{+}17, Wol86, CT96, HH93b, HH93c, LS09, Per94, Vin77a, Vin77b]. Characterisation [KST12]. Characteristic [ISNH94]. Characteristics [HH83]. Characterization [HEWK03, BCG07, IMS97, VW11]. Characters [Aku94, D^{+}23, FMP20, GH82, Aku95, Mha05]. Charleston [ACM93a]. Chart [Mu 95, MuT95, Mun95]. Charts [GM02]. checker [GJS20]. Checking [FF08, Gin67, LSWP19, RHR^{+}21, CGPS13a, HN00, MW94, NRO12]. Chemical [Les79]. Chicago [ACM06]. China [ACM07, B^{+}02]. Chinese [GWvG10, GH82, HZ13, THQ19, XPZ^{+}23]. chip [CDC96, SV87]. Chips [TLC15]. Choices [LD10, LS99]. Choosing [LLL12]. Chord [YJ84]. Chromosome [KS94]. Church [KKM^{+}85, Sha88a]. CIP [BW91]. CIP-Project [BW91]. Ciphers [PP06]. Circuit [IMM^{+}22, PM78, TT20, ADM^{+}13]. Circuits [Brz64a, Brz65, FU82, GHKL18, KBB01, Pie95]. Circular [CKP^{+}21, CHL14, HT17, LA12, Boo80, HOK18a, HOK18b]. cities [Joh01]. ClamAV [OWP16]. Class [CGS17, EU98, Kin92, Kul11, SA96, Sch12, Sch13, BAC06, BRO16, CRV06, Kod79, Pie08, Wal89]. Classes [HHW^{+}99, NR03, Lei85, LS09]. Classic [HSTS01, RB05, MAC14]. Classification [Bon07, BYHT18, Lee91, TD18, WZJH12, XPZ^{+}23, LMT16, TZH^{+}13, WXZY12]. Cleaning [FKRV16, QTO^{+}20]. Clients [CDM11]. CLIs [YC22]. clone [DNR06, Joh94a]. Closed [Kul11]. Closed-Class [Kul11]. Closest [KF91, Sut21]. Closure [LMN16, AS85, Jed87, Lee82, LH03]. Cloud [CFM17, CDM11, SOR16, KLL23]. cloud-edge [KLL23]. clouds [SCF^{+}17]. cluster [MM03]. Clustering [CZW15, LSTW^{+}17, WMGS19, KAT07]. Clustering-Based [WMGS19]. CMA [ZL18]. Co [CDPP23, CA18, CA20]. Co-lexicographically [CDPP23]. Coarse [FL99, TTO^{+}22]. Coarse-Grained [FL99, TTO^{+}22]. Cobol [HHW^{+}99]. Coconut [AK09b]. codata [TCP16]. CODE [BY91, AGT89, Cox12, Fra83, GFH82, GHF83a, GHF83b, Gie90, ND02, RTT02b, SED14, VSM87, WHZ^{+}17, WNL^{+}83, ZWH^{+}21, AG06, BDB90, BY92, BY96b, CLS95, FHP92, Gan89a, GHS82, HV93, MSRR00, MR82, NAR08, OW07, Rém17, UIM22]. code-generator [FHP92]. Coded [BG95, Chu95, RD17, BC95, NH11]. Coder [MP88]. Codes [YK11, Bra90, Mei08]. Coding [CW84, Dav73, JSC83, Kid09, DV21, Ind97, MP88, Shi97]. Cognitive [PW06]. cohabit [Wad87]. coinductive [HN11]. collaboration [Roe25]. Collage [IST05, KMT^{+}01, KMS^{+}03]. Collections [Nav21b, Nav21a, BC13a, CHLS07, CMRV10, FGG^{+}08, HAI02, WL15a]. Colony [ACM83]. Color [Hui92]. coloring [FS13]. Columbia [ACM92d, ACM08, MG94]. Column [SP16]. Com [Lia84]. Com-pu-ter [Lia84]. Combating [KEG^{+}08]. Combinations [Knu05]. combinator [Sta89]. Combinatorial [Ano17, BM08, Cro92a, GIMV03, Mei08, SLTB^{+}06, Val09, WCM^{+}94a, WCM^{+}94b, CDDM05, HLN09, PPPdG20, AL01, AP10, Ano92b, Apo92, Apo93a, AH97, AT02, ACP05, BYCC03, CG94b, FC98, FL08, GU95, GS00, GM11, HM96, KS12a, KU09, LV06, MZ07, PC99, SMD04, Lab12a, Lab12b]. combinators [LT90a]. Combining [Ber00, JA17, JXA20, HBRV10, NR00, THQ19]. Command [Roo99, YC22, Blu08]. Command-line [YC22]. Commands [Lud77]. Comments [Akl78, ZZ12, Gro91a, THQ19]. commerce [ZCT14]. COMMERCIAL [BY91, BY92]. Common [Ale94, ACR20, F^{+}23, FR17, HIRS17, IF94, JN24, KRS19a, KRS19b, KRS23, LJZZ13, DK13, FGKU15, Gra15, Maa06, Mid98, MBH20, Rou21, TU88, TTO^{+}22, Mu 95, MuT95, Mun95]. Commonwealth [ACM89]. Communication [Ayc15, Bao93, HRN^{+}15, HSL10]. Community [LYT^{+}23, LGZ^{+}14]. Commutative [Eke95, HY92]. Compact [Asp12, HAR10, NR01, Ric79, YP12, ZZH16, BFC08, DGM19]. CompactDFA [BBHK14]. compaction [NE93]. compaction-based [NE93]. Comparative [JM85, MSZ17, PSK08]. Comparator [Bur84, Bur82]. Comparing [Hoc19, Hua94]. Comparison [BCT98, JTU96, Lav91, de 82, Bar84a, BCT93, CT96, ECSS88, FBMA05, HA90, KB22, SVS97]. Comparisons [Bre93, CL92, GPR95a, Liu86, Bre96, PW06]. Compatible [Anoxx, LT09]. Competitive [DV21]. Compilation [FU82, KTU87, Ses96, AP90, Dan91, HF13, KGP^{+}05, Sch88b]. Compiler [AJ89, GFH82, Pet92, vNG01, CGZ^{+}13, FKSB06, HWF90, Jør92]. compiler.kit [Abb77]. compilers [BGNP94]. Compiling [AU72, AU73, PS93b, Sch99, GHR^{+}16]. Complement [GN12, Rob79]. Complete [Ano68, BBH^{+}87, Pet02, Sch14, Kin91]. completeness [TCC91]. Complex [Gor00, ZL18, LG16, LR14]. Complexité [Alb89]. Complexity [ABBH^{+}16, BKLP97, BKL^{+}02, BDFW94, BCT94, BCT98, Col94a, CHPZ95, CH97a, EZ74, EZ76, EMTG23, FMMS20, GG91, GG92, GK86, GH15, Hei01, HK11, HU92, HSTS01, KLH16, MNS10, Mor83, NRS18, Prů17, RS98, Akl78, Alb89, AK12, Bar22, BDM19, CGK08, Col90, CH92, CGG90, CS22, FCFM00, FK96, Hun79, KS07, Lei81, LM12, LM13, LMN16, Mag81, Man76, NF04, PS89, PAG09, Sal12, Via04, XH23, Yao79]. components [CFM00]. Composite [XK92]. composition [AGH^{+}17, SV09]. Compositional [LN19, GJS20]. Comprehension [BLS^{+}94]. comprehensions [SVMM17]. Compress [GH82]. Compressed [ABF96, BR09, BA16, BKLP97, BKL^{+}02, BGS23, CHLS07, CLS^{+}10, CHP92, FT98, FV16, FT04, GP01, GP03, Gaw12, Gaw13, GV00, GV05, IST05, Jez15, KTSA99, Kid09, KS05, KS06, LSW08, Loh10, Man94, Man97, MHT09, MHM^{+}01, NR99b, Nav01c, NM07, Nav21b, Rao95, RNOM09, STSA99, TMK^{+}02, YK11, ZMSD93, ABF94b, BCD98, BPMA02, BFG09, BBK12, CP97, FT95, FGG^{+}08, GR99, GO12, HHLS06, KTS^{+}98, KMS^{+}03, NKT^{+}01, NT05, SNZBY00, SLZ^{+}20, SMW^{+}23, TM04, TM05b, TM05a]. compressible [BFKL13]. compressing [WL15a]. Compression [ABM08, BC13b, BK93a, CW84, FG89, GS85, How97, LS94, Man94, Man97, Neu10, RPE81, RT17, Sad96, SKF^{+}00, SC93, SC95, SC96, SC98, SC99, SC01, SC02, SC03, SC04, SC05, SM09, SM10, SM11, ASG99, AGS96, BFNP10, Cha93b, CDC96, CL96, How96, Lar99, OW03, QZC17, RTT02a]. COMPSAC [IEE95b]. Computable [KT23, EH88]. Computation [Bro93, COZ09, Cha86, HU79, HMU01, HMU07, Hua98, Lev95, Ng79, Rao94, SXZ^{+}25, WN90, CCI^{+}13, Far19, Han02, Maa06, NHN^{+}20, NA90, PS93a, QZC17, SY23, SASU13, Sid00, Tak96b, YT03, ACM94b]. Computational [Gus97, Lab12a, Lab12b, Tal81, HN11, Val09, Via04]. Computationally [HT14]. Computations [FKP77, CR91, NEH90, Pra97, PCS99]. Compute [MR11, MS95]. Computer [ACM89, AHU74, Bao93, Cop91, FJ92, Gus97, Hea71, Hwa85, CVP86, IEE89, IEE90, IEE92, IEE93, IEE95a, IEE95b, IEE97, IEE98, IEE09, KL02, Knu05, Kül10, RJK79, Ruc15, SS93a, Coo86, Fat15, II09, Ker04, SS94, VVV04, Win78, iA94, KP15]. Computer-Recognized [RJK79]. Computers [FL99]. Computing [ACM69, ACM74, ACM76, ACM81, ACM84, ACM86, ACM90b, ACM91, ACM92d, ACM93b, ACM94d, ACM95c, ACM97c, ACM99b, ACM00, ACM08, Ano13, BGK^{+}16, CZOdlH17, CFM17, CZ01, Cha94, DT87, DGBH93, FYJ^{+}17, Fra20, HM98, HM87, ISNH94, LK90, Rot91, RW10, Wol90, BGNP94, BC95, IP96, LK88, ZYQ^{+}15]. Concatenation [CGS17]. Concave [KM92, KM95a]. Conception [Hud89]. Concepts [BGJ01]. Conceptual [BK75]. concise [BNSV10, NdMM02a, Yod91]. concrete [JD89]. Concur [SBF80]. Concurrent [GR92, Pel87, SBF80, BFN^{+}09, Gol90, JM90, Rus92, YT03]. condition [Han92, KT90]. Conditional [DJ96]. conditionals [Edw07]. Conduct [NCKL14]. Conference [ACM89, ACM92c, ABB93, AGS93e, Ano87, AAC^{+}01, AOV^{+}99, Bao93, B^{+}02, Bun94, DMVT13, FMA02, CVP86, IEE94a, IEE94b, IEE95b, Kap92, KP15, MG94, NH11, SW94, Sto92, SC93, SC95, SC96, SC98, SC99, SC01, SC02, SC03, SC04, SC05, SM09, SM10, SM11, USE92, DT87, HF13, ACM69, ACM74, ACM76, ACM81, ACM92a, ACM93a, ACM94a, ACM95a, AGS93a, AGS93b, AGS93d, AGS93c]. Conferencing [Sch95]. Configurable [ACF05]. Configuration [Sch95]. Conflicts [YD95]. Congress [FGR72]. congruence [KKH24]. Conjunctive [CDL08, FLS98, Sch22]. connected [Joh01]. connectivity [Sri88]. conquer [KD15, SW12, SHCY93]. consecutive [KKR^{+}13]. Consensus [BDFW94]. Considered [Sym85]. Consistency [ZCS^{+}12, AL08]. Consistent [PW93, MAI^{+}16]. Constant [ABMN20, BGG12, CGG^{+}97, CGH^{+}98, CGR99, Gal95, GPR95a, KRR17, Sto96, BGM13, Gal92]. Constant-Delay [ABMN20]. Constant-Sized [KRR17]. Constant-Space [GPR95a, CGR99, BGM13]. Constant-Time [BGG12, CGG^{+}97, Gal95, Gal92]. constants [KC11]. Constrained [CS11a, CLT07, Jan23, NT20, TTO^{+}22, XJT^{+}04, ZJL14]. Constraint [Coh90, RHR^{+}21, CFK07, Smi91a]. Constraints [BGM19, CCH^{+}23, GRS99, ZGS^{+}15, CDL08, CFLH^{+}22, ETV21, HW09, KS11a, ZXL^{+}13]. Construct [DKA^{+}15]. constructability [Kar82]. Constructing [IY02a, IY02b, Lei80, JRV96, TU88, TTHP05]. Construction [BP11, BP63, EJ23, McC76, MOSZ18, BH96, DKP11, FCFM00, Kos94, Mei08]. Constructions [Ant95, MSZ17, Ant96, Che96]. constructive [Tak96a]. constructors [MME14]. contact [KD15]. Containing [HJ99, CFM00, FSL^{+}15]. Containment [FLS98, CDL08, HN11, SH85]. Content [PW12, Agu23, LMT16, MLC08, TLLL09]. content-based [MLC08]. Context [CK02a, Haz01, Hua94, Kea91a, SBHM94, SA96, KGA^{+}12, Mye95]. Context-Free [SBHM94, KGA^{+}12, Mye95]. Context-Sensitive [SA96]. contextual [JGMP22]. Continental [Bao93]. Continuous [SBF80]. contraction [KPP21]. Control [Bao93, HHW^{+}99, Hoa77, Mu 95, MuT95, FWDL15, Mun95]. Controlled [NAR08]. Convention [ACM89, Bao93]. Convergence [Sid95, Sid99, Sid02]. convergent [Sid95]. Conversion [GJ16, HZ13, Hoa77, Lau00]. convert [ZC99, CM86]. converting [Gal75]. convolution [Ind98]. Convolutions [ALR08, Zha17]. Cookbook [ST03, GL12, Ano12]. coordinated [Mid98]. coordination [CFM00]. copattern [RTO15]. copatterns [AP13, APTS13]. coprocessor [TLLL07]. Copying [JLH18]. Coq [MPdS12, VLP17]. Corasick [CW13, NK07, PLL10, RKM21, TM05b, TZH^{+}13, TVCM12]. Core [LY17, TLC15, JA17, JXA20, MAC14]. Cores [LSTW^{+}17]. corpus [IIK08, II08]. corpus-based [IIK08, II08]. correct [Ryt80]. Correcting [Kuk92]. Correction [And02, Bur84, HOK18a, JP73, KRS19a, KRS23, RJK79, Wag74, BSY00, Mae90, MS95, TIAY90]. CORRECTNESS [BY91, Col91, Sto02, BY92, SBR^{+}07]. correlated [SWZ01]. Correlation [KC99, Sha93, WZJH12, PPZ08]. Correspondence [BYKZ^{+}92, Spi99b]. corresponding [Lif03]. Corrigendum [FLSS93a]. Cortical [TMV^{+}01]. cosine [TP07a, TP07b]. Cost [KLL23, WZL^{+}23, SASU13]. Cost-Efficient [WZL^{+}23]. costs [PW06]. Count [MR11]. Counter [WPKL13]. Counter-Based [WPKL13]. counterparts [BdFED^{+}20]. Counters [LT09]. Counting [CGS17, GGM12, San95, CGPS13a, Gel10, Nic03, THL^{+}20]. counting-set [THL^{+}20]. coupling [All89]. course [GHR14, HR00]. coverage [GJS20]. Covered [Yun12]. Covering [BNV^{+}13, CIK98]. covers [IP96, MS95]. Covert [HL10]. CPM [AL01, AP10, Apo93a, AH97, AT02, ACP05, BYCC03, CG94b, FC98, FL08, GU95, GS00, GM11, HM96, KS12a, KU09, LV06, MZ07, PC99, SMD04]. CPU [LLC17, ZCH23]. CPU/GPU [LLC17]. crash [GHR14]. crash-course [GHR14]. CRCW [Apo93b]. CRCW-PRAM [Apo93b]. CREW [dB93]. Critical [Lut02]. Crochemore [Bre93, Bre96]. Crochemore-Perrin [Bre93, Bre96]. Cross [AS91, FTJ95, KNS12, Sha93, LGZ^{+}14]. Cross-Correlation [Sha93]. Cross-Document [KNS12]. Cross-Interferences [FTJ95]. Crossover [Sut21]. Crowd [CDL^{+}15]. Crowd-Sourcing [CDL^{+}15]. Cryptanalysts [GS93a]. cryptography [DA18]. CS3 [Kah07, Kah09]. CS3/CS4 [Kah09]. CS4 [Kah09]. CSV [AMRV16]. CSV-like [AMRV16]. CTL [MMDdJ11]. CTRL [MMDdJ11]. Cube [ML96a, Dow91, ML96b]. Cubes [CR95c]. Cuckoo [TK07]. CUDAs [KMM15]. curly [HSW97]. curve [SA77]. Curved [LT90b, NA90]. Curves [HHW^{+}99]. Custom [Han01, Vol12]. Cut [AM97, Lud77, CK04]. CVPR [CVP86]. Cycle [KK95]. cycles [Ste12]. Cyclic [KPP21, Mae90].

D [SHCY93, ASG99, BSM^{+}07, BZ98, CJ93, KD15, LT90b, Mun07, TCCK90, ZLN11]. D-pattern [ASG99, KD15]. DAGs [ZZ12]. daisy [SMT^{+}86]. Dallas [IEE95b, NEH90]. Damaged [SP16]. Dark [Hum99]. Darmstadt [AGS93e]. Data [ABM08, AAC^{+}01, BLLW12, B^{+}02, BGNV10, Bon07, CBD^{+}23, CW84, CMNP17, DT87, EF13, FG89, FO76, FBY92, FMA02, FMG23, Gia93, GG97, Gon83, GS85, Har02, JDXD13, KM94, LSW08, LKL02, LM01b, LMV16, LS94, MBP22, MMS14, Neu10, PYY19, Pre99, RPE81, Sad96, SD95, Sli78, SW94, Sto92, SC93, SC95, SC96, SC98, SC99, SC01, SC02, SC03, SM09, SM10, SM11, SOR16, TV14, VMML15, WCM^{+}94b, YDDB15, ABB93, AL08, AMRV16, BGHZ15, BFP^{+}08, BFS00, BC93, Cha93b, CDC96, CDP16b, CD96, FG95b, FG99, FMG22, GMC02, GS22, GW92, GS93b, GS93c, GPN96, GHS12, GS06, HN90, HSL10, HH16, HF13, JO97, JD89, Kra08, Lar99, LWS^{+}16, LRV13, LTV15, MRA^{+}17, MF96, MRR^{+}18, Nil90, ORPF13, OR11, OSSK16, PJP^{+}18, PJMR14, QTO^{+}20, RW93]. data [RM06, SMS15, SG16, SMW^{+}23, TSI13, TG96, VR18, Wad87, WCM^{+}94a, ZBST14, SC04, SC05]. Data-Parallel [VMML15, MMS14, GS93b, GS93c]. data-parallelism [RW93]. Database [ACM83, ACM90a, ACM92b, ACM94c, ACM95b, ACM97a, ACM98, ACM99a, ACM06, ACM07, CCL87, HF13, HAI02, LHMH91, LL08, MNP^{+}23, SK17, VB98, GPTV93, IIK08, KLB12, Len11, OKT92, SHS14, SR16, ZCÖ09, KKP92]. Databases [AAB^{+}17, AOV^{+}99, CCH09, GNU94, Pou93, CN21, FLC^{+}19, SHvR^{+}16, WZS95, ZCT14, ZCÖZ12]. Dataflow [PK85, Ray96]. Datalog [dLFM07]. datatypes [JR15a]. Dates [SM99, HN00]. dbC [GS93b, GS93c]. DC [ACM84]. DCC [SC98, SC04, SC05, SC93, SC95, SC96, SC99, SC01, SC02, SC03, SM09, SM11]. DDC [SM10]. De-Compositional [LN19]. Dead [Cox10c, MD10]. Death [CGP^{+}08]. Debugging [MLM^{+}08]. December [A^{+}08, IEE94a]. decentralized [SMT^{+}86]. Decidability [Asp12, Kar82, San15]. Deciding [CDLM17, Gaá04, MPdS12]. decipherability [AG84]. Decision [MNS10, RS59, CS11b, HW09, TN13, TN15]. Declarations [MGH93]. Declarative [ADR15, FKRV16, Spi99a, Spi99b, KLR^{+}08]. Decoding [Sto96]. Decomposition [AGH^{+}17, KVX12, PS93b]. deduction [Bun94, HR03, Kap92]. Deep [CMS08, LLLL08, LLC17, VWR11, YP13, ARS16, BAC12, NYuR15, PKK18, STKD20, SSYW19]. Defective [BTTF02]. Defending [AV23]. Defense [Bol02]. definability [CDLM17]. Definable [BLSS03, Cho78]. Defined [KMRY20, KMR21, PSK17]. Defining [TBD22]. definition [Yod91]. definitional [CKC07]. Definitions [IEE01a, IEE01b]. deformation [BCWG09]. Degenerate [BGP^{+}22, DW17, IMR08, LJH^{+}17, BPPR20, DGG^{+}19]. Degenerates [GP18]. degree [HY90, LSV08, YH91]. degrees [YH92]. del [ACM69]. Delay [ABMN20]. deletes [Mei15]. Demand [KBB01, FWDL15]. demand-driven [FWDL15]. Demystifying [WBS22]. denial [PM23]. Denmark [AH97]. denoising [CDDM05]. Denotational [Gud92, Mal93]. denoting [HSJ04]. Density [Bur11, Sel84]. Department [Bol02]. Dependence [KK95, AFM94]. dependences [BD98]. dependencies [Cox19, QTO^{+}20, Rai99]. Dependent [CD18, Hua94, CFLH^{+}22, CDP16a, CA18, CA20, Jan23]. Dependently [Xi03, CDP16b, Tej20]. depth [Alb89]. Derivation [Mis03, BGJ89, HR01, Pep91]. Derivative [MNR^{+}23, SL17]. Derivative-Based [SL17]. Derivatives [Ant95, Brz64b, CJM12, CDJM15, SL14, ST19, Urb23, Ant96, ADU16, BRO16, ELF22, KMR21, ORT08, ORT09, Sid00, TN13, TN15]. derived [PS90]. Deriving [TA90]. derivors [Gie90]. Described [KPR97, KPR00, SA96]. DescribeX [CMRV10]. Describing [Pie95]. description [CDL08, FKSB06]. Descriptional [GH15, LMN16, XH23]. Descriptions [KRS95, KRS97, SC88]. Descriptive [GR96]. descriptor [All82]. Design [AHU74, BK75, BCD14, Bir10, Bur84, HL97, OF61, RDRF24, Ski98, VKPI17, ZA87, dLBHC22, Bur82, CMS08, Hur84, KCK93, MI07, OA17, San09, SW12, WKR09]. designers [LS99]. Designing [II09, Mor83, SB09, DC94]. Desktop [Rob99a, Rob99b]. desukutoppu [SM04]. Detect [OWP16]. Detecting [GN19, Mut97, TMV^{+}01, Mar89, Mha05]. Detection [ARM^{+}19, BYHT18, CZCD09, GKW^{+}10, KKK11, KNMH00, Lee91, LY17, Les79, PAMP12, Sli83, TS05, WWW^{+}16, ZL18, ZLN11, ACF05, BKLE18, BAC12, BCD14, DNR06, FNP09, GZ10a, Joh94a, KOI94, KAT07, LHCK04, LGZ^{+}14, OK94, QLY07, SA77, TBS06, VD17, Hig95]. Determining [ZZJC20]. Determinism [GGM12]. Deterministic [BGNV10, BP11, BKW92a, BKW92b, BKW92c, Coo72, FS19, GZ94, GMS12, Ind97, KV15, LY86, Nag21, Ned98, TLLL07, VW11, Vis90, Vis91, XLC19, BS86, BBM21, CDLM17, GHR^{+}16, GM17, Lau00, LMN16, NdMM02a, SV09]. develop [DV21]. developers [LS99]. Developing [MNNS12]. Development [HR01, JLHB92, LHCH93]. Developmental [YCJK08]. Developments [PV91, OKT92]. Devices [HAR10, Sym85, CMS08, Pet07]. DevReplay [UIM22]. DFA [BC13b, CP97, NYuR15, NR01, PW93, TJD^{+}17, VWR11]. DFA’s [CHP92, FDG^{+}11]. Diacritic [JZAA19]. Diacritic-Based [JZAA19]. Diagnosis [SL17]. Dictionaries [Bre94a, Owo93]. Dictionary [And02, Bre94a, Bun95, Gue87, KW19, KS11b, KS12b, AF92, DLF^{+}15, SMW^{+}23]. dictionary-based [DLF^{+}15, SMW^{+}23]. Dictionary-Matching [Bre94a]. Diego [ACM92b, ACM93b, Sto92]. differences [CN21, LV88]. different [YB13]. Differential [FDG^{+}11, PS93a]. Differentiated [ZLN11]. differentiation [PS93a]. Diffusion [MSP^{+}17, Sal01]. Digest [CZCD09]. Digital [AGS93a, AGS93b, AGS93d, AGS93e, AGS93c, JRV96]. Digitized [LV94]. Digree [SK17]. Dijkstra [FvGGM90]. dimension [Bak78]. Dimensional [ABF94a, ADLM96, BYN98, BKLP97, BKL^{+}02, CDEK95, CL95, CHLT14, CR92, CR95b, CGPR95, CGH^{+}98, CIK98, FU98, FNU02, GPP04, Gia93, GG95, GG97, HEWK03, HW12, KPR97, KPR00, KU99, KR94, Les79, Les95, Par96, Prů17, ZT89, AK08, AF92, ABF94b, ABC^{+}04, AKT06, AGM05, ADLM01, BYR93a, BYR93b, Bir77a, CGK08, CGR93, CR94, GP92, HLN09, JKNS00, KM13, Mid96, MPW21, NBY99a, Par98, TIT83, WC14, XMLC11]. dimensions [CCG^{+}93, dRL95]. Directed [Fu95, Fu96, Fu97, Gud92, Kor83, BÖ13, Dan91, Fil21, Nil90, Roe25]. Directly [Man94, Man97]. Directory [ZJP^{+}18, Zve80]. Disambiguation [OS11]. Discontinuities [Lee91]. discourse [Kit94]. Discovering [LSTW^{+}17, LSWP19, SW93]. Discovery [CLST^{+}13, VG01, WCM^{+}94b, MP05, TSI13, WCM^{+}94a, WZS95]. Discrete [ACM97b, Gim73, KC99, Nak14]. Discrimination [KC87]. Disease [TMV^{+}01]. Disease-Specific [TMV^{+}01]. Disjoint [LS10, YD95]. Disjunctive [HR03]. Disk [JDXD13, WHZ^{+}17]. Disk-based [WHZ^{+}17]. disordered [CGM10]. dispatch [MFRW09]. dispatching [FMdB99]. Display [Rei77]. Display-Oriented [Rei77]. Distance [BCP02, CZOdlH17, D’A98, FMP20, KS94, Med23, Ris16, RKH02, TT20, ZCÖ09, AEP06, AK12, AD11, BC95, CM07, EH88, GF08, HKN14, HOK18b, Leu97, PPZ08, Rot91, SKS96, TSI13, ZXL^{+}13, HOK18a]. Distance-join [ZCÖ09]. distances [AAB^{+}09]. Distorted [VB98]. Distortion [KC87]. Distributed [AYS^{+}24, ASA17, BYNTK21, CPW88, IMR08, ML96b, NCKL14, PSP^{+}18, SK17, TVCM12, ZLN11, AC93, DSv94, FWW12, HRN^{+}15, HFFA09, LQL^{+}16, LLC03]. Distributed-Memory [TVCM12]. Distribution [MR11, YJ84]. divergent [Sid95]. Diversified [FWW13a, FWW13b]. Divide [KD15, SW12, SHCY93]. Divide-and-conquer [SHCY93]. Dividing [KKK11]. DNA [BEL17, CLST^{+}13, CN21, HAI02, IMR08, KYG19, KB22, LSTW^{+}17, MT14, NEL17, RLP20, TP97, YT03]. DNA/RNA [IMR08]. Do [JL96, Abb77, GGL94, JL93]. do-it-yourself [Abb77]. d’objets [Alb89]. Document [ABMN20, BK93b, BKW92d, DS19, FKRV15, FRU^{+}20, KNS12, LMNT16, All82, Arn93a, Arn93b, BK93c, BK93d, BK93e, KRML09, WZS95, WCW82]. Domain [CF85, CDC^{+}23, GÁSÁ^{+}13, PIR17, SKS96, SHvR^{+}16]. Domain-Specific [CDC^{+}23, SHvR^{+}16]. Donald [Neu10, Ruc15]. Done [LY86]. Don’t [Aku94, BL94, Aku95, CEPR10, KR97, MBY91, NR17]. DOT [BGWXP22]. doubling [CL09]. down [GOMSJVGP08]. DPI [ABBH^{+}16]. DRAM [WSVS22]. DRAM-CAM [WSVS22]. DREAM [HRN^{+}15]. DReX [ADR15]. Drive [KK02, BC06]. driven [BCD14, FWDL15, GHS82, Mus03, NWE99, Sak21, TJGY22]. Drosophila [YCJK08]. Drum [SP16]. DSL [BCD14]. DSL-driven [BCD14]. DTDs [BNSV10]. Dublin [ABB93]. duel [JHU^{+}24]. duel-and-sweep [JHU^{+}24]. Duration [XJT^{+}04]. Duration-constrained [XJT^{+}04]. during [Sch81]. Dynamic [AGT89, ALLS07, ACR20, BSM^{+}07, BFNP10, CWL^{+}21, CL95, Mye98, Sch95, WBA83, ZLN11, ADM^{+}13, BD98, CHLS07, CGM10, FG95b, FhDAF09, HSL10, JSH09, KT14, LYWL08, Mye99]. Dynamics [JM85, MSP^{+}17].

e-commerce [ZCT14]. early [San09]. East [Sal01]. EBWS [KPP19]. ECG [TZH^{+}13]. Economical [McC76]. Ecosystem [AWD^{+}18]. Edge [FS13, BYK22, KLL23]. edge-centric [BYK22]. Edge-coloring [FS13]. Edinburgh [AOV^{+}99]. Edit [CZOdlH17, JWZ94, Med23, RKH02, AEP06, AK12, BC95, CM07, Leu97, LT97, QWX^{+}13, SKS96, ZXL^{+}13]. Edited [Ano97b]. editing [DOS93]. edition [Ano12]. Editor [Pik87, Pik00, Rei77, Ritxx, Ano17]. Editorial [AGS93a, AGS93b, AGS93c, AGS93d]. Editors [Dav82]. Edsger [FvGGM90]. EDTV [Rei77]. education [Ker04]. effect [Mha05]. effectful [SVMM17]. Effective [AG06, FKSB06, GN19, XLC19, ZGS^{+}15, KC11, PC02, ZKA12]. effectively [ADT15]. effectiveness [BSY00, DNR06]. Efficiency [ALR08, Col91, San15]. Efficient [AC75, ACR01, ALV92, ALLL98a, ALLL98b, Apo93b, BKLE18, BDB90, BC13a, BA15, BC13b, Ben94, BBH^{+}87, BYHT18, BT21, BYK22, Bra94, BC94, BG95, CF06, CCF13, CFP19, COZ09, CGR02, CDDM05, CCI^{+}13, CLT07, CGPS13a, CDC^{+}23, CR95c, DA18, DGG^{+}19, EMC96, FMP20, FMMS20, FRU^{+}20, FT04, FM06, GC01, GP01, GP03, Gaw12, GLS07, GWX^{+}23, Gon83, GM17, GMAS22, Gue90, GZ10a, GS06, HHM^{+}13, HL10, HH16, HW12, KR81a, KR81b, KR87, KRS97, KS94, KT14, KKK11, KRR17, Kos89, LV86a, Lau01, LP13, LKL02, LOZ^{+}24, LTL04, MK90, MC24, MHT09, MOSZ18, NHN^{+}20, NWE99, NdMM02a, NK07, Ode14, Owo93, PAMP12, PDC94, QLY07, RDRF24, SRV^{+}19, SRK20, SA96, SOR16, SvS14, SWW^{+}12, SLZ^{+}20, SMW^{+}23, TZYH14, TJGY22, WZL^{+}23, XQW^{+}13, Yun12, ZJP^{+}18, ZXL^{+}13, AB09]. efficient [Aoe89, CPT92, CGR03, CLZ^{+}15, CW13, CD96, Cox10b, ESL89, FNP09, Fil21, FHP92, GPR95b, GL89, GLS92, LV86b, LVN87, Lee82, Maa06, MRA^{+}17, NAR08, PLL10, QWX^{+}13, SY23, TJD^{+}17, VD17, YKGS11, YB13, YHV^{+}15, ZCT14, ZYQ^{+}15, ZKA12, ZYX^{+}12]. Efficiently [ADR15, DF00, ADM^{+}13, Kim99]. effort [LG16]. egrep [Woo87]. Eighteenth [ACM86, ACM99a]. Eighth [ACM97b, B^{+}02, ACM76]. Einführung [HU92]. Elaborating [CA18, CA20]. Elastic [BGP^{+}22, BPPR20]. Elastic-Degenerate [BGP^{+}22, BPPR20]. Electron [DMWW77]. Electron-Beam [DMWW77]. Element [MGH93]. Elements [UFA^{+}24]. Eleventh [ACM92b]. eliminants [AS85]. Eliminating [CDP16a]. Elimination [Han13b, CK04]. Emacs [MS20]. email [WR15]. Embeddable [Fri97b]. embedded [TLLL07, TLLL09]. Embedding [BDFR08, Fu97, KKM^{+}13, ZCÖZ12]. embeddings [CMO^{+}08]. Emergence [Joh01]. Emilie [Sal01]. Empirical [CL92]. emptiness [CHL^{+}25, Kar82, Rob79]. empty [Zia96]. emulator [VVV04]. Enabling [GWX^{+}23, AB09]. Encoded [CFG12, DS04, KS01]. Encoding [HAR10, KR92, RTT02b, Yun12, FDG^{+}11, KR89, LS09]. Encrypt [BTTF02]. Encrypted [MBP22, SOR16, HH16, MRR^{+}18, OSSK16, WR15]. Encryption [VRC24]. Encryption-based [VRC24]. End [JLH18, JLHB92]. End-to-End [JLH18]. ends [ESL89]. Energy [CDC^{+}23, GMAS22]. Energy-Efficient [CDC^{+}23, GMAS22]. Engine [CZCD09, Hab04, PWW^{+}11, RVV23, SK17, SXZ^{+}25, VCS^{+}12, BC06, CW13, HRN^{+}15, WL15b, dKM04, GWX^{+}23]. Engineering [Bao93, CS18, CFKT17, FHP92, IEE94a, RMK^{+}14]. Engineers [NEH90, Lut02]. Engines [ABBH^{+}16, TBS06, ZV97]. English [Ayc15]. Enhanced [KTY^{+}18]. enhancement [BKM95]. enhancing [FSL^{+}15]. enough [MR09a]. Enriched [MSS^{+}19]. enrichment [LGZ^{+}14]. ensembles [Alb89]. entails [Kar82]. entire [YCJK08]. entity [BDMT16, DLF^{+}15]. entrophy [KS96]. Entropy [YDDB15, CR95a]. Entropy-Scaling [YDDB15]. Enumerating [McI04, PSP^{+}18]. Enumeration [ABMN20, FRU^{+}20, CLZ^{+}15, LQL^{+}16]. Enumerative [JA17, JXA20, SXZ^{+}25, Tan14]. Environment [LHCH93, LZ96, MM02, SBR^{+}07]. Epoch [OSM94a, OSM94b, OSM94c]. Epsilon [GJ16, HSW97]. Epsilon-Free [GJ16, HSW97]. Equality [Gin67]. equally [NCV10]. Equation [CZ01, COZ09, MOSZ18, NC92]. Equational [OND98, AFI98, DL03, NWE99, RW93]. Equations [HOS85b, Ver70a, ZGS^{+}15, HOS85a, KGA^{+}12, Sta89, Ver70b]. Equivalence [Asp12, HSJ04, Hir96, KN12, Liu14, SA77, CS11b, Kap69, MAI^{+}16, SH85, MPdS12]. equivalences [CDP16b]. erasure [Tej20]. Erdos [AW89]. EREW [PDC94]. Ergodic [Shi92, ST96a]. Errata [Ver70a]. Error [RJK79, FSL^{+}15, XQW^{+}13]. error-containing [FSL^{+}15]. Error-Correction [RJK79]. error-tolerant [XQW^{+}13]. errors [AAK^{+}09, AAB^{+}09, BLLP90, CEPR10, KNT11, San95, WM92a]. ESA’93 [Len93]. Espoo [GU95]. Essential [KPP19]. essentially [GHK14]. Essentials [Bal15]. Estimating [TP07a, TP07b]. Estimation [CZW15, KBB01, KC87, JKNS00, KS96, STKD20, TCCK90]. Euclidean [GK86]. Eugene [Hig95]. EUODHILOS [OSM94a, OSM94b, OSM94c]. EUODHILOS-II [OSM94a, OSM94b, OSM94c]. European [Len93]. EUROSAM [Ng79]. EUUG [Ano87]. Evaluating [ADR15, FLC^{+}19, RSG^{+}19, SSSS10, LM12]. Evaluation [BC13b, Cha02a, D’A98, GL01, Reg92, Ses96, VB98, YJ84, ADR03, ADR06, BSY00, Chi17, CD89, DR06, Hur84, Jay92, JLFL14, Jør92, KEG^{+}08, MRA^{+}17, MM03, PSK08, Smi91a]. even [LR14]. Event [CvW18, SGCW14, dH05, CK08, LG16]. Event-based [dH05]. event-processing [CK08]. Events [CEW58, Kle56]. EventScript [CK08]. everyone [Nar91]. Everything [NTS93]. Evolution [CS18, Hud89, MS20]. evolvable [LLC03]. Exact [AOK02, BCT94, CL97, CHL14, CHPZ95, CH97a, FL12a, FNU02, GG91, GG92, IMM^{+}22, KM13, Liu14, MIH17, MA12, PP09, WSVS22, ABH^{+}14, Bak78, CH92, CGG90, ĎHPT10, FL13, FS22, HTX17, KB22, Kar82, Lec07, NF04, QWX^{+}13, SSLL21, Tan14, THG17, TZH^{+}13, YHV^{+}15]. exact-match [Bak78]. examined [ORT09]. examining [BvdM23]. Example [CFCK22, Qui02, Qui00]. Example-based [CFCK22]. Examples [BDD^{+}14, Bra94, BC94, KK08, BGHZ15, GHS12, Kod79, LSO17, SG12, SG16]. Exchange [RMK^{+}14, AL08, HSL10]. excluding [MBH20]. EXE [CGP^{+}08]. Execution [AWD^{+}18, Han92, MZZ10]. exemplars [CJR^{+}21]. Exercise [Wen93]. exercises [BH07]. exhaustive [IM13, KJS17]. Exit [MOG98]. exotic [MR82]. Expanding [Ham88, VHC88]. Expansion [CF85, Gue90, YBTB23]. Expect [Fri97b]. Expected [CZOdlH17, KU99, CL90, Sch88a]. experiment [GHS82, Rus85]. Experimental [ACR01, GIMV03, HBRV10, Lec95, JLFL14]. Experiments [Lec98, MNS84, Smi91b]. Expert [LYT^{+}23, WSS94]. experts [B^{+}07]. Explicit [For02, CFK07]. Exploiting [GKW^{+}10, JDXD13, Kul11, MKF91, KKM^{+}06, Rém17]. exploration [FWDL15, SW12]. explorative [Ker04]. exploratory [ORPF13]. Explore [Cop91]. EXPLORER [FWDL15]. Exploring [CvW18, CMRV10, YB13]. explosion [PLT14]. expressibility [tC09]. Expression [Anoxx, Asp12, BC13b, BP11, Bon07, BTC06, CZ01, CJBW16, CBD^{+}23, CKW09, Cox07, Cox09, Cox10a, Cox12, Dav99, EU98, FC04, GJ16, GRS99, Gib21, Gol93, Han13a, Hol84, Ier09, JLK^{+}20, KM92, KM95a, KN12, LKM23, Lee09, LT16, LN19, MPN^{+}14, Mye92, MOG98, NR99a, NR01, Nav01c, Nav04a, NR04, NR21, NH11, OS11, OWP16, PPA10, RD17, Ric79, Sab76, Sca11, SD95, SM99, SvS14, SL14, SL17, SXZ^{+}25, VCS^{+}12, VB23, WPKL13, WMM95, YP12, YQW^{+}16, ZCH23, vNG01, AI18, Agu23, BAC12, BvdM17, BBvdM21, BFC08, BFG09, BG22, BBRM23, BFS04, BH07, COZ09, CJBW13, CHL^{+}25, Chi17, CLT07, CGPS13b, CS11b, Cox10b, CS22, DF00, FDG^{+}11, Fil21, Fos89, Goo05, GHR14, HN11, Hos06, HVP00, HP01, HP03, HVP05, Hun79, KS08, Kar82, Ker07]. expression [Lee82, Lei80, Lif03, LLL^{+}24, MMI14, Mor90, NNSP22, NT20, OA98, ORT08, ORT09, PIR17, PLT14, PCS99, RTO15, SJ13, SCF^{+}17, Spe85a, Spe85b, Stu03, Stu07, SSYW19, SLZ^{+}20, SMW^{+}23, Tho68, WXZY12, WL15b, WBS22, WW03, WR15, Yam19, YKGS11, YCJK08, YB13, YBTB23, ZZH16, Zia96, ZC99, ZYX^{+}12, dLFM07]. Expressions [ARM^{+}19, AM91, Ano68, Ano12, Anoxx, Ant95, ACT10, Bac94, BDD^{+}14, BR20, Bee13, BF97, Ber00, BGNV10, Bra94, BC94, BMMR19, BK93b, BKW92d, Brz62, BP63, Brz64b, Brz64a, Brz65, CDLV99, CDLV02, Cam99, CSY03, Cha01, Cha02a, CLOZ04, CJM12, CDJM15, CGR02, CHP92, CT23, CC97, CGS17, CDC^{+}23, CDL95, CDL99, Dav03, Dav04, Dav21, Dav22, DM11, FLS98, FU82, Fri02, GHKL18, GGM12, GN12, Ghi62, GS22, Gil70, Gin67, GH13, GH15, HHM^{+}13, Hab04, HM98, Ham88, HWW06, Han13b, HJ99, Hir96, HK11, HSW97, HSW01, Hum99, IY02a, IY02b, KT06, KTU87, Kea91a, KP99b, KP99c, Kin92, KMRY20, KV15, KZ02, KST12, LS99, LS06, LZHZ98, LM01b, LOZ^{+}24, LT09, Loh10, Mad01, MC24, MNS10, MY60, MSZ17, MR09b]. Expressions [MPdS12, MGF97, NM10, OES25, Org03, OF61, Pak91, PM78, Pat71, Pet02, Pik06, Pre99, Ray96, Rez92, SA96, Sch99, SSSS10, Sou99, TV14, TB00, Uma97, Urb23, VHC88, Wen93, WZU14, XK92, XLC19, Yam01, YPG21, ZGS^{+}15, ZMWL20, AFI98, Ano97a, AGM05, AM95, Ant96, AOMC07, ACM02, ADU16, BCG07, BYG96, BRL13, BdFED^{+}20, BTG83, BG91a, BDFR08, BvdM23, BS86, BNSV10, BBM21, BK86, Bra95, BK93f, BK93c, BK93d, BK93e, CGR03, CP97, CX20, Cho78, CK02b, CK08, CGPS13a, CDLM17, DL03, EZ74, EZ76, FL71, FHW10, FS19, Fri97a, Fri06a, GLRÁ11, GR92, Gef03, Gel10, GL03, GL12, GS07, GHR14, GMS12, GM17, GH09, Gue90, HW07, HY90, HWJ03, HSJ04, Hoc19, Hov12, HN00, Jan85, JSH09, Joh69, Kah06, KR22]. expressions [Kap69, KGA^{+}12, Kin91, KMR21, Lar98, Lau00, Lau01, LSO17, Lei81, Lei85, LWS^{+}16, LTV15, LR14, LM12, LM13, LMN16, Lus94, Mag81, MMDdJ11, Mor02, MZZ10, MM89, Nic03, PHXD19, PC02, PM23, PIT^{+}03, PPPdG20, Pra97, RT18, Rob79, Rom14, Ryt89, Sak21, San15, SMS15, Sha88b, SY72, SH85, SM04, Stu07, ST19, SMT^{+}86, TN13, TN15, UIM22, XH23, XJT^{+}04, YH91, YH92, tC09, Hum97a, Hum97b]. Expressive [AGP18, BLLW12, HS08, MFRW09]. Extend [Cal00, dLFM07]. Extended [Ano68, BK93b, CTF^{+}98, Gon02, HY90, HL97, KV15, KZ02, NR98, SvS14, Yam01, YH91, YH92, AM95, BK93c, BK93d, BK93e, CM95, GV00, JM93, RT18, Rob79, SMT^{+}86]. Extendible [vNG01, MKSiA98]. Extending [AS04, DJ96, Jan85, Kea91a, MSRR00, PMS11, WLF14, Bak78]. Extensible [BAC06, KKFI^{+}24, SNM07, BFN^{+}09]. Extension [Liu86, Kau92, SNM07, MMDdJ11]. Extensional [DRW95]. extensions [Mis89a, Mis89b, Wea94, WKR09]. External [GIK97, FG95b, FG99]. extracting [BGHZ15]. Extraction [CBD^{+}23, CT23, FKRV15, FKRV16, HHM^{+}13, KT23, Kea91a, BDMT16, BT21, DLF^{+}15, FKRV13, Kit94, KLR^{+}08]. extractor [Agu23]. extrapolation [Sid95, Sid99, Sid00, Sid02]. extremely [AK08].

F [Ano97a, JGMP22]. FA [CKW09]. faces [KSWC93]. Fact [LSWP19]. Factor [ACR01, CFP19, YQW^{+}16, Bjö93, BH96, HM00, KW05]. factored [Gue90]. factoring [DRSS96]. Factorization [KKP16a, KKP16b]. Factorizations [DW17]. FAdo [MR05]. fails [EMT23]. fairness [MMDdJ11]. Fall [KOI94]. Fall-in [KOI94]. False [Mut97]. Family [KKFI^{+}24]. Fascicle [Knu05]. Fast [ADR03, ADR06, ATX21, BYP92, BYKZ^{+}92, BYR93a, BYR93b, BYP96, BYG96, BYN98, BD80, BLP18, BC13b, BS97, BGP^{+}22, BFC08, BAM^{+}24, BM77, Bre95, BL16, BFK^{+}03, Bun95, CLP98, CR95a, Chu95, Cob94, CP10, Cox07, CCG^{+}99, FL12a, Fen01a, FNU02, Gal76b, GS80, Gia93, Gil85, GWX^{+}23, HAR10, Hor80, HS90, HS91, KBB01, KST94, KKSL01, KMP77, KMP94, KVX12, KNMH00, KRML09, LV88, LV89, Lec07, LT16, LCL06, Man94, Man97, MUHT96, MPN^{+}14, MNS84, Mon17, Mye98, NR98, NBY99a, NR99a, NR00, NR01, NR03, Neb06, Nel96, Ott94, OM88, PPA10, Quo92, Ris16, Ros95, RLP20, Sen00, ST96a, SNZBY00, Smi91b, Sun90a, Sun90b, Tar81a, Vis91, WM92b, WM92a, YKGS11, ZS17, Zha17, ZCH23, AK08, AG84, CDC96, CNPS15, CCG^{+}93, Coo89, Der95, DC94, FDG^{+}11]. fast [IHEH23, II86, KW19, KTP10, LHCK04, Mye99, NBY99c, Nav01b, PS90, RM06, RW10, SW90, Tak93, TLLL09, Vis90, WL15b, dBB08]. Faster [ASM17, ALP04, AKT06, BYN96, BYN99, CH02, DGM90, DGM94, Fre02, GZ94, HGT24, HN02, Ind98, Jez15, LS09, Men89, NKT^{+}01, SB09, WT89, FCLST07, NTS93, WT88, Yam19]. Fastest [Col94b]. fault [BKLE18, BG91a]. fault-tolerant [BG91a]. Feasibility [HTK^{+}21]. feasible [ATdM07]. Feature [Bac94, TBD22, Taf22, WSW16]. Features [OWP16, LR14, Moo12]. February [ACM89, DGBH93]. federated [KLL23]. feed [MA12]. feed-forward [MA12]. Feedback [DKA^{+}15, OES25, Joh95]. fees [SN24]. few [CEPR10, NR17]. FFT [SZ05]. FFT-based [SZ05]. Fgrep [Ash85, KBB01]. Fibonacci [IMS97]. field [WSW16]. fields [CRV06]. Fifth [ACM06, ACM93b, AOV^{+}99]. Fighting [ZGY^{+}16]. File [IK83, Man94, Man97, ZJP^{+}18, All82, KCK93]. Files [ABF96, BH85, BBH^{+}87, Man86, Pol01, TMK^{+}02, ZMSD93, Ayc15, CEMW91, TM05b, TM05a]. Filling [LJZZ13]. Filter [CCH09, FU98, KNMH00, CMS08, ZC89]. Filtering [BAM^{+}24, GMAS22, KVX12, KRML09]. Filters [WZJH12, ZS17, Hos06, MA12]. Filtration [PW95, ST96b, ST96c, LLL13].

Finder [NTS93]. Finding [ATX21, ALLT11, Ben94, Hig86, Iba97, KS11a, KF91, Man75, PRU11, RVV23, VB98, ZD95, BD80, GHST17, LMM17, Lee82, Yam19]. Fingerprint [DS19]. Fingerprint-Based [DS19]. Fingerprinting [LZ18]. Finite [Ant95, BP11, Bow87, CZOdlH17, CLOZ04, CM58, Cho78, FG89, GHKL18, Ghi62, Gol93, GH13, GH15, HSW97, HSW01, JA17, JXA20, KPR97, KPR00, KV15, LY86, Mel95, Pet92, RS59, vNG01, Ant96, BDFR08, BT21, BK93f, Gaá04, HW07, Hur84, Kle56, Kod79, Lei80, Lei81, MMS14, Rou21, Ryt89, SLTB^{+}06, SH85, VHL^{+}12, VW11]. Finite-State [CZOdlH17, vNG01, JXA20, Gaá04, HW07, MMS14]. finitely [AFI98]. Finland [GU95, KS12a]. FIRE [KS08]. FIRE/J [KS08]. Firewall [AV23]. firmware [ZZJC20]. First [BK93g, Len93, Wal89, ACM99b, BLSS03, BD98, BAC06, BRO16, Kau92, Pie08, IEE94a]. First-class [Wal89, BAC06, BRO16, Pie08]. first-order [BLSS03, Kau92]. Fixed [BYCMW94, CF85, KF91, Sut21, ABH^{+}14, GS81a, HOK18a, HOK18b]. fixed-length [HOK18a, HOK18b]. Fixed-Parameter [Sut21]. Fixed-Queries [BYCMW94]. Fixed-Size [CF85]. FL [CVP86]. FlashProfile [PJP^{+}18]. FlashRelate [BGHZ15]. Flexible [LY17, NR02, SvS14, ME97, MM07, NR00, Nav01b, SNZBY00, Lut02]. FLFC [Ode14]. Florida [IEE88, IEE97]. Flow [FO76, HEWK03, FWDL15, MF96]. Fluorescent [HFI^{+}08]. FLUX [Che08]. Fly [PWSG25]. FM [KC21]. FM-index [KC21]. fmpRPMF [LZ18]. Focusing [Kri09, Zei08]. Font [Wol86]. force [GHK14]. Forecast [CDM11]. Forecasting [SF01]. foreign [FF08]. Forest [VRD01]. forests [CADA18]. Foreword [Cro92a]. Form [BMMR19, Lar98]. Formal [AOMC07, BGJ89, CSY03, Coo89, FKRV15, HU92, MGH97, Pag78, ZA87, FKRV13]. formale [HU92]. Formalisation [WZU14]. Formalising [BBvdM21]. formalism [CM90]. Formally [AH14]. formation [Fos89]. Formatter [ZA87]. forms [AOMC07]. formula [CS22, KC11]. formulae [WD99]. Formulating [JM85]. Fortran [PCS99, Wea94]. Forum [Fra83, GHF83a, GHF83b, WNL^{+}83]. Forward [Dur94, MA12]. Forward-Branching [Dur94]. forwarding [SDA17]. FOSS [Bol02]. foundation [Pie08]. Foundations [AAB^{+}17, IEE89, IEE90, IEE92, IEE93, IEE95a, IEE97, IEE98, IEE09, BGWXP22, Win78]. Four [Mye92]. Fourteenth [ACM87, ACM95b]. fourth [ACM92d, A^{+}08]. FPGA [GMAS22, KBB01, LT09, SXZ^{+}25, TK07, YP12, ZCH23]. FPGA-Based [LT09, TK07]. FPGA-CPU [ZCH23]. Fragmentary [HSTS01]. fragments [All82, KMMPN85, OA17]. frame [KB22]. Framework [BGM19, DJ96, KK95, VKPI17, WMGS19, AS04, AMRV16, BKLE18, BYF96, DLF^{+}15, FKRV13, KKM^{+}06, KKM^{+}13, KMS^{+}03, PJP^{+}18, TZH^{+}13, ZKA12]. France [Bun94, KU09, Ng79]. Francisco [ACM92c, ACM95a, DT87, KP15, USE92]. fraud [VD17]. Free [GJ16, HM98, HWW06, Han13a, HSW97, HSW01, Joh69, MFMA15, Nag21, SBHM94, Gef03, HWW07, HY92, HSJ04, KGA^{+}12, Lif03, Mye95, tC09, Bol02, MFMA17]. Freeness [Nag21]. FREME [WL15b]. French [Alb89].

Frequent [BGM19, EASK14]. Friedl [Ano97a]. Friendly [SJNS19]. front [JLHB92]. Frontier [CHZ06, D’A98]. Fukuoka [AT02]. Full [ZJP^{+}18, KWLL08, NM07, OKT92]. Full-Path-Indexed [ZJP^{+}18]. full-text [KWLL08, NM07]. Fully [CDL95, IST05, Jez15, FG95b, GR99]. fully-dynamic [FG95b]. Function [AK09b, BR20, Jon13, MCP17, Ric79, Bar84b, Bjö93, CT96, FF08, GT90]. Functional [ACM92c, Bir10, BK89, Hud89, JR15a, Pou93, Web95, Che08, DA20, Fil21, FHW10, GÁSÁ^{+}13, HJW^{+}92, Mal93, OR11, QTO^{+}20, Rob87, Sch88b, TA90]. Functions [MFMA15, MFMA17, Sch91a, Sch91b, CFLH^{+}22, Dow93, Kod79, Yod91]. functors [dSOMY15]. Fundamental [Sym85]. Further [Gro91a, Sid99, KR22]. Fusing [SVMM17]. Fusion [SdM01]. future [LG16]. fuzzers [CGZ^{+}13]. Fuzzy [GJ16, GN01, RFD23, BC13a, MDLB02, WSS94, WLF14, ZBST14].

G. [Akl78]. Gabriel [Lab12a, Lab12b]. GADT [BGWXP22]. GADT-style [BGWXP22]. GADTs [KSVJ15]. gains [KGP^{+}05]. Galen [Gue87]. Galianosi [Gue87]. Galil [Ano97b, AG86]. Gap [BGJ01, KM92, KM95a, ZKCY07]. Gapped [HGT24, NHN^{+}20]. Gaps [CIM^{+}02, NR03, BGVW12]. GAs [Sut21]. Gases [Mun07]. GateKeeper [BAM^{+}24]. GateKeeper-GPU [BAM^{+}24]. gear [WOQ^{+}07]. gear-shifting [WOQ^{+}07]. Gender [Pak91]. Gene [Bon07, YCJK08]. General [BBRM23, MR92, NR99b, VCS^{+}12, WSVS22, AAB^{+}86, Cha02c, Sch91a, Sch91b, ZHWW12]. General-Purpose [WSVS22, AAB^{+}86, Sch91a, Sch91b]. Generalisation [PWSG25]. Generalization [Shi00, Shi04]. Generalized [Abr87, BK86, GL86, Ham88, Hei01, Hir96, Hol84, MAI^{+}16, OP16, VHC88, Wen93, FL71, Kin91, SW90, Sid95, Sid99, Sid00, Sid02]. Generalizing [SKS96]. Generated [AK09b]. generates [UIM22]. Generating [CGP^{+}08, Jør92, Knu05, WXZY12, BJK^{+}12]. Generation [AGT89, GFH82, GWvG10, HKR92, Ker04, OES25, Pat71, SY72, ZMWL20, BDB90, BA06, FKSB06, Gan89a, GHS82, KKP92, MSRR00, PPPdG20, SMS15]. Generative [TD18, KS08]. Generator [LS79, VSM87, XLC19, AVN22, CLS95, ESL89, FHP92, MR82, RT18]. Generators [Fra83, GHF83a, GHF83b, WNL^{+}83, Gan89b]. Generic [KBB01, WMGS19, ZKA12]. Genesis [SDA17]. Genetic [GC01, RND97, BDMT16, MMDdJ11, Sel84]. GenMatcher [WMGS19]. Genome [WGL^{+}21, SVM14]. Genomic [CCH09]. GenSeq [WGL^{+}21]. geo [ZBST14]. geo-textual [ZBST14]. Geometric [AK09a, CK92, CDEK95, KP96a, Ukk10, AK08, FMdB99, KP96b]. geometrical [Akl78, HLN09, Man76]. geometrically [NA90]. Georgia [ACM99b, IEE09, ACM83]. German [HU92, SM74]. Germany [ACM87, AGS93e, Len93]. Gestalt [RM88]. Gibbs [CRV06]. Given [AW89, Bur11, Lei80]. Glanville [MSRR00]. Global [ZCS^{+}12]. Globbing [HHW^{+}99]. Glossary [ZV97]. Glossary-based [ZV97]. Glushkov [ZC99]. GNU [Hae89, OSM94a, OSM94b, OSM94c]. Go [Far19]. Goal [Gud92, Nil90]. Goal-Directed [Gud92, Nil90]. Gödels [Pux97]. Going [LS06]. Good [BC94, PRT^{+}16, MLM^{+}08, GPTV93, PVA^{+}92]. Google [Cox12]. GOTO [MGF97]. Goyvaerts [Ano12]. GPP [VCS^{+}12]. GPP-Grep [VCS^{+}12]. GPU [BYB^{+}24, BAM^{+}24, DK13, KW19, KMM15, LLC17, MIH17, SR16, TLS16, VB23, ZS13, ZYX^{+}12]. GPU-accelerated [BYB^{+}24, SR16]. GPU-based [ZYX^{+}12]. GPU-to-GPU [ZS13]. GPUs [ARS16, LLCC13, LPJ23, MAC14, VKPI17, YB13]. Grace [HNB^{+}13]. grading [Mor02]. Graham [MSRR00]. Graham-Glanville [MSRR00]. Grails [JNS08]. Grained [FL99, TTO^{+}22]. Gram [ST95, HKN14, KST94, KWL07, KWLL08, KPA10, Sal12]. Gram/2L [KWL07]. GraMi [EASK14]. Grammar [Wat96, Web95, Man06, Wat03]. Grammars [BK93b, BKW92d, Pat71, SBHM94, BK93c, BK93d, BK93e, KGA^{+}12, MMI14, SV09, SH85, Ier09]. Grams [ST04, STK06, Ukk92]. Grand [Bao93]. Graph [AAB^{+}17, AYS^{+}24, BLLW12, BLR14, BYNTK21, BGJ01, CFM17, CMNP17, D’A98, Eke95, FLM^{+}10, Fu97, FMG23, KS93, LYT^{+}23, MNP^{+}23, MSS^{+}19, PJMR14, RKH02, SK17, WHZ^{+}17, Zue96, A^{+}08, BKLE18, BLR11, BSTU08, BYB^{+}24, BCD14, BYK22, EASK14, FWW13a, FWW13b, FWW13c, FMG22, GPTV93, LRV13, MCF^{+}11, MCF^{+}14, SW93, SHvR^{+}16, SGCW14, Sri88, TG96, ZXL^{+}13, ZCÖ09, ZCÖZ12]. Graph- [CMNP17]. Graph-based [PJMR14]. graph-oriented [GPTV93, TG96]. Graph-Structured [BLLW12]. Graphic [LLLC17]. graphical [CMW87, LLS12]. graphics [AK08]. GraphLog [CM90]. Graphs [BJM79, BYNTK21, EMTG23, EMT23, Fu95, Fu96, HPM94, LMV16, LSWP19, MY60, RHR^{+}21, TV14, WD99, BBG13, BÖ13, FFTD15, KC15, QQC^{+}13a, QQC^{+}13b, SVM14, SWW^{+}12, ZYQ^{+}15]. Graspan [WHZ^{+}17, ZWH^{+}21]. great [Sch81]. Greco [Sal01]. Greco-Roman [Sal01]. Greece [Len11]. Greedy [FC04, KR92, TWZ23, TU88, Huc21]. Greek [Gue87]. Greek-Armenian [Gue87]. Grep [Pit98, VCS^{+}12, Goe95, NTS93, Roo99, Tay97, Kah07, Kah09, Sid99, Sid02, Hol84, Bar90, Hae89, Nav01b, Woo86, Hum88b, BK09]. Greppin [Sal01, Gue87]. Grepping [HHW^{+}99]. Greps [Hum88a]. Groovy [JNS08]. ground [KR95]. Group [DT87, KC99, GMC02]. Grouping [OR12]. Groups [Sch22, Joh69]. gset [Liu14]. guarantees [FWW12]. guards [GJS20, JM90, KSVJ15, Rai99]. Gueriguian [Sal01]. Guessing [Pak91]. Guide [GS93a]. guided [FhDAF09, Nav01a]. Guidelines [Anoxx, Dav99]. Guild [B^{+}07].

Hairpin [CDJM15]. Hamiltonian [YT03]. Hamming [HOK18a, AD11, GF08, HOK18b, Ris16, TT20]. HAMPI [KGA^{+}12]. Handwritten [CLP95, SKS96]. hard [LMM17]. Hardware [Bur84, HH83, HKL^{+}14, Lee09, Lut02, MGW14, OWP16, PK85, Rob92, ZS17, AK08, ACF05, Bur82, FNP09, GZ10a, Hur84, KKM^{+}06, MP88, MLC08, TYNM86]. hardware-accelerated [MLC08]. hardware-assist [KKM^{+}06]. Hardware-Based [HH83]. Hardwired [BK75]. harmonic [BCWG09]. Hash [CCH09, Dav73, RJK79, Sch91a, Sch91b, KB22]. Hashing [Bur84, CKW09, CG79a, CG79b, GIG77, Gri79, Har71, LLLC17, TK07, ASM17, Bur82, Kim99, MKSiA98, TLLL07, XMLC11]. Haskell [HJW^{+}92, Jon07, VLP17, Wen93]. Hate [HHW^{+}99]. Hausdorff [Rot91]. Head [LY86, WWW^{+}16]. Head-Modifier [WWW^{+}16]. Heads [JL96, GGL94, JL93]. Healthcare [BTTF02]. height [Tho81]. Help [DKA^{+}15, FR17]. helpful [VVV04]. Helsinki [KS12a]. Hershey [ACM76]. Heterogeneous [MM02, MM03]. heuristic [BCD98, Mus05]. Heuristics [CIL^{+}03, Han13b, KR92]. Hierarchical [GM02, IK83, Loh10, Coo89, KAT07]. hierarchy [Lar98]. High [BYHT18, BTC06, CGM10, Ear74, JLK^{+}20, JGZL12, LK90, Lee09, LPT12, PW12, RSG^{+}19, TS05, VCS^{+}12, WGL^{+}21, Wea94, YP12, ASJDW18, Dit78, HC87, KM84, LK88, LH13a, PLT14, SW12, SNB^{+}19, TLLL07, XMLC11, ZYX^{+}12]. High-level [Wea94, ASJDW18, HC87, SW12]. High-Performance [JLK^{+}20, Lee09, WGL^{+}21, YP12, CGM10, SNB^{+}19]. High-Speed [BYHT18, LK90, PW12, VCS^{+}12, LK88, PLT14, TLLL07, XMLC11]. High-Throughput [BTC06, LPT12]. Higher [HW12, JR15b, KU99, SdM01, Chl08, NRO12, OR11, Pie08, Zei08]. Higher-Dimensional [KU99]. Higher-Order [HW12, JR15b, SdM01, Chl08, NRO12, OR11, Pie08, Zei08]. Highlighting [PWSG25]. Highly [BEL17, BKLP97, BKL^{+}02, GHK^{+}91, Nav21b, Nav21a, NEL17, BFKL13, CDC96, MAC14]. highly-threaded [MAC14]. Hilbert [HHW^{+}99]. History [JDXD13, LG16, Ritxx]. HMM [SB09]. Holes [YGG^{+}23]. Holism [MMZ10]. Holistic [BKS02]. Homology [Zha07]. Hong [B^{+}02]. Honnef [Len93]. Horspool [BYR92, Neb06, Rai92, Smi94, TJMC20]. Host [ZS13]. Host-to-Host [ZS13]. Hotel [ACM83, Bao93]. Hough [KC87, SA77]. HTTP [BBK12]. Huffman [CFG12, DS04, FT04, KS01]. Huge [NR21]. human [KSWC93]. Hy [Lia84]. Hy-phen-a-tion [Lia84]. Hybrid [CLP95, LZHZ98, LLC17, SF01, SW09, VB12, AVN22, Gri85, LLL13, SMW^{+}23]. Hybridizing [RFD23]. HydroJ [LLC03]. hyogen [SM04]. Hypercube [Les94]. Hypermedia [LZ96]. Hypertext [ALL97, ALL00, Nav98, PK95, Nav00, SD91a, SD91b]. Hypothesis [Liu14].

I/O [PSK08, ZYQ^{+}15, dBB08]. IBM [HKL^{+}14, Wei84]. ICL [CPW88]. Icon [Gri83, Gri85, Wal89]. Iconic [GL86]. ID [BCD98]. Ideas [Bee81, Wol90]. Identification [Cob94, LZ18]. Identifying [FLSS93a, FLSS93b]. Identities [McI85b, McI85a]. idf} [TP07a, TP07b]. Idiom [KKM^{+}13, KKM^{+}06]. IDPM [LJH^{+}17]. IEEE [Bao93, CVP86, IEE09]. IFIP [FGR72]. ifs [Edw07]. II [AU73, Nav21b, OSM94a, OSM94b, OSM94c]. Illinois [Hwa85, Hwa85, Rob87]. illustrated [Moy24]. illustrating [HWF90]. Image [DS19, How97, LV94, SN92, VB98, ASG99, AGS96, ZC89]. Images [GR96, HMW^{+}23, KPR97, KPR00, KS06, How96, KS05, YCJK08]. Imaging [AGS93a, AGS93b, AGS93d, AGS93e, AGS93c]. immersion [HFI^{+}08]. impact [MBH20, NEH90, NCV10]. Implement [Cha01, Cha02a, ADM^{+}13]. Implementation [Bar81, DNM00, Gim73, Har71, HOS85a, HOS85b, LZ18, MHT09, NW13, RND97, TT22, VKPI17, Vin77a, Vin77b, Yun12, Aoe89, AG84, Bro77, MK90, NK07, OA17, PLL10, PD12, PCS99, ZYX^{+}12]. implementations [Nak14]. Implementing [AM91, Gri83, LT90a, WT88, WT89, BBvdM21, BD98]. implication [LS10]. Implicit [Cha01, Cha02a]. Imply [Gal76b]. important [Jed87]. Improve [Bon07]. Improved [BY89, BFG09, CMO^{+}08, CM08, GG86, GP90, Han13a, IS86, KV15, KZ02, LSW08, LJH^{+}17, LDI98, LJZZ13, Nav98, Nav00, Pol13, RKM21, RFD23, Tan14, Ayc15, BC95, Oph89, SSYW19]. improvement [Cha87]. Improvements [CK92]. Improving [AYS84, Bir77b, ĎHPT10, Gal79, GKW^{+}10, Hyy08, KCPC13, NBY01, YQW^{+}16, HIEH22]. In- [MPdS12]. in-degree [LSV08]. In-Memory [SRV^{+}19, TT20]. In-place [HTX17]. In-Situ [GMAS22]. In-Storage [JLK^{+}20]. in-vehicle [BKLE18]. Inclusion [CX20, CGPS13b, CGPS13a, Hov12]. Inclusive [MIH17]. Incomplete [NCKL14, Ritxx]. Inconsistencies [FKRV16]. incorporate [SKS96]. Increased [HFN05]. Increasing [HR00]. Incremental [CHL^{+}25, FWW13c, HKR92, Mey85, ISHY88]. Independent [ABF94a, CR95b, KR94, GP92, VS11]. InDesign [Kah06, Kah07, Kah09]. Indeterminate [DW17, SW09, LMS21]. Index [CGR02, CN02, Cox12, Gia93, HTK^{+}21, Man86, Zve80, All82, CGR03, DGM19, HLS^{+}11, KWL07, KWLL08, KC21, KST16, NC06, TPT13, ZCT14]. index-sensitive [TPT13]. Indexed [TCP16, ZJP^{+}18, EMT23, GMC02, GO12, KNT15, Sen00]. Indexes [CLS^{+}10, KRR17, Nav21b, RNOM09, CHLS07, NM07]. Indexing [BGS23, Gib21, GL86, GV05, HGT24, KKSL01, KTY^{+}18, LMRT14, LM01b, NBY99b, Nav21b, Nav21a, SWY75, SVM14, GV00, HAI02, SJ13]. Indianapolis [DGBH93]. indirect [DSv94]. Induction [Zha96]. Inductive [ASJDW21, ASJDW18, CL09]. industrial [Moo12]. industrial-strength [Moo12]. Inference [BNSV10, BGNV10, VB23, JM93, Tej20, Van06]. Infinite [KT06, KST12, SMDS94, APTS13, Sid99]. Infix [HWW06]. Infix-Free [HWW06]. Inflectional [TB00]. Influence [FTJ95]. Inform [Gro91a]. Information [FKRV15, FKRV16, FBY92, IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, KT23, LZ96, MKF91, MSP^{+}17, SD91a, SD91b, Sno01, XPZ^{+}23, FKRV13, Kit94, KLR^{+}08, Lus94, SHS14, SKS96, FGR72]. Informative [IEE01c, IEE01d]. inheritance [Mor02]. Initial [Man75]. initialization [MME14]. Injection [AV23]. Inner [BF97]. Input [TJD^{+}17, Sak21]. input-driven [Sak21]. Inputs [CGP^{+}08]. Inscription [D^{+}23]. Inspection [LLLL08, LLC17, PW12, VWR11, YP13, ARS16, BAC12, NYuR15, NNSP22, SSYW19]. Inspired [GWvG10, Pet92, San09]. Installations [SK17]. instance [FK96]. instant [Abb77]. instead [AGH^{+}17]. instructions [CGT25, KKM^{+}06, MR82]. insufficient [LG16]. Integer [Nav04a]. integers [Mat94]. integral [EF95, SY72]. integrals [Sid99]. Integrated [BGG^{+}94, FU82, WBA83]. Integration [Har79, Fat15, PJMR14]. Intelligence [IEE94b, PJMR14, Rob92]. intelligent [JLHB92]. Interaction [BNV^{+}13, KP96a, KP96b, HR00]. Interactions [BNV^{+}13]. Interactive [Han02, MR05, BKM95, BH07, DKP11, MI07]. interchanges [Chr96]. Interconnect [SRV^{+}19]. interconnection [KRL87]. interdisciplinary [Roe25]. Interest [DT87]. interests [SW93]. Interface [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d]. Interfaces [IEE01g, IEE01h, PW06]. Interferences [FTJ95]. Interleaving [CGS17, CGPS13a, Gel10]. Internal [ACR20, KRRW24]. International [ACM94b, ABB93, AGS93e, AAC^{+}01, AOV^{+}99, Bao93, B^{+}02, Bro93, Bun94, DMVT13, FMA02, IEE95b, Kap92, KP15, Lev95, Ng79, NH11, SW94, Sto92, WN90, A^{+}08, BGNP94, HF13, MG94]. Interpolation [HW12, Lut02]. interpretation [HN11, JP11, NC92, SHCY93]. interpretations [MP09]. Interpreter [HOS85b, Mae94, Eck89, HOS85a, Rob87]. Interprocedural [WHZ^{+}17, ZWH^{+}21, FWDL15]. Intersection [GN12, HL10, Pet02, CS09, CHL^{+}25, CP10, Gel10]. Interval [Via02, Via04, WSS94]. intractable [FLM^{+}10]. Intrinsic [MFMA15, MFMA17]. Introducing [LV86b]. Introduction [Bir77b, CM86, CLR90, GS93a, HU79, HMU01, HMU07, PG90, SC88, HU92]. introductory [LSO17]. Intrusion [BYHT18, CZCD09, Hig95, KKK11, LY17, TS05, ZL18, ACF05, GZ10a, KAT07, LHCK04, TBS06]. Intrusions [KNMH00]. invalid [SMS15]. Invariant [FU98, KMRY20, LT03, ZC89, Deo06, KMR21, MNU05]. Inverse [AAL97b, GIK97, GHW05]. Inversion [KS94, Gie90, GT90]. inversions [CCF13, GFG11]. Inverted [BBH^{+}87, ZMSD93, KWL07]. invisible [EHS07]. invited [Rém17]. Invocation [Kor83]. Involutions [GMMN12]. IO [PSK08]. Ireland [ABB93]. Irregular [PCS99]. ISBN [Ano12]. Island [ACP05]. Isomorphism [BJM79, Gro91a, Gro91b, KSH^{+}15, Mäk89]. Israel [AL01]. ISSAC [Lev95, WN90]. ISSAC’93 [Bro93]. ISSAC’94. [ACM94b]. Issue [Ano17, Cro92a, IEE01a, IEE01g, IEE01e, IEE01b, IEE01h, IEE01f, AGS93a, AGS93b, AGS93d, AGS93c]. issues [BG91a, IS90, San15]. Istanbul [SMD04]. Italy [AAC^{+}01, Apo93a, FL08, GM11]. Iterable [LM02]. iterated [Jan85]. itself [JGMP22].

J [Tal81, KS08]. J2EE [BTTF02]. Jackson [San09]. Jan [Ano12]. January [ACM87, ACM92a, ACM93a, ACM94a, ACM95a, ACM97b, USE92]. Japan [AT02, IEE94a, WN90]. Java [Agu23, Ano96a, Ano96b, Cal00, CGM06, Dwe00, FR00, Hab04, HHW^{+}99, LM02, MFRW09, Mor02, NAR08, NM10, Pit98, Sch14, SM04, Stu07, dH05]. Java-based [Ano96b, Ano96a]. java.util.regex [Hab04]. JavaScript [KT14]. Jeffrey [Ano97a]. Jeju [ACP05]. Jersey [FC98]. Jerusalem [AL01]. Jigsaw [BK93g]. JMatch [LM02]. John [Sal01]. join [WDG^{+}14, WLF14, ZCÖ09]. joins [BKS02, JLFL14]. Jose [ACM95b]. JPEG [KS05, KS06]. jQuery [PIR17]. JTL [CGM06]. July [AL01, AH97, AT02, Bro93, Bun94, Cha86, Cro92a, FC98, GU95, KS12a, KP15, Lev95, LV06, MZ07, PC99, SMD04]. Jumbled [BCFL12, GHLW15, KRR17, BFKL13, GG13]. June [ACM92c, ACM92b, ACM95c, ACM98, ACM99a, ACM06, ACM07, AP10, Apo93a, AH97, ACP05, BYCC03, Bun94, CG94b, FL08, FMA02, GS00, GM11, HM96, HF13, Hwa85, CVP86, Kap92, KU09, LL08, Len11, Ng79, Sto92]. JVM [BFN^{+}09].

K-M-P [RUG97]. Karp [CR91, GBY90a, GBY90b]. Karp-Miller-Rosenberg [CR91]. Karp-Rabin [GBY90a, GBY90b]. Keeper [Wei84]. Kentucky [ACM89]. Kepler [TLS16]. kernel [WKR09]. Kernelization [BCKM15]. Key [CG79a, CG79b, Gri79, RMK^{+}14]. Key-Exchange [RMK^{+}14]. Keys [FFTD15]. Keyword [And02, GN19, ADT15, CWZ10, OSSK16, QQC^{+}13a, QQC^{+}13b, WZ96]. Kiev [Bro93]. kinematics [PS93a]. Kleene [Lee82]. Kleenex [GHR^{+}16]. KMP [Mor90]. Knowledge [LSWP19]. known [KCK93]. Knuth [Bar81, DS04, Hua98, Moy24, PV91, Ruc15, TA90, Ukk10]. Komplexitätstheorie [HU92]. Kong [B^{+}02]. Korea [ACP05]. Korean [JLH18]. Kumar [Hig95].

L [Sal01]. Label [XPZ^{+}23, VR18]. Labeling [FK16]. Labels [KMRY20, KMR21]. Lafayette [Hig95]. Laguna [HM96]. lambda [Dow91]. lambda-calculi [Dow91]. Language [AKW85, ADR15, Ano68, Car77, Fre06, Fre78, GS93a, GP93, GH09, Gud92, GR96, Hir96, KN17, LS99, Lut02, MGH97, NM10, Reg92, Sab76, SBF80, SBM^{+}18, TBD22, Taf22, TB00, VVV04, vNG01, Arn93a, Arn93b, BK86, BFS00, CFM00, CM86, CGM06, CK08, CMW87, FL71, FPD08, Fri97b, HJW^{+}92, Jør92, KH06, KN19, Mal93, MLC08, RT18, RW93, RTO15, SNB^{+}19, SC88, Ste12, Str13, SNM07, YIAS89, Zia96, DMVT13, DWE89]. Languages [ACM92a, ACM93a, ACM94a, ACM95a, AAB^{+}17, AGP18, BLLW12, CM58, CDPP23, EJ23, HWW06, HU79, HU92, HMU01, HMU07, Hud89, KT06, KP99c, KLH16, Kor83, KST12, Nag21, ND02, SA96, Sch12, Sch13, Wag74, ACM87, AGM05, AOMC07, BRL13, BdFED^{+}20, BLSS03, BKW92a, BKW92b, BKW92c, BDM19, CS09, Coh90, Dit78, ETV21, FhDAF09, HWW07, HWJ03, HSJ04, HW09, Kes91, LRV13, McI04, MZZ10, MPW21, Mye95, PP85, Rus88, Sak21, Sch88b, SHvR^{+}16, Smi91a, dLFM07, BGNP94]. Large [AAC^{+}01, AOV^{+}99, BH85, B^{+}02, LP13, LHMH91, Ris16, VB98, WHZ^{+}17, ZMSD93, ZWH^{+}21, ABB93, BC13a, CD96, EASK14, HAI02, LYWL08, Owo93, QQC^{+}13a, QQC^{+}13b, RW10, YHV^{+}15, ZD95, ZCÖ09, ZCÖZ12].

Large-Scale [LP13, WHZ^{+}17, ZWH^{+}21, LYWL08]. Larger [GZ94]. LATA [DMVT13, NH11]. latch [Fos89]. Lattice [Mun07]. Law [AW89]. Layout [JDXD13]. laziness [KSVJ15]. Lazy [KKP16a, KKP16b, Jay92, Jør92, AJ89]. Lazy-ML [AJ89]. LDA [YCJK08]. Learned [HTK^{+}21]. Learning [BGNV10, Bra94, BC94, Bra95, KK08, Kin92, KK02, OES25, Org03, PDL98, RFD23, SPF08, SG12, ZCS^{+}12, BDMT16, BC06, Ker04, MDLB02, PKK18, STKD20, THQ19, VVV04]. learning-based [PKK18]. least [Boo80, DK13]. Left [NWE97, Ned98, CWZ10, CA20, HR03, Tak96b]. Left-to-Right [NWE97, Ned98, Tak96b]. legacy [Joh94b]. Leif [SC88]. Lempel [BK93a, BFG09, FT95, FT98, KKP16a, KKP16b, NR99b, Nav01c, NT05]. Length [BL94, Bre94a, Chu95, LLC17, YJ84, ZGS^{+}15, BFKL13, BGVW12, BC95, CS22, HOK18a, HOK18b, KR97, ZZH10]. Length-Bounded [LLC17]. Lengths [CCH^{+}23, KIH15]. lenses [BFP^{+}08, FPP08]. Less [LMRT14]. Let [ABF96, MW92b]. Lett. [Gro91a]. letter [AGM05, Bar84b]. Letters [Ale94, BTTF02, HHW^{+}99]. Level [JSC83, MFMA17, ASJDW18, Dit78, Ear74, HC87, KWL07, Moo12, dSOMY15, SW12, Wea94]. levels [Lar98]. leverage [LR14]. Leveraging [MGW14]. Levithan [Ano12]. Lex [LS79]. Lexer [SvS14]. Lexical [HKR92, LS79, Nip98, Yan95, ISHY88]. Lexico [KKSL01]. Lexico-Syntactic [KKSL01]. Lexicographically [Boo80, CDPP23]. Lexicons [ZMSD93, ZD95]. Lexing [Urb23, ADU16, ELF22]. Libraries [Ano10]. Library [AK09b, CL95, EU98, Ano01, Cox10b, PSK17]. library-defined [PSK17]. Lie [ABF96]. life [CM90, Ste12]. light [Agu23, SNM^{+}13]. lightweight [BFNP10, DA18, SNM07, GWX^{+}23]. Like [CFG12, GHLW15, Hol84, HK11, AMRV16, BTG83, Mis89a, Mis89b, MPW21, YH91]. Lille [KU09]. Limitation [Kül10]. Limited [HAR10]. limits [Sid00]. Line [FG98, GG97, Lut02, Man75, MM93, Prů17, Sno01, Tak86a, Tak86b, BEL17, Blu08, CLP95, CT96, FG95a, Fre06, Gal75, Joh95, KNT11, NR02, NEL17, Rot91, Shi96, TIT83, YC22]. Linear [BJM79, Brz65, Bur11, CFP19, Cha94, Cha02b, CGS17, CH03, Coo72, CR95b, CGPR95, GS81a, LK90, LO94, Man75, Pat71, PRU11, RPE81, SSSS10, CGPS13b, EH88, ETV88, GFG11, GMS12, HKN14, IKX15, KKR^{+}13, LK88, Rep98, SGYM00]. linear-space [IKX15]. Linear-Time [CR95b, Man75, GS81a, HKN14]. linearised [TJMC20]. linearity [GO80]. Linguistic [Haz01]. Linguistically [GWvG10]. Link [LTL04, KCPC13]. linkage [VRC24]. linked [BAP06]. Linter [UIM22]. Linux [Blu08, Mah07, Qui00, dBB08]. liquid [VLP17]. LISP [ACM92c, JLHB92, Kod79, MS20, Mu 95, MuT95, Mun95]. list [dSOMY15]. list-of-functors [dSOMY15]. listless [Jay92]. Lists [Gil85, BAP06]. Literary [HSTS01]. Literate [Ham88, Pep91, VHC88]. Lithographic [DMWW77]. lithography [SS93b]. Live [YGG^{+}23]. lives [Joh01]. Ljubjana [FGR72]. Load [MM02, MM03]. Local [ABH^{+}14, CM94, DJ96, GHK14, MU02, ZCS^{+}12, GS81a, MPW21]. Locality [TLC15]. Locality-Centric [TLC15]. locally [Mei08]. locating [Mar89]. Location [LYT^{+}23]. Locations [ST95, GS81a]. Log [DJ96]. logarithmic [Sid95, Sid02]. LOGCFL [Pet02]. Logic [Bac94, GHK^{+}91, Mae94, BdFED^{+}20, CDL08, Coh90, HR01, Kun95, Smi91a, TPT13, YIAS89, Tal81]. Logical [CEW58, Wei84, PP85]. logics [LH03, Pel87, tC09]. London [MZ07]. Long [CLP98, KYG19, Kha16, ML96a]. Longest [ACR20, BK93a, Bur11, F^{+}23, FGKU15, HIRS17, JN24, KR92, KRS19b, KRS23, RT17, Rou21, BBHK14, Gra15, NHN^{+}20, TTO^{+}22, HTK^{+}21, KRS19a]. Longest-match [BK93a]. LongestMatch [Huc21]. Look [McC01, Yan95, GPN96]. Look-Ahead [Yan95]. lookahead [BAC12]. Lookaround [MC24]. lookup [KW19]. Loops [BF97, FTJ95, KK95, BK86, RP95]. Lossless [How96, RDRF24, Cha93b]. Lossy [LS94, RT17, How96]. Louis [IEE90]. Louisiana [ACM91, ACM97b]. Louisville [ACM89]. Love [GP18]. low [LH13a]. Lower [BG92, CHPZ95, GG91, GJS20, GK86, AGW13, BCKM15, BG91b, CJPS13, Lif03, LP08, SASU13, SV87]. lower-variance [AGW13]. LR [LK06]. LRPD [RP95]. Lyndon [Fra20, SMDS94]. LZgrep [NT05]. LZSS [LD10]. LZW [GR99, Gaw12, Gaw13, KTS^{+}98, KTSA99, TM04, TM05b, TM05a]. LZW-Compressed [Gaw12, GR99].

M [RUG97, Ram94]. MACHINE [BY91, CG87, Cox09, RFD23, AG84, BY92, Nak14, Ram94, WHZ^{+}17]. Machinery [DT87]. Machines [AYS84, Bow87, BP63, JA17, JXA20, Moo64, OF61, Pux97, YD95, Aoe89, GOMSJVGP08, KAT07, MMS14, Yod91]. macro [Sas79]. Macsyma [JM85]. Madison [FMA02]. Main [BK75]. Maintaining [AS91, GO12]. Maintenance [CS18, MG94, KPA10, PA10]. make [JT94, Mei15]. makematch [Kas08a, Kas08b, Kas08c]. Mäkinen [Gro91a]. Making [ABBH^{+}16, ATdM07, NE93]. Malay [BSY00]. male [KT90]. Malicious [HL10, HT14]. Manacher [Akl78]. Management [BK75, DT87, FMA02, SW94, Sto92, HF13, WXZY12]. Managing [SBM^{+}18]. manga [QPWH08]. Manipulate [Ghi62]. Manipulating [VMML15]. Manipulation [Car77, Hoa77, Ng79, Wea94, GHS12, GS06, Mal93, MR05, RH81, SNM^{+}13]. Manual [AKW85, Mu 95, MuT95, Mun95, Ski98, HA90, TSM88]. Manuscripts [AS91]. Many [JA17, JXA20, TLC15, Wal88, MAC14]. Many-Core [TLC15, MAC14]. Many-Sorted [Wal88]. Many/Multi [JA17, JXA20]. Many/Multi-core [JA17, JXA20]. map [KD15, SASU13, ZBST14]. map-reduce [SASU13]. MAPMaN [HMW^{+}23]. Mapper [GMAS22]. Mapping [BAM^{+}24, CFM17, KYG19]. maps [BCWG09]. March [ACM83, IEE94b, SC93, SC95, SC96, SC98, SC99, SC01, SC03, SC04, SM09, SM10, SM11]. Marina [ACM69]. Markov [KCPC13, KR14, LBK08]. Marseille [Ng79]. Maryland [ACM90b]. Mass [BM08, LZ18]. masses [Vol12]. Massive [BYNTK21, OR12, YDDB15, ZYQ^{+}15]. Massively [CG87]. Master [Tay97]. Mastering [Fri97a, Fri02, Fri06a, LR14, Rom14, Uma97, Ano97a, Hum97a, Hum97b].

Match [GHW05, KR92, LD10, Mor83, Pet92, Ses96, VB98, Zve80, Bak78, BK93a, BBHK14, DWE89, GJS20, JGMP22, KSVJ15, KCK93, Mei15, ZCÖ09, ZCÖZ12, HC87]. Match-Bounds [GHW05]. Matcher [HH83, Sab76, Coo86, Ker07]. Matches [Dav73, KF91, Mut97, MOG98, PRU11, RVV23, GHST17, Mha05, Ukk92, Yam19, ZD95]. Matching [AOK02, Abr87, ABM08, AC75, AGT89, Aku94, AR00, ACR01, ABF94a, ABF96, AAL97b, ALL97, ALLL98a, ALL00, AAL^{+}00, ALR08, AP10, Ano92b, Ano96a, Ano96b, Ano17, AYS84, iA94, AYS^{+}24, AT02, ADLM96, AW89, Ash85, AJS92, ACD01, ASA17, BST^{+}03, BYP92, BYCMW94, BYN96, BY96a, BYN97, BYN98, BYN99, Bak96, BEM^{+}12, BCP02, BLP18, Bee81, Bee13, BEL17, BH02, BH85, BKLP97, BKL^{+}02, BGP^{+}22, BL94, BM00, BBL93, BYNTK21, Bow87, BG92, Bre93, Bre94a, BCT94, BG95, BCT98, BGG12, BG14, BTC06, BL16, BK93g, Bun95, BZ98, BGJ01, BCFL12, BCC^{+}13, CCFG12, CF06, CFP19, CFM17, CDM11, CK02a, CLS^{+}10, CL92, CM94, CL94, CCH09, CL97, CLP98, Cha02b, CN02, CTF^{+}98, CZCD09, CHL14, CJBW16, CWL^{+}21, CK92, CDEK95, CG94a, CLP95, CM08, CL95]. Matching [Chu95, CW84, CHZ06, CJPS12, Cob94, Col94a, CHPZ95, CH97a, CH02, CH03, CHLT14, Col94b, CG79a, CG79b, Cox07, Cox09, Cox10a, Cox12, CP91, Cro92a, CR92, CCG^{+}94, CR95b, CGPR95, CGG^{+}97, CGH^{+}98, CIK98, CIM^{+}02, CIL^{+}03, D’A98, DB86, DLG12, D^{+}23, DN77, DCM15, DGM94, Dwe00, EIV04, ETV88, Eke95, EMC96, EF13, EMTG23, EZYA23, FT98, FL12a, FL12b, FMP20, FMMS20, FG98, FL08, FPT22, FR00, For02, FU98, Fre02, FNU02, FT04, Fre06, FC04, Fu95, Fu96, Fu97, GHLW15, Gal76b, Gal79, GS80, Gal81, GP90, GG91, GG92, Gal95, GPP04, GC01, GPR95a, GIK97, GP01, GP03, GIMV03, Gaw12, Gaw13, GP93, GM02, Gia93, GG95, GG97, GM11, Gib21, GS22, Gil85, GKP19, GZ94, GWX^{+}23, Gon02, GK86, Gri79, Gri83]. Matching [GL01, Gro92, GL86, GV05, GMMN12, GH82, HD80, Han13a, Har02, Har97, HAR10, HL10, HT14, Haz01, Hea71, HEWK03, Hei01, HL97, HUN^{+}19, HH93a, Hig95, HT17, HO82, HMW^{+}23, HSTS01, HGT24, How97, Hui92, HW12, HN02, HN05, IS94, IMP01, IMR08, IST05, IS86, IMM^{+}22, IK83, JZAA19, JLK^{+}20, Jez15, JL96, JGZL12, JN24, JSC83, JTU96, KPR97, KPR00, KU99, KS12a, KR81a, KR81b, KR87, KRS95, KRS97, KO83, KP93, Kes79, Kha16, KTSA99, KMT^{+}01, Kid09, KST94, KS99, KKSL01, KKK11, KS06, KS11b, KS12b, KM92, KM95a, KM95b, KMP77, KLH16, KRR17, KRRW24, Kor83, KMM15, KB18, KK02, KR97, KU09, KNS12, Kül10, KVX12, KNMH00, KC99, Lab12a, Lab12b, LSW08, LV94, Lav91, LP13, Le 91, LM01a, LKM23, Lec95, Lec98, LKL02, Lee09, LY17]. Matching [LT03, Les95, Les94, LV06, LY86, LYT^{+}23, LTL04, LLLL08, LA12, LLCC13, LJH^{+}17, LLC17, LLLC17, LP11, Liu86, Liu88, LM02, LT16, LN19, LGZ^{+}14, LCL06, LLW^{+}15, LS94, Lut02, MZ07, Maa06, MS98, MBP22, MKF91, MU02, MC24, MW92a, MW92b, MGW14, MR11, MNP^{+}23, MSV24, MHT09, MUHT96, McI85b, MPN^{+}14, Mel95, Mey85, MM02, MIH17, Moh97, MS01, Mon17, MNR^{+}23, ML96a, ML96b, Mu 95, MuT95, Mun07, MR92, Mut97, Mut00, Mye92, Mye98, Nao91, NW13, NR98, Nav98, NBY99a, NR99b, NBY99b, NBY01, NR03, Nav04b, NWE97, Ned98, NdMM02b, ND02, NRS18, NCKL14, NR15, NEL17, Ode14, OR12, OS11, OP16, Ott94, OM88, PDL98, PAMP12, PS10, PLL08, PK95, Par96, PV91, PPA10, PW95, Phi94, Pol13, PP09, Pou93, PRT^{+}16, PK85, Prů17, PS93b, RR90]. Matching [RR92, Rao95, RM88, RTT02b, RKM21, RS98, RDRF24, RHR^{+}21, Ric79, Ris16, RKH02, RPE81, Roe25, RMK^{+}14, RT17, RSG^{+}19, RFD23, RNOM09, Sad96, SV94, SMD04, STK10, SCFC94, SN92, SP16, Sca11, Sch95, SRR92, SRR95, SD95, Sha93, STSA99, SKF^{+}00, Shi00, Shi04, Shi92, SSSS10, Sim83, Sim94, SF01, SdM01, Sli78, Sli83, SW09, Som82, Spi99b, Sto96, SXZ^{+}25, ST95, ST96b, ST96c, ST04, TBD22, Taf22, Tak86a, Tak86b, Tak94, TT22, TMK^{+}02, TS05, TZW94, TWZ23, TU93, TP97, TMV^{+}01, TJGY22, TK07, TLC15, TL12a, TL12b, TVCM12, UW93, Ukk10, UFA^{+}24, VSM87, VB12, VWR11, Via02, VG01, VRD01, Vis91, Vis99, VS01, WAH23, WPKL13, WSW16, WMGS19, WZL^{+}23, Wat96, WKA94, WD99, WBA83, Wri94, WM92b, WMM95, WSVS22, Xi03]. Matching [XZL^{+}19, YP12, YP13, YQW^{+}16, YK11, YJ84, YDW18, YGG^{+}23, Yun12, ZZ12, ZS17, ZS13, ZL18, Zha17, ZLN11, ZCH23, ZT89, Zue96, de 82, van14a, van14b, AMB^{+}02, ADR03, ADR06, AK08, AK09a, Akl78, Aku95, ASM17, Alb89, ACF05, ASG99, ALV92, AF92, AFM94, ABF94b, AAL^{+}97a, ALLL98b, AL01, ALP04, ABC^{+}04, AKT06, ALLS07, AAK^{+}09, AAB^{+}09, AEK^{+}11, ABH^{+}14, Ano97b, Ano01, Aoe89, AG84, Apo92, Apo93a, AH97, AG97, ACP05, ADLM01, AGS96, AD11, AGW13, AG06, BFKL13, BKLE18, BYR93a, BYR93b, BYP96, BYCC03, BSY00, Bak78, Bak93, BDB90, BCD98, BEM^{+}13, BSTU08, Bar22, BGFK15a, BGFK15b, BR09, BA15, BA16, BKBB^{+}14, BYB^{+}24, BBvdM21, BCD14, BPPR20, BLLP90, BLPL92, BFC08, BFG09, BGVW12, BG22, Bir77a, BGJ89, BÖ13, BBRM23, BGWXP22, BBL98, BYK22, Bra90]. matching [Bra95, BBK12, BBHK14, BG90, BG91b, BCT93, Bre95, Bre96, BGM13, BKS02, BDM19, BFK^{+}03, BC93, BEL04, CADA18, CGK08, CPT92, CCF13, CS98, CPW88, CF88, CK04, CGM10, CL90, Cha93b, Cha93a, CLZ^{+}15, CKP^{+}21, Cha87, Cha02c, CRV06, CJ93, CR95a, CLS95, CDDM05, CW13, CJBW13, CW18a, CW18b, CN21, CFCK22, CFKT17, CNPS15, CNS18, CH04, CS11a, CR87, CWZ10, CEPR10, CJPS13, CDP14, CDP16a, CA18, CD18, CA20, CP10, Col90, CH92, CCG^{+}93, CH97b, CGG90, Col91, CT96, CD89, Coo89, CM07, Cro92b, CGR93, CG94b, CR94, CL96, CGR99, CCG^{+}99, CKC07, Dai09, DR06, DS04, DA18, DGG^{+}19, DGM19, DOS93, Deo06, Der95, Dij76, Dijxx, Dit78, Dow91, Dow93, DC94, DGM90, DNR06, ĎHPT10, EMT23, FLM^{+}10, FWW13a, FWW13b, FWW13c, FLC^{+}19]. matching [FC98, FCLST07, FT95, FL13, FS22, Fat15, FHV18, Fen01a, Fen01b, FG95a, FMdB99, FDG^{+}11, Fil21, FBMA05, Fre03, FN04, FM06, Fri97b, Gaá04, Gal75, Gal76a, GS81a, GS81b, GS83, Gal84, GG86, GG87, Gal92, GP92, GU95, GPR95b, GR99, GU16, GGL94, GS00, GGF13, GG13, GMC02, GW92, GBY90a, GBY90b, GPN96, GF08, GFG11, GGN06, GL89, GV00, GM17, GZ10a, GS06, HIEH22, HWW07, HY92, HLS07, HFS05, HC87, HR01, HR03, HH93b, HH93c, HM96, HOK18a, HOK18b, HM00, HLS^{+}11, HBRV10, HP01, HP03, HK77, How96, HLN09, HHLS06, HFN05, Hyy08, IIT13a, IHEH23, IS96, Ier09, Ind97, Ind98, IS90, IIK08, IM13, II86, ISHY88, Jan23, JHU^{+}24, JM93, JP11, JL93, Joh95, Joh94a, JU91, KTP10, KSVJ15, KB22, Kas08a]. matching [Kas08b, Kas08c, KN00, Kes91, KTS^{+}98, KMS^{+}03, KST92, Kim99, KWL07, KEF^{+}14, KC21, KKNS23, KKH24, KH25, KNT11, KS01, KS05, KMP94, KS96, Kos89, Kos94, KM13, Kri09, KKR^{+}13, KST16, KGP^{+}05, KPA10, KT90, LMM17, LV86a, LV86b, LVN87, LV88, LV89, Lar99, Lec07, LLC03, LH13a, LH03, LG16, LS10, LLL^{+}24, LS09, LP08, Liu81, LHCK04, LBK08, LO94, LT97, LLL13, MCF^{+}11, MCF^{+}14, MK90, MNU05, Man76, MBY91, MMZ10, Mar07, ME97, MAI^{+}16, MP05, McI85a, MM03, MM07, Mis03, MHM^{+}01, MR09a, MBH20, MA12, MDLB02, MPW21, Mun95, Mus03, Mus05, MM89, Mye95, Mye99, NYuR15, Nak14, NNSP22, Nar91, NBY99c, NR00, Nav00, NKT^{+}01, Nav01a, Nav01b, NR02, NF04, NT05, NC06, Neb06, NWE99, NdMM02a, NC92, NR17, NK07, Nil90, OK94, dSOMY15]. matching [OR11, Oph89, OW03, PS89, PLL10, PPTT15, Par98, PS90, PLT14, PC99, PP94, Per94, Pet07, PMS11, PPZ08, PDC94, PA10, QZC17, Quo92, RM06, RTT02a, RUG97, RTO15, Rus88, RLP20, Sad93, SY23, SVS97, SRK20, STK06, Sal12, SBB19, Sas79, SW90, Sch81, Sch91a, Sch91b, Sch88b, SZ05, Sen00, SS94, SGYM00, ST96a, SN94, SSLL21, Shi97, Sil77, SR16, Smi91a, SDS14, SGCW14, SHCY93, Spe85a, Spe85b, Spi99a, Sri93, SA77, Sto02, SALP20, SWW^{+}12, SLZ^{+}20, SMW^{+}23, SN24, SV87, SNM07, Tak96b, Tak93, TBS06, TZYH14, TJD^{+}17, THQ19, TM04, TM05b, TM05a, THG17, Tej20, Thi93, TIT83, IIT13b, TLS16, TJMC20, TLLL07, TLLL09, TCC91, THL^{+}20, Ukk92, Ukk93, VRC24, Val09, Van06, VLP17, VW11, Via04, Vin77a, Vin77b, Vis90, Vol12, Wad87, WZS95]. matching [WGMH13, WLF14, WC14, WL15b, WZ96, WW03, Wat03, Wea94, XMLC11, YKGS11, Yao79, YT03, YB13, ZMAB03, ZZH10, ZZH16, ZA17, ZdSO18, ZZJC20, ZYX^{+}12, dB93, dRL95, BCKM15, HTK^{+}21, JD89, Neu10]. Matching-Based [CZCD09]. matchings [Iba97, RW10]. matchlib [Ano01]. MatchPy [KB18]. material [RH81]. Mathematica [Har97, Mae94]. Mathematical [Zha96, Rev91, Win78]. Mathematics [HM87, AH14, WSS94]. Matos [Pet95]. Matrices [CIK98, Gia93, PRU11, Lee82]. Matrix [BGP^{+}22, FTJ95, SJNS19, TZW94, Kar82]. Matrix-Vector [FTJ95]. Max [IMP01, WPKL13]. Max-Shift [IMP01]. Maximal [BJM79, Fra20, IF94, IS86, BGK^{+}16, Che96, GHST17, II86, Ukk92, Rep98]. Maximal-munch [Rep98]. Maximally [BNV^{+}13]. Maximize [LJZZ13]. Maximum [ADLM96, ASA17, OP16, ADLM01, LMMN07]. May [ACM69, ACM74, ACM76, ACM81, ACM84, ACM86, ACM90b, ACM91, ACM92d, ACM93b, ACM94c, ACM94d, ACM95b, ACM95c, ACM97a, ACM97c, ACM99a, ACM99b, ACM00, ACM08, Apo92, DT87, KLB12, NH11, SW94].

Mean [KP96a, KP96b, Alb89]. Means [Ray96, SS93a, OW03, WD99]. Measurement [Lee91]. Measures [Nav21a, EZ74, EZ76]. Mechanical [NEH90, Sha88a]. Mechanically [Gol90]. mechanics [NEH90]. Mechanism [JLH18]. mechanized [Chl08]. media [VD17]. Medicine [Sal01]. meet [KSVJ15]. Meeting [NEH90]. Membership [CGS17, GM02, KZ02, Loh10, Pet02, MW94]. Memory [GWX^{+}23, KKK11, LP13, Lec98, SRV^{+}19, TT20, TVCM12, FG99, JSH09, KFG15, LH13a, LMT16, Nak14, PLL10, YKGS11, YIAS89, ZYX^{+}12, OWP16]. MEMORY-Based [OWP16]. Memory-Efficient [GWX^{+}23, KKK11, LP13, PLL10, YKGS11]. Memristive [TT20]. Memristor [BYHT18, dLBHC22]. Memristor-Based [BYHT18]. Merging [Kit94, LK06]. Merlin [SBM^{+}18]. MESA [ZBST14]. Mesh [BM00, CNS18, DKP11]. Meshes [CG94a]. MeshFlow [DKP11]. Message [GM02, GKW^{+}10]. Meta [Moo12, Kes79]. Meta-level [Moo12]. META/REDUCE [Kes79]. Metadata [RHR^{+}21]. Metamathematics [Pux97]. Metamorphosis [KV15]. metaprogramming [JGMP22]. metasystem [GT90]. Method [AYS84, CLP95, Hua94, LPT12, MM93, NBY99b, RJK79, SV94, SD95, Shi96, WZJH12, FMdB99, GPR95b, Ker04, LLL13, MP09, SSYW19, TIAY90]. Methods [BW91, Fal85, GIMV03, SBHM94, BSY00, KNT15, Per94, San09, SHCY93, TTHP05]. Metric [BCP02, CN02, ZCS^{+}12, EH88, Jon13, Mag81, NC06]. Metrics [LP11]. Mexico [ACM92a, BYCC03]. MFC [AS04]. Miami [CVP86, IEE97]. Michoacan [BYCC03]. Microcomputers [ZA87, ZGE85]. microprocessor [BY92, BY96b, BY91]. Middle [Sal01]. Middleboxes [YDW18]. Miller [CR91]. Milwaukee [ACM81, IEE95a]. MIN [WPKL13, YD95]. MIN-MAX [WPKL13]. Minimal [BP11, BNV^{+}13, IF94, Mah07, Ned98, YD95, BGK^{+}16, BH96, HRN^{+}15, KS11a]. Minimization [CMNP18, Moh94, ND02, TZW94, AYCLS02, Kra08]. Minimizing [GS07]. Minimum [ASA17, FR17, PW93, Rot91]. Mining [BGM19, BNV^{+}13, GRS99, MSS^{+}19, NCJF18, ZKCY07, AVN22, EASK14, KD15, MR13]. MINIX [TSM88]. Minneapolis [ACM94c, SW94]. Minnesota [SW94]. MinneSPEC [KL02]. MIPS [ZZJC20]. Miranda [Tur86]. mirror [LG16]. Mismatch [ATX21, AEP06, JN24, Neb06, Sel84]. Mismatches [AW89, AJS92, BST^{+}03, HT17, KRS19b, KRS23, NR15, ALP04, CKP^{+}21, CW18a, CW18b, CGT25, Der95, FGKU15, GG86, GG87, GU16, GGF13, Gra15, GL89, LV86a, NR17, SZ05, KRS19a]. Missing [DCM15]. Mission [Lut02]. Mission-Critical [Lut02]. Missouri [IEE90]. Misuse [Hig95]. Mixed [ASA17]. mixture [Oph89]. MJRTY [AR13]. ML [AJ89, Ses96, SvS14]. MLIR [EZYA23]. MMDBS [DWE89]. MMIX [Ruc15]. MN [ACM94c]. MNCaRT [AWD^{+}18]. mobile [CFM00]. Möbius [Bjö93]. Modal [Yod91, tC09]. Model [CMR18, D^{+}23, FYJ^{+}17, GN19, GWvG10, Hig95, LLS12, MGH93, MGH97, PP09, SWY75, SCFC94, AB89, Liu81, NRO12, Tak96a, TG96].

Model-based [GN19]. Modeling [CLST^{+}13, Haz01, SBHM94, KC15]. Models [BZ98, CDL95, CDL99, A^{+}08, Coo89, CKC07, KCPC13, LT97, Nak14]. Modern [AAB^{+}17]. modification [Sch81, TJMC20]. modified [KJS17, Leu97]. Modifier [WWW^{+}16]. Modular [LP13, dSOMY15, PP06, MRA^{+}17, MFRW09]. Modulated [AEMS14]. Module [ZS17]. Moebius [JGMP22]. monadic [TPT13]. monads [PMS11]. monitors [ATdM07]. monoidal [VLP17]. Monoids [DM11, HY92]. Montreal [ACM94d, GS00, Lev95]. Moore [AG86, BYR92, Rai92, Smi94, Tal81, TJMC20, AR13, BYCG94, BYBDS09, BPMA02, Ber00, BKM95, CFG12, Col90, Col94a, DR06, EMC96, Gal79, Gol90, GO80, HA90, Kau92, KP96a, KP96b, KBN09, Kun95, Lec92, Men89, NT05, Pie95, Rus85, Rus92, Ryt80, STK10, Sch88a, Sto02, Tak96b, TU93, WW03, WW93, Woo86]. Moore-Like [CFG12]. Moore-style [WW03]. Morelia [BYCC03]. MorphJ [HS08]. morphology [Mus03]. morphology-driven [Mus03]. Morris [Bar81, DS04, Hua98, Moy24, TA90, Ukk10, PV91]. most [FL13, GFG11, HY90, LR14, YH91]. Motif [BNV^{+}13, Tan14, YHV^{+}15]. Motion [CZW15, KC99]. motivate [Fla88]. moves [CM07]. movie [WW93]. moyenne [Alb89]. MPI [MM02, MM03, PSK08]. MPI-IO [PSK08]. MR [Gro91a]. MRCSI [WL15a]. MS [VWR11]. MS-DFA [VWR11]. MSO [TN13, TN15]. MT [ZV97]. Mukherjee [Neu10]. Multi [AWD^{+}18, CJ93, FMdB99, GG95, GG97, Har02, HUN^{+}19, HMW^{+}23, LY17, LT03, LT90a, NBY99a, OR12, OSSK16, PLL08, PW12, SOR16, TMK^{+}02, WSW16, Alb89, ARS16, CPT92, CCG^{+}99, ETV88, JKNS00, KTP10, KM84, KPA10, OW03, SDA17, XMLC11, YT03]. Multi-Architecture [AWD^{+}18]. Multi-attribute [Har02]. Multi-byte [TMK^{+}02]. Multi-combinators [LT90a]. Multi-Core [LY17, JA17, JXA20]. Multi-Dimensional [GG95, GG97, NBY99a, JKNS00, XMLC11]. Multi-field [WSW16]. Multi-keyword [OSSK16]. multi-linear [ETV88]. Multi-matching [CJ93]. Multi-method [FMdB99]. multi-pattern [Alb89, CPT92, CCG^{+}99, KTP10]. multi-patterns [KPA10]. multi-resolution [OW03]. Multi-Stage [HMW^{+}23]. Multi-Stride [PW12]. multi-striding [ARS16]. Multi-String [PLL08, YT03]. multi-tenant [SDA17]. multi-term [KM84]. multi-text [YT03]. Multi-Threading [OR12]. Multi-Track [HUN^{+}19, LT03]. Multi-User [SOR16]. Multi-view [CJ93]. Multibox [Dya94]. Multicast [Sch95]. multicharacter [CW13]. multicomputer [TTO^{+}22]. Multicore [VSP08, YP13, ZLN11]. Multidimensional [SN92]. Multidisciplinary [Kni89]. Multidisk [KCK93]. multigraphs [FS13]. multihead [CR87, Pet94]. multilevel [KCPC13]. multilingual [ZV97]. Multimodal [BWG12]. Multipattern [STK06, Yun12, ZS13, BBK12]. Multiple [HR02, BYN97, BLP94, CF06, CJPS12, FL12b, Gaw12, IS94, IS96, KTS^{+}98, KMT^{+}01, KMM15, LLLC17, LT90b, LBK08, MM96, Mut00, OR12, PW95, SVS97, Shi96, SALP20, TM05b, VWR11, CK02b, Dai09, Fen01a, FN04, HFN05, KM13, KGP^{+}05, KIH15, Maa06, Mha05, NF04, PC02, PW06, SSLL21, WL15a, WZ96, ZC89]. Multiple-pattern [TM05b]. Multiple-Stride [VWR11]. multiplexing [Quo92]. Multiplication [BGP^{+}22]. Multiply [FTJ95]. Multiprocessing [WBA83]. multiprocessor [Vin77a, Vin77b]. Multistring [WZL^{+}23]. Multiterm [Bur84, Bur82]. Multithreaded [EGP14]. Multivariate [CvW18, YBTB23]. Multiview [ZCS^{+}12]. munch [Rep98]. Munich [ACM87]. Music [Cop91]. Musical [Cop91]. Mutations [ZJP^{+}18]. Myhill [WZU14].

N [Pux97]. namedCapture [Hoc19]. Names [VB12]. Nancy [Bun94]. nanolithography [SS93b]. narrowing [AEH94]. Nashville [ACM90a]. Natural [Fre06, GR96, vNG01]. Natural-Language [GR96]. Navarro [Hyy08]. Navigational [LRSV18]. near [HFFA09]. near-optimal [HFFA09]. Nearest [CEMW91, QQC^{+}13a, QQC^{+}13b]. Nearest-neighbour [CEMW91]. necessary [KT90]. Need [Gon02]. needed [AEH94]. Negative [YQW^{+}16]. neighborhood [KS11a]. neighbour [CEMW91]. Neighbourhoods [NRS18]. Nerode [WZU14]. nerve [Kle56]. Nested [KKFI^{+}24, Lar98, MZZ10]. Nesting [Jed87]. net [AB89, PP85]. Nets [CEW58, Sim83, BG91a, GR92, Kle56]. Network [BNV^{+}13, CFM17, CFS^{+}89, HMW^{+}23, LY17, LN19, Rei03, SBM^{+}18, TJGY22, VKPI17, ZL18, BKLE18, CMS08, GZ10a, LMMN07, LHCK04, SALP20, SN24, Tak96a, TLLL07, WXZY12]. network-based [BKLE18]. Networks [CLP95, DCM15, JGZL12, KTY^{+}18, MSP^{+}17, Ray96, SF01, TD18, CEMW91, Kin89, KD15, LLL12, LYWL08, MMDdJ11, SD91a, SD91b, SDA17, VD17]. Neural [AB89, CG87, CLP95, SF01, Kin89, PKK18]. neuromata [ŠW98]. neuropsychology [AB89]. Nevada [ACM95c]. next [KKP92]. next-generation [KKP92]. NFA [ARS16, Cha01, CP97, GS07, HM98, Hyy08, Lif03, PD12, YKGS11, ZYX^{+}12]. NFA-based [ARS16]. NFA-OBDDs [YKGS11]. NFA’s [CHP92, GLRÁ11, Lau00, IY02a, IY02b]. ngrep [McC01]. NIDS [TK07]. NIDS/NIPS [TK07]. Nineteenth [ACM92a, IEE95b]. Ninth [ACM90a, ACM97c]. NIPS [TK07]. nm [SS93b]. NMR [SHCY93]. No [CA20, Edw07, Gro91a]. node [SK17, SWW^{+}12]. noise [ZHWW12]. Noisy [Bra94, KO83]. Nominal [KST12]. Non [Ben94, CDM11, Liu14, ABH^{+}14, CCF13, CHL^{+}25, HJW^{+}92, SN94, TZH^{+}13]. non-aligned [SN94]. Non-Central [Liu14]. non-emptiness [CHL^{+}25]. non-exact [TZH^{+}13]. non-fixed [ABH^{+}14]. Non-Overlapping [Ben94, CCF13]. Non-periodic [CDM11]. non-strict [HJW^{+}92]. Nonbacktracking [MNR^{+}23]. Nondeterministic [ABMN20, Cha02a, CDL95, CDL99, DL03, HSW97, HSW01, Nag21, BT21, Gef03, GHR^{+}16, Ryt89]. Nonlinear [MSP^{+}17, RR92, Dan91, NA90]. Nonuniform [Lut02]. Norm [TZW94, WC14]. Normal [BMMR19, JR15b]. normality [Jan23]. normalization [KWLL08]. North [IEE89]. notation [LLS12, Man06, Rev91]. notations [Vol12]. Note [Ano17, BST^{+}03, CR92, Gra15, Gro91b, Ryt89, Sil77, tC09]. Notes [BYKZ^{+}92, BLPL92]. notion [Cha02c]. Novel [LLCC13, RVV23, LS99]. November [A^{+}08, IEE88, IEE89, IEE93, IEE98, NEH90]. Novice [JNS08, MLM^{+}08, PP94]. NP [TCC91]. NP-completeness [TCC91]. NR [Nav01b]. NR-grep [Nav01b]. NUCA [HFFA09]. Nucleic [CCL87]. nucleotide [LVN87]. Number [BM00, GPR95a, GS81a, Kod79, LS09]. Numeric [KAN^{+}17]. Nutshell [Gt92, Rob99a, Rob99b]. NY [AP10, Kap92].

O [PSK08, ZYQ^{+}15, dBB08]. OBDDs [CH04, YKGS11]. obfuscation [OSSK16]. Object [CJ93, GP93, LT90b, BY96b, Coo89, GPTV93, LLC03, MME14, TG96]. Object-Oriented [GP93, LLC03]. Objects [BZ98, SP16, Alb89, BGWXP22, HNB^{+}13, IM13, Mar89, MR05]. Oblivious [FV16, HLS^{+}11]. obliviously [FGG^{+}08]. Observations [Hun79, APTS13]. Obtaining [HW07, DR06]. OCaml [Fri06b]. Occurrence [CIL^{+}03, Cha02c, Mus05]. occurrences [FLSS93a, FLSS93b]. OCR [San95, TIAY90]. October [Bao93, IEE89, IEE90, IEE92, IEE95a, IEE97, IEE09]. off [MNS07]. offer [LS99]. Offers [Fri97b]. Offline [MT14]. offloading [KLL23]. offs [GHST17]. omega [SMT^{+}86]. omega-extended [SMT^{+}86]. Omni [Wol86]. Omni-Font [Wol86]. omnidirectional [KLL23]. On-Line [FG98, GG97, Lut02, Man75, MM93, Prů17, Tak86a, Tak86b, BEL17, CLP95, Fre06, Joh95, KNT11, NEL17, Shi96, CT96, FG95a, Gal75, NR02, TIT83]. On-the-Fly [PWSG25]. One [JKNS00, JL96, LY86, Sch91a, Sch91b, She59, WC14, Alb89, AGM05, Bak78, CX20, CR87, CCG^{+}93, GGL94, JL93, LMNT16]. One-dimensional [JKNS00, WC14]. one-letter [AGM05]. one-unambiguous [CX20]. One-Way [JL96, LY86, Sch91a, Sch91b, She59, CR87, GGL94, JL93]. Online [FL12a, FMP20, LH13b, PS10, CJPS13, FL13, KM13, THQ19]. only [CD18, GS81a]. Ontario [Cha86]. Ontology [RFD23]. Open [AWD^{+}18, Bol02, SDS14, AC93, ZdSO18]. Open-Source [AWD^{+}18, Bol02]. Operating [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d]. Operational [HH83]. Operations [DJ96, AGM05, Ear74, GW92, GH09]. operator [HC87]. Operators [For02, Kea91a, Sym85, MMDdJ11]. Optimal [AOK02, AR00, ABF94b, BH96, CLZ^{+}15, CG94a, CR92, CGH^{+}98, FCLST07, FG95a, FG98, FK16, FNU02, Gal84, Gal95, Gaw13, GG97, Hig86, IY02a, IY02b, KU99, KR94, LMMN07, MS98, MP88, Mor83, Mut00, Nak14, NWE97, NdMM02b, RT17, SN92, BKBB^{+}14, BG90, BKS02, CMO^{+}08, CR94, FN04, GS81b, GS83, Gal92, GR99, GHK14, HFFA09, IKX15, IP96, KR89, KT90, MSRR00, MS95, Neb06, PPTT15, Ryt89, ZC99]. Optimally [CCG^{+}93]. Optimising [Chi17]. Optimization [GC01, HJ99, IMM^{+}22, LT09, Sca11, Spi99b, Web95, All89, CK02b, KLL23, KWLL08, KGP^{+}05, SJ13, Spi99a, VW11]. optimizations [PSK17]. Optimized [AK09b, ELF22]. Optimizing [CJBW13, CJBW16, CMNP17, Kha16, LM01a, KS08]. Optimum [LD10]. Oracle [FPD08, GL03]. oracled [PPPdG20]. oracles [KW05]. Order [GU16, HW12, JR15b, KEF^{+}14, KAN^{+}17, KKNS23, SdM01, Wag74, BLSS03, Bjö93, CFKT17, Chl08, CNPS15, CT96, DGM19, Dow91, Gie90, JHU^{+}24, Kau92, Kes91, NRO12, OR11, OND98, Pie08, TPT13, Zei08]. Order- [Wag74]. Order-preserving [GU16, KEF^{+}14, KKNS23, CFKT17, CNPS15, DGM19, JHU^{+}24]. order-sorted [Gie90, Kes91]. Ordered [ST04, Cro92b, Gro91a, Gro91b, Mäk89]. Ordering [CDPP23]. ordinary [Rev91]. Oregon [ACM94a, ACM00, BGNP94]. O’Reilly [Ano97a, Ano12]. Organization [IK83]. Organizing [CG87]. Orientation [TCCK90]. Oriented [GP93, KS94, Rei77, GPTV93, LLC03, Mus05, TG96]. Orlando [IEE88]. Orleans [ACM91, ACM97b]. Ornaments [Rém17]. Oscillator [FYJ^{+}17]. Oscillator-Based [FYJ^{+}17]. OSN [ZGY^{+}16]. other [Ano97a, Fri97a, Hoc19]. our [FvGGM90]. outbreak [FNP09]. Outerplanar [BJM79]. Output [PM78]. Outsourced [FHV18, MBP22, YDW18]. Overcoming [Kül10]. overlap [KD15, PSK08]. Overlapping [Ben94, BZ98, CCF13]. Overlaps [AGM19]. Overlay [LT16]. overview [PVA^{+}92, Tur86]. Own [Tay97, ZGY^{+}16]. Oxford [Ano97b].

P [RUG97]. P2P [LLL12, LYWL08]. P4Rex [LLL^{+}24]. Pacific [IEE94a]. Package [Liu14, van14a, van14b, Kas08a, Kas08b, Kas08c, Nic03]. packages [Hoc19]. Packed [Zha17, BKBB^{+}14, GGF13]. Packet [BYHT18, LLLL08, LLC17, LMT16, VKPI17, VWR11, YP13, ARS16, BAC12, CMS08, NYuR15, SSYW19]. Padova [Apo93a]. PageRank [LSV08]. pages [Ano97a, Ano97b]. Pair [BNV^{+}13, PPZ08]. Pairs [GLS92]. Paisley [TL12a, TL12b]. Palermo [GM11]. Palindrome [Gal76b, IIT13a, Man75, IIT13b, Gal75, Gal76a]. Palindrome-Recognition [Gal76b]. palindromes [NHN^{+}20]. Palindromic [HIRS17]. Palo [IEE93, IEE98]. PAMA [LCL06]. paper [Pet95]. Papers [Cro92a, Moo64, ACM69, ACM74, ACM76, ACM81, ACM92a, ACM92c, ACM93a, ACM94a, ACM95a, BW91, IEE92, IEE93, A^{+}08]. paradigm [AC93, JHU^{+}24]. PARAID [WOQ^{+}07]. Parallel [Ash85, BL94, BG92, CF06, CG87, Che96, CHL14, CCL87, CR92, CGG^{+}97, CGH^{+}98, DK13, ECSS88, FL99, GG87, Gal95, GJ16, GHK^{+}91, GZ94, GS85, GIG77, GMAS22, HT17, HN02, HN05, IS86, KR94, KKK11, Kül10, LLCC13, LLLC17, MS01, MR09b, Mut00, NR98, RR90, SV94, SN92, TLC15, TVCM12, VMML15, Wei83, BGNP94, BLPL92, BYK22, BG90, BG91b, Bre94b, Bre95, BH96, CL09, CCG^{+}93, CR91, CR94, Gal84, Gal92, GS93b, GS93c, GF08, Hur84, Hyy08, II86, JRV96, JHU^{+}24, KIH15, LV89, MK90, Mis03, MMS14, NYuR15, Ryt89, SY23, TTO^{+}22, TLS16, ZC99]. Parallel-Algorithms [SV94]. Parallelism [JA17, JXA20, MKF91, Wri94, ASM17, CFKT17, HFN05, LV86b, NR00, RW93, SBB19]. Parallelizable [ATX21]. Parallelization [KP93, HA90, NE93, RP95]. Parallelizing [HN90, MIH17]. Parameter [Jok90, Sut21]. Parameterized [Bak96, BRL13, BDFW94, CHLT14, IS94, OP16, PA10, AFM94, Bak93, BA16, CGK08, FM06, HLS07, IS96, KC21, KPA10]. Parameters [CJBW16, CJBW13]. Parametric [Chl08, HPM94, WAH23]. parametricity [Rém17]. parentheses [PDC94]. parentheses-matching [PDC94]. Parenthesis [Sto96]. Paris [Cro92a]. Parity [MSV24]. Park [IEE89]. ParsCit [PKK18]. Parse [Fre78, Kea91a, DF00, Fre78]. Parser [Hol84, TB00, Gan89b, LK06, MLC08, PKK18]. parsers [Dya94]. Parsing [AU72, AU73, BRO16, Cam99, DA20, Gor00, MGH97, NH11, RD17, Rus88, SL14, BvdM17, BG22, BBM21, BBRM23, GHR14, MMI14, Ier09]. Part [CDPP23, JLH18, KP15, Kul11, Nav21a, Nav21b]. Part-of-Speech [JLH18, Kul11]. Partial [Ant95, Ant96, AYS^{+}24, CW84, CD89, GL01, KK08, KMR21, Mor83, Ses96, Smi91a, Zve80, ADR03, ADR06, DR06, HR03, Jør92, KCK93, MR09a, ST19]. Partial-Match [Mor83, Zve80]. Partially [ZMSD93, HY92]. Parties [XZL^{+}19]. Partition [CF85, FR17, WL15b]. Partitioning [Fat15, Kim99, LYWL08, Mid96]. Partitions [Knu05]. partners [LLL12]. Pascal [Liu86, Sha88b]. Paso [ACM97c]. password [KJS17, MW94]. Paste [Lud77, AM97]. patches [TCCK90]. Path [Bac94, BLLW12, CDLV99, CDLV02, HJ99, LM01b, LOZ^{+}24, Sch22, SNM^{+}13, Tar81a, Tar81b, TPT13, ZJP^{+}18, BBG13, CJR^{+}21, Che96, CK02b, LM12, MF96, PC02, SVM14, YCJK08, YT03]. Path- [TPT13]. Path-space [SNM^{+}13]. path-wise [MF96]. Paths [MNP^{+}23, GLS07, LM13]. Pattern [AMB^{+}02, ABM08, AKW85, ABF94a, ABF96, AAL97b, ALL97, AAL^{+}97a, ALLL98a, ALL00, AAL^{+}00, ALR08, AAB^{+}09, AP10, AWS16, Ano92b, Ano96a, Ano96b, Ano17, AYS84, iA94, AYS^{+}24, AG84, AG97, AT02, ADLM96, AW89, Ash85, AJS92, AGS96, AD11, BYN98, Bak96, BCD98, BEM^{+}12, BLP18, Bee81, Bee13, BKLP97, BKL^{+}02, BCKM15, BBL93, BBL98, BYNTK21, Bow87, BTC06, BL16, BGJ01, BCFL12, BC93, BCC^{+}13, CCFG12, CFM17, CS98, CDM11, CG87, CK04, CLST^{+}13, Cha02b, CZCD09, CWL^{+}21, CK92, CDEK95, CG94a, CLP95, CM08, CL95, CHZ06, CEPR10, CJPS12, CDP14, CH03, Col94b, CG79a, CG79b, Cro92a, CR92, CR95b, CGPR95, CL96, CGH^{+}98, D’A98, DB86, DWE89, DLG12, D^{+}23, DN77, Dit78, DCM15, DGM94, Dwe00, EIV04, EF13, EGP14, Far92a, Far92b, FMMS20, FL08, FR00]. Pattern [For02, FNU02, Fu95, Fu96, Fu97, GHLW15, GPP04, GC01, GRS99, GIK97, GP01, GP03, GIMV03, Gaw12, Gaw13, GP93, GM02, Gia93, GG95, GG97, GM11, GMC02, Gil85, GW92, GKP19, GWX^{+}23, GGN06, Gri79, Gri83, GL01, Gro92, GL86, Har02, Har97, HAR10, HH83, HL10, HT14, Haz01, Hea71, HEWK03, Hei01, HL97, HUN^{+}19, Hig95, HO82, HMW^{+}23, HSTS01, HGT24, How97, HW12, CVP86, IMR08, IST05, IMM^{+}22, JLK^{+}20, Jez15, JSC83, KPR97, KPR00, KU99, KS12a, KR81a, KR81b, KR87, KR94, KRS95, KRS97, KN00, KP93, Kes91, Kes79, KTSA99, KMT^{+}01, Kid09, KS99, KKSL01, KKK11, KS01, KS06, KM92, KM95a, KM95b, KMP77, KRR17, KRRW24, Kor83, Kra08, KB18, KK02, KU09, KNS12, Kül10, KVX12, KNMH00, KC99, Lab12a, Lab12b]. Pattern [LV94, Lav91, LP13, LM01a, LKL02, LSTW^{+}17, LY17, LT03, Les95, LV06, LYT^{+}23, LTL04, LA12, LLCC13, LJH^{+}17, LLC17, LP11, Liu86, Liu88, LM02, Lut02, MZ07, MS98, MBP22, MKF91, MU02, MW92a, MW92b, MGW14, MR11, MNP^{+}23, MSV24, MHT09, MUHT96, McI85a, McI85b, MS01, Mon17, Mu 95, MuT95, Mut00, Mye92, Nao91, Nar91, Nav98, NBY99a, NR99b, NBY01, NR03, Nav04b, NWE97, Ned98, NdMM02b, ND02, Neu10, NRS18, NCKL14, OR12, OP16, OW03, Ott94, PDL98, PS10, Par96, PV91, Pet92, PW95, Pit98, PPZ08, PP09, Pou93, PK85, PS93b, QTO^{+}20, RR90, RR92, Rao95, RM88, RKM21, RS98, RDRF24, RHR^{+}21, Ric79, Ris16, RSG^{+}19, SMD04, SCFC94, SN92, SP16, Sch95, SRR92, SRR95, Sel84, Ses96, Sha93, SN94, STSA99, SKF^{+}00, Shi00, Shi04]. Pattern [SSSS10, Sim83, SF01, SdM01, SW09, Som82, Spi99b, SN24, TBD22, Taf22, Tak86a, Tak86b, Tak94, TT22, TMK^{+}02, TM05a, TMV^{+}01, TSI13, TK07, TL12a, TL12b, Ukk10, UFA^{+}24, VSM87, VWR11, Via02, VG01, VRD01, Vis91, Vis99, Vol12, VS01, VB98, WAH23, WCM^{+}94b, WZS95, WSW16, WZL^{+}23, Wat96, WKA94, WD99, WBA83, WM92b, WSVS22, Xi03, YP13, YK11, YDW18, YGG^{+}23, ZZ12, ZZH10, Zha17, ZdSO18, ZLN11, ZT89, Zue96, ADR03, ADR06, AK08, AK09a, Akl78, Alb89, ASG99, AYCLS02, ALV92, ALLL98b, AL01, ABC^{+}04, AKT06, ALLS07, ABH^{+}14, Ano01, Aoe89, Apo92, Apo93a, AH97, ACP05, AP90, ADLM01, AG06, BKLE18, BYR93a, BYR93b, BYCC03, Bak93, BDB90, BEM^{+}13, Bar22, BA15, BA16, BYB^{+}24, BCD14, BPPR20, Bir77a, BGJ89, BÖ13, BGWXP22, BYK22, Bra95]. pattern [BBHK14, BKS02, BDM19, CGK08, CPT92, CPW88, CF88, CGM10, Cha93b, Cha93a, CKP^{+}21, Cha87, Cha02c, CRV06, CR95a, CLS95, CFCK22, CFKT17, CNPS15, CNS18, CS11a, CWZ10, CJPS13, CDP16a, CA18, CD18, CA20, Col90, CCG^{+}93, CH97b, Col91, CT96, CD89, CGR93, CG94b, CR94, CCG^{+}99, CKC07, DS04, DA18, DGG^{+}19, DGM19, Dij76, Dijxx, Dow91, Dow93, DGM90, EASK14, ETV21, FLM^{+}10, FWW13a, FWW13b, FWW13c, FLC^{+}19, FC98, FHV18, Fen01b, FBMA05, Fri97b, Gaá04, GP92, GU95, GR99, GU16, GS00, GGF13, GG13, GPN96, GJS20, GZ10a, GS06, HIEH22, HWW07, HC87, HM96, HBRV10, HP01, HP03, HK77, How96, HLN09, IIT13a, Iba97, IHEH23, Ier09, Ind97, IM13, ISHY88, JGMP22, Jan23, JHU^{+}24, JM93, JP11, Jon07, KTP10, KSVJ15, KS07]. pattern [Kas08a, Kas08b, Kas08c, KTS^{+}98, KMS^{+}03, KCK93, Kim99, KKNS23, KKH24, KH25, KS11a, Kin89, KS05, KMP94, KD15, Kos89, Kos94, KM13, Kri09, KKR^{+}13, KGP^{+}05, LLC03, LH13a, LH03, LS10, LS09, LP08, Liu81, LBK08, LO94, MCF^{+}11, MCF^{+}14, MK90, Man76, MMZ10, Mar07, MAI^{+}16, MP05, MHM^{+}01, MR09a, MR13, MA12, MDLB02, MPW21, Mun95, NYuR15, NNSP22, Nav00, Nav01b, NR02, NWE99, NdMM02a, NR17, NK07, Nil90, OK94, OR11, Oph89, OSSK16, Owo93, PPTT15, Par98, PS90, PC99, Per94, PMS11, Quo92, RM06, Rus88, SRK20, SBB19, Sas79, Sch81, Sch91a, Sch91b, Sch88b, Sen00, SGYM00, SSLL21, Sil77, Smi91a, SDS14, SGCW14, SHCY93, Spe85a, Spe85b, Spi99a, Sri93, Sto02, SALP20, SNM07, TZYH14, THQ19, TM04, TM05b, Tej20, Thi93, TIT83, IIT13b, TLS16]. pattern [TJMC20, Val09, Van06, VW11, Via04, Vin77a, Vin77b, Vis90, Wad87, WCM^{+}94a, WGMH13, WC14, WL15b, WZ96, WW03, Wat03, Wea94, Yao79, YCJK08, YBTB23, ZC89, ZMAB03, ZA17, ZCÖ09, ZCÖZ12, dRL95, JD89, YIAS89, Ano97b]. Pattern-Based [EGP14, Far92a, Far92b, KS07]. Pattern-Directed [Kor83]. Pattern-Match [Pet92, GJS20]. Pattern-Matching [FR00, KPR97, KPR00, KR81a, KR81b, KR87, KRS95, KRS97, KP93, KVX12, LY17, Lut02, MUHT96, NWE97, Ned98, Ott94, Pou93, SCFC94, Sch95, SSSS10, SW09, WM92b, CL96, GMC02, KN00, CF88, Dijxx, Fri97b, Gaá04, Ier09, KSVJ15, LH13a, Nav01b, NWE99, NdMM02a, OR11, Per94, Sch88b, Wea94]. Pattern-Recognition [AWS16]. Patterns [BLR14, BH85, BGS23, CLP98, CMNP17, CMNP18, Gim73, HNB^{+}13, IS94, JGZL12, Kha16, Les79, LSWP19, Prů17, SB09, TMV^{+}01, ADT15, Alb89, AG06, BLR11, BH13, BSM^{+}07, BFS04, Bro77, CP10, Dan91, ETV88, IS96, JSH09, KPA10, KIH15, KRML09, LMM17, MR09a, NdMM02a, Tak93, Ver92, Vou06, Wal89, ZKCY07, ZJL14]. payload [NNSP22]. PCRE [Anoxx]. Pearl [KN12, ADU16, DA20, FHW10, JR15a]. Pearls [Ben86, Ben00, Bir10]. pebbles [EHS07]. peeling [ALLT11]. Peephole [All89, Spi99b, BA06, Spi99a]. peer [AB09]. peer-to-peer [AB09]. Penalties [KM92, KM95a]. Pennsylvania [ACM76, ACM99a, IEE92]. people [Mah07]. pep [Woo86]. Peptide [LZ18, SVS97]. Perfect [LLLC17, XMLC11]. perform [MW92b]. Performance [FWW12, HKL^{+}14, IS90, JLK^{+}20, KNT15, Lee09, MM02, MM03, PRT^{+}16, RSG^{+}19, Sca11, WGL^{+}21, YP12, YK11, YJ84, CGM10, Fen01b, Hur84, LH13a, SNB^{+}19, SWZ01]. perils [Fen01b]. Periodic [Mat94, CDM11, FLSS93a, FLSS93b, ZKCY07]. Periodicities [Sli83]. Periodicity [GPP04, MAI^{+}16]. Perl [Lab12a, Lab12b, Ano97a, Anoxx, Fri97a, Han01, LT09, Mah07, NTS93, SPF08, Sno01, SM04, Stu07, Val09]. Perl-based [NTS93]. Permutation [BL16, MSV24, CNS18, KKR^{+}13, BCKM15]. Permutations [BBL93, BBL98, Chr96, Iba97, Jan23]. Permuted [HUN^{+}19, BEL04]. Perrin [Bre93, Bre96]. Persimmon [KKFI^{+}24]. Personal [VB12]. Pesky [CJBW16, CJBW13]. Petri [GR92, PP85]. Phase [FYJ^{+}17, CK02b]. PhD [HF13]. phen [Lia84]. Phi [TLS16]. Philadelphia [ACM99a]. PHP [B^{+}05, GSL17, ST03, SM04, Stu07]. physically [SNM^{+}13]. physically-based [SNM^{+}13]. Picking [CJBW13, CJBW16]. Piconets [LTL04]. Pictorial [DOS93]. picture [Mar89, MPW21, TCC91]. Pictures [AGM19, JSC83, Tak93]. Pipelined [PLL08, ISHY88, PLL10]. PipesFS [dBB08]. Pisa [FL08]. Piscataway [FC98]. Pitfall [OK94, KOI94]. Pittsburgh [IEE92]. PL [FPD08]. PL/SQL [FPD08]. place [HTX17]. Placed [KBB01]. placement [HFFA09]. Plagiarism [PAMP12]. Planar [CM08, Hig86, TZYH14]. plane [AK09a]. planner [HRN^{+}15]. planted [Tan14]. PLAs [KTU87]. plasma [AP90]. Platform [HZ13, MFMA15, MFMA17, PYY19, ZLN11, FNP09]. Platform-Specific [MFMA15, MFMA17]. play [FHW10]. Plexus [AB09]. plush [II09, MI07]. Plushie [MI07]. PMETA [Kes79]. Pocket [BK09, FPD08, GL03, Stu03, Stu07]. PODS [ACM95b, ACM99a, ACM07, HF13, ACM90a, ACM92b, ACM94c, ACM95b, ACM97a, ACM98]. PODS’08 [LL08]. PODS’11 [Len11]. PODS’12 [KLB12]. PODS’13 [HF13]. Point [CM08, GIMV03, Hig86, MU02, Ukk10, VS01, WKA94, ZHWW12, dRL95, AK09a, ACT10, CGK08, CS98, Rot91, TZYH14, WC14]. Point-Pattern [MU02]. pointer [MF96]. points [Jon13]. Poland [Win78]. Polaris [Wea94]. Policy [LTL04]. Polling [LTL04]. Polymorphic [Vou06]. Polymorphism [KKFI^{+}24, DRW95]. Polynomial [BCC^{+}13, ISNH94, PW93, EMT23, FLM^{+}10, GH09, Liu81]. Polynomial-Time [ISNH94]. polypeptides [SHCY93]. Pool [BTTF02]. POPL [ACM94a, ACM95a, ACM87]. Population [TMV^{+}01]. Population-Based [TMV^{+}01]. Portable [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, ZGE85]. portfolio [LH13b]. Portland [ACM94a, ACM00, BGNP94]. Posets [FK16]. Position [OS11, PRU11, SOR16]. Positions [IY02a, IY02b]. POSIX [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, ADU16, BBvdM21, BT21, SL14, Urb23, Kuk17]. Possibilistic [WD99]. possibilities [LS99]. Post [KGP^{+}05]. Post-compilation [KGP^{+}05]. Postprocessing [RJK79]. PostScript [HV93]. Potential [KNMH00]. Power [Bao93, Wol90, KPP21, WOQ^{+}07]. power-aware [WOQ^{+}07]. PowerEN [HKL^{+}14]. Powerful [Ano97a, Pol01, Fri97a, NTS93]. Practical [ALR08, BYP92, CSY03, DNM00, HHM^{+}13, Hor80, IMM^{+}22, KK95, Lut02, NT20, NR99b, PPTT15, TTHP05, BYP96, CCG^{+}99, ĎHPT10, Maa06, NR02, PSK17, WR15]. Practicality [TT82]. Practice [CCFG12, CJBW16, DGBH93, KP99a, CJBW13, DRSS96, WL15b, Lut02, KKP92]. PRAM [Apo93b, PDC94, dB93]. Pratt [Bar81, DS04, Hua98, Moy24, PV91, TA90, Ukk10]. Pre [BAM^{+}24]. Pre-Alignment [BAM^{+}24]. Prealignment [KYG19]. Precessions [SS93a]. Precise [PIR17]. predicate [BG91a, MFRW09]. predicate/transition [BG91a]. predicates [Gan89b, STKD20, TP07a, TP07b]. Predicting [Git96]. prediction [Fen01b]. Predictive [KKSL01]. Prefetching [JDXD13]. Prefix [BM00, BCT94, BCT98, CIK98, GLS92, HWW07, Han13a, HAR10, HTK^{+}21, RT17, BBHK14, BCT93, Bre95, BFK^{+}03, DK13]. Prefix-Free [Han13a, HWW07]. Prefix-Matching [BCT94, BCT98, BCT93, Bre95]. Preliminary [GS81b, LS94, WCM^{+}94b, Kos94, WCM^{+}94a]. Preprocessing [Nao91, ZCH23, CCG^{+}93, Ryt80, Sto02, Tak96b]. Presence [HT14]. presented [ACM69, ACM74, ACM76, ACM81, ACM92a, ACM92c, ACM93a, ACM94a, ACM95a, NEH90]. Preserving [SJNS19, XZL^{+}19, CFKT17, CNPS15, DGM19, GU16, JHU^{+}24, KEF^{+}14, KKNS23, QPWH08, VRC24]. Press [Ano97b]. Prevention [TS05, TBS06, VS11]. previews [Chi17]. prime [LS09]. Primer [Lut02, YC22]. primitive [Dow93]. Primitives [KM94]. principle [CL09]. principled [Cox10b]. Principles [ACM83, ACM87, ACM90a, ACM92a, ACM92b, ACM93a, ACM94a, ACM94c, ACM95a, ACM95b, ACM97a, ACM98, ACM99a, ACM06, ACM07, AP90, DRSS96, HF13, KLB12, LL08, Len11, MMZ10, KKP92]. priorities [CFLH^{+}22]. Priority [Lav91]. Privacy [MBP22, SJNS19, XZL^{+}19, VRC24]. Privacy-aware [MBP22]. Privacy-Preserving [SJNS19, XZL^{+}19, VRC24]. private [SRK20, WR15]. privatization [RP95]. Probabilistic [AJS92, CZOdlH17, LSV08, MHKR12, Sch91a, Sch91b, TMV^{+}01]. Probabilities [PM78, Neb06]. Probability [SCFC94, SD95]. Problem [BCT94, BCT98, CF06, CK02a, F^{+}23, FR17, HL94, Hui92, KO83, KS11b, KS12b, KZ02, Pet02, PW93, Sut21, Yan95, Akl78, B^{+}05, BSTU08, BC06, BCT93, CNS18, CM07, Dij76, Dijxx, FL13, Goo05, Gra15, Gro91a, GZ10b, Hov12, Kar82, KD15, Lei81, Maa06, Mae90, Mäk89, Man76, RTO15, Rob79, Sch81, SZ05, SSLL21, Tak96a, Tan14, TTO^{+}22, Tho81, YT03, YHV^{+}15, tC09, GLS92, BLPL92]. problem-based [BC06]. problem-solution [B^{+}05, Goo05]. problem-solving [Tak96a]. Problems [CK92, Gon02, Hea71, KPR97, KPR00, Loh10, MNS10, OP16, RS59, SV94, Tar81a, Tar81b, Via02, AI18, FMdB99, Gro91b, GHK14, HY92, HTX17, Hun79, Ind98, Kra08, Mid96, Nic03, Sri93, SH85, Via04]. Procedural [Sym85]. Procedure [Gin67, CS11b, HW09]. procedures [MP88, TN13, TN15]. Proceedings [ACM90a, ACM92b, ACM94b, ACM94c, ACM95b, ACM97a, ACM98, AGS93d, DGBH93, KP15, KLB12, Len93, Len11, Lev95, SC04, SC05, SM09, ACM81, ACM08, ABB93, B^{+}02, Bro93, FGR72, HF13, IEE92, IEE93, IEE94a, IEE98, Kap92, MG94, SC93, SC99, SC01, SC02, SC03, SM10, WN90, Win78, ACM83, ACM84, ACM86, ACM90b, ACM91, ACM92c, ACM92d, ACM93b, ACM94d, ACM95c, ACM97b, ACM97c, ACM99a, ACM99b, ACM00, ACM06, ACM07, AL01, AP10, AGS93a, AGS93b, AGS93e, AGS93c, Ano87, AAC^{+}01, AT02, AOV^{+}99, BYCC03, Bao93, Cha86, DT87, DMVT13, FC98, FL08, FJ92, FMA02, GS00, GM11, Hwa85, CVP86, IEE88, IEE90, IEE94b, IEE95b, IEE09, KS12a, KU09, LL08, LV06, MZ07, PC99, SMD04, SW94, Sto92, SC98, USE92, Apo92, Apo93a, AH97, ACP05, BGNP94, Bun94, CG94b]. proceedings [GU95, HM96]. Proceedings. [BGG^{+}94]. Process [Gro91a, HHW^{+}99, Sid95, Sid99, Sid00, Sid02, VVV04]. Processes [SBF80, AB89]. Processing [AKW85, AWD^{+}18, CCL87, GWX^{+}23, GSL17, JLK^{+}20, LLLC17, LPJ23, Lud77, PYY19, SRV^{+}19, SK17, TMK^{+}02, VCS^{+}12, VKPI17, dLBHC22, vNG01, CL09, CK08, ECSS88, FGR72, Gre88, KNT15, KSH^{+}15, Kit94, QWX^{+}13, ZXL^{+}13]. Processor [HKL^{+}14, RSG^{+}19, TT22, LHCK04, ME97, MM07, Sas79, TLLL07, WKR09, CPW88, Sca11]. processor-based [LHCK04]. Processors [AWS16, LY17, VCS^{+}12, VSP08, YP13, TLLL09, YIAS89]. Production [DWE89, MUHT96]. productivity [AP13]. products [Yod91]. Professional [HZ13, JNS08]. Profile [FhDAF09]. Profile-guided [FhDAF09]. Profiles [SB09, PJP^{+}18]. profondeur [Alb89]. Program [ASJDW21, CDL^{+}15, JP73, Mu 95, MuT95, Wea94, Fil21, HR01, Kod79, KMMPN85, Mag81, MP09, Mun95, Pel87, Pep91, Pra97, WZS95]. Programmable [CFS^{+}89, GHK^{+}91, LK90, OA17, WZL^{+}23, LK88, LLL^{+}24, MM07]. Programmer [AKW85, BF97, HHW^{+}99, Mae94]. Programming [ACM92a, ACM92c, ACM93a, ACM94a, ACM95a, AGT89, AWS16, Ben86, Ben00, Ham88, Hud89, KP99a, KN17, Knu05, Kor83, Mae94, Mor90, Mye98, ND02, NM10, Ruc15, Ste12, Sym85, Tho68, VHC88, Wen93, ACM87, APTS13, AC93, BDMT16, BAC06, BW91, BK89, CM86, Coh90, DV21, Ear74, Fla88, HJW^{+}92, KS08, KN19, Mor02, Mye99, Pie08, Rob87, Smi91a, Str13, YIAS89]. PROGRAMS [BY91, Bir77b, EGP14, RND97, VMML15, Web95, BY92, BK86, Gol90, HN90, Jon07, KOI94, Kap69, NWE99, NAR08, OK94, OR11, Rus92, SLTB^{+}06, dH05]. PROGRES [Zue96]. Progressive [CZW15, XMLC11]. Project [YC22, BW91]. Project-Based [YC22]. Projections [Wei83]. PROLOG [Wei84, CF88, Gan89a]. Proof [Asp12, CD18, Fil21, KN12, Sha88a, WW93, ADU16, CDP16b, GO80, PPTT15, Rus85, Pux97]. Proof-directed [Fil21]. Proof-relevant [CD18, CDP16b]. ProofChecker [SBR^{+}07]. PROOFS [BY91, GHW05, BY92, BY96b, KPP21, SBR^{+}07]. Properties [Sli78, AS85, FWDL15, LMN16]. property [KKM^{+}85, LM13]. Proportion [AW89]. Proposal [Liu88]. Protein [BNV^{+}13, CLST^{+}13, CCL87, LSTW^{+}17, Les95, LZ18, MGW14, NW13, NR03, Pol13, VG01, FLSS93a, FLSS93b, HAI02, KD15]. Protein-DNA [CLST^{+}13, LSTW^{+}17]. Protocol [XZL^{+}19, AB09]. Protocols [GKW^{+}10, HL10, HSL10, SRK20]. Prototype [Mu 95, MuT95, Mun95]. Prototyping [CFS^{+}89]. Provably [GH13]. Prover [Kau92, KP96a, BKM95, Gol90, HA90, KP96b, Moo12, Pie95, Rus85, Rus92, WW93]. provers [VLP17]. providing [BH07]. proving [AGH^{+}17]. proxies [KT14]. Proximity [BYCMW94]. Pruning [ZGS^{+}15, PC02]. PTME [SXZ^{+}25]. pu [Lia84]. publish [ZCT14]. publish/subscribe [ZCT14]. Publisher [KRS23]. PUF [RMK^{+}14]. Punctuation [GWvG10]. Purdue [Hig95]. purely [HJW^{+}92]. Purpose [VCS^{+}12, WSVS22, AAB^{+}86, Sch91a, Sch91b]. push [GOMSJVGP08]. push-down [GOMSJVGP08]. Pushdown [Coo72, MS98]. Puzzle [BK93g]. pyramid [SS94]. Python [DV21, KB18, LR14, Rom14, SM04, Stu07].

Q [MP88]. Q&A [Cal00]. Q-Coder [MP88]. QED [Ritxx]. quadratic [EMT23]. quadtree [SS94]. quadtrees [BK89]. qualitative [MLM^{+}08]. Quality [LLW^{+}15]. quantification [Kau92]. Quantitative [ARM^{+}19, Liu88, Sca11, YPG21, MRA^{+}17]. Quantum [IMM^{+}22, JN24, Mon17]. quasi [Kos94, NCV10]. quasi-equally [NCV10]. quasi-real-time [Kos94]. Quebec [ACM94d]. Queries [AKT20, BYCMW94, BLLW12, CDLV99, CDLV02, FLS98, HJ99, Kin92, KRRW24, LOZ^{+}24, Sch22, WWW^{+}16, ADT15, Arn93a, Arn93b, BBG13, CJR^{+}21, CK02b, FWW12, FLC^{+}19, IKX15, KC15, LMRT14, MRA^{+}17, SVM14, ZXL^{+}13, ZCÖZ12]. Query [AAB^{+}17, FLS98, KM94, LLS^{+}20, AYCLS02, AL08, BLSS03, BFS00, CDL08, CMW87, FWDL15, HRN^{+}15, KNT15, KSH^{+}15, LRV13, MZZ10, PC02, QWX^{+}13, RM06, XQW^{+}13, YBTB23, ZCÖ09]. query- [FWDL15]. Querying [AGP18, BLR11, BLR14, CMNP17, KTY^{+}18, LM01b, LMV16, San15, TV14, CM95, GW92, LLS12]. Question [KKSL01, AFI98, CKC07]. Question-Answering [KKSL01]. queues [SWZ01]. Quick [Rob99a, Rob99b]. Quickscan [Sil77]. QuikFind [Cha91]. quirky [MLM^{+}08]. Quoted [PSK17]. Quotient [FPP08]. QVT [LLS12].

R [Hoc19, KW19, Liu14, Ram94, Val09, van14a, van14b, Lab12a, Lab12b]. R-Trie [KW19]. Rabin [GBY90a, GBY90b]. RAID [WOQ^{+}07]. Ramsey [Kun95]. Random [Ale94, NR21, RKH02, SD95, CADA18, CRV06, Jan23, Rou21, SN94, Yao79]. Randomized [AJS92, ACD01, BST^{+}03, CGG^{+}97, KR81a, KR81b, KR87, TWZ23, AGW13, CH97b, II86]. range [HFI^{+}08]. Rapid [CG79a, CG79b, Gri79, Bak78, AWS16]. Rapidly [Dav73]. RASSA [KYG19]. Raster [AGS93a, AGS93b, AGS93d, AGS93e, AGS93c]. Rationale [IEE01c, IEE01d]. Ratios [Huc21]. ray [SS93b]. RDF [HRN^{+}15, KSH^{+}15, LRV13, LRSV18]. Re [BvdM23, MCP17, ORT09, CGR02, CGR03]. re-examined [ORT09]. Re-examining [BvdM23]. RE-Tree [CGR02, CGR03]. Re-Vectorization [MCP17]. re2 [Cox10b]. reachability [FWW12, GZ10b]. reachability-bound [GZ10b]. REACT [JLK^{+}20]. Reactive [HFFA09]. Read [BAM^{+}24, GMAS22, KYG19, FSL^{+}15]. reading [BWG12]. reads [FSL^{+}15]. Real [BG14, Gal76a, Gal81, HHW^{+}99, Kos94, MNR^{+}23, Sli78, Sli83, VSP08, BGM13, CM90, Gal75, MDLB02, NNSP22]. Real-Time [BG14, Sli78, VSP08, Gal76a, Kos94, BGM13, Gal75, NNSP22]. Real-World [MNR^{+}23]. Realization [CEW58, Kle56, TB00, XK92]. Rearrangement [AAB^{+}09]. Reasoning [ADT15, GNU94, PS88, BGWXP22, KS07, PS90]. Rebus [Gri85]. Recalibration [BM08]. recipes [B^{+}05, Goo05]. Recognition [ACR01, AWS16, BGJ01, CFP19, CG87, CJ93, CTF^{+}98, CLP95, D^{+}23, Gal76b, GS93a, CVP86, LT90b, PDL98, PG90, SA96, WD99, Wol86, Yam01, AAB^{+}86, BWG12, Gal75, Gal76a, KSWC93, KKM^{+}06, KKM^{+}13, Kin89, Sel84, SKS96, ZC89]. recognizable [HY92]. Recognized [RJK79]. recognizers [Fos89]. Recognizing [Ray96]. Recombinant [Cop91]. Recompression [Jez15]. Recomputing [HPM94]. Reconciling [IM13]. Reconfigurable [BM00, MLC08, NNSP22, ZL18, CMS08, Ram94, WKR09]. Reconstructing [Wei83]. Reconstruction [Sha93, Sto96, NCV10]. Record [Wei84, ACM69, ACM74, ACM76, ACM92a, ACM93a, ACM94a, ACM95a, VRC24]. Rectilinear [GK86]. Recursion [Bir77b, AP13, BFS00, CM90, CMW87]. Recursive [FR00, FKP77, JD89, MDLB02, Ode14, Dow93, GPR95b, HN90, Kra08, NTS93]. Recursive-based [Ode14]. Reduce [CKW09, SASU13, Har79, Kes79]. Reduced [TJD^{+}17]. reducibility [KR95]. Reduction [KNMH00, She59, DWE89, RP95, SN24]. redundancy [Joh94b]. Redundant [RJK79]. reexamined [ORT08]. Refactoring [WGMH13]. refactorization [Rém17]. Reference [Rob99a, Rob99b, Sch14, BK09, FPD08, GL03, Mha05, PKK18, SSLL21, Stu03, Stu07, ZZJC20]. References [AS91, Bar90, FS19, WL15a]. referencing [Lar98]. Refined [Pet94, Sch88b]. Refinement [ASJDW21, ASJDW18, Mor90]. reflection [HS08, Mor02]. Reflective [Dwe00]. refunctionalization [RTO15]. Regenerate [RT18]. Regex [Ano10, BDMT16, CFLH^{+}22, Kuk17, MNR^{+}23, RVV23, Sch13, THL^{+}20, CS09, Sch12, BTTF02]. Regex-based [BDMT16]. Regex-dependent [CFLH^{+}22]. regexes [MMI14]. RegeXeX [BH07]. Regexpcount [Nic03]. Region [Bao93]. regions [CM95]. Register [VSM87]. REGISTOR [PYY19]. Registration [DMWW77]. Regular [ARM^{+}19, AM91, ADR15, Ano68, Anoxx, Anoxx, Ant95, ACT10, Asp12, Bac94, BLR14, BTG83, BDD^{+}14, BR20, BC13b, Bee13, BF97, Ber00, BGNV10, BP11, Bra94, BC94, BFS04, BMMR19, BTC06, BK93b, BKW92d, BK93f, Brz62, BP63, Brz64b, Brz64a, Brz65, CDLV99, CDLV02, Cam99, CSY03, CZ01, Cha01, Cha02a, CLOZ04, COZ09, CJM12, CDJM15, CGR02, CJBW16, CBD^{+}23, CHP92, CT23, CC97, CKW09, CGS17, CDC^{+}23, CDL95, CDL99, CDPP23, Cox07, Cox09, Cox10a, Cox12, CS22, Dav99, Dav03, Dav04, Dav21, Dav22, DM11, EJ23, EU98, FRU^{+}20, FLS98, FU82, Fri02, FC04, GJ16, GRS99, GGM12, GN12, Ghi62, Gib21, GS22, Gil70, Gin67, Gol93, Goo05, GL12, GH13, GH15, HHM^{+}13, Hab04, HM98, Ham88, HWW06, Han13a, Han13b]. Regular [HWJ03, HN11, HJ99, Hir96, Hol84, HK11, Hos06, HVP00, HP01, HP03, HVP05, HN00, HSW97, HSW01, Hum97a, Hum97b, Hum99, IY02a, IY02b, JLK^{+}20, KT06, Kap69, KTU87, Kea91a, KP99b, KP99c, Kin92, KM92, KM95a, KLH16, KMRY20, KN12, KV15, KZ02, KST12, LS99, LS06, Lar98, LKM23, Lee09, LZHZ98, LM01b, LOZ^{+}24, LTV15, LT16, LN19, LT09, Loh10, Mad01, MS98, Mag81, MC24, MNS10, MY60, MPN^{+}14, MSZ17, MR09b, MPdS12, MGF97, Mye92, MOG98, NCJF18, NR99a, NR01, Nav01c, Nav04a, NR04, NR21, NH11, NM10, OES25, OS11, OWP16, Org03, OA98, OF61, ORT08, ORT09, Pak91, PM78, PPA10, Pat71, Pet02, Pik06, Pra97, Pre99, Ray96, Rez92, RD17, Ric79, Sab76, SA96, Sca11, Sch22, Sch99, SD95, SS93a, Sou99, Spe85a]. Regular [Spe85b, SM99, Stu03, Stu07, SvS14, SL14, SL17, SXZ^{+}25, TV14, TB00, Uma97, Urb23, VCS^{+}12, VB23, VHC88, Wag74, WPKL13, Wat96, Wen93, WMM95, WZU14, XK92, XLC19, Yam01, YPG21, YP12, YQW^{+}16, ZGS^{+}15, ZMWL20, ZCH23, Zia96, dLFM07, vNG01, AFI98, AI18, Agu23, Ano97a, AGM05, AM95, Ant96, AOMC07, ACM02, ADU16, BCG07, BYG96, BBG13, BAC12, BRL13, BdFED^{+}20, BG91a, BDFR08, BvdM17, BBvdM21, BvdM23, BS86, BNSV10, BFC08, BFG09, BG22, BBM21, BBRM23, BK86, Bra95, BH07, BKW92a, BKW92b, BK93c, BK93d, BKW92c, BK93e, BDM19, CS09, CGR03, CP97, CJR^{+}21, CJBW13, CX20, CHL^{+}25, Chi17, Cho78, CK02b, CLT07, CK08, CGPS13b, CGPS13a, CS11b, Cox10b, CDLM17, DL03, DF00, EZ74, EZ76, FL71, FDG^{+}11, Fil21, FHW10]. regular [Fos89, FS19, Fri97a, Fri06a, GLRÁ11, GR92, Gef03, Gel10, GL03, GS07, GHR14, GMS12, GM17, GH09, HW07, HWW07, HY90, HSJ04, Hoc19, HW09, Hov12, Hun79, Jan85, JSH09, Joh69, Kah06, KR22, KS08, Kar82, Ker07, KGA^{+}12, Kin91, KMR21, Lau00, Lau01, Lee82, LSO17, Lei80, Lei81, Lei85, LWS^{+}16, Lif03, LLL^{+}24, LR14, LM13, LMN16, LMS21, Lus94, MMDdJ11, McI04, MR05, Mor02, MZZ10, MPW21, MM89, NNSP22, NT20, Nic03, PHXD19, PC02, PIR17, PM23, PLT14, PIT^{+}03, PPPdG20, RT18, Rob79, Rom14, Ryt89, Sak21, SCF^{+}17, San15, SMS15, SGYM00, Sha88b, SY72, SH85, SM04, ST19, SSYW19, SLZ^{+}20, SMW^{+}23, SMT^{+}86, TN13, TN15, UIM22, Vou06, WXZY12, WL15b, WBS22, WW03, Wat03, WR15, XH23, XJT^{+}04, Yam19, YKGS11, YH91, YH92, YB13]. regular [YBTB23, ZZH16, ZC99, ZYX^{+}12, tC09, Tho68, Ano12]. Regular-Expression [BTC06, Han13a, YQW^{+}16, ORT08, ORT09, SCF^{+}17, WR15]. Regular-like [BTG83, MPW21]. regulatory [MMDdJ11]. reifiable [dSOMY15]. Reinforcement [KK02]. Related [CHZ06, AS85, Gro91b, Sri93]. Relation [KN12, MR92, Pre99, LSV08]. relational [BGHZ15, HC87, KWLL08, MZZ10, DWE89]. Relations [KAN^{+}17, BLSS03, BÖ13, LLS12, MAI^{+}16]. Relationship [XK92, GR92]. relaxation [SHCY93]. relaxed [BYB^{+}24, BYK22]. relevant [CDP16b, CD18]. Reliability [FO76]. Reliable [KKSL01, CDC96]. Remark [Tho81, Pet95, TCC91]. Remarks [CR87]. REmatch [RVV23]. Rematching [LYT^{+}23]. Remote [HMW^{+}23]. Removal [KK95, MGF97]. Renyi [AW89]. repair [PHXD19, Huc21]. Repairing [CT23, LWS^{+}16]. repeat [KW05]. repeated [LO94, Thi93]. repeating [UIM22]. repetitions [Cro86]. Repetitive [CDM11, Nav21b, Nav21a]. Repetitiveness [Nav21a]. replacement [NAR08]. Replacing [DCM15]. replication [HFFA09]. Report [GS81b, HJW^{+}92]. Reporting [MOG98]. representable [Dow93]. Representation [NR01, YBTB23]. Representations [KAN^{+}17, YB13, ZC89, ZZH16]. Representing [MNP^{+}23]. Required [MW92b]. requirement [LH13a, ZKCY07]. requires [Rob79]. rescuing [FSL^{+}15]. Research [AWD^{+}18, CPW88, IEE89, KL02, Med23, SVM14]. RESeED [SCF^{+}17]. Residue [BM00]. Resilient [ABBH^{+}16, RMK^{+}14]. Resistive [KYG19]. resolution [OW03]. resource [FK96]. resource-bounded [FK96]. resourceful [BFP^{+}08]. Resources [HAR10, SBM^{+}18, LG16, MP09, Roe25]. ResSeq [FSL^{+}15]. Restore [CMR18]. Restricted [Kin92, LOZ^{+}24, OWP16, Lei81]. Results [FMMS20, Lec95, WCM^{+}94b, FL13, Sid99, Sid02, WCM^{+}94a]. Retargetable [GFH82, BDB90, Gan89a, MR82, Fra83, GHF83a, GHF83b, WNL^{+}83].

Retargeting [MFMA15, MFMA17]. RETE [Alb89, MK90]. Retrieval [BBH^{+}87, DS19, FBY92, GR96, LZ96, MKF91, Mor83, Zve80, All82, BSY00, LMNT16, PMD01, SD91a, SD91b, ZKA12]. Reusability [PV91]. reusable [MME14]. Reuse [HL97, HR01, Rém17]. Reverse [RMK^{+}14]. Reverse-Engineering [RMK^{+}14]. Review [Ano97a, Ano97b, Ano12, Gue87, Hig95, Hum97a, Hum97b, Lab12a, Lab12b, Neu10, Pux97, Sal01, Tal81, Uma97, FL13]. revised [A^{+}08]. Revisited [Nag21, AG86, CCI^{+}13, GL01, RUG97]. rewrite [KN00, Ram94]. Rewriting [AM95, CDLV99, CDLV02, Dur94, EZYA23, GHW05, Lav91, ND02, GMC02, KR95, PSK17]. Rewriting-Based [ND02]. REX [Cam99]. Rexx [LS06]. Rey [ACM69]. Rhodopsin [NW13]. rich [ELF22]. Richardson [Sid95, Sid99, Sid00, Sid02]. Richness [QPWH08]. Richness-preserving [QPWH08]. RICS [TJD^{+}17]. RICS-DFA [TJD^{+}17]. Rider [MFMA15, MFMA17]. rifarensu [SM04]. Right [NWE97, Ned98, Pat71, CWZ10, HR03, Tak96b]. Right-Linear [Pat71]. right-to-left [CWZ10, HR03]. Rigid [JM85]. Rigid-Body [JM85]. Rigorous [Taf22]. RISC [TT22]. RISC-V [TT22]. Risk [Lut02]. Risk-Based [Lut02]. RNA [ABH^{+}14, BA15, IMR08, MP05, SBHM94, Shi00, Shi04]. RNAi [QLY07]. Road [KTY^{+}18]. Robert [Tal81]. robot [PS93a]. Robust [BH13, Le 91, LZ18, RMK^{+}14, TD18, WZJH12, YP13, BFN^{+}09, HLN09]. Roll [Tay97]. Roma [AAC^{+}01]. Roman [Sal01]. Root [CHZ06, TLLL07]. root-hashing [TLLL07]. Root-To-Frontier [CHZ06]. Rosenberg [CR91]. Rosser [KKM^{+}85, Sha88a]. Rotation [FU98, HW12, HLN09, TZYH14]. Rotations [FNU02, ABC^{+}04, AKT06]. Route [Les94]. router [MLC08]. Routine [Fre78]. routines [SMS15]. Routing [KBB01, LMMN07]. Rs [GW92]. Rs-operations [GW92]. RSMA [Ode14]. RSMA-FLFC [Ode14]. Ruby [Stu07]. Rule [Han92, PRT^{+}16, RT17, Gre88, Oph89]. rule-based [Gre88]. Rules [CvW18, Ghi62, Lav91, GMC02]. Ruleset [Sca11]. Run [Chu95, MHT09, BFKL13, BC95, RP95]. Run-Length [Chu95, BFKL13, BC95]. Run-Time [MHT09, RP95]. Running [DLG12, Gal79, BYR92, NAR08]. Runs [BL16]. runtime [Rob92, dH05]. Russians [Mye92]. Rust [Bal15, KN17, KN19, YC22].

S [Tal81, BGFK15a, BGFK15b]. safe [HS08]. safer [Rém17]. safety [FF08]. Salomaa [AFI98]. salute [FvGGM90]. sam [Pik87, Pik00]. Samples [GZ94, ST96b, ST96c, Tak94, Kin91]. Sampling [FMP20, GPR95a, Lut02, Vis91, WSW16, CGR93, Vis90, ZHWW12]. San [ACM92c, ACM92b, ACM93b, ACM95a, ACM95b, DT87, IEE94b, KP15, Sto92, USE92]. Sandeep [Hig95]. sanitizer [VS11]. SANTM [TJGY22]. SAR [B^{+}02]. Saratoga [Kap92]. SASL [LT90a]. Satellite [SS93a]. Savage [Sal01]. Savage-Smith [Sal01]. Saving [Bre93, Bre96, GS80]. SAXRegEx [YBTB23]. Scaffold [LJZZ13]. scalability [KNT15]. Scalable [ARS16, BAC12, BBHK14, BTC06, HSL10, JLK^{+}20, LQL^{+}16, LT16, MT14, RHR^{+}21, SRV^{+}19, VWR11, WGL^{+}21, WZL^{+}23, YP13, ZLN11, AB09, BGFK15a, BGFK15b, KTP10, PIR17, TLLL09, VW11, WL15b]. Scale [LP13, Roe25, LYWL08, TZYH14, WHZ^{+}17, ZWH^{+}21]. Scaled [BEL04]. Scaling [HW12, MS01, YDDB15, ALV92, HLN09, KKNS23]. Scan [MIH17, Gre88]. scanner [Hur84, ISHY88]. Scanners [HKR92]. Scanning [AKW85, PWW^{+}11, CWZ10, HFI^{+}08]. SCCs [ZYQ^{+}15]. Scenes [BZ98, BSM^{+}07]. scheduling [LMMN07, Mid98]. schema [IIK08, MNS07, dLFM07]. Schemas [BGNV10, MNS10, KS07]. Scheme [Bur84, FK16, JDXD13, Man94, Man97, Bur82, Kod79, KRL87, LH13a]. Schemes [KK08, RDRF24, Pel87, QWX^{+}13]. School [Cro92a, Ano92b]. Science [ACM89, FJ92, Gus97, IEE89, IEE90, IEE92, IEE93, IEE95a, IEE97, IEE98, IEE09, Ker04, Win78]. Scientific [WCM^{+}94b, ORPF13, WCM^{+}94a, WZS95]. Scolopax [ORPF13]. Score [Ben94]. Scores [CLST^{+}13]. Scoring [KK08, OSSK16]. Scotland [AOV^{+}99]. Scottsdale [KLB12]. Scratchpad [JT94]. screening [QPWH08]. Scrimshaw [Arn93a, Arn93b]. Scripting [Fri97b, RB05, BFN^{+}09, BH13, Blu08, FhDAF09, Han01, LS99]. scsh [Sar02]. SDN [AV23]. SDN-Based [AV23]. Search [AC75, HR02, BYKZ^{+}92, Ber00, BK93g, Cal00, Cha91, EF13, FG98, FL99, GN19, GG97, KR94, Lut02, Man86, MM96, NR99a, NR01, Pol01, RDRF24, Rob88, Smi91b, SED14, SOR16, SJNS19, SB09, Sun90a, Sun90b, Tay97, VB23, WWW^{+}16, WT89, YDDB15, Zha07, ZGS^{+}15, AB09, AVN22, Bar84b, BC13a, BPMA02, BC06, FG95b, FG95a, FG99, Git96, GHK14, HH16, JRV96, KW05, KJS17, KM84, LGZ^{+}14, MRR^{+}18, MKSiA98, NR02, OSSK16, QQC^{+}13a, QQC^{+}13b, QLY07, Rai99, SCF^{+}17, TYNM86, Tan14, Tho68, WDG^{+}14, WT88, YHV^{+}15, YBTB23, ZBST14, dKM04, Cox12]. Search-Based [VB23]. Search-Space [ZGS^{+}15]. Searches [DNM00, GN01, MM93, MT14, Nel96, Shi96, Fen01a, KS08, KBN09, MW92b, Men89, SMS15, WR15]. Searching [BY89, BK93a, BPMA02, BS97, BM77, Bur11, CFG12, CC97, CCL87, CR95c, Dav82, Dav73, EF95, Gon83, HL94, Hor80, HS90, HS91, Knu98, LDI98, Man94, Man97, MNS84, Nav01c, NR03, Nav04a, NR04, PWW^{+}11, PW12, Rai92, Reg92, Roo99, Ros95, Smi94, Ste94, TT82, VSP08, ZMSD93, AG86, Apo93b, AEMS14, Ayc15, BYG92, BYG96, BD80, Bar84a, CGT25, CD96, CEMW91, FGG^{+}08, GO80, HKN14, Han93, IA80, KIH15, Mha05, Mus03, Mus05, Owo93, Per94, Ryt80, SNZBY00, WL15a, WM92a]. Seattle [ACM74, ACM98]. Sebastopol [Ano97a]. Second [ACM83, ACM00, Ano12, ACM90b, Dow91, TPT13]. second-order [Dow91, TPT13]. secondary [BA15, MP05, AVN22]. Sections [DCM15]. Secure [BEM^{+}12, BEM^{+}13, EF13, HT14, SBB19, SJNS19, DA18, SCF^{+}17]. Security [HL10, LN19, MW94, PP06, Rei03, SALP20]. sed [Dou91, Bar90]. seed [CCI^{+}13]. Seeded [LPR^{+}08]. Seeds [Zha07, FCLST07]. SEFT [dKM04]. Segmentation [HMW^{+}23, IIK08]. Segments [EIV04]. Seiki [SM04]. Selected [Cro92a, Moo64, PRT^{+}16, BW91]. Selection [Bon07, CMR18, Gie90, LH13b]. Selections [CvW18]. selectivity [JKNS00, STKD20, TP07a, TP07b]. Self [CG87, TJGY22, AGH^{+}17, CADA18, JP11]. Self-attention-driven [TJGY22]. self-composition [AGH^{+}17]. self-interpretation [JP11]. Self-Organizing [CG87]. self-service [CADA18]. SeLINQ [SHS14]. Semantic [CBD^{+}23, Gan89b, Har97, HMW^{+}23, II08, Coo86, IIK08, MP09, OA17, SG12, AGP18]. Semantica [Har97]. semantically [ELF22]. Semantics [Dan91, Gud92, MNR^{+}23, Pag78, PAG09, BvdM17, Chl08, Mal93, San15]. Semantics-directed [Dan91]. Semi [KV15, LYT^{+}23, TMK^{+}02, BGHZ15, Rob79]. Semi-Extended [KV15]. Semi-structured [TMK^{+}02, BGHZ15]. Semi-Supervised [LYT^{+}23]. semiring [AS85]. Semirings [GHKL18]. semistructured [BFS00]. Sensemaking [LLS^{+}20]. Sensing [HMW^{+}23]. Sensitive [CK02a, SA96, TPT13]. sensor [NCV10]. sensors [BWG12]. Sentiment [THQ19]. Separated [Pol13]. Separating [LMM17]. Separation [Kul11]. September [AAC^{+}01, AOV^{+}99, Len93, MG94, Win78]. Seq [SNB^{+}19]. Sequence [BLP94, BGM19, BDFW94, Bra90, CCL87, GM02, KPP19, KS99, KK08, LPT12, MGW14, NW13, RND97, CLT07, ECSS88, ETV21, GO12, HAI02, MBY91, NT20, SVS97]. sequenced [GW92]. Sequences [BLP18, BEL17, CvW18, Gus97, Hua94, IMR08, LJH^{+}17, MT14, NEL17, DKP11, GGN06, HIEH22, IHEH23, KB22, KR14, LVN87, MZZ10, NR02, Nil90, RLP20, Sel84, SN94, Sid95, Sid00, Sid02, ZKCY07, Lut02]. Sequencing [RM06, WGL^{+}21, FSL^{+}15, KRML09]. sequentiability [Gaá04]. Sequential [BP63, Brz64a, Brz65, Dur94, GRS99, GPR95a, Liu14, Man86, Moh94, MR13, Moo64, OF61, PM78, RS98, Wei83, Zve80, JM90, MMZ10, TTO^{+}22]. SEQUITUR [MHM^{+}01]. Serial [JHU^{+}24, WSVS22, GS93b, GS93c, LV86b, LV89]. Series [LKL02, MR09b, YBTB23]. Series-Parallel [MR09b]. Server [GSL17, SWZ01]. Server-Side [GSL17]. service [CADA18, PM23, ZBST14]. Services [KPP19]. SeSG [AVN22]. Set [HL10, Hig86, Hui92, KF91, CGK08, CP10, KR97, THL^{+}20, WC14, dRL95, TJD^{+}17]. SETH [EMT23]. Sets [AGM19, BNV^{+}13, EIV04, JN24, Pol13, Prů17, Via02, AK09a, Alb89, CS98, Che96, Cho78, Kin91, LMM17, Mat94, Rot91, Sta89, TZYH14]. Setup [SOR16]. Seventeenth [ACM89]. Seventh [LL08, ACM95c, AAC^{+}01]. SGML [BK93b, BKW92d, BK93c, BK93d, BK93e, MGH93, MGH97]. SHA [KJS17]. SHA-1 [KJS17]. shading [BSM^{+}07]. Shallow [Cam99]. Shankar [Pux97]. Shape [BZ98, CTF^{+}98, YJ84, DOS93]. Shaped [HMW^{+}23]. Shapes [HH93a, CT96, HH93b, HH93c]. Shared [TVCM12, VRD01, SW93]. Shared-Forest [VRD01]. sheaf [Sri93]. sheaf-theoretic [Sri93]. shelf [MNS07]. Shell [IEE01e, IEE01f, RB05, Blu08, Wes97]. Shells [Qui02, Qui00]. Shift [Fre03, IMP01, KTSA99, CT96, Per94]. Shift-And [KTSA99]. Shift-or [Fre03, Per94]. shifting [WOQ^{+}07]. Short [Ayc15, BLPL92, BAM^{+}24, Han13b, KRS95, KRS97, MR09b, XPZ^{+}23, BGFK15a, BGFK15b, Che96, FSL^{+}15]. short-read [FSL^{+}15]. Shorter [GH13, HW07, ZZH10]. Shortest [AKT20, Ale94, ATX21, Mid98, TY97, GHST17, HTX17, IKX15, TU88, Yam19]. Shotgun [KS99]. shuffle [Jan85, Jed87, ST19]. SIAM [ACM97b]. Side [GSL17, Hum99]. Sided [CDJM15]. SIG [ACM97a]. SIG-SIGMOD-SIGART [ACM97a]. SIGACT [ACM83, ACM87, ACM90a, ACM92a, ACM92b, ACM93a, ACM94a, ACM94c, ACM95a, ACM95b, ACM98, ACM99a, ACM06, ACM07, KLB12, LL08, Len11]. SIGACT-SIGMOD [ACM83]. SIGACT-SIGMOD-SIGART [ACM90a, ACM92b, ACM94c, ACM95b]. SIGACT-SIGPLAN [ACM87]. SIGAPP [DGBH93]. SIGART [ACM90a, ACM92b, ACM94c, ACM95b, ACM97a, ACM98, ACM99a, ACM06, ACM07, KLB12, LL08, Len11]. SigMatch [KTP10]. SIGMOD [ACM83, ACM90a, ACM92b, ACM94c, ACM95b, ACM97a, ACM98, ACM99a, ACM06, ACM07, FMA02, HF13, KLB12, LL08, Len11, SW94, Sto92, HF13]. SIGMOD-SIGACT-SIGART [ACM99a, ACM06, ACM07, KLB12, LL08, Len11]. SIGMOD/PODS [HF13]. SIGMOD/PODS’13 [HF13]. Sign [BTTF02]. Sign-and-Encrypt [BTTF02]. signal [Nak14, NCV10]. signature [QWX^{+}13, TJD^{+}17]. Signatures [OWP16, TT82, CEMW91, WXZY12]. Significant [PRU11]. SIGPLAN [ACM87, ACM92a, ACM93a, ACM94a, ACM95a]. SIGPLAN-SIGACT [ACM92a, ACM93a, ACM94a, ACM95a]. SIMD [CFKT17, FPT22, JA17, JXA20]. Similar [BEL17, IK83, NEL17, HKN14]. Similar-String [IK83]. Similarity [CM94, MU02, AK09a, BSY00, IIK08, II08, JLFL14, QWX^{+}13, TP07a, TP07b, WDG^{+}14, WLF14, ZXL^{+}13]. Simon [Han93, KKH24]. Simple [AR00, Bra94, BGM13, Cox07, Gaw12, GL89, IST05, LOZ^{+}24, MNS07, Mut00, NR03, Ris16, BBG13, CLS95, Eck89, Fla88, FHP92, IKX15, II86, NBY99c, Pet07]. Simpler [CH02, KC21]. Simplicity [Wol90, Edw07]. simplification [OA98]. Simplified [FYJ^{+}17]. Simplifying [KR22]. Simulation [AYS^{+}24, BÖ13, Coo72, FYJ^{+}17, KL02, FS22, KCPC13, MCF^{+}14, Pet94]. Simulation-Based [KL02]. Simulations [Gal76b]. Simultaneous [PS93a]. Single [AAB^{+}86, Per94, SK17, WHZ^{+}17, EASK14, FN04, NHN^{+}20]. single-arm-gapped [NHN^{+}20]. Single-board [AAB^{+}86]. Single-machine [WHZ^{+}17]. Single-node [SK17]. singleton [AFI98]. singly [BAP06]. singularity [WYA^{+}07]. Situ [GMAS22]. Sixteenth [ACM97a, ACM84]. Sixth [ACM07, ACM74, ACM94d]. Size [BGG12, CF85, CKW09, Hui92, Kül10, BYBDS09, GH09, Lif03, PPTT15]. Sized [KRR17]. sketch [LLS^{+}20]. sketching [SLTB^{+}06]. skills [DV21]. skip [Hua98]. Skolem [Kar82]. Sleeping [ABF96]. sliced [KRL87]. Slicing [DSv94]. Sliding [FL12b]. Small [CLP98, CGPR95, HSW97, HSW01, MPN^{+}14, NTS93, STK10, Sca11]. Small-Ruleset [Sca11]. Smaller [GPR95a, GLRÁ11]. Smallest [Man75]. smart [JR15a]. Smith [Sal01]. smoothed [AK12]. Smoothness [ZCS^{+}12].

Smurf [CADA18]. SNOBOL [Bro77]. SNOBOL4 [Duf82, Gim73, Gri85, Liu86, Sch81, Sil77, Gri83, Pag78]. SNOBOL4/Icon [Gri85]. Snowbird [SC93, SC95, SC96, SC98, SC99, SC01, SC02, SC03, SC04, SM09, SM10, SM11]. Snuba [VR18]. Social [MSP^{+}17, VD17]. Society [CVP86, NEH90]. SODA [BJK^{+}12]. Soft [CKC07, How97, How96]. Software [Bee81, CS18, FO76, IEE94a, IEE95b, KP99c, KPA10, Lut02, MG94, PAMP12, Cox19, Gre88, Joh01, Jon13, KOI94, MP88, PA10, San09, Spe85a, Spe85b, Bol02]. Solaris [Rob99a, Rob99b]. Solution [Hea71, B^{+}05, BLPL92, Goo05, Sch81, YCJK08]. Solutions [Gon02, Sto96, AI18, BBHK14, DL03, Sta89, BGG^{+}94]. solved [SSLL21]. solver [KGA^{+}12]. Solvers [ZGS^{+}15]. Solving [BK93g, CFLH^{+}22, CCH^{+}23, FR17, SED14, Tar81a, CHL^{+}25, Tak96a]. Some [Fen01b, Gal76b, Liu86, MW92b, Sli78, WCM^{+}94b, Wol90, San15, WCM^{+}94a]. Sorted [Wal88, Gie90, Kes91]. Sorting [BS97, CMR18, Chr96, JRV96, Knu98, FCFM00]. sorting-complexity [FCFM00]. Sortledton [FMG22, FMG23]. Sound [PM23]. Source [AWD^{+}18, Bol02, EZYA23, MFMA17, SED14, AG06, Joh94b]. Source-Level [MFMA17]. sources [ST96a]. Sourcing [CDL^{+}15]. South [ACM93a]. Space [BC13b, Ben94, CF06, CZ01, Cha94, CF85, CDEK95, CJPS13, CR95c, CGPR95, GS80, GHST17, GPR95a, GP01, GP03, KC87, LT09, McC76, SY23, SWY75, ZGS^{+}15, AK08, BCWG09, BGM13, CD96, CGR99, GS81b, GS83, GPR95b, GO12, IKX15, KR89, LMRT14, PLT14, Rob79, SW12, SNM^{+}13, TJD^{+}17]. Space-Economical [McC76]. Space-Efficient [BC13b, SY23]. space-optimal [KR89]. Space-Time [CF85, GHST17]. Spaced [Zha07, FCLST07, NCV10]. Spafford [Hig95]. Spain [DMVT13, LV06, NH11]. Spam [ZGY^{+}16, KEG^{+}08]. Spanners [ABMN20, FKRV13, FKRV15, FRU^{+}20]. SPARQL [LM12, LM13, PAG09]. Sparse [WSW16, HSL10, Quo92]. spatial [CS98, CFCK22, FLC^{+}19]. spatio [PMD01]. spatio-temporal [PMD01]. Speaker [PG90]. SPEC [KL02]. Special [ALLL98a, AK09b, AGS93a, AGS93b, AGS93d, AGS93c, Ano17, DT87, ALLL98b, Cro92a]. specialisation [Jon07]. specialization [ADM^{+}13]. specialized [GÁSÁ^{+}13]. Specific [CDC^{+}23, MFMA15, MFMA17, MGW14, MSS^{+}19, TMV^{+}01, SHvR^{+}16, WKR09]. Specification [BG91a, Lut02, Sou99, SMT^{+}86, MRA^{+}17]. Specifications [NCJF18, JM90]. Specified [ZMSD93]. specify [CFM00]. Specifying [Lus94, Lut02]. Spectra [BM08, SHCY93]. spectrum [ZHWW12]. Speculation [JA17, JXA20, SXZ^{+}25]. Speculative [NYuR15, RP95]. Speech [JLH18, Kul11, PG90, RJK79, AAB^{+}86]. Speed [BYHT18, FL12b, JGZL12, JN24, LK90, PWSG25, PW12, VCS^{+}12, ZL18, Git96, KM84, LK88, PLT14, TLLL07, XMLC11, ZYX^{+}12]. Speed-Ups [JN24, PWSG25]. Speeding [CCG^{+}94, Deo06, SKF^{+}00, ACF05]. spelling [AB89, BSY00, TIAY90]. spelling-correction [BSY00]. SPIRIT [GRS99]. Split [KKK11, TBS06]. Splitting [RTT02b]. Sprachen [HU92].

Spreadsheet [GHS12, SG16]. spreadsheets [BGHZ15]. Spring [Ano87]. Springer [Neu10]. Springs [Kap92]. SQL [AV23, BJK^{+}12, FPD08]. Square [ACM83, CIK98]. Squares [CR95c, Rao94, IMS97]. Squib [SM99]. SSD [PYY19]. St [IEE90, TSM88]. stability [Sid99]. Stabilization [CCH^{+}23]. stable [KT90]. Stack [ZZ12]. Stack-based [ZZ12]. STACS [FJ92]. Stage [HMW^{+}23, JGMP22, YCJK08]. staged [PSK17]. Standard [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, XLC19, BR09]. STAP [dLBHC22]. Star [BMMR19, HY90, Tho81, YH91, YH92, tC09]. star-free [tC09]. star-height-problem [Tho81]. starting [Mid98]. State [Bow87, BDM19, CZOdlH17, CM58, Gol93, Han13b, JA17, KLH16, MY60, NRS18, Sut21, WDG^{+}14, Yun12, vNG01, Gaá04, HW07, Hur84, JXA20, MMS14, VHL^{+}12, Yod91]. State-of-the-art [WDG^{+}14]. Stateful [VKPI17]. Statements [JP73]. States [DGBH93, LK06]. Static [Cha02b, HV93, JGZL12, WHZ^{+}17, ZWH^{+}21, ALLS07, Aoe89, FhDAF09, GLS07, HS08, JU91, LYWL08, MP09, PIR17, PM23]. Stationary [KS96, ST96a]. Statistical [BGJ01, GS93a, GWvG10, THQ19]. Statistics [Liu14, Jan23, Maa06, ZZJC20]. Std [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d]. Steady [Sut21]. Steady-State [Sut21]. stem [YHV^{+}15]. step [BD98]. Stereovision [PDL98]. Steven [Ano12]. Stieltjes [KC11]. Still [Gon02, LS06]. STOC [ACM08]. Stochastic [SBHM94, YPG21]. Stone [D^{+}23]. Stopper [RTT02b, RTT02a]. Storage [BK75, JLK^{+}20, PYY19, All82, CDC96, GS81a, SCF^{+}17]. Strategic [Vis99]. Strategies [CJ93, MM02, AVN22, AG86, HBRV10, MM03, MLM^{+}08, PSK08, PCS99, iA94]. Strategy [Bon07, EMC96, AEH94, LLL13]. Stream [PP06, Nil90]. Streamed [DCM15]. Streaming [BG14, GKP19, PP09, BGFK15a, BGFK15b, GHR^{+}16, MRA^{+}17]. StreamQRE [MRA^{+}17]. Streams [CJPS12, DLG12, Har02, CL09, CGM10, San15, SGCW14, TSI13]. strength [Moo12]. strict [HJW^{+}92]. Stride [PW12, VWR11, NYuR15]. stride- [NYuR15]. striding [ARS16]. String [AOK02, Abr87, AC75, Aku94, AR00, ACR01, ADR15, AYS84, iA94, ACD01, BST^{+}03, BY89, BYP92, BYN96, BY96a, BYN97, BYN99, BCP02, Bar84b, Bee13, BK93a, BEL17, BH02, BH85, BGP^{+}22, Ber00, BLLP90, BL94, BM00, BGVW12, BGS23, BM77, BG92, Bre93, BCT94, BG95, BCT98, BGG12, BG14, BK93g, BZ98, Bur84, CZOdlH17, CF06, CFP19, Car77, CF88, CK02a, CLS^{+}10, CL92, Cha93b, CM94, CL94, CCH09, Cha91, CL97, CLP98, CN02, CTF^{+}98, CHL14, CCH^{+}23, CGT25, CH04, Chu95, CW84, Col94a, CHPZ95, CH97a, CH02, CP91, Cro92b, CCG^{+}94, CR95c, CGG^{+}97, CIK98, CIM^{+}02, Dav82, Dav73, DW17, DNM00, EMC96, EMTG23, EZYA23, FT95, FT98, FL12a, FL12b, FMP20, FR17, FG98, FL99, FV16, FPT22, Fra20, FU98]. String [Fre02, FT04, Fre06, Fre78, Gal76b, Gal79, GS80, Gal81, GP90, GG91, GG92, Gal95, GPR95a, GHW05, GZ94, Gon02, GFG11, GV05, GMMN12, GH82, HD80, HL94, HH93a, Hoa77, Hui92, HS90, HS91, HN02, HN05, IMP01, IK83, JL96, JLFL14, JN24, JTU96, Kha16, KST94, KKK11, KS11b, KS12b, KTY^{+}18, KMM15, LSW08, LP13, Le 91, Lec95, Lec98, Les94, LY86, LLLL08, LLLC17, LD10, Liu86, Liu88, LDI98, LCL06, LLW^{+}15, LS94, Man75, MM93, Mel95, Mey85, MM02, MIH17, Moh97, MNS84, ML96a, ML96b, Mun07, MM96, MR92, Mut97, Mye98, Nao91, NW13, NR98, NBY99b, Nav21b, Nav21a, Nel96, NR15, NEL17, Ode14, OM88, PAMP12, PLL08, PWW^{+}11, PW12, PK95, PP94, Pet07, Phi94, PRT^{+}16, Rai92, Rao94, RTT02a, RTT02b, Reg92, RKH02, RPE81]. String [Roe25, Ros95, RFD23, RNOM09, Sad96, SV94, STK10, SD95, Shi96, Shi92, Shi97, Sim94, Sli78, Sli83, Smi94, Spi99b, Ste94, SJNS19, ST95, ST96b, ST96c, ST04, Sut21, TS05, TU93, TP97, TT82, TLC15, TVCM12, UW93, VMML15, VSP08, WT89, Wri94, YP13, YDW18, ZS17, ZS13, ZGS^{+}15, ZMWL20, de 82, van14a, van14b, Aku95, ASM17, ACF05, AVN22, ALP04, AAK^{+}09, AEK^{+}11, Aoe89, AG86, AEMS14, AGW13, Ayc15, BFKL13, BYF96, BYP96, BD80, BSY00, Bak78, Bar84a, BR09, BKBB^{+}14, BLPL92, BFG09, BFP^{+}08, BG90, BG91b, BCT93, Bre94b, Bre95, Bre96, BGM13, Bur82, BEL04, CADA18, CCF13, CL90, Cha93a, CDDM05, CW13, CW18a, CW18b, CN21, CFLH^{+}22, CR87, CH92, CGG90, CD96, CM07, CGR99, Dai09, DR06, Deo06, Der95, DC94, DNR06, ĎHPT10]. string [EMT23, FCLST07, FL13, FS22, Fen01a, FG95a, FMdB99, FG99, FGG^{+}08, FBMA05, Fre03, FN04, FM06, Gal75, Gal76a, GS81a, GS81b, GS83, Gal84, GG86, GG87, Gal92, GPR95b, GGL94, Git96, GBY90a, GBY90b, GF08, GL89, GV00, GO80, GHK14, Han93, HY92, HFS05, HR01, HR03, HH93b, HH93c, HOK18a, HOK18b, HM00, HLS^{+}11, HK77, HHLS06, HFN05, Hyy08, IP96, IMS97, Ind98, IS90, II08, IA80, JL93, Joh95, JU91, KB22, KST92, Kim99, KWL07, KC21, KNT11, KS96, KST16, KPA10, LV86a, LV86b, LVN87, LV88, LV89, Lar99, Lec07, Liu81, LHCK04, LT97, LLL13, Mae90, MNU05, MBY91, ME97, Men89, Mha05, MM03, MM07, Mis03, MS95, Mus03, Mus05, Mye99, Nak14, NHN^{+}20, NBY99c, NR00, NKT^{+}01, Nav01a, NF04, NT05, NC06]. string [Neb06, NC92, PLL10, PA10, PKK18, RUG97, Ryt80, RLP20, Sad93, SY23, STK06, Sal12, SW90, SZ05, SMS15, ST96a, STKD20, SG12, SR16, Spi99a, SV87, TYNM86, Tak96b, TBS06, THG17, TPT13, Ukk92, Ukk93, VRC24, VLP17, Vin77a, Vin77b, WDG^{+}14, WL15a, WLF14, WT88, XMLC11, Yao79, YT03, ZZJC20, dB93]. String-Based [EZYA23]. String-Manipulating [VMML15]. String-Matching [BG14, CCG^{+}94, GS80, Gal95, JL96, Kha16, Les94, LY86, Moh97, Mut97, Sli78, Sli83, CH04, Cro92b, BR09, CCF13, CW13, CR87, CGR99, DR06, Gal75, Gal76a, GS81a, Gal92, GPR95b, HY92, HR01, HR03, JL93, KST92, LHCK04, PLL10, TBS06, Ukk92, Ukk93, dB93]. string-pattern [Kim99]. string-searching [Mha05, Ryt80]. string-similarity [BSY00]. String-to-Dictionary [KS11b, KS12b]. string-to-string [Mae90]. stringdist [van14a, van14b]. Stringlish [Ayc15]. Strings [Ale94, BS97, BCFL12, Chu95, Col94b, FT98, Gaw13, GNU94, GL01, Gus97, HUN^{+}19, HT17, Hor80, Huc21, ISNH94, KRS95, KRS97, KAN^{+}17, KMP77, LT03, Lut02, Shi96, SW09, Ver92, YQW^{+}16, Zha17, ADR03, ADR06, BLSS03, BFK^{+}03, BC95, CD89, CR91, DGG^{+}19, EH88, ETV88, FT95, GO12, JRV96, Jan23, KGA^{+}12, KMP94, KR97, LMM17, LS10, LMS21, McI04, Mei15, NR02]. Strong [BMMR19, GGM12, LS06, MCF^{+}14, WD99, AW89]. Strongly [Dur94]. Strother [Tal81]. Structural [BGJ01, KWLL08, Pik06, Shi00, Shi04, BFS00]. Structure [CGR02, FMG23, Gia93, Les95, Pol13, Sli78, TMV^{+}01, AP90, CGR03, CD96, FG95b, FG99, FLSS93a, FLSS93b, FMG22, KWL07, MP05]. Structured [BLLW12, CMNP17, KM94, KS99, BGHZ15, Fla88, TMK^{+}02]. Structures [Cha01, Cha02a, FBY92, GHLW15, GG97, Gor00, GKW^{+}10, LSW08, Lar99, Lec98, Les79, APTS13, ABH^{+}14, BA15, GMC02, HN90, HTK^{+}21]. stuck [AEK^{+}11]. Students [DKA^{+}15]. Studien [SM74]. Studies [JM85, SM56, SM74, SS93a, AVN22]. Study [CSY03, FTJ95, JM85, KP96a, MM02, MSZ17, OP16, PV91, Sca11, BG91a, Fen01b, KP96b, PSK08, Pep91, SKS96]. Studying [MGH93]. Sturmian [BR09]. Style [Cop91, BGWXP22, WW03]. sub [EMT23, VRC24]. sub-quadratic [EMT23]. sub-string [VRC24]. subexpressions [Fat15]. Subgraph [QZC17, XZL^{+}19, EASK14, KNT15, KSH^{+}15, LQL^{+}16, SWW^{+}12]. Subgraphs [MSS^{+}19, PSP^{+}18]. subject [ETV88, Sch81]. subkeys [BD80]. Sublinear [CL94, FG98, CL90, CWZ10, CGR99, FG95a, WZ96]. sublinearity [Sch88a]. Sublist [Jay92]. Submatch [HHM^{+}13, BT21, Lau01]. Submatching [SvS14]. Suboptimal [Cha94, LS94]. Subquadratic [WMM95]. subscribe [ZCT14]. Subsequence [BGM19, ETV21, HIRS17, TTO^{+}22, ZKA12]. Subsequences [IF94]. Subset [CH03, Kin92, Pag78, AB09, CH97b, HW09]. subshifts [Rou21]. Substitution [For02, JSC83, Sch81]. substitutions [LVN87, Pie08]. Substring [AKT20, ACR20, HR02, BYKZ^{+}92, Bur11, CIL^{+}03, F^{+}23, Har71, JN24, Joh94a, KO83, KW05, KRS19a, KRS19b, KRS23, MRR^{+}18, Rob88, RMK^{+}14, Smi91b, SOR16, Sun90a, Sun90b, Apo93b, BGK^{+}16, BSTU08, ETV21, FG95b, Gra15, HKN14, HTX17, IKX15, JKNS00, Maa06, MAI^{+}16, MKSiA98, Rai99, Rou21, Sto02, TTO^{+}22, Yam19]. substring-preprocessing [Sto02]. Substrings [ATX21, Cob94, Fra20, Boo80, FGKU15, GHST17, LO94]. subtext [BD80]. subtree [Gro91a, Gro91b, Mäk89]. Subtype [WZJH12]. subtypes [JM93]. Succinctness [Gel10, GN12]. sufficient [KT90, MR09a]. Suffix [AOK02, ABM08, DNM00, FL12a, GV05, GLS92, Kid09, LSW08, MM93, McC76, NR98, Nel96, Neu10, OR12, PRT^{+}16, Shi96, Shi00, Shi04, UW93, ACFC^{+}16, BH96, DK13, FCFM00, FS22, GV00, HHLS06, Kos94, NR00, TTHP05, Ukk93]. Suffix-Tree [DNM00]. suffixes [BGK^{+}16]. Suitable [CCL87]. Summary [GH15]. Sums [BM00]. Sup [MP09]. Sup-interpretations [MP09]. Super [Fre02, KM95b, Fre03]. Super-Alphabets [Fre02, Fre03]. Super-Pattern [KM95b]. Supercomputers [RND97]. Supercomputing [IEE88]. superimposed [Ind97]. Superiority [Zha07]. superoptimizers [BA06]. superprimitivity [Bre94b]. Supersequences [IF94]. Superstrings [Ale94, TY97, Che96, Mid98, TU88]. Supervised [LYT^{+}23]. supervision [VR18]. supplement [Ruc15]. Support [GSL17, CL09, GZ10a, Kau92, KAT07, Rob92, ZBST14]. Supporting [SOR16, CMW87]. supports [Nil90, WR15]. surface [TCCK90]. Survey [AKT20, BYNTK21, Brz62, Kni89, HIEH22, LH13b]. Surveyor [Fra83, GHF83a, GHF83b, WNL^{+}83]. Surviving [Cox19]. SVR4 [Rob99a, Rob99b]. Swap [AEP06]. Swaps [ALLL98a, AAL^{+}00, CCFG12, AAL^{+}97a, ALLL98b, Mei15]. SWAR [CL09]. sweep [JHU^{+}24]. Switches [WZL^{+}23, LLL^{+}24]. Switching [Hoa77]. symbol [Rai99]. Symbolic [ACM94b, Bro93, Cha86, GVD15, Har79, Lev95, Ng79, VHL^{+}12, WN90, Fat15, Nic03, NA90, YBTB23, Ng79, NEH90]. Symmetric [Gil70, SS93a]. Symmetries [Hig86]. symmetry [Mar89]. Symposium [ACM69, ACM74, ACM76, ACM81, ACM83, ACM84, ACM86, ACM87, ACM90a, ACM90b, ACM91, ACM92a, ACM92b, ACM92d, ACM93a, ACM93b, ACM94a, ACM94b, ACM94c, ACM94d, ACM95a, ACM95b, ACM95c, ACM97b, ACM97a, ACM97c, ACM98, ACM99a, ACM99b, ACM00, ACM06, ACM07, ACM08, AP10, AH97, AT02, Bro93, Cha86, DGBH93, FC98, FL08, FJ92, GM11, HM96, HF13, Hwa85, IEE89, IEE90, IEE92, IEE93, IEE95a, IEE97, IEE98, IEE09, KS12a, KU09, Len93, LL08, Lev95, LV06, MZ07, Ng79, WN90, Win78, AL01, Apo92, Apo93a, ACP05, BYCC03, CG94b, GU95, GS00, KLB12, Len11, PC99, SMD04]. Symsac [Cha86]. synchronization [JM90]. Synchronized [PIT^{+}03, XH23]. Synchronizing [JN24]. synchronous [Pie95]. Synonyms [LLW^{+}15]. Syntactic [KKSL01, TB00, Wol86]. Syntax [BLS^{+}94, PWSG25, XLC19, AG06, Chl08, Pie08, ZGE85, Zei08]. Synthesis [AI18, BDD^{+}14, CBD^{+}23, CDL^{+}15, BK86, Lei81, SW12, WKR09, ZJL14]. synthesized [Kod79]. Synthesizing [LSO17, PJP^{+}18, SDA17]. System [BM00, CFS^{+}89, DMWW77, Har79, IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, KSWC93, KMT^{+}01, KKSL01, MM02, NCJF18, SF01, Som82, WHZ^{+}17, Wol86, ZJP^{+}18, AAB^{+}86, BAC12, BG91a, BH07, GPTV93, JGMP22, KAT07, KMS^{+}03, KJS17, KLR^{+}08, KPP21, LHCK04, MI07, Rus92, TIAY90, WCW82]. Systematic [KK95, NAR08]. Systemizing [ZWH^{+}21]. Systems [ACM83, ACM90a, ACM92b, ACM94c, ACM95b, ACM97a, ACM98, ACM99a, ACM06, ACM07, Ano68, Dur94, FYJ^{+}17, GHW05, HF13, IST05, JM85, KKK11, Kor83, LHMH91, LY17, LLS^{+}20, LL08, LZ96, Lut02, MUHT96, Mor83, PSP^{+}18, Sar02, WHZ^{+}17, ZWH^{+}21, ADM^{+}13, CDC96, CFM00, DL03, Fat15, JO97, KKM^{+}85, KLL23, KN00, KKP92, KEG^{+}08, KLB12, KR95, LLC03, Len11, Lus94, SD91a, SD91b, WSS94]. SystemT [KLR^{+}08]. systolic [PS88].

t [KPP21]. table [GHS82]. tables [EF95, Hua98, Mus05, Quo92, SDA17]. tagged [Lau00]. Tagging [JLH18, Kul11, KEG^{+}08]. Take [Roo99]. Tale [Hum88a, VLP17]. talk [Rém17]. Taming [CGZ^{+}13, Hab04, KSH^{+}15]. tapes [Cho78]. target [QLY07]. targeting [All89]. Tarragona [NH11]. TASH [Wes97]. Task [YD95]. TAWK [Eck89]. taxonomy [CWZ10, WZ96]. TBNF [Man06]. TCAM [MPN^{+}14, PD12, Yun12]. TCAM-Based [Yun12, PD12]. TCAMs [dLBHC22]. Tcl [Wes97]. Teaching [GOMSJVGP08, Far19]. Technical [BYKZ^{+}92, Spi99b]. Technique [Vis91, ZT89, Bak78, CK02b, Fla88, IHEH23, PC02, Vis90]. Techniques [DCM15, GS93a, GL86, HH93a, Kuk92, Mu 95, MuT95, NR04, Tho68, Ano97a, DOS93, EF95, Fri97a, HH93b, HH93c, MSRR00, Mun95, PPPdG20]. technologies [OKT92]. Technology [IEE01a, IEE01g, IEE01e, IEE01c, IEE01b, IEE01h, IEE01f, IEE01d, THG17]. Template [MME14, SN92, Coo89, FLSS93a, FLSS93b, SS94, SA77]. Templates [HL97, ZGY^{+}16]. temporal [PMD01]. tenant [SDA17]. TENCON [Bao93]. Tennessee [ACM90a]. Tenth [IEE94b]. ter [Lia84]. Term [Dur94, Lav91, Pet92, PS93b, KN00, KM84]. Termination [GHW05, JR15b, AP13]. Terms [Cha02b, ZMSD93]. Ternary [KAN^{+}17]. Tessellation [Prů17, TIT83]. Test [Har71, Liu14, AG84, PPPdG20, RP95, SMS15]. testable [Mei08]. Testing [Bre94b, HHW^{+}99, Hei01, Lut02, ZMWL20, GM17, Han92, KKM^{+}85, MF96]. tests [Thi93]. Texas [ACM97c, NEH90, IEE94b, IEE95b]. Text [BBH^{+}87, CC97, Dav82, DW17, Fal85, FMP20, Gib21, GN01, Gon83, Gor00, GV05, GKW^{+}10, HGT24, How97, KR92, KTSA99, KRRW24, Kuk92, KVX12, Man94, Man97, Nao91, NR99b, Nav01c, Pik87, Pik00, Rei77, Ritxx, STSA99, SKF^{+}00, TMK^{+}02, TT82, TJGY22, XPZ^{+}23, ZA87, AMB^{+}02, AVN22, ALLS07, BYG92, BYG96, BCD98, BGFK15a, BGFK15b, BC13a, BPMA02, BPPR20, BFNP10, CL09, CHLS07, CR95a, CM95, CEMW91, CL96, GGF13, Gre88, GV00, How96, Ier09, II08, KR89, KTS^{+}98, KWLL08, KM84, MW92b, Mus03, Mus05, NKT^{+}01, NT05, NM07, OKT92, RH81, San95, SKS96, SNZBY00, WM92a, YT03, dKM04]. Text-Based [GKW^{+}10]. text-compression [CL96]. Texts [BKLP97, BKL^{+}02, BG95, CFG12, CL95, FT04, Lut02, ML96a, Rao95, TMK^{+}02, BFKL13, BSY00, BFG09, CD96, DS04, JU91, KS01, KBN09, NR02, Sen00]. Textual [BH85, Haz01, Joh94b, ZBST14]. Texture [VB98]. tf [TP07a, TP07b]. tf-idf} [TP07a, TP07b]. Their [Brz62, CJM12, Gim73, HN05, MHKR12, OF61, RS59, BRL13, BdFED^{+}20, CX20, GR92, KSVJ15, Lau00, NEH90, Pel87]. Theorem [GL19, Kau92, KP96a, Sha88a, BKM95, HA90, KP96b, Kun95, Moo12, Pie95, Rus85, WZU14]. theoretic [Pie08, Sri93]. Theoretical [CL92, FJ92, MAC14, Med23, BGWXP22]. Theorie [SM74]. Theory [ACM69, ACM74, ACM76, ACM81, ACM84, ACM86, ACM90b, ACM91, ACM92d, ACM93b, ACM94d, ACM95c, ACM97c, ACM99b, ACM00, ACM08, AU72, AU73, DMVT13, Gim73, HU79, HU92, HMU01, HMU07, Lut02, Pet92, ŠW98, AFI98, Bak93, CD18, CS11b, Far19, Han02, HR00, SBR^{+}07, VVV04].

Thinning [Web95]. third [ACM91, AGS93e, Apo92]. Thirteenth [ACM81, ACM94c]. thirtieth [Len11]. Thirty [ACM00, ACM99b]. thirty-first [ACM99b]. Thorn [BFN^{+}09]. thread [San09]. threaded [MAC14]. Threading [OR12]. Three [Cha02a, GPP04, GGL94, HEWK03, KR94, Les79, Les95, de 82, AK08]. Three-Dimensional [GPP04, HEWK03, Les79, Les95]. Threesomes [GP18]. threshold [BSTU08]. thresholds [AD11, ZA17]. Throughput [BTC06, LPT12, TS05, LH13a, LMMN07]. Thue [KKM^{+}85]. tier [KLL23]. Tietze [Rob88]. TIG [Mu 95, MuT95, Mun95]. Tight [BCT93, Col90, Col94a, SV87]. Tighter [CH92, CHPZ95, CH97a]. Time [BC13b, BGG12, BG14, Bur11, CZ01, CF85, CGS17, CH03, Coo72, CR95b, CR95c, CGPR95, CGG^{+}97, CGH^{+}98, FG98, Gal79, Gal81, GS81b, GS83, Gal95, GP01, GP03, HM98, ISNH94, KU99, LKL02, Man75, MHT09, PRU11, Sli78, Sli83, Sto96, VSP08, WBA83, BYR92, BG90, BGM13, CL90, CNS18, CH97b, CD96, CGR99, EH88, EMT23, FLM^{+}10, FG95a, Gal75, Gal76a, GS81a, Gal92, GHST17, GPR95b, GF08, GFG11, GMS12, HKN14, IKX15, IP96, Kos94, KRL87, KKR^{+}13, Liu81, MDLB02, NNSP22, RP95, Rep98, TJD^{+}17, TSI13, YBTB23]. Time- [BC13b]. time-efficient [TJD^{+}17]. time-optimal [IKX15]. time-sliced [KRL87]. Time-Space [CR95c, GPR95b]. Time-space-optimal [GS81b, GS83]. Time/Space [GP01, GP03]. Timed [ACM02, NCJF18, UFA^{+}24, WAH23]. times [Mid98]. timing [AGH^{+}17]. Timothy [Neu10]. tion [Lia84]. title [LGZ^{+}14]. titles [LGZ^{+}14]. TLA [Lut02, Lut02]. TLex [Kea91b]. TM [BGFK15a, BGFK15b]. token [WLF14]. Tokenization [Kul11, Sca11, Rep98]. Tokyo [IEE94a, WN90]. tolerant [BG91a, WLF14, XQW^{+}13]. Tool [MFMA15, MFMA17, Pol01, SvS14, WM92b, dLBHC22, All89, Ier09, KOI94, Nav01b, NT05, SCF^{+}17]. Toolkit [Lut02, VVV04]. Tools [Lut02, PPA10, CGM06, Fri97a, Han02, Ano97a]. Top [GN19, QQC^{+}13a, QQC^{+}13b, Sca11, CLZ^{+}15, FWW13a, FWW13b, OSM94a, OSM94b, OSM94c]. Top- [GN19, QQC^{+}13a, QQC^{+}13b, CLZ^{+}15, FWW13a, FWW13b]. Top-Performance [Sca11]. Topological [D’A98, Fu97, KKM^{+}13]. Topologies [VG01, NCV10]. Topology [ZJL14, MCF^{+}11, MCF^{+}14]. Topology-constrained [ZJL14]. tour [Nav01a]. toys [II09, MI07]. tr [Bar90]. trace [ATdM07]. Traces [NCJF18]. Track [HUN^{+}19, LT03]. Tracking [CZW15, Joh94a, SHS14]. Tractability [Sut21]. tractable [Lei85]. Trade [Abb94, GHST17]. trade-offs [GHST17]. tradition [dBB08]. traffic [BBK12, SLZ^{+}20, WXZY12]. training [VR18]. Trajectories [GVD15, KTY^{+}18]. transaction [Lus94]. transaction-based [Lus94]. transactional [FMG22]. Transducers [BR20, Cro86, KMRY20, Moh94, CFLH^{+}22, EHS07, Gaá04, GHR^{+}16, VHL^{+}12]. Transducing [KR14]. Transform [ABM08, Neu10, DGG^{+}19, TZYH14, ZMAB03]. Transformation [BCC^{+}13, Gro92, Kha16, MFMA17, AK08, EHS07, Fil21, GT90, KH06, Sid95, Sid99, Sid00, TA90]. Transformational [PV91, PS90].

Transformations [ASJDW21, ADR15, DW17, DN77, JM90, KC87, Rob88, SdM01, AK09a, ASJDW18, Arn93a, Arn93b, ETV88, Ryt89, SG12, dLFM07]. transformed [AMB^{+}02]. Transforming [SG16]. transit [BWG12]. transition [BG91a, CW13, GT90]. Transitions [OS11, Gef03, Lau00]. Transitive [AS85, LOZ^{+}24, LH03]. Translating [HSW97, HSW01, Rev91]. Translation [AU72, AU73, Gef03, Ver70b, Ver70a, Rot91, TZYH14]. translational [Man06]. translocations [GFG11]. Transmission [Jok90]. transport [SNM^{+}13]. Transposition [LT03, MNU05, Deo06]. transposition-invariant [Deo06]. transputer [CEMW91]. traversal [NRO12]. traversal-based [NRO12]. Traversals [Sto96]. TRE [Ano13]. Tree [AGT89, AM91, AYCLS02, Cha02b, Cha02c, CHZ06, CH97b, CH03, CMNP17, CMNP18, DNM00, DGM94, FV16, GHLW15, GL19, JWZ94, Kid09, KM94, KLH16, LPR^{+}08, MS98, McC76, MSZ17, MOSZ18, RR90, Shi00, Shi04, Sto96, BDB90, BTG83, CGR03, CLZ^{+}15, Cha87, CLS95, DF00, DGM90, EHS07, FCFM00, Far92a, Far92b, FG99, KH25, KS11a, Kos89, Mal93, SGYM00, SSLL21, TJMC20, Vou06, CGR02]. Tree-Like [GHLW15]. tree-manipulation [Mal93]. Tree-Structured [CMNP17, KM94]. tree-walking [EHS07]. Trees [BYCMW94, BCP02, FK16, GHLW15, Gol93, Gro92, GV05, Gus97, HO82, JWZ94, Nel96, RR92, SCFC94, Sim83, ACFC^{+}16, CPT92, Gro91a, Gro91b, GV00, JRV96, Kos94, Mäk89, TTHP05, TJMC20, Ukk93, Ver92]. Trial [LRV13, LRSV18]. Triangle [IEE89]. Triangles [GP18]. trichotomy [BBG13]. Tricks [Abb94]. Trie [CCH09, GO12, KW19]. tries [BYG96]. Trigram [Cox12]. Trillion [PSP^{+}18]. Triplestores [LRSV18]. Truly [GP92]. trusted [WXZY12]. Tucson [ACM97a, Apo92]. Tumor [WZJH12]. Tuning [Rai92, Smi94]. Turing [GOMSJVGP08]. Turkey [SMD04]. Tutorial [KP96a, Lut02, KP96b]. Twentieth [ACM93a]. Twenty [ACM06, ACM07, AAC^{+}01, AOV^{+}99, B^{+}02, LL08, ACM90b, ACM91, ACM92d, ACM93b, ACM94d, ACM95c, ACM97c]. Twenty-Eighth [B^{+}02]. Twenty-Fifth [ACM06, AOV^{+}99, ACM93b]. twenty-fourth [ACM92d]. twenty-ninth [ACM97c]. twenty-second [ACM90b]. Twenty-Seventh [LL08, AAC^{+}01, ACM95c]. Twenty-Sixth [ACM07, ACM94d]. Twig [DLG12, BKS02, KRML09, MMZ10]. twigs [RM06]. Two [AF92, ABF94a, ABC^{+}04, Ano68, ADLM96, BYN98, BR20, BKLP97, BKL^{+}02, Bir77a, BGJ01, CDJM15, CL95, CHZ06, CHLT14, Coo72, CP91, CR92, CGR93, CCG^{+}94, CGPR95, CGH^{+}98, CIK98, FU98, FNU02, Gal76b, Gia93, HY92, Hum88a, HW12, JSC83, JU91, KPR00, KU99, LY86, Lud77, Mid96, MPW21, Ott94, Par96, Prů17, She59, TIT83, XZL^{+}19, ZT89, AK08, ABF94b, AKT06, AGM05, ADLM01, BYR93a, BYR93b, Bar84a, CK02b, CP10, CCG^{+}93, CR94, GP92, HY90, HR01, HLN09, KLL23, KWL07, KM13, dSOMY15, Par98, Rot91, SN94, VLP17]. Two- [KU99]. Two-Dimensional [ABF94a, ADLM96, BYN98, BKLP97, BKL^{+}02, CL95, CHLT14, CR92, CGPR95, CGH^{+}98, CIK98, FU98, Gia93, HW12, KPR00, Par96, Prů17, ZT89, AF92, ABC^{+}04, CGR93, Mid96, MPW21, TIT83, ABF94b, AKT06, AGM05, ADLM01, BYR93a, BYR93b, CR94, GP92, HLN09, KM13, Par98]. Two-Head [LY86]. Two-Level [JSC83, KWL07, dSOMY15]. two-patterns [CP10]. two-point [Rot91]. Two-Sided [CDJM15]. two-tier [KLL23]. Two-Way [BR20, Coo72, CP91, She59]. Type [JM93, Sou99, Van06, CD18, CGPS13a, CS11b, FF08, JO97, Nil90, Pie08, Rou21, ZBST14]. type-ahead [ZBST14]. type-checking [CGPS13a]. type-theoretic [Pie08]. Typed [JP11, Xi03, YGG^{+}23, CDP16b, Dow91, Tej20]. Types [ASJDW21, FR00, KKFI^{+}24, Pre99, ASJDW18, BC93, CGPS13b, GLS07, GPN96, GHR14, HVP00, HVP05, JGMP22, JD89, KS93, Kra08, dSOMY15, OR11, SG16, TCP16, Vou06]. typeset [San95]. typing [FhDAF09]. Typography [AGS93a, AGS93b, AGS93d, AGS93e, AGS93c]. typy [OA17].

U [HMW^{+}23]. U-Shaped [HMW^{+}23]. U.S. [Bol02]. UK [AOV^{+}99, PC99]. Ukraine [Bro93]. Ultra [ATX21, KW19]. Ultra-Fast [ATX21]. Unambiguity [BK93b, BK93c, BK93d, BK93e]. Unambiguous [BKW92d, Pre99, CX20, SH85]. Unary [Huc21]. Unavoidability [Hei01]. Unbounded [Bre94a, Nil90]. uncertain [GS22, KC15]. uncovering [Edw07]. Undecidability [Hir96, KR95, Dow93, Leu97, XH23]. Understanding [LLS^{+}20, TWZ23]. Unicode [Anoxx, Chi17, Dav99, Dav03, Dav04, Dav21, Dav22, NK07]. Unification [Kni89, Wal88, CDP16b, CD18, DRSS96, OND98]. Unified [BY96a, BGM19, CLST^{+}13, Tar81b, AP13, DLF^{+}15, GW92, Tak96a]. Unifiers [CDP16b]. Uniform [Bre94a]. Uniform-Length [Bre94a]. Unifying [Wol90, KMS^{+}03, MZZ10]. Union [Nag21]. Union-Free [Nag21]. Union-Freeness [Nag21]. Unique [AKT20, ATX21, AG84, GHST17, HTX17, IKX15, Van06]. Unit [BK75, Les94]. Units [LLLC17, GÁSÁ^{+}13]. Universal [FK16, FMG23, GL19, PS10, Sad96, Apo93b, CDDM05, FMG22, Jon13]. University [Ano97b, Hig95, Hwa85, PC99, HWF90]. UNIX [Ano92a, Gt92, Lud77, Mah07, Qui02, Rei77, Rob99a, Rob99b, Fri97b, Hol84, dBB08]. Unix-Like [Hol84]. unless [EMT23]. Unordered [CGS17]. UnQL [BFS00]. unrestricted [Lei85, Leu97]. Unstructured [Gon83, PYY19]. Unsupervised [WWW^{+}16]. UPaK [WKR09]. Update [FG98, FG95a, NNSP22]. updates [Che08]. Upper [CH97a, GG92, Les94, SASU13]. Ups [JN24, PWSG25]. Urbana [Hwa85]. URL [TD18]. USA [ACM06, AP10, Apo92, BGNP94, CG94b, FC98, KP15, SC04, SM09, SM11, DGBH93, FMA02, HF13, IEE09, Kap92, KLB12]. Use [Bol02, IY02a, IY02b, CC97, WSS94, YIAS89]. used [BY96b, Sch91a, Sch91b]. Usefulness [CR91]. USENIX [USE92]. User [KMRY20, SOR16, KMR21]. User-Defined [KMRY20, KMR21]. users [BJK^{+}12]. Using [AGT89, BYCMW94, BCP02, Bow87, BK93g, BZ98, CvW18, CF85, CHP92, CFM00, CW84, Cop91, Dav73, EZYA23, Far19, Goe95, GHK^{+}91, Gro92, GL86, GH82, HEWK03, How97, JM85, JLH18, KP96a, KKK11, Kin89, LSW08, LY17, LLLL08, LLCC13, MS98, Mar89, MUHT96, MPN^{+}14, Mei15, Mu 95, MuT95, PAMP12, Rez92, SBHM94, Sch95, SHvR^{+}16, STSA99, Spi99b, ST95, TMV^{+}01, TB00, WZJH12, WBA83, Yun12, ZGY^{+}16, AVN22, AG06, BSY00, BDB90, BGHZ15, BCD14, BBHK14, BWG12, CADA18, CPW88, CP97, CGT25, CEMW91, DA18, FL71, GS81a, GHS12, GS06, HTK^{+}21, HM00, HHLS06, II08, JLHB92, JGMP22, JSH09, KCPC13, KW05, KP96b, KKM^{+}13, KT14, KAT07, KST94, Kim99, KWLL08, KJS17, KD15, KM13, KMM15, KST16, Lab12a, Lab12b, LKM23, LS09, LT90a, LMT16, MW92b]. using [MLC08, MDLB02, Mun95, Mus05, NYuR15, Neb06, NK07, OK94, PIR17, San15, SD91a, SD91b, SW93, SMS15, STKD20, SG16, Spi99a, TM05b, Val09, Vol12, Wri94, ZC89, ZMAB03, ZZH10, ZMSD93]. USL [DWE89]. Utah [SC93, SC04, SM09, SM11, SC95, SC96, SC98, SC99, SC01, SC02, SC03, SM10]. Utilities [ASA17, IEE01e, IEE01f]. Utility [Pit98]. Utilizing [XK92, All82].

V [TT22]. valid [SMS15]. validation [SMS15]. validator [dLFM07]. Valiente [Lab12a, Lab12b]. Valuation [DM11]. values [Nil90]. Vancouver [LL08]. Variable [BL94, BGVW12, KR97]. Variables [FMMS20, CFLH^{+}22, KR95, MF96]. variance [AGW13]. Variant [KKFI^{+}24, Neb06]. variation [Lec92]. Variational [BCWG09]. various [KIH15]. VAX [Gre88]. Vector [BH02, CZW15, FTJ95, HEWK03, LKM23, Mye98, NW13, SWY75, KAT07, Mye99, ZJL14]. Vectorization [MCP17, SALP20]. Vegas [ACM95c]. vehicle [BKLE18]. Verbatim [ELF22]. Verifiable [CWL^{+}21, SV09]. Verification [ASJDW18, ASJDW21, EGP14, GMAS22, KP15, YPG21, Rus92, SMT^{+}86, dH05]. Verified [HL97, Nip98, TN13, TN15, AH14, ELF22]. verifier [MR09a]. verify [MMDdJ11]. Verifying [OR11, Gol90, Pie95, VLP17]. Version [Bol02, HHW^{+}99, HJW^{+}92, Kos94]. versus [ETV21, GGM12, HA90, PW06, TLS16]. Vertices [MSS^{+}19]. Very [ABB93, AAC^{+}01, AOV^{+}99, BYKZ^{+}92, B^{+}02, CLP98, Gon83, NBY99c, PPA10, Smi91b, Sun90a, Sun90b, DC94, ABB93]. VF [Kid09, YK11]. VF-Coding [Kid09]. Via [Eke95, YJ84, AS85, BGP^{+}22, BH13, CBD^{+}23, CDL^{+}15, CS22, EF13, GHW05, GZ94, Kin92, KTY^{+}18, KS96, KVX12, MIH17, Mor02, OS11, RHR^{+}21, RPE81, Shi97, SNM07, YCJK08, ZLN11, ZCÖZ12]. VIATRA2 [HBRV10]. Victoria [ACM92d, ACM08, MG94]. video [ASG99]. View [BY96a, CJ93, JR15a]. Views [LT90b, XK92, Wad87]. Virtual [CFM17, Cox09]. Virus [OWP16, PWW^{+}11]. Vision [CVP86]. Visitor [Pit98]. Visual [LLS^{+}20, LZ96, CM90, PW06]. visualization [DKP11, HR00]. Visualizing [Joh94b]. VLDB [ABB93, B^{+}02]. VLDC [VRD01]. VLSI [CF85, Hur84, LHMH91, ME97, TYNM86]. Volume [Knu05]. volumes [Ruc15]. vulnerabilities [PM23].

W [ACM87, FvGGM90]. walking [EHS07]. Walks [RKH02]. want [LLS^{+}20]. Warnings [Mar07, KSVJ15]. Warp [WBA83]. warping [TSI13]. wars [Hum88b]. Warwick [PC99]. Washington [ACM74, ACM84, ACM98, HWF90]. Waterloo [Cha86]. Wavelet [SJNS19, GO12]. WAW [A^{+}08]. Way [BR20, Coo72, CP91, JL96, LY86, Sch91a, Sch91b, She59, Ukk10, CR87, GGL94, JL93, Wad87]. Weak [ACR01, CFP19, For02, GGM12, FS22, VR18]. web [Agu23, A^{+}08, LGZ^{+}14, SMS15, AV23, AOMC07, AGP18, CMRV10, KPP19, LZ18, SWZ01, Tay97]. web-graph [A^{+}08]. web-search [LGZ^{+}14]. WebCollectives [Agu23]. Weight [PRU11]. Weighted [BLP18, CLOZ04, DM11, Rob88, BLLP90]. Weights [Nav04a]. Welding [Mu 95, MuT95, Mun95]. Wellfounded [AP13]. West [Hig95]. WHAM [LPT12]. Wheeler [Neu10, ABM08, DGG^{+}19, ZMAB03]. where [Dow93, JGMP22]. Which [Gal76b, Gon02]. WI [FMA02]. wide [HFS05]. widely [BY96b]. Wild [Cox10a, BvdM17]. wildcard [HH16, LMNT16]. Wildcards [GG95, GG97, GKP19, Zha17, Bar22, DA18, Kas08a, Kas08b, Kas08c, LMRT14, SRK20, ZZH10]. Wilson [Rus85]. window [HFS05, PW06]. Windows [FL12b, FG89]. Winter [NEH90, USE92]. Wire [ZL18]. Wire-Speed [ZL18]. Wireless [DCM15, KCPC13]. Wisconsin [ACM81, IEE95a]. wise [MF96]. Within [Wri94, BDFR08, PW93]. Within-word [Wri94]. without [CDP14, CDP16a, Zia96]. Witnesses [HN05, ALLT11]. Woes [BTTF02]. Word [BGG12, Gil85, Kül10, Lia84, Lud77, Mus05, IIK08, II08, KGA^{+}12, KR95, OND98, SNZBY00, Wri94, Zia96]. Word-oriented [Mus05]. Word-Size [BGG12]. Words [Bun95, Fre78, JZAA19, Kuk92, Roo99, BR09, LTV15, MZZ10, Nic03, SMDS94, TN13, TN15, ZZH10]. work [IP96]. work-time [IP96]. worked [Cox12]. Workload [KL02]. Workloads [GWX^{+}23]. Works [HSTS01]. workshop [A^{+}08, BGNP94]. workstation [ZV97]. workstations [MM03]. World [MNR^{+}23, ZdSO18]. worm [FNP09]. worn [BWG12]. Worse [Gal79]. worst [GF08, KT90]. worst-case [KT90]. Writing [Mis89a, Mis89b, YC22].

X [SS93b]. X-ray [SS93b]. XDuce [Fri06b]. Xeon [TLS16]. XML [B^{+}07, ADT15, AL08, BGNV10, B^{+}07, BKS02, Cam99, Che08, CK02b, CGPS13b, CGPS13a, CMRV10, DLG12, Dwe00, EHS07, FK16, GLS07, Hos06, HVP00, HP01, HP03, HVP05, KS07, KH06, KRML09, LM01b, MNS07, MNS10, MZZ10, RM06, TB00, dLFM07]. XPath [SSSS10]. XSDs [MNNS12]. XSS [VS11].

Yacc [Cox10c, MD10]. YARA [RKM21]. Yates [Hyy08]. years [ACFC^{+}16]. York [AP10, Ano97b, HF13]. yourself [Abb77]. Yugoslavia [FGR72].

Z [ABF96]. Z-Compressed [ABF96]. Zakopane [Win78]. zippers [DA20]. Ziv [BK93a, BFG09, FT95, FT98, KKP16a, KKP16b, NR99b, Nav01c, NT05]. Zooming [PW06, GPR95b]. zur [SM74]. Zvi [Ano97b].

References

Aiello:2008:AMW

[A^{+}08] William Anthony Aiello et al., editors. Algorithms and models for the web-graph: fourth international workshop, WAW 2006, Banff, Canada, November 30–December 1, 2006: revised papers, volume 4936 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-78808-5, 3-540-78807-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2006.

Ackenhusen:1986:SBG

[AAB^{+}86] John G. Ackenhusen, Syed S. Ali, David Bishop, Louis F. Rosa, and Reed Thorkildsen. Single-board general-purpose speech recognition system. AT&T Technical Journal, 65(5): 48–59, September 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Amir:2009:PMA

[AAB^{+}09] Amihood Amir, Yonatan Aumann, Gary Benson, Avivit Levy, Ohad Lipsky, Ely Porat, Steven Skiena, and Uzi Vishne. Pattern matching with address errors: Rearrangement distances. Journal of Computer and System Sciences, 75(6):359–370, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022000009000257.

Angles:2017:FMQ

[AAB^{+}17] Renzo Angles, Marcelo Arenas, Pablo Barceló, Aidan Hogan, Juan Reutter, and Domagoj Vrgoc. Foundations of modern query languages for graph databases. ACM Computing Surveys, 50(5):68:1–68:??, November 2017. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Apers:2001:PTS

[AAC^{+}01] Peter M. G. Apers, Paolo Atzeni, Stefano Ceri, Stefano Paraboschi, Kotagiri Ramamohanarao, and Richard T. Snodgrass, editors. Proceedings of the Twenty-seventh International Conference on Very Large Data Bases: Roma, Italy, 11–14th September, 2001. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2001. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001.

Amir:2009:ASM

[AAK^{+}09] Amihood Amir, Yonatan Aumann, Oren Kapah, Avivit Levy, and Ely Porat. Approximate string matching with address bit errors. Theoretical Computer Science, 410(51):5334–5346, November 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:1997:PMS

[AAL^{+}97a] A. Amir, Y. Aumann, G. M. Landau, M. Lewenstein, and N. Lewenstein. Pattern matching with swaps. In IEEE [IEE97], pages 144–153. CODEN ASFPDV. ISBN 0-8186-8197-7, 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Amir:1997:IPM

[AAL97b] Amihood Amir, Alberto Apostolico, and Moshe Lewenstein. Inverse pattern matching. Journal of Algorithms, 24(2):325– 339, August 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0196677496908500.

Amir:2000:PMS

[AAL^{+}00] Amihood Amir, Yonatan Aumann, Gad M. Landau, Moshe Lewenstein, and Noa Lewenstein. Pattern matching with swaps. Journal of Algorithms, 37(2):247–266, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911209.

Auernheimer:1989:NNM

[AB89] Brent Auernheimer and Alison Butler. Neural net model of the neuropsychology of spelling processes. In ACM [ACM89], pages 430–?? ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989.

Ahmed:2009:PSP

[AB09] Reaz Ahmed and Raouf Boutaba. Plexus: a scalable peerto-peer protocol enabling efficient subset search. IEEE/ACM Transactions on Networking, 17(1):130–143, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abbott:1977:DIY

[Abb77] R. J. Abbott. A do-it-yourself instant compiler.kit. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):53–58, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Agrawal:1993:VLD

[ABB93] Rakesh Agrawal, Sean Baker, and David Bell, editors. Very large data bases, VLDB ’93: proceedings of the 19th International Conference on Very Large Data Bases, August 24– 27, 1993, Dublin, Ireland. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1993. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Abbott:1994:TT

[Abb94] Paul Abbott. Tricks of the trade. Mathematica Journal, 4(1):38–42, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:/

/www.mathematica-journal.com/issue/v4i1/tutorials/
tricks/38-42_tricks41.mj.pdf;   http://www.mathematica-
journal.com/issue/v4i1/tutorials/tricks/index.html.

Abbott:1996:X

[Abb96] Paul Abbott. In[] and Out[]. Mathematica Journal, 6(3):14–21, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:/

/www.mathematica-journal.com/issue/v6i3/tutorials/
inout/contents/63inandout.nb;   http://www.mathematica-
journal.com/issue/v6i3/tutorials/inout/contents/63inout.
pdf;   http://www.mathematica-journal.com/issue/v6i3/
tutorials/inout/index.html.

Afek:2016:MDE

[ABBH^{+}16] Yehuda Afek, Anat Bremler-Barr, Yotam Harchol, David Hay, and Yaron Koral. Making DPI engines resilient to algorithmic complexity attacks. IEEE/ACM Transactions on Networking, 24(6):3262–3275, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Amir:2004:TDP

[ABC^{+}04] Amihood Amir, Ayelet Butman, Maxime Crochemore, Gad M. Landau, and Mary Schaps. Two-dimensional pattern matching with rotations. Theoretical Computer Science, 314 (1–2):173–187, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:1994:AIA

[ABF94a] Amihood Amir, Gary Benson, and Martin Farach. An alphabet independent approach to two-dimensional pattern matching. SIAM Journal on Computing, 23(2):313–323, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/22632.

Amir:1994:OTD

[ABF94b] Amihood Amir, Gary Benson, and Martin Farach. Optimal two-dimensional compressed matching. In Automata, languages and programming (Jerusalem, 1994), volume 820 of Lecture Notes in Comput. Sci., pages 215–226. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Amir:1996:LSF

[ABF96] Amihood Amir, Gary Benson, and Martin Farach. Let sleeping files lie: Pattern matching in Z-compressed files. Journal of Computer and System Sciences, 52(2):299–307, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-

2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022000096900239.

Amit:2014:LEP

[ABH^{+}14] Mika Amit, Rolf Backofen, Steffen Heyne, Gad M. Landau, Mathias Möhl, Christina Otto, and Sebastian Will. Local exact pattern matching for non-fixed RNA structures. IEEE/ ACM Transactions on Computational Biology and Bioinformatics, 11(1):219–230, January 2014. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Adjeroh:2008:BWT

[ABM08] Donald Adjeroh, Tim Bell, and Amar Mukherjee. The Burrows–Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching. Springer Science+Business Media, LLC, Boston, MA, USA, 2008. ISBN 0-387-78908-1, 0-387-78909-X (e-book). xxii + 351 pp. LCCN QA76.9.D33 A35 2008.

Amarilli:2020:CDE

[ABMN20] Antoine Amarilli, Pierre Bourhis, Stefan Mengel, and Matthias Niewerth. Constant-delay enumeration for nondeterministic document spanners. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):25–32, September 2020. CODEN SRECD8. ISSN 0163-5808 (print),

1943-5835 (electronic). URL https://dl.acm.org/doi/10.
1145/3422648.3422655.

Abrahamson:1987:GSM

[Abr87] Karl Abrahamson. Generalized string matching. SIAM Journal on Computing, 16(6):1039–1051, December 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1975:ESM

[AC75] Alfred V. Aho and Margaret J. Corasick. Efficient string matching: an aid to bibliographic search. Communications of the Association for Computing Machinery, 18(6):333–340, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agha:1993:AOD

[AC93] Gul Agha and Christian J. Callsen. ActorSpace: an open distributed programming paradigm. ACM SIGPLAN Notices, 28(7):23–32, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atallah:2001:RAA

[ACD01] M. J. Atallah, F. Chyzak, and P. Dumas. A randomized algorithm for approximate string matching. Algorithmica, 29(3):468–486, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=468.

Aldwairi:2005:CSM

[ACF05] Monther Aldwairi, Thomas Conte, and Paul Franzon. Configurable string matching hardware for speeding up intrusion detection. ACM SIGARCH Computer Architecture News, 33 (1):99–107, March 2005. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Apostolico:2016:YST

[ACFC^{+}16] Alberto Apostolico, Maxime Crochemore, Martin Farach-Colton, Zvi Galil, and S. Muthukrishnan. 40 years of suffix trees. Communications of the Association for Computing Machinery, 59(4):66–73, April 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2016/4/200160/fulltext.

ACM:1969:CRA

[ACM69] ACM, editor. Conference record of ACM Symposium on Theory of Computing; papers presented at the symposium, Marina del Rey, California, May 5, 6, 7, 1969. ACM Press, New York, NY 10036, USA, 1969. LCCN QA75.5 .A22 1969.

ACM:1974:CRS

[ACM74] ACM, editor. Conference record of sixth annual ACM Symposium on Theory of Computing: papers presented at the symposium, Seattle, Washington, April 30–May 2, 1974. ACM Press, New York, NY 10036, USA, 1974. LCCN QA76.6 .A13 1974.

ACM:1976:CRE

[ACM76] ACM, editor. Conference record of the eighth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Hershey, Pennsylvania, May 3–5, 1976. ACM Press, New York, NY 10036, USA, 1976. LCCN QA 76.6 A12 1976.

ACM:1981:CPT

[ACM81] ACM, editor. Conference proceedings of the Thirteenth Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Milwaukee, Wisconsin, May 11–13, 1981. ACM Press, New York, NY 10036, USA, 1981. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

ACM:1983:PSA

[ACM83] Proceedings of the Second ACM SIGACT-SIGMOD Symposium on Principles of Database Systems: 21–23 March 1983, Colony Square Hotel, Atlanta, Georgia. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-097-4. LCCN QA76.9.D3 A15 1983. US$25.00.

ACM:1984:PSA

[ACM84] ACM, editor. Proceedings of the sixteenth annual ACM Symposium on Theory of Computing, Washington, DC, April 30– May 2, 1984. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

ACM:1986:PEA

[ACM86] ACM, editor. Proceedings of the Eighteenth annual ACM Symposium on Theory of Computing, Berkeley, California, May 28–30, 1986. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986. ACM order no. 508860.

ACM:1987:PFA

[ACM87] ACM, editor. POPL ’87. Fourteenth Annual ACM SIGACT-SIGPLAN Symposium on Principles of programming languages, January 21–23, 1987, Munich, W. Germany. ACM Press, New York, NY 10036, USA, 1987. ISBN ????

LCCN ???? URL http://www.acm.org/pubs/contents/
proceedings/plan/41625/index.html.

ACM:1989:SAA

[ACM89] ACM, editor. Seventeenth annual ACM Computer Science Conference, February 21–23, 1989, Commonwealth Convention Center, Louisville, Kentucky. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989.

ACM:1990:PPN

[ACM90a] ACM, editor. PODS ’90. Proceedings of the Ninth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems: April 2–4, 1990, Nashville, Tennessee, volume 51(1) of Journal of Computer and Systems Sciences. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

ACM:1990:PTS

[ACM90b] ACM, editor. Proceedings of the twenty-second annual ACM Symposium on Theory of Computing, Baltimore, Maryland, May 14–16, 1990. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-361-2. LCCN QA76.A15 1990. ACM order no. 508900.

ACM:1991:PTT

[ACM91] ACM, editor. Proceedings of the twenty third annual ACM Symposium on Theory of Computing, New Orleans, Louisiana, May 6–8, 1991. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991. ACM order no. 508910.

ACM:1992:CRN

[ACM92a] ACM, editor. Conference record of the Nineteenth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Albuquerque, New Mexico, January 19–22, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-453-8.

LCCN QA76.7 .A15 1992. URL http://www.acm.org/pubs/
contents/proceedings/plan/143165/index.html. ACM

order number 54990.

ACM:1992:PPE

[ACM92b] ACM, editor. PODS ’92. Proceedings of the Eleventh ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, June 2–4, 1992, San Diego, CA. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26 1992. ACM order number 475920.

ACM:1992:PAC

[ACM92c] ACM, editor. Proceedings of the 1992 ACM Conference on LISP and Functional Programming: papers presented at the conference, San Francisco, California, June 22–24, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. Also published as LISP Pointers, vol. V, no. 1, January-March, 1992. ACM order no. 552920.

ACM:1992:PTF

[ACM92d] ACM, editor. Proceedings of the twenty-fourth annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 4–6, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-511-9. LCCN QA76.A15 1992. ACM order no. 508920.

ACM:1993:CRT

[ACM93a] ACM, editor. Conference record of the Twentieth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Charleston, South Carolina, January 10–13, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7

.A15 1993. URL http://www.acm.org/pubs/contents/
proceedings/plan/158511/index.html. ACM order num-

ber 549930.

ACM:1993:PTF

[ACM93b] ACM, editor. Proceedings of the twenty-fifth annual ACM Symposium on the Theory of Computing, San Diego, California, May 16–18, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

ACM:1994:CRP

[ACM94a] ACM, editor. Conference record of POPL ’94, 21st ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: Portland, Oregon, January 17–21, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-636-0. LCCN QA76.7

.A15 1994. URL http://www.acm.org/pubs/contents/
proceedings/plan/174675/index.html.

ACM:1994:IPI

[ACM94b] ACM, editor. ISSAC’94. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

ACM:1994:PPT

[ACM94c] ACM, editor. PODS ’94. Proceedings of the Thirteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, May 24–26, 1994, Minneapolis, MN, volume 13. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

ACM:1994:PTS

[ACM94d] ACM, editor. Proceedings of the twenty-sixth annual ACM Symposium on the Theory of Computing: Montreal, Quebec, Canada, May 23–25, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. ACM order no. 508930.

ACM:1995:CRP

[ACM95a] ACM, editor. Conference record of POPL ’95, 22nd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: San Francisco, California, January 22–25, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-692-1. LCCN QA

76.7 A11 1995. URL http://www.acm.org/pubs/contents/
proceedings/plan/199448/index.html. ACM order num-

ber: 549950.

ACM:1995:PPF

[ACM95b] ACM, editor. PODS ’95. Proceedings of the Fourteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, PODS 1995, San Jose, California, May 22–25, 1995, volume 14. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

ACM:1995:PTS

[ACM95c] ACM, editor. Proceedings of the twenty-seventh annual ACM Symposium on Theory of Computing: Las Vegas, Nevada, May 29–June 1, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995. ACM order no. 508950.

ACM:1997:PPS

[ACM97a] ACM, editor. PODS ’97. Proceedings of the Sixteenth ACM SIG-SIGMOD-SIGART Symposium on Principles of Database Systems, May 12–14, 1997, Tucson, Arizona. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26 1997.

ACM:1997:PEA

[ACM97b] ACM, editor. Proceedings of the Eighth Annual ACM-SIAM Symposium on Discrete Algorithms, New Orleans, Louisiana, January 5–7, 1997. ACM Press, New York, NY 10036, USA, 1997. CODEN PAAAF2. ISBN 0-89871-390-0. LCCN ????

ACM:1997:PTN

[ACM97c] ACM, editor. Proceedings of the twenty-ninth annual ACM Symposium on the Theory of Computing: El Paso, Texas, May 4–6, 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. ACM order no. 508970.

ACM:1998:PPA

[ACM98] ACM, editor. PODS ’98. Proceedings of the ACM SIGACT– SIGMOD–SIGART Symposium on Principles of Database Systems, June 1–3, 1998, Seattle, Washington. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. ACM order number 475980.

ACM:1999:PEA

[ACM99a] ACM, editor. Proceedings of the Eighteenth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 1999: Philadelphia, Pennsylvania, May 31–June 2, 1999. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296 1999. ACM order number 475990.

ACM:1999:PTF

[ACM99b] ACM, editor. Proceedings of the thirty-first annual ACM Symposium on Theory of Computing: Atlanta, Georgia, May 1–4, 1999. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. ACM order number 508990.

ACM:2000:PTS

[ACM00] ACM, editor. Proceedings of the Thirty Second Annual ACM Symposium on Theory of Computing: Portland, Oregon, May 21–23, [2000]. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-184-4. ACM order number 508000.

Asarin:2002:TRE

[ACM02] Eugene Asarin, Paul Caspi, and Oded Maler. Timed regular expressions. Journal of the Association for Computing Machinery, 49(2):172–206, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

ACM:2006:PTF

[ACM06] ACM, editor. Proceedings of the Twenty-Fifth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems, Chicago, IL, USA June 26–28, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

ACM:2007:PTS

[ACM07] ACM, editor. Proceedings of the Twenty-Sixth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2007, Beijing, China, June 11–13, 2007. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-685-8. LCCN ????

ACM:2008:SPA

[ACM08] ACM, editor. STOC ’08: proceedings of the 40th Annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 17–20, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Apostolico:2005:CPM

[ACP05] Alberto Apostolico, Maxime Crochemore, and Kunsoo Park, editors. Combinatorial pattern matching: 16th annual symposium, CPM 2005, Jeju Island, Korea, June 19–22, 2005, proceedings, volume 3537 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-26201-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/content/978-3-540-26201-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=3537;   http://www.springerlink.
com/openurl.asp?genre=volume&id=doi:10.1007/b137128.

Allauzen:2001:EES

[ACR01] Cyril Allauzen, Maxime Crochemore, and Mathieu Raffinot. Efficient experimental string matching by weak factor recognition. Lecture Notes in Computer Science, 2089:51–??, 2001. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-

6119 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2089/20890051.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2089/20890051.pdf.

Amir:2020:DIL

[ACR20] Amihood Amir, Panagiotis Charalampopoulos, and Jakub Radoszewski. Dynamic and internal longest common substring. Algorithmica, 82(12):3707–3743, December 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00453-020-00744-0.

Asperti:2010:REA

[ACT10] Andrea Asperti, Claudio Sacerdoti Coen, and Enrico Tassi. Regular expressions, au point. arXiv.org, ??(??):12, October

13, 2010. URL https://arxiv.org/abs/1010.2604.

Atallah:2011:PMH

[AD11] Mikhail J. Atallah and Timothy W. Duket. Pattern matching in the Hamming distance with thresholds. Information Processing Letters, 111(14):674–677, July 31, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019011001049.

Arikati:1996:AAM

[ADLM96] S. R. Arikati, A. Dessmark, A. Lingas, and M. Marathe. Approximation algorithms for maximum two-dimensional pattern matching. Lecture Notes in Computer Science, 1075: 348–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arikati:2001:AAM

[ADLM01] Srinivasa R. Arikati, Anders Dessmark, Andrzej Lingas, and Madhav V. Marathe. Approximation algorithms for maximum two-dimensional pattern matching. Theoretical Computer Science, 255(1–2):51–62, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:/

/www.elsevier.nl../23/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/197/21/23/article.pdf.

Abouelella:2013:HEI

[ADM^{+}13] Fatma Abouelella, Tom Davidson, Wim Meeus, Karel Bruneel, and Dirk Stroobandt. How to efficiently implement dynamic circuit specialization systems. ACM Transactions on Design Automation of Electronic Systems (TODAES), 18(3):

35:1–35:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ager:2003:FPE

[ADR03] Mads Sig Ager, Olivier Danvy, and Henning Korsholm Rohde. Fast partial evaluation of pattern matching in strings. ACM SIGPLAN Notices, 38(10):243–249, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ager:2006:FPE

[ADR06] Mads Sig Ager, Olivier Danvy, and Henning Korsholm Rohde. Fast partial evaluation of pattern matching in strings. ACM Transactions on Programming Languages and Systems, 28(4):696–714, July 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Alur:2015:DDL

[ADR15] Rajeev Alur, Loris D’Antoni, and Mukund Raghothaman. DReX: a declarative language for efficiently evaluating regular string transformations. ACM SIGPLAN Notices, 50(1): 125–137, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aksoy:2015:RPE

[ADT15] Cem Aksoy, Aggeliki Dimitriou, and Dimitri Theodoratos. Reasoning with patterns to effectively answer XML keyword queries. VLDB Journal: Very Large Data Bases, 24(3):441– 465, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ausaf:2016:PLD

[ADU16] F. Ausaf, R. Dyckhoff, and C. Urban. POSIX lexing with derivatives of regular expressions (proof pearl). In Jasmin Christian Blanchette and Stephan Merz, editors, Interactive Theorem Proving: 7th International Conference, ITP 2016, Nancy, France, August 22–25, 2016, Proceedings, volume 9807 of Lecture Notes in Computer Science, pages 69–86. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Antoy:1994:NNS

[AEH94] Sergio Antoy, Rachid Echahed, and Michael Hanus. A needed narrowing strategy. In ACM [ACM94a], pages 268–279. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http://www.

acm.org:80/pubs/citations/proceedings/plan/174675/
p268-antoy/.

Amir:2011:ASM

[AEK^{+}11] Amihood Amir, Estrella Eisenberg, Orgad Keller, Avivit Levy, and Ely Porat. Approximate string matching with stuck address bits. Theoretical Computer Science, 412(29): 3537–3544, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:2014:MSS

[AEMS14] Alberto Apostolico, Péter L. Erdős, István Miklós, and Johannes Siemons. Modulated string searching. Theoretical Computer Science, 525(??):23–29, March 13, 2014. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513007652.

Amir:2006:SME

[AEP06] Amihood Amir, Estrella Eisenberg, and Ely Porat. Swap and mismatch edit distance. Algorithmica, 45(1):109–120, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=45&
issue=1&spage=109.

Amir:1992:TDD

[AF92] Amihood Amir and Martin Farach. Two-dimensional dictionary matching. Information Processing Letters, 44(5):233– 239, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aceto:1998:QSE

[AFI98] Luca Aceto, Wan Fokkink, and Anna Ingólfsdóttir. On a question of A. Salomaa: The equational theory of regular expressions over a singleton alphabet is not finitely based. Theoretical Computer Science, 209(1–2):163–178, December 06, 1998.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
tcs/sub/1998/209/1-2/2640.pdf.

Amir:1994:ADP

[AFM94] Amihood Amir, Martin Farach, and S. Muthukrishnan. Alphabet dependence in parameterized matching. Information Processing Letters, 49(3):111–115, February 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1984:PMM

[AG84] A. Apostolico and R. Giancarlo. Pattern matching machine implementation of a fast test for unique decipherability. Information Processing Letters, 18(3):155–158, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1986:BMG

[AG86] Alberto Apostolico and Raffaele Giancarlo. The Boyer– Moore–Galil string searching strategies revisited. SIAM Journal on Computing, 15(1):98–105, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apostolico:1997:PMA

[AG97] Alberto Apostolico and Zvi Galil, editors. Pattern matching algorithms. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-511367-5. 377 pp. LCCN

QA76.9.A43 P38 1997. URL http://www.loc.gov/catdir/
enhancements/fy0603/96049602-d.html;   http://www.
loc.gov/catdir/enhancements/fy0603/96049602-t.html.

Atkinson:2006:EPM

[AG06] Darren C. Atkinson and William G. Griswold. Effective pattern matching of source code using abstract syntax patterns. Software — Practice and Experience, 36(4):413–447, ???? 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www.cse.scu.edu/~

atkinson/papers/spe-06.html.

Antonopoulos:2017:DIS

[AGH^{+}17] Timos Antonopoulos, Paul Gazzillo, Michael Hicks, Eric Koskinen, Tachio Terauchi, and Shiyi Wei. Decomposition instead of self-composition for proving the absence of timing channels. ACM SIGPLAN Notices, 52(6):362–375, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anselmo:2005:NOR

[AGM05] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. New operations and regular expressions for two-dimensional languages over one-letter alphabet. Theoretical Computer Science, 340(2):408–431, June 27, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:2019:SPA

[AGM19] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Sets of pictures avoiding overlaps. International Journal of Foundations of Computer Science (IJFCS), 30 (6–7):875–898, September–November 2019. ISSN 0129-

0541. URL https://www.worldscientific.com/doi/10.
1142/S0129054119400215.

Arenas:2018:ELQ

[AGP18] Marcelo Arenas, Georg Gottlob, and Andreas Pieris. Expressive languages for querying the Semantic Web. ACM Transactions on Database Systems, 43(3):13:1–13:??, November 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Andre:1993:ESIa

[AGS93a] Jacques André, Jakob Gonczarowski, and Richard Southall. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference. Electronic Publishing— Origination, Dissemination, and Design, 6(3):115–116, September 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:1993:ESIb

[AGS93b] Jacques André, Jakob Gonczarowski, and Richard Southall. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference. In Jacques André, Jakob Gonczarowski, and Richard Southall, editors, Proceedings of the Raster Imaging and Digital Typography Conference, volume 6(3) of Electronic Publishing—Origination, Dissemination, and Design, pages 115–116. John Wiley, New York, NY, USA, September 1993. CODEN EPODEU. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN ????

Andre:EPODD-6-3-115

[AGS93c] Jacques André, Jakob Gonczarowski, and Richard Southall. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference. Electronic Publishing— Origination, Dissemination, and Design, 6(3):115–116, September 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:1993:ICR

[AGS93d] Jacques André, Jakob Gonczarowski, and Richard Southall, editors. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference, volume 6(3) of Electronic Publishing—Origination, Dissemination, and Design. John Wiley, New York, NY, USA, September 1993. CO-DEN EPODEU. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN ????

Andre:1993:PTI

[AGS93e] Jacques André, Jakob Gonczarowski, and Richard Southall, editors. Proceedings of the third International Conference on Raster Imaging and Digital Typography, 11–13 April 1994, Darmstadt, Germany, volume 6(3) of Electronic Publishing— Origination, Dissemination, and Design. John Wiley, New York, NY, USA, 1993. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Atallah:1996:PMI

[AGS96] M. Atallah, Y. Genin, and W. Szpankowski. Pattern matching image compression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1996.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=488349. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Aho:1989:CGU

[AGT89] Alfred V. Aho, Mahadevan Ganapathi, and Steven W. K. Tjiang. Code generation using tree matching and dynamic programming. ACM Transactions on Programming Languages and Systems, 11(4):491–516, October 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/75700.html.

Agun:2023:WLR

[Agu23] Hayri Volkan Agun. WebCollectives: a light regular expression based web content extractor in Java. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN

2352-7110. URL http://www.sciencedirect.com/science/
article/pii/S2352711023002650.

Atallah:2013:LVR

[AGW13] Mikhail J. Atallah, Elena Grigorescu, and Yi Wu. A lower-variance randomized algorithm for approximate string matching. Information Processing Letters, 113(18):690– 692, September 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019013001622.

Apostolico:1997:CPM

[AH97] Alberto Apostolico and Jotun Hein, editors. Combinatorial pattern matching: 8th Annual Symposium, CPM 97, Aarhus, Denmark, June 30–July 2, 1997: proceedings, volume 1264 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63220-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 S88

1997. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t1264.htm;   http://www.springerlink.
com/content/978-3-540-63220-7;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
1264.

Avigad:2014:FVM

[AH14] Jeremy Avigad and John Harrison. Formally verified mathematics. Communications of the Association for Computing Machinery, 57(4):66–75, April 2014. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aho:1974:DAC

[AHU74] Alfred V. Aho, John E. Hopcroft, and Jeffrey D. Ullman. The Design and Analysis of Computer Algorithms. Addison-Wes-ley, Reading, MA, USA, 1974. ISBN 0-201-00029-6. x + 470 pp. LCCN QA76.6 .A284 1974.

Ade-Ibijola:2018:SRE

[AI18] Abejide Ade-Ibijola. Synthesis of regular expression problems and solutions. International Journal of Computers and Applications, 42(8):748–764, 2018. CODEN IJ-CAFW. ISSN 1206-212X (print), 1925-7074 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
1206212X.2018.1482398.

Augustsson:1989:CLM

[AJ89] L. Augustsson and T. Johnsson. The Chalmers Lazy-ML compiler. The Computer Journal, 32(2):127–141, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/2/127.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/127.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/128.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/129.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/130.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/131.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/132.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/133.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/134.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/135.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/136.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/137.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/138.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/139.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/140.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/141.tif.

Atallah:1992:PMM

[AJS92] Mikhail J. Atallah, Philippe Jacquet, and Wojciech Szpankowski. Pattern matching with mismatches: a probabilistic analysis and a randomized algorithm. Lecture Notes in Computer Science, 644:27–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aiger:2008:AGH

[AK08] Dror Aiger and Klara Kedem. Applying graphics hardware to achieve extremely fast geometric pattern matching in two and three dimensional transformation space. Information Processing Letters, 105(6):224–230, March 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aiger:2009:GPM

[AK09a] Dror Aiger and Klara Kedem. Geometric pattern matching for point sets in the plane under similarity transformations. Information Processing Letters, 109(16):935–940, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anand:2009:OCS

[AK09b] C. K. Anand and W. Kahl. An optimized Cell BE special function library generated by Coconut. IEEE Transactions on Computers, 58(8):1126–1138, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4731241.

Andoni:2012:SCE

[AK12] Alexandr Andoni and Robert Krauthgamer. The smoothed complexity of edit distance. ACM Transactions on Algorithms, 8(4):44:1–44:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Akl:1978:CGM

[Akl78] S. G. Akl. Comments on: G. Manacher, “An application of pattern matching to a problem in geometrical complexity”. Information Processing Letters, 7(2):86–??, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2006:FTD

[AKT06] Amihood Amir, Oren Kapah, and Dekel Tsur. Faster twodimensional pattern matching with rotations. Theoretical Computer Science, 368(3):196–204, December 10, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abedin:2020:SSU

[AKT20] Paniz Abedin, M. Oguzhan Külekci, and Shama V. Thankachan. A survey on shortest unique substring queries. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-
4893/13/9/224.

Akutsu:1994:ASM

[Aku94] T. Akutsu. Approximate string matching with don’t care characters. Lecture Notes in Computer Science, 807:240–249, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akutsu:1995:ASM

[Aku95] Tatsuya Akutsu. Approximate string matching with don’t care characters. Information Processing Letters, 55(5):235– 239, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aho:1985:APS

[AKW85] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language programmer’s manual. Computing Science Technical Report 118, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 5, 1985. ii ++ 38 pp.

Amir:2001:CPM

[AL01] Amihood Amir and Gad M. Landau, editors. Combinatorial pattern matching: 12th annual symposium, CPM 2001, Jerusalem, Israel, July 1–4, 2001: Proceedings, volume 2089 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-42271-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

2001. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2089.htm;   http://www.springerlink.
com/content/978-3-540-42271-6;   http://www.springerlink.

com/openurl.asp?genre=issue&issn=0302-9743&volume=
2089.

Arenas:2008:XDE

[AL08] Marcelo Arenas and Leonid Libkin. XML data exchange: Consistency and query answering. Journal of the Association for Computing Machinery, 55(2):7:1–7:72, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Albert:1989:CMA

[Alb89] Luc Albert. Complexité en moyenne de l’algorithme de multipattern matching RETE sur des ensembles de patterns et d’objets de profondeur un. (French) [Mean complexity of the multi-pattern matching algorithm RETE on sets of patterns and objects of depth one]. Technical report, Institut National de Recherche en Informatique et en Automatique, Le Chesnay, France, 1989. 34 pp.

Alexander:1994:SCS

[Ale94] K. S. Alexander. Shortest common superstrings for strings of random letters. Lecture Notes in Computer Science, 807: 164–172, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allen:1982:FID

[All82] F. W. Allen. A file index for document storage and retrieval utilizing descriptor fragments. The Computer Journal, 25 (1):2–6, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/1/2.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/2.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/3.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/5.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
6.tif.

Allan:1989:POT

[All89] V. H. Allan. Peephole optimization as a targeting and coupling tool. ACM SIGMICRO Newsletter, 20(3):112–121, August 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.1145/75395.75407.

Amir:1997:PMH

[ALL97] A. Amir, M. Lewenstein, and N. Lewenstein. Pattern matching in hypertext. Lecture Notes in Computer Science, 1272: 160–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2000:PMH

[ALL00] Amihood Amir, Moshe Lewenstein, and Noa Lewenstein. Pattern matching in hypertext. Journal of Algorithms, 35(1):82– 99, April 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0196677499910635.

Amir:1998:ESCa

[ALLL98a] A. Amir, G. M. Landau, M. Lewenstein, and N. Lewenstein. Efficient special cases of pattern matching with swaps. Lecture Notes in Computer Science, 1448:209–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1998:ESCb

[ALLL98b] Amihood Amir, Gad M. Landau, Moshe Lewenstein, and Noa Lewenstein. Efficient special cases of pattern matching with swaps. Information Processing Letters, 68(3): 125–132, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/ipl/cas_
sub/browse/browse.cgi?year=1998&volume=68&issue=3&
aid=2077.

Amir:2007:DTS

[ALLS07] Amihood Amir, Gad M. Landau, Moshe Lewenstein, and Dina Sokol. Dynamic text and static pattern matching. ACM Transactions on Algorithms, 3(2):19:1–19:??, May 2007. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aumann:2011:FWP

[ALLT11] Yonatan Aumann, Moshe Lewenstein, Noa Lewenstein, and Dekel Tsur. Finding witnesses by peeling. ACM Transactions on Algorithms, 7(2):24:1–24:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amir:2004:FAS

[ALP04] Amihood Amir, Moshe Lewenstein, and Ely Porat. Faster algorithms for string matching with k mismatches. Journal of Algorithms, 50(2):257–275, February 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740300097X.

Amir:2008:PEC

[ALR08] Amihood Amir, Avivit Levy, and Liron Reuveni. The practical efficiency of convolutions in pattern matching algorithms. Fundamenta Informaticae, 84(1):1–15, September 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amir:1992:EPM

[ALV92] Amihood Amir, Gad M. Landau, and Uzi Vishkin. Efficient pattern matching with scaling. Journal of Algorithms, 13(1):2–32, March 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749290003U.

Aiken:1991:IRT

[AM91] Alexander Aiken and Brian R. Murphy. Implementing regular tree expressions. Lecture Notes in Computer Science, 523: 427–??, 1991. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Antimirov:1995:RER

[AM95] Valentin M. Antimirov and Peter D. Mosses. Rewriting extended regular expressions. Theoretical Computer Science, 143(1):51–72, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1995&volume=143&issue=1&
aid=1726.

Atzeni:1997:CP

[AM97] Paolo Atzeni and Giansalvatore Mecca. Cut and paste. In ACM [ACM97a], pages 144–153. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26 1997. URL http://

www.acm.org/pubs/articles/proceedings/pods/263661/

p144-atzeni/p144-atzeni.pdf;   http://www.acm.org/
pubs/citations/proceedings/pods/263661/p144-atzeni/
;   http://www.acm.org:80/pubs/citations/proceedings/
pods/263661/p144-atzeni/.

Adjeroh:2002:PMB

[AMB^{+}02] D. Adjeroh, A. Mukherjee, T. Bell, M. Powell, and N. Zhang. Pattern matching in BWT-transformed text. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314. LCCN

QA76.9.D33 D37 2002. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=999988. IEEE Computer

Society Order Number: PR01477.

Arenas:2016:FAC

[AMRV16] Marcelo Arenas, Francisco Maturana, Cristian Riveros, and Domagoj Vrgoc. A framework for annotating CSV-like data. Proceedings of the VLDB Endowment, 9(11):876–887, July 2016. CODEN ???? ISSN 2150-8097.

Andrews:2002:KCD

[And02] Larry Andrews. Keyword correction from a dictionary. C/C++ Users Journal, 20(6):14–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1968:TCA

[Ano68] Anonymous. Two complete axiom systems for the extended language of regular expressions. IEEE Transactions on Computers, C-17(7):700–701, July 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1687431.

Anonymous:1987:ESC

[Ano87] Anonymous, editor. EUUG Spring ’87 Conference Proceedings. EurOpen, Buntingford, Herts, UK, 1987.

Anonymous:1992:AUa

[Ano92a] Anonymous. Answers to UNIX. UNIX/world, 9(10):117–??, October 1992. ISSN 0739-5922.

Anonymous:1992:CPM

[Ano92b] Anonymous. Combinatorial Pattern Matching School. Theoretical Computer Science, 92(1):??, January 06, 1992. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1996:JBP

[Ano96a] Anonymous. Java-based pattern matching. SunExpert Magazine, 7(8):74, August 1996. ISSN 1053-9239. Describes VanillaSearch, a search class for Java developers from Thought, Inc..

Anonymous:1996:JPM

[Ano96b] Anonymous. Java-based pattern matching. SunExpert Magazine, 7(8):74, August 1996. ISSN 1053-9239. Describes VanillaSearch, a search class for Java developers from Thought, Inc..

Anonymous:1997:BRMf

[Ano97a] Anonymous. Book review: Mastering regular expressions: Powerful techniques for perl and other tools: By Jeffrey E. F. Friedl. O’Reilly, Sebastopol, CA. (1997). 342 pages. $29.95. Computers and Mathematics with Applications, 33(5):129, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122197829343.

Anonymous:1997:BRPj

[Ano97b] Anonymous. Book review: Pattern matching algorithms: Edited by Alberto Apostolico and Zvi Galil. Oxford University Press, New York, (1997). 377 pages. $65.00. Computers and Mathematics with Applications, 34(10):138–139, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122197902352.

Anonymous:19xx:URE

[Anoxx] Anonymous. Unicode regular expression guidelines. Unicode Technical Report 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, 19xx. URL http://www.unicode.org/

unicode/reports/techreports.html. In progress.

Anonymous:2001:MLP

[Ano01] Anonymous. matchlib — library for pattern matching. Mu-PAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 4 pp. URL https://www.math.

utah.edu/pub/mupad/doc/pdf_help.zip#matchlib.pdf.

Anonymous:2010:BRL

[Ano10] Anonymous. Benchmark of regex libraries. Web site, July

2010. URL http://lh3lh3.users.sourceforge.net/reb.
shtml.

Anonymous:2012:BRR

[Ano12] Anonymous. Book review: Regular Expressions Cookbook, Jan Goyvaerts and Steven Levithan, Second edition, O’Reilly, ISBN 978-1-4493-1943-4. Network Security, 2012(11):4, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print),

1872-9371 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1353485812701009.

Anonymous:2013:TC

[Ano13] Anonymous. TRE (computing). Encyclopedia article, 2013.

URL https://en.wikipedia.org/wiki/TRE_(computing).

Anonymous:2017:ENS

[Ano17] Anonymous. Editor’s note: Special issue on combinatorial pattern matching. Algorithmica, 79(3):797, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s00453-017-0355-8.pdf.

Anonymous:20xx:PPC

[Anoxx] Anonymous. PCRE — Perl compatible regular expressions. Unmaintained Web site., 20xx. URL http://www.pcre.org/.

Antimirov:1995:PDR

[Ant95] V. Antimirov. Partial derivatives of regular expressions and finite automata constructions. Lecture Notes in Computer Science, 900:455–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Antimirov:1996:PDR

[Ant96] Valentin Antimirov. Partial derivatives of regular expressions and finite automaton constructions. Theoretical Computer Science, 155(2):291–319, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2123.

Aoe:1989:EIS

[Aoe89] J. I. Aoe. An efficient implementation of static string pattern matching machines. IEEE Transactions on Software Engineering, 15(8):1010–1016, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/stamp.jsp?arnumber=31357.

Abouelhoda:2002:OES

[AOK02] Mohamed Ibrahim Abouelhoda, Enno Ohlebusch, and Stefan Kurtz. Optimal exact string matching based on suffix arrays. Lecture Notes in Computer Science, 2476:31–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://link.springer.de/link/service/
series/0558/bibs/2476/24760031.htm;   http://link.
springer.de/link/service/series/0558/papers/2476/24760031.
pdf.

Arbe:2007:FLT

[AOMC07] José Miguel Blanco Arbe, Ana Sánchez Ortega, and Jesús Ibáñez Mártinez-Conde. Formal languages through Web forms and regular expressions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4): 100–104, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Atkinson:1999:PTF

[AOV^{+}99] Malcolm P. Atkinson, Maria E. Orlowska, Patrick Valduriez, Stanley B. Zdonik, and Michael L. Brodie, editors. Proceedings of the Twenty-fifth International Conference on Very Large Databases, Edinburgh, Scotland, UK, 7–10 September, 1999. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1999. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. Also known as VLDB’99.

Arcangeli:1990:PPP

[AP90] J.-P. Arcangeli and C. Pomian. Principles of plasma pattern and alternative structure compilation. Theoretical Computer Science, 71(2):177–191, March 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2010:CPM

[AP10] Amihood Amir and Laxmi Parida, editors. Combinatorial Pattern Matching: 21st Annual Symposium, CPM 2010, New York, NY, USA, June 21–23, 2010. Proceedings, volume 6129 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNCSD9. ISBN 3-642-13508-0 (print), 3-642-13509-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-13509-5.

Abel:2013:WRC

[AP13] Andreas M. Abel and Brigitte Pientka. Wellfounded recursion with copatterns: a unified approach to termination and productivity. ACM SIGPLAN Notices, 48(9):185–196, September 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Apostolico:1992:CPM

[Apo92] Alberto Apostolico, editor. Combinatorial pattern matching: third annual symposium, Tucson, Arizona, USA, April 29–May 1, 1992: proceedings, volume 644 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNCSD9. ISBN 3-540-56024-6 (Berlin), 0-387-56024-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.9.A43 C65 1992. URL http://link.springer-ny.
com/link/service/series/0558/tocs/t0644.htm;   http:
//www.springerlink.com/content/978-3-540-56024-1;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=644.

Apostolico:1993:CPM

[Apo93a] Alberto Apostolico, editor. Combinatorial pattern matching: 4th annual symposium, CPM 93, Padova, Italy, June 1993: proceedings, volume 684 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNCSD9. ISBN 3-540-56764-X (Berlin), 0-387-56764-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43

C653 1993. DM58.00. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0684.htm;   http://
www.springerlink.com/content/978-3-540-56764-6;   http:

//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=684.

Apostolico:1993:ECP

[Apo93b] Alberto Apostolico. Efficient CRCW-PRAM algorithms for universal substring searching. Theoretical Computer Science, 108(2):331–344, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1993&volume=108&issue=2&
aid=1277.

Abel:2013:CPI

[APTS13] Andreas Abel, Brigitte Pientka, David Thibodeau, and Anton Setzer. Copatterns: programming infinite structures by observations. ACM SIGPLAN Notices, 48(1):27–38, January 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allauzen:2000:SOS

[AR00] Cyril Allauzen and Mathieu Raffinot. Simple optimal string matching algorithm. Journal of Algorithms, 36(1):102–116, July 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-

2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910873.

Alonso:2013:ABM

[AR13] Laurent Alonso and Edward M. Reingold. Analysis of Boyer and Moore’s MJRTY algorithm. Information Processing Letters, 113(13):495–497, July 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019013001166.

Abbas:2019:QRE

[ARM^{+}19] Houssam Abbas, Alena Rodionova, Konstantinos Mamouras, Ezio Bartocci, Scott A. Smolka, and Radu Grosu. Quantitative regular expressions for arrhythmia detection. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(5):1586–1597, September 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Arnon:1993:SLD

[Arn93a] Dennis S. Arnon. Scrimshaw: a language for document queries and transformations. Electronic Publishing—Origination, Dissemination, and Design, 6(4):385–396, December 1993. CODEN EPODEU. ISSN 0894-3982.

Arnon:EPODD-6-4-385

[Arn93b] Dennis S. Arnon. Scrimshaw: A language for document queries and transformations. Electronic Publishing— Origination, Dissemination, and Design, 6(4):385–396, December 1993. CODEN EPODEU. ISSN 0894-3982.

Avalle:2016:SAN

[ARS16] Matteo Avalle, Fulvio Risso, and Riccardo Sisto. Scalable algorithms for NFA multi-striding and NFA-based deep packet inspection on GPUs. IEEE/ACM Transactions on Networking, 24(3):1704–1717, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abdali:1985:TCR

[AS85] S. K. Abdali and B. D. Saunders. Transitive closure and related semiring properties via eliminants. Theoretical Computer Science, 40(2-3):257–274, ???? 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aho:1991:MCR

[AS91] Alfred V. Aho and Ravi Sethi. Maintaining cross references in manuscripts. Computing Science Technical Report 127, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 20,

1991. 10 pp. URL https://web.archive.org/web/*/http:
//cm.bell-labs.com/cm/cs/cstr/127.ps.gz.

Archer:2004:EMA

[AS04] Tom Archer and Nishant Sivakumar. Extending MFC applications with the .NET framework. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-17352-X (paperback). xxvii + 626 pp. LCCN QA76.76.A65 A69 2004. URL http://catdir.

loc.gov/catdir/toc/ecip049/2003020714.html.

Azaria:2017:DMM

[ASA17] Amos Azaria, David Sarne, and Yonatan Aumann. Distributed matching with mixed maximum–minimum utilities.

ACM Transactions on Economics and Computation, 5(2): 11:1–11:??, March 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Alzina:1999:PMI

[ASG99] M. Alzina, W. Szpankowski, and A. Grama. 2D-pattern matching image and video compression. In Storer and Cohn [SC99], pages 424–433. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1999.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=755692. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Ashdown:1985:PPM

[Ash85] Ian E. Ashdown. Parallel pattern matching and Fgrep. Dr. Dobb’s Journal of Software Tools, 10(9):46–??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Al-Sibahi:2018:VHL

[ASJDW18] Ahmad Salim Al-Sibahi, Thomas P. Jensen, Aleksandar S. Dimovski, and Andrzej Wasowski. Verification of high-level transformations with inductive refinement types. ACM SIG-PLAN Notices, 53(9):147–160, November 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3393934.3278125.

Al-Sibahi:2021:VPT

[ASJDW21] Ahmad Salim Al-Sibahi, Thomas P. Jensen, Aleksandar S. Dimovski, and Andrzej Wasowski. Verification of program transformations with inductive refinement types. ACM Transactions on Software Engineering and Methodology, 30(1):5:1– 5:33, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.org/

doi/10.1145/3409805.

Al-Ssulami:2017:FSM

[ASM17] Abdulrakeeb M. Al-Ssulami and Hassan Mathkour. Faster string matching based on hashing and bit-parallelism. Information Processing Letters, 123(??):51–55, July 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019017300492.

Asperti:2012:CPD

[Asp12] Andrea Asperti. A compact proof of decidability for regular expression equivalence. Lecture Notes in Computer Science, 7406:283–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32347-8_19/.

Apostolico:2002:CPM

[AT02] Alberto Apostolico and Masayuki Takeda, editors. Combinatorial Pattern Matching: 13th Annual Symposium, CPM 2002 Fukuoka, Japan, July 3–5, 2002. Proceedings, volume 2373 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43862-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43

C65 2002. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t2373.htm;   http://www.springerlink.
com/content/978-3-540-43862-5;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
2373.

Avgustinov:2007:MTM

[ATdM07] Pavel Avgustinov, Julian Tibble, and Oege de Moor. Making trace monitors feasible. ACM SIGPLAN Notices, 42(10): 589–608, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2021:UFP

[ATX21] Daniel R. Allen, Sharma V. Thankachan, and Bojian Xu. An ultra-fast and parallelizable algorithm for finding k-mismatch shortest unique substrings. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 18(1):138–148, January 2021. CODEN ITCBCY. ISSN 1545-5963 (print),

1557-9964 (electronic). URL https://dl.acm.org/doi/10.
1109/TCBB.2020.2968531.

Aho:1972:TPT

[AU72] Alfred V. Aho and Jeffrey D. Ullman. The Theory of Parsing, Translation, and Compiling. I: Parsing, volume I. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1972. ISBN 0-13-914556-7. xviii + 542 pp. LCCN QA76.6 .A286 1972-73.

Aho:1973:TPT

[AU73] Alfred V. Aho and Jeffrey D. Ullman. The Theory of Parsing, Translation, and Compiling. II. Compiling, volume II. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1973. ISBN 0-13-914564-8. xiii + 460 pp. LCCN QA76.6 .A286 1972-73.

Alotaibi:2023:TSB

[AV23] Fahad M. Alotaibi and Vassilios G. Vassilakis. Toward an SDN-based Web application firewall: Defending against SQL injection attacks. Future Internet, 15(5):170, April 29, 2023. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/15/5/170.

Alves:2022:SSS

[AVN22] Leonardo Fuchs Alves, Francisco J. S. Vasconcellos, and Bruno Magalhães Nogueira. SeSG: a search string generator for Secondary Studies with hybrid search strategies using text mining. Empirical Software Engineering, 27(5):??, September 2022. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL https://link.springer.com/
article/10.1007/s10664-021-10084-4.

Arratia:1989:ERS

[AW89] R. Arratia and M. S. Waterman. The Erdős–Renyi Strong Law for pattern matching with a given proportion of mismatches. Annals of Probability, 17(3):1152–1169, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL http://projecteuclid.org/euclid.aop/
1176991262.

Angstadt:2018:MOS

[AWD^{+}18] Kevin Angstadt, Jack Wadden, Vinh Dang, Ted Xie, Dan Kramp, Westley Weimer, Mircea Stan, and Kevin Skadron. MNCaRT: An open-source, multi-architecture automataprocessing research and execution ecosystem. IEEE Computer Architecture Letters, 17(1):84–87, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Angstadt:2016:RPP

[AWS16] Kevin Angstadt, Westley Weimer, and Kevin Skadron. RAPID programming of pattern-recognition processors. Operating Systems Review, 50(2):593–605, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aycock:2015:SCS

[Ayc15] J. Aycock. Short communication: Stringlish: improved English string searching in binary files. Software — Practice and Experience, 45(11):1591–1595, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amer-Yahia:2002:TPQ

[AYCLS02] S. Amer-Yahia, S. Cho, L. V. S. Lakshmanan, and D. Srivastava. Tree pattern query minimization. VLDB Journal: Very Large Data Bases, 11(4):315–331, December 2002. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.de/link/service/
journals/00778/bibs/2011004/20110315.htm;   http:/
/link.springer.de/link/service/journals/00778/papers/
2011004/20110315.pdf.

Aoe:1984:MIS

[AYS84] J. Aoe, Y. Yamamoto, and R. Shimada. A method for improving string pattern matching machines. IEEE Transactions on Software Engineering, SE-10(1):116–120, January/ February 1984. CODEN IESEDJ. ISSN 0098-5589 (print),

1939-3520 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5010205.

Aouar:2024:DPS

[AYS^{+}24] Aissam Aouar, Saïd Yahiaoui, Lamia Sadeg, Nadia Nouali-Taboudjemat, and Kadda Beghdad Bey. Distributed partial simulation for graph pattern matching. The Computer Journal, 67(1):110–126, January 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/academic.oup.com/comjnl/article/67/1/110/6832432.

Bernstein:2002:VPT

[B^{+}02] Philip A. Bernstein et al., editors. VLDB 2002: proceedings of the Twenty-Eighth International Conference on Very Large Data Bases, Hong Kong SAR, China, 20–23 August 2002. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2002. ISBN 1-55860-869-9. LCCN ????

Babin:2005:PRP

[B^{+}05] Lee Babin et al. PHP 5 recipes: a problem-solution approach. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-509-2. xxi + 646 pp. LCCN QA76.73.P224 P455 2005.

Bourret:2007:AXA

[B^{+}07] Ronald Bourret et al. Advanced XML applications from the experts at The XML Guild. Thomson Course Technology, Boston, MA, USA, 2007. ISBN 1-59863-214-0 (paperback). xxii + 362 pp. LCCN QA76.76.H94 A32 2007.

Bansal:2006:AGP

[BA06] Sorav Bansal and Alex Aiken. Automatic generation of peephole superoptimizers. ACM SIGPLAN Notices, 41(11):394– 403, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beal:2015:EPM

[BA15] Richard Beal and Donald Adjeroh. Efficient pattern matching for RNA secondary structures. Theoretical Computer Science, 592(??):59–71, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030439751500420X.

Beal:2016:CPP

[BA16] Richard Beal and Donald Adjeroh. Compressed parameterized pattern matching. Theoretical Computer Science, 609 (part 1)(??):129–142, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397515008440.

Backofen:1994:RPE

[Bac94] Rolf Backofen. Regular path expressions in feature logic. Journal of Symbolic Computation, 17(5):421–455, May 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Blume:2006:EPF

[BAC06] Matthias Blume, Umut A. Acar, and Wonseok Chae. Extensible programming with first-class cases. ACM SIGPLAN Notices, 41(9):239–250, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bando:2012:SLR

[BAC12] Masanori Bando, N. Sertac Artan, and H. Jonathan Chao. Scalable lookahead regular expression detection system for deep packet inspection. IEEE/ACM Transactions on Networking, 20(3):699–714, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baker:1978:TER

[Bak78] Theodore P. Baker. A technique for extending rapid exactmatch string matching to arrays of more than one dimension. SIAM Journal on Computing, 7(4):533–541, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1993:TPP

[Bak93] Brenda S. Baker. A theory of parameterized pattern matching: algorithms and applications. In ACM [ACM93b], pages 71–80. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/pubs/articles/proceedings/
stoc/167088/p71-baker/p71-baker.pdf;   http://www.
acm.org/pubs/citations/proceedings/stoc/167088/p71-

baker/. ACM order no. 508930.

Baker:1996:PPM

[Bak96] Brenda S. Baker. Parameterized pattern matching: Algorithms and applications. Journal of Computer and System Sciences, 52(1):28–42, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000096900033.

Balbaert:2015:RE

[Bal15] Ivo Balbaert. Rust Essentials. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-213-1. x + 161 pp. LCCN QA76.73.R87 B35 2015. URL http://proquest.

safaribooksonline.com/9781785285769.

Bingol:2024:GGF

[BAM^{+}24] Zülal Bingöl, Mohammed Alser, Onur Mutlu, Ozcan Ozturk, and Can Alkan. GateKeeper-GPU: Fast and accurate prealignment filtering in short read mapping. IEEE Transactions on Computers, pages 1–12, 2024. CODEN ITCOB4. ISSN 2326-3814.

Baozong:1993:PTI

[Bao93] Yuan Baozong, editor. Proceedings / TENCON ’93, 1993 IEEE Region 10 Conference on Computer, Communication, Control, and Power Engineering, October 19–21, 1993, Beijing, Beijing International Convention Center, Beijing Continental Grand Hotel. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1233-3. LCCN QA75.5.I155 1993. IEEE Catalog No. 93CH3286-2.

Ben-Amram:2006:BSL

[BAP06] Amir M. Ben-Amram and Holger Petersen. Backing up in singly linked lists. Journal of the Association for Computing Machinery, 53(4):681–705, July 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barth:1981:AIK

[Bar81] Gerhard Barth. An alternative for the implementation of the Knuth–Morris–Pratt algorithm. Information Processing Letters, 13(4–5):134–137, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1984:ACT

[Bar84a] Gerhard Barth. An analytical comparison of two string searching algorithms. Information Processing Letters, 18(5):249– 256, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barton:1984:SSF

[Bar84b] Paul C. Barton. String search function in C [letter]. Dr. Dobb’s Journal of Software Tools, 9(8):8–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Barrett:1990:RGA

[Bar90] Robert Barrett. References on grep, awk, sed, and tr. C Users Journal, 8(3):138–??, March 1990. ISSN 0898-9788.

Barton:2022:ACC

[Bar22] Carl Barton. On the average-case complexity of pattern matching with wildcards. Theoretical Computer Science, 922(??):37–45, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397522002092.

Bagan:2013:TRS

[BBG13] Guillaume Bagan, Angela Bonifati, and Benoit Groz. A trichotomy for regular simple path queries on graphs. In Hull and Fan [HF13], pages 261–272. ISBN 1-4503-2066-X, 1-4503-

2037-6. LCCN ???? URL http://dl.acm.org/citation.
cfm?id=2463664;   http://www.sigmod.org/2013/.

Blumer:1987:CIF

[BBH^{+}87] A. Blumer, J. Blumer, D. Haussler, R. McConnell, and A. Ehrenfeucht. Complete inverted files for efficient text retrieval and analysis. Journal of the Association for Computing Machinery, 34(3):578–595, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/28873.html.

Bremler-Barr:2014:CSP

[BBHK14] Anat Bremler-Barr, David Hay, and Yaron Koral. CompactDFA: Scalable pattern matching using longest prefix match solutions. IEEE/ACM Transactions on Networking, 22 (2):415–428, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bremler-Barr:2012:AMM

[BBK12] Anat Bremler-Barr and Yaron Koral. Accelerating multipattern matching on compressed HTTP traffic. IEEE/ACM Transactions on Networking, 20(3):970–983, June 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bose:1993:PMP

[BBL93] P. Bose, J. F. Buss, and A. Lubiw. Pattern matching for permutations. Lecture Notes in Computer Science, 709:200– ??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bose:1998:PMP

[BBL98] Prosenjit Bose, Jonathan F. Buss, and Anna Lubiw. Pattern matching for permutations. Information Processing Letters, 65(5):277–283, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borsotti:2021:DPA

[BBM21] Angelo Borsotti, Luca Breveglieri, and Angelo Morzenti. A deterministic parsing algorithm for ambiguous regular expressions. Acta Informatica, 58(3):195–229, June 2021. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00236-020-00366-7.

Borsotti:2023:GPR

[BBRM23] Angelo Borsotti, Luca Breveglieri, Stefano Crespi Reghizzi, and Angelo Morzenti. General parsing with regular expression matching. Journal of Computer Languages, 74(??):1– 20, January 2023. CODEN ???? ISSN 2665-9182 (print),

2590-1184 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S2590118422000739.

Berglund:2021:FIB

[BBvdM21] Martin Berglund, Willem Bester, and Brink van der Merwe. Formalising and implementing Boost POSIX regular expression matching. Theoretical Computer Science, 857(??):147–165, February 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397521000232.

Burton:1993:PMA

[BC93] F. Warren Burton and Robert D. Cameron. Pattern matching with abstract data types. Journal of Functional Programming, 3(2):171–190, April 1993. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
04DD26A0E6CA3A1E87E0E6AE8BC02EED.

Brazma:1994:ELR

[BC94] A. Brazma and K. Cerans. Efficient learning of regular expressions from good examples. Lecture Notes in Computer Science, 872:76–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bunke:1995:IAC

[BC95] H. Bunke and J. Csirik. An improved algorithm for computing the edit distance of run-length coded strings. Information Processing Letters, 54(2):93–96, April 28, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bird:2006:BSE

[BC06] Steven Bird and James R. Curran. Building a search engine to drive problem-based learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 153–157, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bast:2013:EFS

[BC13a] Hannah Bast and Marjan Celikik. Efficient fuzzy search in large text collections. ACM Transactions on Information Systems, 31(2):10:1–10:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Becchi:2013:DTS

[BC13b] Michela Becchi and Patrick Crowley. A-DFA: a time- and space-efficient DFA compression algorithm for fast regular expression evaluation. ACM Transactions on Architecture and Code Optimization, 10(1):4:1–4:26, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Butman:2013:PMU

[BCC^{+}13] Ayelet Butman, Peter Clifford, Raphaël Clifford, Markus Jalsenius, Noa Lewenstein, Benny Porat, Ely Porat, and Benjamin Sach. Pattern matching under polynomial transformation. SIAM Journal on Computing, 42(2):611–633, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barcaccia:1998:PMT

[BCD98] P. Barcaccia, A. Cresti, and S. De Agostino. Pattern matching in text compressed with the ID heuristic. In Storer and Cohn [SC98], pages 113–118. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33

D232 1998. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=672137. IEEE Computer Society

Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bernardi:2014:DPD

[BCD14] Mario Luca Bernardi, Marta Cimitile, and Giuseppe Di Lucca. Design pattern detection using a DSL-driven graph matching approach. Journal of Software: Evolution and Process, 26(12): 1233–1266, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Burcsi:2012:AJP

[BCFL12] Péter Burcsi, Ferdinando Cicalese, Gabriele Fici, and Zsuzsanna Lipták. Algorithms for jumbled pattern matching in strings. International Journal of Foundations of Computer Science (IJFCS), 23(2):357–374, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Baeten:2007:CRE

[BCG07] J. C. M. Baeten, F. Corradini, and C. A. Grabmayer. A characterization of regular expressions under bisimulation. Journal of the Association for Computing Machinery, 54(2):6:1–6:28, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bliznets:2015:KLB

[BCKM15] Ivan Bliznets, Marek Cygan, Pawel Komosa, and Lukás Mach. Kernelization lower bound for Permutation Pattern Matching. Information Processing Letters, 115(5):527–531, May 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-

6119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0020019015000101.

Bartolini:2002:SMM

[BCP02] Ilaria Bartolini, Paolo Ciaccia, and Marco Patella. String matching with metric trees using an approximate distance. Lecture Notes in Computer Science, 2476:271–??, 2002. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://link.springer.de/link/service/
series/0558/bibs/2476/24760271.htm;   http://link.
springer.de/link/service/series/0558/papers/2476/24760271.
pdf.

Breslauer:1993:TCB

[BCT93] Dany Breslauer, Livio Colussi, and Laura Toniolo. Tight comparison bounds for the string prefix-matching problem. Information Processing Letters, 47(1):51–57, August 9, 1993.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1994:ECS

[BCT94] D. Breslauer, L. Colussi, and L. Toniolo. On the exact complexity of the string prefix-matching problem. Lecture Notes in Computer Science, 855:483–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1998:CCS

[BCT98] Dany Breslauer, Livio Colussi, and Laura Toniolo. On the comparison complexity of the string prefix-matching problem. Journal of Algorithms, 29(1):18–67, October 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909488.

Ben-Chen:2009:VHM

[BCWG09] Mirela Ben-Chen, Ofir Weber, and Craig Gotsman. Variational harmonic maps for space deformation. ACM Transactions on Graphics, 28(3):34:1–34:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bailey:1980:FSS

[BD80] T. A. Bailey and R. G. Dromey. Fast string searching by finding subkeys in subtext. Information Processing Letters, 11(3):130–133, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bidoit:1998:FST

[BD98] N. Bidoit and S. De Amo. A first step towards implementing dynamic algebraic dependences. Theoretical Computer Science, 190(2):115–149, January 20, 1998. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
tcs/sub/1998/190/2/2624.pdf.

Balachandran:1990:ERC

[BDB90] A. Balachandran, D. M. Dhamdhere, and S. Biswas. Efficient retargetable code generation using bottom-up tree pattern matching. Computer Languages, 15(3):127–140, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bartoli:2014:ASR

[BDD^{+}14] Alberto Bartoli, Giorgio Davanzo, Andrea De Lorenzo, Eric Medvet, and Enrico Sorio. Automatic synthesis of regular expressions from examples. Computer, 47(12):72–80, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://csdl.computer.org/csdl/mags/
co/2014/12/mco2014120072-abs.html.

Barozzini:2020:BRL

[BdFED^{+}20] David Barozzini, David de Frutos-Escrig, Dario Della Monica, Angelo Montanari, and Pietro Sala. Beyond ω-regular languages: ωT -regular expressions and their automata and logic counterparts. Theoretical Computer Science, 813(??):270–304, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397519308114.

Ben-David:2008:EFA

[BDFR08] Shoham Ben-David, Dana Fisman, and Sitvanit Ruah. Embedding finite automata within regular expressions. Theoretical Computer Science, 404(3):202–218, September 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:1994:PCS

[BDFW94] H. Bodlaender, R. G. Downey, M. R. Fellows, and H. T. Wareham. The parameterized complexity of sequence alignment and consensus. Lecture Notes in Computer Science, 807:15– 30, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brzozowski:2019:SCP

[BDM19] Janusz A. Brzozowski, Sylvie Davies, and Abhishek Madan. State complexity of pattern matching in regular languages. Theoretical Computer Science, 777(??):121–131, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397518307357.

Bartoli:2016:RBE

[BDMT16] Alberto Bartoli, Andrea De Lorenzo, Eric Medvet, and Fabiano Tarlao. Regex-based entity extraction with active learning and genetic programming. ACM SIGAPP Applied Computing Review, 16(2):7–15, August 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2993231.2993232.

Beebe:1981:IPM

[Bee81] Nelson H. F. Beebe. Ideas for pattern matching software. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 12, 1981. 3 pp.

Beebe:2013:BPM

[Bee13] Nelson H. F. Beebe. A bibliography on pattern matching, regular expressions, and string matching. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 13, 2013.

292 pp. URL https://www.math.utah.edu/pub/tex/bib/
index-table-s.html#string-matching. This report is up-

dated frequently.

Butman:2004:SPS

[BEL04] Ayelet Butman, Revital Eres, and Gad M. Landau. Scaled and permuted string matching. Information Processing Letters, 92(6):293–297, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

BenNsira:2017:LSM

[BEL17] Nadia Ben Nsira, Mourad Elloumi, and Thierry Lecroq. Online string matching in highly similar DNA sequences. Mathematics in Computer Science, 11(2):113–126, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Baron:2012:SPM

[BEM^{+}12] Joshua Baron, Karim El Defrawy, Kirill Minkovich, Rafail Ostrovsky, and Eric Tressler. 5PM: Secure pattern matching. Lecture Notes in Computer Science, 7485:222–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32928-9_13/.

Baron:2013:SPM

[BEM^{+}13] Joshua Baron, Karim El Defrawy, Kirill Minkovich, Rafail Ostrovsky, and Eric Tressler. 5PM: Secure pattern matching. Journal of Computer Security, 21(5):601–625, ???? 2013.

CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Bentley:1986:PP

[Ben86] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10331-1 (paperback). viii + 195 pp. LCCN QA76.6.B453 1986. Reprinted with corrections.

Benson:1994:SEA

[Ben94] G. Benson. A space efficient algorithm for finding the best non-overlapping alignment score. Lecture Notes in Computer Science, 807:1–14, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:2000:PP

[Ben00] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-65788-0 (paperback), 0-13-449802-X (e-book), 0-13-449805-4. xi + 239 pp. LCCN QA76.6 .B454 2000. US$24.95.

Berry:2000:CBM

[Ber00] David Berry. Combining Boyer–Moore string search with regular expressions. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Bernecky:1997:PBI

[BF97] Robert Bernecky and Michael Fitzpatrick. Programmer’s bookshelf: Inner loops and regular expressions. Dr. Dobb’s Journal of Software Tools, 22(12):141, 143, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Bille:2008:FCR

[BFC08] Philip Bille and Martin Farach-Colton. Fast and compact regular expression matching. Theoretical Computer Science, 409(3):486–496, December 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bille:2009:IAS

[BFG09] Philip Bille, Rolf Fagerberg, and Inge Li Gørtz. Improved approximate string matching and regular expression matching on Ziv–Lempel compressed texts. ACM Transactions on Algorithms, 6(1):3:1–3:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Buchsbaum:2003:FPM

[BFK^{+}03] Adam L. Buchsbaum, Glenn S. Fowler, Balachannder Kirishnamurthy, Kiem-Phong Vo, and Jia Wang. Fast prefix matching of bounded strings. ACM Journal of Experimental Algorithmics, 8:1.3:1–1.3:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Badkobeh:2013:BJS

[BFKL13] Golnaz Badkobeh, Gabriele Fici, Steve Kroon, and Zsuzsanna Lipták. Binary jumbled string matching for highly runlength compressible texts. Information Processing Letters, 113(17):604–608, August 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019013001415.

Bloom:2009:TRC

[BFN^{+}09] Bard Bloom, John Field, Nathaniel Nystrom, Johan Östlund, Gregor Richards, Rok Strniša, Jan Vitek, and Tobias Wrigstad. Thorn: robust, concurrent, extensible scripting on the JVM. ACM SIGPLAN Notices, 44(10):117–136, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brisaboa:2010:DLT

[BFNP10] Nieves Brisaboa, Antonio Fariña, Gonzalo Navarro, and José Paramá. Dynamic lightweight text compression. ACM Transactions on Information Systems, 28(3):10:1–10:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Bohannon:2008:BRL

[BFP^{+}08] Aaron Bohannon, J. Nathan Foster, Benjamin C. Pierce, Alexandre Pilkiewicz, and Alan Schmitt. Boomerang: resourceful lenses for string data. ACM SIGPLAN Notices, 43 (1):407–419, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buneman:2000:UQL

[BFS00] Peter Buneman, Mary F. Fernandez, and Dan Suciu. UnQL: a query language and algebra for semistructured data based on structural recursion. VLDB Journal: Very Large Data Bases, 9(1):76–110, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://ftp.informatik.rwth-aachen.de/dblp/db/
indices/a-tree/b/Buneman:Peter.html;   http://ftp.
informatik.rwth-aachen.de/dblp/db/indices/a-tree/f/
Fernandez:Mary_F=.html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/s/Suciu:Dan.html;
 http://link.springer.de/link/service/journals/00778/
bibs/0009001/00090076.htm;   http://link.springer.
de/link/service/journals/00778/papers/0009001/00090076.

pdf. Electronic edition.

Broberg:2004:REP

[BFS04] Niklas Broberg, Andreas Farre, and Josef Svenningsson. Regular expression patterns. ACM SIGPLAN Notices, 39(9):67– 78, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Breslauer:1990:OTP

[BG90] Dany Breslauer and Zvi Galil. An optimal O(log log n) time parallel string matching algorithm. SIAM Journal on Computing, 19(6):1051–1058, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Belli:1991:SFT

[BG91a] F. Belli and K. E. Grosspietsch. Specification of fault-tolerant system issues by predicate/transition nets and regular expressions — approach and case study. IEEE Transactions on Software Engineering, 17(6):513–526, June 1991. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=87278.

Breslauer:1991:LBP

[BG91b] Dany Breslauer and Zvi Galil. A lower bound for parallel string matching. In ACM [ACM91], pages 439–443. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991. URL

http://www.acm.org/pubs/articles/proceedings/stoc/
103418/p439-breslauer/p439-breslauer.pdf;   http:/
/www.acm.org/pubs/citations/proceedings/stoc/103418/

p439-breslauer/. ACM order no. 508910.

Breslauer:1992:LBP

[BG92] Dany Breslauer and Zvi Galil. A lower bound for parallel string matching. SIAM Journal on Computing, 21(5):856–862, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Breslauer:1995:ESM

[BG95] D. Breslauer and L. Gasieniec. Efficient string matching on coded texts. Lecture Notes in Computer Science, 937:27–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:2014:RTS

[BG14] Dany Breslauer and Zvi Galil. Real-time streaming stringmatching. ACM Transactions on Algorithms, 10(4):22:1– 22:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Bille:2022:REM

[BG22] Philip Bille and Inge Li Gørtz. From regular expression matching to parsing. Acta Informatica, 59(6):709–724, December 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/
10.1007/s00236-022-00420-6.

Basik:2015:STS

[BGFK15a] Fuat Basık, Buğra Gedik, Hakan Ferhatosmanoğlu, and Mert Emin Kalender. S^{33}-TM: scalable streaming short text matching. VLDB Journal: Very Large Data Bases, 24(6): 849–866, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Basik:2015:TSS

[BGFK15b] Fuat Basık, Buğra Gedik, Hakan Ferhatosmanoğlu, and Mert Emin Kalender. S^{33}-TM: scalable streaming short text matching. VLDB Journal: Very Large Data Bases, 24(6): 849–866, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CASCON ’94. Integrated Solutions. National Research Council of Canada, Ottawa, ON, Canada, 1994.

Breslauer:2012:CTW

[BGG12] Dany Breslauer, Leszek Gasieniec, and Roberto Grossi. Constant-time word-size string matching. Lecture Notes in Computer Science, 7354:83–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31265-
6_7/.

Barowy:2015:FER

[BGHZ15] Daniel W. Barowy, Sumit Gulwani, Ted Hart, and Benjamin Zorn. FlashRelate: extracting relational data from semistructured spreadsheets using examples. ACM SIGPLAN Notices, 50(6):218–228, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bird:1989:FDP

[BGJ89] R. S. Bird, J. Gibbons, and G. Jones. Formal derivation of a pattern matching algorithm. Science of Computer Programming, 12(2):93–104, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bunke:2001:TBG

[BGJ01] H. Bunke, S. Günter, and X. Jiang. Towards bridging the gap between statistical and structural pattern recognition: Two new concepts in graph matching. Lecture Notes in Computer Science, 2013:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2013/
20130001.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2013/20130001.pdf.

Babenko:2016:CMM

[BGK^{+}16] Maxim Babenko, Paweł Gawrychowski, Tomasz Kociumaka, Ignat Kolesnichenko, and Tatiana Starikovskaya. Computing minimal and maximal suffixes of a substring. Theoretical Computer Science, 638(??):112–121, July 25, 2016. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397515007707.

Breslauer:2013:SRT

[BGM13] Dany Breslauer, Roberto Grossi, and Filippo Mignosi. Simple real-time constant-space string matching. Theoretical Computer Science, 483(??):2–9, April 29, 2013. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397512010900.

Beedkar:2019:UFF

[BGM19] Kaustubh Beedkar, Rainer Gemulla, and Wim Martens. A unified framework for frequent sequence mining with subsequence constraints. ACM Transactions on Database Systems, 44(3):11:1–11:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3321486.

Banerjee:1994:LCP

[BGNP94] Uptal Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 6th international workshop, Portland, Oregon, USA, August 12– 14, 1993: proceedings, volume 768 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Bex:2010:LDR

[BGNV10] Geert Jan Bex, Wouter Gelade, Frank Neven, and Stijn Vansummeren. Learning deterministic regular expressions for the inference of schemas from XML data. ACM Transactions on the Web (TWEB), 4(4):14:1–14:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bernardini:2022:EDS

[BGP^{+}22] Giulia Bernardini, Paweł Gawrychowski, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Elastic-degenerate string matching via fast matrix multiplication. SIAM Journal on Computing, 51(3):??, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.1137/20M1368033.

Bille:2023:SIC

[BGS23] Philip Bille, Inge Li Gørtz, and Teresa Anna Steiner. String indexing with compressed patterns. ACM Transactions on Algorithms, 19(4):32:1–32:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/10.1145/3607141.

Bille:2012:SMV

[BGVW12] Philip Bille, Inge Li Gørtz, Hjalte Wedel Vildhøj, and David Kofoed Wind. String matching with variable length gaps. Theoretical Computer Science, 443(1):25–34, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397512002915.

Boruch-Gruszecki:2022:CDT

[BGWXP22] Aleksander Boruch-Gruszecki, Radosław Waśko, Yichen Xu, and Lionel Parreaux. A case for DOT: theoretical foundations for objects with pattern matching and GADT-style reasoning. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):179:1–179:??, October 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https://dl.

acm.org/doi/10.1145/3563342.

Berkovich:1985:MSP

[BH85] Simon Y. Berkovich and Abd El Fatah A. Hegazy. Matching String Patterns in Large Textual Files. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0639-8. 122–127 pp. LCCN QA75.5 .I6351 1985. IEEE Service Cent. Piscataway, NJ, USA.

Breslauer:1996:OPC

[BH96] Dany Breslauer and Ramesh Hariharan. Optimal parallel construction of minimal suffix and factor automata. Parallel Processing Letters, 6(1):35–44, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Bergeron:2002:VAA

[BH02] A. Bergeron and S. Hamel. Vector algorithms for approximate string matching. International Journal of Foundations of Computer Science (IJFCS), 13(1):53–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brown:2007:RIS

[BH07] Christopher W. Brown and Eric A. Hardisty. RegeXeX: an interactive system providing regular expression exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):445–449, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bloom:2013:RSP

[BH13] Bard Bloom and Martin J. Hirzel. Robust scripting via patterns. ACM SIGPLAN Notices, 48(2):29–40, February 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bird:1977:TDP

[Bir77a] R. S. Bird. Two dimensional pattern matching. Information Processing Letters, 6(5):168–170, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bird:1977:IPI

[Bir77b] Richard S. Bird. Improving programs by the introduction of recursion. Communications of the Association for Computing Machinery, 20(11):856–863, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bird:2010:PFA

[Bir10] Richard Bird. Pearls of Functional Algorithm Design. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51338-3 (hardcover). xi + 277 pp. LCCN QA76.62 .B57 2010.

Blunschi:2012:SGS

[BJK^{+}12] Lukas Blunschi, Claudio Jossen, Donald Kossmann, Magdalini Mori, and Kurt Stockinger. SODA: generating SQL for business users. Proceedings of the VLDB Endowment, 5 (10):932–943, June 2012. CODEN ???? ISSN 2150-8097.

Beyer:1979:LAI

[BJM79] T. Beyer, W. Jones, and S. Mitchell. Linear algorithms for isomorphism of maximal outerplanar graphs. Journal of the Association for Computing Machinery, 26(4):603–610, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bjorner:1993:MFF

[Bjö93] Anders Björner. The Möbius function of factor order. Theoretical Computer Science, 117(1–2):91–98, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1413.

Berkling:1975:CDH

[BK75] K. Berkling and W. Kluge. Conceptual design of a hardwired main storage management unit. The Computer Journal, 18 (4):360–365, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/360.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_04/tiff/360.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/361.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/363.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/365.tif.

Brazma:1986:GRE

[BK86] A. Brazma and E. Kinber. Generalized regular expressions — a language for synthesis of programs with branching in loops. Theoretical Computer Science, 46(2-3):175–195, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burton:1989:FPQ

[BK89] F. Warren Burton and John G. Kollias. Functional programming with quadtrees. IEEE Software, 6(1):90–97, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bell:1993:LMS

[BK93a] Timothy C. Bell and David Kulp. Longest-match string searching for Ziv–Lempel compression. Software — Practice and Experience, 23(7):757–771, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brueggemann-Klein:1993:UER

[BK93b] A. Brueggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Bruggemann-Klein:1993:UERa

[BK93c] A. Brüggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Bruggemann-Klein:1993:UERb

[BK93d] A. Bruggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

BruggemannKlein:1993:UER

[BK93e] A. Bruggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Bruggemann-Klein:1993:REF

[BK93f] Anne Brüggemann-Klein. Regular expressions into finite automata. Theoretical Computer Science, 120(2): 197–213, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1993&volume=120&issue=2&
aid=1327.

Bunke:1993:JPS

[BK93g] H. Bunke and G. Kaufmann. Jigsaw puzzle solving using approximate string matching and best-first search. Lecture Notes in Computer Science, 719:299–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bambenek:2009:GPR

[BK09] John Bambenek and Agnieszka Klus. grep: Pocket reference. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2009. ISBN 0-596-15360-0 (paperback). v + 75 pp. LCCN QA76.76.U84 B36 2009.

Ben-Kiki:2014:TOP

[BKBB^{+}14] Oren Ben-Kiki, Philip Bille, Dany Breslauer, Leszek Gasieniec, Roberto Grossi, and Oren Weimann. Towards optimal packed string matching. Theoretical Computer Science, 525(??):111–129, March 13, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397513004672.

Berman:2002:CPM

[BKL^{+}02] Piotr Berman, Marek Karpinski, Lawrence L. Larmore, Wojciech Plandowski, and Wojciech Rytter. On the complexity of pattern matching for highly compressed twodimensional texts. Journal of Computer and System Sciences, 65(2):332–350, September 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000002918520.

Baek:2018:EGP

[BKLE18] Sun Geol Baek, Dong Hyun Kang, Sungkil Lee, and Young Ik Eom. Efficient graph pattern matching framework for network-based in-vehicle fault detection. The Journal of Systems and Software, 140(??):17–31, June 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121218300347.

Berman:1997:CPM

[BKLP97] P. Berman, M. Karpinski, L. L. Larmore, and W. Plandowski. On the complexity of pattern matching for highly compressed two-dimensional texts. Lecture Notes in Computer Science, 1264:40–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boyer:1995:BMT

[BKM95] R. S. Boyer, M. Kaufmann, and J. S. Moore. The Boyer– Moore theorem prover and its interactive enhancement. Computers and Mathematics with Applications, 29(2):27–62, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0898122194002157.

Bruno:2002:HTJ

[BKS02] Nicolas Bruno, Nick Koudas, and Divesh Srivastava. Holistic twig joins: optimal XML pattern matching. In Franklin et al. [FMA02], pages 310–321. ISBN ???? LCCN ???? ACM order number 475020.

Bruggemann-Klein:1992:DRLa

[BKW92a] A. Brüggemann-Klein and D. Wood. Deterministic regular languages. In Finkel and Jantzen [FJ92], pages 173–184. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

Bruggemann-Klein:1992:DRLb

[BKW92b] A. Bruggemann-Klein and D. Wood. Deterministic regular languages. In Finkel and Jantzen [FJ92], pages 173–184. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

BruggemannKlein:1992:DRL

[BKW92c] A. Bruggemann-Klein and D. Wood. Deterministic regular languages. In Finkel and Jantzen [FJ92], pages 173–184. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

Bruggemann-Klein:1992:URE

[BKW92d] Anne Brüggemann-Klein and Derick Wood. Unambiguous regular expressions and SGML document grammars. Technical Report 337, Computer Science Department, University of Western Ontario, London, Ontario, Canada, November 1992. ISBN 0-7714-1454-4.

Bertossi:1994:PSM

[BL94] A. A. Bertossi and F. Logi. Parallel string matching with variable length don’t cares. Journal of Parallel and Distributed Computing, 22(2):229–234, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1994.1083/production;

http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1083/production/pdf.

Bruner:2016:FAP

[BL16] Marie-Louise Bruner and Martin Lackner. A fast algorithm for permutation pattern matching based on alternating runs. Algorithmica, 75(1):84–117, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-0013-y.

Bertossi:1990:SMW

[BLLP90] A. A. Bertossi, F. Luccio, E. Lodi, and L. Pagli. String matching with weighted errors. Theoretical Computer Science, 73 (3):319–328, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barcelo:2012:ELP

[BLLW12] Pablo Barceló, Leonid Libkin, Anthony W. Lin, and Peter T. Wood. Expressive languages for path queries over graphstructured data. ACM Transactions on Database Systems, 37(4):31:1–31:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Bafna:1994:AAM

[BLP94] V. Bafna, E. L. Lawler, and P. A. Peuzner. Approximation algorithms for multiple sequence alignment. Lecture Notes in Computer Science, 807:43–53, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barton:2018:FAC

[BLP18] Carl Barton, Chang Liu, and Solon P. Pissis. Fast averagecase pattern matching on weighted sequences. International Journal of Foundations of Computer Science (IJFCS), 29(8): 1331–1343, December 2018. ISSN 0129-0541.

Bertossi:1992:SNP

[BLPL92] A. A. Bertossi, F. Luccio, L. Pagli, and E. Lodi. Short notes: A parallel solution to the approximate string matching problem. The Computer Journal, 35(5):524– 526, October 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/35/5/524.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/Volume_35/
Issue_05/350524.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/Volume_35/Issue_05/tiff/524.tif;
 http://www3.oup.co.uk/computer_journal/Volume_35/
Issue_05/tiff/525.tif;   http://www3.oup.co.uk/computer_
journal/Volume_35/Issue_05/tiff/526.tif.

Barcelo:2011:QGP

[BLR11] Pablo Barceló, Leonid Libkin, and Juan L. Reutter. Querying graph patterns. In Lenzerini [Len11], pages 199–210. ISBN 1-4503-0660-8. LCCN ???? URL http://dl.acm.org/

citation.cfm?id=1989284.

Barcelo:2014:QRG

[BLR14] Pablo Barceló, Leonid Libkin, and Juan L. Reutter. Querying regular graph patterns. Journal of the Association for Computing Machinery, 61(1):8:1–8:??, January 2014. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Buneman:1994:CS

[BLS^{+}94] Peter Buneman, Leonid Libkin, Dan Suciu, Val Tannen, and Limsoon Wong. Comprehension syntax. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (1):87–96, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benedikt:2003:DRF

[BLSS03] Michael Benedikt, Leonid Libkin, Thomas Schwentick, and Luc Segoufin. Definable relations and first-order query languages over strings. Journal of the Association for Computing Machinery, 50(5):694–751, September 2003. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:2008:LCL

[Blu08] Richard Blum. Linux command line and shell scripting bible. John Wiley, New York, NY, USA, 2008. ISBN 0-470-25128-X (paperback). xxx + 809 pp. LCCN QA76.76.O63 B598 2008.

URL http://www.loc.gov/catdir/enhancements/fy0827/
2008012238-d.html;   http://www.loc.gov/catdir/enhancements/
fy0827/2008012238-t.html;   http://www.loc.gov/catdir/
enhancements/fy0828/2008012238-b.html.

Boyer:1977:FSS

[BM77] Robert S. Boyer and J. Strother Moore. A fast string searching algorithm. Communications of the Association for Computing Machinery, 20(10):762–772, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [KMP77, Sun90a, BYG92].

Bertossi:2000:RNS

[BM00] A. A. Bertossi and A. Mei. A residue number system on reconfigurable mesh with applications to prefix sums and approximate string matching. IEEE Transactions on Parallel and Distributed Systems, 11(11):1186–1199, November 2000. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183

(electronic). URL http://dlib.computer.org/td/books/
td2000/pdf/l1186.pdf;   http://ieeexplore.ieee.org/
xpl/tocresult.jsp?isnumber=19224;   http://www.computer.
org/tpds/td2000/l1186abs.htm.

Bocker:2008:CAM

[BM08] Sebastian Bocker and Veli Makinen. Combinatorial approaches for mass spectra recalibration. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 5(1): 91–100, January 2008. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Broda:2019:ABR

[BMMR19] Sabine Broda, António Machiavelo, Nelma Moreira, and Rogério Reis. On average behaviour of regular expressions in strong star normal form. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):899–920, September– November 2019. ISSN 0129-0541. URL https://www.

worldscientific.com/doi/10.1142/S0129054119400227.

Bex:2010:ICR

[BNSV10] Geert Jan Bex, Frank Neven, Thomas Schwentick, and Stijn Vansummeren. Inference of concise regular expressions and DTDs. ACM Transactions on Database Systems, 35(2):11:1– 11:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Boyen:2013:MMM

[BNV^{+}13] Peter Boyen, Frank Neven, Dries Van Dyck, Felipe Valentim, and Aalt van Dijk. Mining minimal motif pair sets maximally covering interactions in a protein–protein interaction network. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(1):73–86, January 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Bjorklund:2013:SRP

[BÖ13] Johanna Björklund and Lars-Daniel Öhman. Simulation relations for pattern matching in directed graphs. Theoretical Computer Science, 485(??):1–15, May 13, 2013. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030439751300234X.

Bollinger:2002:UFO

[Bol02] Terry Bollinger. Use of Free and Open-Source Software (FOSS) in the U.S. Department of Defense: Version: 1.2. Mitre report MP 02 W0000101, MITRE Corporation, Octo-

ber 28, 2002. 160 pp. URL http://www.egovos.org/pdf/
dodfoss.pdf.

Bontempi:2007:BSI

[Bon07] Gianluca Bontempi. A blocking strategy to improve gene selection for classification of gene expression data. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 4(2):293–300, April 2007. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Booth:1980:LLC

[Boo80] Kellogg S. Booth. Lexicographically least circular substrings. Information Processing Letters, 10(4–5):240–242, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bowman:1987:PMU

[Bow87] Charles F. Bowman. Pattern matching using finite state machines. Dr. Dobb’s Journal of Software Tools, 12(10):46–??, October 1987. CODEN DDJOEB. ISSN 0888-3076.

Brzozowski:1963:CSM

[BP63] J. A. Brzozowski and J. F. Poage. On the construction of sequential machines from regular expressions. IEEE Transactions on Electronic Computers, EC-12(4):402–403, August 1963. CODEN IEECA8. ISSN 0367-7508. URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
4037916.

Bhargava:2011:CMD

[BP11] Sanjay Bhargava and G. N. Purohit. Construction of a minimal deterministic finite automaton from a regular expression. International Journal of Computer Applications, 15(??):16–27, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume15/number4/1938-2589/.

Bell:2002:SBC

[BPMA02] T. Bell, M. Powell, A. Mukherjee, and D. Adjeroh. Searching BWT compressed text with the Boyer–Moore algorithm and binary search. In Storer and Cohn [SC02], pages 112– 121. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 2002.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=999949. IEEE Computer Society Order Number: PR01477.

Bernardini:2020:APM

[BPPR20] Giulia Bernardini, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Approximate pattern matching on elasticdegenerate text. Theoretical Computer Science, 812(??):109– 122, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397519305018.

Baturo:2009:CSM

[BR09] Paweł Baturo and Wojciech Rytter. Compressed stringmatching in standard Sturmian words. Theoretical Computer Science, 410(30–32):2804–2810, August 20, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baudru:2020:TWT

[BR20] Nicolas Baudru and Pierre-Alain Reynier. From twoway transducers to regular function expressions. International Journal of Foundations of Computer Science (IJFCS), 31(06):843–873, September 2020. ISSN 0129-

0541. URL https://www.worldscientific.com/doi/10.
1142/S0129054120410087.

Bradford:1990:SMB

[Bra90] James H. Bradford. Sequence matching with binary codes. Information Processing Letters, 34(4):193–196, April 24, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brazma:1994:EAL

[Bra94] A. Brazma. Efficient algorithm for learning simple regular expressions from noisy examples. Lecture Notes in Computer Science, 872:260–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brazma:1995:LRE

[Bra95] A. Brazma. Learning of regular expressions by pattern matching. Lecture Notes in Computer Science, 904:392–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1993:SCC

[Bre93] D. Breslauer. Saving comparisons in the Crochemore-Perrin string matching algorithm. Lecture Notes in Computer Science, 726:61–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1994:DMU

[Bre94a] D. Breslauer. Dictionary-matching on unbounded alphabets: Uniform-length dictionaries. Lecture Notes in Computer Science, 807:184–197, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1994:TSS

[Bre94b] Dany Breslauer. Testing string superprimitivity in parallel. Information Processing Letters, 49(5):235–241, March 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1995:FPS

[Bre95] Dany Breslauer. Fast parallel string prefix-matching. Theoretical Computer Science, 137(2):269–278, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1811.

Breslauer:1996:SCC

[Bre96] Dany Breslauer. Saving comparisons in the Crochemore-Perrin string matching algorithm. Theoretical Computer Science, 158(1–2):177–192, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=158&issue=1-
2&aid=2006.

Barcelo:2013:PRE

[BRL13] Pablo Barceló, Juan Reutter, and Leonid Libkin. Parameterized regular expressions and their languages. Theoretical Computer Science, 474(??):21–45, February 25, 2013. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513000029.

Brownlee:1977:ABI

[Bro77] J. Nevil Brownlee. An Algol-based implementation of SNOBOL 4 patterns. Communications of the Association for Computing Machinery, 20(7):527–529, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. ISSAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6–8, 1993, Kiev, Ukraine. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Brachthauser:2016:PFC

[BRO16] Jonathan Immanuel Brachthäuser, Tillmann Rendel, and Klaus Ostermann. Parsing with first-class derivatives. ACM SIGPLAN Notices, 51(10):588–606, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brzozowski:1962:SRE

[Brz62] Janusz A. Brzozowski. A survey of regular expressions and their applications. IRE Transactions on Electronic Computers, EC-11(3):324–335, June 1962. CODEN IRELAO. ISSN 0367-9950.

Brzozowski:1964:RES

[Brz64a] J. A. Brzozowski. Regular expressions from sequential circuits. IEEE Transactions on Electronic Computers, EC-13(6):741– 744, December 1964. CODEN IEECA8. ISSN 0367-7508.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4038308.

Brzozowski:1964:DRE

[Brz64b] Janusz A. Brzozowski. Derivatives of regular expressions. Journal of the Association for Computing Machinery, 11(4): 481–494, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brzozowski:1965:REL

[Brz65] J. A. Brzozowski. Regular expressions for linear sequential circuits. IEEE Transactions on Electronic Computers, EC-14 (2):148–156, April 1965. CODEN IEECA8. ISSN 0367-7508.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4038393.

Berry:1986:RED

[BS86] G. Berry and R. Sethi. From regular expressions to deterministic automata. Theoretical Computer Science, 48(1):117–126, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bentley:1997:FAS

[BS97] Jon L. Bentley and Robert Sedgewick. Fast algorithms for sorting and searching strings. In ACM [ACM97b], pages 360– 369. CODEN PAAAF2. ISBN 0-89871-390-0. LCCN ????

URL http://www.cs.princeton.edu/~rs/strings/paper.
pdf.

Breslav:2007:DPS

[BSM^{+}07] Simon Breslav, Karol Szerszen, Lee Markosian, Pascal Barla, and Joëlle Thollot. Dynamic 2D patterns for shading 3D scenes. ACM Transactions on Graphics, 26(3):20:1–20:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baba:2003:NRA

[BST^{+}03] K. Baba, A. Shinohara, M. Takeda, S. Inenaga, and S. Arikawa. A note on randomized algorithm for string matching with mismatches. Nordic Journal of Computing, 10(1): 2–??, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Barsky:2008:GAT

[BSTU08] Marina Barsky, Ulrike Stege, Alex Thomo, and Chris Upton. A graph approach to the threshold all-against-all substring matching problem. ACM Journal of Experimental Algorithmics, 12:1.10:1–1.10:??, June 2008. CODEN ???? ISSN 1084-6654.

Bakar:2000:ERE

[BSY00] Zainab Abu Bakar, Tengku Mohd T. Sembok, and Mohammed Yusoff. An evaluation of retrieval effectiveness using spelling-correction and string-similarity matching methods on Malay texts. Journal of the American Society for Information Science, 51(8):691–706, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borsotti:2021:EPS

[BT21] Angelo Borsotti and Ulya Trofimovich. Efficient POSIX submatch extraction on nondeterministic finite automata. Software — Practice and Experience, 51(2):159–192, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brodie:2006:SAH

[BTC06] Benjamin C. Brodie, David E. Taylor, and Ron K. Cytron. A scalable architecture for high-throughput regular-expression pattern matching. ACM SIGARCH Computer Architecture News, 34(2):191–202, 2006. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Barrero:1983:RLT

[BTG83] A. Barrero, M. G. Thomason, and R. C. Gonzalez. Regularlike tree expressions. International Journal of Computer and Information Sciences, 12(1):1–11, February 1983. CODEN IJCIAH. ISSN 0091-7036.

Boyer:2002:LDS

[BTTF02] John Boyer, Andrew D. Todd, Jason Trenough, and Doug Farrell. Letters: Defective sign-and-encrypt and healthcare woes and J2EE cache and pool and Regex++. Dr. Dobb’s Journal of Software Tools, 27(2):10, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Bundy:1994:ADC

[Bun94] Alan Bundy, editor. Automated deduction, CADE-12: 12th International Conference on Automated Deduction, Nancy, France, June 26–July 1, 1994: proceedings, number 814 in Lecture Notes in Computer Science 1994. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Bunke:1995:FAM

[Bun95] H. Bunke. Fast approximate matching of words against a dictionary. Computing: Archiv für informatik und numerik, 55(1):75–89, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/springer.py?Page=10&Key=362&cat=
300607/tocs/springer.py?Page=47&Key=340&cat=3&id_abstract=
265&id_volume=21&id_journal=8.

Burkowski:1982:HHS

[Bur82] Forbes J. Burkowski. A hardware hashing scheme in the design of a multiterm string comparator. IEEE Transactions on Computers, C-31(9):825–834, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1676098. See correction [Bur84].

Burkowski:1984:CHH

[Bur84] F. J. Burkowski. Correction to “A Hardware Hashing Scheme in the Design of a Multiterm String Comparator”. IEEE Transactions on Computers, C-33(4):375, April 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1676447. See [Bur82].

Burton:2011:SBL

[Bur11] Benjamin A. Burton. Searching a bitstream in linear time for the longest substring of any given density. Algorithmica, 61(3):555–579, November 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=3&spage=555.

Berglund:2017:SRE

[BvdM17] Martin Berglund and Brink van der Merwe. On the semantics of regular expression parsing in the wild. Theoretical Computer Science, 679(??):69–82, May 30, 2017. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397516304790.

Berglund:2023:RER

[BvdM23] Martin Berglund and Brink van der Merwe. Re-examining regular expressions with backreferences. Theoretical Computer Science, 940 (part A)(??):66–80, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304397522006570.

Broy:1991:MPS

[BW91] Manfred Broy and Martin Wirsing. Methods of programming. Selected papers on the CIP-Project, volume 544. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNCSD9. ISBN 3-540-54576-X (Berlin), 0-387-54576-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 268 pp. LCCN QA76.6 .M4543 1991,

QA267.A1 L43 no.544. URL http://link.springer-ny.
com/link/service/series/0558/tocs/t0544.htm;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=544.

Bulling:2012:MRR

[BWG12] Andreas Bulling, Jamie A. Ward, and Hans Gellersen. Multimodal recognition of reading activity in transit using bodyworn sensors. ACM Transactions on Applied Perception, 9(1): 2:1–2:??, March 2012. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Braunstein:2008:AB

[BWN08] Roger Braunstein, Mims H. Wright, and Joshua J. Noble. ActionScript 3.0 bible. John Wiley, New York, NY, USA, 2008. ISBN 0-470-13560-3 (paperback). lv + 735 pp. LCCN TR897.7 .B793 2008. URL http://

www.loc.gov/catdir/enhancements/fy0827/2007032141-
b.html;   http://www.loc.gov/catdir/enhancements/fy0827/
2007032141-d.html;   http://www.loc.gov/catdir/enhancements/
fy0827/2007032141-t.html.

Baeza-Yates:1989:ISS

[BY89] Ricardo A. Baeza-Yates. Improved string searching. Software — Practice and Experience, 19(3):257–271, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boyer:1991:ACP

[BY91] Robert S. Boyer and Yuan Yu. AUTOMATED CORRECT-NESS PROOFS OF MACHINE CODE PROGRAMS FOR A COMMERCIAL MICROPROCESSOR. Technical Report TR-91-33, University of Texas, Austin, Austin, TX, USA, November 1991. 15 pp. prize (\ $1.50).

Boyer:1992:ACP

[BY92] R. S. Boyer and Yuan Yu. Automated correctness proofs of machine code programs for a commercial microprocessor. In Kapur [Kap92], pages 416–430. ISBN 3-540-55602-8. LCCN QA76.9.A96I57 1992.

Baeza-Yates:1996:UVS

[BY96a] R. Baeza-Yates. A unified view to string matching algorithms. Lecture Notes in Computer Science, 1175:1–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boyer:1996:APO

[BY96b] Robert S. Boyer and Yuan Yu. Automated proofs of object code for a widely used microprocessor. Journal of the Association for Computing Machinery, 43(1):166–192, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Benachour:2024:GAR

[BYB^{+}24] Amira Benachour, Saïd Yahiaoui, Sarra Bouhenni, Hamamache Kheddouci, and Nadia Nouali-Taboudjemat. GPUaccelerated relaxed graph pattern matching algorithms. The Journal of Supercomputing, 80(15):21811–21836, October 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-024-06283-7.

Baeza-Yates:2009:SBM

[BYBDS09] Ricardo Baeza-Yates, Véronique Bruyère, Olivier Delgrange, and Rodrigo Scheihing. On the size of Boyer–Moore automata. Theoretical Computer Science, 410(43):4432–4443, October 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:2003:CPM

[BYCC03] R. Baeza-Yates, Edgar Chávez, and Maxime Crochemore, editors. Combinatorial pattern matching: 14th annual symposium, CPM 2003, Morelia, Michoacan, Mexico, June 25– 27, 2003: Proceedings, volume 2676 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40311-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

2003. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2676.htm;   http://www.springerlink.
com/content/978-3-540-40311-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=

2676. Also available via the World Wide Web.

Baeza-Yates:1994:BMA

[BYCG94] Ricardo A. Baeza-Yates, Christian Choffrut, and Gaston H. Gonnet. On Boyer–Moore automata. Algorithmica, 12(4–5): 268–292, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baeza-Yates:1994:PMU

[BYCMW94] R. Baeza-Yates, W. Cunto, U. Manber, and S. Wu. Proximity matching using fixed-queries trees. Lecture Notes in Computer Science, 807:198–212, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1996:FAS

[BYF96] Ricardo A. Baeza-Yates and Luis O. Fuentes. A framework to animate string algorithms. Information Processing Letters, 59(5):241–244, September 9, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1992:NAT

[BYG92] Ricardo Baeza-Yates and Gaston H. Gonnet. A new approach to text searching. Communications of the Association for Computing Machinery, 35(10):74–82, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/

0001-0782/135243.html. This paper describes a new lineartime string search algorithm that can handle limited regularexpression pattern matching without backtracking. See also [KMP77], [BM77], [KR81a], [Sun90a], and [WM92a].

Baeza-Yates:1996:FTS

[BYG96] Richardo A. Baeza-Yates and Gaston H. Gonnet. Fast text searching for regular expressions or automaton searching on tries. Journal of the Association for Computing Machinery, 43(6):915–936, November 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/jacm/235810.html.

Bontupalli:2018:EMB

[BYHT18] Venkataramesh Bontupalli, Chris Yakopcic, Raqibul Hasan, and Tarek M. Taha. Efficient memristor-based architecture for intrusion detection and high-speed packet classification. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):41:1–41:??, December 2018. CODEN ????

ISSN 1550-4832. URL https://dl.acm.org/ft_gateway.
cfm?id=3264819.

Bouhenni:2022:EPE

[BYK22] Sarra Bouhenni, Saïd Yahiaoui, and Hamamache Kheddouci. Efficient parallel edge-centric approach for relaxed graph pattern matching. The Journal of Supercomputing, 78(2): 1642–1671, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-021-03938-7.

Baeza-Yates:1992:TCN

[BYKZ^{+}92] Ricardo Baeza-Yates, Fred T. Krogh, Bernard Ziegler, Peter R. Sibbald, and Daniel M. Sunday. Technical correspondence: Notes on a very fast substring search algorithm. Communications of the Association for Computing Machinery, 35 (4):132–137, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baeza-Yates:1996:FAA

[BYN96] R. Baeza-Yates and G. Navarro. A faster algorithm for approximate string matching. Lecture Notes in Computer Science, 1075:1–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1997:MAS

[BYN97] R. Baeza-Yates and G. Navarro. Multiple approximate string matching. Lecture Notes in Computer Science, 1272:174–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1998:FTD

[BYN98] Ricardo Baeza-Yates and Gonzalo Navarro. Fast twodimensional approximate pattern matching. Lecture Notes in Computer Science, 1380:341–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1380/13800341.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1380/13800341.pdf.

Baeza-Yates:1999:FAS

[BYN99] Ricardo A. Baeza-Yates and Gonzalo Navarro. Faster approximate string matching. Algorithmica, 23(2):127–158, February 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.de/link/
service/journals/00453/bibs/23n2p127.html;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=23&issue=2&spage=127.

Bouhenni:2021:SDG

[BYNTK21] Sarra Bouhenni, Saïd Yahiaoui, Nadia Nouali-Taboudjemat, and Hamamache Kheddouci. A survey on distributed graph pattern matching in massive graphs. ACM Computing Surveys, 54(2):36:1–36:35, April 2021. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.

acm.org/doi/10.1145/3439724.

Baeza-Yates:1992:FPA

[BYP92] Ricardo A. Baeza-Yates and Chris H. Perleberg. Fast and practical approximate string matching. Lecture Notes in Computer Science, 644:182–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1996:FPA

[BYP96] Ricardo A. Baeza-Yates and Chris H. Perleberg. Fast and practical approximate string matching. Information Processing Letters, 59(1):21–27, July 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1992:ART

[BYR92] R. A. Baeza-Yates and M. Regnier. Average running time of the Boyer–Moore–Horspool algorithm. Theoretical Computer Science, 92(1):19–31, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:1993:FTD

[BYR93a] Ricardo Baeza-Yates and Mireille Regnier. Fast twodimensional pattern matching. Information Processing Letters, 45(1):51–57, January 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1993:FTP

[BYR93b] Ricardo Baeza-Yates and Mireille Regnier. Fast twodimensional pattern matching. Information Processing Letters, 45(1):51–57, January 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bunke:1998:ADS

[BZ98] H. Bunke and M. Zumbuhl. Acquisition of 2-D shape models from scenes with overlapping objects using string matching. Lecture Notes in Computer Science, 1451:200–??, 1998. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cockx:2018:EDC

[CA18] Jesper Cockx and Andreas Abel. Elaborating dependent (co)pattern matching. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):75:1–75:30, July 2018.

URL https://dl.acm.org/doi/abs/10.1145/3236770.

Cockx:2020:EDC

[CA20] Jesper Cockx and Andreas Abel. Elaborating dependent (co)pattern matching: No pattern left behind. Journal of Functional Programming, 30(??):e2, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic).  URL https://www.cambridge.org/core/
journals/journal-of-functional-programming/article/
elaborating-dependent-copattern-matching-no-pattern-
left-behind/F13CECDAB2B6200135D45452CA44A8B3.

C:2018:SSS

[CADA18] Paul Suganthan G. C., Adel Ardalan, AnHai Doan, and Aditya Akella. Smurf: self-service string matching using random forests. Proceedings of the VLDB Endowment, 12(3): 278–291, November 2018. CODEN ???? ISSN 2150-8097.

Calsavara:2000:JQH

[Cal00] Alexandre Pereira Calsavara. Java Q&A: How can I extend Java’s search capabilities? Dr. Dobb’s Journal of Software Tools, 25(12):141–142, 144, 146, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

ftp/2000/2000_12/jqa0012.zip.

Cameron:1999:RXS

[Cam99] Robert D. Cameron. REX: XML shallow parsing with regular expressions. Markup languages: theory & practice, 1(3):61–88, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Carr:1977:CSM

[Car77] R. G. Carr. Character string manipulation in the C language. Technical Memorandum 1208, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1977. ?? pp.

Chen:2023:DES

[CBD^{+}23] Qiaochu Chen, Arko Banerjee, Çağatay Demiralp, Greg Durrett, and Işıl Dillig. Data extraction via semantic regular expression synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA2):287:1–287:??, October 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.1145/3622863.

Clarke:1997:URE

[CC97] Charles L. A. Clarke and Gordon V. Cormack. On the use of regular expressions for searching text. ACM Transactions on Programming Languages and Systems, 19(3):413–426, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593

(electronic). URL http://www.acm.org/pubs/citations/
journals/toplas/1997-19-3/p413-clarke/.

Cantone:2013:ESM

[CCF13] Domenico Cantone, Salvatore Cristofaro, and Simone Faro. Efficient string-matching allowing for non-overlapping inversions. Theoretical Computer Science, 483(??):85–95, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397512005452.

Campanelli:2012:PMS

[CCFG12] Matteo Campanelli, Domenico Cantone, Simone Faro, and Emanuele Giaquinta. Pattern matching with swaps in practice. International Journal of Foundations of Computer Science (IJFCS), 23(2):323–342, February 2012. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cole:1993:OFP

[CCG^{+}93] R. Cole, M. Crochemore, Z. Galil, L. Gasieniec, R. Eariharan, S. Muthukrishnan, K. Park, and W. Rytter. Optimally fast parallel algorithms for preprocessing and pattern matching in one and two dimensions. In IEEE [IEE93], pages 248–258. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Crochemore:1994:STS

[CCG^{+}94] Maxime Crochemore, Artur Czumaj, Leszek Gasieniec, Stefan Jarominek, Thierry Lecroq, Wojciech Plandowski, and Wojciech Rytter. Speeding up two string-matching algorithms. Algorithmica, 12(4–5):247–267, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Crochemore:1999:FPM

[CCG^{+}99] Maxime Crochemore, A. Czumaj, L. Gsaieniec, T. Lecroq, W. Plandowski, and W. Rytter. Fast practical multipattern matching. Information Processing Letters, 71(3– 4):107–113, August 27, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.elsevier.nl/gej-ng/10/23/20/49/28/26/abstract.
html;   http://www.elsevier.nl/gej-ng/10/23/20/49/
28/26/article.pdf.

Chang:2009:HTF

[CCH09] Ye-In Chang, Jiun-Rung Chen, and Min-Tze Hsu. A hash trie filter approach to approximate string matching for genomic databases. Lecture Notes in Computer Science, 5579: 816–825, 2009. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-642-02568-6_83.

Chen:2023:SSC

[CCH^{+}23] Yu-Fang Chen, David Chocholatý, Vojtěch Havlena, Lukáš Holík, Ondřej Lengál, and Juraj Síč. Solving string constraints with lengths by stabilization. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA2):296:1– 296:??, October 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3622872.

Christou:2013:ESC

[CCI^{+}13] M. Christou, M. Crochemore, C. S. Iliopoulos, M. Kubica, S. P. Pissis, J. Radoszewski, W. Rytter, B. Szreder, and T. Waleń. Efficient seed computation revisited. Theoretical Computer Science, 483(??):171–181, April 29, 2013. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397511010632.

Coulson:1987:PNA

[CCL87] A. F. W. Coulson, J. F. Collins, and A. Lyall. Protein and nucleic acid sequence database searching: a suitable case for parallel processing. The Computer Journal, 30(5):420–424, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/420.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/420.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/421.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/422.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/423.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/424.tif.

Consel:1989:PEP

[CD89] Charles Consel and Olivier Danvy. Partial evaluation of pattern matching in strings. Information Processing Letters, 30 (2):79–86, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colussi:1996:TSE

[CD96] Livio Colussi and Alessia De Col. A time and space efficient data structure for string searching on large texts. Information Processing Letters, 58(5):217–222, June 10, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cockx:2018:PRU

[CD18] Jesper Cockx and Dominique Devriese. Proof-relevant unification: Dependent pattern matching with only the axioms of your type theory. Journal of Functional Programming, 28:e12, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/journals/journal-of-functional-
programming/article/proofrelevant-unification-dependent-
pattern-matching-with-only-the-axioms-of-your-type-
theory/E54D56DC3F5D5361CCDECA824030C38E.

Cheng:1996:FHR

[CDC96] J.-M. Cheng, L. M. Duyanovich, and D. J. Craft. A fast, highly reliable data compression chip and algorithm for storage systems. IBM Journal of Research and Development, 40(6):603–613, November 1996. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

almaden.ibm.com/journal/rd40-6.html#two.

Conficconi:2023:EED

[CDC^{+}23] Davide Conficconi, Emanuele Del Sozzo, Filippo Carloni, Alessandro Comodi, Alberto Scolari, and Marco Domenico Santambrogio. An energy-efficient domain-specific architecture for regular expressions. IEEE Transactions on Emerging Topics in Computing, 11(1):3–17, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2005:ESM

[CDDM05] S. Chen, S. Diggavi, S. Dusad, and S. Muthukrishnan. Efficient string matching algorithms for combinatorial universal denoising. In Storer and Cohn [SC05], pages 153– 162. ISBN 0-7695-2309-9. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402176.

Chew:1995:GPM

[CDEK95] L. P. Chew, D. Dor, A. Efrat, and K. Kedem. Geometric pattern matching in d-dimensional space. Lecture Notes in Computer Science, 979:264–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Champarnaud:2015:TSD

[CDJM15] Jean-Marc Champarnaud, Jean-Philippe Dubernard, Hadrien Jeanne, and Ludovic Mignot. Two-sided derivatives for regular expressions and for hairpin expressions. Fundamenta Informaticae, 137(4):425–455, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Corradini:1995:FAM

[CDL95] F. Corradini, R. De Nicola, and A. Labella. Fully abstract models for nondeterministic regular expressions. Lecture Notes in Computer Science, 962:130–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Corradini:1999:MNR

[CDL99] Flavio Corradini, Rocco De Nicola, and Anna Labella. Models of nondeterministic regular expressions. Journal of Computer and System Sciences, 59(3):412–449, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000099916367.

Calvanese:2008:CQC

[CDL08] Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Conjunctive query containment and answering under description logic constraints. ACM Transactions on Computational Logic, 9(3):22:1–22:??, June 2008. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Cochran:2015:PBP

[CDL^{+}15] Robert A. Cochran, Loris D’Antoni, Benjamin Livshits, David Molnar, and Margus Veanes. Program boosting: Program synthesis via crowd-sourcing. ACM SIGPLAN Notices, 50(1): 677–688, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Czerwinski:2017:DDD

[CDLM17] Wojciech Czerwiński, Claire David, Katja Losemann, and Wim Martens. Deciding definability by deterministic regular expressions. Journal of Computer and System Sciences, 88(??):75–89, September 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000017300405.

Calvanese:1999:RRE

[CDLV99] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. Rewriting of regular expressions and regular path queries. In ACM [ACM99a], pages 194–204. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296 1999. URL

http://www.acm.org/pubs/articles/proceedings/pods/
303976/p194-calvanese/p194-calvanese.pdf;   http:/
/www.acm.org/pubs/citations/proceedings/pods/303976/

p194-calvanese/. ACM order number 475990.

Calvanese:2002:RRE

[CDLV02] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. Rewriting of regular expressions and regular path queries. Journal of Computer and System Sciences, 64(3):443–465, May 2002. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000001918057.

Caron:2011:PMB

[CDM11] Eddy Caron, Frédéric Desprez, and Adrian Muresan. Pattern matching based forecast of non-periodic repetitive behavior for cloud clients. Journal of Grid Computing, 9(1):49–64, March 2011. CODEN ???? ISSN 1570-7873 (print), 1572-

9184 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1570-7873&volume=9&
issue=1&spage=49.

Cockx:2014:PMK

[CDP14] Jesper Cockx, Dominique Devriese, and Frank Piessens. Pattern matching without K. ACM SIGPLAN Notices, 49(9): 257–268, September 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cockx:2016:EDP

[CDP16a] Jesper Cockx, Dominique Devriese, and Frank Piessens. Eliminating dependent pattern matching without K. Journal of Functional Programming, 26:e16, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic).  URL https://www.cambridge.org/core/
journals/journal-of-functional-programming/article/
eliminating-dependent-pattern-matching-without-k/4BC4EA2D02D801E5ABED264FE5FB177A.

Cockx:2016:UEP

[CDP16b] Jesper Cockx, Dominique Devriese, and Frank Piessens. Unifiers as equivalences: proof-relevant unification of dependently typed data. ACM SIGPLAN Notices, 51(9):270–283, September 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cotumaccio:2023:CLO

[CDPP23] Nicola Cotumaccio, Giovanna D’Agostino, Alberto Policriti, and Nicola Prezza. Co-lexicographically ordering automata and regular languages — part I. Journal of the Association for Computing Machinery, 70(4):27:1–27:??, August 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3607471.

Cringean:1991:NNS

[CEMW91] Janey K. Cringean, Roger England, Gordon A. Manson, and Peter Willett. Nearest-neighbour searching in files of text signatures using transputer networks. Electronic Publishing— Origination, Dissemination, and Design, 4(4):185–203, December 1991. CODEN EPODEU. ISSN 0894-3982.

Clifford:2010:PMD

[CEPR10] Raphaël Clifford, Klim Efremenko, Ely Porat, and Amir Rothschild. Pattern matching with don’t cares and few errors. Journal of Computer and System Sciences, 76(2):115–124, March 2010. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022000009000567.

Copi:1958:REL

[CEW58] Irving M. Copi, Calvin C. Elgot, and Jesse B. Wright. Realization of events by logical nets. Journal of the Association for Computing Machinery, 5(2):181–196, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cheng:1985:APF

[CF85] H. D. Cheng and K. S. Fu. Algorithm partition for a fixed-size VLSI architecture using space-time domain expansion. In Hwang [Hwa85], pages 126–132. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985. URL http://

www.acsel-lab.com/arithmetic/arith7/papers/ARITH7_

Cheng_Fu.pdf. IEEE catalog number 85CH2146-9. IEEE Computer Society order number 632.

Casanova:1988:SPM

[CF88] Marco A. Casanova and Antonio L. Furtado. String patternmatching in Prolog. Computer Languages, 13(3-4):149–170, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cantone:2006:SEB

[CF06] Domenico Cantone and Simone Faro. A space efficient bitparallel algorithm for the multiple string matching problem. International Journal of Foundations of Computer Science (IJFCS), 17(6):1235–1251, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2022:EBS

[CFCK22] Yue Chen, Kaiyu Feng, Gao Cong, and Han Mao Kiah. Example-based spatial pattern matching. Proceedings of the VLDB Endowment, 15(11):2572–2584, July 2022. CODEN

???? ISSN 2150-8097. URL https://dl.acm.org/doi/10.
14778/3551793.3551815.

Cantone:2012:ABM

[CFG12] Domenico Cantone, Simone Faro, and Emanuele Giaquinta. Adapting Boyer–Moore-like algorithms for searching Huffman encoded texts. International Journal of Foundations of Computer Science (IJFCS), 23(2):343–356, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cirstea:2007:CEC

[CFK07] Horatiu Cirstea, Germain Faure, and Claude Kirchner. A ρcalculus of explicit constraint application. Higher-Order and Symbolic Computation, 20(1–2):37–72, June 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1388-3690&volume=20&issue=1&spage=37.

Chhabra:2017:EOP

[CFKT17] Tamanna Chhabra, Simone Faro, M. Oğuzhan Külekci, and Jorma Tarhio. Engineering order-preserving pattern matching with SIMD parallelism. Software — Practice and Experience, 47(5):731–739, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2022:SSC

[CFLH^{+}22] Taolue Chen, Alejandro Flores-Lamas, Matthew Hague, Zhilei Han, Denghang Hu, Shuanglong Kan, Anthony W. Lin, Philipp Rümmer, and Zhilin Wu. Solving string constraints with regex-dependent functions through transducers with priorities and variables. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):45:1–45:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https://

dl.acm.org/doi/10.1145/3498707.

Ciancarini:2000:UCL

[CFM00] P. Ciancarini, F. Franzé, and C. Mascolo. Using a coordination language to specify and analyze systems containing mobile components. ACM Transactions on Software Engineering and Methodology, 9(2):167–198, April 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392

(electronic). URL http://www.acm.org/pubs/articles/
journals/tosem/2000-9-2/p167-ciancarini/p167-ciancarini.
pdf;   http://www.acm.org/pubs/citations/journals/
tosem/2000-9-2/p167-ciancarini/.

Cao:2017:VNM

[CFM17] Yang Cao, Wenfei Fan, and Shuai Ma. Virtual network mapping in cloud computing: a graph pattern matching approach. The Computer Journal, 60(3):60–??, March 2017. CODEN CMPJA6. ISSN ???? URL https:/

/academic.oup.com/comjnl/article/60/3/287/2608063/
Virtual-Network-Mapping-in-Cloud-Computing-A-Graph.

Cantone:2019:LES

[CFP19] Domenico Cantone, Simone Faro, and Arianna Pavone. Linear and efficient string matching algorithms based on weak factor recognition. ACM Journal of Experimental Algorithmics, 24(1):1.8:1–1.8:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3301295.

Cieslak:1989:PNP

[CFS^{+}89] R. Cieslak, A. Fawaz, S. Sachs, P. Varaiya, J. Walrand, and A. Li. The programmable network prototyping system. Computer, 22(5):67–76, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cowan:1979:HKRa

[CG79a] Richard M. Cowan and Martin L. Griss. Hashing — the key to rapid pattern matching. In Ng [Ng79], pages 266–278. ISBN 0-387-09519-5. LCCN QA155.7.E4 I57 1979.

Cowan:1979:HKRb

[CG79b] Richard M. Cowan and Martin L. Griss. Hashing — the key to rapid pattern matching. In Proc. EUROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 266–278. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

Carpenter:1987:MPA

[CG87] Gail A. Carpenter and Stephen Grossberg. A massively parallel architecture for a self-organizing neural pattern recognition machine. Computer Vision, Graphics, and Image Processing, 37(1):54–115, January 1987. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic). CARPENTER87.

Chlebus:1994:OPM

[CG94a] B. S. Chlebus and L. Gasieniec. Optimal pattern matching on meshes. Lecture Notes in Computer Science, 775:213–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1994:CPM

[CG94b] Maxime Crochemore and Dan Gusfield, editors. Combinatorial pattern matching: 5th annual symposium, CPM 94, Asilomar, CA, USA, June 5–8, 1994: proceedings, volume 807 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-58094-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

1994. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t0807.htm;   http://www.springerlink.
com/content/978-0-387-58094-4;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
807.

Colussi:1990:ECS

[CGG90] L. Colussi, Z. Galil, and R. Giancarlo. On the exact complexity of string matching. In IEEE [IEE90], pages 135–144. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Crochemore:1997:CTR

[CGG^{+}97] Maxime Crochemore, Zvi Galil, Leszek Ga̧sieniec, Kunsoo Park, and Wojciech Rytter. Constant-time randomized parallel string matching. SIAM Journal on Computing, 26(4): 950–960, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/28007.

Crochemore:1998:CTO

[CGH^{+}98] Maxime Crochemore, Leszek Gasieniec, Ramesh Hariharan, S. Muthukrishnan, and Wojciech Rytter. A constant time optimal parallel algorithm for two-dimensional pattern matching. SIAM Journal on Computing, 27(3):668–681, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/28006.

Cabello:2008:PCD

[CGK08] Sergio Cabello, Panos Giannopoulos, and Christian Knauer. On the parameterized complexity of d-dimensional point set pattern matching. Information Processing Letters, 105(2):73– 77, January 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:2006:JJTa

[CGM06] Tal Cohen, Joseph (Yossi) Gil, and Itay Maman. JTL: the Java tools language. ACM SIGPLAN Notices, 41(10):89–108, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandramouli:2010:HPD

[CGM10] Badrish Chandramouli, Jonathan Goldstein, and David Maier. High-performance dynamic pattern matching over disordered streams. Proceedings of the VLDB Endowment, 3(1– 2):220–231, September 2010. CODEN ???? ISSN 2150-8097.

Cadar:2008:EAG

[CGP^{+}08] Cristian Cadar, Vijay Ganesh, Peter M. Pawlowski, David L. Dill, and Dawson R. Engler. EXE: Automatically generating inputs of death. ACM Transactions on Information and System Security, 12(2):10:1–10:??, December 2008. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Crochemore:1995:TDP

[CGPR95] M. Crochemore, L. Gasieniec, W. Plandowski, and W. Rytter. Two-dimensional pattern matching in linear time and small space. Lecture Notes in Computer Science, 900:181–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colazzo:2013:EAI

[CGPS13a] D. Colazzo, G. Ghelli, L. Pardini, and C. Sartiani. Efficient asymmetric inclusion of regular expressions with interleaving and counting for XML type-checking. Theoretical Computer Science, 492(??):88–116, June 24, 2013. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513003319.

Colazzo:2013:ALI

[CGPS13b] Dario Colazzo, Giorgio Ghelli, Luca Pardini, and Carlo Sartiani. Almost-linear inclusion for XML regular expression types. ACM Transactions on Database Systems, 38(3):15:1– 15:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Crochemore:1993:TDP

[CGR93] Maxime Crochemore, Leszek Gasieniec, and Wojciech Rytter. Two-dimensional pattern matching by sampling. Information Processing Letters, 46(4):159–162, June 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1999:CSS

[CGR99] Maxime Crochemore, Leszek Gaşieniec, and Wojciech Rytter. Constant-space string-matching in sublinear average time. Theoretical Computer Science, 218(1):197–203, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
tcs/sub/1999/218/1/3068.pdf.

Chan:2002:RTE

[CGR02] Chee Yong Chan, Minos N. Garofalakis, and Rajeev Rastogi. RE-Tree: An efficient index structure for regular expressions. In Bernstein et al. [B^{+}02], pages 263–274. ISBN 1-55860-

869-9. LCCN ???? URL http://www.vldb.org/conf/2002/
S08P02.pdf.

Chan:2003:RTE

[CGR03] Chee-Yong Chan, Minos Garofalakis, and Rajeev Rastogi. RE-tree: an efficient index structure for regular expressions. VLDB Journal: Very Large Data Bases, 12(2):102–119, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Colazzo:2017:LTM

[CGS17] Dario Colazzo, Giorgio Ghelli, and Carlo Sartiani. Linear time membership in a class of regular expressions with counting, interleaving, and unordered concatenation. ACM Transactions on Database Systems, 42(4):24:1–24:??, November 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Chhabra:2025:SSM

[CGT25] Tamanna Chhabra, Sukhpal Singh Ghuman, and Jorma Tarhio. String searching with mismatches using AVX2 and AVX-512 instructions. Information Processing Letters, 189(??):??, March 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019025000018.

Chen:2013:TCF

[CGZ^{+}13] Yang Chen, Alex Groce, Chaoqiang Zhang, Weng-Keen Wong, Xiaoli Fern, Eric Eide, and John Regehr. Taming compiler fuzzers. ACM SIGPLAN Notices, 48(6):197–208, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cole:1992:TBE

[CH92] R. Cole and R. Hariharan. Tighter bounds on the exact complexity of string matching. In IEEE [IEE92], pages 600–609. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Cole:1997:TUB

[CH97a] Richard Cole and Ramesh Hariharan. Tighter upper bounds on the exact complexity of string matching. SIAM Journal on Computing, 26(3):803–856, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/24694.

Cole:1997:TPM

[CH97b] Richard Cole and Ramesh Hariharan. Tree pattern matching and subset matching in randomized O(n log^{3} m) time.

In ACM [ACM97c], pages 66–75. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997. URL http://www.

acm.org/pubs/articles/proceedings/stoc/258533/p66-
cole/p66-cole.pdf;   http://www.acm.org/pubs/citations/
proceedings/stoc/258533/p66-cole/. ACM order no.

508970.

Cole:2002:ASM

[CH02] Richard Cole and Ramesh Hariharan. Approximate string matching: a simpler faster algorithm. SIAM Journal on Computing, 31(6):1761–1782, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/37052.

Cole:2003:TPM

[CH03] Richard Cole and Ramesh Hariharan. Tree pattern matching to subset matching in linear time. SIAM Journal on Computing, 32(4):1056–1066, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/38270.

Choffrut:2004:SMO

[CH04] Ch. Choffrut and Y. Haddad. String-matching with OB-DDs. Theoretical Computer Science, 320(2–3):187–198, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Char:1986:PSS

[Cha86] Bruce W. Char, editor. Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. ACM order no. 505860.

Chase:1987:IBT

[Cha87] D. R. Chase. An improvement to bottom-up tree pattern matching. In ACM [ACM87], pages 168–177. ISBN

???? LCCN ???? URL http://www.acm.org:80/pubs/
citations/proceedings/plan/41625/p168-chase/.

Chapman:1991:QSS

[Cha91] Rob Chapman. QuikFind string search. Forth Dimensions, 13(4):21–??, November 1, 1991. CODEN FODMD5. ISSN 0884-0822.

Chang:1993:SPMb

[Cha93a] Daniel K. Chang. A string pattern—matching algorithm. The Journal of Systems and Software, 22(3):207–216, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1993:SPMa

[Cha93b] Daniel Kuo-Yee Chang. String pattern matching and lossless data compression. Ph.D. thesis, City University of New York, New York, NY, USA, 1993. 110 pp. URL http://search.

proquest.com/docview/304028781.

Chao:1994:CAS

[Cha94] K.-M. Chao. Computing all suboptimal alignments in linear space. Lecture Notes in Computer Science, 807:31–42, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Champarnaud:2001:ISI

[Cha01] Jean-Marc Champarnaud. Implicit structures to implement NFA’s from regular expressions. Lecture Notes in Computer Science, 2088:80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2088/
20880080.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2088/20880080.pdf.

Champarnaud:2002:ETI

[Cha02a] J.-M. Champarnaud. Evaluation of three implicit structures to implement nondeterministic automata from regular expressions. International Journal of Foundations of Computer Science (IJFCS), 13(1):99–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chauve:2002:TPMa

[Cha02b] Cedric Chauve. Tree pattern matching for linear static terms. Lecture Notes in Computer Science, 2476:160–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.de/link/service/
series/0558/bibs/2476/24760160.htm;   http://link.
springer.de/link/service/series/0558/papers/2476/24760160.
pdf.

Chauve:2002:TPMb

[Cha02c] Cedric Chauve. Tree pattern matching with a more general notion of occurrence of the pattern. Information Processing Letters, 82(4):197–201, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1996:PCM

[Che96] Zhi-Zhong Chen. Parallel constructions of maximal path sets and applications to short superstrings. Theoretical Computer Science, 161(1–2):1–21, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2051.

Cheney:2008:FFU

[Che08] James Cheney. FLUX: functional updates for XML. ACM SIGPLAN Notices, 43(9):3–14, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chivers:2017:OUR

[Chi17] Howard Chivers. Optimising Unicode regular expression evaluation with previews. Software — Practice and Experience, 47(5):669–688, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chlipala:2008:PHO

[Chl08] Adam Chlipala. Parametric higher-order abstract syntax for mechanized semantics. ACM SIGPLAN Notices, 43(9):143– 156, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2014:BPA

[CHL14] Kuei-Hao Chen, Guan-Shieng Huang, and Richard Chia-Tung Lee. Bit-parallel algorithms for exact circular string matching. The Computer Journal, 57(5):731–743, May 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/57/5/
731.full.pdf+html.

Chen:2025:IAS

[CHL^{+}25] Haiming Chen, Hong Huang, Rongchen Li, Chengyao Peng, and Weihao Su. Incremental algorithms for solving regular expression intersection non-emptiness. Theoretical Computer Science, 1043(??):??, July 30, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304397525001872.

Chan:2007:CID

[CHLS07] Ho-Leung Chan, Wing-Kai Hon, Tak-Wah Lam, and Kunihiko Sadakane. Compressed indexes for dynamic text collections. ACM Transactions on Algorithms, 3(2):21:1–21:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cole:2014:TDP

[CHLT14] Richard Cole, Carmit Hazay, Moshe Lewenstein, and Dekel Tsur. Two-dimensional parameterized matching. ACM Transactions on Algorithms, 11(2):12:1–12:??, October 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Choueka:1978:FAD

[Cho78] Yaacov Choueka. Finite automata, definable sets, and regular expressions over ω^{n}-tapes. Journal of Computer and System Sciences, 17(1):81–97, August 1978. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0022000078900363.

Chia-Hsiang:1992:RED

[CHP92] Chang Chia-Hsiang and Robert Paige. From regular expressions to DFA’s using compressed NFA’s. Lecture Notes in Computer Science, 644:88–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cole:1995:TLB

[CHPZ95] Richard Cole, Ramesh Hariharan, Mike Paterson, and Uri Zwick. Tighter lower bounds on the exact complexity of string matching. SIAM Journal on Computing, 24(1):30–45, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/24582.

Christie:1996:SPB

[Chr96] David A. Christie. Sorting permutations by block-interchanges. Information Processing Letters, 60(4):165–169, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1995:FSM

[Chu95] K.-L. Chung. Fast string matching algorithms for runlength coded strings. Computing: Archiv für informatik und numerik, 54(2):119–125, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/springer.py?Page=10&Key=362&cat=
300607/tocs/springer.py?Page=47&Key=340&cat=3&id_abstract=
244&id_volume=18&id_journal=8.

Cleophas:2006:TRA

[CHZ06] Loek Cleophas, Kees Hemerik, and Gerard Zwaan. Two related algorithms for root-to-frontier tree pattern matching. International Journal of Foundations of Computer Science (IJFCS), 17(6):1253–1272, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Crochemore:1998:TDP

[CIK98] Maxime Crochemore, Costas S. Iliopoulos, and M. Korda. Two-dimensional prefix string matching and covering on square matrices. Algorithmica, 20(4):353–373, April 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=20&issue=4&
spage=353.

Crochemore:2003:OSH

[CIL^{+}03] Maxime Crochemore, Costas S. Iliopoulos, Thierry Lecroq, Yoan J. Pinzon, Wojciech Plandowski, and Wojciech Rytter. Occurrence and substring heuristics for δ-matching. Fundamenta Informaticae, 56(1–2):1–21, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crochemore:2002:ASM

[CIM^{+}02] M. Crochemore, C. Iliopoulos, C. Makris, W. Rytter, A. Tsakalidis, and K. Tsichlas. Approximate string matching with gaps. Nordic Journal of Computing, 9(1):54–??, Spring 2002. CODEN NJCOFR. ISSN 1236-6064.

Chen:1993:SMV

[CJ93] Sei-Wang W. Chen and Anil K. Jain. Strategies of multi-view and multi-matching for 3D object recognition. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):121–130, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

URL http://www.idealibrary.com/links/artid/ciun.
1993.1008/production;   http://www.idealibrary.com/
links/artid/ciun.1993.1008/production/pdf;   http:
//www.idealibrary.com/links/artid/cviu.1993.1008/production;
 http://www.idealibrary.com/links/artid/cviu.1993.
1008/production/pdf.

Chen:2013:PPP

[CJBW13] X. Chen, B. Jones, M. Becchi, and T. Wolf. Picking pesky parameters: Optimizing regular expression matching in practice. In Walid A. Najjar, editor, 2013 ACM/IEEE Symposium on Architectures for Networking and Communications Systems (ANCS), pages 203–213. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4799-1640-4, 1-4799-1639-0. LCCN QA76.9.A73.

Chen:2016:PPP

[CJBW16] Xinming Chen, Brandon Jones, Michela Becchi, and Tilman Wolf. Picking pesky parameters: Optimizing regular expression matching in practice. IEEE Transactions on Parallel and Distributed Systems, 27(5):1430–1442, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.org/csdl/trans/td/2016/05/
07152954-abs.html.

Champarnaud:2012:ARE

[CJM12] Jean-Marc Champarnaud, Hadrien Jeanne, and Ludovic Mignot. Approximate regular expressions and their derivatives. Lecture Notes in Computer Science, 7183:179–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-28332-1_16/.

Clifford:2012:PMM

[CJPS12] Raphaël Clifford, Markus Jalsenius, Ely Porat, and Benjamin Sach. Pattern matching in multiple streams. Lecture Notes in Computer Science, 7354:97–109, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31265-6_8/.

Clifford:2013:SLB

[CJPS13] Raphaël Clifford, Markus Jalsenius, Ely Porat, and Benjamin Sach. Space lower bounds for online pattern matching. Theoretical Computer Science, 483(??):68–74, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397512005816.

Chauhan:2021:ARP

[CJR^{+}21] Komal Chauhan, Kartik Jain, Sayan Ranu, Srikanta Bedathur, and Amitabha Bagchi. Answering regular path queries through exemplars. Proceedings of the VLDB Endowment, 15(2):299–311, October 2021. CODEN ???? ISSN 2150-

8097. URL https://dl.acm.org/doi/10.14778/3489496.
3489510.

Chew:1992:IGP

[CK92] L. P. Chew and K. Kedem. Improvements on geometric pattern matching problems. Lecture Notes in Computer Science, 621:318–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakaravarthy:2002:PCS

[CK02a] Venkatesan T. Chakaravarthy and Rajasekar Krishnamurthy. The problem of context sensitive string matching. Lecture Notes in Computer Science, 2373:64–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-

6119 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2373/23730064.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2373/23730064.pdf.

Chung:2002:TPO

[CK02b] Tae-Sun Chung and Hyoung-Joo Kim. A two phase optimization technique for XML queries with multiple regular path expressions. The Journal of Systems and Software, 64(3):183– 193, December 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cerrito:2004:PMC

[CK04] Serenella Cerrito and Delia Kesner. Pattern matching as cut elimination. Theoretical Computer Science, 323(1–3):71–127, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2008:EEP

[CK08] Norman H. Cohen and Karl Trygve Kalleberg. EventScript: an event-processing language based on regular expressions with actions. ACM SIGPLAN Notices, 43(7):111–120, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cui:2007:SPM

[CKC07] Hang Cui, Min-Yen Kan, and Tat-Seng Chua. Soft pattern matching models for definitional question answering. ACM Transactions on Information Systems, 25(2):8:1–8:??, April 2007. CODEN ATISET. ISSN 1046-8188.

Charalampopoulos:2021:CPM

[CKP^{+}21] Panagiotis Charalampopoulos, Tomasz Kociumaka, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, Juliusz Straszyński, Tomasz Waleń, and Wiktor Zuba. Circular pattern matching with k mismatches. Journal of Computer and System Sciences, 115(??):73–85, February 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000020300702.

Coetser:2009:REH

[CKW09] Wikus Coetser, Derrick G. Kourie, and Bruce W. Watson. On regular expression hashing to reduce FA size. International Journal of Foundations of Computer Science (IJFCS), 20(6): 1069–1086, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chang:1990:ASM

[CL90] W. I. Chang and E. L. Lawler. Approximate string matching in sublinear expected time. In IEEE [IEE90], pages 116– 124. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chang:1992:TEC

[CL92] William I. Chang and Jordan Lampe. Theoretical and empirical comparisons of approximate string matching algorithms. Lecture Notes in Computer Science, 644:172–??, 1992. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1994:SAS

[CL94] William I. Chang and Eugene L. Lawler. Sublinear approximate string matching and biological applications. Algorithmica, 12(4–5):327–344, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Choi:1995:TDP

[CL95] Y. Choi and T. W. Lam. Two-dimensional pattern matching on a dynamic library of texts. Lecture Notes in Computer Science, 959:530–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1996:PMT

[CL96] Maxime Crochemore and Thierry Lecroq. Pattern-matching and text-compression algorithms. ACM Computing Surveys, 28(1):39–41, March 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.

org/pubs/citations/journals/surveys/1996-28-1/p39-
crochemore/;   http://www.acm.org/pubs/toc/Abstracts/
surveys/234331.html.

Charras:1997:ESM

[CL97] Christian Charras and Thierry Lecroq. Exact string matching algorithms. Web site., January 14, 1997. URL http://www-

igm.univ-mlv.fr/~lecroq/string/index.html.

Cameron:2009:ASS

[CL09] Robert D. Cameron and Dan Lin. Architectural support for SWAR text processing with parallel bit streams: the inductive doubling principle. ACM SIGPLAN Notices, 44(3):337– 348, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Champarnaud:2004:RWE

[CLOZ04] Jean-Marc Champarnaud, Éric Laugerotte, Faissal Ouardi, and Djelloul Ziadi. From regular weighted expressions to finite automata. International Journal of Foundations of Computer Science (IJFCS), 15(5):687–??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cho:1995:LHC

[CLP95] J.-W. Cho, S.-Y. Lee, and C. H. Park. On-line handwritten character recognition by a hybrid method based on neural networks and pattern matching. Lecture Notes in Computer Science, 930:926–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Charras:1998:VFS

[CLP98] C. Charras, T. Lecroq, and J. D. Pehoushek. A very fast string matching algorithm for small alphabets and long patterns. Lecture Notes in Computer Science, 1448:55–??, 1998. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cormen:1990:IA

[CLR90] Thomas H. Cormen, Charles E. (Eric) Leiserson, and Ronald L. Rivest. Introduction to Algorithms. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-03141-8, 0-07-013143-0 (McGraw-Hill). xvii + 1028 pp. LCCN QA76.6 .C662 1990.

Chen:1995:STP

[CLS95] Tzer-Shyong Chen, Feipei Lai, and Rung-Ji Shang. A simple tree pattern matching algorithm for code generator. In IEEE [IEE95b], pages 162–167. ISBN 0-8186-7119-X. LCCN ???? IEEE Catalog No. 95CB35838.

Chan:2010:CIA

[CLS^{+}10] Ho-Leung Chan, Tak-Wah Lam, Wing-Kin Sung, Siu-Lung Tam, and Swee-Seong Wong. Compressed indexes for approximate string matching. Algorithmica, 58(2):263–281, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=58&
issue=2&spage=263.

Chan:2013:MAP

[CLST^{+}13] Tak-Ming Chan, Leung-Yau Lo, Ho-Yin Sze-To, Kwong-Sak Leung, Xinshu Xiao, and Man-Hon Wong. Modeling associated protein-DNA pattern discovery with unified scores. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(3):696–707, May 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Chung:2007:EAR

[CLT07] Yun-Sheng Chung, Chin Lung Lu, and Chuan Yi Tang. Efficient algorithms for regular expression constrained sequence alignment. Information Processing Letters, 103(6):240–246, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2015:OEE

[CLZ^{+}15] Lijun Chang, Xuemin Lin, Wenjie Zhang, Jeffrey Xu Yu, Ying Zhang, and Lu Qin. Optimal enumeration: efficient top-k tree matching. Proceedings of the VLDB Endowment, 8(5): 533–544, January 2015. CODEN ???? ISSN 2150-8097.

Chomsky:1958:FSL

[CM58] Noam Chomsky and George A. Miller. Finite state languages. Information and Control, 1(2):91–112, May 1958. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995858900822. This paper show the

equivalence of regular languages and regular sets.

Cisneros:1986:IPL

[CM86] G. Cisneros and H. V. McIntosh. Introduction to the programming language Convert. ACM SIGPLAN Notices, 21(4):

48–57, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Consens:1990:GVF

[CM90] Mariano P. Consens and Alberto O. Mendelzon. GraphLog: a visual formalism for real life recursion. In ACM [ACM90a], pages 404–416. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990.

URL http://www.acm.org/pubs/articles/proceedings/
pods/298514/p404-consens/p404-consens.pdf;   http:
//www.acm.org/pubs/citations/proceedings/pods/298514/

p404-consens/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Chang:1994:ASM

[CM94] W. I. Chang and T. G. Marr. Approximate string matching and local similarity. Lecture Notes in Computer Science, 807: 259–273, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Consens:1995:AQT

[CM95] Mariano P. Consens and Tova Milo. Algebras for querying text regions (extended abstract). In ACM [ACM95b], pages 11–22. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

URL http://www.acm.org/pubs/articles/proceedings/
pods/212433/p11-consens/p11-consens.pdf;   http://
www.acm.org/pubs/citations/proceedings/pods/212433/
p11-consens/;   http://www.acm.org:80/pubs/citations/
proceedings/pods/212433/p11-consens/.

Cormode:2007:SED

[CM07] Graham Cormode and S. Muthukrishnan. The string edit distance matching problem with moves. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cho:2008:IAB

[CM08] Minkyoung Cho and David M. Mount. Improved approximation bounds for planar point pattern matching. Algorithmica, 50(2):175–207, February 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=2&spage=175.

Czerwinski:2017:OTP

[CMNP17] Wojciech Czerwinski, Wim Martens, Matthias Niewerth, and Pawel Parys. Optimizing tree patterns for querying graphand tree-structured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):15–22, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Czerwinski:2018:MTP

[CMNP18] Wojciech Czerwiński, Wim Martens, Matthias Niewerth, and Pawel Parys. Minimization of tree patterns. Journal of the Association for Computing Machinery, 65(4):26:1–26:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandran:2008:IAO

[CMO^{+}08] Nishanth Chandran, Ryan Moriarty, Rafail Ostrovsky, Omkant Pandey, Mohammad Ali Safari, and Amit Sahai. Improved algorithms for optimal embeddings. ACM Transactions on Algorithms, 4(4):45:1–45:14, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2018:SSR

[CMR18] Timothy M. Chan, J. Ian Munro, and Venkatesh Raman. Selection and sorting in the “Restore” model. ACM Transactions on Algorithms, 14(2):11:1–11:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Consens:2010:EXW

[CMRV10] Mariano P. Consens, Renée J. Miller, Flavio Rizzolo, and Alejandro A. Vaisman. Exploring XML Web collections with DescribeX. ACM Transactions on the Web (TWEB), 4(3): 11:1–11:??, July 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cho:2008:DNP

[CMS08] Young H. Cho and William H. Mangione-Smith. Deep network packet filter design for reconfigurable devices. ACM Transactions on Embedded Computing Systems, 7(2):21:1–21:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cruz:1987:GQL

[CMW87] Isabel F. Cruz, Alberto O. Mendelzon, and Peter T. Wood. A graphical query language supporting recursion. In Dayal and Traiger [DT87], pages 323–330. ISBN 0-89791-236-

5. LCCN QA 76.9 D3 P76 1987. URL http://www.acm.

org/pubs/articles/proceedings/mod/38713/p323-cruz/
p323-cruz.pdf;   http://www.acm.org/pubs/citations/
proceedings/mod/38713/p323-cruz/. ACM order number

472870.

Chavez:2002:MIA

[CN02] Edgar Chávez and Gonzalo Navarro. A metric index for approximate string matching. Lecture Notes in Computer Science, 2286:181–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2286/
22860181.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2286/22860181.pdf.

Chen:2021:SMK

[CN21] Yangjun Chen and Hoang Hai Nguyen. On the string matching with k differences in DNA databases. Proceedings of the VLDB Endowment, 14(6):903–915, February 2021. CODEN

???? ISSN 2150-8097. URL https://dl.acm.org/doi/10.
14778/3447689.3447695.

Cho:2015:FAO

[CNPS15] Sukhyeun Cho, Joong Chae Na, Kunsoo Park, and Jeong Seop Sim. A fast algorithm for order-preserving pattern matching. Information Processing Letters, 115(2):397–402, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-

6119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0020019014002336.

Cho:2018:TAB

[CNS18] Sukhyeun Cho, Joong Chae Na, and Jeong Seop Sim. An O(n^{2} log m)-time algorithm for the boxed-mesh permutation pattern matching problem. Theoretical Computer Science, 710(??):35–43, February 1, 2018. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030439751730155X.

Cobbs:1994:FIA

[Cob94] A. L. Cobbs. Fast identification of approximately matching substrings. Lecture Notes in Computer Science, 807:64–74, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1990:CLP

[Coh90] Jacques Cohen. Constraint logic programming languages. Communications of the Association for Computing Machinery, 33(7):52–68, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0001-0782/79209.html.

Cole:1990:TBC

[Col90] R. Cole. Tight bounds on the complexity of the Boyer–Moore pattern matching algorithm. Report 512, Computer Science Dept, New York University, New York, NY, SA, June 1990.

Colussi:1991:CEP

[Col91] Livio Colussi. Correctness and efficiency of pattern matching algorithms. Information and Computation, 95(2):225–251, 1991. CODEN INFCEC. ISSN 0890-5401 (print), 1090-2651

(electronic). URL https://doi.org/10.1016/0890-5401(
91)90046-5.

Cole:1994:TBC

[Col94a] Richard Cole. Tight bounds on the complexity of the Boyer– Moore string matching algorithm. SIAM Journal on Computing, 23(5):1075–1091, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/19554.

Colussi:1994:FPM

[Col94b] L. Colussi. Fastest pattern matching in strings. Journal of Algorithms, 16(2):163–189, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S019667748471008X.

Cook:1972:LTS

[Coo72] Stephen A. Cook. Linear time simulation of deterministic twoway pushdown automata. In Freiman et al. [FGR72], pages 75–80. ISBN 0-7204-2063-6. LCCN QA75.5 .I532 1971. URL

https://books.google.com/books?id=PifeQwAACAAJ.

Cooperman:1986:SMC

[Coo86] Gene Cooperman. A semantic matcher for computer algebra. In Char [Cha86], pages 132–134. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/
32439/p132-cooperman/. ACM order no. 505860.

Cooper:1989:FHO

[Coo89] M. C. Cooper. Formal hierarchical object models for fast template matching. The Computer Journal, 32(4):351–361, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/351.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/352.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/353.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/354.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/355.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/356.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/358.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/359.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/360.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/361.tif.

Cope:1991:RMU

[Cop91] David Cope. Recombinant music: Using the computer to explore musical style. Computer, 24(7):22–28, July 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cox:2007:REM

[Cox07] Russ Cox. Regular expression matching can be simple and fast. Report, swtch.com, Cambridge, MA, USA, January

2007. URL http://swtch.com/~rsc/regexp/regexp1.html.

See also [Tho68, KP99c, Cox09, Cox10a, Cox12].

Cox:2009:REM

[Cox09] Russ Cox. Regular expression matching: the virtual machine approach. Report, swtch.com, Cambridge, MA, USA, Decem-

ber 2009. URL http://swtch.com/~rsc/regexp/regexp2.

html. See also [Tho68, KP99c, Cox07, Cox10a, Cox12].

Cox:2010:REM

[Cox10a] Russ Cox. Regular expression matching in the wild. Report, swtch.com, Cambridge, MA, USA, March 2010. URL http:

//swtch.com/~rsc/regexp/regexp3.html. See also [Tho68,

KP99c, Cox07, Cox09, Cox12].

Cox:2010:REP

[Cox10b] Russ Cox. re2: an efficient, principled regular expression library. Google Code project, March 2, 2010. URL http://

code.google.com/p/re2/.

Cox:2010:YD

[Cox10c] Russ Cox. Yacc is not dead. Web site., December 6, 2010.

URL http://research.swtch.com/yaccalive.

Cox:2012:REM

[Cox12] Russ Cox. Regular expression matching with a trigram index, or how Google Code Search worked. Report, swtch.com, Cambridge, MA, USA, January 2012. URL http://swtch.

com/~rsc/regexp/regexp4.html. See also [Tho68, KP99c,

Cox07, Cox09, Cox10a].

Cox:2019:SSD

[Cox19] Russ Cox. Surviving software dependencies. Communications of the Association for Computing Machinery, 62(9):36– 43, September 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://cacm.acm.

org/magazines/2019/9/238968/fulltext.

Champarnaud:2009:ECE

[COZ09] Jean-Marc Champarnaud, Faissal Ouardi, and Djelloul Ziadi. An efficient computation of the equation K-automaton of a regular K-expression. Fundamenta Informaticae, 90(1–2): 1–16, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crochemore:1991:TWS

[CP91] Maxime Crochemore and Dominique Perrin. Two-way string matching. Journal of the Association for Computing Machinery, 38(3):651–675, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/116845.html.

Chang:1997:RED

[CP97] Chia-Hsiang Chang and Robert Paige. From regular expressions to DFA’s using compressed NFA’s. Theoretical Computer Science, 178(1–2):1–36, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2343.

Cohen:2010:FSI

[CP10] Hagai Cohen and Ely Porat. Fast set intersection and twopatterns matching. Theoretical Computer Science, 411(40– 42):3795–3800, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:1992:MEB

[CPT92] J. Cai, R. Paige, and R. Tarjan. More efficient bottom-up multi-pattern matching in trees. Theoretical Computer Science, 106(1):21–60, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carroll:1988:RBP

[CPW88] David M. Carroll, Christine A. Pogue, and Peter Willett. Research: Bibliographic pattern matching using the ICL Distributed Array Processor. Journal of the American Society for Information Science, 39(6):390–399, November 1988. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chrobak:1987:RSM

[CR87] Marek Chrobak and Wojciech Rytter. Remarks on stringmatching and one-way multihead automata. Information Processing Letters, 24(5):325–329, March 16, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1991:UKM

[CR91] M. Crochemore and W. Rytter. Usefulness of the Karp-Miller-Rosenberg algorithm in parallel computations on strings and arrays. Theoretical Computer Science, 88(1):59–82, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1992:NTD

[CR92] M. Crochemore and W. Rytter. Note on two-dimensional pattern matching by optimal parallel algorithms. Lecture Notes in Computer Science, 654:100–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1994:TDP

[CR94] Maxime Crochemore and Wojciech Rytter. On twodimensional pattern matching by optimal parallel algorithms. Theoretical Computer Science, 132(1–2):403– 414, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1994&volume=132&issue=1-
2&aid=1653.

Chen:1995:FPM

[CR95a] Shenfeng Chen and J. H. Reif. Fast pattern matching for entropy bounded text. In Storer and Cohn [SC95], pages 282–291. ISBN 0-8186-7012-6. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515518. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Crochemore:1995:LTA

[CR95b] M. Crochemore and W. Rytter. On linear-time alphabetindependent 2-dimensional pattern matching. Lecture Notes in Computer Science, 911:220–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1995:SCT

[CR95c] Maxime Crochemore and Wojciech Rytter. Squares, cubes, and time-space efficient string searching. Algorithmica, 13(5): 405–425, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Crochemore:1986:TR

[Cro86] M. Crochemore. Transducers and repetitions. Theoretical Computer Science, 45(1):63–86, ???? 1986. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1992:FSI

[Cro92a] M. Crochemore. Foreword to the Special Issue on Selected Papers of the Combinatorial Pattern Matching School, Paris, July 1990. Theoretical Computer Science, 92(1):1–??, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1992:SMO

[Cro92b] M. Crochemore. String-matching on ordered alphabets. Theoretical Computer Science, 92(1):33–47, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chazottes:2006:APM

[CRV06] Jean-Rene Chazottes, Frank Redig, and Evgeny Verbitskiy. On approximate pattern matching for a class of Gibbs random fields. Annals of Applied Probability, 16(2):670–684, May 2006. CODEN ???? ISSN 1050-5164. URL http://

projecteuclid.org/euclid.aoap/1151592247.

Cardoze:1998:PMS

[CS98] D. E. Cardoze and L. J. Schulman. Pattern matching for spatial point sets. In IEEE [IEE98], pages 156–165. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Campeanu:2009:IRL

[CS09] Cezar Câmpeanu and Nicolae Santean. On the intersection of regex languages with regular languages. Theoretical Computer Science, 410(24–25):2336–2344, May 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choi:2011:CPM

[CS11a] Yongwook Choi and Wojciech Szpankowski. Constrained pattern matching. ACM Transactions on Algorithms, 7(2):25:1– 25:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Coquand:2011:DPR

[CS11b] T. Coquand and V. Siles. A decision procedure for regular expression equivalence in type theory. In Jean-Pierre Jouannaud and Zhong Shao, editors, Proceedings of the 1st International Conference on Certified Programs and Proofs (CPP), volume 7086 of Lecture Notes in Computer Science, pages 119–134. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011.

Carver:2018:SME

[CS18] J. C. Carver and A. Serebrenik. Software maintenance and evolution and automated software engineering. IEEE Software, 35(2):102–104, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cseresnyes:2022:REL

[CS22] Ehud Cseresnyes and Hannes Seiwert. Regular expression length via arithmetic formula complexity. Journal of Computer and System Sciences, 125(??):1–24, May 2022. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002200002100101X.

Campeanu:2003:FSP

[CSY03] Cezar Câmpeanu, Kai Salomaa, and Sheng Yu. A formal study of practical regular expressions. International Journal of Foundations of Computer Science (IJFCS), 14(6):1007–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Colussi:1996:HCC

[CT96] Livio Colussi and Laura Toniolo. How the character comparison order shapes the shift function of on-line pattern matching algorithms. Theoretical Computer Science, 163 (1–2):117–144, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=163&issue=1-
2&aid=2078.

Chida:2023:RRE

[CT23] Nariyoshi Chida and Tachio Terauchi. Repairing regular expressions for extraction. Proceedings of the ACM on Programming Languages (PACMPL), 7(PLDI):173:1–173:??, June 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.1145/3591287.

Chen:1998:EAS

[CTF^{+}98] S. W. Chen, S. T. Tung, C. Y. Fang, Shen Cherng, and Anil K. Jain. Extended attributed string matching for shape recognition. Computer Vision and Image Understanding: CVIU, 70(1):36–50, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

URL http://www.idealibrary.com/links/artid/cviu.
1998.0599/production;   http://www.idealibrary.com/
links/artid/cviu.1998.0599/production/pdf;   http:
//www.idealibrary.com/links/artid/cviu.1998.0599/production/
ref.

IEEE:1986:PCI

[CVP86] Proceedings, CVPR ’86 (IEEE Computer Society Conference on Computer Vision and Pattern Recognition, Miami Beach, FL, June 22–26, 1986), IEEE Publ. 86CH2290-5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0721-1. LCCN TA1632 .I36 1986.

Cappers:2018:EME

[CvW18] B. C. M. Cappers and J. J. van Wijk. Exploring multivariate event sequences using rules, aggregations, and selections. IEEE Transactions on Visualization and Computer Graphics, 24(1):532–541, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cleary:1984:DCU

[CW84] John G. Cleary and I. H. Witten. Data compression using adaptive coding and partial string matching. IEEE Transactions on Communications, COM-32(4):396–402, April 1984. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Chen:2013:EMT

[CW13] Chien-Chi Chen and Sheng-De Wang. An efficient multicharacter transition string-matching engine based on the Aho– Corasick algorithm. ACM Transactions on Architecture and Code Optimization, 10(4):25:1–25:??, December 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2018:SMMa

[CW18a] Yangjun Chen and Yujia Wu. On the string matching with k mismatches. Theoretical Computer Science, 726(??):5–29, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397518300781.

Chen:2018:SMMb

[CW18b] Yangjun Chen and Yujia Wu. On the string matching with k mismatches. Theoretical Computer Science, 726(??):5–29, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397518300781.

Chen:2021:TDV

[CWL^{+}21] Fei Chen, Donghong Wang, Qiuzhen Lin, Jianyong Chen, Zhong Ming, Wei Yu, and Jing Qin. Towards dynamic verifiable pattern matching. IEEE Transactions on Big Data, 7 (2):421–435, June 2021. ISSN 2332-7790.

Cleophas:2010:NTS

[CWZ10] Loek Cleophas, Bruce W. Watson, and Gerard Zwaan. A new taxonomy of sublinear right-to-left scanning keyword pattern matching algorithms. Science of Computer Programming, 75 (11):1095–1112, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:2020:IAO

[CX20] Haiming Chen and Zhiwu Xu. Inclusion algorithms for oneunambiguous regular expressions and their applications. Science of Computer Programming, 193(??):??, July 1, 2020.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016764231830296X.

Champarnaud:2001:CEA

[CZ01] Jean-Marc Champarnaud and Djelloul Ziadi. Computing the equation automaton of a regular expression in O(s^{2}) space and time. Lecture Notes in Computer Science, 2089:157–168, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2089/20890157.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2089/20890157.pdf.

Chen:2009:DPM

[CZCD09] Zhongqiang Chen, Yuan Zhang, Zhongrong Chen, and Alex Delis. A digest and pattern matching-based intrusion detection engine. The Computer Journal, 52 (6):699–723, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/cgi/content/abstract/52/
6/699;   http://comjnl.oxfordjournals.org/cgi/reprint/
52/6/699.

Calvo-Zaragoza:2017:CEE

[CZOdlH17] Jorge Calvo-Zaragoza, Jose Oncina, and Colin de la Higuera. Computing the expected edit distance from a string to a probabilistic finite-state automaton. International Journal of Foundations of Computer Science (IJFCS), 28(5):603–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Chen:2015:PMV

[CZW15] Ke Chen, Zhong Zhou, and Wei Wu. Progressive motion vector clustering for motion estimation and auxiliary tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):33:1–33:??, January 2015. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Devi:2023:PMM

[D^{+}23] K. Durga Devi et al. Pattern matching model for recognition of stone inscription characters. The Computer Journal, 66(3): 554–564, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://academic.oup.

com/comjnl/article/66/3/554/6424463.

DAndrea:1998:DEP

[D’A98] G. D’Andrea. Distance evaluation in pattern matching based on frontier topological graph. Lecture Notes in Computer Science, 1451:132–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darivandpour:2018:ESP

[DA18] Javad Darivandpour and Mikhail J. Atallah. Efficient and secure pattern matching with wildcards using lightweight cryptography. Computers & Security, 77(??):666–674, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481830021X.

Darragh:2020:PZF

[DA20] Pierce Darragh and Michael D. Adams. Parsing with zippers (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):108:1–108:28, August 2020.

URL https://dl.acm.org/doi/10.1145/3408990.

Dai:2009:AAM

[Dai09] Liuling Dai. An aggressive algorithm for multiple string matching. Information Processing Letters, 109(11):553–559, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danvy:1991:SDC

[Dan91] Olivier Danvy. Semantics-directed compilation of nonlinear patterns. Information Processing Letters, 37(6):315–322, March 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davison:1973:RSC

[Dav73] G. A. Davison. Rapidly searching for character string matches using hash coding. IBM Technical Disclosure Bulletin, 16(1): ??, June 1973. CODEN IBMTAA. ISSN 0018-8689.

Davies:1982:SST

[Dav82] D. Julian M. Davies. String searching in text editors. Software — Practice and Experience, 12(8):709–717, August 1982.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davis:1999:URE

[Dav99] Mark Davis. Unicode regular expression guidelines. Unicode Technical Report 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, November 23, 1999. URL http://www.

unicode.org/unicode/reports/tr18/.

Davis:2003:URE

[Dav03] Mark Davis. Unicode regular expressions. Proposed Update Unicode Technical Standard 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, August 13, 2003. URL http:

//www.unicode.org/reports/tr18/tr18-8.html.

Davis:2004:URE

[Dav04] Mark Davis. Unicode regular expressions. Unicode Technical Standard 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, January 9, 2004. URL http://www.unicode.

org/unicode/reports/tr18/.

Davis:2021:URE

[Dav21] Mark Davis. Unicode regular expressions. Unicode Technical Report 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, February 16, 2021. URL http://www.unicode.

org/unicode/reports/tr18/tr18-22.html.

Davis:2022:URE

[Dav22] Mark Davis. Unicode regular expressions. Unicode Technical Report 18, The Unicode Consortium, San Jose, CA 95170-0519, USA, February 8, 2022. URL https://www.unicode.

org/reports/tr18/tr18-23.html.

Davies:1986:APM

[DB86] G. Davies and S. Bowsher. Algorithms for pattern matching. Software — Practice and Experience, 16(6):575–601, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deAlmeida:1993:SMA

[dB93] Nalvo F. de Almeida, Jr. and Valmir C. Barbosa. A stringmatching algorithm for the CREW PRAM. Information Processing Letters, 47(5):257–259, October 8, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deBruijn:2008:PFL

[dBB08] Willem de Bruijn and Herbert Bos. PipesFS: fast Linux I/O in the Unix tradition. Operating Systems Review, 42(5):55– 63, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Du:1994:ADV

[DC94] M.-W. Du and S. C. Chang. Approach to designing very fast approximate string matching algorithms. IEEE Transactions on Knowledge and Data Engineering, 6(4):620–633, August 1994. CODEN ITKEEH. ISSN 1041-4347 (print), 1558-2191 (electronic).

Doherty:2015:PMT

[DCM15] Jonathan Doherty, Kevin Curran, and Paul McKevitt. Pattern matching techniques for replacing missing sections of audio streamed across wireless networks. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):25:1–25:??, April 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

deVSmit:1982:CTS

[de 82] G. de V. Smit. A comparison of three string matching algorithms. Software — Practice and Experience, 12(1):57–66, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deorowicz:2006:STI

[Deo06] Sebastian Deorowicz. Speeding up transposition-invariant string matching. Information Processing Letters, 100(1):14– 20, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dermouche:1995:FAS

[Der95] A. Dermouche. A fast algorithm for string matching with mismatches. Information Processing Letters, 55(2):105–110, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dube:2000:EBP

[DF00] Danny Dubé and Marc Feeley. Efficiently building a parse tree from a regular expression. Acta Informatica, 37 (2):121–144, September 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00236/
bibs/0037002/00370121.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/0037002/
00370121.pdf.

Deaton:1993:ACS

[DGBH93] E. Deaton, K. M. George, H. Bergel, and G. Hedrick, editors. Applied Computing: States of the Art and Practice — 1993 Proceedings of the 1993 ACM/SIGAPP Symposium on Applied Computing — February 1993, Indianapolis, IN, USA, Applied Computing — Symposium 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-567-4. LCCN QA76.76.A65 S95 1993.

Daykin:2019:EPM

[DGG^{+}19] J. W. Daykin, R. Groult, Y. Guesnet, T. Lecroq, A. Lefebvre, M. Léonard, L. Mouchard, É. Prieur-Gaston, and B. Watson. Efficient pattern matching in degenerate strings with the Burrows–Wheeler transform. Information Processing Letters, 147(??):82–87, July 2019. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019019300535.

Dubiner:1990:FTP

[DGM90] M. Dubiner, Z. Galil, and E. Magen. Faster tree pattern matching. In IEEE [IEE90], pages 145–150. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Dubiner:1994:FTP

[DGM94] Moshe Dubiner, Zvi Galil, and Edith Magen. Faster tree pattern matching. Journal of the Association for Computing Machinery, 41(2):205–213, March 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174653.html.

Decaroli:2019:CIO

[DGM19] Gianni Decaroli, Travis Gagie, and Giovanni Manzini. A compact index for order-preserving pattern matching. Software — Practice and Experience, 49(6):1041–1051, June 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

dAmorim:2005:EBR

[dH05] Marcelo d’Amorim and Klaus Havelund. Event-based runtime verification of Java programs. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Durian:2010:IPE

[ĎHPT10] Branislav Ďurian, Jan Holub, Hannu Peltola, and Jorma Tarhio. Improving practical exact string matching. Information Processing Letters, 110(4):148–152, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1976:PMP

[Dij76] Edsger W. Dijkstra. The pattern matching problem, chapter 18. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:20xx:PMP

[Dijxx] Edsger W. Dijkstra. The pattern-matching problem. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd04xx/EWD459.PDF.

Ditzel:1978:PMH

[Dit78] D. R. Ditzel. Pattern matching for high level languages. ACM SIGPLAN Notices, 13(5):46–55, May 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeBosschere:1996:EFL

[DJ96] Koen De Bosschere and Jean-Marie Jacquet. Extending the μLog framework with local and conditional blackboard operations. Journal of Symbolic Computation, 21(4/5/6):669–698 (or 669–697??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Deo:2013:PSA

[DK13] Mrinal Deo and Sean Keely. Parallel suffix array and least common prefix for the GPU. ACM SIGPLAN Notices, 48(8): 197–206, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

DAntoni:2015:HCA

[DKA^{+}15] Loris D’Antoni, Dileep Kini, Rajeev Alur, Sumit Gulwani, Mahesh Viswanathan, and Björn Hartmann. How can automatic feedback help students construct automata? ACM Transactions on Computer-Human Interaction, 22(2):9:1– 9:??, April 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

deKretser:2004:SSE

[dKM04] Owen de Kretser and Alistair Moffat. SEFT: a search engine for text. Software — Practice and Experience, 34(10): 1011–1023, August 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denning:2011:MIV

[DKP11] Jonathan D. Denning, William B. Kerr, and Fabio Pellacini. MeshFlow: interactive visualization of mesh construction sequences. ACM Transactions on Graphics, 30(4):66:1–66:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeNicola:2003:NRE

[DL03] Rocco De Nicola and Anna Labella. Nondeterministic regular expressions as solutions of equational systems. Theoretical Computer Science, 302(1–3):179–189, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLima:2022:SAD

[dLBHC22] João Paulo Cardoso de Lima, Marcelo Brandalero, Michael Hübner, and Luigi Carro. STAP: an architecture and design tool for automata processing on memristor TCAMs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):39:1–39:22, April 2022. CODEN ????

ISSN 1550-4832. URL https://dl.acm.org/doi/10.1145/
3450769.

Deng:2015:UFA

[DLF^{+}15] Dong Deng, Guoliang Li, Jianhua Feng, Yi Duan, and Zhiguo Gong. A unified framework for approximate dictionary-based entity extraction. VLDB Journal: Very Large Data Bases, 24 (1):143–167, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

daLuz:2007:RET

[dLFM07] Robson da Luz, Mírian Halfeld Ferrari, and Martin A. Musicante. Regular expression transformations to extend regular languages (with application to a Datalog XML schema validator). Journal of Algorithms, 62(3–4):148–167, July/ October 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0196677407000326.

Deng:2012:TPM

[DLG12] Ziqiang Deng, Husheng Liao, and Hongyu Gao. Twig pattern matching running on XML streams. Lecture Notes in Computer Science, 7234:35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29426-6_6/.

Droste:2011:WAR

[DM11] Manfred Droste and Ingmar Meinecke. Weighted automata and regular expressions over valuation monoids. International Journal of Foundations of Computer Science (IJFCS), 22(8): 1829–1844, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dediu:2013:LAT

[DMVT13] Adrian-Horia Dediu, Carlos Martín-Vide, and Bianca Truthe, editors. Language and Automata Theory and Applications:

7th International Conference, LATA 2013, Bilbao, Spain, April 2–5, 2013. Proceedings, volume 7810 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-37063-2 (print), 3-642-37064-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5-76.95.

Davis:1977:ARE

[DMWW77] D. E. Davis, R. D. Moore, M. C. Williams, and O. C. Woodard. Automatic registration in an electron-beam lithographic system. IBM Journal of Research and Development, 21(6):498–505, November 1977. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Dirilten:1977:PMU

[DN77] H. Dirilten and T. G. Newman. Pattern matching under affine transformations. IEEE Transactions on Computers, C-26(3): 314–317, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1674832.

Dorohonceanu:2000:AAP

[DNM00] Bogdan Dorohonceanu and Craig Nevill-Manning. Algorithm alley: A practical suffix-tree implementation for string searches. Dr. Dobb’s Journal of Software Tools, 25(7):133– 136, 140, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/ftp/2000/2000_07/aa700.txt;
 http://www.ddj.com/ftp/2000/2000_07/aa700.zip.

Ducasse:2006:ECD

[DNR06] Stéphane Ducasse, Oscar Nierstrasz, and Matthias Rieger. On the effectiveness of clone detection by string matching. Journal of Software Maintenance and Evolution: Research and Practice, 18(1):37–58, January 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DellaVentura:1993:PES

[DOS93] A. Della Ventura, P. Ongaro, and R. Schettini. Pictorial editing by shape matching techniques. Computer Graphics Forum, 12(2):111–122, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dougherty:1991:SA

[Dou91] Dale Dougherty. sed & awk. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 1991. ISBN 0-937175-59-5. xxii + 394 pp. LCCN QA76.76.U84 D69 1991.

URL http://www.oreilly.com/catalog/9780937175590.

Dowek:1991:SOP

[Dow91] G. Dowek. A second-order pattern matching algorithm for the cube of typed lambda-calculi. Lecture Notes in Computer Science, 520:151–??, 1991. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dowek:1993:UPM

[Dow93] Gilles Dowek. The undecidability of pattern matching in calculi where primitive recursive functions are representable. Theoretical Computer Science, 107(2):349–356, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1295.

Danvy:2006:OBM

[DR06] Olivier Danvy and Henning Korsholm Rohde. On obtaining the Boyer–Moore string-matching algorithm by partial evaluation. Information Processing Letters, 99(4):158–162, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deRezende:1995:PSP

[dRL95] P. J. de Rezende and D. T. Lee. Point set pattern matching in d-dimensions. Algorithmica, 13(4):387–404, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dawson:1996:PPU

[DRSS96] Steven Dawson, C. R. Ramakrishnan, Steven Skiena, and Terrance Swift. Principles and practice of unification factoring. ACM Transactions on Programming Languages and Systems, 18(5):528–563, September 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/0164-0925/232722.
html;   http://www.acm.org/pubs/toc/Abstracts/toplas/
232722.html;   http://www.acm.org/pubs/toc/Abstracts/
toplas/235455.html.

Dubois:1995:EP

[DRW95] Catherine Dubois, François Rouaix, and Pierre Weis. Extensional polymorphism. In ACM [ACM95a], pages 118–129. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995.  URL http://www.acm.org:80/pubs/citations/
proceedings/plan/199448/p118-dubois/.  ACM order

number: 549950.

Daptardar:2004:AKM

[DS04] A. Daptardar and D. Shapira. Adapting the Knuth–Morris– Pratt algorithm for pattern matching in Huffman encoded texts. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314. LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281511. IEEE

Computer Society Order Number: P2082.

Dixit:2019:FBD

[DS19] Umesh D. Dixit and M. S. Shirdhonkar. Fingerprintbased document image retrieval. International Journal of Image and Graphics (IJIG), 19(2):??, 2019. ISSN 0219-

4678. URL https://www.worldscientific.com/doi/10.
1142/S0219467819500086.

Oliveira:2015:MRM

[dSOMY15] Bruno C. d. S. Oliveira, Shin-Cheng Mu, and Shu-Hung You. Modular reifiable matching: a list-of-functors approach to two-level types. ACM SIGPLAN Notices, 50(12):82–93, December 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Das:1994:SAI

[DSv94] R. Das, J. Saltz, and R. von Hanxleden. Slicing analysis and indirect accesses to distributed arrays. In Banerjee et al. [BGNP94], pages 152–168. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Dayal:1987:PAC

[DT87] Umeshwar Dayal and Irv Traiger, editors. Proceedings of Association for Computing Machinery Special Interest Group on Management of Data 1987 annual conference, San Francisco, May 27–29, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76 1987. ACM order number 472870.

Duff:1982:CBS

[Duf82] Steven G. Duff. The case for BUFFERS in SNOBOL4. ACM SIGPLAN Notices, 17(11):24–30, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Durand:1994:BSS

[Dur94] Iréne Durand. Bounded, strongly sequential and forwardbranching term rewriting systems. Journal of Symbolic Computation, 18(4):319–352, October 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Durr:2021:CPP

[DV21] Christoph Dürr and Jill-Jênn Vie. Competitive programming in Python: 128 algorithms to develop your coding skills. Cambridge University Press, Cambridge, UK, 2021. ISBN 1-108-71682-2 (paperback), 1-108-59192-2 (e-pub).

LCCN QA76.73.P98. URL https://assets.cambridge.
org/97811087/16826/toc/9781108716826_toc.pdf. Trans-

lation to English from the French original by Greg Gibbons and Danièle Gibbons.

Daykin:2017:ISF

[DW17] Jacqueline W. Daykin and Bruce Watson. Indeterminate string factorizations and degenerate text transformations. Mathematics in Computer Science, 11(2):209–218, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11786-016-0285-x.pdf.

Delcambre:1989:PMR

[DWE89] L. M. L. Delcambre, J. Waramahaputi, and J. N. Etheredge. Pattern match reduction for Relational Production Language in the USL MMDBS. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):59–67, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dwelly:2000:XRP

[Dwe00] Andrew Dwelly. XML, reflective pattern matching, and Java. Dr. Dobb’s Journal of Software Tools, 25(6):46, 49– 52, 54, June 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/ftp/2000/2000_06/marius05.
zip;   http://www.ddj.com/ftp/2000/2000_06/xmljava.
txt.

Dyadkin:1994:MP

[Dya94] Lev J. Dyadkin. Multibox parsers. ACM SIGSOFT Software Engineering Notes, 19(3):23–25, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Earley:1974:HLO

[Ear74] J. Earley. High level operations in automatic programming. ACM SIGPLAN Notices, 9(4):34–42, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elseidy:2014:GFS

[EASK14] Mohammed Elseidy, Ehab Abdelhamid, Spiros Skiadopoulos, and Panos Kalnis. GraMi: frequent subgraph and pattern mining in a single large graph. Proceedings of the VLDB Endowment, 7(7):517–528, March 2014. CODEN ???? ISSN 2150-8097.

Eckel:1989:TSI

[Eck89] Bruce Eckel. TAWK: a simple interpreter in C++. Dr. Dobb’s Journal of Software Tools, 14(5):50–60, 98, 100–102, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Edmiston:1988:PPB

[ECSS88] Elizabeth W. Edmiston, Nolan G. Core, Joel H. Saltz, and Roger M. Smith. Parallel processing of biological sequence comparison algorithms. International Journal of Parallel Programming, 17(3):259–275, June 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Edwards:2007:NIA

[Edw07] Jonathan Edwards. No ifs, ands, or buts: uncovering the simplicity of conditionals. ACM SIGPLAN Notices, 42(10): 639–658, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Einwohner:1995:STI

[EF95] T. H. Einwohner and Richard J. Fateman. Searching techniques for integral tables. In Levelt [Lev95], pages 133–139.

ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/

220346/p133-einwohner/. ACM order number: 505950.

ElDefrawy:2013:BDS

[EF13] Karim El Defrawy and Sky Faber. Blindfolded data search via secure pattern matching. Computer, 46(12):68–75, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Esparza:2014:PBV

[EGP14] Javier Esparza, Pierre Ganty, and Tomás Poch. Patternbased verification for multithreaded programs. ACM Transactions on Programming Languages and Systems, 36(3):9:1–9:??, September 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ehrenfeucht:1988:NDM

[EH88] A. Ehrenfeucht and D. Haussler. A new distance metric on strings computable in linear time. Discrete Applied Mathematics, 20(3):191–203, July 1988. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Engelfriet:2007:XTT

[EHS07] Joost Engelfriet, Hendrik Jan Hoogeboom, and Bart Samwel. XML transformation by tree-walking transducers with invisible pebbles. In ACM [ACM07], pages 63–72. ISBN 1-59593-685-8. LCCN ????

Efrat:2004:PMS

[EIV04] Alon Efrat, Piotr Indyk, and Suresh Venkatasubramanian. Pattern matching for sets of segments. Algorithmica, 40(3):147–160, August 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=3&spage=147.

Edixhoven:2023:BCR

[EJ23] Luc Edixhoven and Sung-Shik Jongmans. Balanced-by-construction regular and ω-regular languages. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):117–144, February–April 2023. ISSN 0129-0541.

URL https://www.worldscientific.com/doi/10.1142/
S0129054122440026.

Eker:1995:ACM

[Eke95] S. M. Eker. Associative-commutative matching via bipartite graph matching. The Computer Journal, 38(5): 381–399, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/38/5/381.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/Volume_38/
Issue_05/Vol38_05.body.html#AbstractEker.

Egolf:2022:VVO

[ELF22] Derek Egolf, Sam Lasser, and Kathleen Fisher. Verbatim++: verified, optimized, and semantically rich lexing with derivatives. In Andrei Popescu and Steve Zdancewic, editors, Proceedings of the 11th ACM SIGPLAN International Conference on Certified Programs and Proofs, pages 27–39. ACM Press, New York, NY 10036, USA, January 2022. ISBN 1-4503-9182-

6. LCCN ???? URL https://doi.org/10.1145/3497775.
3503694.

El-Mabrouk:1996:BMS

[EMC96] N. El-Mabrouk and M. Crochemore. Boyer–Moore strategy to efficient approximate string matching. Lecture Notes in Computer Science, 1075:24–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Equi:2023:GCI

[EMT23] Massimo Equi, Veli Mäkinen, and Alexandru I. Tomescu. Graphs cannot be indexed in polynomial time for subquadratic time string matching, unless SETH fails. Theoretical Computer Science, 975(??):??, October 9, 2023. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397523004413.

Equi:2023:CSM

[EMTG23] Massimo Equi, Veli Mäkinen, Alexandru I. Tomescu, and Roberto Grossi. On the complexity of string matching for graphs. ACM Transactions on Algorithms, 19(3):21:1–21:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3588334.

Emmelmann:1989:BGE

[ESL89] Helmut Emmelmann, Friedrich-Wilhelm Schröer, and Rudolf Landwehr. BEG: a generator for efficient back ends. ACM SIGPLAN Notices, 24(7):227–237, July 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-

1160 (electronic). URL http://www.acm.org:80/pubs/
citations/proceedings/pldi/73141/p227-emmelmann/.

Eilam-Tzoreff:1988:MPS

[ETV88] Tali Eilam-Tzoreff and Uzi Vishkin. Matching patterns in strings subject to multi-linear transformations. Theoretical Computer Science, 60(3):231–254, December 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engels:2021:SVS

[ETV21] Steven Engels, Tony Tan, and Jan Van den Bussche. Subsequence versus substring constraints in sequence pattern languages. Acta Informatica, 58(1–2):35–56, April 2021. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00236-019-00347-5.

Ellis:1998:REC

[EU98] Duncan Ellis and Sameer Udeshi. A regular expression class library. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Ehrenfeucht:1974:CMR

[EZ74] Andrzej Ehrenfeucht and Paul Zeiger. Complexity measures for regular expressions. In ACM [ACM74], pages 75–79. LCCN QA76.6 .A13 1974.

Ehrenfeucht:1976:CMR

[EZ76] Andrzej Ehrenfeucht and Paul Zeiger. Complexity measures for regular expressions. Journal of Computer and System Sciences, 12(2):134–146, April 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000076800347.

Espindola:2023:SMR

[EZYA23] Vinicius Espindola, Luciano Zago, Hervé Yviquel, and Guido Araujo. Source matching and rewriting for MLIR using stringbased automata. ACM Transactions on Architecture and Code Optimization, 20(2):22:1–22:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/3571283.

Faro:2023:LCC

[F^{+}23] Simone Faro et al. On the longest common Cartesian substring problem. The Computer Journal, 66(4):907–923, April 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.oup.com/comjnl/article/
66/4/907/6500728.

Faloutsos:1985:AMT

[Fal85] Christos Faloutsos. Access methods for text. ACM Computing Surveys, 17(1):49–74, March 1985. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0360-0300/4080.html. Also

published in/as: “Multiattribute Hashing Using Gray Codes”, ACM SIGMOD, 1986.

Farnum:1992:PBT

[Far92a] Charles Farnum. Pattern-based tree attribution. In ACM [ACM92a], pages 211–222. ISBN 0-89791-453-8. LCCN

QA76.7 .A15 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/plan/143165/p211-farnum/. ACM

order number 54990.

Farnum:1992:PTA

[Far92b] Charles Farnum. Pattern-based tree attribution. In ACM [ACM92a], pages 211–222. ISBN 0-89791-453-8. LCCN

QA76.7 .A15 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/plan/143165/p211-farnum/. ACM

order number 54990.

Farr:2019:UGT

[Far19] Graham Farr. Using Go in teaching the theory of computation. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 50(1):65–78, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fateman:2015:PAS

[Fat15] Richard Fateman. Partitioning of algebraic subexpressions in computer algebra systems: an alternative to matching with an application to symbolic integration. ACM Communications in Computer Algebra, 49(2):38–47, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Firth:2005:CBA

[FBMA05] Andrew Firth, Tim Bell, Amar Mukherjee, and Don Adjeroh. A comparison of BWT approaches to string pattern matching. Software — Practice and Experience, 35(13):1217–1258, November 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frakes:1992:IRD

[FBY92] William B. Frakes and Ricardo Baeza-Yates, editors. Information Retrieval: Data Structures and Algorithms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-463837-9. viii + 504 pp. LCCN QA76.9.D351543 1992. US$56.00.

Farach-Colton:1998:CPM

[FC98] Martin Farach-Colton, editor. Combinatorial pattern matching: 9th Annual Symposium, CPM 98, Piscataway, New Jersey, USA, July 20–22, 1998: Proceedings, volume 1448 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CO-DEN LNCSD9. ISBN 3-540-64739-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

1998. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t1448.htm;   http://www.springerlink.
com/content/978-3-540-64739-3;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
1448.

Frisch:2004:GRE

[FC04] Alain Frisch and Luca Cardelli. Greedy regular expression matching. In Josep Díaz, Juhani Karhumäki, Arto Lepistö, and Donald Sannella, editors, Automata, Languages and Programming: 31st International Colloquium, ICALP 2004, Turku, Finland, July 12–16, 2004, Proceedings, volume 3142 of Lecture Notes in Computer Science, pages 618–629. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004.

Farach-Colton:2000:SCS

[FCFM00] Martin Farach-Colton, Paolo Ferragina, and S. Muthukrishnan. On the sorting-complexity of suffix tree construction. Journal of the Association for Computing Machinery, 47(6):987–1011, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org/pubs/citations/journals/jacm/2000-47-
6/p987-farach-colton/.

Farach-Colton:2007:OSS

[FCLST07] Martin Farach-Colton, Gad M. Landau, S. Cenk Sahinalp, and Dekel Tsur. Optimal spaced seeds for faster approximate string matching. Journal of Computer and System Sciences, 73(7):1035–1044, November 2007. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000007000256.

Ficara:2011:DED

[FDG^{+}11] Domenico Ficara, Andrea Di Pietro, Stefano Giordano, Gregorio Procissi, Fabio Vitucci, and Gianni Antichi. Differential encoding of DFAs for fast regular expression matching. IEEE/ ACM Transactions on Networking, 19(3):683–694, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fenwick:2001:FSM

[Fen01a] Peter Fenwick. Fast string matching for multiple searches. Software — Practice and Experience, 31(9):815–833, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/78505028/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
78505028&PLACEBO=IE.pdf.

Fenwick:2001:SPP

[Fen01b] Peter Fenwick. Some perils of performance prediction: a case study on pattern matching. Software — Practice and Experience, 31(9):835–843, July 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/78505029/START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?ID=78505029&PLACEBO=IE.
pdf.

Furr:2008:CTS

[FF08] Michael Furr and Jeffrey S. Foster. Checking type safety of foreign function calls. ACM Transactions on Programming Languages and Systems, 30(4):18:1–18:63, July 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fan:2015:KG

[FFTD15] Wenfei Fan, Zhe Fan, Chao Tian, and Xin Luna Dong. Keys for graphs. Proceedings of the VLDB Endowment, 8(12):1590– 1601, August 2015. CODEN ???? ISSN 2150-8097.

Fiala:1989:DCF

[FG89] E. R. Fiala and D. H. Greene. Data compression with finite Windows. Communications of the Association for Computing Machinery, 32(4):490–505, April 1989. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/63341.html.

Ferragina:1995:OLS

[FG95a] P. Ferragina and R. Grossi. Optimal on-line search and sublinear time update in string matching. In IEEE [IEE95a], pages 604–612. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Ferragina:1995:FDD

[FG95b] Paolo Ferragina and Roberto Grossi. A fully-dynamic data structure for external substring search. In ACM [ACM95c], pages 693–702. ISBN 0-89791-718-9. LCCN QA 76.6

A13 1995. URL http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p693-ferragina/p693-ferragina.
pdf;   http://www.acm.org/pubs/citations/proceedings/
stoc/225058/p693-ferragina/. ACM order no. 508950.

Ferragina:1998:OLS

[FG98] Paolo Ferragina and Roberto Grossi. Optimal on-line search and sublinear time update in string matching. SIAM Journal on Computing, 27(3):713–736, June 1998. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/28611.

Ferragina:1999:SBT

[FG99] Paolo Ferragina and Roberto Grossi. The string B-tree: a new data structure for string search in external memory and its applications. Journal of the Association for Computing Machinery, 46(2):236–280, March 1999. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org:80/pubs/citations/journals/
jacm/1999-46-2/p236-ferragina/.

Ferragina:2008:SCS

[FGG^{+}08] Paolo Ferragina, Roberto Grossi, Ankur Gupta, Rahul Shah, and Jeffrey Scott Vitter. On searching compressed string collections cache-obliviously. In Lenzerini and Lembo [LL08], pages 181–190. ISBN 1-60560-932-3. LCCN ????

Flouri:2015:LCS

[FGKU15] Tomas Flouri, Emanuele Giaquinta, Kassian Kobert, and Esko Ukkonen. Longest common substrings with k mismatches. Information Processing Letters, 115(6–8):643– 647, June/August 2015. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019015000459.

Freiman:1972:IPP

[FGR72] C. V. Freiman, J. E. Griffith, and J. L. Rosenfeld, editors. Information processing 71: proceedings of IFIP Congress 71, Ljubjana, Yugoslavia, August 23–28, 1971, volume 1 of IFIP congress series. North-Holland, Amsterdam, The Netherlands, 1972. ISBN 0-7204-2063-6. LCCN QA75.5 .I532 1971.

URL https://dblp.org/db/conf/ifip/ifip71-1.html.

Furr:2009:PGS

[FhDAF09] Michael Furr, Jong hoon (David) An, and Jeffrey S. Foster. Profile-guided static typing for dynamic scripting languages. ACM SIGPLAN Notices, 44(10):283–300, October 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1992:ESE

[FHP92] Christopher W. Fraser, David R. Hanson, and Todd A. Proebsting. Engineering a simple, efficient code-generator generator. ACM Letters on Programming Languages and Systems, 1(3):213–226, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:/

/storage.webhop.net/documents/iburg.pdf;   http://
www.acm.org/pubs/toc/Abstracts/1057-4514/151642.html;
 http://www.cs.princeton.edu/software/iburg/.

Faust:2018:OPM

[FHV18] Sebastian Faust, Carmit Hazay, and Daniele Venturi. Outsourced pattern matching. International Journal of Information Security, 17(3):327–346, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL http://link.

springer.com/article/10.1007/s10207-017-0374-0.

Fischer:2010:PRE

[FHW10] Sebastian Fischer, Frank Huch, and Thomas Wilke. A play on regular expressions: functional pearl. ACM SIGPLAN Notices, 45(9):357–368, September 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filinski:2021:PDP

[Fil21]  Andrzej Filinski.  Proof-directed program transforma-

tion: a functional account of efficient regular expression matching. Journal of Functional Programming, 31 (??):e12, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/journals/journal-of-functional-
programming/article/proofdirected-program-transformation-
a-functional-account-of-efficient-regular-expression-
matching/454BB5CD9B0B056FA91957F2F9CC3EC5.

Finkel:1992:SAS

[FJ92] A. Finkel and M. Jantzen, editors. STACS 92. 9th Annual Symposium on Theoretical Aspects of Computer Science. Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

Fortnow:1996:RBI

[FK96] Lance Fortnow and Martin Kummer. On resource-bounded instance complexity. Theoretical Computer Science, 161 (1–2):123–140, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=161&issue=1-
2&aid=2038.

Fraigniaud:2016:OAL

[FK16] Pierre Fraigniaud and Amos Korman. An optimal ancestry labeling scheme with applications to XML trees and universal posets. Journal of the Association for Computing Machinery, 63(1):6:1–6:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Francez:1977:BRC

[FKP77] Nissim Francez, Boris Klebansky, and Amir Pnueli. Backtracking in recursive computations. Acta Informatica, 8(2): 125–144, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fagin:2013:SFF

[FKRV13] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Spanners: a formal framework for information extraction. In Hull and Fan [HF13], pages 37–48. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http:/

/dl.acm.org/citation.cfm?id=2463664;   http://www.
sigmod.org/2013/.

Fagin:2015:DSF

[FKRV15] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Document spanners: a formal approach to information extraction. Journal of the Association for Computing Machinery, 62(2):12:1–12:??, May 2015. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fagin:2016:DCI

[FKRV16] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Declarative cleaning of inconsistencies in information extraction. ACM Transactions on Database Systems, 41(1):6:1–6:??, April 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Farfeleder:2006:ECG

[FKSB06] Stefan Farfeleder, Andreas Krall, Edwin Steiner, and Florian Brandner. Effective compiler generation by architecture description. ACM SIGPLAN Notices, 41(7):145–152, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferguson:1971:GAL

[FL71] W. A. Ferguson and G. J. Lipovski. A generalized assembly language using regular expressions. ACM SIGPLAN Notices, 6(2):217–236, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferragina:1999:SSC

[FL99] Paolo Ferragina and Fabrizio Luccio. String search in coarsegrained parallel computers. Algorithmica, 24(3–4):177–194, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p177.html;
 http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=3&spage=177.

Ferragina:2008:CPM

[FL08] Paolo Ferragina and Gad M. Landau, editors. Combinatorial Pattern Matching: 19th Annual Symposium, CPM 2008, Pisa, Italy, June 18–20, 2008 Proceedings, volume 5029 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-69066-2 (print), 3-540-69068-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-540-69068-9.

Faro:2012:FSA

[FL12a] Simone Faro and Thierry Lecroq. A fast suffix automata based algorithm for exact online string matching. Lecture Notes in Computer Science, 7381:149–158, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_13/.

Faro:2012:MSW

[FL12b] Simone Faro and Thierry Lecroq. A multiple sliding Windows approach to speed up string matching algorithms. Lecture Notes in Computer Science, 7276:172–183, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30850-5_16/.

Faro:2013:EOS

[FL13] Simone Faro and Thierry Lecroq. The exact online string matching problem: a review of the most recent results. ACM Computing Surveys, 45(2):13:1–13:42, February 2013. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Flaherty:1988:STM

[Fla88] Terry Flaherty. A simple technique to motivate structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(1):153–155, February 1988. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fang:2019:EPM

[FLC^{+}19] Yixiang Fang, Yun Li, Reynold Cheng, Nikos Mamoulis, and Gao Cong. Evaluating pattern matching queries for spatial databases. VLDB Journal: Very Large Data Bases, 28(5): 649–673, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://link.

springer.com/article/10.1007/s00778-019-00550-3.

Fan:2010:GPM

[FLM^{+}10] Wenfei Fan, Jianzhong Li, Shuai Ma, Nan Tang, Yinghui Wu, and Yunpeng Wu. Graph pattern matching: from intractable to polynomial time. Proceedings of the VLDB Endowment, 3 (1–2):264–275, September 2010. CODEN ???? ISSN 2150-8097.

Florescu:1998:QCC

[FLS98] Daniela Florescu, Alon Levy, and Dan Suciu. Query containment for conjunctive queries with regular expressions. In ACM [ACM98], pages 139–148. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. URL http://

www.acm.org/pubs/articles/proceedings/pods/275487/
p139-florescu/p139-florescu.pdf;   http://www.acm.
org/pubs/citations/proceedings/pods/275487/p139-florescu/

. ACM order number 475980.

Fischetti:1993:CIP

[FLSS93a] Vincent A. Fischetti, Gad M. Landau, Jeanette P. Schmidt, and Peter H. Sellers. Corrigendum: “Identifying periodic occurrences of a template with applications to protein structure”. Information Processing Letters, 46(3):157, June 11, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [FLSS93b].

Fischetti:1993:IPO

[FLSS93b] Vincent A. Fischetti, Gad M. Landau, Peter H. Sellers, and Jeanette P. Schmidt. Identifying periodic occurrences of a template with applications to protein structure. Information Processing Letters, 45(1):11–18, January 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [FLSS93a].

Fredriksson:2006:EPS

[FM06] Kimmo Fredriksson and Maxim Mozgovoy. Efficient parameterized string matching. Information Processing Letters, 100 (3):91–96, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franklin:2002:PAS

[FMA02] Michael Franklin, Bongki Moon, and Anastassia Ailamaki, editors. Proceedings of the ACM SIGMOD International Conference on Management of Data, June 3–6, 2002, Madison, WI, USA. ACM Press, New York, NY 10036, USA, 2002. ISBN ???? LCCN ???? ACM order number 475020.

Ferragina:1999:MMD

[FMdB99] Paolo Ferragina, S. Muthukrishnan, and Mark de Berg. Multimethod dispatching: a geometric approach with applications to string matching problems. In ACM [ACM99b], pages 483–

491. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. URL

http://www.acm.org/pubs/articles/proceedings/stoc/
301250/p483-ferragina/p483-ferragina.pdf;   http:/
/www.acm.org/pubs/citations/proceedings/stoc/301250/

p483-ferragina/. ACM order number 508990.

Fuchs:2022:SUT

[FMG22] Per Fuchs, Domagoj Margan, and Jana Giceva. Sortledton: a universal, transactional graph data structure. Proceedings of the VLDB Endowment, 15(6):1173–1186, February 2022. CODEN ???? ISSN 2150-8097. URL https://dl.acm.org/

doi/10.14778/3514061.3514065.

Fuchs:2023:SUG

[FMG23] Per Fuchs, Domagoj Margan, and Jana Giceva. Sortledton: a universal graph data structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):17– 25, March 2023. CODEN SRECD8. ISSN 0163-5808 (print),

1943-5835 (electronic). URL https://dl.acm.org/doi/10.
1145/3604437.3604442.

Fernau:2020:PMV

[FMMS20] Henning Fernau, Florin Manea, Robert Mercas, and Markus L. Schmid. Pattern matching with variables: Efficient algorithms and complexity results. ACM Transactions on Computation Theory, 12(1):6:1–6:37, February 2020. CODEN ???? ISSN 1942-3454 (print), 1942-3462 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3369935.

Faro:2020:EOS

[FMP20] Simone Faro, Francesco Pio Marino, and Arianna Pavone. Efficient online string matching based on characters distance text sampling. Algorithmica, 82(11):3390–3412, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00732-4.

Fredriksson:2004:AOS

[FN04] Kimmo Fredriksson and Gonzalo Navarro. Average-optimal single and multiple approximate string matching. ACM Journal of Experimental Algorithmics, 9:1.4:1–1.4:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Faezipour:2009:HPE

[FNP09] Miad Faezipour, Mehrdad Nourani, and Rina Panigrahy. A hardware platform for efficient worm outbreak detection. ACM Transactions on Design Automation of Electronic Systems (TODAES), 14(4):49:1–49:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fredriksson:2002:OEF

[FNU02] Kimmo Fredriksson, Gonzalo Navarro, and Esko Ukkonen. Optimal exact and fast approximate two dimensional pattern matching allowing rotations. Lecture Notes in Computer Science, 2373:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2373/
23730235.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2373/23730235.pdf.

Fosdick:1976:DFA

[FO76] Lloyd D. Fosdick and Leon J. Osterweil. Data flow analysis in software reliability. ACM Computing Surveys, 8(3):305–330, September 1976. CODEN CMSVAN. ISSN 0010-4892.

Forest:2002:WCE

[For02] Julien Forest. A weak calculus with explicit operators for pattern matching and substitution. Lecture Notes in Computer Science, 2378:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2378/
23780174.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2378/23780174.pdf.

Foster:1989:ALF

[Fos89] M. J. Foster. Avoiding latch formation in regular expression recognizers. IEEE Transactions on Computers, 38(5):754– 756, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=24279.

Feuerstein:2008:OPS

[FPD08] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/ SQL language: pocket reference. O’Reilly Media, Sebastopol, CA, USA, fourth edition, 2008. ISBN 0-596-51404-2 (paperback). vii + 170 pp. LCCN QA76.9.D3. URL http://www.

oreilly.com/catalog/9780596514044.

Foster:2008:QL

[FPP08] J. Nathan Foster, Alexandre Pilkiewicz, and Benjamin C. Pierce. Quotient lenses. ACM SIGPLAN Notices, 43(9):383– 396, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fiori:2022:ASM

[FPT22] Fernando J. Fiori, Waltteri Pakalén, and Jorma Tarhio. Approximate string matching with SIMD. The Computer Journal, 65(6):1472–1488, June 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/article/65/6/1472/6134013.

Forax:2000:RTP

[FR00] Rémi Forax and Gilles Roussel. Recursive types and patternmatching in Java. Lecture Notes in Computer Science, 1799: 147–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/1799/17990147.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/1799/17990147.pdf.

Ferdous:2017:SMC

[FR17] S. M. Ferdous and M. Sohel Rahman. Solving the minimum common string partition problem with the help of ants. Mathematics in Computer Science, 11(2):233–249, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Fraser:1983:SFR

[Fra83] Christopher W. Fraser. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):281–283, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, GHF83a, GHF83b].

Franek:2020:CML

[Fra20] Frantisek Franek. Computing maximal Lyndon substrings of a string. Algorithms (Basel), 13(11):??, November 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https:/

/www.mdpi.com/1999-4893/13/11/294.

Freeman:1978:PCL

[Fre78] R. B. Freeman. Parse() — a C language routine to parse a string into words. Technical Memorandum 1271, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 15, 1978. ?? pp.

Fredriksson:2002:FSM

[Fre02] Kimmo Fredriksson. Faster string matching with superalphabets. Lecture Notes in Computer Science, 2476:44– ??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://link.springer.de/

link/service/series/0558/bibs/2476/24760044.htm;
http://link.springer.de/link/service/series/0558/papers/
2476/24760044.pdf.

Fredriksson:2003:SSM

[Fre03] Kimmo Fredriksson. Shift-or string matching with superalphabets. Information Processing Letters, 87(4):201–204, August 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2006:LAS

[Fre06] Kimmo Fredriksson. On-line approximate string matching in natural language. Fundamenta Informaticae, 72(4):453–466, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). URL http://content.iospress.

com/articles/fundamenta-informaticae/fi72-4-02;   http:
//www.cs.uku.fi/~fredriks/pub/papers/fi06.pdf.

Friedl:1997:MRE

[Fri97a] Jeffrey E. F. Friedl. Mastering regular expressions: powerful techniques for Perl and other tools. A Nutshell handbook. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 1997. ISBN 1-56592-257-3. xxiv + 342 pp. LCCN QA76.73.P22 F75 1997; QA76.9.D3G728 1997. US$29.95.

URL http://www.ora.com/catalog/regex/;   http://
www.oreilly.com/catalog/9781565922570;   http://www.
oreilly.com/catalog/regex.

Friesenhahn:1997:EOU

[Fri97b] Bob Friesenhahn. Expect offers Unix scripting — embeddable language offers pattern-matching capabilities. Byte Magazine, 22(4):63–??, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friedl:2002:MRE

[Fri02] Jeffrey E. F. Friedl. Mastering Regular Expressions. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, second edition, 2002. ISBN 0-596-00289-0. xxii + 460 pp. LCCN QA76.73.P22 F75 2002; QA76.9.T48 F75 2002. US$39.95. URL http:

//www.oreilly.com/catalog/9780596002893;   http://
www.oreilly.com/catalog/regex2.

Friedl:2006:MRE

[Fri06a] Jeffrey E. F. Friedl. Mastering regular expressions. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, third edition, 2006. ISBN 0-596-52812-4 (paperback). xxiv + 515 pp. LCCN QA76.9.T48 F75 2006. URL http://

www.loc.gov/catdir/enhancements/fy0715/2007272426-
d.html;   http://www.oreilly.com/catalog/9780596528126.

Frisch:2006:OX

[Fri06b] Alain Frisch. OCaml + XDuce. ACM SIGPLAN Notices, 41 (9):192–200, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Florenzano:2020:EEA

[FRU^{+}20] Fernando Florenzano, Cristian Riveros, Martín Ugarte, Stijn Vansummeren, and Domagoj Vrgoc. Efficient enumeration algorithms for regular document spanners. ACM Transactions on Database Systems, 45(1):3:1–3:42, March 2020. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3351451.

Feder:2013:ECA

[FS13] Tomás Feder and Carlos Subi. Edge-coloring almost bipartite multigraphs. Information Processing Letters, 113(18): 685–689, September 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019013001737.

Freydenberger:2019:DRE

[FS19] Dominik D. Freydenberger and Markus L. Schmid. Deterministic regular expressions with back-references. Journal of Computer and System Sciences, 105(??):1–39, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-

2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022000018301818.

Faro:2022:WAS

[FS22] Simone Faro and Stefano Scafiti. A weak approach to suffix automata simulation for exact and approximate string matching. Theoretical Computer Science, 933(??):88–103, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397522005138.

Feng:2015:RES

[FSL^{+}15] Weixing Feng, Peichao Sang, Deyuan Lian, Yansheng Dong, Fengfei Song, Meng Li, Bo He, Fenglin Cao, and Yunlong Liu. ResSeq: enhancing short-read sequencing alignment by rescuing error-containing reads. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(4):795–798, July 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Farach:1995:SML

[FT95] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. In ACM [ACM95c], pages 703–712. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/pubs/articles/proceedings/
stoc/225058/p703-farach/p703-farach.pdf;   http://
www.acm.org/pubs/citations/proceedings/stoc/225058/

p703-farach/. ACM order no. 508950.

Farach:1998:SML

[FT98] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. Algorithmica, 20(4):388–404, April 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=20&
issue=4&spage=388.

Fredriksson:2004:ESM

[FT04] Kimmo Fredriksson and Jorma Tarhio. Efficient string matching in Huffman compressed texts. Fundamenta Informaticae, 63(1):1–16, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fricker:1995:ICI

[FTJ95] Christine Fricker, Olivier Temam, and William Jalby. Influence of cross-interferences on blocked loops: a case study with matrix-vector multiply. ACM Transactions on Programming Languages and Systems, 17(4):561–575, July 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/210185.html.

Floyd:1982:CRE

[FU82] Robert W. Floyd and Jeffrey D. Ullman. The compilation of regular expressions into integrated circuits. Journal of the Association for Computing Machinery, 29(3):603–622, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fu:1995:PMD

[Fu95] J. Fu. Pattern matching in directed graphs. Lecture Notes in Computer Science, 937:64–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fu:1996:APM

[Fu96] J. J. Fu. Approximate pattern matching in directed graphs. Lecture Notes in Computer Science, 1075:373–??, 1996. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fu:1997:DGP

[Fu97] James Jianghai Fu. Directed graph pattern matching and topological embedding. Journal of Algorithms, 22(2):372–391, February 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0196677496908184.

Fredriksson:1998:RIF

[FU98] K. Fredriksson and E. Ukkonen. A rotation invariant filter for two-dimensional string matching. Lecture Notes in Computer Science, 1448:118–??, 1998. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ferragina:2016:CCO

[FV16] Paolo Ferragina and Rossano Venturini. Compressed cacheoblivious string B-tree. ACM Transactions on Algorithms, 12 (4):52:1–52:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, D. Gries, and J. Misra, editors. Beauty is our business: a birthday salute to Edsger W. Dijkstra. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326 1990.

Feng:2015:EQD

[FWDL15] Yu Feng, Xinyu Wang, Isil Dillig, and Calvin Lin. EX-PLORER : query- and demand-driven exploration of interprocedural control flow properties. ACM SIGPLAN Notices, 50(10):520–534, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fan:2012:PGD

[FWW12] Wenfei Fan, Xin Wang, and Yinghui Wu. Performance guarantees for distributed reachability queries. Proceedings of the VLDB Endowment, 5(11):1304–1316, July 2012. CODEN ???? ISSN 2150-8097.

Fan:2013:DTG

[FWW13a] Wenfei Fan, Xin Wang, and Yinghui Wu. Diversified top-k graph pattern matching. Proceedings of the VLDB Endowment, 6(13):1510–1521, August 2013. CODEN ???? ISSN 2150-8097.

Fan:2013:DTK

[FWW13b] Wenfei Fan, Xin Wang, and Yinghui Wu. Diversified top-k graph pattern matching. Proceedings of the VLDB Endowment, 6(13):1510–1521, August 2013. CODEN ???? ISSN 2150-8097.

Fan:2013:IGP

[FWW13c] Wenfei Fan, Xin Wang, and Yinghui Wu. Incremental graph pattern matching. ACM Transactions on Database Systems, 38(3):18:1–18:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Fang:2017:SPM

[FYJ^{+}17] Yan Fang, Victor V. Yashin, Brandon B. Jennings, Donald M. Chiarulli, and Steven P. Levitan. A simplified phase model for simulation of oscillator-based computing systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):14:1–14:??, March 2017. CODEN ???? ISSN 1550-4832.

Gaal:2004:DSF

[Gaá04] Tamás Gaál. Deciding sequentiability of finite-state transducers by finite-state pattern-matching. Theoretical Computer Science, 313(1):105–117, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1975:CLA

[Gal75] Zvi Galil. On converting on-line algorithms into real-time and on real-time algorithms for string-matching and palindrome recognition. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 7(4):26–30, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1976:RTA

[Gal76a] Zvi Galil. Real-time algorithms for string-matching and palindrome recognition. In ACM [ACM76], pages 161–173. LCCN QA 76.6 A12 1976.

Galil:1976:TFS

[Gal76b] Zvi Galil. Two fast simulations which imply some fast string matching and palindrome-recognition algorithms. Information Processing Letters, 4(4):85–87, January ??, 1976. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1979:IWC

[Gal79] Zvi Galil. On improving the worse case running time of the Boyer–Moore string matching algorithm. Communications of the Association for Computing Machinery, 22(9):505–508, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Galil:1981:SMR

[Gal81] Z. Galil. String matching in real time. Journal of the Association for Computing Machinery, 28(1):134–149, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1984:OPA

[Gal84] Zvi Galil. Optimal parallel algorithms for string matching. In ACM [ACM84], pages 240–248. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Galil:1992:CTO

[Gal92] Zvi Galil. A constant-time optimal parallel string-matching algorithm. In ACM [ACM92d], pages 69–76. ISBN 0-89791-511-9. LCCN QA76.A15 1992. URL http://www.

acm.org/pubs/articles/proceedings/stoc/129712/p69-
galil/p69-galil.pdf;   http://www.acm.org/pubs/citations/
proceedings/stoc/129712/p69-galil/. ACM order no.

508920.

Galil:1995:CTO

[Gal95] Zvi Galil. A constant-time optimal parallel string-matching algorithm. Journal of the Association for Computing Machinery, 42(4):908–918, July 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/210341.html.

Ganapathi:1989:PBR

[Gan89a] Mahadevan Ganapathi. Prolog based retargetable code generation. Computer Languages, 14(3):193–204, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ganapathi:1989:SPP

[Gan89b] Mahadevan Ganapathi. Semantic predicates in parser generators. Computer Languages, 14(1):25–33, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gonzalez-Alvarez:2013:AAD

[GÁSÁ^{+}13] Cecilia González-Álvarez, Jennifer B. Sartor, Carlos Álvarez, Daniel Jiménez-González, and Lieven Eeckhout. Accelerating an application domain with specialized functional units. ACM Transactions on Architecture and Code Optimization, 10(4): 47:1–47:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gawrychowski:2012:SEL

[Gaw12] Paweł Gawrychowski. Simple and efficient LZW-compressed multiple pattern matching. Lecture Notes in Computer Science, 7354:232–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_19/.

Gawrychowski:2013:OPM

[Gaw13] Pawel Gawrychowski. Optimal pattern matching in LZW compressed strings. ACM Transactions on Algorithms, 9(3): 25:1–25:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gonnet:1990:AKR

[GBY90a] Gaston H. Gonnet and Ricardo A. Baeza-Yates. An analysis of the Karp-Rabin string matching algorithm. Information Processing Letters, 34(5):271–274, May 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonnet:1990:AKS

[GBY90b] Gaston H. Gonnet and Ricardo A. Baeza-Yates. An analysis of the Karp-Rabin string matching algorithm. Information Processing Letters, 34(5):271–274, May 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garai:2001:CGA

[GC01] Gautam Garai and B. B. Chaudhuri. A cascaded genetic algorithm for efficient optimization and pattern matching. Lecture Notes in Computer Science, 2013:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2013/20130032.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2013/20130032.pdf.

Geffert:2003:TBR

[Gef03] Viliam Geffert. Translation of binary regular expressions into nondeterministic ϵ-free automata with O(n log n) transitions. Journal of Computer and System Sciences, 66(3):451– 472, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022000003000369.

Gelade:2010:SRE

[Gel10] Wouter Gelade. Succinctness of regular expressions with interleaving, intersection and counting. Theoretical Computer Science, 411(31–33):2987–2998, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grabowski:2008:BPS

[GF08] Szymon Grabowski and Kimmo Fredriksson. Bit-parallel string matching under Hamming distance in O(n⌈m/w⌉) worst case time. Information Processing Letters, 105(5):182– 187, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grabowski:2011:SMI

[GFG11] Szymon Grabowski, Simone Faro, and Emanuele Giaquinta. String matching with inversions and translocations in linear average time (most of the time). Information Processing Letters, 111(11):516–520, May 15, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ganapathi:1982:RCC

[GFH82] Mahadevan Ganapathi, Charles N. Fischer, and John L. Hennessy. Retargetable compiler code generation. ACM Computing Surveys, 14(4):573–592, December 1982. CODEN CMSVAN. ISSN 0010-4892. See also [WNL^{+}83, GHF83a, Fra83, GHF83b].

Galil:1986:ISM

[GG86] Z. Galil and R. Giancarlo. Improved string matching with k mismatches. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 17(4):52–54, Spring 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1987:PSM

[GG87] Z. Galil and R. Giancarlo. Parallel string matching with k mismatches. Theoretical Computer Science, 51(3):341–348, ???? 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1991:ECS

[GG91] Zvi Galil and Raffaele Giancarlo. On the exact complexity of string matching: Lower bounds. SIAM Journal on Computing, 20(6):1008–1020, December 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1992:ECS

[GG92] Zvi Galil and Raffaele Giancarlo. On the exact complexity of string matching: Upper bounds. SIAM Journal on Computing, 21(3):407–437, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Giancarlo:1995:MDP

[GG95] R. Giancarlo and R. Grossi. Multi-dimensional pattern matching with dimensional wildcards. Lecture Notes in Computer Science, 937:90–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giancarlo:1997:MDP

[GG97] Raffaele Giancarlo and Roberto Grossi. Multi-dimensional pattern matching with dimensional wildcards: Data structures and optimal on-line search algorithms. Journal of Algorithms, 24(2):223–265, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0196677496908445.

Giaquinta:2013:NAB

[GG13] Emanuele Giaquinta and Szymon Grabowski. New algorithms for binary jumbled pattern matching. Information Processing Letters, 113(14–16):538–542, July/August 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019013001300.

Giaquinta:2013:APM

[GGF13] Emanuele Giaquinta, Szymon Grabowski, and Kimmo Fredriksson. Approximate pattern matching with kmismatches in packed text. Information Processing Letters, 113(19–21):693–697, September/October 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019013001919.

Gereb-Graus:1994:TOW

[GGL94] Mihály Geréb-Graus and Ming Li. Three one-way heads cannot do string matching. Journal of Computer and System Sciences, 48(1):1–8, February 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000005800200.

Gelade:2012:REC

[GGM12] Wouter Gelade, Marc Gyssens, and Wim Martens. Regular expressions with counting: Weak versus strong determinism. SIAM Journal on Computing, 41(1):160–190, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.siam.org/sicomp/resource/
1/smjcat/v41/i1/p160_s1.

Gramm:2006:PMA

[GGN06] Jens Gramm, Jiong Guo, and Rolf Niedermeier. Pattern matching for arc-annotated sequences. ACM Transactions on Algorithms, 2(1):44–65, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Guoan:1982:USM

[GH82] Gu Guoan and John Hobby. Using string matching to compress Chinese characters. Report STAN-CS-82-914, Stanford University, Department of Computer Science, Stanford, CA, USA, 1982.

Gruber:2009:LOR

[GH09] Hermann Gruber and Markus Holzer. Language operations with regular expressions of polynomial size. Theoretical Computer Science, 410(35):3281–3289, August 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gruber:2013:PSR

[GH13] Hermann Gruber and Markus Holzer. Provably shorter regular expressions from finite automata. International Journal of Foundations of Computer Science (IJFCS), 24(8):1255–??, December 2013. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gruber:2015:FAR

[GH15] Hermann Gruber and Markus Holzer. From finite automata to regular expressions and back — a summary on descriptional complexity. International Journal of Foundations of Computer Science (IJFCS), 26(8):1009–??, December 2015.

CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ganapathi:1983:SFRa

[GHF83a] M. Ganapathi, J. L. Hennessy, and C. N. Fischer. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):280–281, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, Fra83, GHF83b].

Ganapathi:1983:SFRb

[GHF83b] M. Ganapathi, J. L. Hennessy, and C. N. Fischer. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):283–284, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, GHF83a, Fra83].

Ghiron:1962:RMR

[Ghi62] Hugo Ghiron. Rules to manipulate regular expressions of finite automata. IRE Transactions on Electronic Computers, EC-11 (4):574–575, August 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5219404.

Gokhale:1991:BUH

[GHK^{+}91] Maya Gokhale, William Holmes, Andrew Kopser, Sara Lucas, Ronald Minnich, Douglas Sweely, and Daniel Lopresti. Building and using a highly parallel programmable logic array. Computer, 24(1):81–89, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guo:2014:LSS

[GHK14] Jiong Guo, Danny Hermelin, and Christian Komusiewicz. Local search for string problems: Brute-force is essentially optimal. Theoretical Computer Science, 525(??):30–41, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397513003575.

Ganardi:2018:CEF

[GHKL18] Moses Ganardi, Danny Hucke, Daniel König, and Markus Lohrey. Circuits and expressions over finite semirings. ACM Transactions on Computation Theory, 10(4):15:1–15:??, October 2018. CODEN ???? ISSN 1942-3454 (print), 1942-3462 (electronic).

Gagie:2015:BJP

[GHLW15] Travis Gagie, Danny Hermelin, Gad M. Landau, and Oren Weimann. Binary jumbled pattern matching on trees and tree-like structures. Algorithmica, 73(3):571–588, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/
article/10.1007/s00453-014-9957-6.

Grathwohl:2014:CCR

[GHR14] Niels Bjørn Bugge Grathwohl, Fritz Henglein, and Ulrik Terp Rasmussen. A crash-course in regular expression parsing and regular expressions as types. Technical report ??, Department of Computer Science (DIKU), University of Copenhagen, Copenhagen, Denmark, August 21, 2014. 40 pp. URL https:

//di.ku.dk/kmc/documents/AiPL-CrashCourse.pdf.

Grathwohl:2016:KCN

[GHR^{+}16] Bjørn Bugge Grathwohl, Fritz Henglein, Ulrik Terp Rasmussen, Kristoffer Aalund Søholm, and Sebastian Paaske Tørholm. Kleenex: compiling nondeterministic transducers to deterministic streaming transducers. ACM SIGPLAN Notices, 51(1):284–297, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1982:ETD

[GHS82] Susan L. Graham, Robert R. Henry, and Robert A. Schulman. An experiment in table driven code generation. ACM SIGPLAN Notices, 17(6):32–43, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gulwani:2012:SDM

[GHS12] Sumit Gulwani, William R. Harris, and Rishabh Singh. Spreadsheet data manipulation using examples. Communications of the Association for Computing Machinery, 55(8): 97–105, August 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ganguly:2017:STT

[GHST17] Arnab Ganguly, Wing-Kai Hon, Rahul Shah, and Sharma V. Thankachan. Space-time trade-offs for finding shortest unique substrings and maximal unique matches. Theoretical Computer Science, 700(??):75–88, November 14, 2017. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397517305832.

Geser:2005:TPS

[GHW05] Alfons Geser, Dieter Hofbauer, and Johannes Waldmann. Termination proofs for string rewriting systems via inverse match-bounds. Journal of Automated Reasoning, 34(4):365– 385, May 2005. CODEN JAREEW. ISSN 0168-7433 (print),

1573-0670 (electronic). URL http://link.springer.com/
article/10.1007/s10817-005-9024-8.

Giancarlo:1993:IDS

[Gia93] R. Giancarlo. An index data structure for matrices, with applications to fast two-dimensional pattern matching. Lecture Notes in Computer Science, 709:337–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gibney:2021:TIR

[Gib21] Daniel Gibney. Text indexing for regular expression matching. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/133.

Giegerich:1990:CSI

[Gie90] R. Giegerich. Code selection by inversion of order-sorted derivors. Theoretical Computer Science, 73(2):177–211, June 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goto:1977:PHA

[GIG77] E. Goto, T. Ida, and T. Gunji. Parallel hashing algorithms. Information Processing Letters, 6(1):8–13, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasieniec:1997:EIP

[GIK97] L. Gasieniec, P. Indyk, and P. Krysta. External inverse pattern matching. Lecture Notes in Computer Science, 1264:90– ??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gill:1970:SAR

[Gil70] A. Gill. Symmetric and antisymmetric regular expressions. SIAM Journal on Applied Mathematics, 18(3):539–557, May 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gillogly:1985:FPM

[Gil85] James J. Gillogly. Fast pattern matching for word lists. Cryptologia, 9(1):55–62, January 1985. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http://www.

informaworld.com/smpp/content~content=a741902694~db=
all~order=page.

Gimpel:1973:TDP

[Gim73] James F. Gimpel. A theory of discrete patterns and their implementation in SNOBOL4. Communications of the Association for Computing Machinery, 16(2):91–100, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gavrilov:2003:CEM

[GIMV03] Martin Gavrilov, Piotr Indyk, Rajeev Motwani, and Suresh Venkatasubramanian. Combinatorial and experimental methods for approximate point pattern matching. Algorithmica, 38(1):59–90, October 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=1&spage=59.

Ginzburg:1967:PCE

[Gin67] A. Ginzburg. A procedure for checking equality of regular expressions. Journal of the Association for Computing Machinery, 14(2):355–362, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gittleman:1996:PSS

[Git96] Arthur Gittleman. Predicting string search speed. ACM Journal of Experimental Algorithmics, 1:2:1–2:??, ???? 1996. CO-DEN ???? ISSN 1084-6654.

Garhwal:2016:PFR

[GJ16] Sunita Garhwal and Ram Jiwari. Parallel fuzzy regular expression and its conversion to epsilon-free fuzzy automaton. The Computer Journal, 59(9):1383–1391, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/59/9/1383.

Graf:2020:LYG

[GJS20] Sebastian Graf, Simon Peyton Jones, and Ryan G. Scott. Lower your guards: a compositional pattern-match coverage checker. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):107:1–107:30, August 2020.

URL https://dl.acm.org/doi/10.1145/3408989.

Grigoriadis:1986:LBC

[GK86] M. D. Grigoriadis and B. Kalantari. A lower bound to the complexity of Euclidean and rectilinear matching algorithms. Information Processing Letters, 22(2):73–76, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golan:2019:SPM

[GKP19] Shay Golan, Tsvi Kopelowitz, and Ely Porat. Streaming pattern matching with d wildcards. Algorithmica, 81(5):1988– 2015, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Guthle:2010:IAD

[GKW^{+}10] Martin Güthle, Jochen Kögel, Stefan Wahl, Matthias Kaschub, and Christian M. Mueller. Improving anomaly detection for text-based protocols by exploiting message structures. Future Internet, 2(4):662–669, December 21, 2010. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/2/4/662.

Grosky:1986:IIU

[GL86] William I. Grosky and Yi Lu. Iconic indexing using generalized pattern matching techniques. Computer Vision, Graphics, and Image Processing, 35(3):383–403, September 1986. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Grossi:1989:SES

[GL89] R. Grossi and F. Luccio. Simple and efficient string matching with k mismatches. Information Processing Letters, 33(3): 113–120, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grobauer:2001:PEP

[GL01] Bernd Grobauer and Julia L. Lawall. Partial evaluation of pattern matching in strings, revisited. Nordic Journal of Computing, 8(4):437–462, Winter 2001. CODEN NJCOFR.

ISSN 1236-6064. URL http://www.cs.helsinki.fi/njc/
References/grobauer2001:437.html.

Gennick:2003:ORE

[GL03] Jonathan Gennick and Peter Linsley. Oracle regular expressions: pocket reference. O’Reilly Media, Sebastopol, CA, USA, 2003. ISBN 0-596-00601-2. iv + 60 pp. LCCN

QA76.9.T48 G46 2003. URL http://www.oreilly.com/
catalog/9780596006013.

Goyvaerts:2012:REC

[GL12] Jan Goyvaerts and Steven Levithan. Regular expressions cookbook. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, second edition, 2012. ISBN 1-4493-1943-2 (paperback). xiv + 594 pp. LCCN NLS PB8.212.757/6.

Ganardi:2019:UTB

[GL19] Moses Ganardi and Markus Lohrey. A universal tree balancing theorem. ACM Transactions on Computation Theory, 11 (1):1:1–1:??, January 2019. CODEN ???? ISSN 1942-3454 (print), 1942-3462 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3278158.

Garcia:2011:RES

[GLRÁ11] Pedro García, Damián López, José Ruiz, and Gloria I. Álvarez. From regular expressions to smaller NFAs. Theoretical Computer Science, 412(41):5802–5807, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gusfield:1992:EAA

[GLS92] Dan Gusfield, Gad M. Landau, and Baruch Schieber. An efficient algorithm for the All Pairs Suffix — Prefix Problem. Information Processing Letters, 41(4):181–185, March 18, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geneves:2007:ESA

[GLS07] Pierre Genevès, Nabil Layaïda, and Alan Schmitt. Efficient static analysis of XML paths and types. ACM SIGPLAN Notices, 42(6):342–351, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Genest:2002:PMM

[GM02] Blaise Genest and Anca Muscholl. Pattern matching and membership for hierarchical message sequence charts. Lecture Notes in Computer Science, 2286:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2286/22860326.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2286/22860326.pdf.

Giancarlo:2011:CPM

[GM11] Raffaele Giancarlo and Giovanni Manzini, editors. Combinatorial Pattern Matching: 22nd Annual Symposium, CPM 2011, Palermo, Italy, June 27–29, 2011. Proceedings, volume 6661 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-21457-6 (print), 3-642-21458-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-21458-5.

Groz:2017:ETM

[GM17] B. Groz and S. Maneth. Efficient testing and matching of deterministic regular expressions. Journal of Computer and System Sciences, 89(??):372–399, November 2017. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000017300843.

Gudur:2022:FBE

[GMAS22] Venkateshwarlu Yellaswamy Gudur, Sidharth Maheshwari, Amit Acharyya, and Rishad Shafik. An FPGA based energy-efficient read mapper with parallel filtering and insitu verification. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 19(5):2697–2711, 2022. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TCBB.
2021.3106311.

Giavitto:2002:PMR

[GMC02] Jean-Louis Giavitto, Olivier Michel, and Julien Cohen. Pattern-matching and rewriting rules for group indexed data structures. ACM SIGPLAN Notices, 37(12):76–87, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grozea:2012:SMI

[GMMN12] Cristian Grozea, Florin Manea, Mike Müller, and Dirk Nowotka. String matching with involutions. Lecture Notes in Computer Science, 7445:106–117, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32894-7_11/.

Groz:2012:DRE

[GMS12] Benotît Groz, Sebastian Maneth, and Slawek Staworko. Deterministic regular expressions in linear time. In Krötzsch et al. [KLB12], pages 49–60. ISBN ???? LCCN ???? URL

http://www.sigmod.org/2012/.

Golomshtok:2001:FTS

[GN01] Alex Golomshtok and Yefim Nodelman. Fuzzy text searches with agrep and afind. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):14, 16, 18, 20, 22, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Gelade:2012:SCI

[GN12] Wouter Gelade and Frank Neven. Succinctness of the complement and intersection of regular expressions. ACM Transactions on Computational Logic, 13(1):4:1–4:??, January 2012. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Ghanbarpour:2019:MBK

[GN19] Asieh Ghanbarpour and Hassan Naderi. A model-based keyword search approach for detecting top-k effective answers. The Computer Journal, 62(3):377–393, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.oup.com/comjnl/article/
62/3/377/5033366.

Grahne:1994:RAS

[GNU94] Gösta Grahne, Matti Nykänen, and Esko Ukkonen. Reasoning about strings in databases. In ACM [ACM94c], pages 303–312. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

URL http://www.acm.org/pubs/articles/proceedings/
pods/182591/p303-grahne/p303-grahne.pdf;   http://
www.acm.org/pubs/citations/proceedings/pods/182591/
p303-grahne/;   http://www.acm.org:80/pubs/citations/
proceedings/pods/182591/p303-grahne/.

Guibas:1980:NPL

[GO80] Leo J. Guibas and Andrew M. Odlyzko. A new proof of the linearity of the Boyer–Moore string searching algorithm. SIAM Journal on Computing, 9(4):672–682, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grossi:2012:WTM

[GO12] Roberto Grossi and Giuseppe Ottaviano. The wavelet trie: maintaining an indexed sequence of strings in compressed space. In Krötzsch et al. [KLB12], pages 203–214. ISBN ????

LCCN ???? URL http://www.sigmod.org/2012/.

Goedelbecker:1995:UG

[Goe95] Eric Goedelbecker. Using grep. Linux Journal, 18:??, October 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Goldschlag:1990:MVC

[Gol90] D. M. Goldschlag. Mechanically verifying concurrent programs with the Boyer–Moore prover. IEEE Transactions on Software Engineering, 16(9):1005–1023, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=58787.

Goldberg:1993:FSA

[Gol93] Robert R. Goldberg. Finite state automata from regular expression trees. The Computer Journal, 36(7):623– 630, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/36/7/623.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/Volume_36/
Issue_07/Vol36_07.body.html#AbstractGoldberg.

Garcia-Osorio:2008:TPA

[GOMSJVGP08] César García-Osorio, Iñigo Mediavilla-Sáiz, Javier Jimeno-Visitación, and Nicolás García-Pedrajas. Teaching push-down automata and Turing machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 316, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Gonnet:1983:UDB

[Gon83] Gaston H. Gonnet. Unstructured data bases or very efficient text searching. In ACM-PODS’83 [ACM83], pages 117–124. ISBN 0-89791-097-4. LCCN QA76.9.D3 A15 1983. US$25.00. ACM (order n 475830). Baltimore, MD, USA.

Gonnet:2002:SMP

[Gon02] Gaston H. Gonnet. String matching problems from bioinformatics which still need better solutions extended abstract. Lecture Notes in Computer Science, 2476:27–??, 2002. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://link.springer.de/link/service/

series/0558/bibs/2476/24760027.htm;   http://link.
springer.de/link/service/series/0558/papers/2476/24760027.
pdf.

Good:2005:RER

[Goo05] Nathan A. Good. Regular expression recipes: a problemsolution approach. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-441-X (paperback). xxix + 289 pp. LCCN QA76.9.T48 G66 2005.

Gorman:2000:PCT

[Gor00] Ian E. Gorman. Parsing complex text structures. Dr. Dobb’s Journal of Software Tools, 25(6):90, 92–98, June 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2000/2000_06/parse.txt;   http://
www.ddj.com/ftp/2000/2000_06/parse.zip.

Galil:1990:IAA

[GP90] Zvi Galil and Kunsoo Park. An improved algorithm for approximate string matching. SIAM Journal on Computing, 19 (6):989–999, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1992:TAI

[GP92] Z. Galil and K. Park. Truly alphabet-independent twodimensional pattern matching. In IEEE [IEE92], pages 247–

256. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Gemis:1993:OOP

[GP93] Marc Gemis and Jan Paredaens. An object-oriented pattern matching language. Lecture Notes in Computer Science, 742: 339–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasieniec:2001:TSE

[GP01] Leszek Gasieniec and Igor Potapov. Time/space efficient compressed pattern matching. Lecture Notes in Computer Science, 2138:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2138/

21380138.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2138/21380138.pdf.

Gasieniec:2003:TSE

[GP03] Leszek Gasieniec and Igor Potapov. Time/space efficient compressed pattern matching. Fundamenta Informaticae, 56(1–2): 137–154, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gronlund:2018:TDL

[GP18] Allan Grønlund and Seth Pettie. Threesomes, degenerates, and love triangles. Journal of the Association for Computing Machinery, 65(4):22:1–22:??, August 2018. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gostanza:1996:NLP

[GPN96] Pedro Palao Gostanza, Ricardo Peña, and Manuel Núñez. A new look at pattern matching in abstract data types. ACM SIGPLAN Notices, 31(6):110–121, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Galil:2004:TDP

[GPP04] Zvi Galil, Jong Geun Park, and Kunsoo Park. Threedimensional periodicity and its application to pattern matching. SIAM Journal on Discrete Mathematics, 18(2):362–381, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146

(electronic). URL http://epubs.siam.org/sam-bin/dbq/
article/39030.

Gasieniec:1995:CSS

[GPR95a] L. Gasieniec, W. Plandowski, and W. Rytter. Constant-space string matching with smaller number of comparisons: Sequential sampling. Lecture Notes in Computer Science, 937:78–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasieniec:1995:ZMR

[GPR95b] Leszek Gasieniec, Wojciech Plandowski, and Wojciech Rytter. The zooming method: a recursive approach to timespace efficient string-matching. Theoretical Computer Science, 147(1–2):19–30, August 07, 1995. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1995&volume=147&issue=1-
2&aid=1883.

Gemis:1993:GGO

[GPTV93] Marc Gemis, Jan Paredaens, Inge Thyssens, and Jan Van den Bussche. GOOD: a graph-oriented object database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):505–510, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garg:1992:CRE

[GR92] V. K. Garg and M. T. Ragunath. Concurrent regular expressions and their relationship to Petri nets. Theoretical Computer Science, 96(2):285–304, April 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guglielmo:1996:NLR

[GR96] Eugene J. Guglielmo and Neil C. Rowe. Natural-language retrieval of images based on descriptive captions. ACM Transactions on Information Systems, 14(3):237–267, July 1996. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.

org:80/tois/abstracts/guglielmo.html.

Gasieniec:1999:AOF

[GR99] L. Gasieniec and W. Rytter. Almost-optimal fully LZWcompressed pattern matching. In Storer and Cohn [SC99], pages 316–325. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1999.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=755681. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Grabowski:2015:NLC

[Gra15] Szymon Grabowski. A note on the longest common substring with k-mismatches problem. Information Processing Letters, 115(6–8):640–642, June/August 2015. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019015000411.

Greenwood:1988:VSR

[Gre88] S. R. Greenwood. VAX SCAN: rule-based text processing software. Digital Technical Journal of Digital Equipment Corporation, 1(6):40–50, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Griss:1979:HKR

[Gri79] M. L. Griss. Hashing—the key to rapid pattern matching. Technical Report UUCS-79-108, Computer Science Department, University of Utah, 1979. ?? pp.

Griswold:1983:ISP

[Gri83] Ralph E. Griswold. Implementing Snobol4 pattern matching in Icon. Computer Languages, 8(2):77–92, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Griswold:1985:RSI

[Gri85] Ralph E. Griswold. Rebus — a SNOBOL4/Icon hybrid. ACM SIGPLAN Notices, 20(2):7–16, February 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grossi:1991:FCS

[Gro91a] Roberto Grossi. Further comments on the subtree isomorphism for ordered trees: “On the subtree isomorphism problem for ordered trees” [Inform. Process. Lett. 32 (1989), no. 5, 271–273; MR 90k:68139] by E. Mäkinen. Information Processing Letters, 40(5):255–256, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mäk89].

Grossi:1991:NSI

[Gro91b] Roberto Grossi. A note on the subtree isomorphism for ordered trees and related problems. Information Processing Letters, 39(2):81–84, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grosch:1992:TAT

[Gro92] Josef Grosch. Transformation of attributed trees using pattern matching. Lecture Notes in Computer Science, 641:1–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garofalakis:1999:SSP

[GRS99] Minos N. Garofalakis, Rajeev Rastogi, and Kyuseok Shim. SPIRIT: Sequential pattern mining with regular expression constraints. In Atkinson et al. [AOV^{+}99], pages 223–234. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/dblp/db/conf/vldb/
GarofalakisRS99.html. Also known as VLDB’99.

Galil:1980:SSF

[GS80] Zvi Galil and Joel Seiferas. Saving space in fast stringmatching. SIAM Journal on Computing, 9(2):417–438, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1981:LTS

[GS81a] Z. Galil and J. Seireras. Linear-time string-matching using only a fixed number of local storage locations. Theoretical Computer Science, 13(3):331–336, March 1981. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1981:TSO

[GS81b] Zvi Galil and Joel Seiferas. Time-space-optimal string matching (preliminary report). In ACM [ACM81], pages 106–113. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Galil:1983:TSO

[GS83] Zvi Galil and Joel Seiferas. Time-space-optimal string matching. Journal of Computer and System Sciences, 26(3):280– 294, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0022000083900028.

GonzalezSmith:1985:PAD

[GS85] M. E. Gonzalez Smith and J. A. Storer. Parallel algorithms for data compression. Journal of the Association for Computing Machinery, 32(2):344–373, April 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3152.html.

Ganesan:1993:STL

[GS93a] Ravi Ganesan and Alan T. Sherman. Statistical techniques for language recognition: An introduction and guide for cryptanalysts. Cryptologia, 17(4):321–366, October 1993. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586

(electronic). URL http://www.informaworld.com/smpp/
content~content=a748639241~db=all~order=page. Pre-

liminary version available as Technical Report CS-TR-3036/UMIACS-TR-93-16, University of Maryland College Park (February 1993), and as Technical Report TR CS-93-02, University of Maryland Baltimore County (February 28, 1993).

Gokhale:1993:DBC

[GS93b] Maya B. Gokhale and Judith D. Schlesinger. A data-parallel bit-serial C (dbC). Technical report SRC-TR-93-096, Supercomputing Research Center: IDA, Lanham, MD, USA, May 1993. 14 pp.

Gokhale:1993:DPB

[GS93c] Maya B. Gokhale and Judith D. Schlesinger. A data-parallel bit-serial C (dbC). Technical report SRC-TR-93-096, Supercomputing Research Center: IDA, Lanham, MD, USA, May 1993. 14 pp.

Giancarlo:2000:CPM

[GS00] Raffaele Giancarlo and David Sankoff, editors. Combinatorial pattern matching: 11th annual symposium, CPM 2000, Montréal, Canada, June 21–23, 2000: Proceedings, volume 1848 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNCSD9. ISBN 3-540-67633-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1

L43 no.1848. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1848.htm;   http://www.springerlink.
com/content/978-3-540-67633-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
1848.

Gustafsson:2006:EMB

[GS06] Per Gustafsson and Konstantinos Sagonas. Efficient manipulation of binary data using pattern matching. Journal of Functional Programming, 16(1):35–74, January 2006. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
5FC43829285903225B3BD0A52B1AD975.

Gramlich:2007:MNR

[GS07] Gregor Gramlich and Georg Schnitger. Minimizing NFA’s and regular expressions. Journal of Computer and System Sciences, 73(6):908–923, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000006001735.

Gil:2022:MRE

[GS22] José Arturo Gil and Simone Santini. Matching regular expressions on uncertain data. Algorithmica, 84(2):532–564, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/
article/10.1007/s00453-021-00906-8.

Gope:2017:ASS

[GSL17] Dibakar Gope, David J. Schlais, and Mikko H. Lipasti. Architectural support for server-side PHP processing. ACM SIGARCH Computer Architecture News, 45(2):507–520, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Glueck:1990:AMT

[GT90] R. Glueck and V. F. Turchin. Application of metasystem transition to function inversion and transformation. In Watanabe and Nagata [WN90], pages 286–287. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p286-glueck/.

Gilly:1992:UN

[Gt92] Daniel Gilly and the staff of O’Reilly and Associates. UNIX in a Nutshell. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, second edition, 1992. ISBN 1-56592-001-5. LCCN QA76.76.O63 G55 1992.

Galil:1995:CPM

[GU95] Zvi Galil and E. Ukkonen, editors. Combinatorial pattern matching: 6th annual symposium, CPM 95, Espoo, Finland, July 5–7, 1995: proceedings, volume 937 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60044-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

1995. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t0937.htm;   http://www.springerlink.
com/content/978-3-540-60044-2;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
937.

Gawrychowski:2016:OPP

[GU16] Pawel Gawrychowski and Przemyslaw Uznański. Orderpreserving pattern matching with k mismatches. Theoretical Computer Science, 638(??):136–144, July 25, 2016. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397515007690.

Gudeman:1992:DSG

[Gud92] David A. Gudeman. Denotational semantics of a goaldirected language. ACM Transactions on Programming Languages and Systems, 14(1):107–125, January 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/104659.html.

Gueriguian:1987:BRB

[Gue87] John L. Gueriguian. Book review: Bark’ Galianosi: The Greek-Armenian Dictionary to Galen by John A. C. Greppin. Isis, 78(2):300–301, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231579.

Guenther:1990:EEF

[Gue90] G. R. Guenther. Efficient expansion of factored expressions. Information Processing Letters, 35(2):69–72, June 29, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:1997:AST

[Gus97] Dan Gusfield. Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58519-8 (hardcover). xviii + 534 pp. LCCN QA76.9.A43 G87 1997. US$64.95.

Grossi:2000:CSA

[GV00] Roberto Grossi and Jeffrey Scott Vitter. Compressed suffix arrays and suffix trees with applications to text indexing and string matching (extended abstract). In ACM [ACM00], pages 397–406. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/proceedings/stoc/335305/
p397-grossi/p397-grossi.pdf;   http://www.acm.org/
pubs/citations/proceedings/stoc/335305/p397-grossi/

. ACM order number 508000.

Grossi:2005:CSA

[GV05] Roberto Grossi and Jeffrey Scott Vitter. Compressed suffix arrays and suffix trees with applications to text indexing and string matching. SIAM Journal on Computing, 35(2):378– 407, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guting:2015:ST

[GVD15] Ralf Hartmut Güting, Fabio Valdés, and Maria Luisa Damiani. Symbolic trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2):7:1–7:51, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL http://dl.acm.org/citation.cfm?id=2786756.

Ginsburg:1992:PMR

[GW92] Seymour Ginsburg and Xiaoyang Wang. Pattern matching by Rs-operations: towards a unified approach to querying sequenced data. In ACM [ACM92b], pages 293–

300. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26 1992. URL http://

www.acm.org/pubs/articles/proceedings/pods/137097/
p293-ginsburg/p293-ginsburg.pdf;   http://www.acm.
org/pubs/citations/proceedings/pods/137097/p293-ginsburg/
;   http://www.acm.org:80/pubs/citations/proceedings/
pods/137097/p293-ginsburg/. ACM order number 475920.

Guo:2010:LIS

[GWvG10] Yuqing Guo, Haifeng Wang, and Josef van Genabith. A linguistically inspired statistical model for Chinese punctuation generation. ACM Transactions on Asian Language Information Processing, 9(2):6:1–6:??, June 2010. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Gong:2023:EFM

[GWX^{+}23] Lei Gong, Chao Wang, Haojun Xia, Xianglan Chen, Xi Li, and Xuehai Zhou. Enabling fast and memory-efficient acceleration for pattern matching workloads: The Lightweight Automata Processing Engine. IEEE Transactions on Computers, 72(4): 1011–1025, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Goldberg:1994:FPS

[GZ94] T. Goldberg and U. Zwick. Faster parallel string matching via larger deterministic samples. Journal of Algorithms, 16 (2):295–308, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710145.

Guinde:2010:EHS

[GZ10a] Nitesh B. Guinde and Sotirios G. Ziavras. Efficient hardware support for pattern matching in network intrusion detection. Computers & Security, 29(7):756–769, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404810000350.

Gulwani:2010:RBP

[GZ10b] Sumit Gulwani and Florian Zuleger. The reachability-bound problem. ACM SIGPLAN Notices, 45(6):292–304, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1990:CAV

[HA90] Luddy Harrison and Zahira Ammarguellat. A comparison of automatic versus manual parallelization of the Boyer–Moore theorem prover. Technical Report CSRD 960, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1990. 17 pp.

Habibi:2004:JRE

[Hab04] Mehran Habibi. Java Regular Expressions: Taming the java.util.regex Engine. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-107-0. 255 (est.) pp. LCCN ????

Haertel:1989:GG

[Hae89] M. Haertel. GNU e?grep. Usenet archive comp.sources.unix, February 1989.

Hunt:2002:DIL

[HAI02] Ela Hunt, Malcolm P. Atkinson, and Robert W. Irving. Database indexing for large DNA and protein sequence collections. VLDB Journal: Very Large Data Bases, 11(3):256–271, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://link.springer.de/

link/service/journals/00778/bibs/2011003/20110256.
htm;   http://link.springer.de/link/service/journals/
00778/papers/2011003/20110256.pdf. Special issue VLDB

best papers 2001.

Hamilton:1988:LPE

[Ham88] Eric Hamilton. Literate programming—expanding generalized regular expressions. Communications of the Association for Computing Machinery, 31(12):1376–1385, December 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1992:RCT

[Han92] Eric N. Hanson. Rule condition testing and action execution in Ariel. In Stonebraker [Sto92], pages 49–58. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ????

URL http://www.acm.org/pubs/articles/proceedings/
mod/130283/p49-hanson/p49-hanson.pdf;   http://www.
acm.org/pubs/citations/proceedings/mod/130283/p49-
hanson/.

Hancart:1993:SSS

[Han93] Christophe Hancart. On Simon’s string searching algorithm. Information Processing Letters, 47(2):95–99, August 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanegan:2001:CCS

[Han01] Kevin Hanegan. Custom CGI scripting with Perl. John Wiley, New York, NY, USA, 2001. ISBN 0-471-01379-X (e-book), 0-471-39597-8. xxvii + 276 pp. LCCN QA76.73.P22 H36 2001eb.

Hannay:2002:ITC

[Han02] David G. Hannay. Interactive tools for computation theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):68–70, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Han:2013:IPF

[Han13a] Yo-Sub Han. An improved prefix-free regular-expression matching. International Journal of Foundations of Computer Science (IJFCS), 24(5):679–??, August 2013. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2013:SEH

[Han13b] Yo-Sub Han. State elimination heuristics for short regular expressions. Fundamenta Informaticae, 128(4):445–462, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Harrison:1971:IST

[Har71] Malcolm C. Harrison. Implementation of the substring test by hashing. Communications of the Association for Computing Machinery, 14(12):777–779, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [TT82].

Harrington:1979:NSI

[Har79] Steven J. Harrington. A new symbolic integration system in REDUCE. The Computer Journal, 22(2):127–131, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/2/127.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/127.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/128.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/129.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/130.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/131.tif.

Harris:1997:SSP

[Har97] Jason Harris. Semantica: Semantic pattern matching in Mathematica. Mathematica Journal, 7(3):??, ???? 1997. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Harada:2002:PMM

[Har02] Lilian Harada. Pattern matching over multi-attribute data streams. Lecture Notes in Computer Science, 2476:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.de/

link/service/series/0558/bibs/2476/24760187.htm;
http://link.springer.de/link/service/series/0558/papers/
2476/24760187.pdf.

Harrusi:2010:FCP

[HAR10] S. Harrusi, A. Averbuch, and N. Rabin. A fast compact prefix encoding for pattern matching in limited resources devices. In Storer and Marcellin [SM10], page 533. ISBN 0-7695-3994-7. ISSN 1068-0314. LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453510.

Hazez:2001:MTC

[Haz01] Slim Ben Hazez. Modeling textual context in linguistic pattern matching. Lecture Notes in Computer Science, 2004:93– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2004/20040093.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2004/20040093.pdf.

Horvath:2010:EAC

[HBRV10] Ákos Horváth, Gábor Bergmann, István Ráth, and Dániel Varró. Experimental assessment of combining pattern matching strategies with VIATRA2. International Journal on Software Tools for Technology Transfer (STTT), 12(3–4):211–230, July 2010. CODEN ???? ISSN 1433-2779 (print), 1433-

2787 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=1433-2779&volume=12&
issue=3&spage=211.

Held:1987:MNH

[HC87] James P. Held and John V. Carlis. MATCH: a new highlevel relational operator for pattern matching. Communications of the Association for Computing Machinery, 30(1): 62–75, January 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/7889.html.

Hall:1980:ASM

[HD80] Patrick A. V. Hall and Geoff R. Dowling. Approximate string matching. ACM Computing Surveys, 12(4):381–402, December 1980. CODEN CMSVAN. ISSN 0010-4892.

Hearn:1971:CSA

[Hea71] Anthony C. Hearn. The computer solution of algebraic problems by pattern matching. In Proceedings of the Second Colloquium on Advanced Computing Methods in Theoretical Physics, CNRS, Marseilles, pages I–45–I–57. ????, ????, 1971.

Heitsch:2001:GPM

[Hei01] Christine E. Heitsch. Generalized pattern matching and the complexity of unavoidability testing. Lecture Notes in Computer Science, 2089:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2089/
20890219.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2089/20890219.pdf.

Heiberg:2003:TDF

[HEWK03] Einar Heiberg, Tino Ebbers, Lars Wigström, and Matts Karlsson. Three-dimensional flow characterization using vector pattern matching. IEEE Transactions on Visualization and Computer Graphics, 9(3):313–319, July/September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL http://csdl.computer.org/

comp/trans/tg/2003/03/v0313abs.htm;   http://csdl.
computer.org/dl/trans/tg/2003/03/v0313.pdf.

Hull:2013:SPC

[HF13] Richard Hull and Wenfei Fan, editors. SIGMOD/PODS’13: compilation proceedings of the 2013 ACM Symposium on Principles of Database Systems, ACM SIGMOD international conference on management of data, and SIGMOD/PODS 2013 PhD symposium: June 22–27, 2013, New York, New York, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http:/

/dl.acm.org/citation.cfm?id=2463664;   http://www.
sigmod.org/2013/.

Hardavellas:2009:RNN

[HFFA09] Nikos Hardavellas, Michael Ferdman, Babak Falsafi, and Anastasia Ailamaki. Reactive NUCA: near-optimal block placement and replication in distributed caches. ACM SIGARCH Computer Architecture News, 37(3):184–195, June 2009. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hullin:2008:FIR

[HFI^{+}08] Matthias B. Hullin, Martin Fuchs, Ivo Ihrke, Hans-Peter Seidel, and Hendrik P. A. Lensch. Fluorescent immersion range scanning. ACM Transactions on Graphics, 27(3):87:1–87:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hyyro:2005:IBP

[HFN05] Heikki Hyyrö, Kimmo Fredriksson, and Gonzalo Navarro. Increased bit-parallelism for approximate and multiple string matching. ACM Journal of Experimental Algorithmics, 10: 2.6:1–2.6:??, ???? 2005. CODEN ???? ISSN 1084-6654.

He:2005:WWS

[HFS05] Longtao He, Binxing Fang, and Jie Sui. The wide window string matching algorithm. Theoretical Computer Science, 332 (1–3):391–404, February 28, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hossen:2024:TIF

[HGT24] Md Helal Hossen, Daniel Gibney, and Sharma V. Thankachan. Text indexing for faster gapped pattern matching. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-
4893/17/12/537.

Haskin:1983:OCH

[HH83] Roger L. Haskin and Lee A. Hollaar. Operational characteristics of a hardware-based pattern matcher. ACM Transactions on Database Systems, 8(1):15–40, March 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644

(electronic). URL http://www.acm.org/pubs/articles/
journals/tods/1983-8-1/p15-haskin/p15-haskin.pdf;
http://www.acm.org/pubs/citations/journals/tods/1983-
8-1/p15-haskin/.

Herz:1993:ACSa

[HH93a] J. Herz and R. D. Hersch. Analysing character shapes by string matching techniques. In André et al. [AGS93e], pages 261–272. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Herz:1993:ACSb

[HH93b] Jacky Herz and Roger D. Hersch. Analysing character shapes by string matching techniques. Electronic Publishing—Origination, Dissemination, and Design, 6(3): 261–272, September 1993. CODEN EPODEU. ISSN 0894-3982.

Herz:EPODD-6-3-261

[HH93c] Jacky Herz and Roger D. Hersch. Analysing character shapes by string matching techniques. Electronic Publishing—Origination, Dissemination, and Design, 6(3): 261–272, September 1993. CODEN EPODEU. ISSN 0894-3982.

Hu:2016:EWS

[HH16] Changhui Hu and Lidong Han. Efficient wildcard search over encrypted data. International Journal of Information Security, 15(5):539–547, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL http://link.

springer.com/article/10.1007/s10207-015-0302-0.

Huynh:2006:ASM

[HHLS06] Trinh N. D. Huynh, Wing-Kai Hon, Tak-Wah Lam, and Wing-Kin Sung. Approximate string matching using compressed suffix arrays. Theoretical Computer Science, 352(1–3):240– 249, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haber:2013:ESE

[HHM^{+}13] Stuart Haber, William Horne, Pratyusa Manadhata, Miranda Mowbray, and Prasad Rao. Efficient submatch extraction for practical regular expressions. In Dediu et al. [DMVT13], pages 323–334. ISBN 3-642-37063-2 (print), 3-642-37064-0 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5-76.95.

Harrison:1999:LRP

[HHW^{+}99] Brian Harrison, Philip Hendrickson, Murel Warren, Jr., Lee Kamentsky, Ron Gutman, Brenton Hoff, Martin Handwerker, Tom Culliton, and Aspi Havewala. Letters: Real programmer’s hate Cobol; 1984; Hilbert curves; grepping and globbing; testing Java classes; the version control process. Dr. Dobb’s Journal of Software Tools, 24(9):10, 12, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Hamed:2022:SIP

[HIEH22] Belal A. Hamed, Osman Ali Sadek Ibrahim, and Tarek Abd El-Hafeez. A survey on improving pattern matching algorithms for biological sequences. Concurrency and Computation: Practice and Experience, 34(26):e7292:1–e7292:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Highnam:1986:OAF

[Hig86] P. T. Highnam. Optimal algorithms for finding the symmetries of a planar point set. Information Processing Letters, 22(5): 219–222, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Highland:1995:BRP

[Hig95] Harold Joseph Highland. Book review: A Pattern Matching Model for Misuse Intrusion Detection: Sandeep Kumar and Eugene Spafford, Purdue University, West Lafayette, IN. Computers & Security, 14(1):28, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489596997H.

Hirshfeld:1996:ULE

[Hir96] Yoram Hirshfeld. Undecidability of language equivalence for generalized regular expressions. Fundamenta Informaticae, 26 (1):95–102, January 1996. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hasan:2017:PSA

[HIRS17] Md. Mahbubul Hasan, A. S. M. Sohidull Islam, M. Sohel Rahman, and Ayon Sen. Palindromic subsequence automata and longest common palindromic subsequence. Mathematics in Computer Science, 11(2):219–232, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Henrich:1999:OQC

[HJ99] A. Henrich and S. Jamin. On the optimization of queries containing regular path expressions. Lecture Notes in Computer Science, 1649:58–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudak:1992:RPL

[HJW^{+}92] Paul Hudak, Simon Peyton Jones, Philip Wadler, Brian Boutel, Jon Fairbairn, Joseph Fasel, María M. Guzmán, Kevin Hammond, John Hughes, Thomas Johnsson, Dick Kieburtz, Rishiyur Nikhil, Will Partain, and John Peterson. Report on the programming language Haskell: a non-strict, purely functional language (version 1.2). ACM SIGPLAN Notices, 27 (5):Ri–Rx, R1–R163, May 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Housden:1977:CSP

[HK77] R. J. W. Housden and N. Kotarski. Character string pattern matching in Algol 68. ACM SIGPLAN Notices, 12(6):144– 152, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holzer:2011:CRL

[HK11] Markus Holzer and Martin Kutrib. The complexity of regular(-like) expressions. International Journal of Foundations of Computer Science (IJFCS), 22(7):1533–1548, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Heil:2014:APH

[HKL^{+}14] Timothy Heil, Anil Krishna, Nicholas Lindberg, Farnaz Toussi, and Steven Vanderwiel. Architecture and performance of the hardware accelerators in IBM’s PowerEN processor. ACM Transactions on Parallel Computing (TOPC), 1(1):5:1–5:??, September 2014. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Hanada:2014:ACL

[HKN14] Hiroyuki Hanada, Mineichi Kudo, and Atsuyoshi Nakamura. Average-case linear-time similar substring searching by the qgram distance. Theoretical Computer Science, 530(??):23–41, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397514001285.

Heering:1992:IGL

[HKR92] J. Heering, P. Klint, and J. Rekers. Incremental generation of lexical scanners. ACM Transactions on Programming Languages and Systems, 14(4):490–520, October 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/133240.html.

Havas:1994:NPS

[HL94] G. Havas and Jin Xian Lian. A new problem in string searching. Lecture Notes in Computer Science, 834:660–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemer:1997:RVD

[HL97] D. Hemer and P. A. Lindsay. Reuse of verified design templates through extended pattern matching. Lecture Notes in Computer Science, 1313:495–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hazay:2010:EPS

[HL10]  Carmit Hazay and Yehuda Lindell.  Efficient proto-

cols for set intersection and pattern matching with security against malicious and covert adversaries. Journal of Cryptology, 23(3):422–456, July 2010. CODEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0933-2790&volume=23&issue=3&spage=422.

Hundt:2009:CGA

[HLN09] Christian Hundt, Maciej Liśkiewicz, and Ragnar Nevries. A combinatorial geometrical approach to two-dimensional robust pattern matching with scaling and rotation. Theoretical Computer Science, 410(51):5317–5333, November 28, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hazay:2007:APM

[HLS07] Carmit Hazay, Moshe Lewenstein, and Dina Sokol. Approximate parameterized matching. ACM Transactions on Algorithms, 3(3):29:1–29:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hon:2011:COI

[HLS^{+}11] Wing-Kai Hon, Tak-Wah Lam, Rahul Shah, Siu-Lung Tam, and Jeffrey Scott Vitter. Cache-oblivious index for approximate string matching. Theoretical Computer Science, 412(29): 3579–3588, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoffman:1987:MC

[HM87] Alan J. Hoffman and Willard L. Miranker, editors. Mathematics and Computing, volume 31(2) of IBM Journal of Research and Development. IBM Corporation, Armonk, NY, USA / Poughkeepsie, NY, USA / San Jose, CA, USA / Yorktown Heights, NY, USA, March 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Hirschberg:1996:CPM

[HM96] Dan Hirschberg and Gene Myers, editors. Combinatorial pattern matching: 7th Annual Symposium, CPM 96, Laguna Beach, California, June 10–12, 1996: proceedings, volume 1075 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61258-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

1996. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t1075.htm;   http://www.springerlink.
com/content/978-3-540-61258-2;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=

Hagenah:1998:CFN

[HM98] Christian Hagenah and Anca Muscholl. Computing ϵfree NFA from regular expressions in O(n log^{2}(n)) time. Lecture Notes in Computer Science, 1450:277–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1450/14500277.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1450/14500277.pdf.

Holub:2000:ASM

[HM00] Jan Holub and Bořivoj Melichar. Approximate string matching using factor automata. Theoretical Computer Science, 249(2):305–311, October 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-ng/10/41/16/184/23/
24/abstract.html;   http://www.elsevier.nl/gej-ng/
10/41/16/184/23/24/article.pdf.

Hopcroft:2001:IAT

[HMU01] John E. (John Edward) Hopcroft, Rajeev Motwani, and Jeffrey D. (Jeffrey David) Ullman. Introduction to Automata Theory, Languages, and Computation. Addison-Wesley, Reading, MA, USA, second edition, 2001. ISBN 0-201-44124-1 (paperback), 0-321-21029-8 (International edition paperback). xiv + 521 pp. LCCN QA267 .H56 2001.

Hopcroft:2007:IAT

[HMU07] John E. (John Edward) Hopcroft, Rajeev Motwani, and Jeffrey D. (Jeffrey David) Ullman. Introduction to Automata Theory, Languages, and Computation. Pearson Education, Boston, MA, USA, third edition, 2007. ISBN 0-321-45536-3 (hardcover), 0-321-45537-1 (student access kit), 0-321-47617-4 (paperback), 0-321-51448-3 (complete edition). xvii + 535 pp. LCCN QA76.9 .A73H67.

Hong:2023:MMS

[HMW^{+}23] T. Hong, X. Ma, X. Wang, R. Che, C. Hu, T. Feng, and W. Zhang. MAPMaN: Multi-stage U-shaped adaptive pattern matching network for semantic segmentation of remote sensing images. Computer Graphics Forum, 42(7):e14978:1– e14978:??, October 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hendren:1990:PPR

[HN90] Laurie J. Hendren and Alexandru Nicolau. Parallelizing programs with recursive data structures. IEEE Transactions on Parallel and Distributed Systems, 1(1):35–47, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Howland:2000:REC

[HN00] Eric Howland and David Niergarth. Regular expressions for checking dates. Markup languages: theory & practice, 2(2): 126–132, Spring 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http://mitpress.mit.

edu/catalog/item/default.asp?sid=3108FA5E-91BB-480F-
9930-C1C27725EAB8&ttype=6&tid=6855.

Hyyro:2002:FBP

[HN02] Heikki Hyyrö and Gonzalo Navarro. Faster bit-parallel approximate string matching. Lecture Notes in Computer Science, 2373:203–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2373/
23730203.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2373/23730203.pdf.

Hyyro:2005:BPW

[HN05] Heikki Hyyrö and Gonzalo Navarro. Bit-parallel witnesses and their applications to approximate string matching. Algorithmica, 41(3):203–231, January 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=3&spage=203.

Henglein:2011:REC

[HN11] Fritz Henglein and Lasse Nielsen. Regular expression containment: coinductive axiomatization and computational interpretation. ACM SIGPLAN Notices, 46(1):385–398, January 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Homer:2013:POG

[HNB^{+}13] Michael Homer, James Noble, Kim B. Bruce, Andrew P. Black, and David J. Pearce. Patterns as objects in Grace.

ACM SIGPLAN Notices, 48(2):17–28, February 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoffmann:1982:PMT

[HO82] Christoph M. Hoffmann and Michael J. O’Donnell. Pattern matching in trees. Journal of the Association for Computing Machinery, 29(1):68–95, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoalst:1977:SMC

[Hoa77] B. C. Hoalst. String manipulation and conversion for switching control center. Technical Memorandum 1211 (TM 77-5222-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 20, 1977. ?? pp.

Hocking:2019:CNO

[Hoc19] Toby Dylan Hocking. Comparing namedCapture with other R packages for regular expressions. The R Journal, 11(2): 328–346, December 2019. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2019/RJ-2019-050.

Ho:2018:CNA

[HOK18a] ThienLuan Ho, Seung-Rohk Oh, and HyunJin Kim. Correction to: New algorithms for fixed-length approximate string matching and approximate circular string matching under the Hamming distance. The Journal of Supercomputing, 74(5):1835, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s11227-018-

2324-7.pdf. See [HOK18b].

Ho:2018:NAF

[HOK18b] ThienLuan Ho, Seung-Rohk Oh, and HyunJin Kim. New algorithms for fixed-length approximate string matching and approximate circular string matching under the Hamming distance. The Journal of Supercomputing, 74(5):1815–1834, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [HOK18a].

Holub:1984:GCU

[Hol84] Allen I. Holub. Grep C — a Unix-like, generalized, regular expression parser in C. Dr. Dobb’s Journal of Software Tools, 9(10):50–??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Horspool:1980:PFS

[Hor80] R. Nigel Horspool. Practical fast searching in strings. Software — Practice and Experience, 10(6):501–506, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffman:1985:IIA

[HOS85a] C. M. Hoffman, M. J. O’Donnell, and R. I. Strandh. Implementation of an interpreter for abstract equations. Software — Practice and Experience, 15(12):1204–1185, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffmann:1985:IIA

[HOS85b] Christoph M. Hoffmann, Michael J. O’Donnell, and Robert I. Strandh. Implementation of an interpreter for abstract equations. Software — Practice and Experience, 15(12):1185–1204, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hosaya:2006:REF

[Hos06] Haruo Hosaya. Regular expression filters for XML. Journal of Functional Programming, 16(6):711–750, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
F27B4BABC2160A54AF14D9B76D6CC999.

Hovland:2012:IPR

[Hov12] Dag Hovland. The inclusion problem for regular expressions. Journal of Computer and System Sciences, 78(6):1795–1813, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022000011001486.

Howard:1996:LLC

[How96] P. G. Howard. Lossless and lossy compression of text images by soft pattern matching. In Storer and Cohn [SC96], pages 210–219. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488326. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Howard:1997:TIC

[How97] P. G. Howard. Text image compression using soft pattern matching. The Computer Journal, 40(2–3):146–156, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/40/2_and_3/146.full.pdf+html;   http:/
/www.oup.co.uk/computer_journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractHoward;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_03.
body.html#AbstractHoward.

Hosoya:2001:REP

[HP01] Haruo Hosoya and Benjamin Pierce. Regular expression pattern matching for XML. ACM SIGPLAN Notices, 36(3):67– 80, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

acm.org/pubs/articles/proceedings/plan/360204/p67-
hosoya/p67-hosoya.pdf;   http://www.acm.org/pubs/citations/
proceedings/plan/360204/p67-hosoya/.

Hosoya:2003:REP

[HP03] Haruo Hosoya and Benjamin C. Pierce. Regular expression pattern matching for XML. Journal of Functional Programming, 13(6):961–1004, November 2003. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
C845B41E6B150FBD7731EE396FBCB911.

Huang:1994:PRA

[HPM94] X. Huang, P. A. Pevzner, and W. Miller. Parametric recomputing in alignment graphs. Lecture Notes in Computer Science, 807:87–101, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hung:2000:IVI

[HR00] Ted Hung and Susan H. Rodger. Increasing visualization and interaction in the automata theory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):6–10, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hernadez:2001:DRL

[HR01] Manuel Hernádez and David A. Rosenblueth. Development reuse and the logic program derivation of two string-matching algorithms. In Rocco De Nicola and Harald Søndergaard, editors, Proceedings of the 3rd ACM SIGPLAN international conference on Principles and practice of declarative programming, Florence, Italy, September 5–7, 2001, PPDP01, pages 38–48. ACM Press, New York, NY 10036, USA, September 2001. ISBN 1-58113-388-X (paperback). LCCN QA76.615 .P57 2001.

Anonymous:2002:MSS

[HR02] Moishe Halibard and Moshe Rubin. A multiple substring search algorithm. C/C++ Users Journal, 20(6):6–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Hernandez:2003:DPD

[HR03] Manuel Hernández and David A. Rosenblueth. Disjunctive partial deduction of a right-to-left string-matching algorithm. Information Processing Letters, 87(5):235–241, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hammoud:2015:DDR

[HRN^{+}15] Mohammad Hammoud, Dania Abed Rabbou, Reza Nouri, Seyed-Mehdi-Reza Beheshti, and Sherif Sakr. DREAM: distributed RDF engine with adaptive query planner and minimal communication. Proceedings of the VLDB Endowment, 8 (6):654–665, February 2015. CODEN ???? ISSN 2150-8097.

Hume:1990:FSS

[HS90] Andrew Hume and Daniel Sunday. Fast string searching. Computing Science Technical Report 156, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990. ?? pp.

Hume:1991:FSS

[HS91] Andrew Hume and Daniel Sunday. Fast string searching. Software — Practice and Experience, 21(11):1221–1248, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [BM77, KMP77, Sun90a].

Huang:2008:ESS

[HS08] Shan Shan Huang and Yannis Smaragdakis. Expressive and safe static reflection with MorphJ. ACM SIGPLAN Notices, 43(6):79–89, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hashiguchi:2004:ERB

[HSJ04] Kosaburo Hashiguchi, Naoto Sakakibara, and Shuji Jimbo. Equivalence of regular binoid expressions and regular expressions denoting binoid languages over free binoids. Theoretical Computer Science, 312(2–3):251–266, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoefler:2010:SCP

[HSL10] Torsten Hoefler, Christian Siebert, and Andrew Lumsdaine. Scalable communication protocols for dynamic sparse data exchange. ACM SIGPLAN Notices, 45(5):159–168, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hori:2001:FPM

[HSTS01] Hideaki Hori, Shinichi Shimozono, Masayuki Takeda, and Ayumi Shinohara. Fragmentary pattern matching: Complexity, algorithms and applications for analyzing classic literary works. Lecture Notes in Computer Science, 2223:719–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2223/22230719.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2223/22230719.pdf.

Hromkovic:1997:TRE

[HSW97] J. Hromkovic, S. Seibert, and T. Wilke. Translating regular expressions into small curly epsilon-free nondeterministic finite automata. Lecture Notes in Computer Science, 1200:55– ??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hromkovic:2001:TRE

[HSW01] Juraj Hromkovic, Sebastian Seibert, and Thomas Wilke. Translating regular expressions into small ϵ-free nondeterministic finite automata. Journal of Computer and System Sciences, 62(4):565–588, June 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000001917489.

Hazay:2014:CSP

[HT14] Carmit Hazay and Tomas Toft. Computationally secure pattern matching in the presence of malicious adversaries. Journal of Cryptology, 27(2):358–395, April 2014. CO-DEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s00145-013-9147-8.

Hirvola:2017:BPA

[HT17] Tommi Hirvola and Jorma Tarhio. Bit-parallel approximate matching of circular strings with k mismatches. ACM Journal of Experimental Algorithmics, 22(??):1.5:1–1.5:??, 2017. CODEN ???? ISSN 1084-6654.

Higuchi:2021:FLP

[HTK^{+}21] Shunsuke Higuchi, Junji Takemasa, Yuki Koizumi, Atsushi Tagami, and Toru Hasegawa. Feasibility of Longest Prefix Matching using Learned Index Structures. ACM SIGMET-RICS Performance Evaluation Review, 48(4):45–48, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3466826.
3466842.

Hon:2017:PAE

[HTX17] Wing-Kai Hon, Sharma V. Thankachan, and Bojian Xu. In-place algorithms for exact and approximate shortest unique substring problems. Theoretical Computer Science, 690(??):12–25, August 22, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397517304760.

Hopcroft:1979:IAT

[HU79] John E. Hopcroft and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation. Addison-Wesley Series in Computer Science. Addison-Wesley, Reading, MA, USA, 1979. ISBN 0-201-02988-X. 418 pp. LCCN QA267 .H56.

Hopcroft:1992:EAF

[HU92] John E. Hopcroft and Jeffrey D. Ullman. Einführung in die Automatentheorie, formale Sprachen und Komplexitätstheorie. (German) [Introduction to Automata Theory, Formal Languages and Complexity Theory]. Internationale Computer-Bibliothek. Addison-Wesley, Reading, MA, USA,

1. nachdr. edition, 1992. ISBN 0-201-02988-X, 3-89319-181-X. ix + 461 pp. LCCN ????

Huang:1994:CDM

[Hua94] X. Huang. A context dependent method for comparing sequences. Lecture Notes in Computer Science, 807:54–63, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1998:CKM

[Hua98] Andy Huang. Computation of the Knuth–Morris–Pratt skip tables. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 29(3):59–61, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hucke:2021:ARR

[Huc21] Danny Hucke. Approximation ratios of RePair, LongestMatch and Greedy on unary strings. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/2/65.

Hudak:1989:CEA

[Hud89] Paul Hudak. Conception, evolution, and application of functional programming languages. ACM Computing Surveys, 21 (3):359–411, September 1989. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0360-0300/72554.html.

Hui:1992:CSS

[Hui92] Lucas C. K. Hui. Color set size problem with applications to string matching. Lecture Notes in Computer Science, 644: 227–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hume:1988:TTG

[Hum88a] Andrew Hume. A tale of two greps. Software — Practice and Experience, 18(11):1063–1072, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:1988:GW

[Hum88b] Andrew Hume. Grep wars. In ????, editor, 1988 Spring (London) EUUG Conference Proceedings, pages 237–245. ????, ????, 1988.

Hume:1997:BRMa

[Hum97a] Andrew Hume. Book review: Mastering Regular Expressions. ;login:, 22(5):67–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Hume:1997:BRMb

[Hum97b] Andrew Hume. Book review: Mastering Regular Expressions. ;login:, 22(5):67–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Hume:1999:DSR

[Hum99] Andrew Hume. The dark side of regular expressions. ;login:, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/publications/login/1999-
4/reg_exp.html.

Hunt:1979:OCR

[Hun79] H. B. Hunt. Observations on the complexity of regular expression problems. Journal of Computer and System Sciences, 19(3):222–236, December 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0022000079900023.

Hendrian:2019:PPM

[HUN^{+}19] Diptarama Hendrian, Yohei Ueki, Kazuyuki Narisawa, Ryo Yoshinaka, and Ayumi Shinohara. Permuted pattern matching algorithms on multi-track strings. Algorithms (Basel), 12 (4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/4/73.

Hurson:1984:VDP

[Hur84] A. R. Hurson. A VLSI design for the parallel finite state automaton and its performance evaluation as a hardware scanner. International Journal of Computer and Information Sciences, 13(6):491–508, December 1984. CODEN IJCIAH. ISSN 0091-7036.

Horspool:1993:SAP

[HV93] R. Nigel Horspool and Jan Vitek. Static analysis of PostScript code. Computer Languages, 19(2):65–78, April 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hosoya:2000:RET

[HVP00] Haruo Hosoya, Jérôme Vouillon, and Benjamin C. Pierce. Regular expression types for XML. ACM SIGPLAN Notices, 35(9):11–22, September 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/pubs/articles/proceedings/
fp/351240/p11-hosoya/p11-hosoya.pdf;   http://www.
acm.org/pubs/citations/proceedings/fp/351240/p11-hosoya/

.

Hosoya:2005:RET

[HVP05] Haruo Hosoya, Jérôme Vouillon, and Benjamin C. Pierce. Regular expression types for XML. ACM Transactions on Programming Languages and Systems, 27(1):46–90, January 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Han:2007:OSR

[HW07] Yo-Sub Han and Derick Wood. Obtaining shorter regular expressions from finite-state automata. Theoretical Computer Science, 370(1–3):110–120, February 12, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hooimeijer:2009:DPS

[HW09] Pieter Hooimeijer and Westley Weimer. A decision procedure for subset constraints over regular languages. ACM SIGPLAN Notices, 44(6):188–198, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hundt:2012:ETD

[HW12] Christian Hundt and Florian Wendland. Efficient twodimensional pattern matching with scaling and rotation and higher-order interpolation. Lecture Notes in Computer Science, 7354:124–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_10/.

Hwang:1985:PSC

[Hwa85] Kai Hwang, editor. Proceedings: 7th Symposium on Computer Arithmetic, June 4–6, 1985, University of Illinois, Urbana, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985. IEEE catalog number 85CH2146-9. IEEE Computer Society order number 632.

Henry:1990:UWI

[HWF90] Robert R. Henry, Kenneth M. Whaley, and Bruce Forstall. The University of Washington illustrating compiler. ACM SIGPLAN Notices, 25(6):223–233, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org:80/pubs/citations/proceedings/pldi/93542/p223-
henry/.

Hashiguchi:2003:RBE

[HWJ03] Kosaburo Hashiguchi, Yoshito Wada, and Shuji Jimbo. Regular binoid expressions and regular binoid languages. Theoretical Computer Science, 304(1–3):291–313, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2006:IFR

[HWW06] Yo-Sub Han, Yajun Wang, and Derick Wood. Infix-free regular expressions and languages. International Journal of Foundations of Computer Science (IJFCS), 17(2):379–??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2007:PFR

[HWW07] Yo-Sub Han, Yajun Wang, and Derick Wood. Prefix-free regular languages and pattern matching. Theoretical Computer Science, 389(1–2):307–317, December 10, 2007. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1990:ERE

[HY90] K. Hashiguchi and H. Yoo. Extended regular expressions of star degree at most two. Theoretical Computer Science, 76 (2-3):273–284, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1992:TRS

[HY92] K. Hashiguchi and K. Yamada. Two recognizable stringmatching problems over free partially commutative monoids. Theoretical Computer Science, 92(1):77–86, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hyyro:2008:IBP

[Hyy08] Heikki Hyyrö. Improving the bit-parallel NFA of Baeza-Yates and Navarro for approximate string matching. Information Processing Letters, 108(5):313–319, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hao:2013:TPP

[HZ13] Tianyong Hao and Chunshen Zhu. Toward a professional platform for Chinese character conversion. ACM Transactions on Asian Language Information Processing, 12(1):1:1–1:??, March 2013. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Iyengar:1980:SSA

[IA80] S. Sitharama Iyengar and Vincent Alia. A string searching algorithm. Applied Mathematics and Computation, 6 (2):123–131, March 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0096300380900375.

Aoe:1994:CAS

[iA94] Jun ichi Aoe, editor. Computer Algorithms: String Pattern Matching Strategies. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5461-9 (microfiche), 0-8186-5462-7 (hardcover), 0-8186-5460-0 (paperback). LCCN QA76.9.A43 C67 1994. US$56.00. IEEE Computer Society Press order number: 5462-05. IEEE catalog number: 94EH0389-7.

Ibarra:1997:FPM

[Iba97] Louis Ibarra. Finding pattern matchings for permutations. Information Processing Letters, 61(6):293–295, April 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

IEEE:1988:PSN

[IEE88] IEEE, editor. Proceedings, Supercomputing ’88: November 14–18, 1988, Orlando, Florida, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

IEEE:1989:ASF

[IEE89] IEEE, editor. 30th annual Symposium on Foundations of Computer Science, October 30–November 1, 1989, Research Triangle Park, North Carolina. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

IEEE:1990:PAS

[IEE90] IEEE, editor. Proceedings: 31st Annual Symposium on Foundations of Computer Science: October 22–24, 1990, St. Louis, Missouri. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

IEEE:1992:ASF

[IEE92] IEEE, editor. 33rd Annual Symposium on Foundations of Computer Science: October 24–27, 1992, Pittsburgh, Pennsylvania: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

IEEE:1993:ASF

[IEE93] IEEE, editor. 34th Annual Symposium on Foundations of Computer Science: November 3–5, 1993, Palo Alto, California: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

IEEE:1994:FAP

[IEE94a] IEEE, editor. First Asia-Pacific Software Engineering Conference: proceedings, December 7–9, 1994, Tokyo, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6960-8 (paper), 0-8186-6961-6 (microfiche). LCCN QA76.758.A77 1994.

IEEE:1994:PTC

[IEE94b] IEEE, editor. Proceedings / the Tenth Conference on Artificial Intelligence for Applications, March 1–4, 1994, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5550-X. LCCN Q 334 C66 1994. IEEE Catalog No. 94CH3421-5.

IEEE:1995:ASF

[IEE95a] IEEE, editor. 36th Annual Symposium on Foundations of Computer Science: October 23–25, 1995, Milwaukee, Wisconsin. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

IEEE:1995:PNA

[IEE95b] IEEE, editor. Proceedings: the nineteenth annual International Computer Software and Applications Conference (COMPSAC ’95): August 9–11, 1995, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7119-X. LCCN QA76.6 .I5 1995. IEEE Catalog No. 95CB35838.

IEEE:1997:ASF

[IEE97] IEEE, editor. 38th Annual Symposium on Foundations of Computer Science: October 20–22, 1997, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ASFPDV. ISBN 0-8186-8197-7, 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

IEEE:1998:ASF

[IEE98] IEEE, editor. 39th Annual Symposium on Foundations of Computer Science: proceedings: November 8–11, 1998, Palo Alto, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

IEEE:2001:ISRa

[IEE01a] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Base Definitions, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3047-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xliv + 448 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISSa

[IEE01b] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Base Definitions, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3047-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xliv + 448 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRd

[IEE01c] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Rationale (Informative). IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3048-6 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxiv + 310 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISSd

[IEE01d] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Rationale (Informative). IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3048-6 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxiv + 310 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRc

[IEE01e] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Shell and Utilities, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3050-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxii + 1090 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISSc

[IEE01f] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Shell and Utilities, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3050-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxii + 1090 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992 Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRb

[IEE01g] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) System Interfaces, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3094-4 (invalid checksum??) (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxx + 1690 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992, Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISSb

[IEE01h] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) System Interfaces, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3094-X (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxx + 1690 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992, Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2009:PAI

[IEE09] IEEE, editor. Proceedings of the 50th Annual IEEE Symposium on Foundations of Computer Science: October 25–27, 2009, Atlanta, Georgia, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Ierusalimschy:2009:TPM

[Ier09] Roberto Ierusalimschy. A text pattern-matching tool based on Parsing Expression Grammars. Software — Practice and Experience, 39(3):221–258, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Irving:1994:MCS

[IF94] R. W. Irving and C. B. Fraser. Maximal common subsequences and minimal common supersequences. Lecture Notes in Computer Science, 807:173–183, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibrahim:2023:NFT

[IHEH23] Osman Ali Sadek Ibrahim, Belal A. Hamed, and Tarek Abd El-Hafeez. A new fast technique for pattern matching in biological sequences. The Journal of Supercomputing, 79(1): 367–388, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-022-04673-3.

Israeli:1986:FSR

[II86] Amos Israeli and A. Itai. A fast and simple randomized parallel algorithm for maximal matching. Information Processing Letters, 22(2):77–80, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Islam:2008:STS

[II08] Aminul Islam and Diana Inkpen. Semantic text similarity using corpus-based word similarity and string similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 2 (2):10:1–10:??, July 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Igarashi:2009:DPT

[II09] Yuki Igarashi and Takeo Igarashi. Designing plush toys with a computer. Communications of the Association for Computing Machinery, 52(12):81–88, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Islam:2008:ACB

[IIK08] Aminul Islam, Diana Inkpen, and Iluju Kiringa. Applications of corpus-based semantic similarity and word segmentation to database schema matching. VLDB Journal: Very Large Data Bases, 17(5):1293–1320, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

I:2013:PPM

[IIT13a] Tomohiro I, Shunsuke Inenaga, and Masayuki Takeda. Palindrome pattern matching. Theoretical Computer Science, 483(??):162–170, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397512001041.

Tomohiro:2013:PPM

[IIT13b] Tomohiro I, Shunsuke Inenaga, and Masayuki Takeda. Palindrome pattern matching. Theoretical Computer Science, 483(??):162–170, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397512001041.

Ito:1983:HFO

[IK83] Tetsuro Ito and Makoto Kizawa. Hierarchical file organization and its application to similar-string matching. ACM Transactions on Database Systems, 8(3):410–433, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.acm.org/pubs/

articles/journals/tods/1983-8-3/p410-ito/p410-ito.
pdf;   http://www.acm.org/pubs/citations/journals/
tods/1983-8-3/p410-ito/.

Ileri:2015:SYT

[IKX15] Atalay Mert Ileri, M. Oguzhan Külekci, and Bojian Xu. A simple yet time-optimal and linear-space algorithm for shortest unique substring queries. Theoretical Computer Science, 562(??):621–633, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397514008470.

Isradisaikul:2013:REP

[IM13] Chinawat Isradisaikul and Andrew C. Myers. Reconciling exhaustive pattern matching with objects. ACM SIGPLAN Notices, 48(6):343–354, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iten:2022:EPP

[IMM^{+}22] Raban Iten, Romain Moyard, Tony Metger, David Sutter, and Stefan Woerner. Exact and practical pattern matching for quantum circuit optimization. ACM Transactions on Quantum Computing (TQC), 3(1):4:1–4:41, March 2022. CODEN ???? ISSN 2643-6809 (print), 2643-6817 (electronic). URL

https://dl.acm.org/doi/10.1145/3498325.

Iliopoulos:2001:MSA

[IMP01] Costas S. Iliopoulos, Laurent Mouchard, and Yoan J. Pinzon. The max-shift algorithm for approximate string matching. Lecture Notes in Computer Science, 2141:13–??, 2001. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-

6119 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2141/21410013.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2141/21410013.pdf.

Iliopoulos:2008:NAP

[IMR08] Costas S. Iliopoulos, Laurent Mouchard, and M. Sohel Rahman. A new approach to pattern matching in degenerate DNA/RNA sequences and distributed pattern matching. Mathematics in Computer Science, 1(4):557–569, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1661-8270&volume=1&issue=4&spage=557.

Iliopoulos:1997:CSF

[IMS97] Costas S. Iliopoulos, Dennis Moore, and W. F. Smyth. A characterization of the squares in a Fibonacci string. Theoretical Computer Science, 172(1–2):281–291, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2344.

Indyk:1997:DSC

[Ind97] P. Indyk. Deterministic superimposed coding with applications to pattern matching. In IEEE [IEE97], pages 127–136.

CODEN ASFPDV. ISBN 0-8186-8197-7, 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Indyk:1998:FAS

[Ind98] P. Indyk. Faster algorithms for string matching problems: matching the convolution bound. In IEEE [IEE98], pages 166– 173. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Iliopoulos:1996:WTO

[IP96] Costas S. Iliopoulos and Kunsoo Park. A work-time optimal algorithm for computing all string covers. Theoretical Computer Science, 164(1–2):299–310, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2255.

Israeli:1986:IPA

[IS86] Amos Israeli and Y. Shiloach. An improved parallel algorithm for maximal matching. Information Processing Letters, 22(2): 57–60, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Isenman:1990:PAI

[IS90] M. E. Isenman and D. E. Shasha. Performance and architectural issues for string matching. IEEE Transactions on Computers, 39(2):238–250, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=45209.

Idury:1994:MMP

[IS94] R. M. Idury and A. A. Schaeffer. Multiple matching of parameterized patterns. Lecture Notes in Computer Science, 807: 226–239, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Idury:1996:MMP

[IS96] Ramana M. Idury and Alejandro A. Schäffer. Multiple matching of parameterized patterns. Theoretical Computer Science, 154(2):203–224, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1904.

Itano:1988:IPM

[ISHY88] Kozo Itano, Yutaka Sato, Hidemi Hirai, and Tomoyoshi Yamagata. An incremental pattern matching algorithm for the pipelined lexical scanner. Information Processing Letters, 27 (5):253–258, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ito:1994:PTA

[ISNH94] M. Ito, K. Shimizu, M. Nakanishi, and A. Hashimoto. Polynomial-time algorithms for computing characteristic strings. Lecture Notes in Computer Science, 807:274–288, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inenaga:2005:FCP

[IST05] Shunsuke Inenaga, Ayumi Shinohara, and Masayuki Takeda. A fully compressed pattern matching algorithm for simple collage systems. International Journal of Foundations of Computer Science (IJFCS), 16(6):1155–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ilie:2002:CNOa

[IY02a] Lucian Ilie and Sheng Yu. Constructing NFAs by optimal use of positions in regular expressions. Lecture Notes in Computer Science, 2373:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2373/
23730279.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2373/23730279.pdf.

Ilie:2002:CNOb

[IY02b] Lucian Ilie and Sheng Yu. Constructing NFAs by optimal use of positions in regular expressions. Lecture Notes in Computer Science, 2373:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2373/
23730279.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2373/23730279.pdf.

Jiang:2017:CSM

[JA17] Peng Jiang and Gagan Agrawal. Combining SIMD and many/ multi-core parallelism for finite state machines with enumerative speculation. ACM SIGPLAN Notices, 52(8):179–191, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jantzen:1985:ERE

[Jan85] M. Jantzen. Extending regular expressions with iterated shuffle. Theoretical Computer Science, 38(2-3):223–247, June 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janson:2023:AND

[Jan23] Svante Janson. Asymptotic normality for m -dependent and constrained U -statistics, with applications to pattern matching in random strings and permutations. Advances in Applied Probability, 55(3):841–894, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://www.cambridge.

org/core/journals/advances-in-applied-probability/
article/asymptotic-normality-for-boldsymbolmdependent-
and-constrained-boldsymbolustatistics-with-applications-
to-pattern-matching-in-random-strings-and-permutations/
1CF7F29EE91684C5AD9391696C90B973.

Jayaraman:1992:SAL

[Jay92] Bharat Jayaraman. Sublist assertions for listless and lazy evaluation. Computer Languages, 17(2):133–146, April 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jouvelot:1989:RPM

[JD89] P. Jouvelot and B. Dehbonei. Recursive Pattern Matching on concrete data types. ACM SIGPLAN Notices, 24(11):84–93, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:2013:PSE

[JDXD13] Song Jiang, Xiaoning Ding, Yuehai Xu, and Kei Davis. A prefetching scheme exploiting both data layout and access history on disk. ACM Transactions on Storage, 9(3):10:1–10:??, August 2013. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Jpolhkedrzejowicz:1987:NSC

[Jed87] Joanna Jedrzejowicz. Nesting of shuffle closure is important. Information Processing Letters, 25(6):363–367, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jez:2015:FFC

[Jez15] Artur Jez. Faster fully compressed pattern matching by recompression. ACM Transactions on Algorithms, 11(3):20:1– 20:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jang:2022:MMU

[JGMP22] Junyoung Jang, Samuel Gélineau, Stefan Monnier, and Brigitte Pientka. Moebius: metaprogramming using contextual types: the stage where system f can pattern match on itself. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):39:1–39:27, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https://dl.acm.

org/doi/10.1145/3498700.

Jiang:2012:SPM

[JGZL12] Kunpeng Jiang, Huifang Guo, Shengping Zhu, and Julong Lan. Static patterns matching for high speed networks. Lecture Notes in Computer Science, 7473:15–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-34062-8_2/.

Jargalsaikhan:2024:SPA

[JHU^{+}24] Davaajav Jargalsaikhan, Diptarama Hendrian, Yohei Ueki, Ryo Yoshinaka, and Ayumi Shinohara. Serial and parallel algorithms for order-preserving pattern matching based on the duel-and-sweep paradigm. Acta Informatica, 61(4):415–444, December 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/
article/10.1007/s00236-024-00464-w.

Jagadish:2000:ODM

[JKNS00] H. V. Jagadish, Olga Kapitskaia, Raymond T. Ng, and Divesh Srivastava. One-dimensional and multi-dimensional substring selectivity estimation. VLDB Journal: Very Large Data Bases, 9(3):214–230, December 2000. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.de/link/service/
journals/00778/bibs/0009003/00090214.htm;   http:/
/link.springer.de/link/service/journals/00778/papers/
0009003/00090214.pdf.

Jiang:1993:OWH

[JL93] Tao Jiang and Ming Li. k one-way heads cannot do stringmatching. In ACM [ACM93b], pages 62–70. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. URL http://www.

acm.org/pubs/articles/proceedings/stoc/167088/p62-
jiang/p62-jiang.pdf;   http://www.acm.org/pubs/citations/
proceedings/stoc/167088/p62-jiang/. ACM order no.

508930.

Jiang:1996:OWH

[JL96] Tao Jiang and Ming Li. k one-way heads cannot do string-matching. Journal of Computer and System Sciences, 53(3):513–524, December 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000096900847.

Jiang:2014:SSJ

[JLFL14] Yu Jiang, Guoliang Li, Jianhua Feng, and Wen-Syan Li. String similarity joins: an experimental evaluation. Proceedings of the VLDB Endowment, 7(8):625–636, April 2014. CO-DEN ???? ISSN 2150-8097.

Jung:2018:EEK

[JLH18] Sangkeun Jung, Changki Lee, and Hyunsun Hwang. End-to-end Korean part-of-speech tagging using copying mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):19:1–19:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jambunathan:1992:DIF

[JLHB92] K. Jambunathan, E. Lai, S. L. Hartle, and B. L. Button. Development of an intelligent front end using LISP. Applications of Artificial Intelligence in Engineering, pages 228–243, 1992. CODEN AAIEEO.

Jeong:2020:RSH

[JLK^{+}20] W. S. Jeong, C. Lee, K. Kim, M. K. Yoon, W. Jeon, M. Jung, and W. W. Ro. REACT: Scalable and high-performance regular expression pattern matching accelerator for in-storage processing. IEEE Transactions on Parallel and Distributed Systems, 31(5):1137–1151, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ju:1985:CSF

[JM85] M. S. Ju and J. M. Mansour. Comparative studies of formulating the dynamics of rigid-body systems using Macsyma — a case study. Developments in Mechanics, 13:185–186, 1985. CODEN DEMEAX. ISSN 0419-0262.

Janicki:1990:TSS

[JM90] R. Janicki and T. Muldner. Transformations of sequential specifications into concurrent specifications by synchronization guards. Theoretical Computer Science, 77(1-2):97–129, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jategaonkar:1993:TIE

[JM93] Lalita A. Jategaonkar and John C. Mitchell. Type inference with extended pattern matching and subtypes. Fundamenta Informaticae, 19(1–2):127–165, September/October 1993. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jin:2024:QSU

[JN24] Ce Jin and Jakob Nogler. Quantum speed-ups for string synchronizing sets, longest common substring, and k-mismatch matching. ACM Transactions on Algorithms, 20(4):32:1– 32:??, October 2024. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/10.
1145/3672395.

Judd:2008:BGG

[JNS08] Christopher Judd, Joseph Faisal Nusairat, and James Shingler. Beginning Groovy and Grails: from Novice to Professional. Expert’s voice in open source. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-1045-1 (paperback). xxvi + 413 pp. LCCN QA76.73.G23 J84 2008.

Jouannaud:1997:ADT

[JO97] Jean-Pierre Jouannaud and Mitsuhiro Okada. Abstract data type systems. Theoretical Computer Science, 173(2): 349–391, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1997&volume=173&issue=2&
aid=2374.

Johansen:1969:FGR

[Joh69] Peter Johansen. Free groups and regular expressions. In ACM [ACM69], pages 113–128. LCCN QA75.5 .A22 1969.

Johnson:1994:SMC

[Joh94a] J. H. Johnson. Substring matching for clone detection and change tracking. In Muller and Georges [MG94], pages 120– 126. ISBN 0-8186-6330-8. LCCN QA76.76.S64I58 1994.

Johnson:1994:VTR

[Joh94b] J. H. Johnson. Visualizing textual redundancy in legacy source. In Botsford et al. [BGG^{+}94], pages 9–18.

Johansen:1995:LSM

[Joh95] Peter Johansen. On-line string matching with feedback. Theoretical Computer Science, 141(1–2):53–67, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1852.

Johnson:2001:ECL

[Joh01] Steven Johnson. Emergence: the connected lives of ants, brains, cities, and software. Scribner, New York, NY, USA, 2001. ISBN 0-684-86875-X, 0-684-86876-8 (paperback). 288 pp. LCCN Q325 .J65 2001.

Jokinen:1990:PTA

[Jok90] M. O. Jokinen. Parameter transmission abstractions. The Computer Journal, 33(2):133–139, April 1990. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/33/2/133.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/133.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/
tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/135.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/
tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/137.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/139.tif.

Jones:2007:CPS

[Jon07] Simon Peyton Jones. Call-pattern specialisation for Haskell programs. ACM SIGPLAN Notices, 42(9):327–337, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2013:FPU

[Jon13] Capers Jones. Function points as a universal software metric. ACM SIGSOFT Software Engineering Notes, 38(4):1–27, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jorgensen:1992:GCL

[Jør92] Jesper Jørgensen. Generating a compiler for a lazy language by partial evaluation. In ACM [ACM92a], pages 258–268. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992.  URL http://www.acm.org:80/pubs/citations/
proceedings/plan/143165/p258-jorgensen/. ACM order

number 54990.

James:1973:ACP

[JP73] E. B. James and Derek P. Partridge. Adaptive correction of program statements. Communications of the Association for Computing Machinery, 16(1):27–37, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jay:2011:TSI

[JP11] Barry Jay and Jens Palsberg. Typed self-interpretation by pattern matching. ACM SIGPLAN Notices, 46(9):247–258, September 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ICFP ’11 conference proceedings.

Jaskelioff:2015:FPS

[JR15a] Mauro Jaskelioff and Exequiel Rivas. Functional pearl: a smart view on datatypes. ACM SIGPLAN Notices, 50(9): 355–361, September 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jouannaud:2015:NHO

[JR15b] Jean-Pierre Jouannaud and Albert Rubio. Normal higherorder termination. ACM Transactions on Computational Logic, 16(2):13:1–13:??, March 2015. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

JaJa:1996:SSC

[JRV96] Joseph F. JáJá, Kwan Woo Ryu, and Uzi Vishkin. Sorting strings and constructing digital search trees in parallel. Theoretical Computer Science, 154(2):225–245, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1897.

Johnsen:1983:CTL

[JSC83] O. Johnsen, J. Segen, and G. L. Cash. Coding of two-level pictures by pattern matching and substitution. The Bell System Technical Journal, 62(8):2513–2545, October 1983. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-labs.

com/BSTJ/images/Vol62/bstj62-8-2513.pdf;   http://
www.alcatel-lucent.com/bstj/vol62-1983/articles/bstj62-
8-2513.pdf.

Jeon:2009:AAP

[JSH09] Jinseong Jeon, Keoncheol Shin, and Hwansoo Han. Abstracting access patterns of dynamic memory using regular expressions. ACM Transactions on Architecture and Code Optimization, 5(4):18:1–18:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jenks:1994:HMA

[JT94] Richard D. Jenks and Barry M. Trager. How to make AX-IOM into a Scratchpad. In ACM [ACM94b], pages 32–40. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/
190347/p32-jenks/.

Jokinen:1996:CAS

[JTU96] Petteri Jokinen, Jorma Tarhio, and Esko Ukkonen. A comparison of approximate string matching algorithms. Software — Practice and Experience, 26(12):1439–1458, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=16782.

Jokinen:1991:TAA

[JU91] P. Jokinen and E. Ukkonen. Two algorithms for approximate string matching in static texts. Lecture Notes in Computer Science, 520:240–??, 1991. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1994:ATA

[JWZ94] T. Jiang, L. Wang, and K. Zhang. Alignment of trees — an alternative to tree edit. Lecture Notes in Computer Science, 807:75–86, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2020:CSM

[JXA20] Peng Jiang, Yang Xia, and Gagan Agrawal. Combining SIMD and many/multi-core parallelism for finite-state machines with enumerative speculation. ACM Transactions on Parallel Computing (TOPC), 7(3):15:1–15:26, August 2020. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3399714.

Jarrar:2019:DBM

[JZAA19] Mustafa Jarrar, Fadi Zaraket, Rami Asia, and Hamzeh Amayreh. Diacritic-based matching of Arabic words. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):10:1–10:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3242177.

Kahrel:2006:AIR

[Kah06] Peter Kahrel. Automating InDesign with regular expressions. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2006. ISBN 0-596-52937-6. LCCN Z253.532.A34; Z253.532.A34 K34 2006eb. URL http://www.oreilly.com/

catalog/9780596529376.

Kahrel:2007:GIC

[Kah07] Peter Kahrel. GREP in InDesign CS3. O’Reilly shortcuts. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2007. ISBN 0-596-51706-8. LCCN Z253.532.A34; Z253.532.A34 K34 2007eb. URL http://www.oreilly.com/

catalog/9780596517069.

Kahrel:2008:GIC

[Kah09] Peter Kahrel. GREP in InDesign CS3/CS4. O’Reilly Media, Sebastopol, CA, USA, 2009. ISBN 0-596-15717-7. 53 pp. LCCN Z253.532.A34 K34 2008.

Kim:2017:RTO

[KAN^{+}17] Jinil Kim, Amihood Amir, Joong Chae Na, Kunsoo Park, and Jeong Seop Sim. On representations of ternary order relations in numeric strings. Mathematics in Computer Science, 11(2): 127–136, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kaplan:1969:REE

[Kap69] Donald M. Kaplan. Regular expressions and the equivalence of programs. Journal of Computer and System Sciences, 3(4):361–386, November 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000069800279.

Kapur:1992:ADC

[Kap92] D. Kapur, editor. Automated deduction, CADE-11: 11th International Conference on Automated Deduction, Saratoga Springs, NY, USA, June 15–18, 1992: proceedings, volume 607 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55602-8. LCCN QA76.9.A96I57 1992.

Karpinski:1982:DSM

[Kar82] M. Karpinski. Decidability of ‘Skolem matrix emptiness problem’ entails constructability of exact regular expression. Theoretical Computer Science, 17(1):99–102, January 1982. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kastrup:2008:MLP

[Kas08a] David Kastrup. makematch, a L^{A}TEX package for pattern matching with wildcards. TUGboat, 29(1):190–192, 2008.

ISSN 0896-3207. URL http://www.tug.org/TUGboat/tb29-
1/tb91kastrup-match.pdf.

Kastrup:2008:PML

[Kas08b] David Kastrup. makematch, a L^{A}TEX package for pattern matching with wildcards. TUGboat, 29(1):190–192, 2008.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91kastrup-match.pdf.

Kastrup:TB29-1-190

[Kas08c] David Kastrup. makematch, a L^{A}TEX package for pattern matching with wildcards. TUGboat, 29(1):190–192, 2008.

ISSN 0896-3207. URL https://tug.org/TUGboat/tb29-1/
tb91kastrup-match.pdf.

Khan:2007:NID

[KAT07] Latifur Khan, Mamoun Awad, and Bhavani Thuraisingham. A new intrusion detection system using support vector machines and hierarchical clustering. VLDB Journal: Very Large Data Bases, 16(4):507–521, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kaufmann:1992:EBM

[Kau92] Matt Kaufmann. An extension of the Boyer–Moore theorem prover to support first-order quantification. Journal of Automated Reasoning, 9(3):355–372, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670

(electronic). URL http://link.springer.com/article/
10.1007/BF00245295.

Krebber:2018:PMP

[KB18] Manuel Krebber and Henrik Barthels. MatchPy: Pattern matching in Python. Journal of Open Source Software, 3(26):

670:1–670:2, June 2018. CODEN ???? ISSN 2475-9066. URL

http://joss.theoj.org/papers/10.21105/joss.00670.

Karcioglu:2022:QFH

[KB22] Abdullah Ammar Karcioglu and Hasan Bulut. q-frame hash comparison based exact string matching algorithms for DNA sequences. Concurrency and Computation: Practice and Experience, 34(9):e6505:1–e6505:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kannan:2001:FFG

[KBB01] Parivallal Kannan, Shankar Balachandran, and Dinesh Bhatia. fGREP — fast generic routing demand estimation for placed FPGA circuits. Lecture Notes in Computer Science, 2147:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2147/21470037.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/2147/21470037.pdf.

Klein:2009:ABM

[KBN09] Shmuel T. Klein and Miri Kopel Ben-Nissan. Accelerating Boyer–Moore searches on binary texts. Theoretical Computer Science, 410(37):3563–3571, September 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krishnapuram:1987:HST

[KC87] Raghuram Krishnapuram and David Casasent. Hough space transformations for discrimination and distortion estimation. Computer Vision, Graphics, and Image Processing, 38(3): 299–316, June 1987. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Kyatkin:1999:PMC

[KC99] Alexander B. Kyatkin and Gregory S. Chirikjian. Pattern matching as a correlation on the discrete motion group. Computer Vision and Image Understanding: CVIU, 74(1):22–35, April 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http://www.idealibrary.

com/links/artid/cviu.1999.0745/production;   http:
//www.idealibrary.com/links/artid/cviu.1999.0745/production/
pdf;   http://www.idealibrary.com/links/artid/cviu.
1999.0745/production/ref.

Knessl:2011:EAF

[KC11] Charles Knessl and Mark W. Coffey. An effective asymptotic formula for the Stieltjes constants. Mathematics of Computation, 80(273):379–386, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.ams.org/journals/mcom/2011-80-273/S0025-5718-
2010-02390-7/;   http://www.ams.org/journals/mcom/
2011-80-273/S0025-5718-2010-02390-7/home.html;   http:
//www.ams.org/journals/mcom/2011-80-273/S0025-5718-
2010-02390-7/S0025-5718-2010-02390-7.pdf.

Khan:2015:UGM

[KC15] Arijit Khan and Lei Chen. On uncertain graphs modeling and queries. Proceedings of the VLDB Endowment, 8(12): 2042–2043, August 2015. CODEN ???? ISSN 2150-8097.

Kim:2021:SFI

[KC21] Sung-Hwan Kim and Hwan-Gue Cho. Simpler FM-index for parameterized string matching. Information Processing Letters, 165(??):Article 106026, January 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019020301137.

Kim:1993:MPM

[KCK93] Jeong Uk Kim, Ho Chang, and Tag Gon Kim. Multidisk partial match file design with known access pattern. Information Processing Letters, 45(1):33–39, January 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamthe:2013:IWL

[KCPC13] Ankur Kamthe, Miguel Á Carreira-Perpiñán, and Alberto E. Cerpa. Improving wireless link simulation using multilevel Markov models. ACM Transactions on Sensor Networks, 10 (1):17:1–17:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Koneru:2015:DCA

[KD15] Suvarna Vani Koneru and Bhavani S. Durga. Divide and conquer approach to contact map overlap problem using 2Dpattern mining of protein contact networks. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(4):729–737, July 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Kearns:1991:ERE

[Kea91a] Steven M. Kearns. Extending regular expressions with context operators and parse extraction. Software — Practice and Experience, 21(8):787–804, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kearns:1991:T

[Kea91b] Steven M. Kearns. TLex. Software — Practice and Experience, 21(8):805–821, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2014:OPM

[KEF^{+}14] Jinil Kim, Peter Eades, Rudolf Fleischer, Seok-Hee Hong, Costas S. Iliopoulos, Kunsoo Park, Simon J. Puglisi, and Takeshi Tokuyama. Order-preserving matching. Theoretical Computer Science, 525(??):68–79, March 13, 2014. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513007585.

Koutrika:2008:CST

[KEG^{+}08] Georgia Koutrika, Frans Adjie Effendi, Zolt´ n Gyöngyi, Paul Heymann, and Hector Garcia-Molina. Combating spam in tagging systems: an evaluation. ACM Transactions on the Web (TWEB), 2(4):22:1–22:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kerren:2004:GME

[Ker04] Andreas Kerren. Generation as method for explorative learning in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):77–81, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kernighan:2007:REM

[Ker07] Brian W. Kernighan. A regular expression matcher. In Oram and Wilson [OW07], pages 1–8. ISBN 0-596-51004-7 (paperback). LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http://www.oreilly.com/

catalog/9780596510046.

Kessler:1979:PPM

[Kes79] R. Kessler. PMETA — pattern matching META/REDUCE. Report USCG Op. Note No. 40, Univ. of Utah, CS Dept., Salt Lake City, UT, USA, January 1979.

Kesner:1991:PMO

[Kes91] D. Kesner. Pattern matching in order-sorted languages. Lecture Notes in Computer Science, 520:267–??, 1991. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klier:1991:FCB

[KF91] Peter Klier and Richard J. Fateman. On finding the closest bitwise matches in a fixed set. ACM Transactions on Mathematical Software, 17(1):88–97, March 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1991-17-1/p88-klier/.

Kocberber:2015:AMA

[KFG15] Onur Kocberber, Babak Falsafi, and Boris Grot. Asynchronous memory access chaining. Proceedings of the VLDB Endowment, 9(4):252–263, December 2015. CODEN ???? ISSN 2150-8097.

Kiezun:2012:HSW

[KGA^{+}12] Adam Kiezun, Vijay Ganesh, Shay Artzi, Philip J. Guo, Pieter Hooimeijer, and Michael D. Ernst. HAMPI: a solver for word equations over strings, regular expressions, and contextfree grammars. ACM Transactions on Software Engineering and Methodology, 21(4):25:1–25:??, November 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kumar:2005:PCO

[KGP^{+}05] Rajeev Kumar, Amit Gupta, B. S. Pankaj, Mrinmoy Ghosh, and P. P. Chakrabarti. Post-compilation optimization for multiple gains with pattern matching. ACM SIGPLAN Notices, 40(12):14–23, December 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawanaka:2006:BBT

[KH06] Shinya Kawanaka and Haruo Hosoya. biXid: a bidirectional transformation language for XML. ACM SIGPLAN Notices, 41(9):201–214, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2025:ACT

[KH25] Sungmin Kim and Yo-Sub Han. Approximate Cartesian tree pattern matching. Theoretical Computer Science, 1056(??):??, November 21, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030439752500444X.

Khan:2016:TOS

[Kha16] Minhaj Ahmad Khan. A transformation for optimizing stringmatching algorithms for long patterns. The Computer Journal, 59(12):1749–1759, December 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/59/12/1749.

Kida:2009:STB

[Kid09] T. Kida. Suffix tree based VF-coding for compressed pattern matching. In Storer and Marcellin [SM09], page 449. ISBN 0-7695-3592-5. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976503. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Kusudo:2015:BPA

[KIH15] Ko Kusudo, Fumihiko Ino, and Kenichi Hagihara. A bitparallel algorithm for searching multiple patterns with various lengths. Journal of Parallel and Distributed Computing, 76(??):49–57, February 2015. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151400210X.

Kim:1999:NSP

[Kim99] Sun Kim. A new string-pattern matching algorithm using partitioning and hashing efficiently. ACM Journal of Experimental Algorithmics, 4:2:1–2:??, ???? 1999. CODEN ???? ISSN 1084-6654.

King:1989:UNN

[Kin89] Todd King. Using neural networks for pattern recognition. Dr. Dobb’s Journal of Software Tools, 14(1):14–28, 90–95, January 1989. CODEN DDJOEB. ISSN 0888-3076.

Kinber:1991:CSS

[Kin91] E. B. Kinber. On complete sets of samples for generalized regular expressions. Theoretical Computer Science, 91(1):101– 117, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kinber:1992:LCR

[Kin92] Efim Kinber. Learning a class of regular expressions via restricted subset queries. Lecture Notes in Computer Science, 642:232–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kitani:1994:MID

[Kit94] T. Kitani. Merging information by discourse processing for information extraction. In IEEE [IEE94b], pages 412–418. ISBN 0-8186-5550-X. LCCN Q 334 C66 1994. IEEE Catalog No. 94CH3421-5.

Kim:2017:MES

[KJS17] Minchul Kim, Younghoon Jung, and Junghwan Song. A modified exhaustive search on a password system using SHA-1. International Journal of Information Security, 16(3):263– 269, June 2017. CODEN ???? ISSN 1615-5262 (print),

1615-5270 (electronic). URL http://link.springer.com/
article/10.1007/s10207-016-0332-2.

Konda:1995:SFD

[KK95] Venkat Konda and Anup Kumar. A systematic framework for the dependence cycle removal in practical loops. Journal of Parallel and Distributed Computing, 27(2):157–171, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1995.1079/production;
http://www.idealibrary.com/links/doi/10.1006/jpdc.
1995.1079/production/pdf.

Krodel:2002:RLD

[KK02] M. Krödel and K.-D. Kuhnert. Reinforcement learning to drive a car by pattern matching. Lecture Notes in Computer Science, 2449:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/0558/bibs/2449/24490322.
htm;   http://link.springer.de/link/service/series/
0558/papers/2449/24490322.pdf.

Kim:2008:LSS

[KK08] Eagu Kim and John Kececioglu. Learning scoring schemes for sequence alignment from partial examples. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 5(4):546–556, October 2008. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Kravchuk-Kirilyuk:2024:PNF

[KKFI^{+}24] Anastasiya Kravchuk-Kirilyuk, Gary Feng, Jonas Iskander, Yizhou Zhang, and Nada Amin. Persimmon: Nested family polymorphism with extensible variant types. Proceedings of the ACM on Programming Languages (PACMPL), 8 (OOPSLA1):119:1–119:??, April 2024. CODEN ???? ISSN

2475-1421 (electronic). URL https://dl.acm.org/doi/10.
1145/3649836.

Kim:2024:SCP

[KKH24] Sungmin Kim, Sang-Ki Ko, and Yo-Sub Han. Simon’s congruence pattern matching. Theoretical Computer Science, 994(??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304397524000938.

Kim:2011:MEB

[KKK11] Hyun Jin Kim, Hong-Sik Kim, and Sungho Kang. A memoryefficient bit-split parallel string matching using pattern dividing for intrusion detection systems. IEEE Transactions on Parallel and Distributed Systems, 22(11):1904–1911, November 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kapur:1985:ATC

[KKM^{+}85] D. Kapur, M. S. Krishnamoorthy, R. McNaughton, Narendran, and P. An O(modT mod ^{3}) algorithm for testing the Church–Rosser property of Thue systems. Theoretical Computer Science, 35(1):109–114, January 1985. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawahito:2006:NIR

[KKM^{+}06] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. A new idiom recognition framework for exploiting hardware-assist instructions. ACM SIGPLAN Notices, 41(11):382–393, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawahito:2013:IRF

[KKM^{+}13] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. Idiom recognition framework using topological embedding. ACM Transactions on Architecture and Code Optimization, 10(3):13:1–13:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2023:OPP

[KKNS23] Youngho Kim, Munseong Kang, Joong Chae Na, and Jeong Seop Sim. Order-preserving pattern matching with scaling. Information Processing Letters, 180(??):Article 106333, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print),

1872-6119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0020019022000904.

Kim:1992:DSN

[KKP92] Won Kim, Y. Kambayashi, and In Sup Paik, editors. Database systems for next-generation applications: principles and practice, volume 1 of Advanced Database Research and Development Series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Karkkainen:2016:LLZ

[KKP16a] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, 2016. CODEN ???? ISSN 1084-6654.

Karkkainen:2017:LLZ

[KKP16b] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, 2016. CODEN ???? ISSN 1084-6654.

Kubica:2013:LTA

[KKR^{+}13] M. Kubica, T. Kulczyński, J. Radoszewski, W. Rytter, and T. Waleń. A linear time algorithm for consecutive permutation pattern matching. Information Processing Letters, 113(12):430–433, June 30, 2013. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019013000926.

Kim:2001:FRQ

[KKSL01] Harksoo Kim, Kyungsun Kim, Jungyun Seo, and Gary Geunbae Lee. A fast and reliable question-answering system based on predictive answer indexing and lexico-syntactic pattern matching. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4):341–??, 2001. CO-DEN ???? ISSN 0219-4279.

KleinOsowski:2002:MNS

[KL02] A. J. KleinOsowski and D. J. Lilja. MinneSPEC: A new SPEC benchmark workload for simulation-based computer architecture research. IEEE Computer Architecture Letters, 1(1):7, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Krotzsch:2012:PPA

[KLB12] Markus Krötzsch, Maurizio Lenzerini, and Michael Benedikt, editors. PODS’12: Proceedings of the 31st ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems: May 20–24, 2012, Scottsdale, AZ, USA. ACM Press, New York, NY 10036, USA, 2012. ISBN ???? LCCN ????

URL http://www.sigmod.org/2012/.

Kleene:1956:REN

[Kle56] Stephen C. Kleene. Realization of events in nerve nets and finite automata. In Shannon and McCarthy [SM56], pages 3–42. CODEN ANMAAH. ISBN 0-691-07916-1. ISSN 0066-2313. German translation in [SM74].

Ko:2016:SCR

[KLH16] Sang-Ki Ko, Ha-Rim Lee, and Yo-Sub Han. State complexity of regular tree languages for tree matching. International Journal of Foundations of Computer Science (IJFCS), 27(8): 965–980, December 2016. CODEN IFCSEN. ISSN 0129-0541.

Kar:2023:COO

[KLL23] Binayak Kar, Ying-Dar Lin, and Yuan-Cheng Lai. Cost optimization of omnidirectional offloading in two-tier cloudedge federated systems. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804523000498.

Krishnamurthy:2008:SSD

[KLR^{+}08] Rajasekar Krishnamurthy, Yunyao Li, Sriram Raghavan, Frederick Reiss, Shivakumar Vaithyanathan, and Huaiyu Zhu. SystemT: a system for declarative information extraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):7–13, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kowalski:1984:NMT

[KM84] G. Kowalski and A. Meltzer. New multi-term high speed text search algorithms. In IEEE, editor, First International Conference on Computer Applications (June 20–22, 1984), pages 514–522. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984.

Knight:1992:ARE

[KM92] James R. Knight and Eugene W. Myers. Approximate regular expression pattern matching with concave gap penalties. Lecture Notes in Computer Science, 644:66–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kilpelainen:1994:QPT

[KM94] P. Kilpelainen and H. Mannila. Query primitives for treestructured data. Lecture Notes in Computer Science, 807: 213–225, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knight:1995:ARE

[KM95a] James R. Knight and Eugene W. Myers. Approximate regular expression pattern matching with concave gap penalties. Algorithmica, 14(1):85–121, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Knight:1995:SPM

[KM95b] James R. Knight and Eugene W. Myers. Super-pattern matching. Algorithmica, 13(1–2):211–243, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kouzinopoulos:2013:EOT

[KM13] Charalampos S. Kouzinopoulos and Konstantinos G. Margaritis. Exact online two-dimensional pattern matching using multiple pattern matching algorithms. ACM Journal of Experimental Algorithmics, 18(1):2.4:1–2.4:??, December 2013. CODEN ???? ISSN 1084-6654.

Kouzinopoulos:2015:MSM

[KMM15] Charalampos S. Kouzinopoulos, Panagiotis D. Michailidis, and Konstantinos G. Margaritis. Multiple string matching on a GPU using CUDAs. Scalable Computing: Practice and Experience, 16(2):121–138, ???? 2015. CODEN ???? ISSN

1895-1767. URL https://www.scpe.org/index.php/scpe/
article/view/1085.

Kristensen:1985:APF

[KMMPN85] Bent Bruun Kristensen, Ole Lehrmann Madsen, Birger Møller-Pedersen, and Kristen Nygaard. An algebra for program fragments. ACM SIGPLAN Notices, 20(7):161–170, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knuth:1977:FPM

[KMP77] Donald E. Knuth, James H. Morris, Jr., and Vaughan R. Pratt. Fast pattern matching in strings. SIAM Journal on Computing, 6(2):323–350, June 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Errata, see [Ryt80]. See also [BM77, Ryt80, Sun90a, HS91, BYG92].

Knuth:1994:FPM

[KMP94] Donald E. Knuth, James H. Morris, Jr., and Vaughan R. Pratt. Fast pattern matching in strings. In ichi Aoe [iA94], pages 8–35. ISBN 0-8186-5461-9 (microfiche), 0-8186-5462-7 (hardcover), 0-8186-5460-0 (paperback). LCCN QA76.9.A43 C67 1994. US$56.00. IEEE Computer Society Press order number: 5462-05. IEEE catalog number: 94EH0389-7.

Konstantinidis:2021:PDR

[KMR21] Stavros Konstantinidis, Nelma Moreira, and Rogério Reis. Partial derivatives of regular expressions over alphabetinvariant and user-defined labels. Theoretical Computer Science, 870(??):103–120, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397520307544.

Konstantinidis:2020:RET

[KMRY20] Stavros Konstantinidis, Nelma Moreira, Rogério Reis, and Joshua Young. Regular expressions and transducers over alphabet-invariant and user-defined labels. International Journal of Foundations of Computer Science (IJFCS), 31(08):983–1019, December 2020. ISSN 0129-

0541. URL https://www.worldscientific.com/doi/10.
1142/S0129054120420010.

Kida:2003:CSU

[KMS^{+}03] Takuya Kida, Tetsuya Matsumoto, Yusuke Shibata, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Collage system: a unifying framework for compressed pattern matching. Theoretical Computer Science, 298(1):253–272, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kida:2001:MPM

[KMT^{+}01] Takuya Kida, Tetsuya Matsumoto, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Multiple pattern matching algorithms on collage system. Lecture Notes in Computer Science, 2089:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2089/
20890193.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2089/20890193.pdf.

Katoen:2000:PMA

[KN00] Joost-Pieter Katoen and Albert Nymeyer. Pattern-matching algorithms based on term rewrite systems. Theoretical Computer Science, 238(1–2):439–464, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-ng/10/41/16/172/21/
34/abstract.html;   http://www.elsevier.nl/gej-ng/
10/41/16/172/21/34/article.pdf.

Krauss:2012:PPR

[KN12] Alexander Krauss and Tobias Nipkow. Proof pearl: Regular expression equivalence and relation algebra. Journal of Automated Reasoning, 49(1):95–106, June 2012. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10817-011-9223-4.

Klabnik:2017:RPL

[KN17] Steve Klabnik and Carol Nichols. The Rust Programming Language. No Starch Press, San Francisco, CA, USA, 2017. ISBN 1-59327-828-4 (paperback), 1-59327-851-9 (e-pub). xxvii + 519 pp. LCCN QA76.73.R87 K53 2018.

Klabnik:2019:RPL

[KN19] Steve Klabnik and Carol Nichols. The Rust programming language. No Starch Press, San Francisco, CA, USA, second edition, 2019. ISBN 1-0981-2253-4, 1-71850-044-0 (paperback). xxix + 526 pp. LCCN QA76.73.R87. URL http://proquest.

safaribooksonline.com/?fpi=9781098122539;   https:
//nostarch.com/download/samples/RustProgrammingLanguage2018_
Sample_ToC.pdf;   https://nostarch.com/Rust2018.

Knight:1989:UMS

[Kni89] Kevin Knight. Unification: a multidisciplinary survey. ACM Computing Surveys, 21(1):93–124, March 1989. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0360-0300/62030.html.

Kuri:2000:PMB

[KNMH00] Josué Kuri, Gonzalo Navarro, Ludovic Mé, and Laurent Heye. A pattern matching based filter for audit reduction and fast detection of potential intrusions.

Lecture Notes in Computer Science, 1907:17–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1907/19070017.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1907/19070017.pdf.

Kucherov:2012:CDP

[KNS12] Gregory Kucherov, Yakov Nekrich, and Tatiana Starikovskaya. Cross-document pattern matching. Lecture Notes in Computer Science, 7354:196–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_16/.

Kiwi:2011:LAS

[KNT11] Marcos Kiwi, Gonzalo Navarro, and Claudio Telha. Online approximate string matching with bounded errors. Theoretical Computer Science, 412(45):6359–6370, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397511006669.

Katsarou:2015:PSI

[KNT15] Foteini Katsarou, Nikos Ntarmos, and Peter Triantafillou. Performance and scalability of indexed subgraph query processing methods. Proceedings of the VLDB Endowment, 8(12): 1566–1577, August 2015. CODEN ???? ISSN 2150-8097.

Knuth:1998:SS

[Knu98] Donald E. Knuth. Sorting and Searching, volume 3 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-89685-0. xiv + 780 pp. LCCN QA76.6.K64 1997. US$49.95.

Knuth:2005:ACPb

[Knu05] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 3. Generating All Combinations and Partitions. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-85394-9. iv + 150 pp. LCCN QA76.6 .K64 2005.

Kashyap:1983:NSM

[KO83] R. L. Kashyap and B. J. Oommen. The noisy substring matching problem. IEEE Transactions on Software Engineering, SE-9(3):365–370, May/June 1983. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1703065.

Kodratoff:1979:CFS

[Kod79] Yves Kodratoff. A class of functions synthesized from a finite number of examples and a LISP program scheme. International Journal of Computer and Information Sciences, 8(6): 489–521, December 1979. CODEN IJCIAH. ISSN 0091-7036.

Kakeshita:1994:FCS

[KOI94] T. Kakeshita, M. Oda, and Y. Imamura. Fall-in C: a software tool for pitfall detection in C programs. In IEEE [IEE94a], pages 256–265. ISBN 0-8186-6960-8 (paper), 0-8186-6961-6 (microfiche). LCCN QA76.758.A77 1994.

Kornman:1983:PMP

[Kor83] Brent D. Kornman. Pattern matching and pattern-directed invocation in systems programming languages. The Journal of Systems and Software, 3(1):95–102, March 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kosaraju:1989:ETP

[Kos89] S. R. Kosaraju. Efficient tree pattern matching. In IEEE [IEE89], pages 178–183. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kosaraju:1994:RTP

[Kos94] S. Rao Kosaraju. Real-time pattern matching and quasireal-time construction of suffix trees (preliminary version). In ACM [ACM94d], pages 310–316. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. URL http://

www.acm.org/pubs/articles/proceedings/stoc/195058/
p310-kosaraju/p310-kosaraju.pdf;   http://www.acm.
org/pubs/citations/proceedings/stoc/195058/p310-kosaraju/

. ACM order no. 508930.

Kebler:1993:APP

[KP93] C. W. Kebler and W. J. Paul. Automatic parallelization by pattern-matching. Lecture Notes in Computer Science, 734: 166–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaufmann:1996:IBMa

[KP96a] Matt Kaufmann and Paolo Pecchiari. Interaction with the Boyer–Moore theorem prover: a tutorial study using the arithmetic–geometric mean theorem. In Zhang [Zha96], pages 181–222. ISBN 94-010-7250-7 (print), 94-009-1675-2 (e-

book). LCCN Q334-342. URL http://link.springer.com/
chapter/10.1007/978-94-009-1675-3_6.

Kaufmann:1996:IBMb

[KP96b] Matt Kaufmann and Paolo Pecchiari. Interaction with the Boyer–Moore theorem prover: A tutorial study using the arithmetic–geometric mean theorem. Journal of Automated Reasoning, 16(1–2):181–222, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00244463.

Kernighan:1999:PP

[KP99a] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61586-X. xii + 267 pp. LCCN QA76.6 .K48 1999. US$24.95,

CAN$37.50. URL http://cm.bell-labs.com/cm/cs/tpop/
code.html;   http://cseng.aw.com/bookdetail.qry?ISBN=
0-201-61586-X&ptype=0;   http://tpop.awl.com.

Kernighan:1999:RE

[KP99b] Brian W. Kernighan and Rob Pike. Regular expressions. In The Practice of Programming [KP99a], chapter 9.2, pages 222–227. ISBN 0-201-61586-X. LCCN QA76.6 .K48 1999. US$24.95, CAN$37.50. URL http:

//cm.bell-labs.com/cm/cs/tpop/code.html;   http://
cseng.aw.com/bookdetail.qry?ISBN=0-201-61586-X&ptype=
0;   http://tpop.awl.com.

Kernighan:1999:REL

[KP99c] Brian W. Kernighan and Rob Pike. Regular expressions: Languages, algorithms, software. Dr. Dobb’s Journal of Software Tools, 24(4):19–22, April 1999. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/ftp/1999/1999_04/
regexp.txt;   http://www.ddj.com/ftp/1999/1999_04/
regexp.zip;   http://www.drdobbs.com/architecture-
and-design/regular-expressions/184410904;   http:/
/www.drdobbs.com/architecture-and-design/regular-expressions/
architecture-and-design/sourcecode/regular-expressions/
30200909;   http://www.drdobbs.com/architecture-and-
design/regular-expressions/architecture-and-design/
sourcecode/regular-expressions/30200910. See also

[Tho68, Cox07, Cox09, Cox10a, Cox12].

Kroening:2015:CAV

[KP15] Daniel Kroening and Corina S. Păsăreanu, editors. Computer Aided Verification: 27th International Conference, CAV 2015, San Francisco, CA, USA, July 18–24, 2015, Proceedings, Part I, volume 9206 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-21689-9. URL http:/

/link.springer.com/book/10.1007/978-3-319-21690-4.

Kumar:2010:SMM

[KPA10] Krishna Kumar, Rajesh Prasad, and Suneeta Agarwal. Software maintenance by multi-patterns parameterized string matching with q-gram. ACM SIGSOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaloudas:2019:EEB

[KPP19] Dimitrios Kaloudas, Nikolet Pavlova, and Robert Penchovsky. EBWS: Essential bioinformatics Web services for sequence analyses. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(3):942–953, May 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Kuperberg:2021:CPS

[KPP21] Denis Kuperberg, Laureline Pinault, and Damien Pous. Cyclic proofs, system t, and the power of contraction. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):

1:1–1:28, January 2021. URL https://dl.acm.org/doi/10.
1145/3434282.

Karhumaeki:1997:PMP

[KPR97] J. Karhumaeki, W. Plandowski, and W. Rytter. Patternmatching problems for 2-dimensional images described by finite automata. Lecture Notes in Computer Science, 1279: 245–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karhumaki:2000:PMP

[KPR00] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. Pattern-matching problems for two-dimensional images described by finite automata. Nordic Journal of Computing, 7(1):1–??, Spring 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.fi/njc/References/
karhumakipr2000:1.html.

Karp:1981:ERPa

[KR81a] R. M. Karp and M. O. Rabin. Efficient randomized patternmatching algorithms. Technical report TR-31-81, Harvard University, Cambridge, MA, USA, ?? 1981. ?? pp. An incremental hash function is described for application to the string search problem. See [BYG92].

Karp:1981:ERPb

[KR81b] R. M. Karp and M. O. Rabin. Efficient randomized patternmatching algorithms. Technical Report TR-31-81, Harvard University, Cambridge, MA, USA, 1981. See [BYG92].

Karp:1987:ERP

[KR87] Richard M. Karp and Michael O. Rabin. Efficient randomized pattern-matching algorithms. IBM Journal of Research and Development, 31(2):249–260, March 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Katajainen:1989:AAS

[KR89] J. Katajainen and T. Raita. An approximation algorithm for space-optimal encoding of a text. The Computer Journal, 32(3):228–237, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/3/228.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/228.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_03/tiff/229.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_03/tiff/231.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_03/tiff/233.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/234.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_03/tiff/235.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_03/tiff/237.tif.

Katajainen:1992:ALM

[KR92] Jyrki Katajainen and Timo Raita. An analysis of the longest match and the greedy heuristics in text encoding. Journal of the Association for Computing Machinery, 39(2):281–294, April 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0004-5411/128751.html.

Karpinski:1994:AIO

[KR94] M. Karpinski and W. Rytter. An alphabet-independent optimal parallel search for three dimensional pattern. Lecture Notes in Computer Science, 807:125–135, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucherov:1995:UGR

[KR95] Gregory Kucherov and Michaël Rusinowitch. Undecidability of ground reducibility for word rewriting systems with variables. Information Processing Letters, 53(4):209–215, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucherov:1997:MSS

[KR97] Gregory Kucherov and Michaël Rusinowitch. Matching a set of strings with variable length don’t cares. Theoretical Computer Science, 178(1–2):129–154, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2297.

Kimelfeld:2014:TMS

[KR14] Benny Kimelfeld and Christopher Ré. Transducing Markov sequences. Journal of the Association for Computing Machinery, 61(5):32:1–32:??, August 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kahrs:2022:SRE

[KR22] Stefan Kahrs and Colin Runciman. Simplifying regular expressions further. Journal of Symbolic Computation, 109(??):124–143, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0747717121000572.

Krauss:2008:PMP

[Kra08] Alexander Krauss. Pattern minimization problems over recursive data types. ACM SIGPLAN Notices, 43(9):267–274, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaswami:2009:FPM

[Kri09] Neelakantan R. Krishnaswami. Focusing on pattern matching. ACM SIGPLAN Notices, 44(1):366–378, January 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kovaleski:1987:AIS

[KRL87] A. Kovaleski, S. Ratheal, and F. Lombardi. An architecture and an interconnection scheme for time-sliced buses. Journal of Parallel and Distributed Computing, 4(2):209–229, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kwon:2009:FXD

[KRML09] Joonho Kwon, Praveen Rao, Bongki Moon, and Sukho Lee. Fast XML document filtering by sequencing twig patterns. ACM Transactions on Internet Technology (TOIT), 9(4): 13:1–13:??, September 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kociumaka:2017:EIJ

[KRR17] Tomasz Kociumaka, Jakub Radoszewski, and Wojciech Rytter. Efficient indexes for jumbled pattern matching with constant-sized alphabet. Algorithmica, 77(4):1194–1215, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s00453-016-0140-0.pdf.

Kociumaka:2024:IPM

[KRRW24] Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Internal pattern matching queries in a text and applications. SIAM Journal on Computing, 53(5): 1524–1577, ???? 2024. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https://epubs.siam.

org/doi/10.1137/23M1567618.

Karpinski:1995:PMS

[KRS95] M. Karpinski, W. Rytter, and A. Shinohara. Patternmatching for strings with short descriptions. Lecture Notes in Computer Science, 937:205–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karpinski:1997:EPM

[KRS97] Marek Karpinski, Wojciech Rytter, and Ayumi Shinohara. An efficient pattern-matching algorithm for strings with short descriptions. Nordic Journal of Computing, 4(2): 172–186, Summer 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.fi/njc/References/
karpinskirs1997:172.html.

Kociumaka:2019:CLC

[KRS19a] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Correction to: Longest Common Substring with Approximately k Mismatches. Algorithmica, 81(7):3074, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00560-

1.pdf. See [KRS19b].

Kociumaka:2019:LCS

[KRS19b] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Longest common substring with approximately k mismatches. Algorithmica, 81(6):2633–2652, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kociumaka:2023:PCL

[KRS23] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Publisher correction: Longest common substring with approximately k mismatches. Algorithmica, 85 (10):3323, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01119-x. See

[KRS19b].

Klarlund:1993:GT

[KS93] Nils Klarlund and Michael I. Schwartzbach. Graph types. In ACM [ACM93a], pages 196–205. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7

.A15 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/plan/158511/p196-klarlund/. ACM order

number 549930.

Kececioglu:1994:EBO

[KS94] J. Kececioglu and D. Sankoff. Efficient bounds for oriented chromosome inversion distance. Lecture Notes in Computer Science, 807:307–325, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kontoyiannis:1996:SEE

[KS96] I. Kontoyiannis and Y. M. Suhov. Stationary entrophy estimation via string matching. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1996.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=488376. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Kim:1999:ASP

[KS99] Sun Kim and Alberto Maria Segre. AMASS: a structured pattern matching approach to shotgun sequence assembly. Journal of Computational Biology, 6(2):163–186, January 1999. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.liebertpub.com/doi/abs/10.

1089/cmb.1999.6.163;   https://www.liebertpub.com/
doi/pdf/10.1089/cmb.1999.6.163.

Klein:2001:PMH

[KS01] S. T. Klein and D. Shapira. Pattern matching in Huffman encoded texts. In Storer and Cohn [SC01], pages 449–458. ISBN 0-7695-1031-0. ISSN 1068-0314. LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917176. IEEE CSP number 01PR1031.

Klein:2005:CPM

[KS05] S. T. Klein and D. Shapira. Compressed pattern matching in JPEG images. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314. LCCN ???? URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402223.

Klein:2006:CPM

[KS06] Shmuel T. Klein and Dana Shapira. Compressed pattern matching in JPEG images. International Journal of Foundations of Computer Science (IJFCS), 17(6):1297–1306, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kasneci:2007:CRA

[KS07] Gjergji Kasneci and Thomas Schwentick. The complexity of reasoning about pattern-based XML schemas. In ACM [ACM07], pages 155–164. ISBN 1-59593-685-8. LCCN ????

Karakoidas:2008:FJO

[KS08] Vassilios Karakoidas and Diomidis Spinellis. FIRE/J — optimizing regular expression searches with generative programming. Software — Practice and Experience, 38(6):557–573, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kimelfeld:2011:FMT

[KS11a] Benny Kimelfeld and Yehoshua Sagiv. Finding a minimal tree pattern under neighborhood constraints. In Lenzerini [Len11], pages 235–246. ISBN 1-4503-0660-8. LCCN ???? URL http:

//dl.acm.org/citation.cfm?id=1989284.

Klein:2011:SDM

[KS11b] S. T. Klein and D. Shapira. The string-to-dictionary matching problem. In Storer and Marcellin [SM11], pages 143– 152. ISBN 1-61284-279-8. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749472. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Karkkainen:2012:CPM

[KS12a] Juha Kärkkäinen and Jens Stoye, editors. Combinatorial Pattern Matching: 23rd Annual Symposium, CPM 2012, Helsinki, Finland, July 3–5, 2012. Proceedings, volume 7354 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-31264-0 (print), 3-642-31265-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31265-6.

Klein:2012:SDM

[KS12b] Shmuel T. Klein and Dana Shapira. The string-to-dictionary matching problem. The Computer Journal, 55 (11):1347–1356, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/55/11/1347.full.
pdf+html.

Kim:2015:TSI

[KSH^{+}15] Jinha Kim, Hyungyu Shin, Wook-Shin Han, Sungpack Hong, and Hassan Chafi. Taming subgraph isomorphism for RDF query processing. Proceedings of the VLDB Endowment, 8 (11):1238–1249, July 2015. CODEN ???? ISSN 2150-8097.

Kim:1992:ASM

[KST92] J. Y. Kim and J. Shawe-Taylor. An approximate stringmatching algorithm. Theoretical Computer Science, 92(1): 107–117, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1994:FSM

[KST94] Jong Yong Kim and John Shawe-Taylor. Fast string matching using an n-gram algorithm. Software — Practice and Experience, 24(1):79–88, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kurz:2012:CLI

[KST12] Alexander Kurz, Tomoyuki Suzuki, and Emilio Tuosto. A characterisation of languages on infinite alphabets with nominal regular expressions. Lecture Notes in Computer Science, 7604:193–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33475-7_14/.

Kucherov:2016:ASM

[KST16] Gregory Kucherov, Kamil Salikhov, and Dekel Tsur. Approximate string matching using a bidirectional index. Theoretical Computer Science, 638(??):145–158, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397515009494.

Karachalias:2015:GMT

[KSVJ15] Georgios Karachalias, Tom Schrijvers, Dimitrios Vytiniotis, and Simon Peyton Jones. GADTs meet their match: patternmatching warnings that account for GADTs, guards, and laziness. ACM SIGPLAN Notices, 50(9):424–436, September 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamel:1993:SRH

[KSWC93] M. S. Kamel, H. C. Shen, A. K. C. Wong, and R. I. Campeanu. System for the recognition of human faces. IBM Systems Journal, 32(2):307–320, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5515.

Kuo:1990:NSC

[KT90] R. T. Kuo and S. S. Tseng. The necessary and sufficient condition for the worst-case male optimal stable matching. Information Processing Letters, 34(5):261–263, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaminski:2006:REL

[KT06] Michael Kaminski and Tony Tan. Regular expressions for languages over infinite alphabets. Fundamenta Informaticae, 69 (3):301–318, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keil:2014:EDA

[KT14] Matthias Keil and Peter Thiemann. Efficient dynamic access analysis using JavaScript proxies. ACM SIGPLAN Notices, 49(2):49–60, February 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Kassaie:2023:ACI

[KT23] Besat Kassaie and Frank Wm. Tompa. Autonomously computable information extraction. Proceedings of the VLDB Endowment, 16(10):2431–2443, June 2023. CODEN ????

ISSN 2150-8097. URL https://dl.acm.org/doi/10.14778/
3603581.3603585.

Kandhan:2010:SFS

[KTP10] Ramakrishnan Kandhan, Nikhil Teletia, and Jignesh M. Patel. SigMatch: fast and scalable multi-pattern matching. Proceedings of the VLDB Endowment, 3(1–2):1173–1184, September 2010. CODEN ???? ISSN 2150-8097.

Kida:1998:MPM

[KTS^{+}98] T. Kida, M. Takeda, A. Shinohara, M. Miyazaki, and S. Arikawa. Multiple pattern matching in LZW compressed text. In Storer and Cohn [SC98], pages 103–112. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672136. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kida:1999:SAP

[KTSA99] Takuya Kida, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Shift-and approach to pattern matching in LZW compressed text. Lecture Notes in Computer Science, 1645: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1645/16450001.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1645/16450001.pdf.

Karlin:1987:ACR

[KTU87] Anna R. Karlin, Howard W. Trickey, and Jeffrey D. Ullman. Algorithms for the compilation of regular expressions into PLAs. Algorithmica, 2(3):283–314, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Koide:2018:EIQ

[KTY^{+}18] Satoshi Koide, Yukihiro Tadokoro, Takayoshi Yoshimura, Chuan Xiao, and Yoshiharu Ishikawa. Enhanced indexing and querying of trajectories in road networks via string algorithms. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(1):3:1–3:??, June 2018. CODEN ???? ISSN

2374-0353. URL https://dl.acm.org/citation.cfm?id=
3200200.

Karkkainen:1999:THD

[KU99] Juha Kärkkäinen and Esko Ukkonen. Two- and higherdimensional pattern matching in optimal expected time. SIAM Journal on Computing, 29(2):571–589, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111

(electronic). URL http://epubs.siam.org/sam-bin/dbq/
article/27587.

Kucherov:2009:CPM

[KU09] Gregory Kucherov and Esko Ukkonen, editors. Combinatorial Pattern Matching: 20th Annual Symposium, CPM 2009 Lille, France, June 22–24, 2009 Proceedings, volume 5577 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNCSD9. ISBN 3-642-02440-8 (print), 3-642-02441-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-642-02441-2.

Kukich:1992:TAC

[Kuk92] Karen Kukich. Techniques for automatically correcting words in text. ACM Computing Surveys, 24(4):377–439, December 1992. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0360-0300/146380.html.

Kuklewicz:2017:RP

[Kuk17] Chris Kuklewicz. Regex Posix. Web site, March 10,

2017. URL https://wiki.haskell.org/Regex_Posix. The

site compares several implementations of regular expression matching against the POSIX specification, and documents errors in all of the implementations.

Kulekci:2010:BNB

[Kül10] M. Oguzhan Külekci. BLIM: a new bit-parallel pattern matching algorithm overcoming computer word size limitation. Mathematics in Computer Science, 3(4):407–420, June 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1661-8270&volume=3&issue=4&spage=407.

Kulick:2011:ESC

[Kul11] Seth Kulick. Exploiting separation of closed-class categories for Arabic tokenization and part-of-speech tagging. ACM Transactions on Asian Language Information Processing, 10 (1):4:1–4:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Kunen:1995:RTB

[Kun95] Kenneth Kunen. A Ramsey theorem in Boyer–Moore logic. Journal of Automated Reasoning, 15(2):217–235, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670

(electronic). URL http://link.springer.com/article/
10.1007/BF00881917.

Kumar:2015:IAM

[KV15] Ajay Kumar and Anil Kumar Verma. An improved algorithm for the metamorphosis of semi-extended regular expressions to deterministic finite automata. The Computer Journal, 58 (3):448–456, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/58/3/448.

Kulekci:2012:FPM

[KVX12] M. Oğuzhan Külekci, Jeffrey Scott Vitter, and Bojian Xu. Fast pattern-matching via k-bit filtering based text decomposition. The Computer Journal, 55(1):62–68, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/55/1/62.full.pdf+html.

Kato:2005:SSR

[KW05] Ryoichi Kato and Osamu Watanabe. Substring search and repeat search using factor oracles. Information Processing Letters, 93(6):269–274, March 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaczmarski:2019:GRT

[KW19] Krzysztof Kaczmarski and Albert Wolant. GPU R-Trie: Dictionary with ultra fast lookup. Concurrency and Computation: Practice and Experience, 31(19):e5027:1–e5027:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2007:GAT

[KWL07] Min-Soo Kim, Kyu-Young Whang, and Jae-Gil Lee. n-Gram/ 2L-approximation: a two-level n-gram inverted index structure for approximate string matching. International Journal of Computer Systems Science and Engineering, 22(6):??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Kim:2008:SOF

[KWLL08] Min-Soo Kim, Kyu-Young Whang, Jae-Gil Lee, and Min-Jae Lee. Structural optimization of a full-text n-gram index using relational normalization. VLDB Journal: Very Large Data Bases, 17(6):1485–1507, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kaplan:2019:RRP

[KYG19] R. Kaplan, L. Yavits, and R. Ginosar. RASSA: Resistive prealignment accelerator for approximate DNA long read mapping. IEEE Micro, 39(4):44–54, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kupferman:2002:IAM

[KZ02] Orna Kupferman and Sharon Zuhovitzky. An improved algorithm for the membership problem for extended regular expressions. Lecture Notes in Computer Science, 2420: 446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.de/

link/service/series/0558/bibs/2420/24200446.htm;
http://link.springer.de/link/service/series/0558/papers/
2420/24200446.pdf.

Lin:2012:AAA

[LA12] Jie Lin and Don Adjeroh. All-against-all circular pattern matching. The Computer Journal, 55(7):897–906, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/55/7/897.full.pdf+html. Special Focus on

the Centenary of Alan Turing.

Labarre:2012:RBP

[Lab12a] Anthony Labarre. Review of Combinatorial Pattern Matching Algorithms in Computational Biology using Perl and R, by Gabriel Valiente. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 43(3):48–50, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Labarre:2012:RCP

[Lab12b] Anthony Labarre. Review of Combinatorial Pattern Matching Algorithms in Computational Biology using Perl and R, by Gabriel Valiente. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 43(3):48–50, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Larsen:1998:REN

[Lar98] Kim S. Larsen. Regular expressions with nested levels of back referencing form a hierarchy. Information Processing Letters, 65(4):169–172, February 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Larsson:1999:SSM

[Lar99] N. Jesper Larsson. Structures of string matching and data compression. Ph.D. thesis, Lunds Universitet, Lund, Sweden,

1999. 130 pp. URL http://search.proquest.com/docview/
304568808.

Laurikari:2000:NTT

[Lau00] V. Laurikari. NFAs with tagged transitions, their conversion to deterministic automata and application to regular expressions. In Seventh International Symposium on String Processing and Information Retrieval, 2000. SPIRE 2000. 27–29 September 2000, A Curuña, Spain. Proceedings, pages 181– 187. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0746-8, 0-7695-0747-6 (case), 0-7695-0748-4 (microfiche). LCCN QA76.9.T48 I59 2000.

Laurikari:2001:ESA

[Lau01] V. Laurikari. Efficient submatch addressing for regular expressions. Master’s thesis, Helsinki University of Technology, Helsinki, Finland, 2001.

Laville:1991:CPR

[Lav91] Alain Laville. Comparison of priority rules in pattern matching and term rewriting. Journal of Symbolic Computation, 11 (4):321–348 (or 321–347??), April 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lladser:2008:MPM

[LBK08] Manuel E. Lladser, M. D. Betterton, and Rob Knight. Multiple pattern matching: a Markov chain approach. Journal of Mathematical Biology, 56(1–2):51–92, January 2008. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s00285-007-0109-3.pdf.

Lu:2006:PFS

[LCL06] Shiyong Lu, Feng Cao, and Yi Lu. PAMA: a fast string matching algorithm. International Journal of Foundations of Computer Science (IJFCS), 17(2):357–378, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Little:2010:OSM

[LD10] G. Little and J. Diamond. Optimum string match choices in LZSS. In Storer and Marcellin [SM10], page 538. ISBN 0-7695-3994-7. ISSN 1068-0314. LCCN ???? URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453493.

Liu:1998:IAS

[LDI98] Z. Liu, X. Du, and N. Ishii. An improved adaptive string searching algorithm. Software — Practice and Experience, 28 (2):191–198, February 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=1775;
http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=1775&PLACEBO=IE.pdf.

LeBret:1991:RSM

[Le 91] Christophe Le Bret. Robust string matching. Computer Language Magazine, 8(12):71–??, December 1991. CODEN COM-LEF. ISSN 0749-2839.

Lecroq:1992:VBM

[Lec92] T. Lecroq. A variation on the Boyer–Moore algorithm. Theoretical Computer Science, 92(1):119–144, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lecroq:1995:ERS

[Lec95] Thierry Lecroq. Experimental results on string matching algorithms. Software — Practice and Experience, 25(7):727–765, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lecroq:1998:ESM

[Lec98] Thierry Lecroq. Experiments on string matching in memory structures. Software — Practice and Experience, 28(5):561– 568, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=1796;   http://www3.
interscience.wiley.com/cgi-bin/fulltext?ID=1796&PLACEBO=
IE.pdf.

Lecroq:2007:FES

[Lec07] Thierry Lecroq. Fast exact string matching algorithms. Information Processing Letters, 102(6):229–235, June 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1982:EAF

[Lee82] Edward T. Lee. An efficient algorithm for finding Kleene closure of regular expression matrices. International Journal of Computer and Information Sciences, 11(6):409–415, December 1982. CODEN IJCIAH. ISSN 0091-7036.

Lee:1991:DCM

[Lee91] David Lee. Detection, classification, and measurement of discontinuities. SIAM Journal on Scientific and Statistical Computing, 12(2):311–341, March 1991. CODEN SIJCD4. ISSN 0196-5204.

Lee:2009:HAH

[Lee09] Tsern-Huei Lee. Hardware architecture for high-performance regular expression matching. IEEE Transactions on Computers, 58(7):984–993, July 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
4599575.

Leiss:1980:CFA

[Lei80] Ernst Leiss. Constructing a finite automaton for a given regular expression. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 12(3):81–87, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Leiss:1981:CRR

[Lei81] Ernst Leiss. The complexity of restricted regular expressions and the synthesis problem for finite automata. Journal of Computer and System Sciences, 23(3):348–354, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0022000081900702.

Leiss:1985:CTU

[Lei85] E. Leiss. On classes of tractable unrestricted regular expressions. Theoretical Computer Science, 35(2-3):313–327, February 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lengauer:1993:AEF

[Len93] T. Lengauer, editor. Algorithms — ESA’93. First Annual European Symposium Proceedings: September 1993, Bad Honnef, Germany, volume 726 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Lenzerini:2011:PPT

[Len11] Maurizio Lenzerini, editor. PODS’11: Proceedings of the thirtieth ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems: June 13–15, 2011, Athens, Greece. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-

0660-8. LCCN ???? URL http://dl.acm.org/citation.
cfm?id=1989284.

Lesk:1979:DTD

[Les79] Arthur M. Lesk. Detection of three-dimensional patterns of atoms in chemical structures. Communications of the Association for Computing Machinery, 22(4):219–224, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lestree:1994:URU

[Les94] L. Lestree. Unit route upper bound for string-matching on hypercube. Lecture Notes in Computer Science, 807:136–145, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lesk:1995:TDP

[Les95] A. M. Lesk. Three-dimensional pattern matching in protein structure analysis. Lecture Notes in Computer Science, 937: 248–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1997:UUM

[Leu97] Vitus J. Leung. The undecidability of the unrestricted modified edit distance. Theoretical Computer Science, 180 (1–2):203–215, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_

sub/browse/browse.cgi?year=1997&volume=180&issue=1-
2&aid=2399.

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montréal, Canada, ISSAC - PROCEEDINGS- 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Li:2016:HMF

[LG16] Zheng Li and Tingjian Ge. History is a mirror to the future: best-effort approximate complex event matching with insufficient resources. Proceedings of the VLDB Endowment, 10(4): 397–408, November 2016. CODEN ???? ISSN 2150-8097.

Londhe:2014:MTC

[LGZ^{+}14] Nikhil Londhe, Vishrawas Gopalakrishnan, Aidong Zhang, Hung Q. Ngo, and Rohini Srihari. Matching titles with cross title web-search enrichment and community detection. Proceedings of the VLDB Endowment, 7(12):1167–1178, August 2014. CODEN ???? ISSN 2150-8097.

Lemstrom:2003:APM

[LH03] Kjell Lemström and Lauri Hella. Approximate pattern matching and transitive closure logics. Theoretical Computer Science, 299(1–3):387–412, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2013:PMS

[LH13a] Tsern-Huei Lee and Nai-Lun Huang. A pattern-matching scheme with high throughput performance and low memory requirement. IEEE/ACM Transactions on Networking, 21(4): 1104–1116, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:OPS

[LH13b] Bin Li and Steven C. H. Hoi. Online portfolio selection: a survey. ACM Computing Surveys, 46(3):35:1–35:??, January 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Lai:1993:AAD

[LHCH93] Feipei Lai, Shu-Lin Hwang, Tzer-Shyong Chen, and Chia-Rung Hsieh. Arden — architecture development environment. In Baozong [Bao93], pages 5–9 (vol.1). ISBN 0-7803-1233-3. LCCN QA75.5.I155 1993. IEEE Catalog No. 93CH3286-2.

Liu:2004:FSM

[LHCK04] Rong-Tai Liu, Nen-Fu Huang, Chih-Hao Chen, and Chia-Nan Kao. A fast string-matching algorithm for network processorbased intrusion detection system. ACM Transactions on Embedded Computing Systems, 3(3):614–633, August 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:1991:VAL

[LHMH91] Kuo Chu Lee, Takako Matoba Hickey, Victor W. Mak, and Gary E. Herman. VLSI accelerators for large database systems. IEEE Micro, 11(6):8–20, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liang:1984:WHP

[Lia84] Franklin Mark Liang. Word Hy-phen-a-tion by Com-pu-ter. Ph.D. dissertation, Computer Science Department, Stanford University, Stanford, CA, USA, March 1984. 92

pp. URL http://www.tug.org/docs/liang/;   http://
wwwlib.umi.com/dissertations/fullcit/8329742.

Lifshits:2003:LBS

[Lif03] Yuri Lifshits. A lower bound on the size of ϵ-free NFA corresponding to a regular expression. Information Processing Letters, 85(6):293–299, March 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1981:SPM

[Liu81] Ken Chih Liu. On string pattern matching: a new model with a polynomial time algorithm. SIAM Journal on Computing, 10(1):118–140, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liu:1986:SPM

[Liu86] Ken-Chih Liu. A string pattern matching extension to Pascal and some comparisons with SNOBOL4. Software — Practice and Experience, 16(6):541–548, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:1988:SPM

[Liu88] Ken-Chih Liu. On string pattern matching: a quantitative analysis and a proposal. Computer Languages, 13(1):23–29, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liu:2014:PRP

[Liu14] Fang Liu. gset: An R package for exact sequential test of equivalence hypothesis based on bivariate non-central tstatistics. The R Journal, 6(2):174–184, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.r-

project.org/archive/2014-2/RJournal_2014-2_liu.pdf.

Lin:2017:IID

[LJH^{+}17] Jie Lin, Yue Jiang, E. James Harner, Bing-Hua Jiang, and Don Adjeroh. IDPM: An improved degenerate pattern matching algorithm for biological sequences. International Journal of Foundations of Computer Science (IJFCS), 28(7):889–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Liu:2013:IAA

[LJZZ13] Nan Liu, Haitao Jiang, Daming Zhu, and Binhai Zhu. An improved approximation algorithm for scaffold filling to maximize the common adjacencies. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(4):905–913, July 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Lee:1988:HSC

[LK88] Peizong Lee and Z. M. Kedem. On high-speed computing with a programmable linear array. In IEEE [IEE88], pages 425–432. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Lee:1990:HSC

[LK90] Peizong Lee and Zvi M. Kedem. On high-speed computing with a programmable linear array. The Journal of Supercomputing, 4(3):223–249, September 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=4&issue=3&spage=223.

Lenka:2006:SML

[LK06] Dillip Kumar Lenka and Pawan Kumar. States merging in LR parser. ACM SIGPLAN Notices, 41(4):24–29, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2002:EPM

[LKL02] Sangjun Lee, Dongseop Kwon, and Sukho Lee. Efficient pattern matching of time series data. Lecture Notes in Computer Science, 2358:586–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2358/
23580586.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2358/23580586.pdf.

LeGlaunec:2023:REM

[LKM23] Alexis Le Glaunec, Lingkun Kong, and Konstantinos Mamouras. Regular expression matching using bit vector automata. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):92:1–92:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL https://dl.acm.

org/doi/10.1145/3586044.

Lenzerini:2008:PTS

[LL08] Maurizio Lenzerini and Domenico Lembo, editors. Proceedings of the Twenty-Seventh ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS’08, Vancouver, BC, Canada, June 9–11, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-60560-932-3. LCCN ????

Lee:2003:HOO

[LLC03] Keunwoo Lee, Anthony LaMarca, and Craig Chambers. HydroJ: object-oriented pattern matching for evolvable distributed systems. ACM SIGPLAN Notices, 38(11):205–223, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2017:LBH

[LLC17] Yi-Shan Lin, Chun-Liang Lee, and Yaw-Chung Chen. Lengthbounded hybrid CPU/GPU pattern matching algorithm for deep packet inspection. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/1/16.

Lin:2013:APM

[LLCC13] Cheng-Hung Lin, Chen-Hsiung Liu, Lung-Sheng Chien, and Shih-Chieh Chang. Accelerating pattern matching using a novel parallel algorithm on GPUs. IEEE Transactions on Computers, 62(10):1906–1916, October 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

LeBlond:2012:CPB

[LLL12] Stevens Le Blond, Fabrice Le Fessant, and Erwan Le Merrer. Choosing partners based on availability in P2P networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):25:1–25:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lu:2013:NFM

[LLL13] Chia Wei Lu, Chin Lung Lu, and R. C. T. Lee. A new filtration method and a hybrid strategy for approximate string matching. Theoretical Computer Science, 481(??):9–17, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397513001540.

Lin:2024:PAR

[LLL^{+}24] Jing Lin, Weiwei Lin, Hang Lin, Longlong Zhu, Dong Zhang, and Chunming Wu. P4Rex: Accelerating regular expression matching with programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128624004948.

Lin:2017:PHB

[LLLC17] Cheng-Hung Lin, Jin-Cheng Li, Chen-Hsiung Liu, and Shih-Chieh Chang. Perfect hashing based parallel algorithms for multiple string matching on graphic processing units. IEEE Transactions on Parallel and Distributed Systems, 28(9):2639–2650, September 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/09/07864442-abs.html.

Lin:2008:USM

[LLLL08] Po-Ching Lin, Ying-Dar Lin, Yuan-Cheng Lai, and Tsern-Huei Lee. Using string matching for deep packet inspection. Computer, 41(4):23–28, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2012:MQG

[LLS12] Dan Li, Xiaoshan Li, and Volker Stolz. Model querying with graphical notation of QVT relations. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:2020:YCA

[LLS^{+}20] D. J. Lee, J. Lee, T. Siddiqui, J. Kim, K. Karahalios, and A. Parameswaran. You can’t always sketch what you want: Understanding sensemaking in visual query systems. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1267–1277, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2015:BQA

[LLW^{+}15] Jiaheng Lu, Chunbin Lin, Wei Wang, Chen Li, and Xiaokui Xiao. Boosting the quality of approximate string matching by synonyms. ACM Transactions on Database Systems, 40 (3):15:1–15:??, October 2015. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

LeFessant:2001:OPM

[LM01a] Fabrice Le Fessant and Luc Maranget. Optimizing pattern matching. ACM SIGPLAN Notices, 36(10):26–37, October 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://cristal.

inria.fr/ICFP2001/Abstracts/8.html.

Li:2001:IQX

[LM01b] Quanzhong Li and Bongki Moon. Indexing and querying XML data for regular path expressions. In Apers et al. [AAC^{+}01], pages 361–370. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/conf/2001/P361.pdf.

Liu:2002:JIA

[LM02] J. Liu and A. C. Myers. JMatch: Iterable abstract pattern matching for Java. Lecture Notes in Computer Science, 2562: 110–127, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Losemann:2012:CEP

[LM12] Katja Losemann and Wim Martens. The complexity of evaluating path expressions in SPARQL. In Krötzsch et al. [KLB12], pages 101–112. ISBN ???? LCCN ???? URL

http://www.sigmod.org/2012/.

Losemann:2013:CRE

[LM13] Katja Losemann and Wim Martens. The complexity of regular expressions and property paths in SPARQL. ACM Transactions on Database Systems, 38(4):24:1–24:??, November 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lancia:2017:SSS

[LMM17] Giuseppe Lancia, Luke Mathieson, and Pablo Moscato. Separating sets of strings by finding matching patterns is almost always hard. Theoretical Computer Science, 665(??):73–86, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397516307460.

Leonardi:2007:OSR

[LMMN07] Emilio Leonardi, Marco Mellia, Marco Ajmone Marsan, and Fabio Neri. Optimal scheduling and routing for maximum network throughput. IEEE/ACM Transactions on Networking, 15(6):1541–1554, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Losemann:2016:CPD

[LMN16] Katja Losemann, Wim Martens, and Matthias Niewerth. Closure properties and descriptional complexity of deterministic regular expressions. Theoretical Computer Science, 627(??):54–70, May 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397516001584.

Lewenstein:2016:DRO

[LMNT16] Moshe Lewenstein, J. Ian Munro, Yakov Nekrich, and Sharma V. Thankachan. Document retrieval with one wildcard. Theoretical Computer Science, 635(??):94–101, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397516301608.

Lewenstein:2014:LSI

[LMRT14] Moshe Lewenstein, J. Ian Munro, Venkatesh Raman, and Sharma V. Thankachan. Less space: Indexing for queries with wildcards. Theoretical Computer Science, 557(??):120–127, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397514006562.

Louza:2021:NAR

[LMS21] Felipe A. Louza, Neerja Mhaskar, and W. F. Smyth. A new approach to regular & indeterminate strings. Theoretical Computer Science, 854(??):105–115, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397520307118.

Liu:2016:PCU

[LMT16] Alex X. Liu, Chad R. Meiners, and Eric Torng. Packet classification using binary content addressable memory. IEEE/ ACM Transactions on Networking, 24(3):1295–1307, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Libkin:2016:QGD

[LMV16] Leonid Libkin, Wim Martens, and Domagoj Vrgoc. Querying graphs with data. Journal of the Association for Computing Machinery, 63(2):14:1–14:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:2019:CAR

[LN19] Alex X. Liu and Eric Norige. A de-compositional approach to regular expression matching for network security. IEEE/ ACM Transactions on Networking, 27(6):2179–2191, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/abs/10.
1109/TNET.2019.2941920.

Lopez-Ortiz:1994:LPM

[LO94] Alejandro López-Ortiz. Linear pattern matching of repeated substrings. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 25(3):114– 121, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lohrey:2010:CMP

[Loh10] Markus Lohrey. Compressed membership problems for regular expressions and hierarchical automata. International Journal of Foundations of Computer Science (IJFCS), 21(5):817–841, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liang:2024:ERS

[LOZ^{+}24] Qi Liang, Dian Ouyang, Fan Zhang, Jianye Yang, Xuemin Lin, and Zhihong Tian. Efficient regular simple path queries under transitive restricted expressions. Proceedings of the VLDB Endowment, 17(7):1710–1722, March 2024. CODEN

???? ISSN 2150-8097. URL https://dl.acm.org/doi/10.
14778/3654621.3654636.

Lipsky:2008:PML

[LP08] Ohad Lipsky and Ely Porat. L_{1} pattern matching lower bound. Information Processing Letters, 105(4):141–143, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipsky:2011:APM

[LP11] Ohad Lipsky and Ely Porat. Approximate pattern matching with the L_{1}, L_{2} and L_{∞} metrics. Algorithmica, 60(2):335–348, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=60&
issue=2&spage=335.

Le:2013:MEM

[LP13] Hoang Le and Viktor K. Prasanna. A memory-efficient and modular approach for large-scale string pattern matching. IEEE Transactions on Computers, 62(5):844–857, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:AAP

[LPJ23] Hongyuan Liu, Sreepathi Pai, and Adwait Jog. Asynchronous automata processing on GPUs. ACM SIGMETRICS Performance Evaluation Review, 51(1):23–24, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL

https://dl.acm.org/doi/10.1145/3606376.3593524.

Lozano:2008:STA

[LPR^{+}08] Antoni Lozano, Ron Y. Pinter, Oleg Rokhlenko, Gabriel Valiente, and Michal Ziv-Ukelson. Seeded tree alignment. IEEE/ ACM Transactions on Computational Biology and Bioinformatics, 5(4):503–513, October 2008. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Li:2012:WHT

[LPT12] Yinan Li, Jignesh M. Patel, and Allison Terrell. WHAM: a high-throughput sequence alignment method. ACM Transactions on Database Systems, 37(4):28:1–28:??, December 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Lai:2016:SDS

[LQL^{+}16] Longbin Lai, Lu Qin, Xuemin Lin, Ying Zhang, Lijun Chang, and Shiyu Yang. Scalable distributed subgraph enumeration. Proceedings of the VLDB Endowment, 10(3):217–228, November 2016. CODEN ???? ISSN 2150-8097.

Lopez:2014:MPR

[LR14] Félix López and Víctor Romero. Mastering Python regular expressions: leverage regular expressions in Python even for the most complex features. Community experience distilled. Packt Pub., Birmingham, UK, 2014. ISBN 1-78328-315-7 (paperback), 1-78328-316-5 (e-book). 110 pp. LCCN QA76.73.P98

L67 2014. URL http://alltitles.ebrary.com/Doc?id=
10842105;   http://proquest.safaribooksonline.com/

?fpi=9781783283156;   http://proquest.tech.safaribooksonline.
de/9781783283156.

Libkin:2018:TNA

[LRSV18] Leonid Libkin, Juan L. Reutter, Adrián Soto, and Domagoj Vrgoc. TriAL: a navigational algebra for RDF triplestores. ACM Transactions on Database Systems, 43(1):5:1– 5:??, April 2018. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Libkin:2013:TRA

[LRV13] Leonid Libkin, Juan Reutter, and Domagoj Vrgoc. Trial for RDF: adapting graph query languages for RDF data. In Hull and Fan [HF13], pages 201–212. ISBN 1-4503-2066-X, 1-4503-

2037-6. LCCN ???? URL http://dl.acm.org/citation.
cfm?id=2463664;   http://www.sigmod.org/2013/.

Lesk:1979:LLAa

[LS79] Michael E. Lesk and E. Schmidt. Lex — a lexical analyzer generator. Technical Memorandum 1061, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. 13?? pp.

Luczak:1994:LDC

[LS94] T. Luczak and W. Szpankowski. A lossy data compression based on string matching: Preliminary analysis and suboptimal algorithms. Lecture Notes in Computer Science, 807: 102–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laird:1999:REN

[LS99] Cameron Laird and Kathryn Soraiz. Regular expressions: New choices for scripting: Language designers offer novel possibilities for scripting developers. Sunworld Online, 5 (2):??, February 1999. ISSN 1091-8914. URL http://www.

sunworld.com/swol-02-1999/swol-02-regex.html?0202a.

Discusses new scripting languages: Ficl, FIJI, Rebol, Ruby and WebL. Ficl is a descendant of Forth. FIJI is real Forth, but with objects that are Java objects, and with full access to Java.

Laird:2006:RER

[LS06] Cameron Laird and Kathryn Soraiz. Regular expressions: Rexx still going strong. UNIX Review, ??(??):

??, January 2006. CODEN UNRED5. ISSN 0742-

3136. URL http://www.unixreview.com/documents/s=
9953/ur0601h/ur0601h.html.

Linhart:2009:FPM

[LS09] Chaim Linhart and Ron Shamir. Faster pattern matching with character classes using prime number encoding. Journal of Computer and System Sciences, 75(3):155–162, May 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-

2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022000008000755.

Libkin:2010:DPM

[LS10] Leonid Libkin and Cristina Sirangelo. Disjoint pattern matching and implication in strings. Information Processing Letters, 110(4):143–147, January 16, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2017:SRE

[LSO17] Mina Lee, Sunbeom So, and Hakjoo Oh. Synthesizing regular expressions from examples for introductory automata assignments. ACM SIGPLAN Notices, 52(3):70–80, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2017:DPD

[LSTW^{+}17] En-Shiun Annie Lee, Ho-Yin Antonio Sze-To, Man-Hon Wong, Kwong-Sak Leung, Terrence Chi-Kong Lau, and Andrew K. C. Wong. Discovering protein-DNA binding cores by aligned pattern clustering. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 14(2):254–263, March 2017. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Litvak:2008:PRB

[LSV08] Nelly Litvak, Werner R. W. Scheinhardt, and Yana Volkovich. Probabilistic relation between in-degree and PageRank. In Aiello et al. [A^{+}08], pages 72–83. ISBN 3-540-78808-5, 3-540-78807-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2006.

Lam:2008:IAS

[LSW08] Tak-Wah Lam, Wing-Kin Sung, and Swee-Seong Wong. Improved approximate string matching using compressed suffix data structures. Algorithmica, 51(3):298–314, July 2008. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=51&issue=3&
spage=298.

Lin:2019:DPF

[LSWP19] Peng Lin, Qi Song, Yinghui Wu, and Jiaxing Pi. Discovering patterns for fact checking in knowledge graphs. Journal of Data and Information Quality (JDIQ), 11(3):13:1–13:??, July 2019. CODEN ???? ISSN 1936-1955. URL https://dl.acm.

org/ft_gateway.cfm?id=3286488.

Lins:1990:ISU

[LT90a] Rafael D. Lins and Simon J. Thompson. Implementing SASL using categorical multi-combinators. Software — Practice and Experience, 20(11):1137–1165, November 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:1990:COR

[LT90b] Cheng-Hsiung Liu and Wen-Hsiang Tsai. 3D curved object recognition from multiple 2D camera views. Computer Vision, Graphics, and Image Processing, 50(2):177–187, May 1990. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Lopresti:1997:BEM

[LT97] Daniel Lopresti and Andrew Tomkins. Block edit models for approximate string matching. Theoretical Computer Science, 181(1):159–179, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1997&volume=181&issue=1&
aid=2470.

Lemstrom:2003:TIP

[LT03] K. Lemström and J. Tarhio. Transposition invariant pattern matching for multi-track strings. Nordic Journal of Computing, 10(3):185–??, Fall 2003. CODEN NJCOFR. ISSN 1236-

Lo:2009:SOC

[LT09] Chia-Tien Dan Lo and Yi-Gang Tai. Space optimization on counters for FPGA-based Perl compatible regular expressions. ACM Transactions on Reconfigurable Technology and Systems, 2(4):23:1–23:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Liu:2016:OAA

[LT16] Alex X. Liu and Eric Torng. Overlay automata and algorithms for fast and scalable regular expression matching. IEEE/ACM Transactions on Networking, 24(4):2400–2415, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2004:ELP

[LTL04] Ting-Yu Lin, Yu-Chee Tseng, and Yuan-Ting Lu. An efficient link polling policy by pattern matching for Bluetooth Piconets. The Computer Journal, 47(2):169–178, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/470169.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_47/Issue_
02/pdf/470169.pdf.

Libkin:2015:RED

[LTV15] Leonid Libkin, Tony Tan, and Domagoj Vrgoc. Regular expressions for data words. Journal of Computer and System Sciences, 81(7):1278–1297, November 2015. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000015000276.

Luderer:1977:CPT

[Lud77] G. W. R. Luderer. Cut and paste — two UNIX word processing commands. Technical Memorandum 1205 (MF 77-8234-086), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1977. ?? pp.

Lustman:1994:STB

[Lus94] F. Lustman. Specifying transaction-based information systems with regular expressions. IEEE Transactions on Software Engineering, 20(3):207–217, March 1994. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=268922.

Lutz:2002:BRB

[Lut02] Michael J. Lutz. Bookshelf: Risk-based approach to mission-critical software testing; the TLA+ specification language primer [Specifying Systems: The TLA+ Language and Tools for Hardware and Software Engineers]; interpolation tutorial [Nonuniform Sampling: Theory and Practice]; pattern-matching toolkit [Flexible Pattern Matching in Strings: Practical On-Line Search Algorithms for Texts and Biological Sequences]. Computer, 35(9):81, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://csdl.computer.org/dl/
mags/co/2002/09/r9081.htm;   http://csdl.computer.
org/dl/mags/co/2002/09/r9081.pdf.

Landau:1986:ESM

[LV86a] G. M. Landau and U. Vishkin. Efficient string matching with k mismatches. Theoretical Computer Science, 43(2-3):239– 249, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Landau:1986:IEP

[LV86b] G. M. Landau and U. Vishkin. Introducing efficient parallelism into approximate string matching and a new serial algorithm. In ACM [ACM86], pages 220–230. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986. URL http://www.

acm.org/pubs/articles/proceedings/stoc/12130/p220-
landau/p220-landau.pdf;   http://www.acm.org/pubs/
citations/proceedings/stoc/12130/p220-landau/. ACM

order no. 508860.

Landau:1988:FSM

[LV88] Gad M. Landau and Uzi Vishkin. Fast string matching with k differences. Journal of Computer and System Sciences, 37(1):63–78, August 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0022000088900451.

Landau:1989:FPS

[LV89] Gad M. Landau and Uzi Vishkin. Fast parallel and serial approximate string matching. Journal of Algorithms, 10(2):157– 169, June 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0196677489900102.

Landau:1994:PMD

[LV94] Gad M. Landau and Uzi Vishkin. Pattern matching in a digitized image. Algorithmica, 12(4–5):375–408, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lewenstein:2006:CPM

[LV06] Moshe Lewenstein and Gabriel Valiente, editors. Combinatorial Pattern Matching: 17th Annual Symposium, CPM 2006, Barcelona, Spain, July 5–7, 2006. Proceedings, volume 4009 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-35455-7 (print), 3-540-35461-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-540-35461-1.

Landau:1987:ESM

[LVN87] Gad M. Landau, Uzi Vishkin, and Ruth Nussinov. An efficient string matching algorithm with K substitutions for nucleotide and amino acid sequences. Journal of Theoretical Biology, 126(4):483–490, June 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801534.

Li:2016:RDT

[LWS^{+}16] Zeyu Li, Hongzhi Wang, Wei Shao, Jianzhong Li, and Hong Gao. Repairing data through regular expressions. Proceedings of the VLDB Endowment, 9(5):432–443, January 2016. CODEN ???? ISSN 2150-8097.

Li:1986:SMC

[LY86] Ming Li and Yaacov Yesha. String-matching cannot be done by a two-head one-way deterministic finite automaton. Information Processing Letters, 22(5):231–235, April ??, 1986.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2017:FPM

[LY17] Chun-Liang Lee and Tzu-Hao Yang. A flexible patternmatching algorithm for network intrusion detection systems using multi-core processors. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/2/58.

Li:2023:SSG

[LYT^{+}23] Lei Li, Mengjiao Yan, Zhenchao Tao, Huanhuan Chen, and Xindong Wu. Semi-supervised graph pattern matching and rematching for expert community location. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):6:1– 6:??, January 2023. CODEN ???? ISSN 1556-4681 (print),

1556-472X (electronic). URL https://dl.acm.org/doi/10.
1145/3532623.

Leonard:2008:SDP

[LYWL08] Derek Leonard, Zhongmei Yao, Xiaoming Wang, and Dmitri Loguinov. On static and dynamic partitioning behavior of large-scale P2P networks. IEEE/ACM Transactions on Networking, 16(6):1475–1488, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lucarella:1996:VRE

[LZ96] Dario Lucarella and Antonella Zanzi. A visual retrieval environment for hypermedia information systems. ACM Transactions on Information Systems, 14(1):3–29, January 1996. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.

org:80/tois/abstracts/lucarella.html.

Li:2018:FWI

[LZ18] Youyuan Li and Yingping Zhuang. fmpRPMF: a Web implementation for protein identification by robust peptide mass fingerprinting. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(5):1728–1731, September 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Li:1998:HRE

[LZHZ98] X. Li, T. Zheng, J. Hou, and J. Zhao. Hybrid regular expressions. Lecture Notes in Computer Science, 1386:384–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moraru:2012:EPM

[MA12] Iulian Moraru and David G. Andersen. Exact pattern matching with feed-forward Bloom filters. ACM Journal of Experimental Algorithmics, 17(??):3.4:1–3.4:??, 2012. CODEN ???? ISSN 1084-6654.

Maass:2006:MSE

[Maa06] Moritz G. Maaß. Matching statistics: efficient computation and a new practical algorithm for the multiple common substring problem. Software — Practice and Experience, 36(3): 305–331, March 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ma:2014:TAC

[MAC14] Lin Ma, Kunal Agrawal, and Roger D. Chamberlain. Theoretical analysis of classic algorithms on highly-threaded manycore GPUs. ACM SIGPLAN Notices, 49(8):391–392, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maddock:2001:REC

[Mad01] John Maddock. Regular expressions in C++. Dr. Dobb’s Journal of Software Tools, 26(10):21–22, 24, 26, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/2001/2001_10/regexpp3.txt;   http:
//www.ddj.com/ftp/2001/2001_10/regexpp3.zip.

Maes:1990:CSS

[Mae90] Maurice Maes. On a cyclic string-to-string correction problem. Information Processing Letters, 35(2):73–78, June 29, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maeder:1994:MPL

[Mae94] Roman E. Maeder. The Mathematica programmer: Logic programming I: The interpreter. Mathematica Journal, 4 (1):53–63, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http://www.

mathematica-journal.com/issue/v4i1/columns/maeder/
53-63_Roman41.mj.pdf;   http://www.mathematica-journal.
com/issue/v4i1/columns/maeder/index.html.

Magel:1981:REP

[Mag81] Kenneth Magel. Regular expressions in a program complexity metric. ACM SIGPLAN Notices, 16(7):61–65, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maher:2007:MPU

[Mah07] Tim Maher. Minimal Perl: for UNIX and Linux people. Manning Publications, Greenwich, CT, USA, 2007. ISBN 1-932394-50-8 (paperback). xxxviii + 450 pp. LCCN QA76.73.P22 M34 2007.

Matsuoka:2016:GPM

[MAI^{+}16] Yoshiaki Matsuoka, Takahiro Aoki, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Generalized pattern matching and periodicity under substring consistent equivalence relations. Theoretical Computer Science, 656 (Part B)(??):225–233, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397516001390.

Makinen:1989:SIP

[Mäk89] Erkki Mäkinen. On the subtree isomorphism problem for ordered trees. Information Processing Letters, 32(5):271–273, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Gro91a].

Malton:1993:DSF

[Mal93] Andrew Malton. The denotational semantics of a functional tree-manipulation language. Computer Languages, 19(3):157– 168, July 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Manacher:1975:NLT

[Man75] Glenn Manacher. A new linear-time “On-Line” algorithm for finding the smallest initial palindrome of a string. Journal of the Association for Computing Machinery, 22(3):346–351, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manacher:1976:APM

[Man76] G. K. Manacher. An application of pattern matching to a problem in geometrical complexity. Information Processing Letters, 5(1):6–7, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manolopoulos:1986:BSI

[Man86] Yannis Manolopoulos. Batched search of index sequential files. Information Processing Letters, 22(5):267–272, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manber:1994:TCS

[Man94] U. Manber. A text compression scheme that allows fast searching directly in the compressed file. Lecture Notes in Computer Science, 807:113–124, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manber:1997:TCS

[Man97] Udi Manber. A text compression scheme that allows fast searching directly in the compressed file. ACM Transactions on Information Systems, 15(2):124–136, April 1997. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.org:80/

tois/abstracts/manber.html.

Mann:2006:TBG

[Man06] Paul B. Mann. A translational BNF grammar notation (TBNF). ACM SIGPLAN Notices, 41(4):16–23, April 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marola:1989:USD

[Mar89] Giovanni Marola. Using symmetry for detecting and locating objects in a picture. Computer Vision, Graphics, and Image Processing, 46(2):179–195, May 1989. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Maranget:2007:WPM

[Mar07] Luc Maranget. Warnings for pattern matching. Journal of Functional Programming, 17(3):387–421, May 2007. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
3165B75113781E2431E3856972940347.

Matos:1994:PSI

[Mat94] Armando B. Matos. Periodic sets of integers. Theoretical Computer Science, 127(2):287–312, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1479. See remark [Pet95].

Moia:2020:IEC

[MBH20] Vitor Hugo Galhardo Moia, Frank Breitinger, and Marco Aurélio Amaral Henriques. The impact of excluding common blocks for approximate matching. Computers & Security, 89(??):Article 101676, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302159.

Mainardi:2022:PAC

[MBP22] Nicholas Mainardi, Alessandro Barenghi, and Gerardo Pelosi. Privacy-aware character pattern matching over outsourced encrypted data. Digital Threats: Research and Practice (DTRAP), 3(1):7:1–7:38, March 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https://dl.

acm.org/doi/10.1145/3462333.

Manber:1991:ASM

[MBY91] Udi Manber and Ricardo Baeza-Yates. An algorithm for string matching with a sequence of don’t cares. Information Processing Letters, 37(3):133–136, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mamouras:2024:EMR

[MC24] Konstantinos Mamouras and Agnishom Chattopadhyay. Efficient matching of regular expressions with lookaround assertions. Proceedings of the ACM on Programming Languages (PACMPL), 8(POPL):92:1–92:??, January 2024. CODEN ???? ISSN 2475-1421 (electronic). URL https://dl.acm.

org/doi/10.1145/3632934.

McCreight:1976:SES

[McC76] Edward M. McCreight. A space-economical suffix tree construction algorithm. Journal of the Association for Computing Machinery, 23(2):262–272, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McCarty:2001:LN

[McC01] Ron McCarty. A look at ngrep. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):75–76, May 2001. CO-DEN SYADE7. ISSN 1061-2688. URL http://www.samag.

com/.

Ma:2011:CTG

[MCF^{+}11] Shuai Ma, Yang Cao, Wenfei Fan, Jinpeng Huai, and Tianyu Wo. Capturing topology in graph pattern matching. Proceedings of the VLDB Endowment, 5(4):310–321, December 2011. CODEN ???? ISSN 2150-8097.

Ma:2014:SSC

[MCF^{+}14] Shuai Ma, Yang Cao, Wenfei Fan, Jinpeng Huai, and Tianyu Wo. Strong simulation: Capturing topology in graph pattern matching. ACM Transactions on Database Systems, 39(1): 4:1–4:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

McIsaac:1985:PMA

[McI85a] Kevin McIsaac. Pattern matching algebraic identities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(2):4–13, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McIsaac:85

[McI85b] Kevin McIsaac. Pattern matching algebraic identities. SIGSAM Bulletin, 19(2):4–13, May 1985.

McIlroy:2004:ESR

[McI04] M. Douglas McIlroy. Enumerating the strings of regular languages. Journal of Functional Programming, 14 (5):503–518, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL http:

//www.cs.dartmouth.edu/~doug/nfa.ps.gz;   https://
www.cambridge.org/core/product/1D46239B6CC6299AA385B3094EBC80E1.

Moreira:2017:FCR

[MCP17] Rubens E. A. Moreira, Sylvain Collange, and Fernando Magno Quintão Pereira. Function call re-vectorization. ACM SIGPLAN Notices, 52(8):313–326, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Might:2010:YD

[MD10] Matthew Might and David Darais. Yacc is dead. arxiv.org, ??(??):1–18, October 24, 2010. URL http://arxiv.org/abs/

1010.5023;   http://www.ucombinator.org/projects/parsing/

.

Mouchaweh:2002:RLR

[MDLB02] Moamar Sayed Mouchaweh, Arnaud Devillez, Gerard Villermain Lecolier, and Patrice Billaudel. Recursive learning in real time using fuzzy pattern matching. Mathematics and Computers in Simulation, 60(3–5):209–216, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0378475402000150.

Margaritis:1997:VPA

[ME97] K. G. Margaritis and D. J. Evans. A VLSI processor array for flexible string matching. Parallel Algorithms and Applications, 11(1–2):45–60, May 1997. CODEN PAAPEC. ISSN 1063-

7192. URL http://www.informaworld.com/smpp/content~
content=a772742761.

Medvedev:2023:RTA

[Med23] Paul Medvedev. Research: Theoretical analysis of edit distance algorithms. Communications of the Association for Computing Machinery, 66(12):64–71, December 2023. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3582490.

Meir:2008:CCL

[Mei08] Or Meir. Combinatorial construction of locally testable codes. In ACM [ACM08], pages 285–294. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Meister:2015:USD

[Mei15] Daniel Meister. Using swaps and deletes to make strings match. Theoretical Computer Science, 562(??):606–620, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397514008457.

Melichar:1995:ASM

[Mel95] B. Melichar. Approximate string matching by finite automata. Lecture Notes in Computer Science, 970:342–349, 1995. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Menico:1989:FSS

[Men89] Costas Menico. Faster string searches (Boyer–Moore algorithm). Dr. Dobb’s Journal of Software Tools, 14(7):74–75, 98–99, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Meyer:1985:ISM

[Mey85] Bertrand Meyer. Incremental string matching. Information Processing Letters, 21(5):219–227, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marx:1996:PWA

[MF96] Delia I. S. Marx and Phyllis G. Frankl. The path-wise approach to data flow testing with pointer variables. ACM SIG-SOFT Software Engineering Notes, 21(3):135–146, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Manilov:2015:FRT

[MFMA15] Stanislav Manilov, Björn Franke, Anthony Magrath, and Cedric Andrieu. Free rider: a tool for retargeting platformspecific intrinsic functions. ACM SIGPLAN Notices, 50(5): 5:1–5:??, May 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manilov:2017:FRS

[MFMA17] Stanislav Manilov, Björn Franke, Anthony Magrath, and Cedric Andrieu. Free Rider: a source-level transformation tool for retargeting platform-specific intrinsic functions. ACM Transactions on Embedded Computing Systems, 16(2):38:1– 38:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Millstein:2009:EMP

[MFRW09] Todd Millstein, Christopher Frost, Jason Ryder, and Alessandro Warth. Expressive and modular predicate dispatch for Java. ACM Transactions on Programming Languages and Systems, 31(2):7:1–7:54, February 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Muller:1994:ICS

[MG94] Hausi A. Muller and Mari Georges, editors. International Conference on Software Maintenance: proceedings, September 19–23, 1994, Victoria, British Columbia, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6330-8. LCCN QA76.76.S64I58 1994.

Morris:1997:GRB

[MGF97] Paul H. Morris, Ronald A. Gray, and Robert E. Filman. GOTO removal based on regular expressions. Journal of Software Maintenance: Research and Practice, 9(1):47–66, January 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Matzen:1993:MSA

[MGH93] R. W. Matzen, K. M. George, and G. E. Hedrick. A model for studying ambiguity in SGML element declarations. In Deaton et al. [DGBH93], pages 668–676. ISBN 0-89791-567-4. LCCN QA76.76.A65 S95 1993.

Matzen:1997:FLM

[MGH97] R. W. Matzen, K. M. George, and G. E. Hedrick. A formal language model for parsing SGML. The Journal of Systems and Software, 36(2):147–166, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manole:2014:PSP

[MGW14] Sagi Manole, Amit Golander, and Shlomo Weiss. Protein sequence pattern matching: Leveraging application specific hardware accelerators. IEEE Transactions on Computers, 63 (2):448–460, February 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mhashi:2005:EMR

[Mha05] Mahmoud Moh’d Mhashi. The effect of multiple reference characters on detecting matches in string-searching algorithms. Software — Practice and Experience, 35(13):1299– 1315, November 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marschall:2012:PAA

[MHKR12] Tobias Marschall, Inke Herms, Hans-Michael Kaltenbach, and Sven Rahmann. Probabilistic arithmetic automata and their applications. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(6):1737–1750, November 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Mitarai:2001:CPM

[MHM^{+}01] S. Mitarai, M. Hirao, T. Matsumoto, A. Shinohara, M. Takeda, and S. Arikawa. Compressed pattern matching for SEQUITUR. In Storer and Cohn [SC01], pages 469–478. ISBN 0-7695-1031-0. ISSN 1068-0314. LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917178. IEEE CSP number 01PR1031.

Matsumoto:2009:RTE

[MHT09] Tetsuya Matsumoto, Kazuhito Hagio, and Masayuki Takeda. A run-time efficient implementation of compressed pattern matching automata. International Journal of Foundations of Computer Science (IJFCS), 20(4):717–733, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mori:2007:PID

[MI07] Yuki Mori and Takeo Igarashi. Plushie: an interactive design system for plush toys. ACM Transactions on Graphics, 26(3): 45:1–45:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Middendorf:1996:TDP

[Mid96] Martin Middendorf. Two-dimensional partitioning problems. Theoretical Computer Science, 164(1–2):73–106, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.elsevier.com/cgi-
bin/cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2126.

Middendorf:1998:SCS

[Mid98] Martin Middendorf. Shortest common superstrings and scheduling with coordinated starting times. Theoretical Computer Science, 191(1–2):205–214, January 30, 1998. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
tcs/sub/1998/191/1-2/2669.pdf.

Mitani:2017:PEA

[MIH17] Yasuaki Mitani, Fumihiko Ino, and Kenichi Hagihara. Parallelizing exact and approximate string matching via inclusive scan on a GPU. IEEE Transactions on Parallel and Distributed Systems, 28(7):1989–2002, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/07/07797444-abs.html.

Mischel:1989:WAE

[Mis89a] James Mischel. Writing AWK-like extensions to C. Dr. Dobb’s Journal of Software Tools, 14(6):64–66, 68, 70, 94, 96, 98, 103– 104, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Mischel:1989:WAL

[Mis89b] James Mischel. Writing AWK-like extensions to C. Dr. Dobb’s Journal of Software Tools, 14(6):64–66, 68, 70, 94, 96, 98, 103– 104, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Misra:2003:DPS

[Mis03] Jayadev Misra. Derivation of a parallel string matching algorithm. Information Processing Letters, 85(5):255–260, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahajan:1990:EPI

[MK90] Milind Mahajan and V. K. Prasanna Kumar. Efficient parallel implementation of RETE pattern matching. Computer Systems Science and Engineering, 5(3):187–192, July 1990. CODEN CSSEEI. ISSN 0267-6192.

Mak:1991:EPP

[MKF91] Victor Wing-Kit Mak, Chu Lee Kuo, and Ophir Frieder. Exploiting parallelism in pattern matching: An information retrieval application. ACM Transactions on Information Systems, 9(1):52–74, January 1991. CODEN ATISET. ISSN

1046-8188. URL http://www.acm.org:80.

Mochizuki:1998:SSA

[MKSiA98] H. Mochizuki, M. Koyama, M. Shishibori, and J. i. Aoe. A substring search algorithm in extendible hashing. Information Sciences, 108(1):13–30, July 1998. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Mossouni:1996:CSM

[ML96a] F. Mossouni and C. Lavault. N-cube string matching algorithm with long texts. Lecture Notes in Computer Science, 1120:328–340, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moussouni:1996:DSM

[ML96b] F. Moussouni and C. Lavault. Distributed string matching algorithm on the N-cube. Lecture Notes in Computer Science, 1123:832–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moscola:2008:RCB

[MLC08] James Moscola, John W. Lockwood, and Young H. Cho. Reconfigurable content-based router using hardware-accelerated language parser. ACM Transactions on Design Automation of Electronic Systems (TODAES), 13(2):28:1–28:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Murphy:2008:DGB

[MLM^{+}08] Laurie Murphy, Gary Lewandowski, Renée McCauley, Beth Simon, Lynda Thomas, and Carol Zander. Debugging: the good, the bad, and the quirky — a qualitative analysis of novices’ strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):163–167, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Myers:1989:AMR

[MM89] Eugene W. Myers and Webb Miller. Approximate matching of regular expressions. Bulletin of Mathematical Biology, 51(1):5–37, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02458834.

Manber:1993:SAN

[MM93] Udi Manber and Gene Myers. Suffix arrays: a new method for on-line string searches. SIAM Journal on Computing, 22(5): 935–948, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Muth:1996:AMS

[MM96] R. Muth and U. Manber. Approximate multiple string search. Lecture Notes in Computer Science, 1075:75–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michailidis:2002:PSL

[MM02] Panagiotis D. Michailidis and Konstantinos G. Margaritis. A performance study of load balancing strategies for approximate string matching on an MPI heterogeneous system environment. Lecture Notes in Computer Science, 2474: 432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.de/

link/service/series/0558/bibs/2474/24740432.htm;
http://link.springer.de/link/service/series/0558/papers/
2474/24740432.pdf.

Michailidis:2003:PEL

[MM03] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Performance evaluation of load balancing strategies for approximate string matching application on an MPI cluster of heterogeneous workstations. Future Generation Computer Systems, 19(7):1075–1104, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Michailidis:2007:PAP

[MM07] Panagiotis D. Michailidis and Konstantinos G. Margaritis. A programmable array processor architecture for flexible approximate string matching algorithms. Journal of Parallel and Distributed Computing, 67(2):131–141, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mateescu:2011:CEC

[MMDdJ11] Radu Mateescu, Pedro T. Monteiro, Estelle Dumas, and Hidde de Jong. CTRL: Extension of CTL with regular expressions and fairness operators to verify genetic regulatory networks. Theoretical Computer Science, 412(26):2854–2883, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2014:TCR

[MME14] Marko Martin, Mira Mezini, and Sebastian Erdweg. Template constructors for reusable object initialization. ACM SIGPLAN Notices, 49(3):43–52, March 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Medeiros:2014:RPE

[MMI14] Sérgio Medeiros, Fabio Mascarenhas, and Roberto Ierusalimschy. From regexes to parsing expression grammars. Science of Computer Programming, 93 (part A)(??):3–18, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-

7964 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167642312002171.

Mytkowicz:2014:DPF

[MMS14] Todd Mytkowicz, Madanlal Musuvathi, and Wolfram Schulte. Data-parallel finite-state machines. ACM SIGARCH Computer Architecture News, 42(1):529–542, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mandreoli:2010:PHS

[MMZ10] Federica Mandreoli, Riccardo Martoglia, and Pavel Zezula. Principles of Holism for sequential twig pattern matching. VLDB Journal: Very Large Data Bases, 18(6):1369–1392, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Martens:2012:DAX

[MNNS12] Wim Martens, Matthias Niewerth, Frank Neven, and Thomas Schwentick. Developing and analyzing XSDs through BonXai.

Proceedings of the VLDB Endowment, 5(12):1994–1997, August 2012. CODEN ???? ISSN 2150-8097.

Martens:2023:RPG

[MNP^{+}23] Wim Martens, Matthias Niewerth, Tina Popp, Carlos Rojas, Stijn Vansummeren, and Domagoj Vrgoc. Representing paths in graph database pattern matching. Proceedings of the VLDB Endowment, 16(7):1790–1803, March 2023. CODEN

???? ISSN 2150-8097. URL https://dl.acm.org/doi/10.
14778/3587136.3587151.

Moseley:2023:DBN

[MNR^{+}23] Dan Moseley, Mario Nishio, Jose Perez Rodriguez, Olli Saarikivi, Stephen Toub, Margus Veanes, Tiki Wan, and Eric Xu. Derivative based nonbacktracking real-world regex matching with backtracking semantics. Proceedings of the ACM on Programming Languages, 7-PLDI:1026–1049, June 2023.

Moller-Nielsen:1984:EFS

[MNS84] Peter Moller-Nielsen and Jorgen Staunstrup. Experiments with a fast string searching algorithm. Information Processing Letters, 18(3):129–135, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martens:2007:SSA

[MNS07] Wim Martens, Frank Neven, and Thomas Schwentick. Simple off the shelf abstractions for XML schema. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(3): 15–22, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martens:2010:CDP

[MNS10] Wim Martens, Frank Neven, and Thomas Schwentick. Complexity of decision problems for XML schemas and chain regular expressions. SIAM Journal on Computing, 39(4):1486– 1530, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Makinen:2005:TIS

[MNU05] Veli Mäkinen, Gonzalo Navarro, and Esko Ukkonen. Transposition invariant string matching. Journal of Algorithms, 56(2):124–153, August 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001427.

Myers:1998:REA

[MOG98] E. G. Myers, P. Oliva, and K. Guimaraes. Reporting exit and approximate regular expression matches. Lecture Notes in Computer Science, 1448:91–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohri:1994:MST

[Moh94] M. Mohri. Minimization of sequential transducers. Lecture Notes in Computer Science, 807:151–163, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mohri:1997:SMA

[Moh97] Mehryar Mohri. String-matching with automata. Nordic Journal of Computing, 4(2):217–231, Summer 1997. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.helsinki.

fi/njc/References/mohri1997:217.html.

Montanaro:2017:QPM

[Mon17] Ashley Montanaro. Quantum pattern matching fast on average. Algorithmica, 77(1):16–39, January 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s00453-015-0060-4.

Moore:1964:SMS

[Moo64] Edward F. Moore. Sequential Machines: Selected Papers. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, 1964. v + 266 pp. LCCN QA76.5 .M57.

Moore:2012:MLF

[Moo12] J. Strother Moore. Meta-level features in an industrialstrength theorem prover. ACM SIGPLAN Notices, 47(1): 425–426, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moran:1983:CDO

[Mor83] Shlomo Moran. On the complexity of designing optimal partial-match retrieval systems. ACM Transactions on Database Systems, 8(4):543–551, December 1983.

CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644

(electronic). URL http://www.acm.org/pubs/articles/
journals/tods/1983-8-4/p543-moran/p543-moran.pdf;
http://www.acm.org/pubs/citations/journals/tods/1983-
8-4/p543-moran/.

Morris:1990:PER

[Mor90] Joseph M. Morris. Programming by expression refinement: the KMP algorithm. In Feijen et al. [FvGGM90], chapter 37, pages 327–338. ISBN 0-387-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990.

Morris:2002:AGJ

[Mor02] D. S. Morris. Automatically grading Java programming assignments via reflection, inheritance, and regular expressions. Frontiers in Education Conference, 1(??):T3G–22–??, 2002. CODEN PFECDR. ISSN 0190-5848.

Mignot:2018:EAC

[MOSZ18] Ludovic Mignot, Nadia Ouali-Sebti, and Djelloul Ziadi. An efficient algorithm for the construction of the equation tree automaton. International Journal of Foundations of Computer Science (IJFCS), 29(6):951–978, September 2018. CO-DEN IFCSEN. ISSN 0129-0541. URL https://www.

worldscientific.com/doi/10.1142/S0129054118500156.

Moy:2024:KMP

[Moy24] Cameron Moy. Knuth–Morris–Pratt illustrated. Journal of Functional Programming, 34:??, ???? 2024. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/knuthmorrispratt-illustrated/
8EFA77D663D585B68630E372BCE1EBA4.

Mitchell:1988:OHS

[MP88] Joan L. Mitchell and William B. Pennebaker. Optimal hardware and software arithmetic coding procedures for the Q-Coder. IBM Journal of Research and Development, 32(6): 727–736, November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Mauri:2005:APM

[MP05] Giancarlo Mauri and Giulio Pavesi. Algorithms for pattern matching and discovery in RNA secondary structure. Theoretical Computer Science, 335(1):29–51, May 20, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marion:2009:SIS

[MP09] Jean-Yves Marion and Romain Péchoux. Sup-interpretations, a semantic method for static analysis of program resources. ACM Transactions on Computational Logic, 10(4):27:1–27:??, August 2009. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Moreira:2012:DRE

[MPdS12] Nelma Moreira, David Pereira, and Simão Melo de Sousa. Deciding regular expressions (in-)Equivalence in Coq. Lecture Notes in Computer Science, 7560:98–113, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33314-9_7/.

Meiners:2014:FRE

[MPN^{+}14] Chad R. Meiners, Jignesh Patel, Eric Norige, Alex X. Liu, and Eric Torng. Fast regular expression matching using small TCAM. IEEE/ACM Transactions on Networking, 22(1):94– 109, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mraz:2021:TDP

[MPW21] František Mráz, Daniel Průša, and Michael Wehar. Twodimensional pattern matching against local and regular-like picture languages. Theoretical Computer Science, 870(??): 137–152, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304397520307490.

Morgan:1982:AEI

[MR82] Thomas M. Morgan and Lawrence A. Rowe. Analyzing exotic instructions for a retargetable code generator. ACM SIG-PLAN Notices, 17(6):197–204, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muthukrishnan:1992:SMU

[MR92] S. Muthukrishnan and H. Ramesh. String matching under a general matching relation. Lecture Notes in Computer Science, 652:356–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moreira:2005:IMR

[MR05] Nelma Moreira and Rogério Reis. Interactive manipulation of regular objects with FAdo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 335–339, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitchell:2009:APE

[MR09a] Neil Mitchell and Colin Runciman. Not all patterns, but enough: an automatic verifier for partial but sufficient pattern matching. ACM SIGPLAN Notices, 44(2):49–60, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moreira:2009:SPA

[MR09b] Nelma Moreira and Rogério Reis. Series-parallel automata and short regular expressions. Fundamenta Informaticae, 91 (3–4):611–629, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marschall:2011:ACC

[MR11] Tobias Marschall and Sven Rahmann. An algorithm to compute the character access count distribution for pattern matching algorithms. Algorithms (Basel), 4(4):285–306, December 2011. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/4/4/285.

Mooney:2013:SPM

[MR13] Carl H. Mooney and John F. Roddick. Sequential pattern mining — approaches and algorithms. ACM Computing Surveys, 45(2):19:1–19:39, February 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Mamouras:2017:SMS

[MRA^{+}17] Konstantinos Mamouras, Mukund Raghothaman, Rajeev Alur, Zachary G. Ives, and Sanjeev Khanna. StreamQRE: modular specification and efficient evaluation of quantitative queries over streaming data. ACM SIGPLAN Notices, 52 (6):693–708, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moataz:2018:SSE

[MRR^{+}18] Tarik Moataz, Indrajit Ray, Indrakshi Ray, Abdullatif Shikfa, Frédéric Cuppens, and Nora Cuppens. Substring search over encrypted data. Journal of Computer Security, 26(1):1–30, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Moore:1995:COA

[MS95] Dennis Moore and W. F. Smyth. A correction to “An optimal algorithm to compute all the covers of a string”. Information Processing Letters, 54(2):101–103, April 28, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madhavan:1998:ORT

[MS98] M. Madhavan and P. Shankar. Optimal regular tree pattern matching using pushdown automata. Lecture Notes in Computer Science, 1530:122–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mongelli:2001:PPM

[MS01] H. Mongelli and S. W. Song. Parallel pattern matching with scaling. Parallel Processing Letters, 11(1):125–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Monnier:2020:EEL

[MS20] Stefan Monnier and Michael Sperber. Evolution of Emacs Lisp. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):74:1–74:55, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/3386324.

Matsubara:2017:NDI

[MSP^{+}17] Yasuko Matsubara, Yasushi Sakurai, B. Aditya Prakash, Lei Li, and Christos Faloutsos. Nonlinear dynamics of information diffusion in social networks. ACM Transactions on the Web (TWEB), 11(2):11:1–11:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Madhavan:2000:EGG

[MSRR00] Maya Madhavan, Priti Shankar, Siddhartha Rai, and U. Ramakrishna. Extending Graham-Glanville techniques for optimal code generation. ACM Transactions on Programming Languages and Systems, 22(6):973–1001, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593

(electronic). URL http://www.acm.org/pubs/articles/
journals/toplas/2000-22-6/p973-madhavan/p973-madhavan.
pdf;   http://www.acm.org/pubs/citations/journals/
toplas/2000-22-6/p973-madhavan/.

Meysman:2019:MES

[MSS^{+}19] Pieter Meysman, Yvan Saeys, Ehsan Sabaghian, Wout Bittremieux, Yves Van de Peer, Bart Goethals, and Kris Laukens. Mining the enriched subgraphs for specific vertices in a biological graph. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(5):1496–1507, September 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Martinez:2024:PPP

[MSV24] Virginia Ardévol Martínez, Florian Sikora, and Stéphane Vialette. Parity permutation pattern matching. Algorithmica, 86(8):2605–2624, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01237-
0.

Mignot:2017:TAC

[MSZ17] Ludovic Mignot, Nadia Ouali Sebti, and Djelloul Ziadi. Tree automata constructions from regular expressions: a comparative study. Fundamenta Informaticae, 156(1):69–94, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mohanty:2014:SOS

[MT14] Pragyan (Sheela) Mohanty and Spyros Tragoudas. Scalable offline searches in DNA sequences. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):18:1– 18:??, November 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

MuQqoz:1995:MTW

[Mu 95] J. Mu Qq oz. Manual TIG welding and control chart advisor: a prototype program in Common Lisp using pattern matching techniques. Quality Engineering, 7(2):277–??, 1995. CODEN QUENE7. ISSN 0898-2112.

Makinen:2002:LSB

[MU02] Veli Mäkinen and Esko Ukkonen. Local similarity based pointpattern matching. Lecture Notes in Computer Science, 2373: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2373/23730115.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/2373/23730115.pdf.

Matsushita:1996:FPM

[MUHT96] M. Matsushita, M. Umano, I. Hatono, and H. Tamura. A fast pattern-matching algorithm using matching candidates for production systems. Lecture Notes in Computer Science, 1114:646–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Munoz:1994:MTW

[Mun95] J. Munoz. Manual TIG welding and control chart advisor. A prototype program in Common Lisp using pattern matching techniques. Quality Engineering, 7(2):277–??, 1994/1995. CODEN QUENE7. ISSN 0898-2112.

Mungan:2007:SML

[Mun07] Muhittin Mungan. String matching and 1D lattice gases. Journal of Statistical Physics, 126(1):207–242, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-006-9247-z.

Mustafa:2003:MDS

[Mus03] Suleiman H. Mustafa. A morphology-driven string matching approach to Arabic text searching. The Journal of Systems and Software, 67(2):77–87, August 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mustafa:2005:WOA

[Mus05] Suleiman H. Mustafa. Word-oriented approximate string matching using occurrence heuristic tables: a heuristic for searching Arabic text. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1504– 1511, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

MuThoz:1995:MTW

[MuT95] J. MuThoz. Manual TIG welding and control chart advisor: a prototype program in Common Lisp using pattern matching techniques. Quality engineering, 7(2):277–??, ???? 1995. ISSN 0898-2112.

Muthukrishnan:1997:DFM

[Mut97] S. Muthukrishnan. Detecting false matches in string-matching algorithms. Algorithmica, 18(4):512–520, August 1997. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=18&issue=4&
spage=512.

Muthukrishnan:2000:SOP

[Mut00] S. Muthukrishnan. Simple optimal parallel multiple pattern matching. Journal of Algorithms, 34(1):1–13, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-

2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910155.

Manber:1992:APM

[MW92a] Udi Manber and Sun Wu. Approximate pattern matching. Byte Magazine, 17(??):??, ?? 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manber:1992:SAR

[MW92b] Udi Manber and Sun Wu. Some assembly required. approximate pattern matching: Agrep’s algorithms let you perform text searches using an approximate pattern. Byte Magazine, 17(12):281–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manber:1994:AAM

[MW94] Udi Manber and Sun Wu. An algorithm for approximate membership checking with application to password security. Information Processing Letters, 50(4):191–197, May 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McNaughton:1960:RES

[MY60] R. McNaughton and H. Yamada. Regular expressions and state graphs for automata. IRE Transactions on Electronic Computers, EC-9(1):39–47, March 1960. CODEN IRE-

LAO. ISSN 0367-9950. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5221603. Reprinted in [?,

157–174].

Myers:1992:FRA

[Mye92] Gene Myers. A four Russians algorithm for regular expression pattern matching. Journal of the Association for Computing Machinery, 39(2):430–448, April 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/128755.html.

Myers:1995:AMC

[Mye95] Gene Myers. Approximately matching context-free languages. Information Processing Letters, 54(2):85–92, April 28, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Myers:1998:FBV

[Mye98] G. Myers. A fast bit-vector algorithm for approximate string matching based on dynamic programming. Lecture Notes in Computer Science, 1448:1–??, 1998. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Myers:1999:FBV

[Mye99] Gene Myers. A fast bit-vector algorithm for approximate string matching based on dynamic programming. Journal of the Association for Computing Machinery, 46(3):395–415, May 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1999-46-3/p395-myers/.

Ma:2007:CPM

[MZ07] Bin Ma and Kaizhong Zhang, editors. Combinatorial Pattern Matching: 18th Annual Symposium, CPM 2007, London, Canada, July 9–11, 2007. Proceedings, volume 4580 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-73436-8 (print), 3-540-73437-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/
978-3-540-73437-6.

Mozafari:2010:REN

[MZZ10] Barzan Mozafari, Kai Zeng, and Carlo Zaniolo. From regular expressions to nested words: unifying languages and query execution for relational and XML sequences. Proceedings of the VLDB Endowment, 3(1–2):150–161, September 2010. CO-DEN ???? ISSN 2150-8097.

Noor:1990:ASC

[NA90] Ahmed K. Noor and Carl M. Anderson. Application of symbolic computation to geometrically nonlinear analysis of curved beams. In Noor et al. [NEH90], pages 115–148. CO-DEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Nagy:2021:UFR

[Nag21] Benedek Nagy. Union-freeness revisited — between deterministic and nondeterministic union-free languages. International Journal of Foundations of Computer Science (IJFCS), 32(05): 551–573, August 2021. ISSN 0129-0541. URL https://www.

worldscientific.com/doi/10.1142/S0129054121410070.

Nakano:2014:OIA

[Nak14] Koji Nakano. Optimal implementations of the approximate string matching and the approximate discrete signal matching on the memory machine models. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(2): 104–118, 2014.

Naor:1991:SMP

[Nao91] M. Naor. String matching with preprocessing of text and pattern. Lecture Notes in Computer Science, 510:739–??, 1991.

CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narayanan:1991:PME

[Nar91] V. Narayanan. Pattern matching for everyone. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):27–30, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nicoara:2008:CSE

[NAR08] Angela Nicoara, Gustavo Alonso, and Timothy Roscoe. Controlled, systematic, and efficient code replacement for running Java programs. Operating Systems Review, 42(4):233– 246, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Navarro:1998:IAP

[Nav98] Gonzalo Navarro. Improved approximate pattern matching on hypertext. Lecture Notes in Computer Science, 1380:352– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1380/13800352.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1380/13800352.pdf.

Navarro:2000:IAP

[Nav00] Gonzalo Navarro. Improved approximate pattern matching on hypertext. Theoretical Computer Science, 237(1–2):455–463, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/45/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/171/21/45/article.pdf.

Navarro:2001:GTA

[Nav01a] Gonzalo Navarro. A guided tour to approximate string matching. ACM Computing Surveys, 33(1):31–88, March 2001. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341

(electronic). URL http://www.acm.org/pubs/articles/
journals/surveys/2001-33-1/p31-navarro/p31-navarro.
pdf;   http://www.acm.org/pubs/citations/journals/
surveys/2001-33-1/p31-navarro/.

Navarro:2001:NGF

[Nav01b] Gonzalo Navarro. NR-grep: a fast and flexible patternmatching tool. Software — Practice and Experience, 31(13): 1265–1312, November 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/85512195/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=85512195&PLACEBO=IE.pdf.

Navarro:2001:RES

[Nav01c] Gonzalo Navarro. Regular expression searching over Ziv– Lempel compressed text. Lecture Notes in Computer Science, 2089:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2089/20890001.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/2089/20890001.pdf.

Navarro:2004:ARE

[Nav04a] G. Navarro. Approximate regular expression searching with arbitrary integer weights. Nordic Journal of Computing, 11 (4):356–??, Winter 2004. CODEN NJCOFR. ISSN 1236-6064.

Navarro:2004:PM

[Nav04b] Gonzalo Navarro. Pattern matching. Journal of Applied Statistics, 31(8):925–949, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Navarro:2021:IHRb

[Nav21a] Gonzalo Navarro. Indexing highly repetitive string collections, part i: Repetitiveness measures. ACM Computing Surveys, 54 (2):29:1–29:31, April 2021. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.

org/doi/10.1145/3434399.

Navarro:2021:IHRa

[Nav21b] Gonzalo Navarro. Indexing highly repetitive string collections, Part II: Compressed indexes. ACM Computing Surveys, 54(2): 26:1–26:32, April 2021. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

doi/10.1145/3432999.

Navarro:1999:FMD

[NBY99a] Gonzalo Navarro and Ricardo Baeza-Yates. Fast multidimensional approximate pattern matching. Lecture Notes in Computer Science, 1645:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1645/16450243.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1645/16450243.pdf.

Navarro:1999:NIM

[NBY99b] Gonzalo Navarro and Ricardo Baeza-Yates. A new indexing method for approximate string matching. Lecture Notes in Computer Science, 1645:163–??, 1999. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1645/16450163.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1645/16450163.pdf.

Navarro:1999:VFS

[NBY99c] Gonzalo Navarro and Ricardo Baeza-Yates. Very fast and simple approximate string matching. Information Processing Letters, 72(1–2):65–70, October 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.elsevier.nl/gej-ng/29/28/24/60/27/33/abstract.
html;   http://www.elsevier.nl/gej-ng/29/28/24/60/
27/33/article.pdf.

Navarro:2001:IAA

[NBY01] G. Navarro and R. Baeza-Yates. Improving an algorithm for approximate pattern matching. Algorithmica, 30(4):473–502, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=

30&issue=4&spage=473. Algorithm engineering.

Neraud:1992:SMI

[NC92] J. Neraud and M. Crochemore. A string matching interpretation of the equation x^{m}y^{n} = z^{p}. Theoretical Computer Science, 92(1):145–164, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Navarro:2006:MIA

[NC06] Gonzalo Navarro and Edgar Chávez. A metric index for approximate string matching. Theoretical Computer Science, 352(1–3):266–279, March 7, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narayan:2018:MTR

[NCJF18] Apurva Narayan, Greta Cutulenco, Yogi Joshi, and Sebastian Fischmeister. Mining timed regular specifications from system traces. ACM Transactions on Embedded Computing Systems, 17(2):46:1–46:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ni:2014:HCD

[NCKL14] Lionel Ni, Lei Chen, Lei Kang, and Siyuan Liu. How to conduct distributed incomplete pattern matching. IEEE Transactions on Parallel and Distributed Systems, 25(4):982–992, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nordio:2010:IQE

[NCV10] Alessandro Nordio, Carla-Fabiana Chiasserini, and Emanuele Viterbo. The impact of quasi-equally spaced sensor topologies on signal reconstruction. ACM Transactions on Sensor Networks, 6(2):11:1–11:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nedjah:2002:PMC

[ND02] Nadia Nedjah and Luiza De Macedo Mourelle. Pattern matching code minimization in rewriting-based programming languages. International Journal of Foundations of Computer Science (IJFCS), 13(6):873–??, December 2002. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nedjah:2002:ECD

[NdMM02a] Nadia Nedjah and Luiza de Macedo Mourelle. Efficient concise deterministic pattern-matching automata for ambiguous patterns. ACM SIGPLAN Notices, 37(2):57–67, February 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nedjah:2002:OAP

[NdMM02b] Nadia Nedjah and Luiza de Macedo Mourelle. Optimal adaptive pattern matching. Lecture Notes in Computer Science, 2358:768–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2358/
23580768.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2358/23580768.pdf.

Nakatani:1993:MCB

[NE93] Toshio Nakatani and Kemal Ebcioğlu. Making compactionbased parallelization affordable. IEEE Transactions on Parallel and Distributed Systems, 4(9):1014–1029, September 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nebel:2006:FSM

[Neb06] Markus E. Nebel. Fast string matching by using probabilities: On an optimal mismatch variant of Horspool’s algorithm. Theoretical Computer Science, 359(1–3):329–343, August 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nedjah:1998:MDL

[Ned98] Nadia Nedjah. Minimal deterministic left-to-right patternmatching automata. ACM SIGPLAN Notices, 33(1):40–47, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noor:1990:SCT

[NEH90] Ahmed K. Noor, Isaac Elishakoff, and Greg Hulbert, editors. Symbolic computations and their impact on mechanics: presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Dallas, Texas, November 25–30, 1990, volume 205 of PVP. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1990. CODEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Nelson:1996:AAF

[Nel96] Mark R. Nelson. Algorithm alley: Fast string searches with suffix trees. Dr. Dobb’s Journal of Software Tools, 21(8):115– ??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Nsira:2017:LSM

[NEL17] Nadia Ben Nsira, Mourad Elloumi, and Thierry Lecroq. Online string matching in highly similar DNA sequences. Mathematics in Computer Science, 11(2):113–126, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Neuburger:2010:BRB

[Neu10] Shoshana Neuburger. Book review: The Burrows–Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching, by Donald Adjeroh, Timothy Bell and Amar Mukherjee Springer, 2008. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 41(1): 21–24, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [ABM08].

Navarro:2004:ACE

[NF04] Gonzalo Navarro and Kimmo Fredriksson. Average complexity of exact and approximate multiple string matching. Theoretical Computer Science, 321(2–3):283–290, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ng:1979:SAC

[Ng79] Edward W. Ng, editor. Symbolic and Algebraic Computation: EUROSAM ’79, an International Symposium on Symbolic and Algebraic Manipulation, Marseille, France, June 1979, volume 72 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-09519-5. LCCN QA155.7.E4 I57 1979.

Nielsen:2011:BCR

[NH11] Lasse Nielsen and Fritz Henglein. Bit-coded regular expression parsing: 5th International Conference, LATA 2011, Tarragona, Spain, May 26–31, 2011. In Carlos Martín-Vide, Shunsuke Inenaga, and Adrian-Horia Dediu, editors, Language and Automata Theory and Applications, volume 6638 of Lecture Notes in Computer Science, pages 402–413. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-21253-0 (soft cover), 3-642-21254-9 (e-book). LCCN QA267 .L38 2011.

Narisada:2020:ECL

[NHN^{+}20] Shintaro Narisada, Diptarama Hendrian, Kazuyuki Narisawa, Shunsuke Inenaga, and Ayumi Shinohara. Efficient computation of longest single-arm-gapped palindromes in a string. Theoretical Computer Science, 812(??):160–173, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397519306541.

Nicodeme:2003:RSP

[Nic03] Pierre Nicodème. Regexpcount, a symbolic package for counting problems on regular expressions and words. Fundamenta Informaticae, 56(1–2):71–88, January 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nilsen:1990:SDT

[Nil90] Kelvin Nilsen. A stream data type that supports goal-directed pattern matching on unbounded sequences of values. Computer Languages, 15(1):41–54, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nipkow:1998:VLA

[Nip98] Tobias Nipkow. Verified lexical analysis. In Jim Grundy and Malcolm Newey, editors, Theorem Proving in Higher Order Logics: 11th International Conference, TPHOLs’98, Canberra, Australia, September 27–October 1, 1998, Proceedings, volume 1479 of Lecture Notes in Computer Science, pages 1–15. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64987-5 (softcover), 3-540-49801-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T655 1998.

Nieminen:2007:EIA

[NK07] Janne Nieminen and Pekka Kilpeläinen. Efficient implementation of Aho–Corasick pattern matching automata using Unicode. Software — Practice and Experience, 37(6):669–690, May 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Navarro:2001:FAS

[NKT^{+}01] G. Navarro, T. Kida, M. Takeda, A. Shinohara, and S. Arikawa. Faster approximate string matching over compressed text. In Storer and Cohn [SC01], pages 459–468. ISBN 0-7695-1031-0. ISSN 1068-0314. LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917177. IEEE CSP number 01PR1031.

Navarro:2007:CFT

[NM07] Gonzalo Navarro and Veli Mäkinen. Compressed full-text indexes. ACM Computing Surveys, 39(1):2:1–2:61, April 2007. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Nourie:2010:REJ

[NM10] D. Nourie and M. McCloskey. Regular expressions and the Java programming language. Web report, 2010. URL http://

java.sun.com/developer/technicalArticles/releases/
1.4regex.

Nam:2022:RRE

[NNSP22] Jaehyun Nam, Seung Ho Na, Seungwon Shin, and Taejune Park. Reconfigurable regular expression matching architecture for real-time pattern update and payload inspection. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804522001497.

Navarro:1998:BPA

[NR98] G. Navarro and M. Raffinot. A bit-parallel approach to suffix automata: Fast extended string matching. Lecture Notes in Computer Science, 1448:14–??, 1998. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Navarro:1999:FRE

[NR99a] Gonzalo Navarro and Mathieu Raffinot. Fast regular expression search. Lecture Notes in Computer Science, 1668:198– 212, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1668/16680198.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1668/16680198.pdf.

Navarro:1999:GPA

[NR99b] Gonzalo Navarro and Mathieu Raffinot. A general practical approach to pattern matching over Ziv–Lempel compressed text. Lecture Notes in Computer Science, 1645:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1645/16450014.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1645/16450014.pdf.

Navarro:2000:FFS

[NR00] Gonzalo Navarro and Mathieu Raffinot. Fast and flexible string matching by combining bit-parallelism and suffix automata. ACM Journal of Experimental Algorithmics, 5:4:1– 4:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Navarro:2001:CDR

[NR01] Gonzalo Navarro and Mathieu Raffinot. Compact DFA representation for fast regular expression search. Lecture Notes in Computer Science, 2141:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2141/
21410001.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2141/21410001.pdf.

Navarro:2002:FPM

[NR02] Gonzalo Navarro and Mathieu Raffinot. Flexible pattern matching in strings: practical on-line search algorithms for texts and biological sequences. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-81307-7 (hardcover). x + 221 pp. LCCN QA76.9.A43

N38 2002. URL http://www.dcc.uchile.cl/~gnavarro/
FPMbook/;   http://www.loc.gov/catdir/description/
cam022/2001043704.html;   http://www.loc.gov/catdir/
toc/cam023/2001043704.html.

Navarro:2003:FSC

[NR03] Gonzalo Navarro and Mathieu Raffinot. Fast and simple character classes and bounded gaps pattern matching, with applications to protein searching. Journal of Computational Biology, 10(6):903–923, December 2003. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.liebertpub.com/doi/abs/10.
1089/106652703322756140;   https://www.liebertpub.
com/doi/pdf/10.1089/106652703322756140.

Navarro:2004:NTR

[NR04] Gonzalo Navarro and Mathieu Raffinot. New techniques for regular expression searching. Algorithmica, 41(2):89–116, November 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=
41&issue=2&spage=89.

Nicolae:2015:SMM

[NR15] Marius Nicolae and Sanguthevar Rajasekaran. On string matching with mismatches. Algorithms (Basel), 8(2):248–270, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/8/2/248.

Nicolae:2017:PMM

[NR17] Marius Nicolae and Sanguthevar Rajasekaran. On pattern matching with k mismatches and few don’t cares. Information Processing Letters, 118(??):78–82, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019016301442.

Nicaud:2021:RRE

[NR21] Cyril Nicaud and Pablo Rotondo. Random regular expression over huge alphabets. International Journal of Foundations of Computer Science (IJFCS), 32(05):419–438, August 2021.

ISSN 0129-0541. URL https://www.worldscientific.com/
doi/10.1142/S012905412141001X.

Neatherway:2012:TBA

[NRO12] Robin P. Neatherway, Steven J. Ramsay, and Chih-Hao Luke Ong. A traversal-based algorithm for higher-order model checking. ACM SIGPLAN Notices, 47(9):353–364, September 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ng:2018:SCN

[NRS18] Timothy Ng, David Rappaport, and Kai Salomaa. State complexity of neighbourhoods and approximate pattern matching. International Journal of Foundations of Computer Science (IJFCS), 29(2):315–??, February 2018. CODEN IFC-SEN. ISSN 0129-0541.

Navarro:2005:LBM

[NT05] Gonzalo Navarro and Jorma Tarhio. LZgrep: a Boyer–Moore string matching tool for Ziv–Lempel compressed text. Software — Practice and Experience, 35(12):1107–1130, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X

(electronic). URL http://www.dcc.uchile.cl/~gnavarro/
pubcode/.

Navarrete:2020:PRE

[NT20] Lise Rommel Romero Navarrete and Guilherme P. Telles. Practical regular expression constrained sequence alignment. Theoretical Computer Science, 815(??):95–108, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397520301018.

Nance:1993:SBE

[NTS93] Barry Nance, Tom Thompson, and Ben Smith. A small browser with everything: A powerful DOS browser, a faster finder, and Perl-based recursive grep. Byte Magazine, 18(2): 235–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280.

Nataliani:2013:IBV

[NW13] Yessica Nataliani and Theophilus Wellem. Implementation of bit-vector algorithm for approximate string matching on rhodopsin protein sequence. International Journal of Computer Applications, 72(??):34–38, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number14/12565-9214/.

Nedjah:1997:OLR

[NWE97] N. Nedjah, C. D. Walter, and S. E. Eldridge. Optimal left-to-right pattern-matching automata. Lecture Notes in Computer Science, 1298:273–??, 1997. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nedjah:1999:EAD

[NWE99] Nadia Nedjah, Colin D. Walter, and Stephen E. Eldridge. Efficient automata-driven pattern-matching for equational programs. Software — Practice and Experience, 29 (9):793–813, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
62501864;   http://www3.interscience.wiley.com/cgi-
bin/fulltext?ID=62501864&PLACEBO=IE.pdf.

Najam:2015:SPP

[NYuR15] Maleeha Najam, Usman Younis, and Raihan ur Rasool. Speculative parallel pattern matching using stride-k DFA for deep packet inspection. Journal of Network and Computer Applications, 54(??):78–87, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804515000867.

Ortiz:1998:RES

[OA98] Alejandro A. R. Trejo Ortiz and Guillermo Fernández Anaya. Regular expression simplification. Mathematics and Computers in Simulation, 45(1–2):59–71, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0378475497000864.

Omar:2017:PSF

[OA17] Cyrus Omar and Jonathan Aldrich. Programmable semantic fragments: the design and implementation of typy. ACM SIGPLAN Notices, 52(3):81–92, March 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odeh:2014:NER

[Ode14] Jehad Q. Odeh. New and efficient recursive-based string matching algorithm (RSMA-FLFC). International Journal of Computer Applications, 86(??):1–6, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume86/number15/15058-3190/.

Okuboyejo:2025:AFG

[OES25] Olaperi Okuboyejo, Sigrid Ewert, and Ian Sanders. Automatic feedback generation for the learning of regular expressions. ACM Transactions on Computing Education, 25(3): 36:1–36:26, September 2025. CODEN ???? ISSN 1946-6226.

Ott:1961:DSM

[OF61] Gene Ott and Neil H. Feinstein. Design of sequential machines from their regular expressions. Journal of the Association for Computing Machinery, 8(4):585–600, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Oda:1994:PDC

[OK94] M. Oda and T. Kakeshita. Pitfall detection of C programs using pattern matching. Transactions of the Information Processing Society of Japan, 35(11):2427–2436, November 1994. CODEN JSGRD5. ISSN 0387-5806.

Ogawa:1992:RDF

[OKT92] R. Ogawa, Y. Kikuchi, and K. Takahashi. Recent developments in full text database technologies. Journal of the Information Processing Society of Japan = Joho Shori, 33(4): 404–412, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Owolabi:1988:FAS

[OM88] O. Owolabi and D. R. McGregor. Fast approximate string matching. Software — Practice and Experience, 18(4):387– 393, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Otto:1998:EUW

[OND98] Friedrich Otto, Paliath Narendran, and Daniel J. Dougherty. Equational unification and word unification, and 2nd-order equational unification. Theoretical Computer Science, 198(1– 2):1–47, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.

com/cas/tree/store/tcs/sub/1998/198/1-2/2639.pdf.

Ordyniak:2016:PSM

[OP16] Sebastian Ordyniak and Alexandru Popa. A parameterized study of maximum generalized pattern matching problems. Algorithmica, 75(1):1–26, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-0008-8.

Ophel:1989:IMR

[Oph89] John Ophel. An improved mixture rule for pattern matching. ACM SIGPLAN Notices, 24(6):91–96, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ong:2011:VHO

[OR11] C.-H. Luke Ong and Steven James Ramsay. Verifying higherorder functional programs with pattern-matching algebraic data types. ACM SIGPLAN Notices, 46(1):587–598, January 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oh:2012:MTS

[OR12] Doohwan Oh and Won W. Ro. Multi-threading and suffix grouping on massive multiple pattern matching algorithm. The Computer Journal, 55(11):1331–1346, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/55/11/1331.full.pdf+html.

Organtini:2003:LRE

[Org03] Giovanni Organtini. Learning regular expressions. Linux Journal, 109:91–93, May 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Okcan:2013:SEA

[ORPF13] Alper Okcan, Mirek Riedewald, Biswanath Panda, and Daniel Fink. Scolopax: exploratory analysis of scientific data. Proceedings of the VLDB Endowment, 6(12):1298–1301, August 2013. CODEN ???? ISSN 2150-8097.

Owens:2008:RED

[ORT08] Scott Owens, John Reppy, and Aaron Turon. Regularexpression derivatives reexamined. Report, University of Cambridge and University of Chicago and Northeastern University, Cambridge, UK; Chicago, IL, USA; Boston, MA, USA, August 12, 2008. 18 pp. URL http://www.ccs.neu.

edu/home/turon/re-deriv.pdf.

Owens:2009:RED

[ORT09] Scott Owens, John Reppy, and Aaron Turon. Regularexpression derivatives re-examined. Journal of Functional Programming, 19(2):173–190, March 2009. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
E5734B86DEB96C61C69E5CF3C4FB0AFA.

Okui:2011:DRE

[OS11] Satoshi Okui and Taro Suzuki. Disambiguation in regular expression matching via position automata with augmented transitions. In Michael Domaratzki and Kai Salomaa, editors, Implementation and Application of Automata: 15th International Conference, CIAA 2010, Manitoba, Canada, August 12-15, 2010. Revised Selected Papers, volume 6482 of Lecture Notes in Computer Science, pages 231–240. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011.

Ohtani:1994:EITa

[OSM94a] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU Epoch. In Bundy [Bun94], pages 816–820. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Ohtani:1994:EITb

[OSM94b] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU Epoch. In Bundy [Bun94], pages 816–820. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Ohtani:1994:ETG

[OSM94c] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU Epoch. In Bundy [Bun94], pages 816–820. ISBN 3-540-58156-1, 0-387-58156-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D337I58 1994.

Orencik:2016:MKS

[OSSK16] Cengiz Orencik, Ayse Selcuk, Erkay Savas, and Murat Kantarcioğlu. Multi-keyword search over encrypted data with scoring and search pattern obfuscation. International Journal of Information Security, 15(3):251–269, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270

(electronic). URL http://link.springer.com/article/
10.1007/s10207-015-0294-9.

Otto:1994:TFP

[Ott94] Erick Otto. Two fast pattern-matching algorithms. C Users Journal, 12(2):39–??, February 1994. ISSN 0898-9788.

Orpaz:2003:PMM

[OW03] A. Orpaz and S. Weiss. Pattern matching by means of multi-resolution compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314. LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194060. IEEE Computer

Society Order number PR01896.

Oram:2007:BC

[OW07] Andrew Oram and Greg Wilson, editors. Beautiful code. Theory in practice. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2007. ISBN 0-596-51004-7 (paperback). xxi + 593 pp. LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http://www.oreilly.

com/catalog/9780596510046.

Owolabi:1993:EPS

[Owo93] Olumide Owolabi. Efficient pattern searching over large dictionaries. Information Processing Letters, 47(1):17–21, August 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Or:2016:MBH

[OWP16] N. L. Or, X. Wang, and D. Pao. MEMORY-based hardware architectures to detect ClamAV virus signatures with restricted regular expression features. IEEE Transactions on Computers, 65(4):1225–1238, April 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Prasad:2010:PSM

[PA10] Rajesh Prasad and Suneeta Agarwal. Parameterized string matching: an application to software maintenance. ACM SIGSOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pagan:1978:FSS

[Pag78] F. G. Pagan. Formal semantics of a Snobol4 subset. Computer Languages, 3(1):13–30, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Perez:2009:SCS

[PAG09] Jorge Pérez, Marcelo Arenas, and Claudio Gutierrez. Semantics and complexity of SPARQL. ACM Transactions on Database Systems, 34(3):16:1–16:??, August 2009. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Pakin:1991:REG

[Pak91] Scott Pakin. Regular expressions and gender guessing. Computer Language Magazine, 8(12):59–??, December 1991. CO-DEN COMLEF. ISSN 0749-2839.

Pandey:2012:PDS

[PAMP12] Kusum Lata Pandey, Suneeta Agarwal, Sanjay Misra, and Rajesh Prasad. Plagiarism detection in software using efficient string matching. Lecture Notes in Computer Science, 7336: 147–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-31128-4_11/.

Park:1996:ATD

[Par96] K. Park. Analysis of two-dimensional approximate pattern matching algorithms. Lecture Notes in Computer Science, 1075:335–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1998:ATD

[Par98] Kunsoo Park. Analysis of two-dimensional approximate pattern matching algorithms. Theoretical Computer Science, 201(1–2):263–273, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/tcs/sub/1998/201/
1-2/2812.pdf.

Patel:1971:GRL

[Pat71] A. R. Patel. Generation of right-linear grammars from regular expressions. IEEE Transactions on Computers, C-20(4):472– 473, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1671864.

Paterson:1999:CPM

[PC99] Michael S. Paterson and Maxime Crochemore, editors. Combinatorial pattern matching: 10th annual symposium, CPM 99, Warwick University, UK, July 22–24, 1999: Proceedings, volume 1645 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISBN 3-540-66278-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.9.A43 C65 1999. URL http://link.springer-ny.
com/link/service/series/0558/tocs/t1645.htm;   http:
//www.springerlink.com/content/978-3-540-66278-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=1645.

Park:2002:EQP

[PC02] Chang-Won Park and Chin-Wan Chung. An effective query pruning technique for multiple regular path expressions. The Journal of Systems and Software, 64(3):219–233, December 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Prins:1999:ICF

[PCS99] Jan F. Prins, Siddhartha Chatterjee, and Martin Simons. Irregular computations in Fortran — expression and implementation strategies. Scientific Programming, 7(3–4):313–326, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X

(electronic). URL http://iospress.metapress.com/app/
home/contribution.asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=issue%2C10%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%2C1%2C1.

Peng:2012:TBN

[PD12] Kunyang Peng and Qunfeng Dong. TCAM-based NFA implementation. ACM SIGMETRICS Performance Evaluation Review, 40(1):379–380, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Prasad:1994:EEP

[PDC94] Sushil K. Prasad, Sajal K. Das, and Calvin C.-Y. Chen. Efficient EREW PRAM algorithms for parentheses-matching. IEEE Transactions on Parallel and Distributed Systems, 5(9): 995–1008, September 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pajares:1998:PRL

[PDL98] G. Pajares, J. M. De la Cruz, and J. A. Lopez. Pattern recognition learning applied to stereovision matching. Lecture Notes in Computer Science, 1451:997–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peleg:1987:CPS

[Pel87] D. Peleg. Concurrent program schemes and their logics. Theoretical Computer Science, 55(1):1–45, November 1987. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pepper:1991:LPD

[Pep91] P. Pepper. Literate program derivation: a case study. In Methods of programming. Selected papers on the CIP-Project [BW91], pages 101–124. CODEN LNCSD9. ISBN 3-540-54576-X (Berlin), 0-387-54576-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M4543 1991,

QA267.A1 L43 no.544. URL http://link.springer-ny.
com/link/service/series/0558/tocs/t0544.htm;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=544.

Perleberg:1994:SCS

[Per94] Chris H. Perleberg. Single character searching methods and the shift-or pattern-matching algorithm. Information Processing Letters, 50(5):269–275, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettersson:1992:TPM

[Pet92] Mikael Pettersson. A term pattern-match compiler inspired by finite automata theory. Lecture Notes in Computer Science, 641:258–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:1994:RSM

[Pet94] H. Petersen. Refined simulation of multihead automata. Information Processing Letters, 52(5):229–233, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:1995:RPB

[Pet95] H. Petersen. A remark on a paper by A. B. Matos. Theoretical Computer Science, 141(1–2):329–330, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1851. See [Mat94].

Petersen:2002:MPR

[Pet02] Holger Petersen. The membership problem for regular expressions with intersection is complete in LOGCFL. Lecture Notes in Computer Science, 2285:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2285/22850513.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2285/22850513.pdf.

Petersen:2007:SMS

[Pet07] Holger Petersen. String matching with simple devices. Information Processing Letters, 105(1):32–34, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peacocke:1990:ISS

[PG90] Richard D. Peacocke and Daryl H. Graf. An introduction to speech and speaker recognition. Computer, 23(8):26–33, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Phillips:1994:ASM

[Phi94] Thomas Phillips. Approximate string matching. C Users Journal, 12(4):77–??, April 1994. ISSN 0898-9788.

Pan:2019:ARR

[PHXD19] Rong Pan, Qinheping Hu, Gaowei Xu, and Loris D’Antoni. Automatic repair of regular expressions. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 139:1–139:29, October 2019. URL https://dl.acm.org/

doi/abs/10.1145/3360565.

Pierre:1995:DVS

[Pie95] L. Pierre. Describing and verifying synchronous circuits with the Boyer–Moore theorem prover. Lecture Notes in Computer Science, 987:35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pientka:2008:TTF

[Pie08] Brigitte Pientka. A type-theoretic foundation for programming with higher-order abstract syntax and first-class substitutions. ACM SIGPLAN Notices, 43(1):371–382, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pike:1987:TES

[Pik87] Rob Pike. The text editor sam. Software — Practice and Experience, 17(11):813–845, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:2000:TES

[Pik00] Rob Pike. The text editor sam. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, June 7, 2000. 30 pp. URL http://plan9.

bell-labs.com/sys/doc/sam/sam.pdf.

Pike:2006:SRE

[Pik06] Rob Pike. Structural regular expressions. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 23, 2006. 7

pp. URL https://doc.cat-v.org/bell_labs/structural_
regexps/se.pdf.

Park:2017:PSS

[PIR17] Changhee Park, Hyeonseung Im, and Sukyoung Ryu. Precise and scalable static analysis of jQuery using a regular expression domain. ACM SIGPLAN Notices, 52(2):25–36, February 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pitt:1998:VPJ

[Pit98] W. David Pitt. The visitor pattern and a Java grep utility. Dr. Dobb’s Journal of Software Tools, 23(6):30, 32, 93, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Penna:2003:SRE

[PIT^{+}03] Giuseppe Della Penna, Benedetto Intrigila, Enrico Tronci, et al. Synchronized regular expressions. Acta Informatica, 39 (1):31–70, January 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Petermann:2014:GBD

[PJMR14] André Petermann, Martin Junghanns, Robert Müller, and Erhard Rahm. Graph-based data integration and business intelligence with BIIIG. Proceedings of the VLDB Endowment, 7 (13):1577–1580, August 2014. CODEN ???? ISSN 2150-8097.

Padhi:2018:FFS

[PJP^{+}18] Saswat Padhi, Prateek Jain, Daniel Perelman, Oleksandr Polozov, Sumit Gulwani, and Todd Millstein. FlashProfile: a framework for synthesizing data profiles. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 150:1–150:28, October 2018. URL https://dl.acm.org/

doi/abs/10.1145/3276520.

Pramanik:1985:HPM

[PK85] Sakti Pramanik and Chung-Ta King. A hardware pattern matching algorithm on a dataflow. The Computer Journal, 28(3):264–269, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/28/3/264.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_03/tiff/265.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/266.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_03/tiff/267.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_03/tiff/269.tif.

Park:1995:SMH

[PK95] K. Park and Dong Kyue Kim. String matching in hypertext. Lecture Notes in Computer Science, 937:318–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Prasad:2018:NPD

[PKK18] Animesh Prasad, Manpreet Kaur, and Min-Yen Kan. Neural ParsCit: a deep learning-based reference string parser. International Journal on Digital Libraries, 19(4):323–337, November 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.com/article/
10.1007/s00799-018-0242-1.

Pao:2008:PAM

[PLL08] D. Pao, W. Lin, and B. Liu. Pipelined architecture for multistring matching. IEEE Computer Architecture Letters, 7(2): 33–36, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pao:2010:MEP

[PLL10] Derek Pao, Wei Lin, and Bin Liu. A memory-efficient pipelined implementation of the Aho–Corasick string-matching algorithm. ACM Transactions on Architecture and Code Optimization, 7(2):10:1–10:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Patel:2014:BSE

[PLT14] Jignesh Patel, Alex X. Liu, and Eric Torng. Bypassing space explosion in high-speed regular expression matching. IEEE/ ACM Transactions on Networking, 22(6):1701–1714, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Parker:1978:SCO

[PM78] K. P. Parker and E. J. McCluskey. Sequential circuit output probabilities from regular expressions. IEEE Transactions on Computers, C-27(3):222–231, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1675075.

Parolini:2023:SSA

[PM23] Francesco Parolini and Antoine Miné. Sound static analysis of regular expressions for vulnerabilities to denial of service attacks. Science of Computer Programming, 229(??): ??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print),

1872-7964 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167642323000424.

Papadias:2001:AST

[PMD01] Dimitris Papadias, Nikos Mamoulis, and Vasilis Delis. Approximate spatio-temporal retrieval. ACM Transactions on Information Systems, 19(1):53–96, 2001. CODEN

ATISET. ISSN 1046-8188. URL http://www.acm.org/pubs/
articles/journals/tois/2001-19-1/p53-papadias/p53-
papadias.pdf;   http://www.acm.org/pubs/citations/
journals/tois/2001-19-1/p53-papadias/.

Petricek:2011:EMP

[PMS11] Tomas Petricek, Alan Mycroft, and Don Syme. Extending monads with pattern matching. ACM SIGPLAN Notices, 46 (12):1–12, December 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pol:2001:PST

[Pol01] Koos Pol. A powerful search tool for ASCII files. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):49–51, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http:

//www.samag.com/.

Poleksic:2013:IAM

[Pol13] Aleksandar Poleksic. Improved algorithms for matching rseparated sets with applications to protein structure alignment. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(1):226–229, January 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Poulovassilis:1993:PMA

[Pou93] A. Poulovassilis. A pattern-matching algorithm for functional databases. The Computer Journal, 36(2):195– 199, April 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/36/2/195.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/Volume_36/
Issue_02/Vol36_02.body.html#AbstractPoulovassilis.

Parigot:1985:LAP

[PP85] M. Parigot and E. Pelz. A logical approach of Petri net languages. Theoretical Computer Science, 39(2-3):155–169, August 1985. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Percus:1994:SMN

[PP94] Ora E. Percus and Jerome K. Percus. String matching for the novice. American Mathematical Monthly, 101(10):944– 947, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Paul:2006:SSC

[PP06] Souradyuti Paul and Bart Preneel. On the (in)security of stream ciphers based on arrays and modular addition. Lecture Notes in Computer Science, 4284:69–83, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/11935230_5.pdf.

Porat:2009:EAP

[PP09] B. Porat and E. Porat. Exact and approximate pattern matching in the streaming model. In IEEE [IEE09], pages 315–323. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Pasetto:2010:TVF

[PPA10] Davide Pasetto, Fabrizio Petrini, and Virat Agarwal. Tools for very fast regular expression matching. Computer, 43(3):50– 58, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Polo:2020:AGO

[PPPdG20] Macario Polo, Oscar Pedreira, Ángeles S. Places, and Ignacio García Rodríguez de Guzmán. Automated generation of oracled test cases with regular expressions and combinatorial techniques. Journal of Software: Evolution and Process, 32 (12):e2273:1–e2273:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Papadopoulos:2015:PAP

[PPTT15] Dimitrios Papadopoulos, Charalampos Papamanthou, Roberto Tamassia, and Nikos Triandopoulos. Practical authenticated pattern matching with optimal proof size. Proceedings of the VLDB Endowment, 8(7):750–761, February 2015. CODEN ???? ISSN 2150-8097.

Porat:2008:PMP

[PPZ08] Benny Porat, Ely Porat, and Asaf Zur. Pattern matching with pair correlation distance. Theoretical Computer Science, 407 (1–3):587–590, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prather:1997:REP

[Pra97] Ronald E. Prather. Regular expressions for program computations. American Mathematical Monthly, 104(2):120–130, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.org/pubs/

monthly_feb97_toc.html.

Preoteasa:1999:RBU

[Pre99] Viorel Preoteasa. A relation between unambiguous regular expressions and abstract data types. Fundamenta Informaticae, 40(1):53–77, October 1999. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Prajapati:2016:PAS

[PRT^{+}16] G. L. Prajapati, Abhijeet Singh Rathore, Bhavana Tanwar, Surbhi Bhadviy, and Tushar Jain. Performance analysis of selected string matching algorithms based on good suffix and bad character rule. International Journal of Computer Applications, 140(??):28–37, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number9/24623-2016909445/.

Pizzi:2011:FSM

[PRU11] Cinzia Pizzi, Pasi Rastas, and Esko Ukkonen. Finding significant matches of position weight matrices in linear time. IEEE/ ACM Transactions on Computational Biology and Bioinformatics, 8(1):69–79, January 2011. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Prusa:2017:CMS

[Prů17] Daniel Průša. Complexity of matching sets of two-dimensional patterns by two-dimensional on-line tessellation automaton. International Journal of Foundations of Computer Science (IJFCS), 28(5):623–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Purushothaman:1988:RAS

[PS88] S. Purushothaman and P. A. Subrahmanyam. Reasoning about systolic algorithms. Journal of Parallel and Distributed Computing, 5(6):669–699, December 1988. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Padberg:1989:CMB

[PS89] Manfred Padberg and Antonio Sassano. The complexity of matching with bonds. Information Processing Letters, 32(6): 297–300, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Partsch:1990:FPM

[PS90] H. A. Partsch and F. A. Stomp. A fast pattern matching algorithm derived by transformational and assertional reasoning. Formal Aspects of Computing, 2(1):109–122, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X

(electronic). URL http://link.springer.com/article/
10.1007/BF01888219.

Pai:1993:SCR

[PS93a] Dinesh K. Pai and Tony H. S. Ser. Simultaneous computation of robot kinematics and differential kinematics with automatic differentiation. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0823-9. 775–780 pp. LCCN TJ210.3.I447 1993. IEEE catalog number 93CH3213-6.

Puel:1993:CPM

[PS93b] Laurence Puel and Ascánder Suárez. Compiling pattern matching by term decomposition. Journal of Symbolic Computation, 15(1):1–26, January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pandurangan:2010:UOC

[PS10] Gopal Pandurangan and Wojciech Szpankowski. A universal online caching algorithm based on pattern matching. Algorithmica, 57(1):62–73, May 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=1&spage=62.

Patrick:2008:CEO

[PSK08] Christina M. Patrick, SeungWoo Son, and Mahmut Kandemir. Comparative evaluation of overlap strategies with study of I/O overlap in MPI-IO. Operating Systems Review, 42(6):43–49, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parreaux:2017:QSR

[PSK17] Lionel Parreaux, Amir Shaikhha, and Christoph E. Koch. Quoted staged rewriting: a practical approach to librarydefined optimizations. ACM SIGPLAN Notices, 52(12):131– 145, December 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2018:ETS

[PSP^{+}18] Ha-Myung Park, Francesco Silvestri, Rasmus Pagh, Chin-Wan Chung, Sung-Hyon Myaeng, and U. Kang. Enumerating trillion subgraphs on distributed systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6): 71:1–71:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Puxang:1997:BRB

[Pux97] K. Puxang. Book review: Metamathematics, Machines, and Gödels Proof, by N. Shankar. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 28 (4):16–19, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [?].

Partsch:1991:ACS

[PV91] H. A. Partsch and N. Volker. Another case study on reusability of transformational developments pattern matching according to Knuth, Morris, and Pratt. Lecture Notes in Computer Science, 544:35–??, 1991. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paredaens:1992:OG

[PVA^{+}92] Jan Paredaens, Jan Van den Bussche, Marc Andries, Marc Gemis, Marc Gyssens, Inge Thyssens, Dirk Van Gucht, Vijay Sarathy, and Lawrence Saxton. An overview of GOOD. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(1):25–31, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pitt:1993:MCD

[PW93] Leonard Pitt and Manfred K. Warmuth. The minimum consistent DFA problem cannot be approximated within any polynomial. Journal of the Association for Computing Machinery, 40(1):95–142, January 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/138042.html.

Pevzner:1995:MFA

[PW95] Pavel A. Pevzner and M. S. Waterman. Multiple filtration and approximate pattern matching. Algorithmica, 13(1–2): 135–154, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Plumlee:2006:ZVM

[PW06] Matthew D. Plumlee and Colin Ware. Zooming versus multiple window interfaces: Cognitive costs of visual comparisons. ACM Transactions on Computer-Human Interaction, 13(2):179–209, June 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pao:2012:MSS

[PW12] Derek Pao and Xing Wang. Multi-stride string searching for high-speed content inspection. The Computer Journal, 55 (10):1216–1231, October 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/55/10/1216.full.
pdf+html.

Palma:2025:FSH

[PWSG25] Marco Edoardo Palma, Alex Wolf, Pasquale Salza, and Harald C. Gall. On-the-fly syntax highlighting: Generalisation and speed-ups. IEEE Transactions on Software Engineering, 51(2):355–370, February 2025. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Pao:2011:SSE

[PWW^{+}11] Derek Pao, Xing Wang, Xiaoran Wang, Cong Cao, and Yuesheng Zhu. String searching engine for virus scanning.

IEEE Transactions on Computers, 60(11):1596–1609, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5669261.

Pei:2019:RPU

[PYY19] Shuyi Pei, Jing Yang, and Qing Yang. REGISTOR: a platform for unstructured data processing inside SSD storage. ACM Transactions on Storage, 15(1):7:1–7:??, April 2019. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3310149.

Qiu:2007:ESA

[QLY07] Shibin Qiu, Terran Lane, and Cundong Yang. Efficient search algorithms for RNAi target detection. The Journal of Supercomputing, 42(3):303–319, December 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=42&issue=3&spage=303.

Qu:2008:RPM

[QPWH08] Yingge Qu, Wai-Man Pang, Tien-Tsin Wong, and Pheng-Ann Heng. Richness-preserving manga screening. ACM Transactions on Graphics, 27(5):155:1–155:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qiao:2013:TKN

[QQC^{+}13a] Miao Qiao, Lu Qin, Hong Cheng, Jeffrey Xu Yu, and Wentao Tian. Top-k nearest keyword search on large graphs. Proceedings of the VLDB Endowment, 6(10):901–912, August 2013. CODEN ???? ISSN 2150-8097.

Qiao:2013:TNK

[QQC^{+}13b] Miao Qiao, Lu Qin, Hong Cheng, Jeffrey Xu Yu, and Wentao Tian. Top-k nearest keyword search on large graphs. Proceedings of the VLDB Endowment, 6(10):901–912, August 2013. CODEN ???? ISSN 2150-8097.

Qahtan:2020:PFD

[QTO^{+}20] Abdulhakim Qahtan, Nan Tang, Mourad Ouzzani, Yang Cao, and Michael Stonebraker. Pattern functional dependencies for data cleaning. Proceedings of the VLDB Endowment, 13(5):684–697, January 2020. CODEN ???? ISSN

2150-8097. URL https://dl.acm.org/doi/abs/10.14778/
3377369.3377377.

Quigley:2000:LSE

[Qui00] Ellie Quigley. Linux shells by example. Open source technology series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-014711-7. xviii + 761 pp. LCCN QA76.76.O63 Q538 2000.

Quigley:2002:USE

[Qui02] Ellie Quigley. UNIX Shells by Example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2002. ISBN 0-13-066538-X (paperback). xix + 1015 pp. LCCN QA76.76.O63 Q54 2002.

Quong:1992:FAC

[Quo92] R. W. Quong. Fast average-case pattern matching by multiplexing sparse tables. Theoretical Computer Science, 92(1): 165–179, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Qin:2013:ASS

[QWX^{+}13] Jianbin Qin, Wei Wang, Chuan Xiao, Yifei Lu, Xuemin Lin, and Haixun Wang. Asymmetric signature schemes for efficient exact edit similarity query processing. ACM Transactions on Database Systems, 38(3):16:1–16:??, August 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Qiao:2017:SMC

[QZC17] Miao Qiao, Hao Zhang, and Hong Cheng. Subgraph matching: on compression and computation. Proceedings of the VLDB Endowment, 11(2):176–188, October 2017. CODEN ???? ISSN 2150-8097.

Raita:1992:TBM

[Rai92] Timo Raita. Tuning the Boyer–Moore–Horspool string searching algorithm. Software — Practice and Experience, 22(10): 879–884, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raita:1999:GSD

[Rai99] Timo Raita. On guards and symbol dependencies in substring search. Software — Practice and Experience, 29 (11):931–941, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
63501200;   http://www3.interscience.wiley.com/cgi-
bin/fulltext?ID=63501200&PLACEBO=IE.pdf.

Ramesh:1994:RMR

[Ram94] R. Ramesh. R^{2}-M: a reconfigurable rewrite machine. Parallel Processing Letters, 4(1-2):171–180, June 1994. CODEN PPLTEE. ISSN 0129-6264.

RaoKosaraju:1994:CSS

[Rao94] S. Rao Kosaraju. Computation of squares in a string. Lecture Notes in Computer Science, 807:146–150, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

RaoKosaraju:1995:PMC

[Rao95] S. Rao Kosaraju. Pattern matching in compressed texts. Lecture Notes in Computer Science, 1026:349–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raymond:1996:RRE

[Ray96] P. Raymond. Recognizing regular expressions by means of dataflow networks. Lecture Notes in Computer Science, 1099: 336–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robbins:2005:CSS

[RB05] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Sebastopol, CA, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005. URL

http://www.oreilly.com/catalog/shellsrptg/.

Ribeiro:2017:CBC

[RD17] Rodrigo Ribeiro and André Du Bois. Certified bit-coded regular expression parsing. In Fabio Mascarenhas, editor, Proceedings of the 21st Brazilian Symposium on Programming Languages: Fortaleza CE, Brazil, September 21–22, 2017, pages 1–8. ACM Press, New York, NY 10036, USA, September 2017. ISBN 1-4503-5389-4.

Renders:2024:LAP

[RDRF24] Luca Renders, Lore Depuydt, Sven Rahmann, and Jan Fostier. Lossless approximate pattern matching: Automated design of efficient search schemes. Journal of Computational Biology, 31(10):975–989, October 2024. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https:/

/www.liebertpub.com/doi/abs/10.1089/cmb.2024.0664;
 https://www.liebertpub.com/doi/reader/10.1089/cmb.
2024.0664.

Regnier:1992:LAS

[Reg92] Mireille Regnier. A language approach to string searching evaluation. Lecture Notes in Computer Science, 644:15–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiser:1977:EDO

[Rei77] John F. Reiser. EDTV — a display-oriented text editor for UNIX. Technical Memorandum 1177 (TM 77-1353-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1977. ?? pp.

Reid:2003:BNS

[Rei03] Paul Reid. Biometrics and Network security. Prentice Hall series in computer networking and distributed systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-101549-4. 320 pp. LCCN TK7882.B56 R45 2004.

Remy:2017:OEP

[Rém17] Didier Rémy. Ornaments: exploiting parametricity for safer, more automated code refactorization and code reuse (invited talk). ACM SIGPLAN Notices, 52(10):1, October 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reps:1998:MMT

[Rep98] Thomas Reps. “Maximal-munch” tokenization in linear time. ACM Transactions on Programming Languages and Systems, 20(2):259–273, March 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/journals/toplas/1998-20-
2/p259-reps/.

Revesz:1991:TOM

[Rev91] G. E. Revesz. On translating ordinary mathematical notation. Structured programming, 12(3):115–122, 1991. CODEN STPGEM. ISSN 0935-1183.

Reznick:1992:URE

[Rez92] Larry Reznick. Using regular expressions. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):59–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Rudwan:2023:HFS

[RFD23] Mohammed Suleiman Mohammed Rudwan, , and Jean Vincent Fonou-Dombeu. Hybridizing fuzzy string matching and machine learning for improved ontology alignment. Future Internet, 15(7):229–??, July 2023. CODEN ???? ISSN 1999-

5903. URL https://www.mdpi.com/1999-5903/15/7/229.

Reid:1981:ABB

[RH81] Brian K. Reid and David Hanson. An annotated bibliography of background material on text manipulation. ACM SIG-PLAN Notices, 16(6):157–160, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reza:2021:SPM

[RHR^{+}21] Tahsin Reza, Hassan Halawa, Matei Ripeanu, Geoffrey Sanders, and Roger A. Pearce. Scalable pattern matching in metadata graphs via constraint checking. ACM Transactions on Parallel Computing (TOPC), 8(1):2:1–2:45, April 2021. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3434391.

Richards:1979:CFR

[Ric79] Martin Richards. A compact function for regular expression pattern matching. Software — Practice and Experience, 9 (7):527–534, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ristov:2016:FSP

[Ris16] Strahil Ristov. A fast and simple pattern matching with Hamming distance on large alphabets. Journal of Computational Biology, 23(11):874–876, November 2016. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https:/

/www.liebertpub.com/doi/abs/10.1089/cmb.2016.0020;
 https://www.liebertpub.com/doi/pdf/10.1089/cmb.2016.
0020.

Ritchie:19xx:IHQ

[Ritxx] Dennis Ritchie. An incomplete history of the QED text editor. Report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx. URL http://plan9.bell-

labs.com/who/dmr/qed.html.

Reuhkala:1979:RHA

[RJK79] Erkki Reuhkala, Matti Jalanko, and Teuvo Kohonen. Redundant hash addressing method adapted for the postprocessing and error-correction of computer-recognized speech. Record — IEEE International Conference on Acoustics, Speech & Signal Processing, ??(??):591–594, ???? 1979. CODEN RIIPDR.

Robles-Kelly:2002:SED

[RKH02] Antonio Robles-Kelly and Edwin R. Hancock. String edit distance, random walks and graph matching. Lecture Notes in Computer Science, 2396:104–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
2396/23960104.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2396/23960104.pdf.

Regeciova:2021:PMY

[RKM21] Dominika Regeciova, Dusan Kolar, and Marek Milkovic. Pattern matching in YARA: Improved Aho–Corasick algorithm. IEEE Access, 9:62857–62866, 2021. ISSN 2169-3536.

Ryu:2020:FSM

[RLP20] Cheol Ryu, Thierry Lecroq, and Kunsoo Park. Fast string matching for DNA sequences. Theoretical Computer Science, 812(??):137–148, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397519305821.

Ratcliff:1988:PMG

[RM88] John W. Ratcliff and David E. Metzener. Pattern matching: The gestalt approach. Dr. Dobb’s Journal of Software Tools, 13(7):46, 47, 59–51, 68–72, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Rao:2006:SXD

[RM06] Praveen Rao and Bongki Moon. Sequencing XML data and query twigs for fast pattern matching. ACM Transactions on Database Systems, 31(1):299–345, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Rostami:2014:RRE

[RMK^{+}14] Masoud Rostami, Mehrdad Majzoobi, Farinaz Koushanfar, Dan S. Wallach, and Srinivas Devadas. Robust and reverseengineering resilient PUF authentication and key-exchange by substring matching. IEEE Transactions on Emerging Topics in Computing, 2(1):37–49, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ropelewski:1997:IGS

[RND97] Alexander J. Ropelewski, Hugh B. Nicholas, Jr., and David W. Deerfield, II. Implementation of genetic sequence alignment programs on supercomputers. The Journal of Supercomputing, 11(3):237–253, November 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=11&issue=3&spage=237;
 http://www.wkap.nl/oasis.htm/144236.

Russo:2009:ASM

[RNOM09] Luís M. S. Russo, Gonzalo Navarro, Arlindo L. Oliveira, and Pedro Morales. Approximate string matching with compressed indexes. Algorithms (Basel), 2(3):1105–1136, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/2/3/1105.

Robson:1979:ECP

[Rob79] J. M. Robson. The emptiness of complement problem for semi extended regular expressions requires c^{n} space. Information Processing Letters, 9(5):220–222, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robison:1987:IFP

[Rob87] A. D. Robison. The Illinois functional programming interpreter. ACM SIGPLAN Notices, 22(7):64–73, July 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print),

1558-1160 (electronic). URL http://www.acm.org:80/pubs/
citations/proceedings/plan/29650/p64-robison/.

Robertson:1988:TTW

[Rob88] Edmund F. Robertson. Tietze transformations with weighted substring search. Journal of Symbolic Computation, 6(1):59– 64, August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Robinson:1992:HSR

[Rob92] Ian N. Robinson. Hardware to support runtime intelligence. Computer, 25(5):63–66, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robbins:1999:UND

[Rob99a] Arnold Robbins. UNIX in a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http://

www.oreilly.com/catalog/unixnut3.

Robbins:1999:UNS

[Rob99b] Arnold Robbins. UNIX in a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http://

www.oreilly.com/catalog/unixnut3.

Roellke:2025:SMS

[Roe25] Dennis Roellke. String matching at scale: a call for interdisciplinary collaboration and better-directed resources. ACM Queue: Tomorrow’s Computing Today, 23(1):??, January 2025. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Romero:2014:MPR

[Rom14] Victor Romero. Mastering Python regular expressions. Shroff Publishers, ????, 2014. ISBN 93-5110-550-4. LCCN ????

Rooijackers:1999:TCG

[Roo99] Jan Rooijackers. Take command: grep: Searching for words. Linux Journal, 60:??, April 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Ross:1995:FSS

[Ros95] John W. Ross. Fast string searching. C/C++ Users Journal, 13(7):63–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Rote:1991:CMH

[Rot91] Günter Rote. Computing the minimum Hausdorff distance between two-point sets on a line under translation. Information Processing Letters, 38(3):123–127, May 17, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rousseau:2021:LCS

[Rou21] Jérôme Rousseau. Longest common substring for random subshifts of finite type. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1768–1785, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Longest-common-substring-
for-random-subshifts-of-finite-type/10.1214/20-AIHP1130.
full.

Rauchwerger:1995:LTS

[RP95] Lawrence Rauchwerger and David Padua. The LRPD test: speculative run-time parallelization of loops with privatization and reduction parallelization. ACM SIGPLAN Notices, 30(6):218–232, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p218-rauchwerger/.

Rodeh:1981:LAD

[RPE81] Michael Rodeh, Vaughan R. Pratt, and Shimon Even. Linear algorithm for data compression via string matching. Journal of the Association for Computing Machinery, 28(1):16–24, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramesh:1990:PTP

[RR90] R. Ramesh and I. V. Ramakrishnan. Parallel tree pattern matching. Journal of Symbolic Computation, 9(4):485–502 (or 485–501??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ramesh:1992:NPM

[RR92] R. Ramesh and I. V. Ramakrishnan. Nonlinear pattern matching in trees. Journal of the Association for Computing Machinery, 39(2):295–316, April 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/128752.html.

Rabin:1959:FAT

[RS59] M. O. Rabin and D. Scott. Finite automata and their decision problems. IBM Journal of Research and Development, 3 (2):114–125, April 1959. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5392601. This

paper shows the equivalence of deterministic and nondeterministic finite automata.

Regnier:1998:CSP

[RS98] Mireille Régnier and Wojciech Szpankowski. Complexity of sequential pattern matching algorithms. Lecture Notes in Computer Science, 1518:187–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1518/
15180187.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/1518/15180187.pdf.

Roy:2019:EHP

[RSG^{+}19] I. Roy, A. Srivastava, M. Grimm, M. Nourian, M. Becchi, and S. Aluru. Evaluating high performance pattern matching on the automata processor. IEEE Transactions on Computers, 68(8):1201–1212, August 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rottenstreich:2017:ORC

[RT17] Ori Rottenstreich and Janos Tapolcai. Optimal rule caching and lossy compression for longest prefix matching. IEEE/ ACM Transactions on Networking, 25(2):864–878, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Radanne:2018:RLG

[RT18] Gabriel Radanne and Peter Thiemann. Regenerate: a language generator for extended regular expressions. ACM SIG-PLAN Notices, 53(9):202–214, November 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3393934.3278133.

Rendel:2015:ARL

[RTO15] Tillmann Rendel, Julia Trieflinger, and Klaus Ostermann. Automatic refunctionalization to a language with copattern matching: with applications to the expression problem. ACM SIGPLAN Notices, 50(9):269–279, September 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rautio:2002:SMSa

[RTT02a] J. Rautio, J. Tanninen, and J. Tarhio. String matching with stopper compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 2002.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1000012. IEEE Computer Society Order Number: PR01477.

Rautio:2002:SMSb

[RTT02b] Jussi Rautio, Jani Tanninen, and Jorma Tarhio. String matching with stopper encoding and code splitting. Lecture Notes in Computer Science, 2373:42–??, 2002. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2373/
23730042.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2373/23730042.pdf.

Ruckert:2015:MSS

[Ruc15] Martin Ruckert. The MMIX supplement: supplement to The Art of Computer Programming, volumes 1, 2, 3 by Donald E. Knuth. Addison-Wesley, Reading, MA, USA, 2015. ISBN 0-13-399231-4 (paperback), 0-13-399289-6. xxi + 193 pp. LCCN QA76.6 .K64 2005 Suppl. 1. URL http://mmix.cs.hm.edu/.

Reingold:1997:KMP

[RUG97] Edward M. Reingold, Kenneth J. Urban, and David Gries. K-M-P string matching revisited. Information Processing Letters, 64(5):217–223, December 23, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Russinoff:1985:EBM

[Rus85] David M. Russinoff. An experiment with the Boyer–Moore theorem prover: A proof of Wilson’s theorem. Journal of Automated Reasoning, 1(2):121–139, June 1985. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF00244993.

Rus:1988:PLP

[Rus88] T. Rus. Parsing languages by pattern matching. IEEE Transactions on Software Engineering, 14(4):498–511, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=4672.

Russinoff:1992:VSC

[Rus92] David M. Russinoff. A verification system for concurrent programs based on the Boyer–Moore prover. Formal Aspects of Computing, 4(1S):597–611, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X

(electronic). URL http://link.springer.com/article/
10.1007/BF03180564.

Riveros:2023:RNR

[RVV23] Cristian Riveros, Nicolás Van Sint Jan, and Domagoj Vrgoc. REmatch: a novel regex engine for finding all matches. Proceedings of the VLDB Endowment, 16(11):2792–2804, July 2023. CODEN ???? ISSN 2150-8097. URL https://dl.

acm.org/doi/10.14778/3611479.3611488.

Rao:1993:ELD

[RW93] Pushpa Rao and Clifford Walinsky. An equational language for data-parallelism. ACM SIGPLAN Notices, 28(7):112–118, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rutter:2010:CLM

[RW10] Ignaz Rutter and Alexander Wolff. Computing large matchings fast. ACM Transactions on Algorithms, 7(1):1:1–1:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rytter:1980:CPA

[Ryt80] Wojciech Rytter. A correct preprocessing algorithm for Boyer–Moore string-searching. SIAM Journal on Computing, 9(3):509–512, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rytter:1989:NOP

[Ryt89] Wojciech Rytter. A note on optimal parallel transformations of regular expressions to nondeterministic finite automata. Information Processing Letters, 31(2):103–109, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stockman:1977:EHC

[SA77] G. C. Stockman and Ashok K. Agrawala. Equivalence of Hough curve detection to template matching. Communications of the Association for Computing Machinery, 20(11): 820–822, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sanfeliu:1996:ERC

[SA96] A. Sanfeliu and R. Alquezar. Efficient recognition of a class of context-sensitive languages described by augmented regular expressions. Lecture Notes in Computer Science, 1121:1–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sabsevitz:1976:REM

[Sab76] A. L. Sabsevitz. A regular expression matcher for the C language. Technical Memorandum 1101 (TM 76-9141-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 2, 1976. ?? pp.

Sadeh:1993:ASM

[Sad93] I. Sadeh. On approximate string matching. In Storer and Cohn [SC93], pages 148–157. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314. LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253135. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Sadeh:1996:UDC

[Sad96] Ilan Sadeh. Universal data compression algorithm based on approximate string matching. Probability in the Engineering and Informational Sciences, 10(4):465–486, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F03252060CD00649B02534E0CF74CD68.

Sakharov:2021:ARE

[Sak21] Alexander Sakharov. Annotated regular expressions and input-driven languages. Information Processing Letters, ?? (??):Article 105958, ???? 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019020300454.

Saliba:2001:BRB

[Sal01] George Saliba. Book review: The Diffusion of Greco-Roman Medicine into the Middle East and the Caucasus by John A. C. Greppin; Emilie Savage-Smith; John L. Gueriguian. Isis, 92(4):763–764, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080358.

Salmela:2012:ACB

[Sal12]  Leena Salmela.  Average complexity of backward q-

gram string matching algorithms. Information Processing Letters, 112(11):433–437, June 15, 2012. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019012000592.

Stylianopoulos:2020:MPM

[SALP20] Charalampos Stylianopoulos, Magnus Almgren, Olaf Landsiedel, and Marina Papatriantafilou. Multiple pattern matching for network security applications: Acceleration through vectorization. Journal of Parallel and Distributed Computing, 137(??):34–52, March 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731519301984.

Sandberg:1995:COE

[San95] Jonathan S. Sandberg. Counting OCR errors in typeset text. Proceedings of SPIE — The International Society for Optical Engineering, 2422:184–195, 1995. CODEN PSISDG. ISBN 0-8194-1769-6. ISSN 0277-786X (print), 1996-756X (electronic).

Sanden:2009:ISD

[San09] Bo Sandén. Inspired software design early Jackson methods to thread architectures. ACM SIGSOFT Software Engineering Notes, 34(4):1–6, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Santini:2015:QSU

[San15] Simone Santini. Querying streams using regular expressions: some semantics, decidability, and efficiency issues. VLDB Journal: Very Large Data Bases, 24(6):801–821, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sarmiento:2002:SAS

[Sar02] Evan Sarmiento. Systems administration with scsh. Sys Admin: The Journal for UNIX Systems Administrators, 11(1): 16, 18–22, January 2002. CODEN SYADE7. ISSN 1061-2688.

Sassa:1979:PMM

[Sas79] Masataka Sassa. A pattern matching macro processor. Software — Practice and Experience, 9(6):439–456, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sarma:2013:ULB

[SASU13] Anish Das Sarma, Foto N. Afrati, Semih Salihoglu, and Jeffrey D. Ullman. Upper and lower bounds on the cost of a map-reduce computation. Proceedings of the VLDB Endowment, 6(4):277–288, February 2013. CODEN ???? ISSN 2150-8097.

Sun:2009:DPP

[SB09] Yanni Sun and Jeremy Buhler. Designing patterns and profiles for faster HMM search. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 6(2):232–243, April 2009. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Samadani:2019:SPM

[SBB19] Mohammad Hasan Samadani, Mehdi Berenjkoob, and Marina Blanton. Secure pattern matching based on bit parallelism. International Journal of Information Security, 18(3): 371–391, June 2019. CODEN ???? ISSN 1615-5262 (print),

1615-5270 (electronic). URL http://link.springer.com/
article/10.1007/s10207-018-0410-8.

Salter:1980:CLC

[SBF80] Richard M. Salter, Terence J. Brennan, and Daniel P. Friedman. Concur: a language for continuous, concurrent processes. Computer Languages, 5(3-4):163–189, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sakakibara:1994:RMR

[SBHM94] Y. Sakakibara, M. Brown, R. Hughey, and I. S. Mian. Recent methods for RNA modeling using stochastic context-free grammars. Lecture Notes in Computer Science, 807:289–306, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Soule:2018:MLM

[SBM^{+}18] Robert Soule, Shrutarshi Basu, Parisa Jalili Marandi, Fernando Pedone, Robert Kleinberg, Emin Gun Sirer, and Nate Foster. Merlin: a language for managing network resources. IEEE/ACM Transactions on Networking, 26(5):2188–2201, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stallmann:2007:PAE

[SBR^{+}07] Matthias F. Stallmann, Suzanne P. Balik, Robert D. Rodman, Sina Bahram, Michael C. Grace, and Susan D. High. ProofChecker: an accessible environment for automata theory correctness proofs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):48–52, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Smith:1988:ILL

[SC88] W. W. Smith and R. H. Campbell. Introduction to leif language descriptions. Technical Report UIUCDCS-R-88-1444, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, July 1988. 78 pp.

Storer:1993:DDC

[SC93] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’93: Data Compression Conference: [March 30–April 2, 1993, Snowbird, Utah: proceedings]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314. LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=452. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Storer:1995:DDC

[SC95] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’95, Data Compression Conference: March 28– 30, 1995, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7012-6. ISSN 1068-0314. LCCN

???? URL http://ieeexplore.ieee.org/servlet/opac?

punumber=3874. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Storer:1996:DDC

[SC96] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’96: Data Compression Conference, March 31–April 3, 1996, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33

D37 1996. URL http://ieeexplore.ieee.org/servlet/

opac?punumber=3509. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Storer:1998:PDD

[SC98] James A. (James Andrew) Storer and Martin Cohn, editors. Proceedings: DCC ’98: Data Compression Conference: March 30–April 1, 1998, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672118. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Storer:1999:DPD

[SC99] James A. Storer and Martin Cohn, editors. DCC ’99: proceedings: Data Compression Conference: March 29–31, 1999, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755647. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Storer:2001:DPD

[SC01] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2001: proceedings, Data Compression Conference, March 27–29, 2001, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1031-0. ISSN 1068-0314. LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917130. IEEE

CSP number 01PR1031.

Storer:2002:DPD

[SC02] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2002: proceedings: Data Compression Conference: April 2–4, 2002, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314. LCCN

QA76.9.D33 D37 2002. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=999937. IEEE Computer

Society Order Number: PR01477.

Storer:2003:DPD

[SC03] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2003: proceedings: Data Compression Conference: March 25–27, 2003, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1896-6. ISSN 1068-0314. LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1193990. IEEE

Computer Society Order number PR01896.

Storer:2004:DCC

[SC04] James A. (James Andrew) Storer and Martin Cohn, editors. Data Compression Conference, 2004. Proceedings. DCC 2004: Snowbird, Utah, USA: 23–25 March 2004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2082-0. ISSN 1068-

0314. LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281440. IEEE Computer

Society Order Number: P2082.

Storer:2005:DCC

[SC05] James A. (James Andrew) Storer and Martin Cohn, editors. Data Compression Conference, 2005. Proceedings. DCC 2005. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2309-9. ISSN 1068-0314. LCCN

???? URL http://ieeexplore.ieee.org/servlet/opac?
punumber=9633;   http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402158.

Scarpazza:2011:TPT

[Sca11] Daniele Paolo Scarpazza. Top-performance tokenization and small-ruleset regular expression matching: a quantitative performance analysis and optimization study on the Cell/B.E. Processor. International Journal of Parallel Programming, 39(1):3–32, February 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7458&volume=39&issue=1&spage=3.

Salehi:2017:RSR

[SCF^{+}17] Mohsen Amini Salehi, Thomas Caldwell, Alejandro Fernandez, Emmanuel Mickiewicz, Eric W. D. Rozier, Saman Zonouz, and David Redberg. RESeED: a secure regularexpression search tool for storage clouds. Software — Practice and Experience, 47(9):1221–1241, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sanchez-Couso:1994:ACA

[SCFC94] J. R. Sanchez-Couso and M. I. Fernandez-Camacho. Averagecase analysis of pattern-matching in trees under the BST probability model. Lecture Notes in Computer Science, 820: 178–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schilit:1981:SGB

[Sch81] Bill N. Schilit. A solution to the great big substitution problem: subject modification during pattern matching in SNOBOL4. ACM SIGPLAN Notices, 16(8):41–49, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schaback:1988:ESB

[Sch88a] R. Schaback. On the expected sublinearity of the Boyer– Moore algorithm. SIAM Journal on Computing, 17(4):648– 658, August 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schnoebelen:1988:RCP

[Sch88b] Ph. Schnoebelen. Refined compilation of pattern-matching for functional languages. Science of Computer Programming, 11 (2):133–159, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schneier:1991:OHF

[Sch91a] Bruce Schneier. One-way hash functions: Probabilistic algorithms can be used for general-purpose pattern matching. Dr. Dobb’s Journal of Software Tools, 16(9):148–151, September 1, 1991. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1991:OWH

[Sch91b] Bruce Schneier. One-way hash functions: Probabilistic algorithms can be used for general-purpose pattern matching. Dr. Dobb’s Journal of Software Tools, 16(9):148–151, September 1, 1991. CODEN DDJOEB. ISSN 1044-789X.

Schulzrinne:1995:DCC

[Sch95] H. Schulzrinne. Dynamic configuration of conferencing applications using pattern-matching multicast. Lecture Notes in Computer Science, 1018:216–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwartz:1999:CRE

[Sch99] Randal Schwartz. Compiling regular expressions. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):39–40, 42, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Schmid:2012:ICR

[Sch12] Markus L. Schmid. Inside the class of REGEX languages. Lecture Notes in Computer Science, 7410:73–84, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31653-1_8/.

Schmid:2013:ICR

[Sch13] Markus L. Schmid. Inside the class of regex languages. International Journal of Foundations of Computer Science (IJFCS), 24(7):1117–??, November 2013. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schildt:2014:JCRb

[Sch14] Herbert Schildt, editor. Java: The Complete Reference. Mc-Graw-Hill, New York, NY, USA, ninth edition, 2014. ISBN 0-07-180855-8 (paperback), 0-07-180925-2, 0-07-180856-6. xxxiv + 1274 pp. LCCN QA76.73.J38 S332 2014eb.

Schmid:2022:CRP

[Sch22] Markus L. Schmid. Conjunctive regular path queries with capture groups. ACM Transactions on Database Systems, 47 (2):5:1–5:52, June 2022. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL https://dl.acm.org/

doi/10.1145/3514230.

Savoy:1991:IRH

[SD91a] Jacques Savoy and Daniel Desbois. Information retrieval in hypertext systems: an approach using Bayesian networks. Electronic Publishing—Origination, Dissemination, and Design, 4(2):87–108, June 1991. CODEN EPODEU. ISSN 0894-3982.

Savoy:EPODD-4-2-87

[SD91b] Jacques Savoy and Daniel Desbois. Information retrieval in hypertext systems: an approach using Bayesian networks. Electronic Publishing—Origination, Dissemination, and Design, 4(2):87–108, June 1991. CODEN EPODEU. ISSN 0894-3982.

Sewell:1995:MCP

[SD95] Roger F. Sewell and Richard Durbin. Method for calculation of probability of matching a bounded regular expression in a random data string. Journal of Computational Biology, 2(1):25–31, January 1995. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https:/

/www.liebertpub.com/doi/abs/10.1089/cmb.1995.2.25;
 https://www.liebertpub.com/doi/pdf/10.1089/cmb.1995.
2.25.

Subramanian:2017:GSF

[SDA17] Kausik Subramanian, Loris D’Antoni, and Aditya Akella. Genesis: synthesizing forwarding tables in multi-tenant networks. ACM SIGPLAN Notices, 52(1):572–585, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sittampalam:2001:HOP

[SdM01] Ganesh Sittampalam and Oege de Moor. Higher-order pattern matching for automatically applying fusion transformations. Lecture Notes in Computer Science, 2053:218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2053/20530218.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2053/20530218.pdf.

Solodkyy:2014:OPM

[SDS14] Yuriy Solodkyy, Gabriel Dos Reis, and Bjarne Stroustrup. Open pattern matching for C++. ACM SIGPLAN Notices, 49(3):33–42, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sedgewick:1983:A

[Sed83] Robert Sedgewick. Algorithms. Addison-Wesley, Reading, MA, USA, 1983. ISBN 0-201-06672-6. viii + 551 pp. LCCN QA76.6 .S435 1983.

Sedgewick:1990:AC

[Sed90] Robert Sedgewick. Algorithms in C. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51425-7. xii + 657 pp. LCCN QA76.73.C15 S43 1990.

Sedgewick:1992:AC

[Sed92] Robert Sedgewick. Algorithms in C++. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-36118-3, 0-201-51059-6. xiv + 656 pp. LCCN QA76.73.C153 S38 1992.

Stolee:2014:SSS

[SED14] Kathryn T. Stolee, Sebastian Elbaum, and Daniel Dobos. Solving the search for source code. ACM Transactions on Software Engineering and Methodology, 23(3):26:1–26:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sellers:1984:PRG

[Sel84] Peter H. Sellers. Pattern recognition in genetic sequences by mismatch density. Bulletin of Mathematical Biology, 46 (4):501–514, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459499.

Senellart:2000:FPM

[Sen00] Jean Senellart. Fast pattern matching in indexed texts. Theoretical Computer Science, 237(1–2):239–262, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.nl/gej-ng/10/41/16/
171/21/33/abstract.html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/33/article.pdf.

Sestoft:1996:MPM

[Ses96] P. Sestoft. ML pattern match compilation and partial evaluation. Lecture Notes in Computer Science, 1110:446–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singh:2001:PMN

[SF01] Sameer Singh and Jonathan Fieldsend. Pattern matching and neural networks based hybrid forecasting system. Lecture Notes in Computer Science, 2013:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2013/20130072.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2013/20130072.pdf.

Singh:2012:LSS

[SG12] Rishabh Singh and Sumit Gulwani. Learning semantic string transformations from examples. Proceedings of the VLDB Endowment, 5(8):740–751, April 2012. CODEN ???? ISSN 2150-8097.

Singh:2016:TSD

[SG16] Rishabh Singh and Sumit Gulwani. Transforming spreadsheet data types using examples. ACM SIGPLAN Notices, 51(1): 343–356, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2014:EPM

[SGCW14] Chunyao Song, Tingjian Ge, Cindy Chen, and Jie Wang. Event pattern matching over graph streams. Proceedings of the VLDB Endowment, 8(4):413–424, December 2014. CO-DEN ???? ISSN 2150-8097.

Shankar:2000:NAL

[SGYM00] Priti Shankar, Amitranjan Gantait, A. R. Yuvaraj, and Maya Madhavan. A new algorithm for linear regular tree pattern matching. Theoretical Computer Science, 242(1–2):125–142, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.elsevier.nl/gej-
ng/10/41/16/176/21/28/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/176/21/28/article.pdf.

Stearns:1985:ECP

[SH85] R. E. Stearns and H. B. Hunt, III. On the equivalence and containment problems for unambiguous regular expressions, regular grammars and finite automata. SIAM Journal on Computing, 14(3):598–611, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shankar:1988:MPC

[Sha88a] N. Shankar. A mechanical proof of the Church–Rosser theorem. Journal of the Association for Computing Machinery, 35 (3):475–522, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/44484.html.

Sharpe:1988:ARE

[Sha88b] Daniel Sharpe. Adding regular expressions to Pascal. ACM SIGPLAN Notices, 23(12):125–133, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shapiro:1993:CCR

[Sha93] V. Shapiro. Cross-correlation with reconstruction: a new approach to pattern matching. Lecture Notes in Computer Science, 719:548–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Soo:1993:DCP

[SHCY93] Von-Wun Soo, Jan-Fu Hwang, Tung-Bo Chen, and Chin Yu. Divide-and-conquer, pattern matching, and relaxation methods in interpretation of 2-D NMR spectra of polypeptides. Journal of Computational Chemistry, 14(10):1164–1171, October 1993. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shepherdson:1959:RTW

[She59] J. C. Shepherdson. The reduction of two-way automata to one-way automata. IBM Journal of Research and Development, 3(2):198–200, April 1959. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5392614.

Shields:1992:SME

[Shi92] Paul C. Shields. String matching: The ergodic case. Annals of Probability, 20(3):1199–1203, July 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176989686.

Shi:1996:SAM

[Shi96] F. Shi. Suffix arrays for multiple strings: a method for on-line multiple string searches. Lecture Notes in Computer Science, 1179:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shields:1997:SMB

[Shi97] Paul C. Shields. String matching bounds via coding. Annals of Probability, 25(1):329–336, January 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1024404290.

Shibuya:2000:GST

[Shi00] Tetsuo Shibuya. Generalization of a suffix tree for RNA structural pattern matching. Lecture Notes in Computer Science, 1851:393–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1851/
18510393.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/1851/18510393.pdf.

Shibuya:2004:GST

[Shi04] Tetsuo Shibuya. Generalization of a suffix tree for RNA structural pattern matching. Algorithmica, 39(1):1–19, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=39&
issue=1&spage=1.

Schoepe:2014:STI

[SHS14] Daniel Schoepe, Daniel Hedin, and Andrei Sabelfeld. SeLINQ: tracking information across application-database boundaries. ACM SIGPLAN Notices, 49(9):25–38, September 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sevenich:2016:UDS

[SHvR^{+}16] Martin Sevenich, Sungpack Hong, Oskar van Rest, Zhe Wu, Jayanta Banerjee, and Hassan Chafi. Using domain-specific languages for analytic graph databases. Proceedings of the VLDB Endowment, 9(13):1257–1268, September 2016. CO-DEN ???? ISSN 2150-8097.

Sidi:1995:CAG

[Sid95] Avram Sidi. Convergence analysis for a generalized Richardson extrapolation process with an application to the d^{(1)}- transformation on convergent and divergent logarithmic sequences. Mathematics of Computation, 64(212):1627–1657, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1999:FCS

[Sid99] Avram Sidi. Further convergence and stability results for the generalized Richardson extrapolation process GREP ^{(1)} with an application to the D ^{(1)}-transformation for infinite integrals. Journal of Computational and Applied Mathematics, 112(1–2):269–290, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042799902261.

Sidi:2000:GRE

[Sid00] Avram Sidi. The generalized Richardson extrapolation process GREP^{(1)} and computation of derivatives of limits of sequences with applications to the d^{(1)}-transformation. Journal of Computational and Applied Mathematics, 122(1–2): 251–273, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042700003629.

Sidi:2002:NCR

[Sid02] Avram Sidi. New convergence results on the generalized Richardson extrapolation process GREP^{(1)} for logarithmic sequences. Mathematics of Computation, 71 (240):1569–1596, October 2002. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journal-getitem?pii=S0025-5718-
01-01384-9;   http://www.ams.org/mcom/2002-71-240/
S0025-5718-01-01384-9/S0025-5718-01-01384-9.dvi;

http://www.ams.org/mcom/2002-71-240/S0025-5718-01-
01384-9/S0025-5718-01-01384-9.pdf;   http://www.ams.
org/mcom/2002-71-240/S0025-5718-01-01384-9/S0025-5718-
01-01384-9.ps;   http://www.ams.org/mcom/2002-71-240/
S0025-5718-01-01384-9/S0025-5718-01-01384-9.tex.

Silverston:1977:NPM

[Sil77] Stefan M. Silverston. A note on pattern matching under Quickscan in SNOBOL4. ACM SIGPLAN Notices, 12(10): 70–74, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simon:1983:PMT

[Sim83] Hans-Ulrich U. Simon. Pattern matching in trees and nets. Acta Informatica, 20(3):227–248, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Simon:1994:SMA

[Sim94] I. Simon. String matching algorithms and automata. Lecture Notes in Computer Science, 812:386–395, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sadoghi:2013:AOB

[SJ13] Mohammad Sadoghi and Hans-Arno Jacobsen. Analysis and optimization for Boolean expression indexing. ACM Transactions on Database Systems, 38(2):8:1–8:??, June 2013. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Sudo:2019:SWM

[SJNS19] Hiroki Sudo, Masanobu Jimbo, Koji Nuida, and Kana Shimizu. Secure wavelet matrix: Alphabet-friendly privacypreserving string search for bioinformatics. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(5):1675–1684, September 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Spyropoulos:2017:DBD

[SK17] Vasilis Spyropoulos and Yannis Kotidis Kotidis. Digree: Building a distributed graph processing engine out of singlenode graph database installations. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):22–27, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shibata:2000:SPM

[SKF^{+}00] Yusuke Shibata, Takuya Kida, Shuichi Fukamachi, Masayuki Takeda, Ayumi Shinohara, Takeshi Shinohara, and Setsuo Arikawa. Speeding up pattern matching by text compression. Lecture Notes in Computer Science, 1767:306–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1767/17670306.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1767/17670306.pdf.

Skiena:1998:ADM

[Ski98] Steven S. Skiena. The Algorithm Design Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94860-0. xvi + 486 pp. LCCN QA76.9.A43S55 1997. US$54.95.

Seni:1996:GED

[SKS96] Giovanni Seni, V. Kripasundar, and Rohini K. Srihari. Generalizing edit distance to incorporate domain information: handwritten text recognition as a case study. Pattern Recognition, 29(3):405–414, March 1996. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Sulzmann:2014:PRE

[SL14] Martin Sulzmann and Kenny Zhuo Ming Lu. POSIX regular expression parsing with derivatives. In Michael Codish and Eijiro Sumii, editors, Functional and Logic Programming: 12th International Symposium, FLOPS 2014, Kanazawa, Japan, June 4–6, 2014. Proceedings, volume 8475 of Lecture Notes in Computer Science, pages 203–220. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014.

Sulzmann:2017:DBD

[SL17] Martin Sulzmann and Kenny Zhuo Ming Lu. Derivative-based diagnosis of regular expression ambiguity. International Journal of Foundations of Computer Science (IJFCS), 28(5):543– ??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Slissenko:1978:SMR

[Sli78] A. Slissenko. String-matching in real-time: Some properties of the data structure. In Winkowski [Win78], pages 493– 496. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1978.

Slissenko:1983:DPS

[Sli83] A. Slissenko. Detection of periodicities and string-matching in real time. Journal of Soviet Mathematics, 22(3):1316–1386, 1983. CODEN JSOMAR. ISSN 0090-4104 (print), 2376-5798 (electronic). Russian original in: Zapiski Nauchnykh Seminarov LOMI, 105:62–173, 1981.

Solar-Lezama:2006:CSF

[SLTB^{+}06] Armando Solar-Lezama, Liviu Tancau, Rastislav Bodik, Sanjit Seshia, and Vijay Saraswat. Combinatorial sketching for finite programs. ACM SIGPLAN Notices, 41(11):404–415, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sun:2020:ERE

[SLZ^{+}20] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Zheng Peng, and Chengchen Hu. Efficient regular expression matching over compressed traffic. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):Article 106996, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128618311939.

Shannon:1956:AS

[SM56] C. E. Shannon and J. McCarthy, editors. Automata Studies, volume AM-34 of Annals of Mathematics Studies. Princeton University Press, Princeton, NJ, USA, 1956. CODEN AN-MAAH. ISBN 0-691-07916-1. ISSN 0066-2313. ix + 285 pp. German translation in [SM74].

Shannon:1974:STA

[SM74] C. E. Shannon and J. McCarthy, editors. Studien zur Theorie der Automaten. (German) Automata Studies. Rogner & Bernhard, Munich, West Germany, 1974. ISBN 3-8077-0001-3, 0-398-03003-0. xxxiii + 452 pp. LCCN ???? Edited by Franz Kaltenbeck and Peter Weibel with notes by Dieter Roth.

Sperberg-McQueen:1999:SRE

[SM99] C. M. Sperberg-McQueen. Squib: Regular expression for dates. Markup languages: theory & practice, 1(4):20–26, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Stubblebine:2004:SHD

[SM04] Tony Stubblebine and Junko Mishima. Seiki hyogen desukutoppu rifarensu: regular expressions for Perl, C, PHP, Python, Java, and .NET. Orairi Japan, Tokyo, Japan, 2004. ISBN 4-87311-170-6. vi + 96 pp. LCCN ????

Storer:2009:DPD

[SM09] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2009: Proceedings Data Compression Conference: Snowbird, Utah, USA, 16–28 March 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3592-5. ISSN 1068-0314. LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976436. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Storer:2010:DPD

[SM10] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DDC 2010: proceedings: Data Compression Conference: 24–26 March 2010, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-3994-7. ISSN 1068-0314.

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453521.

Storer:2011:DDC

[SM11] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2011: Data Compression Conference: Snowbird, Utah, USA, 29–31 March 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-61284-279-8. ISSN 1068-0314. LCCN

???? URL http://ieeexplore.ieee.org/servlet/opac?

punumber=5749456. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Sahinalp:2004:CPM

[SMD04] Suleyman Cenk Sahinalp, S. Muthukrishnan, and Ugur Dogrusoz, editors. Combinatorial Pattern Matching: 15th annual symposium, CPM 2004, Istanbul, Turkey, July 5– 7, 2004: Proceedings, volume 3109 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22341-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 C65

2004. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t3109.htm;   http://www.springerlink.
com/content/978-3-540-22341-2;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
3109;   http://www.springerlink.com/openurl.asp?genre=
volume&id=doi:10.1007/b98377.

Siromoney:1994:ILW

[SMDS94] Rani Siromoney, Lisa Mathew, V. R. Dare, and K. G. Subramanian. Infinite Lyndon words. Information Processing Letters, 50(2):101–104, April 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Smith:1991:PEP

[Smi91a] Donald A. Smith. Partial evaluation of pattern matching in constraint logic programming languages. ACM SIGPLAN Notices, 26(9):62–71, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1991:EVF

[Smi91b] P. D. Smith. Experiments with a very fast substring search algorithm. Software — Practice and Experience, 21(10):1065– 1074, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1994:TBM

[Smi94] P. D. Smith. On tuning the Boyer–Moore–Horspool string searching algorithm. Software — Practice and Experience, 24 (4):435–436, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shahbaz:2015:AGV

[SMS15] Muzammil Shahbaz, Phil McMinn, and Mark Stevenson. Automatic generation of valid and invalid test data for string validation routines using web searches and regular expressions. Science of Computer Programming, 97 (part 4)(??):405–425, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print),

1872-7964 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167642314001725.

Suzuki:1986:SVD

[SMT^{+}86] I. Suzuki, Y. Motohashi, K. Taniguchi, T. Kasami, and T. Okamoto. Specification and verification of decentralized daisy chain arbiters with omega-extended regular expressions. Theoretical Computer Science, 43(2-3):277–291, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sun:2023:ERE

[SMW^{+}23] Xiuwen Sun, Da Mo, Di Wu, Chunhui Ye, Qingying Yu, Jie Cui, and Hong Zhong. Efficient regular expression matching over hybrid dictionary-based compressed data. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804523000541.

Saoudi:1992:OPA

[SN92] A. Saoudi and M. Nivat. Optimal parallel algorithms for multidimensional image template matching and pattern matching. Lecture Notes in Computer Science, 654:240–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheng:1994:PMB

[SN94] Ke-Ning Sheng and Joseph I. Naus. Pattern matching between two non-aligned random sequences. Bulletin of Mathematical Biology, 56(6):1143–1162, November 1994. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02460290.

Susik:2024:PMA

[SN24] Robert Susik and Robert Nowotniak. Pattern matching algorithms in blockchain for network fees reduction. The Journal of Supercomputing, 80(12):17741–17759, August 2024. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-024-06115-8.

Shajii:2019:SHP

[SNB^{+}19] Ariya Shajii, Ibrahim Numanagić, Riyadh Baghdadi, Bonnie Berger, and Saman Amarasinghe. Seq: a high-performance language for bioinformatics. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):125:1–125:29,

October 2019. URL https://dl.acm.org/doi/abs/10.
1145/3360551.

Syme:2007:EPM

[SNM07] Don Syme, Gregory Neverov, and James Margetson. Extensible pattern matching via a lightweight language extension. ACM SIGPLAN Notices, 42(9):29–40, September 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:2013:PSM

[SNM^{+}13] Thorsten-Walther Schmidt, Jan Novák, Johannes Meng, Anton S. Kaplanyan, Tim Reiner, Derek Nowrouzezahrai, and Carsten Dachsbacher. Path-space manipulation of physicallybased light transport. ACM Transactions on Graphics, 32(4): 129:1–129:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Snow:2001:IAL

[Sno01] Ray Snow. An information assembly line in Perl. Dr. Dobb’s Journal of Software Tools, 26(6):72, 76, 78, 80, 82, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.

com/ftp/2001/2001_06/infoline.txt.

SilvadeMoura:2000:FFW

[SNZBY00] Edleno Silva de Moura, Gonzalo Navarro, Nivio Ziviani, and Ricardo Baeza-Yates. Fast and flexible word searching on compressed text. ACM Transactions on Information Systems, 18(2):113–139, April 2000. CODEN ATISET. ISSN

1046-8188. URL http://www.acm.org/pubs/citations/
journals/tois/2000-18-2/p113-silva_de_moura/.

Sommerville:1982:PMS

[Som82] Ian Sommerville. A pattern matching system. Software — Practice and Experience, 12(6):517–530, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Strizhov:2016:SPS

[SOR16] Mikhail Strizhov, Zachary Osman, and Indrajit Ray. Substring position search over encrypted cloud data supporting efficient multi-user setup. Future Internet, 8(3):28, July 04, 2016. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/8/3/28.

Soufi:1999:TSR

[Sou99]  L. Soufi.  Type specification by regular expressions.

J.UCS: Journal of Universal Computer Science, 5(9):622–631, September 28, 1999. CODEN ???? ISSN 0948-695X (print),

0948-6968 (electronic). URL http://www.jucs.org/jucs_5_
9/type_specification_by_regular.

Sapirstein:2016:PMA

[SP16] Philip Sapirstein and Eric Psota. Pattern matching and the analysis of damaged ancient objects: The case of the column drum. Journal on Computing and Cultural Heritage (JOCCH), 9(3):13:1–13:??, November 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Spencer:1985:REP

[Spe85a] Henry Spencer. Regular expression pattern matching software. Usenet mod.sources and net.sources archives, November 1985.

Spencer:regexp

[Spe85b] Henry Spencer. Regular expression pattern matching software. Usenet mod.sources and net.sources archives, November 1985.

Schwartz:2008:LP

[SPF08] Randal L. Schwartz, Tom Phoenix, and Brian D. Foy. Learning Perl. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, fifth edition, 2008. ISBN 0-596-52010-7. xviii + 328 pp. LCCN QA76.73.P22 S37 2008.

Spinellis:1999:DPO

[Spi99a] Diomidis Spinellis. Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34(2): 47–50, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1999:TCD

[Spi99b] Diomidis Spinellis. Technical correspondence: Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34(2):47–51, February 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.informatik.

rwth-aachen.de/dblp/db/indices/a-tree/s/Spinellis:
Diomidis.html.

Sitaridi:2016:GAS

[SR16] Evangelia A. Sitaridi and Kenneth A. Ross. GPU-accelerated string matching for database applications. VLDB Journal: Very Large Data Bases, 25(5):719–740, October 2016. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sridhar:1988:CBG

[Sri88] M. A. Sridhar. On the connectivity of the De Bruijn graph. Information Processing Letters, 27(6):315–318, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srinivas:1993:STA

[Sri93] Yellamraju V. Srinivas. A sheaf-theoretic approach to pattern matching and related problems. Theoretical Computer Science, 112(1):53–97, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1993&volume=112&issue=1&
aid=1301.

Saha:2020:EPP

[SRK20] Tushar Kanti Saha, Deevashwer Rathee, and Takeshi Koshiba. Efficient protocols for private wildcards pattern matching. Journal of Information Security and Applications (JISA), 55(??):??, December 2020. CODEN ???? ISSN

2214-2126. URL http://www.sciencedirect.com/science/
article/pii/S2214212620307742.

Sekar:1992:APM

[SRR92] R. C. Sekar, R. Ramesh, and I. V. Ramakrishnan. Adaptive pattern matching. Lecture Notes in Computer Science, 623: 247–??, 1992. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sekar:1995:APM

[SRR95] R. C. Sekar, R. Ramesh, and I. V. Ramakrishnan. Adaptive pattern matching. SIAM Journal on Computing, 24(6):1207– 1234, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/24625.

Sadredini:2019:SEM

[SRV^{+}19] E. Sadredini, R. Rahimi, V. Verma, M. Stan, and K. Skadron. A scalable and efficient in-memory interconnect architecture for automata processing. IEEE Computer Architecture Letters, 18(2):87–90, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Shevchenko:1993:SRP

[SS93a] Ivan I. Shevchenko and Andrej G. Sokolsky. Studies of regular precessions of a symmetric satellite by means of computer algebra. In Bronstein [Bro93], pages 65–67. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/

164081/p65-shevchenko/. ACM order number: 505930.

Smith:1993:XRL

[SS93b] Henry I. Smith and M. L. Schattenburg. X-ray lithography, from 500 to 30 nm: X-ray nanolithography. IBM Journal of Research and Development, 37(3):319–329, May 1993. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Senoussi:1994:QAT

[SS94] H. Senoussi and A. Saoudi. A quadtree algorithm for template matching on a pyramid computer. Theoretical Computer Science, 136(2):387–417, December 29, 1994.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1661.

Shieh:2021:EMP

[SSLL21] Yi-Kung Shieh, Shyong Jian Shyu, Chin Lung Lu, and Richard Chia-Tung Lee. The exact multiple pattern matching problem solved by a reference tree approach. Theoretical Computer Science, 882(??):29–48, August 23, 2021. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397521003480.

Silvasti:2010:ELX

[SSSS10] P. Silvasti, S. Sippu, and E. Soisalon-Soininen. Evaluating linear XPath expressions by pattern-matching automata. J.UCS: Journal of Universal Computer Science, 16(5):833–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-

6968 (electronic). URL http://www.jucs.org/jucs_16_5/
evaluating_linear_xpath_expressions.

Sun:2019:IMD

[SSYW19] Ruxia Sun, Lingfeng Shi, Chunyong Yin, and Jin Wang. An improved method in deep packet inspection based on regular expression. The Journal of Supercomputing, 75(6):3317–3333, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sutinen:1995:UGL

[ST95] E. Sutinen and J. Tarhio. On using q-gram locations in approximate string matching. Lecture Notes in Computer Science, 979:327–??, 1995. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shawe-Taylor:1996:FSM

[ST96a] John Shawe-Taylor. Fast string matching in stationary ergodic sources. Combinatorics, Probability and Computing, 5(4):415– 427, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sutinen:1996:FQS

[ST96b] E. Sutinen and J. Tarhio. Filtration with q-samples in approximate string matching. Lecture Notes in Computer Science, 1075:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutinen:1996:FSA

[ST96c] E. Sutinen and J. Tarhio. Filtration with q-samples in approximate string matching. Lecture Notes in Computer Science, 1075:50–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sklar:2003:PC

[ST03] David Sklar and Adam Trachtenberg. PHP Cookbook. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2003. ISBN 1-56592-681-1. xxi + 608 pp. LCCN QA76.73.P224 S56 2003. US$39.95. URL http:

//www.oreilly.com/catalog/9781565926813;   http://
www.oreilly.com/catalog/phpckbk.

Sutinen:2004:ASM

[ST04] E. Sutinen and J. Tarhio. Approximate string matching with ordered q-grams. Nordic Journal of Computing, 11(4):321–??, Winter 2004. CODEN NJCOFR. ISSN 1236-6064.

Sulzmann:2019:DPD

[ST19] Martin Sulzmann and Peter Thiemann. Derivatives and partial derivatives for regular shuffle expressions. Journal of Computer and System Sciences, 104(??):323–341, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-

2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022000016301325.

Statman:1989:SSC

[Sta89] R. Statman. On sets of solutions to combinator equations. Theoretical Computer Science, 66(1):99–104, August 2, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stephen:1994:SSA

[Ste94] Graham A. Stephen. String Searching Algorithms, volume 3 of Lecture notes series on computing. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1829-X (hardcover). xii + 243 pp. LCCN QA76.9.A43 S73 1994.

Steele:2012:PLL

[Ste12] Guy L. Steele, Jr. Programming language life cycles. ACM SIGADA Ada Letters, 32(3):95–96, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Salmela:2006:MSM

[STK06] Leena Salmela, Jorma Tarhio, and Jari Kytöjoki. Multipattern string matching with q-grams. ACM Journal of Experimental Algorithmics, 11:1.1:1–1.1:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Salmela:2010:ABM

[STK10] Leena Salmela, Jorma Tarhio, and Petri Kalsi. Approximate Boyer–Moore string matching for small alphabets. Algorithmica, 58(3):591–609, November 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=3&spage=591.

Shetiya:2020:AAS

[STKD20] Suraj Shetiya, Saravanan Thirumuruganathan, Nick Koudas, and Gautam Das. Astrid: accurate selectivity estimation for string predicates using deep learning. Proceedings of the VLDB Endowment, 14(4):471–484, December 2020. CODEN

???? ISSN 2150-8097. URL https://dl.acm.org/doi/10.
14778/3436905.3436907.

Stonebraker:1992:PAS

[Sto92] Michael Stonebraker, editor. Proceedings of the 1992 ACM SIGMOD International Conference on Management of Data, San Diego, California, June 2–5, 1992, volume 21(2) of SIG-MOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ????

Stojmenovic:1996:CTB

[Sto96] Ivan Stojmenovic. Constant time BSR solutions to parenthesis matching, tree decoding, and tree reconstruction from its traversals. IEEE Transactions on Parallel and Distributed Systems, 7(2):218–224, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/tpds/td1996/l0218abs.htm.

Stomp:2002:CSP

[Sto02] Frank Stomp. Correctness of substring-preprocessing in Boyer–Moore’s pattern matching algorithm. Theoretical Computer Science, 290(1):59–78, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stroustrup:2013:CPL

[Str13] Bjarne Stroustrup. The C++ programming language. Addison-Wesley, Reading, MA, USA, fourth edition, 2013. ISBN 0-321-56384-0 (paperback), 0-321-95832-2 (hardcover), 0-13-352283-0 (e-book). xiv + 1346 pp. LCCN QA76.73.C153 S77 2013.

Shibata:1999:PMT

[STSA99] Yusuke Shibata, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Pattern matching in text compressed by using antidictionaries. Lecture Notes in Computer Science, 1645: 37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1645/16450037.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1645/16450037.pdf.

Stubblebine:2003:REP

[Stu03] Tony Stubblebine. Regular expression pocket reference. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, 2003. ISBN 0-596-00415-X. vi + 93 pp. LCCN QA76.9.T48 S78 2003. URL http://www.oreilly.

com/catalog/9780596004156.

Stubblebine:2007:REP

[Stu07] Tony Stubblebine. Regular expression pocket reference: Regular expressions for Perl, Ruby, PHP, Python, C, Java, and .NET. O’Reilly & Associates, Sebastopol, CA, USA, and Cambridge, MA, USA, second edition, 2007. ISBN 0-596-51427-1 (paperback). vii + 117 pp. LCCN QA76.9.T48

S78 2007. URL http://www.loc.gov/catdir/toc/fy0802/
2007281074.html.

Sunday:1990:VFS

[Sun90a] Daniel M. Sunday. A very fast substring search algorithm. Communications of the Association for Computing Machinery, 33(8):132–142, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-0782/79184.html. See

also [BM77, KMP77, BYG92].

Sunday:string-search

[Sun90b] Daniel M. Sunday. A very fast substring search algorithm. Communications of the Association for Computing Machinery, 33(8):132–142, August 1990. See also [?, ?, ?].

Sutton:2021:FPT

[Sut21] Andrew M. Sutton. Fixed-parameter tractability of crossover: Steady-state GAs on the closest string problem. Algorithmica, 83(4):1138–1163, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00809-8.

Sykora:1987:TCA

[SV87] O. Sýkora and I. Vř’to. Tight chip area lower bounds for string matching. Information Processing Letters, 26(3):117– 119, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sahinalp:1994:PAM

[SV94] S. C. Sahinalp and U. Vishkin. On a parallel-algorithms method for string matching problems. Lecture Notes in Computer Science, 778:22–??, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwerdfeger:2009:VCD

[SV09] August C. Schwerdfeger and Eric R. Van Wyk. Verifiable composition of deterministic grammars. ACM SIGPLAN Notices, 44(6):199–210, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siren:2014:IGP

[SVM14] Jouni Sirén, Niko Välimäki, and Veli Mäkinen. Indexing graphs for path queries with applications in genome research. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 11(2):375–388, March 2014. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Saarikivi:2017:FEC

[SVMM17] Olli Saarikivi, Margus Veanes, Todd Mytkowicz, and Madan Musuvathi. Fusing effectful comprehensions. ACM SIGPLAN Notices, 52(6):17–32, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sagot:1997:MSC

[SVS97] Marie-France Sagot, Alain Viari, and Henri Soldano. Multiple sequence comparison — a peptide matching approach. Theoretical Computer Science, 180(1–2):115–137, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2340.

Sulzmann:2014:FEM

[SvS14] Martin Sulzmann and Pippijn van Steenhoven. A flexible and efficient ML lexer tool based on extended regular expression submatching. In Albert Cohen, editor, Compiler Construction: 23rd International Conference, CC 2014, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2014, Grenoble, France, April 5–13, 2014, Proceedings, volume 8409 of Lecture Notes in Computer Science, pages 174–191. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014.

Saxton:1990:FGA

[SW90] Lawrence V. Saxton and Nalin Wijesinghe. A fast generalized approximate string matching algorithm. Congressus Numerantium, 78:199–206, 1990. ISSN 0384-9864. Proceedings of the Twenty-first Southeastern Conference on Combinatorics, Graph Theory, and Computing (Boca Raton, FL, 1990).

Schwartz:1993:DSI

[SW93] Michael F. Schwartz and David C. M. Wood. Discovering shared interests using graph analysis. Communications of the Association for Computing Machinery, 36(8):78–89, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/163402.html.

Snodgrass:1994:PAS

[SW94] Richard T. Snodgrass and Marianne Winslett, editors. Proceedings of the 1994 ACM SIGMOD International Conference on Management of Data / SIGMOD ’94, Minneapolis, Minnesota, May 24–27, 1994, volume 23(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

Sima:1998:TN

[ŠW98] Jiří Šíma and Jiří Wiedermann. Theory of neuromata. Journal of the Association for Computing Machinery, 45(1): 155–178, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org:80/pubs/citations/journals/jacm/1998-45-1/
p155-scaronima/.

Smyth:2009:AHP

[SW09] W. F. Smyth and Shu Wang. An adaptive hybrid patternmatching algorithm on indeterminate strings. International Journal of Foundations of Computer Science (IJFCS), 20(6): 985–1004, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schafer:2012:DCH

[SW12] Benjamin Carrion Schafer and Kazutoshi Wakabayashi. Divide and conquer high-level synthesis design space exploration. ACM Transactions on Design Automation of Electronic Systems (TODAES), 17(3):29:1–29:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sun:2012:ESM

[SWW^{+}12] Zhao Sun, Hongzhi Wang, Haixun Wang, Bin Shao, and Jianzhong Li. Efficient subgraph matching on billion node graphs. Proceedings of the VLDB Endowment, 5(9):788–799, May 2012. CODEN ???? ISSN 2150-8097.

Salton:1975:VSM

[SWY75] Gerard Salton, A. Wong, and C. S. Yang. A vector space model for automatic indexing. Communications of the Association for Computing Machinery, 18(11):613–620, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Squillante:2001:AQU

[SWZ01] Mark S. Squillante, Baffelly Woo, and Li Zhang. Analysis of queues under correlated arrivals with applications to Web server performance. ACM SIGMETRICS Performance Evaluation Review, 28(4):41–43, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sun:2025:PRE

[SXZ^{+}25] Mingqian Sun, Guangwei Xie, Fan Zhang, Wei Guo, Xitian Fan, Tianyang Li, Li Chen, and Jiayu Du. PTME: a regular expression matching engine based on speculation and enumerative computation on FPGA. ACM Transactions on Reconfigurable Technology and Systems, 18(1):3:1–3:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3655626.

Smith:1972:GRE

[SY72] L. W. Smith and S. S. Yau. Generation of regular expressions for automata by the integral of regular expressions. The Computer Journal, 15(3):222–228, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/3/222.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150222.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/222.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/224.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/226.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/228.tif.

Sadiq:2023:SEC

[SY23] Muhammad Umair Sadiq and Muhammad Murtaza Yousaf. Space-efficient computation of parallel approximate string matching. The Journal of Supercomputing, 79(8):9093–9126, May 2023. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-022-05038-6.

Symes:1985:POC

[Sym85] D. Michael Symes. Procedural operators considered as fundamental programming devices. Computer Languages, 10(2):75– 89, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schoenmeyr:2005:FBA

[SZ05] Tor Schoenmeyr and David Yu Zhang. FFT-based algorithms for the string matching with mismatches problem. Journal of Algorithms, 57(2):130–139, November 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677405000180.

Takeichi:1990:DFK

[TA90] Masato Takeichi and Yoji Akama. Deriving a functional Knuth–Morris–Pratt algorithm by transformation. Journal of Information Processing, 13(4):522–528, 1990. CODEN JIPRDE. ISSN 0387-6101.

Taft:2022:RPM

[Taf22] S. Tucker Taft. Rigorous pattern matching as a language feature. ACM SIGADA Ada Letters, 42(2):69–74, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-

9476 (electronic). URL https://dl.acm.org/doi/10.1145/
3591335.3591342.

Takaoka:1986:LPM

[Tak86a] Tadao Takaoka. An on-line pattern matching algorithm. Information Processing Letters, 22(6):329–330, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takaoka:1986:OPM

[Tak86b] Tadao Takaoka. An on-line pattern matching algorithm. Information Processing Letters, 22(6):329–330, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takeda:1993:FMA

[Tak93] Masayuki Takeda. A fast matching algorithm for patterns with pictures. Bull. Inform. Cybernet., 25(3–4):137–153, 1993. ISSN 0286-522X.

Takaoka:1994:APM

[Tak94] T. Takaoka. Approximate pattern matching with samples. Lecture Notes in Computer Science, 834:234–??, 1994. CO-DEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takahashi:1996:UCN

[Tak96a] Yoshikane Takahashi. A unified constructive network model for problem-solving. Theoretical Computer Science, 156(1– 2):217–261, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=156&issue=1-
2&aid=2075.

Takaoka:1996:LRP

[Tak96b] Tadao Takaoka. A left-to-right preprocessing computation for the Boyer–Moore string matching algorithm. The Computer Journal, 39(5):413–416, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/39/5/413.full.
pdf+html;   http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_05/390413.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/Volume_39/Issue_
05/Vol39_05.body.html#AbstractTakaoka.

Talcott:1981:BRC

[Tal81] Carolyn Talcott. Book review: A Computational Logic (Robert S. Boyer and J. Strother Moore). SIAM Review, 23 (2):264–266, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tanaka:2014:IEE

[Tan14] Shunji Tanaka. Improved exact enumerative algorithms for the planted (l, d)-motif search problem. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 11(2): 361–374, March 2014. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Tarjan:1981:FAS

[Tar81a] Robert Endre Tarjan. Fast algorithms for solving path problems. Journal of the Association for Computing Machinery, 28 (3):594–614, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1981:UAP

[Tar81b] Robert Endre Tarjan. A unified approach to path problems. Journal of the Association for Computing Machinery, 28(3): 577–593, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Taylor:1997:WMR

[Tay97] Dave Taylor. The Web master: Roll your own search with grep. ;login:, 22(4):37–41, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Trabalka:2000:RSP

[TB00] Marek Trabalka and Mária Bieliková. Realization of syntactic parser for inflectional language using XML and regular expressions. Lecture Notes in Computer Science, 1902:63–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1902/19020063.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1902/19020063.pdf.

Taft:2022:DPM

[TBD22] S. Tucker Taft, Stephen Baird, and Claire Dross. Defining a pattern matching language feature for Ada. ACM SIGADA Ada Letters, 42(1):79, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://dl.

acm.org/doi/10.1145/3577949.3577963.

Tan:2006:BSS

[TBS06] Lin Tan, Brett Brotherton, and Timothy Sherwood. Bit-split string-matching engines for intrusion detection and prevention. ACM Transactions on Architecture and Code Optimization, 3(1):3–34, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

tenCate:2009:NEP

[tC09] Balder ten Cate. A note on the expressibility problem for modal logics and star-free regular expressions. Information Processing Letters, 109(10):509–513, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tucci:1991:RNC

[TCC91] Maurizio Tucci, Gennaro Costagliola, and Shi-Kuo Chang. A remark on NP-completeness of picture matching. Information Processing Letters, 39(5):241–243, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsui:1990:OES

[TCCK90] Hung-Tat Tsui, Ming-Hong Chan, Kin-Cheong Chu, and Shao-Hua Kong. Orientation estimation of 3D surface patches. Computer Vision, Graphics, and Image Processing, 50(1): 112–124, April 1990. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Thibodeau:2016:ICT

[TCP16] David Thibodeau, Andrew Cave, and Brigitte Pientka. Indexed codata types. ACM SIGPLAN Notices, 51(9):351–363, September 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trevisan:2018:RUC

[TD18] Martino Trevisan and Idilio Drago. Robust URL classification with generative adversarial networks. ACM SIGMET-RICS Performance Evaluation Review, 46(3):143–146, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tejiscak:2020:DTC

[Tej20] Matús Tejiscák. A dependently typed calculus with pattern matching and erasure inference. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):91:1–

91:29, August 2020. URL https://dl.acm.org/doi/10.
1145/3408973.

Tuijn:1996:CCG

[TG96] Chris Tuijn and Marc Gyssens. CGOOD, a categorical graph-oriented object data model. Theoretical Computer Science, 160(1–2):217–239, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=160&issue=1-
2&aid=2030.

Tarhio:2017:TBA

[THG17] Jorma Tarhio, Jan Holub, and Emanuele Giaquinta. Technology beats algorithms (in exact string matching). Software — Practice and Experience, 47(12):1877–1885, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thiemann:1993:ART

[Thi93] P. Thiemann. Avoiding repeated tests in pattern matching. Lecture Notes in Computer Science, 724:141–??, 1993. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turonova:2020:RMC

[THL^{+}20] Lenka Turoňová, Lukáš Holík, Ondřej Lengál, Olli Saarikivi, Margus Veanes, and Tomáš Vojnar. Regex matching with counting-set automata. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):218:1–218:30,

November 2020. URL https://dl.acm.org/doi/10.1145/
3428286.

Thompson:1968:PTR

[Tho68] Ken Thompson. Programming techniques: Regular expression search algorithm. Communications of the Association for Computing Machinery, 11(6):419–422, June 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://patft.uspto.gov/. See also [KP99c, Cox07, Cox09, Cox10a, Cox12].

Thomas:1981:RSH

[Tho81] W. Thomas. Remark on the star-height-problem. Theoretical Computer Science, 13(2):231–237, February 1981. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tang:2019:SAO

[THQ19] Junjian Tang, Shunli Hao, and Wenqi Qu. Sentiment analysis of online Chinese comments based on statistical learning combining with pattern matching. Concurrency and Computation: Practice and Experience, 31(10):e4765:1–e4765:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Takahashi:1990:SCM

[TIAY90] H. Takahashi, N. Itoh, T. Amano, and A. Yamashita. A spelling correction method and its application to an OCR system. Pattern Recognition, 23(3-4):363–377, 1990. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Toda:1983:TDP

[TIT83] M. Toda, K. Inoue, and I. Takanami. Two-dimensional pattern matching by two-dimensional on-line tessellation acceptors. Theoretical Computer Science, 24(2):179–194, July 1983. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tang:2017:RDS

[TJD^{+}17] Qiu Tang, Lei Jiang, Qiong Dai, Majing Su, Hongtao Xie, and Binxing Fang. RICS-DFA: a space and time-efficient signature matching algorithm with Reduced Input Character Set. Concurrency and Computation: Practice and Experience, 29(20): ??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tiwari:2022:SES

[TJGY22] Prayag Tiwari, Amit Kumar Jaiswal, Sahil Garg, and Ilsun You. SANTM: Efficient self-attention-driven network for text matching. ACM Transactions on Internet Technology (TOIT), 22(3):55:1–55:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https://dl.

acm.org/doi/10.1145/3426971.

Travnicek:2020:MBM

[TJMC20] Jan Trávníček, Jan Janoušek, Bořivoj Melichar, and Loek Cleophas. On modification of Boyer–Moore–Horspool’s algorithm for tree pattern matching in linearised trees. Theoretical Computer Science, 830–831(??):60–90, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397520302425.

Tran:2007:FBC

[TK07] Thinh Ngoc Tran and Surin Kittitornkun. FPGA-based cuckoo hashing for pattern matching in NIDS/NIPS. Lecture Notes in Computer Science, 4773:334–343, 2007. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-540-75476-3_34.

TranconyWidemann:2012:PPMa

[TL12a] Baltasar Trancón y Widemann and Markus Lepper. Paisley: Pattern matching à la carte. Lecture Notes in Computer Science, 7307:240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30476-7_16/.

TranconyWidemann:2012:PPMb

[TL12b] Baltasar Trancón y Widemann and Markus Lepper. Paisley: Pattern matching à la carte. Lecture Notes in Computer Science, 7307:240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30476-7_16/.

Tran:2015:CLC

[TLC15] Nhat-Phuong Tran, Myungho Lee, and Dong Hoon Choi. Cache locality-centric parallel string matching on many-core accelerator chips. Scientific Programming, 2015(??):937694:1– 937694:20, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://www.hindawi.

com/journals/sp/2015/937694/.

Tseng:2007:DHS

[TLLL07] Kuo-Kun Tseng, Ying-Dar Lin, Tsern-Huei Lee, and Yuan-Cheng Lai. Deterministic high-speed root-hashing automaton matching coprocessor for embedded network processor. ACM SIGARCH Computer Architecture News, 35(3):36–43, June 2007. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tseng:2009:FSA

[TLLL09] Kuo-Kun Tseng, Yuan-Cheng Lai, Ying-Dar Lin, and Tsern-Huei Lee. A fast scalable automaton-matching accelerator for embedded content processors. ACM Transactions on Embedded Computing Systems, 8(3):19:1–19:??, April 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tran:2016:BPA

[TLS16] Tuan Tu Tran, Yongchao Liu, and Bertil Schmidt. Bitparallel approximate pattern matching: Kepler GPU versus Xeon Phi. Parallel Computing, 54(??):128–138, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-

7336 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167819115001477.

Tao:2004:LBC

[TM04] T. Tao and A. Mukherjee. LZW based compressed pattern matching. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314. LCCN ???? URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281544. IEEE Computer Society Order Number: P2082.

Tao:2005:PML

[TM05a] T. Tao and Amar Mukherjee. Pattern matching in LZW compressed files. IEEE Transactions on Computers, 54(8):929– 938, August 2005. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1453495.

Tao:2005:MPM

[TM05b] Tao Tao and A. Mukherjee. Multiple-pattern matching in LZW compressed files using Aho–Corasick algorithm. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN

1068-0314. LCCN ???? URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=1402239.

Takeda:2002:PTF

[TMK^{+}02] Masayuki Takeda, Satoru Miyamoto, Takuya Kida, Ayumi Shinohara, Shuichi Fukamachi, Takeshi Shinohara, and Setsuo Arikawa. Processing text files as is: Pattern matching over compressed texts, multi-byte character texts, and semistructured texts. Lecture Notes in Computer Science, 2476: 170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.de/

link/service/series/0558/bibs/2476/24760170.htm;

http://link.springer.de/link/service/series/0558/papers/
2476/24760170.pdf.

Thompson:2001:DDS

[TMV^{+}01] Paul M. Thompson, Michael S. Mega, Christine Vidal, Judith L. Rapoport, and Arthur W. Toga. Detecting diseasespecific patterns of brain structure using cortical pattern matching and a population-based probabilistic brain atlas. Lecture Notes in Computer Science, 2082:488–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2082/20820488.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2082/20820488.pdf.

Traytel:2013:VDP

[TN13] Dmitriy Traytel and Tobias Nipkow. Verified decision procedures for MSO on words based on derivatives of regular expressions. ACM SIGPLAN Notices, 48(9):3–12, September 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Traytel:2015:VDP

[TN15] Dmitriy Traytel and Tobias Nipkow. Verified decision procedures for MSO on words based on derivatives of regular expressions. Journal of Functional Programming, 25:e18, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/journals/journal-of-functional-
programming/article/verified-decision-procedures-for-
mso-on-words-based-on-derivatives-of-regular-expressions/
18DCED718D5D525252C97EFA3501B4A4.

Tarhio:1997:SMD

[TP97] Jorma Tarhio and Hannu Peltola. String matching in the DNA alphabet. Software — Practice and Experience, 27(7):851– 861, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=7320;   http://www3.
interscience.wiley.com/cgi-bin/fulltext?ID=7320&PLACEBO=
IE.pdf.

Tata:2007:ESTa

[TP07a] Sandeep Tata and Jignesh M. Patel. Estimating the selectivity of tf-idf based cosine similarity predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2): 7–12, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tata:2007:ESTb

[TP07b] Sandeep Tata and Jignesh M. Patel. Estimating the selectivity of tf-idf based cosine similarity predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4): 75–80, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tateishi:2013:PIS

[TPT13] Takaaki Tateishi, Marco Pistoia, and Omer Tripp. Pathand index-sensitive string analysis based on monadic secondorder logic. ACM Transactions on Software Engineering and Methodology, 22(4):33:1–33:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tan:2005:HTS

[TS05] Lin Tan and Timothy Sherwood. A high throughput string matching architecture for intrusion detection and prevention. ACM SIGARCH Computer Architecture News, 33(2):112–122, May 2005. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Toyoda:2013:PDD

[TSI13] Machiko Toyoda, Yasushi Sakurai, and Yoshiharu Ishikawa. Pattern discovery in data streams under the time warping distance. VLDB Journal: Very Large Data Bases, 22(3):295– 318, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tanenbaum:1988:MAS

[TSM88] Andrew S. Tanenbaum, Johan W. Stevenson, and Jost Muller. MINIX for the ATARI ST and MINIX manual for the ATARI ST. Prentice-Hall, Upper Saddle River, NJ 07458, USA, version 1.1. edition, 1988. ISBN 0-13-584392-8 (disks), 0-13-584434-7 (manual). LCCN QA76.76.O63. 9 computer disks.

Tharp:1982:PTS

[TT82] Alan L. Tharp and Kuo-Chung Tai. The practicality of text signatures for accelerating string searching. Software — Practice and Experience, 12(1):35–44, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Expands Harrison’s work [Har71].

Taha:2020:AMM

[TT20] Mohammad M. A. Taha and Christof Teuscher. Approximate memristive in-memory Hamming distance circuit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):18:1–18:14, April 2020. CODEN ???? ISSN 1550-4832.

URL https://dl.acm.org/doi/abs/10.1145/3371391.

Takayama:2022:IPM

[TT22] Riku Takayama and Jubee Tada. An implementation of a pattern matching accelerator on a RISC-V processor. In IEEE, editor, 2022 Tenth International Symposium on Computing and Networking Workshops (CANDARW), pages 273– 275. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Tian:2005:PMC

[TTHP05] Yuanyuan Tian, Sandeep Tata, Richard A. Hankins, and Jignesh M. Patel. Practical methods for constructing suffix trees. VLDB Journal: Very Large Data Bases, 14(3):281–299, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0938-1287&volume=
14&issue=3&spage=281.

Tchendji:2022:CGM

[TTO^{+}22] Vianney Kengne Tchendji, Hermann Bogning Tepiele, Mathias Akong Onabid, Jean Frédéric Myoupo, and Jerry Lacmou Zeutouo. A coarse-grained multicomputer parallel algorithm for the sequential substring constrained longest common subsequence problem. Parallel Computing, 111(??):??, July 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-

7336 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016781912200028X.

Tarhio:1988:GAA

[TU88] Jorma Tarhio and Esko Ukkonen. A greedy approximation algorithm for constructing shortest common superstrings. Theoretical Computer Science, 57(1):131–145, April 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tarhio:1993:ABM

[TU93] Jorma Tarhio and Esko Ukkonen. Approximate Boyer–Moore string matching. SIAM Journal on Computing, 22(2):243– 260, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Turner:1986:OM

[Tur86] D. Turner. An overview of Miranda. ACM SIGPLAN Notices, 21(12):158–166, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tan:2014:REQ

[TV14] Tony Tan and Domagoj Vrgoč. Regular expressions for querying data graphs. International Journal of Foundations of Computer Science (IJFCS), 25(8):971–??, December 2014. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tumeo:2012:ACS

[TVCM12] Antonino Tumeo, Oreste Villa, and Daniel G. Chavarria-Miranda. Aho–Corasick string matching on shared and distributed-memory parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 23(3):436–443, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2023:TBU

[TWZ23] Zhihao Gavin Tang, Xiaowei Wu, and Yuhao Zhang. Toward a better understanding of randomized greedy matching. Journal of the Association for Computing Machinery, 70(6):39:1– 39:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3614318.

Teng:1997:ASS

[TY97] Shang-Hua Teng and Frances F. Yao. Approximating shortest superstrings. SIAM Journal on Computing, 26(2):410–417, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/28612.

Takahashi:1986:NSS

[TYNM86] K. Takahashi, H. Yamada, H. Nagai, and K. Matsumi. A new string search hardware architecture for VLSI. ACM SIGARCH Computer Architecture News, 14(2):20–27, June 1986. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Tseng:2013:NNE

[TZH^{+}13] Kuo-Kun Tseng, Fu-Fu Zeng, Huang-Nan Huang, Yiming Liu, Jeng-Shyang Pan, W. H. Ip, and C. H. Wu. A new nonexact Aho–Corasick framework for ECG classification. ACM SIGARCH Computer Architecture News, 41(2):41–46, May 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tang:1994:MMN

[TZW94] S. Tang, K. Zhang, and X. Wu. Matching with matrix norm minimization. Lecture Notes in Computer Science, 807:250– 258, 1994. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tang:2014:EPP

[TZYH14] Zhenjun Tang, Xianquan Zhang, Chunqiang Yu, and Dan He. Efficient point pattern matching algorithm for planar point sets under transform of translation, rotation and scale. Applied Mathematics and Computation, 232(??):624–631, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-

5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300314001246.

Ulus:2024:ETP

[UFA^{+}24] Dogan Ulus, Thomas Ferrère, Eugene Asarin, Dejan Nickovic, and Oded Maler. Elements of timed pattern matching. ACM Transactions on Embedded Computing Systems, 23(4):59:1– 59:??, July 2024. CODEN ???? ISSN 1539-9087 (print),

1558-3465 (electronic). URL https://dl.acm.org/doi/10.
1145/3645114.

Ueda:2022:DLG

[UIM22] Yuki Ueda, Takashi Ishio, and Kenichi Matsumoto. DevReplay: Linter that generates regular expressions for repeating code changes. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167642322000909.

Ukkonen:1992:ASM

[Ukk92] E. Ukkonen. Approximate string-matching with q-grams and maximal matches. Theoretical Computer Science, 92(1):191– 211, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ukkonen:1993:ASMa

[Ukk93] Esko Ukkonen. Approximate string-matching over suffix trees. In Combinatorial pattern matching (Padova, 1993), volume 684 of Lecture Notes in Computer Science, pages 228–242. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993.

Ukkonen:2010:GPP

[Ukk10] E. Ukkonen. Geometric point pattern matching in the Knuth– Morris–Pratt way. J.UCS: Journal of Universal Computer Science, 16(14):1902–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://www.jucs.

org/jucs_16_14/geometric_point_pattern_matching.

Umali:1997:BRM

[Uma97] Rick Umali. Book review: Mastering regular expressions. ;login:, 22(5):66–67, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Urban:2023:PLD

[Urb23] Christian Urban. POSIX lexing with derivatives of regular expressions. Journal of Automated Reasoning, 67(3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print),

1573-0670 (electronic). URL https://link.springer.com/
article/10.1007/s10817-023-09667-1.

USENIX:1992:PWU

[USE92] USENIX, editor. Proceedings of the Winter 1992 USENIX Conference: January 20 — January 24, 1992, San Francisco, California. USENIX, Berkeley, CA, USA, 1992.

Ukkonen:1993:ASMb

[UW93] Esko Ukkonen and Derick Wood. Approximate string matching with suffix automata. Algorithmica, 10(5):353–364, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Valiente:2009:CPM

[Val09] Gabriel Valiente. Combinatorial pattern matching algorithms in computational biology using Perl and R. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, Boca Raton, FL, USA, 2009. ISBN 1-4200-6973-X (hardcover), 1-4200-6974-8 (e-book). 352 pp. LCCN QH324.2 .V35 2009.

Vansummeren:2006:TIU

[Van06] Stijn Vansummeren. Type inference for unique pattern matching. ACM Transactions on Programming Languages and Systems, 28(3):389–428, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

vanderLoo:2014:PRP

[van14a] Mark van der Loo. stringdist: an R package for approximate string matching. The R Journal, 6(1):111– 122, June 2014. CODEN ???? ISSN 2073-4859.

URL http://journal.r-project.org/archive/2014-1/
RJournal_2014-1_loo.pdf.

vanderLoo:2014:PSR

[van14b] Mark van der Loo. stringdist: an R package for approximate string matching. The R Journal, 6(1):111– 122, June 2014. CODEN ???? ISSN 2073-4859.

URL http://journal.r-project.org/archive/2014-1/
RJournal_2014-1_loo.pdf.

Vujovic:1998:EAF

[VB98] N. Vujovic and D. Brzakovic. Evaluation of an algorithm for finding a match of a distorted texture pattern in a large image database. ACM Transactions on Information Systems, 16(1): 31–60, January 1998. CODEN ATISET. ISSN 1046-8188.

Varol:2012:HMA

[VB12] Cihan Varol and Coskun Bayrak. Hybrid matching algorithm for personal names. Journal of Data and Information Quality (JDIQ), 3(4):8:1–8:??, September 2012. CODEN ???? ISSN 1936-1955.

Valizadeh:2023:SBR

[VB23] Mojtaba Valizadeh and Martin Berger. Search-based regular expression inference on a GPU. Proceedings of the ACM on Programming Languages (PACMPL), 7(PLDI):160:1–160:??, June 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.1145/3591274.

Valgenti:2012:GGH

[VCS^{+}12] Victor C. Valgenti, Jatin Chhugani, Yan Sun, Nadathur Satish, Min Sik Kim, and Changkyu Kim. GPP-grep: Highspeed regular expression processing engine on general purpose processors. Lecture Notes in Computer Science, 7462: 334–353, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-33338-5_17/.

Viswam:2017:EBF

[VD17] A. Viswam and G. Darsan. An efficient Bitcoin fraud detection in social media networks. In 2017 International Conference on Circuit, Power and Computing Technologies (ICCPCT), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2017.

Vere:1970:TEE

[Ver70a] S. Vere. Translation equations (errata). Communications of the Association for Computing Machinery, 13(5):286, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ver70b].

Vere:1970:TE

[Ver70b] Steven Vere. Translation equations. Communications of the Association for Computing Machinery, 13(2):83–89, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [Ver70a].

Verma:1992:STP

[Ver92] Rakesh M. Verma. Strings, trees, and patterns. Information Processing Letters, 41(3):157–161, March 6, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Viksna:2001:PMP

[VG01] Juris Vīksna and David Gilbert. Pattern matching and pattern discovery algorithms for protein topologies. Lecture Notes in Computer Science, 2149:98–111, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
2149/21490098.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2149/21490098.pdf.

VanWyk:1988:LPE

[VHC88] Christopher J. Van Wyk, Eric Hamilton, and Don Colner. Literate programming: Expanding generalized regular expressions. Communications of the Association for Computing Machinery, 31(12):1376–1385, December 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Veanes:2012:SFS

[VHL^{+}12] Margus Veanes, Pieter Hooimeijer, Benjamin Livshits, David Molnar, and Nikolaj Bjorner. Symbolic finite state transducers: algorithms and applications. ACM SIGPLAN Notices, 47(1):137–150, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vialette:2002:PMP

[Via02] Stéphane Vialette. Pattern matching problems over 2interval sets. Lecture Notes in Computer Science, 2373:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2373/23730053.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2373/23730053.pdf.

Vialette:2004:CCI

[Via04] Stéphane Vialette. On the computational complexity of 2interval pattern matching problems. Theoretical Computer Science, 312(2–3):223–249, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vineberg:1977:ICSa

[Vin77a] Maniel Vineberg. Implementation of character string pattern matching on a multiprocessor. ACM SIGARCH Computer Architecture News, 6(2):1–7, May 1977. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Vineberg:1977:ICSb

[Vin77b] Maniel Vineberg. Implementation of character string pattern matching on a multiprocessor. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):1–7, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vishkin:1990:DSN

[Vis90] U. Vishkin. Deterministic sampling — a new technique for fast pattern matching. In ACM [ACM90b], pages 170–180. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/pubs/citations/proceedings/
stoc/100216/p170-vishkin/. ACM order no. 508900.

Vishkin:1991:DSN

[Vis91] Uzi Vishkin. Deterministic sampling — a new technique for fast pattern matching. SIAM Journal on Computing, 20(1): 22–40, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Visser:1999:SPM

[Vis99] Eelco Visser. Strategic pattern matching. Lecture Notes in Computer Science, 1631:30–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1631/
16310030.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/1631/16310030.pdf.

Vasiliadis:2017:DIS

[VKPI17] Giorgos Vasiliadis, Lazaros Koromilas, Michalis Polychronakis, and Sotiris Ioannidis. Design and implementation of a stateful network packet processing framework for GPUs. IEEE/ACM Transactions on Networking, 25(1):610–623, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vazou:2017:TTP

[VLP17] Niki Vazou, Leonidas Lampropoulos, and Jeff Polakow. A tale of two provers: verifying monoidal string matching in liquid Haskell and Coq. ACM SIGPLAN Notices, 52(10):63– 74, October 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Veanes:2015:DPS

[VMML15] Margus Veanes, Todd Mytkowicz, David Molnar, and Benjamin Livshits. Data-parallel string-manipulating programs. ACM SIGPLAN Notices, 50(1):139–152, January 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanNoord:2001:ERE

[vNG01] Gertjan van Noord and Dale Gerdemann. An extendible regular expression compiler for finite-state approaches in natural language processing. Lecture Notes in Computer Science, 2214:122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2214/
22140122.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2214/22140122.pdf.

Volanschi:2012:PMM

[Vol12] Nic Volanschi. Pattern matching for the masses using custom notations. Science of Computer Programming, 77(5):609–635, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print),

1872-7964 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167642311002243.

Vouillon:2006:PRT

[Vou06] Jérôme Vouillon. Polymorphic regular tree types and patterns. ACM SIGPLAN Notices, 41(1):103–114, January 2006. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Varma:2018:SAW

[VR18] Paroma Varma and Christopher Ré. Snuba: automating weak supervision to label training data. Proceedings of the VLDB Endowment, 12(3):223–236, November 2018. CODEN ???? ISSN 2150-8097.

Vaiwsri:2024:EBS

[VRC24] Sirintra Vaiwsri, Thilina Ranbaduge, and Peter Christen. Encryption-based sub-string matching for privacy-preserving record linkage. Journal of Information Security and Applications (JISA), 81(??):??, March 2024. CODEN ???? ISSN

2214-2126. URL http://www.sciencedirect.com/science/
article/pii/S2214212624000152.

Vilares:2001:AVP

[VRD01] Manuel Vilares, Francisco J. Ribadas, and Victor M. Darriba. Approximate VLDC pattern matching in sharedforest. Lecture Notes in Computer Science, 2004:483–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2004/20040483.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2004/20040483.pdf.

Voss:2001:APP

[VS01] K. Voss and H. Suesse. Affine point pattern matching. Lecture Notes in Computer Science, 2191:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2191/21910155.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2191/21910155.pdf.

V:2011:BBI

[VS11] Sharath Chandra V. and S. Selvakumar. BIXSAN: browser independent XSS sanitizer for prevention of XSS attacks. ACM SIGSOFT Software Engineering Notes, 36(5):1–7, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanBiljon:1987:RAP

[VSM87] W. R. Van Biljon, D. A. Sewry, and M. A. Mulders. Register allocation in a pattern matching code generator. Software — Practice and Experience, 17(8):521–531, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Villa:2008:ART

[VSP08] Oreste Villa, Daniele Paolo Scarpazza, and Fabrizio Petrini. Accelerating real-time string searching with multicore processors. Computer, 41(4):42–50, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vieira:2004:LEH

[VVV04] Luiz Filipe M. Vieira, Marcos Augusto M. Vieira, and Newton J. Vieira. Language emulator, a helpful toolkit in the learning process of computer theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):135–139, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vespa:2011:DFA

[VW11] Lucas Vespa and Ning Weng. Deterministic finite automata characterization and optimization for scalable pattern matching. ACM Transactions on Architecture and Code Optimization, 8(1):4:1–4:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vespa:2011:MDM

[VWR11] Lucas Vespa, Ning Weng, and Ramaswamy Ramaswamy. MS-DFA: Multiple-stride pattern matching for scalable deep packet inspection. The Computer Journal, 54(2):285– 303, February 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/54/2/285.full.pdf+html.

Wadler:1987:VWP

[Wad87] P. Wadler. Views: a way for pattern matching to cohabit with data abstraction. In ACM [ACM87], pages 307–313.

ISBN ???? LCCN ???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/41625/p307-wadler/.

Wagner:1974:OCR

[Wag74] Robert A. Wagner. Order-n correction for regular languages. Communications of the Association for Computing Machinery, 17(5):265–268, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waga:2023:PTP

[WAH23] Masaki Waga, Étienne André, and Ichiro Hasuo. Parametric timed pattern matching. ACM Transactions on Software Engineering and Methodology, 32(1):10:1–10:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3517194.

Walther:1988:MSU

[Wal88] Christoph Walther. Many-sorted unification. Journal of the Association for Computing Machinery, 35(1):1–17, January 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/45071.html.

Walker:1989:FCP

[Wal89] Kenneth W. Walker. First-class patterns for Icon. Computer Languages, 14(3):153–163, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Watson:1996:NRG

[Wat96] B. W. Watson. A new regular grammar pattern matching algorithm. Lecture Notes in Computer Science, 1136:364–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watson:2003:NRG

[Wat03] Bruce W. Watson. A new regular grammar pattern matching algorithm. Theoretical Computer Science, 299(1–3):509–521, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weste:1983:DTW

[WBA83] N. Weste, D. J. Burr, and B. D. Ackland. Dynamic time Warp pattern matching using an integrated multiprocessing array. IEEE Transactions on Computers, C-32(8):731–744, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1676311.

Wang:2022:DRE

[WBS22] Peipei Wang, Chris Brown, and Kathryn T. Stolee. Demystifying regular expression bugs. Empirical Software Engineering, 27(1):??, January 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://link.

springer.com/article/10.1007/s10664-021-10033-1.

Wang:2014:ODA

[WC14] Hung-Lung Wang and Kuan-Yu Chen. One-dimensional approximate point set pattern matching with L_{p}-norm. Theoretical Computer Science, 521(??):42–50, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513008669.

Wang:1994:CPDa

[WCM^{+}94a] Jason Tsong-Li Wang, Gung-Wei Chirn, Thomas G. Marr, Bruce Shapiro, Dennis Shasha, and Kaizhong Zhang. Combinatorial pattern discovery for scientific data: some preliminary results. In Snodgrass and Winslett [SW94], pages 115–125. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

URL http://www.acm.org/pubs/articles/proceedings/
mod/191839/p115-wang/p115-wang.pdf;   http://www.acm.
org/pubs/citations/proceedings/mod/191839/p115-wang/

.

Wang:1994:CPDb

[WCM^{+}94b] Jason Tsong-Li Wang, Gung-Wei Chirn, Thomas G. Marr, Bruce Shapiro, Dennis Shasha, and Kaizhong Zhang. Combinatorial pattern discovery for scientific data: Some preliminary results. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):115–125, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wong:1982:DAS

[WCW82] K. Y. Wong, R. G. Casey, and F. M. Wahl. Document Analysis System. IBM Journal of Research and Development, 26(6): 647–656, November 1982. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Wendling:1999:PRS

[WD99] L. Wendling and J. Desachy. Pattern recognition of strong graphs based on possibilistic c-means and k-formulae matching. Lecture Notes in Computer Science, 1566:180–189, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wandelt:2014:SAS

[WDG^{+}14] Sebastian Wandelt, Dong Deng, Stefan Gerdjikov, Shashwat Mishra, Petar Mitankin, Manish Patil, Enrico Siragusa, Alexander Tiskin, Wei Wang, Jiaying Wang, and Ulf Leser. State-of-the-art in string similarity search and join. SIG-MOD Record (ACM Special Interest Group on Management of Data), 43(1):64–76, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weatherford:1994:HLP

[Wea94] Stephen Andrew Weatherford. High-level pattern-matching extensions to C++ for Fortran program manipulation in Polaris. Thesis (m.s), University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1994. viii + 104 pp.

Webber:1995:OFP

[Web95] Adam Webber. Optimization of functional programs by grammar thinning. ACM Transactions on Programming Languages and Systems, 17(2):293–330, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/0164-0925/201067.
html;   https://www.math.utah.edu/pub/tex/bib/fibquart.
bib.

Weiser:1983:RSB

[Wei83] Mark Weiser. Reconstructing sequential behavior from parallel behavior projections. Information Processing Letters, 17 (3):129–135, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weiner:1984:LRK

[Wei84] J. L. Weiner. The logical record keeper: PROLOG on the IBM. Byte Magazine, 9(9):125–31, September 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). QA76.5B98.

Wentworth:1993:GRE

[Wen93] E. P. Wentworth. Generalized regular expressions — a programming exercise in Haskell. ACM SIGPLAN Notices, 28(5):

49–54, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Westly:1997:TTA

[Wes97] T. J. Westly. TASH: Tcl Ada SHell, an Ada/Tcl binding. ACM SIGADA Ada Letters, 17(2):82–91, March/April 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wang:2021:GSH

[WGL^{+}21] Chao Wang, Lei Gong, Shiming Lei, Haijie Fang, Xi Li, Aili Wang, and Xuehai Zhou. GenSeq+: a scalable highperformance accelerator for genome sequencing. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 18(4):1512–1523, July/August 2021. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic). URL https://dl.

acm.org/doi/10.1109/TCBB.2019.2947059.

Wang:2013:RPM

[WGMH13] Meng Wang, Jeremy Gibbons, Kazutaka Matsuda, and Zhenjiang Hu. Refactoring pattern matching. Science of Computer Programming, 78(11):2216–2242, November 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642312001426.

Wang:2017:GSM

[WHZ^{+}17] Kai Wang, Aftab Hussain, Zhiqiang Zuo, Guoqing Xu, and Ardalan Amiri Sani. Graspan: a single-machine disk-based graph system for interprocedural static analyses of large-scale systems code. ACM SIGARCH Computer Architecture News, 45(1):389–404, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Winkowski:1978:MFC

[Win78] Jozef Winkowski, editor. Mathematical foundations of computer science, 1978: proceedings, 7th Symposium, Zakopane, Poland, September 4–8, 1978, volume 64 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1978.

Weber:1994:APP

[WKA94] G. Weber, L. Knipping, and H. Alt. An application of point pattern matching in astronautics. Journal of Symbolic Computation, 17(4):321–340, April 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wolinski:2009:ADA

[WKR09] Christophe Wolinski, Krzysztof Kuchcinski, and Erwan Raffin. Automatic design of application-specific reconfigurable processor extensions with UPaK synthesis kernel. ACM Transactions on Design Automation of Electronic Systems (TODAES), 15(1):1:1–1:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wandelt:2015:MCS

[WL15a] Sebastian Wandelt and Ulf Leser. MRCSI: compressing and searching string collections with multiple references. Proceedings of the VLDB Endowment, 8(5):461–472, January 2015. CODEN ???? ISSN 2150-8097.

Wang:2015:FPPb

[WL15b] Kai Wang and Jun Li. FREME: a pattern partition based engine for fast and scalable regular expression matching in practice. Journal of Network and Computer Applications, 55(??):154–169, September 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804515001198.

Wang:2014:ESS

[WLF14] Jiannan Wang, Guoliang Li, and Jianhua Feng. Extending string similarity join to tolerant fuzzy token matching. ACM Transactions on Database Systems, 39(1):7:1–7:??, January 2014. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wu:1992:FTS

[WM92a] Sun Wu and Udi Manber. Fast text searching allowing errors. Communications of the Association for Computing Machinery, 35(10):83–91, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/0001-0782/135244.

html. This algorithm in this paper is implemented in the agrep program, publicly available via ANONYMOUS FTP to cs.arizona.edu in the agrep subdirectory. See also [BYG92].

Wu:1992:AFA

[WM92b] Sun Wu and Udi Manber. agrep — a fast approximate pattern-matching tool. In USENIX [USE92], pages 153–162.

Wang:2019:GGC

[WMGS19] Ping Wang, Luke Mchale, Paul V. Gratz, and Alex Sprintson. GenMatcher: a generic clustering-based arbitrary matching framework. ACM Transactions on Architecture and Code Optimization, 15(4):51:1–51:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:1995:SAA

[WMM95] S. Wu, U. Manber, and E. Myers. A subquadratic algorithm for approximate regular expression matching. Journal of Algorithms, 19(3):346–360, November 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710413.

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Wulf:1983:SFR

[WNL^{+}83] William A. Wulf, Joe Newcomer, Bruce Leverett, Rick Cattell, and Paul Knueven. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):279–280, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, GHF83a, Fra83, GHF83b].

Wolberg:1986:SOF

[Wol86] George Wolberg. A syntactic omni-font character recognition system. In CVPR 1986 [CVP86], pages 168–173. ISBN 0-8186-0721-1. LCCN TA1632 .I36 1986. IEEE Service Cent. Piscataway, NJ, USA.

Wolff:1990:SPS

[Wol90] J. G. Wolff. Simplicity and power — some unifying ideas in computing. The Computer Journal, 33(6):518–534, December 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/33/6/518.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/518.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/519.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/520.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/521.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/522.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/523.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/524.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/525.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/526.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/527.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/528.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/529.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/530.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/531.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/532.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_06/tiff/533.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_06/
tiff/534.tif.

Woods:1986:MPB

[Woo86] J. A. Woods. More pep for Boyer–Moore grep. Usenet netnews group net.unix, March 18, 1986.

Woods:1987:E

[Woo87] J. A. Woods. egrep. Usenet netnews group net.unix, March 1987.

Weddle:2007:PGS

[WOQ^{+}07] Charles Weddle, Mathew Oldham, Jin Qian, An-I Andy Wang, Peter Reiher, and Geoff Kuenning. PARAID: a gearshifting power-aware RAID. ACM Transactions on Storage, 3 (3):13:1–13:??, October 2007. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Wang:2013:MMC

[WPKL13] Hao Wang, Shi Pu, Gabe Knezek, and Jyh-Charn Liu. MIN-MAX: A counter-based algorithm for regular expression matching. IEEE Transactions on Parallel and Distributed Systems, 24(1):92–103, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2015:TPE

[WR15] Lei Wei and Michael K. Reiter. Toward practical encrypted email that supports private, regular-expression searches. International Journal of Information Security, 14(5):397–416, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL http://

link.springer.com/article/10.1007/s10207-014-0268-
3;   http://link.springer.com/content/pdf/10.1007/
s10207-014-0268-3.pdf.

Wright:1994:ASM

[Wri94] Alden H. Wright. Approximate string matching using withinword parallelism. Software — Practice and Experience, 24 (4):337–362, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wagman:1994:UIM

[WSS94] D. Wagman, M. Schneider, and E. Shnaider. On the use of interval mathematics in fuzzy expert systems. International Journal of Intelligent Systems, 9(2):241–259, February 1994. CODEN IJISED. ISSN 0884-8173 (print), 1098-111x (electronic).

Wu:2022:DCG

[WSVS22] Lingxi Wu, Rasool Sharifi, Ashish Venkat, and Kevin Skadron. DRAM-CAM: General-purpose bit-serial exact pattern matching. IEEE Computer Architecture Letters, 21(2): 89–92, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2016:MFP

[WSW16] Z. Wang, H. Seidel, and T. Weinkauf. Multi-field pattern matching based on sparse feature sampling. IEEE Transactions on Visualization and Computer Graphics, 22(1):807– 816, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wood:1988:IFS

[WT88] David P. Wood and David Turcaso. Implementing a faster string search algorithm in Ada. ACM SIGADA Ada Letters, 8(3):87–97, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1989:IFS

[WT89] P. Wood and D. Turcaso. Implementing a faster string search algorithm in Ada. ACM SIGADA Ada Letters, 8(3):87–97, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weber-Wulff:1993:PMP

[WW93] Debora Weber-Wulff. Proof movie — a proof with the Boyer– Moore prover. Formal Aspects of Computing, 5(2):121–151, March 1993. CODEN FACME5. ISSN 0934-5043 (print),

1433-299X (electronic). URL http://link.springer.com/
article/10.1007/BF01211302.

Watson:2003:BMS

[WW03] Bruce W. Watson and Richard E. Watson. A Boyer–Moorestyle algorithm for regular expression pattern matching. Science of Computer Programming, 48(2–3):99–117, August/ September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wang:2016:UHM

[WWW^{+}16] Zhongyuan Wang, Fang Wang, Haixun Wang, Zhirui Hu, Jun Yan, Fangtao Li, Ji-Rong Wen, and Zhoujun Li. Unsupervised head-modifier detection in search queries. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 19:1–19:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2012:GRE

[WXZY12] Yu Wang, Yang Xiang, Wanlei Zhou, and Shunzheng Yu. Generating regular expression signatures for network traffic classification in trusted network management. Journal of Network and Computer Applications, 35(3):992–1000, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804511000713.

Wobber:2007:AAS

[WYA^{+}07] Ted Wobber, Aydan Yumerefendi, Martín Abadi, Andrew Birrell, and Daniel R. Simon. Authorizing applications in singularity. Operating Systems Review, 41(3):355–368, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Watson:1996:TSM

[WZ96] B. W. Watson and G. Zwaan. A taxonomy of sublinear multiple keyword pattern matching algorithms. Science of Computer Programming, 27(2):85–118, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wang:2012:RCM

[WZJH12] Shu-Lin Wang, Yi-Hai Zhu, Wei Jia, and De-Shuang Huang. Robust classification method of tumor subtype by using correlation filters. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(2):580–591, March 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Wang:2023:BSC

[WZL^{+}23] Shicheng Wang, Menghao Zhang, Guanyu Li, Chang Liu, Zhiliang Wang, Ying Liu, and Mingwei Xu. Bolt: Scalable and cost-efficient multistring pattern matching with programmable switches. IEEE/ACM Transactions on Networking, 31(2):846–861, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.acm.

org/doi/10.1109/TNET.2022.3202523.

Wang:1995:PMP

[WZS95] Jason T. L. Wang, Kaizhong Zhang, and Dennis Shasha. Pattern matching and pattern discovery in scientific, program, and document databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):487, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835

(electronic). URL http://www.acm.org/pubs/articles/
proceedings/mod/223784/p487-wang/p487-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/mod/223784/
p487-wang/.

Wu:2014:FMN

[WZU14] Chunhan Wu, Xingyuan Zhang, and Christian Urban. A formalisation of the Myhill–Nerode Theorem based on regular expressions. Journal of Automated Reasoning, 52(4):451–480, April 2014. CODEN JAREEW. ISSN 0168-7433 (print),

1573-0670 (electronic). URL http://link.springer.com/
article/10.1007/s10817-013-9297-2.

Xie:2023:UDC

[XH23] Jingnan Xie and Harry B. Hunt III. On the undecidability and descriptional complexity of synchronized regular expressions. Acta Informatica, 60(3):257–278, September 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00236-023-00439-3.

Xi:2003:DTP

[Xi03] H. Xi. Dependently typed pattern matching. J.UCS: Journal of Universal Computer Science, 9(8):851–872, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-

6968 (electronic). URL http://www.jucs.org/jucs_9_8/
dependently_typed_pattern_matching.

Xuandong:2004:DCR

[XJT^{+}04] Li Xuandong, Zhao Jianhua, Zheng Tao, Li Yong, and Zheng Guoliang. Duration-constrained regular expressions. Formal Aspects of Computing, 16(2):155–163, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X

(electronic). URL http://link.springer.com/article/
10.1007/s00165-004-0033-x.

Xu:1992:RCR

[XK92] H. Xu and Y. Kambayashi. Realization of composite relationship views utilizing regular expressions. In Kim et al. [KKP92], pages 79–87. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Xu:2019:TES

[XLC19] Zhiwu Xu, Ping Lu, and Haiming Chen. Towards an effective syntax and a generator for deterministic standard regular expressions. The Computer Journal, 62(9):1322–1341, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://academic.oup.com/comjnl/
article/62/9/1322/5165111.

Xu:2011:MDP

[XMLC11] Y. Xu, L. Ma, Z. Liu, and H. J. Chao. A multi-dimensional progressive perfect hashing for high-speed string matching. In ACM, editor, ANCS’11: Proceedings of the 17th ACM/IEEE Symposium on Architectures for Networking and Communications Systems, Brooklyn, NY, USA, October 3–4, 2011, pages 167–177. ACM Press, New York, NY 10036, USA, 2011. ISBN 0-7695-4521-1.

Xu:2023:STC

[XPZ^{+}23] Qianqian Xu, Junjie Peng, Cangzhi Zheng, Shuhua Tan, Fen Yi, and Feng Cheng. Short text classification of Chinese with label information assisting. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):119:1–119:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3582301.

Xiao:2013:EET

[XQW^{+}13] Chuan Xiao, Jianbin Qin, Wei Wang, Yoshiharu Ishikawa, Koji Tsuda, and Kunihiko Sadakane. Efficient error-tolerant query autocompletion. Proceedings of the VLDB Endowment, 6(6):373–384, April 2013. CODEN ???? ISSN 2150-8097.

Xu:2019:PPS

[XZL^{+}19] Zifeng Xu, Fucai Zhou, Yuxi Li, Jian Xu, and Qiang Wang. Privacy-preserving subgraph matching protocol for two parties. International Journal of Foundations of Computer Science (IJFCS), 30(4):571–588, June 2019. ISSN 0129-

0541. URL https://www.worldscientific.com/doi/10.
1142/S0129054119400136.

Yamamoto:2001:NRA

[Yam01] Hiroaki Yamamoto. A new recognition algorithm for extended regular expressions. Lecture Notes in Computer Science, 2223:

257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2223/22230257.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/2223/22230257.pdf.

Yamamoto:2019:FAF

[Yam19] Hiroaki Yamamoto. A faster algorithm for finding shortest substring matches of a regular expression. Information Processing Letters, 143(??):56–60, March 2019. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019018302333.

Yang:1995:LAP

[Yan95] Wuu Yang. On the look-ahead problem in lexical analysis. Acta Informatica, 32(5):459–476, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL http://cissun51.cis.nctu.edu.tw/~wuuyang/lookahead.
ps.Z;   http://link.springer-ny.com/link/service/journals/
00236/bibs/5032005/50320459.htm.

Yao:1979:CPM

[Yao79] Andrew Chi Chih Yao. The complexity of pattern matching for a random string. SIAM Journal on Computing, 8(3):368– 387, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yu:2013:EDA

[YB13] Xiaodong Yu and Michela Becchi. Exploring different automata representations for efficient regular expression matching on GPUs. ACM SIGPLAN Notices, 48(8):287–288, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Yu:2023:SMT

[YBTB23] Yuncong Yu, Tim Becker, Le Minh Trinh, and Michael Behrisch. SAXRegEx: Multivariate time series pattern search with symbolic representation, regular expression, and query expansion. Computers and Graphics, 112(??):13–21, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-

7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849323000316.

Youens-Clark:2022:CLR

[YC22] Ken Youens-Clark. Command-line Rust: a Project-Based Primer for Writing Rust CLIs. O’Reilly Media, Sebastopol, CA, USA, 2022. ISBN 1-0981-0943-0 (paperback). xviii + 377 pp. LCCN QA76.73.R87 Y68 2022.

Ye:2008:DSA

[YCJK08] Jieping Ye, Jianhui Chen, Ravi Janardan, and Sudhir Kumar. Developmental stage annotation of Drosophila gene expression pattern images via an entire solution path for LDA. ACM Transactions on Knowledge Discovery from Data (TKDD), 2 (1):4:1–4:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yu:1995:DTA

[YD95] Chansu Yu and Chita R. Das. Disjoint task allocation algorithms for MIN machines with minimal conflicts. IEEE Transactions on Parallel and Distributed Systems, 6(4):373–387, April 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-

2183 (electronic). URL http://www.computer.org/tpds/
td1995/l0373abs.htm.

Yu:2015:ESS

[YDDB15] William Yu, Noah M. Daniels, David Christian Danko, and Bonnie Berger. Entropy-scaling search of massive biological data. Cell Systems, 1(2):130–140, August 26,

2015. ISSN 2405-4712. URL http://gems.csail.mit.edu/;
 http://www.sciencedirect.com/science/article/pii/
S2405471215000587.

Yuan:2018:ASP

[YDW18] Xingliang Yuan, Huayi Duan, and Cong Wang. Assuring string pattern matching in outsourced middleboxes. IEEE/ ACM Transactions on Networking, 26(3):1362–1375, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2023:LPM

[YGG^{+}23] Yongwei Yuan, Scott Guest, Eric Griffis, Hannah Potter, David Moon, and Cyrus Omar. Live pattern matching with typed holes. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):96:1–96:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL https://

dl.acm.org/doi/10.1145/3586048.

Yoo:1991:EAL

[YH91] H. Yoo and K. Hashiguchi. Extended automata-like regular expressions of star degree at most (2, 1). Theoretical Computer Science, 88(2):351–363, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoo:1992:ERE

[YH92] H. Yoo and K. Hashiguchi. Extended regular expressions of arbitrary star degrees. Theoretical Computer Science, 97(2): 217–231, April 27, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yu:2015:EEA

[YHV^{+}15] Qiang Yu, Hongwei Huo, Jeffrey Scott Vitter, Jun Huan, and Yakov Nekrich. An efficient exact algorithm for the motif stem search problem over large alphabets. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(2): 384–397, March 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Yasuda:1989:PAM

[YIAS89] Takuya Yasuda, Satoshi Igarashi, Oichi Atoda, and Nobuo Saito. ‘Pattern associative memory’ and its use in logic programming language processors. Systems and Computers in Japan, 20(11):11–20, November 1989. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

You:1984:PES

[YJ84] Zhisheng You and Anil K. Jain. Performance evaluation of shape matching via chord length distribution. Computer Vision, Graphics, and Image Processing, 28(2):185–198, November 1984. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Yoshida:2011:PCP

[YK11] S. Yoshida and T. Kida. On performance of compressed pattern matching on VF codes. In Storer and Marcellin [SM11], page 486. ISBN 1-61284-279-8. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749543. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Yang:2011:FME

[YKGS11] Liu Yang, Rezwana Karim, Vinod Ganapathy, and Randy Smith. Fast, memory-efficient regular expression matching with NFA-OBDDs. Computer Networks (Amsterdam, Netherlands: 1999), 55(15):3376–3393, October 27, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128611002489.

Yodaiken:1991:MFC

[Yod91] Victor Yodaiken. Modal functions for concise definition of state machines and products. Information Processing Letters, 40(2):65–72, October 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2012:HPC

[YP12] Yi-Hua Edward Yang and Viktor K. Prasanna. Highperformance and compact architecture for regular expression matching on FPGA. IEEE Transactions on Computers, 61 (7):1013–1025, July 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2013:RSS

[YP13] Yi-Hua E. Yang and Viktor K. Prasanna. Robust and scalable string pattern matching for deep packet inspection on multicore processors. IEEE Transactions on Parallel and Distributed Systems, 24(11):2283–2292, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yaman:2021:QVS

[YPG21] Sinem Getir Yaman, Esteban Pavese, and Lars Grunske. Quantitative verification of stochastic regular expressions. Fundamenta Informaticae, 179(2):135–163, ???? 2021. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2016:NFI

[YQW^{+}16] Xiaochun Yang, Tao Qiu, Bin Wang, Baihua Zheng, Yaoshu Wang, and Chen Li. Negative factor: Improving regularexpression matching in strings. ACM Transactions on Database Systems, 40(4):25:1–25:46, January 2016. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Yeh:2003:CMS

[YT03] Yi-Shiung Yeh and Ta-Shan Tsui. A concurrent multi-string matching from multi-text algorithm based on the algorithm of Hamiltonian path problem and DNA computation. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):71– 96, 2003. CODEN ???? ISSN 0972-0529.

Yun:2012:ETB

[Yun12] SangKyun Yun. An efficient TCAM-based implementation of multipattern matching using covered state encoding. IEEE Transactions on Computers, 61(2):213–221, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zaki:1987:FDA

[ZA87] M. Zaki and Al. H. Albarhamtoshy. Formal design of an Arabic text formatter for microcomputers. Computer Languages, 12(2):123–143, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zhang:2017:APM

[ZA17] Peng Zhang and Mikhail J. Atallah. On approximate pattern matching with thresholds. Information Processing Letters, 123(??):21–26, July 2017. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019017300376.

Zheng:2014:MMS

[ZBST14] Yuxin Zheng, Zhifeng Bao, Lidan Shou, and Anthony K. H. Tung. MESA: a map service to support fuzzy type-ahead search over geo-textual data. Proceedings of the VLDB Endowment, 7(13):1545–1548, August 2014. CODEN ???? ISSN 2150-8097.

Zetzsche:1989:IPR

[ZC89] Christoph Zetzsche and Terry Caelli. Invariant pattern recognition using multiple filter image representations. Computer Vision, Graphics, and Image Processing, 45(2):251–262, February 1989. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Ziadi:1999:OPA

[ZC99] Djelloul Ziadi and Jean-Marc Champarnaud. An optimal parallel algorithm to convert a regular expression into its Glushkov automaton. Theoretical Computer Science, 215 (1–2):69–87, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:/

/www.elsevier.com/cas/tree/store/tcs/sub/1999/215/
1-2/2667.pdf.

Zhong:2023:FCA

[ZCH23] Jincheng Zhong, Shuhui Chen, and Biao Han. FPGA-CPU architecture accelerated regular expression matching with fast preprocessing. The Computer Journal, 66(12):2928–2947, December 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://academic.oup.com/

comjnl/article/66/12/2928/6770084.

Zou:2009:DJP

[ZCÖ09] Lei Zou, Lei Chen, and M. Tamer Özsu. Distance-join: pattern match query in a large graph database. Proceedings of the VLDB Endowment, 2(1):886–897, August 2009. CODEN ???? ISSN 2150-8097.

Zou:2012:APM

[ZCÖZ12] Lei Zou, Lei Chen, M. Tamer Özsu, and Dongyan Zhao. Answering pattern match queries in large graph databases via graph embedding. VLDB Journal: Very Large Data Bases, 21(1):97–120, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhai:2012:MML

[ZCS^{+}12] Deming Zhai, Hong Chang, Shiguang Shan, Xilin Chen, and Wen Gao. Multiview metric learning with global consistency and local smoothness. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):53:1–53:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2014:EPS

[ZCT14] Dongxiang Zhang, Chee-Yong Chan, and Kian-Lee Tan. An efficient publish/subscribe index for e-commerce databases. Proceedings of the VLDB Endowment, 7(8):613–624, April 2014. CODEN ???? ISSN 2150-8097.

Zobel:1995:FAM

[ZD95] J. Zobel and P. Dart. Finding approximate matches in large lexicons. Software — Practice and Experience, 25(3):331–345, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2018:PMO

[ZdSO18] Weixin Zhang and Bruno C. d. S. Oliveira. Pattern matching in an open world. ACM SIGPLAN Notices, 53(9):134–146, November 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3393934.3278124.

Zeilberger:2008:FHO

[Zei08] Noam Zeilberger. Focusing and higher-order abstract syntax. ACM SIGPLAN Notices, 43(1):359–369, January 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zaki:1985:PSA

[ZGE85] Mohamed Zaki and S. A. Gamal-Eldin. A portable syntax analyzer for microcomputers. Computer Languages, 10(2): 127–146, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zheng:2015:ESS

[ZGS^{+}15] Yunhui Zheng, Vijay Ganesh, Sanu Subramanian, Omer Tripp, Julian Dolby, and Xiangyu Zhang. Effective searchspace pruning for solvers of string equations, regular expressions and length constraints. In Kroening and Păsăreanu [KP15], pages 235–254. ISBN 3-319-21689-9. URL http://

link.springer.com/chapter/10.1007/978-3-319-21690-
4_14.

Zhu:2016:BAC

[ZGY^{+}16] Tiantian Zhu, Hongyu Gao, Yi Yang, Kai Bu, Yan Chen, Doug Downey, Kathy Lee, and Alok N. Choudhary. Beating the artificial chaos: Fighting OSN spam using its own templates. IEEE/ACM Transactions on Networking, 24(6):3856– 3869, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:1996:AMI

[Zha96] H. (Hantao) Zhang, editor. Automated Mathematical Induction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 94-010-7250-7 (print), 94-009-1675-2 (e-book). 224 pp. LCCN Q334-342. URL http:

//public.eblib.com/choice/publicfullrecord.aspx?p=
3102529.

Zhang:2007:SSS

[Zha07] Louxin Zhang. Superiority of spaced seeds for homology search. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 4(3):496–505, July 2007. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Zhang:2017:FCP

[Zha17] Meng Zhang. Fast convolutions of packed strings and pattern matching with wildcards. International Journal of Foundations of Computer Science (IJFCS), 28(3):289–??, April 2017. CODEN IFCSEN. ISSN 0129-0541.

Zhou:2012:PSG

[ZHWW12] Yahan Zhou, Haibin Huang, Li-Yi Wei, and Rui Wang. Point sampling with general noise spectrum. ACM Transactions on Graphics, 31(4):76:1–76:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ziadi:1996:REL

[Zia96] Djelloul Ziadi. Regular expression for a language without empty word. Theoretical Computer Science, 163(1– 2):309–315, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_
sub/browse/browse.cgi?year=1996&volume=163&issue=1-
2&aid=2231.

Zhou:2014:TCS

[ZJL14] Shizhe Zhou, Changyun Jiang, and Sylvain Lefebvre. Topology-constrained synthesis of vector patterns. ACM Transactions on Graphics, 33(6):215:1–215:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhan:2018:EDM

[ZJP^{+}18] Yang Zhan, Yizheng Jiao, Donald E. Porter, Alex Conway, Eric Knorr, Martin Farach-Colton, Michael A. Bender, Jun Yuan, William Jannen, and Rob Johnson. Efficient directory mutations in a full-path-indexed file system. ACM Transactions on Storage, 14(3):22:1–22:??, November 2018. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Zhu:2012:GFE

[ZKA12] Haohan Zhu, George Kollios, and Vassilis Athitsos. A generic framework for efficient and effective subsequence retrieval. Proceedings of the VLDB Endowment, 5(11):1579–1590, July 2012. CODEN ???? ISSN 2150-8097.

Zhang:2007:MPP

[ZKCY07] Minghua Zhang, Ben Kao, David W. Cheung, and Kevin Y. Yip. Mining periodic patterns with gap requirement from sequences. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(2):7:1–7:??, August 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zha:2018:CRC

[ZL18] Yue Zha and Jing Li. CMA: A reconfigurable complex matching accelerator for wire-speed network intrusion detection. IEEE Computer Architecture Letters, 17(1):33–36, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zheng:2011:SPM

[ZLN11] Kai Zheng, Hongbin Lu, and Erich Nahum. Scalable pattern matching on multicore platform via dynamic differentiated distributed detection (D^{4}). IEEE Transactions on Computers, 60(3):346–359, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2003:APM

[ZMAB03] Nan Zhang, A. Mukherjee, D. Adjeroh, and T. Bell. Approximate pattern matching using the Burrows–Wheeler transform. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314. LCCN QA76.9.D33 D37 2003.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1194077. IEEE Computer Society Order number PR01896.

Zobel:1993:SLL

[ZMSD93] Justin Zobel, Alistair Moffat, and Ron Sacks-Davis. Searching large lexicons for partially specified terms using compressed inverted files. In Agrawal et al. [ABB93], pages 290–301. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/ZobelMS93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Zheng:2020:SGT

[ZMWL20] Lixiao Zheng, Shuai Ma, Yuanyang Wang, and Gang Lin. String generation for testing regular expressions. The Computer Journal, 63(1):41–65, January 2020. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/academic.oup.com/comjnl/article/63/1/41/5288328.

Zha:2013:GGH

[ZS13] Xinyan Zha and Sartaj Sahni. GPU-to-GPU and host-to-host multipattern string matching on a GPU. IEEE Transactions on Computers, 62(6):1156–1169, June 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zengin:2017:FAH

[ZS17] Salih Zengin and Ece Guran Schmidt. A fast and accurate hardware string matching module with Bloom filters. IEEE Transactions on Parallel and Distributed Systems, 28(2):305–317, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2017/02/07485864-
abs.html.

Zhu:1989:TTD

[ZT89] Rui Feng Zhu and Tadao Takaoka. A technique for twodimensional pattern matching. Communications of the Association for Computing Machinery, 32(9):1110–1120, September 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/66459.html.

Zuendorf:1996:GPM

[Zue96] A. Zuendorf. Graph pattern matching in PROGRES. Lecture Notes in Computer Science, 1073:454–??, 1996. CODEN LNCSD9. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zajac:1997:GBM

[ZV97] R. Zajac and M. Vanni. Glossary-based MT engines in a multilingual analyst’s workstation architecture. Machine Translation, 12(1–2):131–151, ???? 1997. CODEN MACTEZ. ISSN 0922-6567 (print), 1573-0573 (electronic).

Zvegintzov:1980:PMR

[Zve80] N. Zvegintzov. Partial-match retrieval in an index sequential directory. The Computer Journal, 23(1):37–40, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/37.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_01/tiff/37.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
39.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/40.tif.

Zuo:2021:SIS

[ZWH^{+}21] Zhiqiang Zuo, Kai Wang, Aftab Hussain, Ardalan Amiri Sani, Yiyu Zhang, Shenming Lu, Wensheng Dou, Linzhang Wang, Xuandong Li, Chenxi Wang, and Guoqing Harry Xu. Systemizing interprocedural static analysis of large-scale systems code with Graspan. ACM Transactions on Computer Systems, 38(1–2):4:1–4:39, July 2021. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL https://dl.

acm.org/doi/10.1145/3466820.

Zhao:2013:EPG

[ZXL^{+}13] Xiang Zhao, Chuan Xiao, Xuemin Lin, Wei Wang, and Yoshiharu Ishikawa. Efficient processing of graph similarity queries with edit distance constraints. VLDB Journal: Very Large Data Bases, 22(6):727–752, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2015:ECS

[ZYQ^{+}15] Zhiwei Zhang, Jeffrey Xu Yu, Lu Qin, Lijun Chang, and Xuemin Lin. I/O efficient: computing SCCs in massive graphs. VLDB Journal: Very Large Data Bases, 24(2):245– 270, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zu:2012:GBN

[ZYX^{+}12] Yuan Zu, Ming Yang, Zhonghu Xu, Lin Wang, Xin Tian, Kunyang Peng, and Qunfeng Dong. GPU-based NFA implementation for memory efficient high speed regular expression matching. ACM SIGPLAN Notices, 47(8):129–140, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Zeng:2012:CSB

[ZZ12] Qiang Zeng and Hai Zhuge. Comments on “Stack-based Algorithms for Pattern Matching on DAGs”. Proceedings of the VLDB Endowment, 5(7):668–679, March 2012. CODEN ???? ISSN 2150-8097.

Zhang:2010:PMW

[ZZH10] Meng Zhang, Yi Zhang, and Liang Hu. Pattern matching with wildcards using words of shorter length. Information Processing Letters, 110(24):1099–1102, November 30, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2016:CRA

[ZZH16] Meng Zhang, Yi Zhang, and Chen Hou. Compact representations of automata for regular expression matching. Information Processing Letters, 116(12):750–756, December 2016. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019016301016.

Zhu:2020:DBA

[ZZJC20] Xiaodong Zhu, Yi Zhang, Liehui Jiang, and Rui Chang. Determining the base address of MIPS firmware based on absolute address statistics and string reference matching. Computers & Security, 88(??):Article 101504, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819300860.